A Complete Bibliography of Publications in Computer Networks (Amsterdam, Netherlands: 2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

14 August 2020 Version 1.105

Title word cross-reference

2 [1672, 2377]. 2.5 [122]. 3 [1204, 1489, 1722, 2377, 2496, 2757, 2885, 2925]. 3.5 [122]. 4 [1000]. 5 [2309, 2748, 2880, 2926, 2936, 2956, 3009]. 6 [2992]. ^{∗} [636, 1879]. A^{3} [1114]. Δ [1675]. K [143, 428, 669, 833]. m [80, 1568]. M^{2} [305]. N [1156]. p [57, 102, 224]. Q [2922]. S [2655].

-box [1879]. -coverage [669]. -cycle [57, 305]. -cycles [102, 224]. -Dense [833]. -Fault [428]. -frame [1672]. -learning [2922]. -month [2992]. -PONs [2993]. -trail [57]. -trails [1568]. -Tree [143].

.nl [2532]. .nl-domain [2532].

/1609 [436].

1 [1800, 1826]. 10 [38, 756]. 106854 [2997]. 162 [2997].

2 [1189, 1755]. 2017 [2376]. 2019

[2974, 2998]. 2167 [697]. 24 [2998]. 2VC [2655].

3-ent [2344]. 3DVS [2885]. 3G

[531, 926, 1382, 1954]. 3G/4G [1954]. 3GPP [527, 1054, 1139, 1379, 1662].

40 [2798]. 4G [1397, 1484, 1954, 2079]. 4PR [2045].

56 [697]. 5G [1564, 1709, 1916, 2055, 2092, 2119, 2175, 2251, 2264, 2318, 2353–2357, 2459, 2600, 2614, 2657, 2662, 2683, 2684, 2689, 2798]. 5G-IoT [2318].

60 [1113, 2019]. 65 [1783]. 6LoWPAN [2396, 2537]. 6Tree [2832].

79 [2127].

802.11 [137, 198, 230, 296, 566, 1377, 1480, 1489, 1551, 1573, 1920, 2523, 2615]. 802.11-based [439, 597, 1573]. 802.11/LTE [1920]. 802.11/LTE-Advanced [1920]. 802.11a [51]. 802.11ac [1829]. 802.11ah [1898, 1984]. 802.11e [184, 376, 718, 803]. 802.11g [128]. 802.11n [738]. 802.11p [522, 1953]. 802.11p/a [522]. 802.11s [1637]. 802.15.4 [1103, 1780, 2696]. 802.15.4-based [1780]. 802.16 [164, 372, 515, 517, 518]. 802.16e [456, 513, 523, 793, 930]. 802.16j [516, 529, 992, 1259, 1262]. 802.16m [519, 521]. 802.21 [364].

9 [2974].

A-CACHE [1581]. AAoT [2444]. Abacus [351]. abandonment [1930]. ABC [2478]. ABC-AFS [2478]. abnormal [2143]. ABS [190, 2744]. ABSCEV [2515]. absolute [238]. academic [583]. accelerated [1039]. Access [3, 15, 44, 46, 84, 127, 163, 197, 222, 264, 267, 280, 283, 286, 290, 362, 404, 454, 526, 550, 584, 620, 674, 678, 681, 689, 712, 722, 734, 763, 819, 828, 832, 841, 903, 904, 945, 1023, 1062, 1091, 1101, 1129, 1149, 1191, 1205, 1208, 1231, 1252, 1266, 1283, 1302, 1350, 1351, 1418, 1422, 1459, 1496, 1557, 1596, 1627, 1632, 1683, 1711, 1729, 1750, 1754, 1755, 1758, 1792, 1809, 1858, 1931, 1948, 2038, 2074, 2079, 2086, 2122, 2145, 2219, 2344, 2426, 2434, 2452, 2503, 2529, 2543, 2562, 2592, 2632, 2684, 2733, 2740, 2747, 2751, 2807, 2907, 2947, 2977, 2989]. access [559, 624, 1829, 1910, 2064, 2367]. access-based [362]. access/multi [264]. accessibility [2793]. accessing [2398]. accountable [1606]. Accuracy [336, 561, 1032, 2457, 2473, 2758]. accuracy/performance [2473]. Accurate [76, 296, 351, 499, 652, 1248, 1641, 1913, 1954, 2105, 2331, 2528, 2859, 2969]. achievable [1294]. achieve [36, 186, 2126]. Achieving [479, 761, 1169, 1383, 2078, 2264, 2290, 2487, 2593]. ACK [184]. acknowledged [2699]. acknowledging [1540]. ACO [2276, 2899]. ACO-based [2899]. ACO-inspired [2276]. acoustic [1007, 2564, 2897]. across [325, 1409, 1690, 2427]. ACST [2933]. act [784]. action [488, 2475]. activation [2211]. Active [50, 449, 596, 864, 897, 950, 1147, 1440, 1647, 1884, 1952, 2832]. activities [539, 1769, 2097]. Activity [845, 1939, 2497, 2646, 2694]. actor [1681]. actuator [2803]. actuators [2456]. Ad [8, 46, 64, 78, 92, 112, 113, 130, 143, 146, 271, 306, 388, 416, 423, 425, 464, 467, 477, 484, 496, 572, 588, 598, 601, 628, 653, 720, 784, 789, 805, 838, 942, 990, 1001, 1071, 1096, 1177, 1284, 1343, 1386, 1398, 1399, 1432, 1495, 1500, 1501, 1526, 1550, 1595, 1639, 1652, 1831, 1905, 1951, 1964, 1980, 1997, 2028, 2077, 2102, 2150, 2154, 2185, 2193, 2204, 2220, 2230, 2270, 2379, 2390, 2397, 2485, 2757, 2794, 2892, 2983]. Ad-hoc [477, 601, 942, 1495, 1501, 1595, 1905, 1951, 2028, 2150, 2204, 2485]. ad-hoc-network [306]. Adaptable [1910]. Adaptation [221, 230, 375, 515, 560, 588, 899, 1229, 1367, 1516, 1551, 1852, 2091, 2324, 2421, 2491]. adaption [2501]. Adaptive [86, 101, 190, 194, 226, 338, 370, 403, 426, 434, 521, 552, 575, 581, 657, 665, 701, 717, 739, 741, 771, 843, 880, 925, 992, 1028, 1069, 1103, 1210, 1217, 1280, 1297, 1318, 1350, 1424, 1433, 1503, 1516, 1532, 1539, 1674, 1699, 1715, 1781, 1906, 1909, 1917, 1930, 1947, 1997, 2011, 2119, 2211, 2238, 2263, 2358, 2421, 2487, 2489, 2495, 2496, 2550, 2562, 2577, 2584, 2606, 2626, 2654, 2655, 2665, 2681, 2740, 2754, 2782, 2794, 2813, 2859, 2907, 2926, 2942, 2957, 2961, 2994]. add [915]. add/drop [915]. Adding [1248]. additional [1440]. additive [703]. Address [24, 388, 553, 1205, 2161, 2203, 2537, 2832]. Address-free [1205]. addresses [100, 347, 1820, 2832]. Addressing [2037]. adjacency [1470]. adjustment [1052, 2417]. Admission [3, 535, 906, 1000, 1064, 1262, 2584]. admitting [2590]. Adopting [2852]. Adoption [970, 1408, 1854]. ADSL2 [709]. advance [2521]. Advanced [958, 1095, 1153, 1159, 1287, 1336, 1420, 1603, 1920, 1956, 2003, 2394, 2452, 2658, 2728, 2777, 2906]. Advances [42, 356, 1312, 1991, 2036, 2588]. advantage [1815]. Adversarial [2602, 2887]. adverse [397, 2912]. advertising [578, 1466, 1719]. Advice [216, 2352]. aerial [1805, 2845, 3002]. AF [1104, 1872]. AFAR [1624]. affected [2698]. Affinity [157]. AFS [2478]. after [560, 599]. against [170, 200, 234, 245, 522, 530, 677, 879, 1122, 1343, 1679, 1814, 1978, 1997, 2206, 2486, 2496, 2511, 2970, 2981, 3010]. Age [1243, 2900]. aged [2347]. agenda [2114]. Agent [231, 1368, 1846, 2515]. Agent-based [231, 1368, 2515]. agents [1383, 2455]. aggravation [2953]. aggregate [185]. Aggregated [1169, 1395, 1435, 2772]. aggregates [344]. Aggregation [268, 298, 377, 738, 923, 1030, 1082, 1090, 1200, 1226, 1231, 1429, 1438, 1621, 1678, 1684, 1829, 1878, 1904, 2227, 2326, 2328, 2361, 2467, 2529, 2700, 2901]. Agile [32, 1821]. agility [835]. agreement [228, 483, 1061, 1102, 1757, 2338, 2362, 2398, 2546, 2572, 2707, 2958]. ahead [506, 1423, 1476]. AI [3008]. aided [1075, 2144, 2674, 2924]. AIMD [83]. AIMD/RED [83]. AIR [636]. airborne [2344]. Airtime [1727]. AKA [1061]. AL [2296]. AL-FEC [2296]. alarm [2621]. alert [907]. Alerting [2024]. alerts [410, 2397]. algebra [420]. Algorithm [58, 64, 75, 123, 126, 136, 192, 213, 226, 273, 305, 309, 334, 427, 429, 456, 485, 514, 553, 575, 620, 629, 721, 747, 785, 793, 917, 934, 1009, 1010, 1015, 1149, 1180, 1226, 1259, 1316, 1357, 1365, 1370, 1372, 1386, 1475, 1494, 1557, 1653, 1759, 1765, 1836, 1840, 1848, 1867, 1898, 1944, 1985, 2021, 2077, 2135, 2141, 2149, 2165, 2168, 2179, 2182, 2201, 2205, 2235, 2238, 2243, 2269, 2271, 2278, 2286, 2287, 2313, 2336, 2340, 2347, 2377, 2465, 2478, 2493, 2513, 2561, 2574, 2622, 2642, 2653, 2654, 2659, 2662, 2709, 2737, 2774, 2781, 2782, 2796, 2809, 2840, 2884, 2887, 2922, 2928, 2932, 3013]. Algorithm [764, 1073, 1847]. Algorithmic [787, 1431]. Algorithms [91, 142, 225, 247, 275, 301, 398, 457, 484, 513, 601, 619, 622, 670, 749, 751, 763, 768, 804, 869, 1004, 1101, 1186, 1332, 1467, 1537, 1681, 1715, 1806, 1824, 1857, 1876, 2049, 2111, 2118, 2203, 2230, 2258, 2405, 2421, 2442, 2521, 2575, 2614, 2687, 2783, 2863, 2864, 2869, 2931, 2943, 2964]. aligned [1586]. alignment [2792]. all-optical [32, 305, 394, 721, 1100]. All-Path [1473]. All-to-all [1366]. alleviate [599, 1898]. Alleviating [1660]. allocate [2401]. allocating [453, 2952]. allocation [33, 36, 73, 131, 171, 338, 440, 513, 516, 665, 692, 725, 736, 780, 998, 1013, 1017, 1024, 1031, 1068, 1095, 1104, 1127, 1209, 1219, 1230, 1240, 1262, 1274, 1304, 1320, 1335, 1337, 1338, 1370, 1438, 1475, 1481, 1491, 1512, 1550, 1586, 1599, 1629, 1651, 1653, 1667, 1692, 1759, 1785, 1819, 1823, 1826, 1832, 1872, 1873, 1876, 1888, 1963, 2120, 2138, 2162, 2210, 2222, 2235, 2255, 2279, 2346, 2389, 2422, 2448, 2469, 2518, 2537, 2543, 2565, 2584, 2592, 2626, 2658, 2727, 2768, 2774, 2809, 2834, 2863, 2882, 2888, 2918, 3009]. allocations [1272, 1534, 2902]. allowances [534]. ALM [913]. ALOC [692]. ALOHA [1666, 2831]. alone [1125]. ALPS [741]. altering [2700]. alternate [594, 721]. alternates [100, 1777]. alternative [1389, 2204]. altruistic [677]. always [949]. always-on [949]. Amazon [1735]. ambient [217, 658, 1643]. ambiguity [921]. amenable [324]. America [472]. AMGT [2418]. AMI [2412]. among [180, 362, 758, 1273, 1438, 2027, 2774, 2800]. amplification [2131]. amusement [202]. An-Improved-Route [636]. analog [2064]. Analysing [1522, 2380, 2418]. Analysis [9, 115, 131, 165, 167, 176, 178, 184, 196, 198, 259, 304, 329, 331, 353, 379, 412, 433, 518, 568, 593, 611, 643, 688, 700, 710, 711, 723, 845, 851, 857, 881, 895, 901, 904, 922, 963, 989, 1003, 1019, 1054, 1055, 1093, 1125, 1141, 1142, 1172, 1219, 1316, 1336, 1411, 1417, 1453, 1455, 1495, 1506, 1538, 1553, 1626, 1661, 1710, 1728, 1751, 1753, 1794, 1821, 1935, 1955, 1975, 2003, 2017, 2094, 2151, 2261, 2321, 2345, 2378, 2416, 2477, 2526, 2539, 2542, 2650, 2673, 2704, 2717, 2733, 2741, 2745, 2779, 2900, 2936, 2960, 2964, 2979, 2982, 2990, 2992, 2993]. analysis [1, 46, 67, 89, 102, 182, 462, 547, 559, 676, 705, 860, 899, 931, 996, 1027, 1333, 1406, 1922, 1953, 2101, 2162, 2311, 2587, 2668, 2867, 2876, 2884]. Analyst [1272]. analytic [538, 865]. Analytical [25, 239, 250, 435, 966, 1369, 1440, 1586, 1666, 2047, 2547, 2862]. analytics [1839, 2364, 2413]. analyze [1049]. analyzer [1780]. analyzers [2455]. Analyzing [930, 964, 2090]. anatomy [809]. anchor [1581]. anchor-based [1581]. anchoring [1318]. Android [1225, 1783, 2084]. Android-based [2084]. ANDSF [1920]. ANDSF-Assisted [1920]. ANEL [2983]. ANFIS [1490]. ANFIS-based [1490]. annealing [725]. annotation [1621]. annulus [2674]. Anomaly [135, 499, 549, 922, 1058, 1148, 1174, 1210, 1347, 1626, 1883, 1938, 2297, 2453, 2478, 2560, 2617, 2865, 2961]. anomaly-based [549, 922, 1883, 2297]. anonymity [31, 266, 479, 1114, 1260, 1334, 1838]. anonymity-preserving [1838]. Anonymous [145, 498, 1005, 1356, 1511, 1707, 2349, 2362, 2546, 2724, 2881]. Ant [220, 606, 701, 1166, 2622]. Ant-based [220]. AntBot [391]. ante [2323]. antenna [990, 1057, 2446, 2741]. antennas [173, 634, 1095]. Anti [145, 391, 953, 968, 2737]. anti-collision [953]. Anti-localization [145]. anti-matchability [2737]. anti-patterns [968]. Anti-pollution [391]. any [2537]. any-To-any [2537]. anycast [9, 994, 1475, 2118, 2665]. anycast-based [2118]. AOA [381]. AP [430, 2407]. API [2856]. Aplasia [1012]. appending [1007]. applicability [2270]. Application [22, 200, 252, 283, 379, 483, 489, 561, 726, 761, 875, 913, 923, 1025, 1044, 1045, 1109, 1114, 1274, 1276, 1277, 1371, 1447, 1553, 1580, 1647, 1662, 1890, 1936, 2084, 2183, 2331, 2410, 2459, 2580, 2588, 2797, 2822, 2830, 2867, 2884]. Application-aware [22, 1114]. application-based [561, 1662]. application-defined [1580]. application-layer [1890]. Application-Level [489]. application-specific [1277, 1371]. Applications [153, 227, 243, 248, 269, 349, 354, 376, 382, 408, 417, 475, 522, 531, 550, 557, 612, 614, 656, 658, 739, 803, 887, 935, 1055, 1069, 1098, 1193, 1197, 1207, 1223, 1253, 1285, 1309, 1312, 1359, 1479, 1553, 1594, 1600, 1606, 1615, 1624, 1693, 1817, 1834, 1879, 1886, 1891, 1903, 1943, 2046– 2048, 2051, 2065, 2315, 2419, 2438, 2443, 2464, 2641, 2710, 2762, 2772, 2778, 2799, 2811, 2817, 2852, 2881, 2936]. applied [630, 693, 1027, 2653]. applying [1467]. Approach [21, 54, 72, 99, 114, 210, 227, 238, 241, 250, 258, 293, 548, 608, 733, 739, 750, 772, 777, 839, 922, 946, 956, 977, 987, 998, 1048, 1058, 1083, 1086, 1089, 1101, 1149, 1158, 1296, 1315, 1319, 1338, 1376, 1400, 1443, 1444, 1467, 1472, 1521, 1541, 1547, 1576, 1577, 1589, 1590, 1610, 1625, 1637, 1652, 1658, 1666, 1694, 1699, 1748, 1766, 1776, 1778, 1864, 1865, 1870, 1935, 1958, 2008, 2027, 2028, 2039, 2060, 2067, 2124, 2185, 2215, 2360, 2381, 2406, 2441, 2451, 2527, 2538, 2634, 2694, 2708, 2751, 2793, 2823, 2836, 2849, 2857, 2858, 2862, 2873, 2906, 2913, 3009]. approach [921, 1090, 1317, 1406, 1425, 1719, 2403, 2487, 2883, 2930]. approaches [74, 92, 346, 686, 905, 1285, 1332, 1364, 1817, 2076, 2181, 2350, 2390, 2535, 2650, 2702, 2775]. approximate [225]. Approximating [2573]. approximation [561, 626, 888, 1024, 1767]. approximations [1537]. AQM [753, 1236]. AQMs [1609]. ARBAT [2689]. arbitrary [1396]. architectural [1042, 1762, 2646]. Architecture [6, 19, 82, 87, 263, 282, 319, 358, 362, 364, 424, 510, 565, 726, 790, 911, 969, 974, 1080, 1096, 1153, 1176, 1185, 1195, 1266, 1281, 1363, 1447, 1483, 1564, 1755, 1756, 1825, 1853, 1874, 1956, 2107, 2208, 2288, 2291, 2299, 2354, 2359, 2394, 2502, 2536, 2672, 2689, 2748, 2821, 2856, 2950]. Architectures [315, 408, 507, 1136, 1139, 1286, 1773, 2191, 2464]. archival [2241]. Area [149, 201, 263, 415, 450, 519, 521, 548, 694, 877, 1293, 1553, 1802, 1998, 2019, 2063, 2274, 2362, 2529, 2546, 2595, 2929, 2961]. areas [2170, 2822]. AREOR [2942]. ARP [1086]. ARQ [403, 657]. ARQ-HARQ [657]. arrivals [2395]. art [1098, 1328, 2770]. Artificial [258, 1494, 2605]. artworks [1621]. AS-level [183, 833]. AS-paths [972]. aspects [913, 1042]. ASSCA [2856]. assembly [2868]. Assessing [1292, 1877, 1941, 2397, 2409]. Assessment [24, 228, 317, 462, 478, 993, 1076, 1305, 1954, 2433, 2798, 2910, 2950]. assign [2427]. assignment [75, 254, 333, 411, 554, 558, 662, 1048, 1104, 1486, 1758, 2029, 2138, 2274, 2334, 2342, 2739, 2989]. assistance [473, 2641]. Assisted [303, 554, 675, 957, 1037, 1199, 1397, 1478, 1822, 1842, 1849, 1920, 1923, 1980, 2011, 2136, 2319, 2365, 2423, 2425, 2506, 2548, 2552, 2560, 2766, 2770, 2977, 2979, 3002]. associated [336]. Association [685, 1079, 2503, 2639, 2727, 2842, 2907]. assurance [357, 1373, 2327, 2553, 2816]. Astraea [2830]. Asymmetric [483, 1603, 2859]. asymmetry [840]. Asymptotic [546, 1220, 1823]. asynchronous [146, 165, 825, 1050, 1597]. Atlas [2409]. attachment [102, 1427]. Attack [86, 260, 328, 410, 412, 770, 965, 1105, 1327, 1515, 1761, 1814, 1882, 1998, 2090, 2554, 2760, 2970]. attacker [2418]. Attacker-Manager [2418]. Attackers [1646]. Attacking [2841]. Attacks [170, 200, 203, 256, 304, 530, 549, 717, 774, 783, 879, 1040, 1086, 1334, 1343, 1468, 1490, 1679, 1792, 1978, 1979, 2005, 2131, 2137, 2183, 2185, 2206, 2361, 2481, 2486, 2492, 2511, 2671, 2757, 2792, 2981, 3010]. attention [2597]. attenuator [328]. attestation [1511, 2444, 2725]. Attribute [31, 2425, 2426, 2548, 2772, 2807]. Attribute-based [31, 2426]. attribute-hiding [2807]. Auction [1252, 1496, 1763, 2120, 2469, 2726, 2834]. Auction-based [1252, 1763, 2120]. auctions [414]. audio [630]. Audit [2933]. Audit-based [2933]. augmented [814]. Augmenting [16, 2752]. Authenticated [744, 1102, 1838, 2000, 2538]. Authenticating [576]. Authentication [117, 185, 266, 443, 523, 556, 682, 777, 997, 1005, 1054, 1061, 1139, 1334, 1344, 1387, 1470, 1594, 1757, 1813, 1891, 2321, 2322, 2349, 2362, 2398, 2437, 2444, 2546, 2706, 2707, 2724, 2881, 2929, 2983, 2984]. authority [2425, 2718]. autoconfiguration [24, 388]. autoforwarding [1012]. automata [64, 181, 1235, 2231]. Automated [1431, 1589, 2759, 2811]. Automatic [166, 338, 373, 850, 1129, 2284, 2420, 2706, 2716, 2910]. automation [985, 2865]. Autonomic [74, 361, 544, 870]. Autonomous [525, 885, 1084, 1972, 2465, 2466, 2530]. AV [1172]. AV-based [1172]. Availability [252, 265, 313, 459, 758, 1213, 1270, 1355, 1426, 1519, 1559, 1671, 2232, 2647, 2833]. Availability-driven [2833]. available [76, 399, 766, 823, 1111, 2615, 2643, 2644]. average [1767]. avoidance [720, 904, 1947, 2323]. Awarding [1540]. Aware [22, 47, 58, 92, 113, 214, 219, 238, 259, 268, 360, 405, 426, 668, 670, 684, 697, 714, 716, 838, 919, 925, 932, 937, 939, 940, 954, 995, 1029, 1126, 1186, 1267, 1279, 1295, 1318, 1359, 1364, 1365, 1368, 1422, 1456, 1600, 1601, 1633, 1655, 1665, 1713, 1739, 1758, 1764, 1772, 1789, 1809, 1832, 1853, 1868, 1911, 1966, 1984, 2018, 2030, 2055, 2092, 2138, 2140, 2151, 2155, 2238, 2250, 2283, 2289, 2309, 2340, 2344, 2355, 2382, 2452, 2466, 2480, 2527, 2545, 2581, 2592, 2611, 2656, 2661, 2686, 2689, 2701, 2801, 2816, 2860, 2863, 2875, 2880, 2881, 2907, 2940, 3001, 3013]. aware [157, 171, 221, 310, 342, 344, 374, 393, 484, 647, 662, 673, 706, 707, 761, 1018, 1114, 1116, 1121, 1407, 1464, 1466, 1505, 1636, 1654, 1738, 1795, 1802, 1837, 1892, 2038, 2075, 2084, 2121, 2143, 2145, 2172, 2220, 2476, 2662, 2720, 2727, 2781, 2805, 2842]. Awareness [416, 660, 843, 942, 1482, 1616, 1974, 2025, 2048, 2501]. away [1405]. AWG [754, 2635]. AWG-based [754, 2635].

backbone [64, 129, 389, 583, 645, 764, 800, 884, 1021, 1171, 1206, 1347, 1453, 2301, 2423, 2733, 2773]. backbone-assisted [2423]. backbones [6, 820, 2659]. Background [943]. backhaul [584, 643, 2438, 2529, 3013]. backlight [2091]. backoff [1377]. Backpressure [740]. Backscatter [1041]. backup [1539, 2121, 2729, 2900]. backups [1736]. backward [2338]. bacteria [307]. Bad [1609, 1940]. Badumna [151]. BailighPulse [1282]. balance [156, 1241, 2509]. balanced [553, 579, 684, 697, 1894, 2200, 2773]. Balancing [231, 281, 452, 619, 686, 870, 902, 1271, 1280, 1325, 1330, 1350, 1383, 1421, 1561, 1677, 1764, 1826, 1925, 1932, 2207, 2300, 2343, 2442, 2494, 2556, 2859, 2880, 2943, 2994]. balancing-cell [1826]. band [122, 222, 237]. BandAMC [514]. Bandwidth [32, 36, 76, 338, 354, 399, 414, 449, 532, 534, 554, 678, 766, 823, 824, 952, 958, 1030, 1062, 1077, 1111, 1217, 1219, 1262, 1270, 1320, 1332, 1337, 1543, 1653, 1674, 1684, 1705, 1832, 1977, 2083, 2138, 2188, 2290, 2401, 2467, 2678, 2687, 2754, 2816, 3003]. Bandwidth-aware [1832]. bandwidth-based [2188]. bandwidth-delay [449, 1217]. Bandwidth-efficiency-oriented [1543]. Bandwidth-efficient [2687]. BANETs [2432]. bang [2131]. Banking [1263]. bargaining [1332, 1547, 2403]. barrier [302, 1180, 2078, 2410, 2967]. barriers [970]. barring [1350, 2562, 2740]. barycentric [1074]. base [540, 558, 926, 1047, 1124, 1450, 1459, 2059, 2240, 2845, 2880]. base-station [2240]. baseband [2776]. Based [24, 54, 90, 140, 178, 213, 220, 346, 353, 388, 418, 426, 563, 593, 642, 717, 722, 748, 763, 782, 828, 832, 843, 869, 873, 884, 900, 922, 954, 966, 967, 983, 1049, 1050, 1058, 1064, 1078, 1080, 1127, 1164, 1168, 1252, 1268, 1316, 1320, 1338, 1361, 1479, 1503, 1508, 1510, 1541, 1887, 1900, 1915, 2077, 2132, 2188, 2207, 2282, 2323, 2336, 2373, 2379, 2421, 2425, 2441, 2472, 2479, 2481, 2503, 2535, 2581, 2625–2627, 2635, 2641, 2649, 2672, 2756, 2759, 2796, 2828, 2836, 2849, 2858, 2871, 2935, 2940, 2942, 2958, 2961, 2984, 3009]. based [5, 86, 98, 138, 158, 181, 371, 380, 381, 389, 402, 407, 543, 549, 597, 612, 689, 692, 695, 712, 741, 754, 769, 803, 854, 858, 874, 883, 902, 1009, 1038, 1073, 1074, 1150, 1203, 1218, 1234, 1271, 1327, 1368, 1371, 1432, 1433, 1454, 1465, 1490, 1516, 1533, 1558, 1573, 1627, 1658, 1716, 1752, 1759, 1763, 1809, 1810, 1826, 1828, 1836, 1847, 1883, 1884, 1912, 1970, 1989, 2047, 2080, 2102, 2109, 2120, 2164, 2231, 2232, 2238, 2259, 2288, 2321, 2339, 2377, 2426, 2470, 2478, 2491, 2515, 2548, 2569, 2646, 2663, 2755, 2777, 2800, 2807, 2810, 2865, 2870, 2933, 2982, 3001]. based [3, 4, 114, 146, 161, 197, 202, 239, 303, 345, 384, 433, 439, 498, 531, 534, 560, 561, 615, 629, 648, 670, 696, 701, 739, 830, 853, 864, 877, 880, 887, 890, 958, 974, 1025, 1051, 1052, 1068, 1069, 1079, 1086, 1159, 1166, 1167, 1189, 1262, 1287, 1293, 1300, 1305, 1319, 1324, 1332, 1445, 1544, 1549, 1598, 1643, 1644, 1675, 1676, 1696, 1801, 1813, 1823, 1830, 1840, 1843, 1860, 1864, 1866, 1867, 1875, 1965, 2021, 2055, 2060, 2081, 2138, 2147, 2193, 2210, 2222, 2249, 2284, 2290, 2387, 2438, 2455, 2484, 2499, 2561, 2604, 2687, 2726, 2774, 2784, 2824, 2827, 2879, 2889, 2899, 2904, 2917, 2918, 2920, 2980, 2991]. based [9, 52, 55, 64, 84, 134, 157, 162, 177, 231, 243, 255, 299, 316, 322, 434, 446, 447, 494, 553, 649, 650, 716, 750, 772, 792, 821, 822, 852, 907, 942, 1017, 1027, 1143, 1172, 1209, 1341, 1348, 1379, 1407, 1443, 1481, 1512, 1528, 1555, 1584, 1594, 1605, 1608, 1625, 1636, 1657, 1662, 1692, 1765– 1767, 1780, 1839, 1870, 1903, 1958, 1960, 1967, 2039, 2099, 2118, 2135, 2153, 2162, 2203, 2324, 2335, 2340, 2406, 2459, 2501, 2508, 2526, 2566, 2578, 2621, 2648, 2711, 2725, 2732, 2737, 2738, 2741, 2781, 2802, 2803, 2815, 2840, 2844, 2859, 2910, 2948, 2949, 2971, 2978, 2997]. based [25, 31, 45, 112, 120, 141, 204, 226, 295, 309, 340, 352, 358, 362, 436, 442, 462, 485, 525, 538, 582, 586, 594, 624, 625, 639, 705, 721, 731, 744, 789, 841, 1011, 1063, 1106, 1109, 1110, 1129, 1169, 1222, 1240, 1290, 1301, 1344, 1367, 1382, 1387, 1496, 1511, 1543, 1574, 1581, 1585, 1588, 1646, 1682, 1699, 1718, 1719, 1771, 1802, 1846, 1852, 1862, 1894, 1924, 1944, 2037, 2040, 2071, 2084, 2088, 2101, 2124, 2150, 2179, 2182, 2201, 2263, 2271, 2279, 2292, 2297, 2311, 2322, 2328, 2348, 2356, 2361, 2408, 2422, 2423, 2512, 2559, 2593, 2645, 2675, 2694, 2703, 2727, 2734, 2752, 2775, 2826, 2843, 2850, 2866, 2922, 2932, 2957]. based [918, 994, 1416, 1439, 1679, 2183, 2316, 2585, 2655, 2663, 2668, 2778]. BasisEvolution [2453]. batch [768, 2869]. BATON [2895, 2896]. battery [2087, 2382]. battery-powered [2382]. Bayes [2957]. Bayesian [2231, 2607, 2777]. BBS [466]. BBU [2614, 2989]. BCDN [2924]. be [316, 347, 1540]. beacons [474]. beam [634, 1997, 2496]. beamforming [2212, 2907]. BeaQoS [1925]. bearer [2926]. bearing [2124]. before [2306, 2677]. Behavior [4, 568, 900, 922, 1676, 2004, 2104, 2439]. behavioral [351, 852, 863, 2021]. behaviors [677, 1137, 2640]. behaviour [2968]. beings [1750]. belief [1584]. beneficial [1967]. Benefits [1521, 1617, 2207, 2657, 2776]. BER [2296]. Best [48, 180, 606, 1038, 1617, 2016, 2131, 2796]. best-effort [1038]. best-fit [2796]. Better [324, 342, 986, 2012]. Between [4, 156, 216, 299, 682, 740, 1118, 1222, 1715, 1877, 1887, 2105, 2190, 2418, 2580, 2964, 2973]. Beyond [847, 974, 1802, 2950]. beyond-wireless [1802]. BGP [297, 392, 688, 885, 972, 1179, 1235, 1671, 1690, 2017, 2206, 2273]. BGP-XM [885]. bi [1787]. bi-connectivity [1787]. BiCo [396]. bidirectional [124, 711, 1097, 2272, 2779]. Big [1088, 1839, 1846, 1934, 2002, 2074, 2079, 2100, 2279, 2364, 2413, 2524, 2779, 2973]. Bilateral [396, 1775]. bilingual [1406]. billing [141]. bills [700]. binary [1467, 1608, 1830]. binaural [1618]. binning [2754]. Binomial [1707]. bio [70]. bio-inspired [70]. biogeography [1658]. biogeography-based [1658]. Biologically [71, 495]. biometric [2958]. biometrics [863, 2572]. bipartite [978]. bipolar [1299]. bird [900]. bit [1286]. BITE [1439]. bitrate [1358, 1930]. BitTorrent [291, 339, 610, 615, 672, 783, 905, 1123, 1131, 1137, 1143, 1356, 2160]. BitTorrent-based [615]. BitTorrent-like [905, 1131]. black [466, 1343]. blackhole [2090]. BLE [2969]. blended [814]. Blind [195, 1571]. Block [184, 345, 863, 882, 1050, 1758, 1926, 1963, 2162, 2739]. Block-level [882]. block-wise [1926]. BlockA [2984]. blockchain [2648, 2810, 2924, 2941, 2984]. blockchain-aided [2924]. blocker [2985]. blocker-enabled [2985]. blocking [915, 1642, 2217]. Blog [863]. Bloom [13, 245, 349, 382, 595, 1098]. bluetooth [2969]. bluetooth/BLE [2969]. Board [14, 28, 41, 53, 68, 81, 95, 103, 118, 132, 147, 160, 175, 191, 206, 215, 232, 277, 288, 314, 327, 355, 368, 395, 419, 432, 441, 451, 469, 480, 508, 524, 536, 570, 602, 617, 641, 651, 664, 679, 690, 698, 715, 730, 745, 759, 775, 788, 794, 806, 817, 844, 859, 876, 892, 910, 924, 936, 947, 960, 975, 988, 1002, 1016, 1033, 1278, 1289, 1313, 1322, 1354, 1362, 1374, 1403, 1412, 1434, 1448, 1449, 1487, 1497, 1517, 1602, 2414, 2431, 2474, 2488, 2504, 2522, 2534]. Board [1529, 1556, 1563, 1569, 1575, 1582, 1623, 1673, 1686, 1704, 1732, 1733, 1745, 1774, 1784, 1790, 1799, 1808, 1818, 1833, 1850, 1863, 1880, 1896, 1914, 1933, 1945, 1962, 1990, 2001, 2013, 2014, 2042, 2056, 2082, 2103, 2117, 2129, 2139, 2148, 2157, 2166, 2180, 2196, 2214, 2229, 2244, 2260, 2281, 2307, 2325, 2351, 2374, 2375, 2388, 2399, 2551, 2570, 2591, 2612, 2630, 2652, 2670, 2690, 2721, 2742, 2764, 2785, 2786, 2806, 2819, 2829, 2839, 2847, 2861, 2878, 2893, 2911, 2937, 2959, 2975, 2999]. Body [1087, 1293, 1488, 1553, 1776, 1802, 2353, 2362, 2372, 2484, 2546, 2595, 2607, 2929, 2961]. Body-to-Body [1776]. Boltzmann [2617]. BOND [2722]. Bonded [500]. bonding [2790]. BonFIRE [1163]. books [1540]. boosting [372]. border [639]. borrowing [529]. bot [854, 999]. BotCluster [2645]. both [252, 2693]. Botnet [485, 845, 847, 851, 857, 858, 1793, 2645]. botnets [391, 539, 846, 849, 856, 1308]. bots [863, 1910]. bots-infested [1910]. bottleneck [1111, 1644, 1788]. bottleneck-free [1644]. bottlenecks [2953]. bottom [2487]. bottom-up [2487]. Bound [2457, 2711]. Bound-based [2711]. boundary [126, 1296, 1820, 2948, 2997]. Bounded [1198, 1430, 1919, 2639]. Bounds [83, 546, 592]. box [782, 1879, 2514]. Branch [1768, 2648]. branch-and-price [1768]. branching [1101]. Breach [35]. break [2677]. breathing [2442]. Bridging [6, 1473, 2631]. bring [1722]. Bringing [1671]. broadband [404, 819, 1062, 1157, 1302, 1647, 1779, 1940, 2684, 2808]. broadcast [27, 63, 425, 429, 483, 574, 1173, 1595, 1953, 2133, 2606, 2828]. broadcasting [92, 620, 944, 1028, 2274, 2487]. broker [1304]. brokers [2908]. browsing [578, 1923, 2267, 2992]. BS [276, 1462, 1822]. BS/BS [276]. BSC [1124]. BSC-MAC [1124]. BSDE [839]. BSM [565]. BSS [2683]. bucket [1583]. buckets [632, 2519]. budget [675, 2768]. buffer [190, 667, 1441, 1672, 2022, 2341]. bufferbloat [1236]. Buffering [782, 1955, 2685]. buffers [1198]. Building [24, 150, 628, 684, 697, 974, 1057, 1791, 1839, 1969]. builds [1498]. built [2126, 2342]. built-in [2126]. Bulk [723, 1273, 1379, 2987]. bundle [1066, 1580]. bundled [939]. bundles [1540]. bundling [2187]. Burst [56, 263, 466, 2213]. bursty [766, 1336]. bus [1176, 1541]. bus-based [1541]. business [1688, 2112]. busy [402]. busy-hour [402]. buyer [1558]. buyers [2577]. buying [2726]. bypass [1206]. bypasses [2420]. bypassing [2343]. byte [223, 2131].

C [2469, 2510, 2614, 3009]. C-RAN

[2469, 2510, 2614, 3009]. C1G2 [556]. CA [25, 435, 734, 767, 1003, 1352]. cable [1647]. CAC [88]. cache [59, 678, 832, 1040, 1121, 1234, 1397, 1422, 1490, 1559, 1581, 1873, 2015, 2159, 2257, 2258, 2389, 2435, 2892]. cache-aware [1422]. cache-enabled [2892]. caches [1635, 2682]. Caching [93, 881, 989, 1036, 1037, 1078, 1220, 1397, 1444, 1451, 1536, 1542, 1581, 1659, 1710, 2025, 2035, 2037, 2245, 2259, 2277, 2389, 2541, 2585, 2649, 2846, 2862, 2875, 3012]. calculus [194, 824]. calendaring [2947]. call [643, 1064, 1262]. calls [633, 1674]. camera [2424, 2520]. camera-enabled [2520]. campus [6, 1244, 1452, 1473]. Can [246, 316, 347]. cancelable [2572]. cancelation [2831]. cancellation [460, 2423, 2683]. cancellation-based [2423]. cancer [2891]. candidate [457]. Cap [534]. capability [56, 1248, 1531, 1769, 2284, 2850]. capable [394, 420, 1876]. Capacity [23, 73, 110, 164, 353, 366, 404, 411, 494, 515, 643, 800, 824, 872, 931, 1013, 1091, 1111, 1279, 1325, 1375, 1583, 1802, 2138, 2529, 2729, 2979]. capacity-based [494]. capillary [2815]. capture [59, 1610, 2969]. capturing [539]. car [476, 2483]. cards [1102, 2798]. care [972]. Carlo [1011]. CARRADS [86]. carrier [11, 796, 1360, 1432, 1689, 1936, 2029]. carrier-grade [1936]. cars [1239]. cascade [2217]. cascades [611]. cascading [1410]. case [416, 493, 578, 583, 914, 970, 1735, 1825, 2315, 2505, 2532, 3008]. case-study [2315]. cases [1107, 2440]. cast [2169]. casting [2218]. Catching [106, 639]. Categories [1829]. causal [2008]. CCIPCA [564]. CCIPCA-OPCSC [564]. CCMAC [49]. CCN [1762, 2585, 2604]. CCN-based [2585]. CCNxTomcat [1392]. CCS [1926]. CDF [967]. CDMA [72, 222, 940]. CDN [765, 1941, 2008, 2095, 2168, 2924]. CDNs [2118, 2820]. CDS [582, 634, 2459, 2773]. CDS-MEC [2459]. CDS-mode [582, 634]. CDS-SDS [2773]. Cell [891, 1064, 1462, 1826, 2212, 2346, 2371, 2442, 2739, 2920, 3013]. cells [1484, 1824, 2119, 2536]. Cellular [23, 267, 280, 558, 866, 991, 995, 1047, 1129, 1209, 1378, 1391, 1457, 1459, 1461, 1525, 1542, 1561, 1709, 1778, 1822, 1917, 1952, 1955, 2119, 2162, 2264, 2327, 2367, 2411, 2422, 2442, 2650, 2683, 2704, 2882, 2916, 2979, 3001]. cellular-mode [2411]. cellular/WiFi [2327]. CeMon [1680]. Censorship [1554]. Center [408, 714, 865, 902, 911, 938, 1018, 1113, 1232, 1267, 1268, 1337, 1426, 1473, 1483, 1578, 1591, 1600, 1656, 1660, 1842, 1868, 1902, 1909, 1969, 2061, 2080, 2179, 2211, 2252, 2253, 2288, 2306, 2332, 2380, 2517, 2672]. centers [212, 683, 772, 831, 1060, 1534, 1651, 1655, 1748, 1928, 1977, 1983, 2018, 2172, 2381, 2383, 2626, 2698]. central [877, 1649, 2265]. central-based [877]. centralised [1715]. CEntrality [605, 993, 1765, 1881, 2335]. Centrality-based [2335]. centralization [1690]. Centralized [597, 1364, 2122, 2299, 2310]. centrally [914, 1532]. centre [1929]. centres [2946]. Centric [13, 319, 552, 799, 989, 1034– 1038, 1041–1043, 1045, 1077, 1121, 1193, 1247, 1310, 1323, 1364, 1368, 1392, 1422, 1437, 1451, 1483, 1513, 1530, 1559, 1644, 1663, 1696, 1753, 1869, 1893, 1912, 1950, 1951, 1960, 2015, 2023, 2025, 2027, 2035, 2058, 2093, 2126, 2254, 2255, 2276, 2304, 2326, 2403, 2541, 2547, 2557, 2569, 2723, 2813, 2986, 3012]. Centron [2485]. certificate [1233]. certificate-chains [1233]. certificateless [185]. Certification [2054]. CGR [2335]. chain [1702, 1703, 1971, 1984, 2427]. chaining [1741, 2111, 2415, 2869, 2903]. chains [1233, 2952, 2978]. challenge [2618]. Challenges [477, 506, 742, 746, 834, 909, 973, 1312, 1328, 1402, 1416, 1476, 1482, 1599, 1612, 1620, 1688, 1717, 2074, 2337, 2363, 2629, 2657, 2769, 2770, 2822, 2973, 3005]. Change [2352, 2439]. changes [538, 583]. Channel [37, 110, 154, 307, 340, 346, 353, 405, 407, 411, 468, 500, 599, 635, 662, 671, 864, 937, 991, 1024, 1028, 1048, 1082, 1108, 1205, 1214, 1222, 1360, 1384, 1603, 1806, 1884, 1955, 1994, 2030, 2064, 2107, 2342, 2422, 2446, 2467, 2494, 2510, 2554, 2790, 2809, 2837, 2888, 2918, 3003]. channel-aware [937]. channels [21, 201, 448, 696, 851, 899, 2000, 2874]. Chaotic [1819, 1992]. character [1069]. characterisation [2445]. Characteristic [2093, 2802]. characteristics [329, 827, 875, 1125, 2115, 2201, 2477]. Characterization [307, 339, 402, 671, 790, 804, 1501, 2046, 2163, 2483, 2876, 2990]. Characterizing [234, 242, 1879, 2104, 2131]. Chargers [1492, 1796]. charging [527, 981, 1379, 1662, 1715, 1836, 2327, 2515, 2697]. cheater [1573]. check [610]. checking [1646]. chicken [2826, 2964]. children [202]. Chinese [1406, 1631]. chip [2867]. choice [180, 272, 2189]. CHOKe [829]. Chord [459, 563]. Chord-PKI [563]. chunk [1026, 1659]. churn [55]. CINs [964]. ciphertext [2426]. ciphertext-policy [2426]. circuit [1453]. circuit-switched [1453]. circuits [1247]. circulant [1543]. circular [182, 1442]. circular-shaped [182]. Circumventing [1917]. cities [1839, 2178, 2315, 2448, 2647, 2949]. city [1165, 2618, 2844]. CLAMP [177]. class [177, 275, 518, 958, 1350, 2562, 2740, 2978]. class-based [177, 958]. classes [593, 2062]. classification [82, 331, 351, 371, 389, 499, 531, 555, 661, 747, 748, 956, 1006, 1419, 1553, 1817, 1874, 1943, 2153, 2158, 2164, 2203, 2287, 2339, 2390, 2406, 2419, 2470, 2566, 2694, 2709, 2834, 2855, 2873, 3007]. classifier [1640, 1669, 2004, 2627, 2705, 2780, 2980]. classify [2891]. Classifying [2955]. classless [800]. Client [117, 678, 1261, 1825]. Client-server [1261]. clinical [2100]. cloak [174]. clock [917]. CLONE [673]. Cloning [303]. Closed [1150, 1758, 1835]. Closed-loop [1835]. Closeness [993]. Cloud [332, 831, 873, 959, 1163, 1164, 1171, 1185, 1199, 1228, 1265– 1267, 1269, 1272, 1275, 1276, 1295, 1309, 1365, 1445, 1507, 1515, 1577, 1592, 1650, 1651, 1655, 1656, 1730, 1734, 1735, 1737, 1740, 1742– 1744, 1748, 1763, 1792, 1835, 1842, 1849, 1900, 1928, 1942, 1969, 1983, 2005, 2007, 2063, 2071, 2079, 2122, 2155, 2172, 2199, 2207, 2222, 2233, 2319, 2329, 2340, 2365, 2387, 2398, 2419, 2425, 2433, 2434, 2464, 2506, 2508, 2543, 2548, 2590, 2647, 2683, 2695, 2741, 2746, 2751, 2768, 2770, 2778, 2780, 2807, 2898, 2907, 2915, 2917, 2930, 2943, 2947, 2989, 3008, 3011]. Cloud-assisted [1199, 1842, 1849, 2319, 2365, 2506]. Cloud-based [873, 1164, 1445, 2071, 2387, 2508, 2807]. Cloud-BSS [2683]. cloud-enabled [2419, 2433]. Cloud-Integrated [1266]. Cloud-oriented [2233]. Cloud-Radio [2543]. cloudlet [2383]. cloudlets [2988]. clouds [1274, 1736, 1982, 2075, 2427, 2634, 2669]. clubs [2160]. ClusteR [692, 1636, 1810, 2060, 2569, 2683, 2781, 2815, 2870]. ClusteR-based [692, 1636, 2569, 2815]. Cluster-level [1810]. Clustered [77, 779, 1495, 1926]. Clustering [381, 452, 647, 649, 728, 851, 852, 1004, 1027, 1049, 1083, 1119, 1371, 1386, 1787, 1845, 1869, 2146, 2192, 2518, 2527, 2641, 2645, 2663, 2759, 2872, 2884, 2899]. clustering-based [1049]. clusters [155, 2320]. CMAC [2808]. CMIMO [2815]. CMT [2081]. Co [418, 1317, 1715, 1913, 2446, 2625]. co-based [418]. co-channel [2446]. co-exist [1317]. CO-OFDM [2625]. co-operative [1913]. Co-simulation [1715]. coal [1891, 2928]. coalition [1115, 1340, 2055, 2815]. coalitional [1031, 1209, 1444, 1576, 2634]. Coalitionally [1798]. coarse [135]. coarse-grained [135]. Cocoon [2051]. CoCoSpot [851]. Cod [1797]. code [977, 1580, 1830, 2153]. code-based [1830]. Coded [1301, 1782, 2892]. codes [447, 544, 748, 1005]. Coding [5, 17, 140, 342, 378, 403, 424, 431, 544, 582, 742, 780, 882, 887, 896, 918, 1046, 1097, 1142, 1366, 1376, 1381, 1437, 1523, 1552, 1555, 1585, 1646, 1668, 1762, 1782, 1788, 1795, 1810, 1814, 1870, 1903, 1905, 1907, 1911, 1967, 1987, 2015, 2040, 2088, 2126, 2263, 2386, 2511, 2631, 2704, 2828, 2945, 2971]. Coding-aware [1911]. coding-based [140, 887, 1646, 1870, 2040, 2828]. coding/ARQ [403]. Coercion [1498]. coevolution [1819]. Coexistence [1236, 1415, 1807, 2181, 2450, 2523, 2979]. coexisting [1023, 2774]. coflows [2877]. Cognitive [48, 73, 188, 189, 274, 345, 375, 439, 526, 576, 671, 677, 722, 725, 771, 797, 926, 937, 945, 1008, 1023, 1024, 1052, 1075, 1106, 1127, 1167, 1320, 1370, 1400, 1428, 1485, 1496, 1499, 1550, 1557, 1628, 1631, 1636, 1807, 1819, 1871, 1872, 1922, 2021, 2127, 2128, 2212, 2215, 2231, 2305, 2432, 2467, 2493, 2497, 2565, 2568, 2571, 2573, 2584, 2587, 2673, 2803, 2837, 2851, 2918, 2970]. collaboration [149, 984]. collaborations [2095]. Collaborative [148, 150, 153, 157, 158, 182, 481, 482, 549, 929, 1076, 1086, 1105, 1215, 1391, 1638, 1710, 1796, 1841, 2027, 2161, 2186, 2206, 2355, 2579, 2695, 2731, 2862]. Collaboratively [2397]. collapse [772]. Collateral [2005]. collecting [2273]. collection [912, 1173, 1187, 1301, 1341, 1368, 1649, 1926, 2200, 2298, 2586, 2718, 2843, 3002]. collectors [2843]. Collision [221, 953, 1103, 1984, 2079]. Collision-Aware [221]. collisions [904, 2165]. colluded [2142]. Collusion [501, 1882]. collusion-attack-resilient [1882]. Colony [606, 701, 1166, 2622]. coloring [2774]. combat [2876]. combination [100, 223, 1702, 1938, 2930]. Combined [535, 895, 903, 2856]. Combing [1762]. Combining [155, 1005, 1174, 1200, 2259, 2945]. command [851, 858]. commodity [2101]. common [1205, 1221, 1622, 1691]. common-friend [1221]. CommonFinder [1221]. communicate [279]. Communication [55, 97, 169, 353, 374, 434, 468, 470, 471, 476, 479, 481, 486, 494, 507, 630, 635, 653, 695, 699, 826, 862, 940, 964, 1022, 1025, 1085, 1103, 1238, 1240, 1253, 1270, 1281, 1310, 1312, 1353, 1356, 1555, 1624, 1707, 1801, 1907, 1980, 1994, 2068, 2107, 2119, 2178, 2187, 2222, 2225, 2235, 2313, 2349, 2357, 2439, 2446, 2542, 2618, 2651, 2796, 2799, 2872, 2874, 2890, 2963, 3004]. communication-oriented [55]. Communications [474, 729, 875, 982, 984, 1115, 1172, 1258, 1265, 1350, 1379, 1460, 1509, 1571, 1599, 1605, 1734, 1813, 1893, 1968, 2088, 2133, 2293, 2367, 2394, 2411, 2422, 2441, 2476, 2483, 2552, 2600, 2689, 2728, 2739, 2767, 2769, 2834, 2916]. communications-enabled [984]. communities [609, 833, 929, 1137, 1722]. Community [24, 229, 1349, 1721, 1723–1728, 1730, 1731, 1944, 2049, 2051, 2075, 2221, 2405, 2686, 2986]. community-aware [2075]. community-oriented [2051]. COMNET [697, 818, 1747, 2376, 2691]. CoMP [2212]. compact [2533]. Comparative [165, 749, 800, 905, 1552, 2390, 2964]. compare [1806]. Comparing [128, 346, 1441]. Comparison [176, 390, 444, 457, 522, 686, 1285, 1419, 2287]. compatible [1966, 1968, 2595]. compensation [869]. competing [1394, 2908]. competition [363, 1304, 1558, 1590, 1775, 2190]. competitive [2217]. compiler [1959]. Complete [409]. completion [491, 1771, 2877, 2987]. Complex [603, 606, 623, 750, 993, 1070]. Complexity [18, 1669, 2192, 2423, 2542]. compliant [1748, 3004]. component [676, 1039, 1361, 2029]. components [1655]. COMPOSER [2533]. composite [2218]. composition [323, 542, 987, 1193, 1249, 1703]. Comprehensive [875, 1333, 1469, 1558, 1878, 2160, 2216, 2430, 2613, 2909, 2965]. compressed [491, 2195]. Compressing [1919]. Compressive [1926, 1987, 2664, 2824]. comprising [2779]. compromised [2498, 2933]. computation [225, 496, 1076, 1532, 1966, 1988, 2064, 2360, 2977]. computational [2048]. compute [1101]. computed [2789]. Computer [42, 67, 71, 94, 491, 612, 699, 839, 845, 961, 1238, 1243, 1246, 1610, 1783, 2057, 2127, 2691, 2894, 2996, 2997]. computing [210, 1228, 1309, 1363, 1365, 1515, 1524, 1594, 1651, 1656, 1737, 1744, 1835, 1969, 2005, 2140, 2340, 2360, 2464, 2609, 2633, 2636, 2738, 2763, 2768, 2770, 2772, 2788, 2858, 2870, 2931, 2977, 2989, 3011]. Concealed [377, 1878]. concept [282, 790, 1519, 2163, 2924]. concepts [98]. concerns [1555]. concurrent [326, 354, 622, 666, 2081, 2628]. Condition [631, 1384, 1835, 2024, 2701]. condition-aware [2701]. conditional [2265, 2437]. Conditioning [1134]. conditions [296, 397, 1141, 1342, 2912]. conductance [1877]. conducting [1191, 1406]. CONFab [1361]. conferencing [935]. confidence [2731]. confidentiality [674, 1422, 2338]. configuration [6, 257, 373, 384, 474, 680, 708, 926, 1307, 1431]. configurations [762, 1021, 2755]. configuring [1060, 2744]. conflicts [985]. Congestion [39, 105, 120, 253, 255, 292, 293, 343, 390, 421, 564, 625, 630, 868, 900, 980, 1059, 1143, 1217, 1236, 1257, 1342, 1371, 1389, 1674, 1964, 2023, 2234, 2250, 2299, 2380, 2439, 2443, 2513, 2562, 2765, 2800, 2859, 2968]. Congestion-aware [2250]. Congestion-proportionality [2800]. Connected [352, 1201, 1386, 2224, 2476, 2566, 2573, 2734, 2753, 2759, 2773, 2996]. connectedness [1404]. connecting [732]. connection [138, 183, 731, 954, 1000, 1291]. connection-oriented [183, 954]. connectionless [550]. connections [265, 758, 998, 1523, 1743, 2653]. Connectivity [258, 281, 335, 455, 654, 686, 710, 942, 1388, 1628, 1751, 1787, 1960, 2053, 2301, 2302, 2313, 2318, 2382, 2581, 2801, 2926]. conquer [234, 921, 1089]. conscious [2233, 2240]. consensus [1338, 1467]. consensus-based [1338]. CONSERT [1752]. conservation [261, 2290]. consideration [2210]. Considerations [1330]. Considering [250, 986, 1603, 1712, 1744, 2026]. consistency [656, 1166, 1426, 2841]. consistent [1694, 2699, 2835]. consolidated [1228, 2626]. Consolidation [2252, 2337, 2381]. constant [2343]. constrained [225, 264, 316, 757, 1524, 1624, 1813, 2226, 2230, 2450, 2460, 2463, 2568, 2575, 2624]. Constraint [1011, 1457, 1657, 1677, 1993, 2150]. Constraint-based [1657]. constraints [73, 258, 289, 369, 561, 592, 703, 1369, 1384, 1425, 1572, 1860, 2072, 2138, 2711, 2768, 2863].

Constructing [585, 1752, 2586, 2659].

construction [488, 489, 798, 934, 1173, 1180, 1387, 2071, 2224]. consumer [1974]. consumers [933]. consumption [712, 738, 830, 1206, 1244, 1367, 1384, 1388, 1420, 1453, 1750, 2370, 2371, 2664, 2923]. contact [605, 1201, 2047, 2991]. contact-based [2047]. contacts [2991]. container [2988]. Containing [1372]. containment [963, 2345]. Content [134, 152, 157, 578, 593, 675, 681, 883, 951, 989, 1035, 1043, 1078, 1089, 1218, 1247, 1323, 1349, 1363, 1375, 1392, 1409, 1422, 1444, 1451, 1530, 1537, 1549, 1638, 1659, 1696, 1753, 1841, 1868, 1873, 1903, 1912, 1917, 1927, 1949, 1950, 1952, 2035, 2037, 2136, 2173, 2178, 2190, 2192, 2256, 2258, 2274, 2317, 2541, 2544, 2545, 2547, 2557, 2575, 2610, 2624, 2723, 2737, 2752, 2793, 2813]. Content-aware [2545]. content-based [134, 157, 883, 1078, 1549, 2737, 2752]. Content-centric [989, 1043, 1323, 1392, 1422, 1451, 1753, 2813]. Content-Delivery [1753, 2192]. content-dependent [1903]. content-oriented [1444]. content-sharing [951]. content/service [1363]. content/service-oriented [1363]. Contention [599, 624, 895, 942, 1129, 1205, 1637, 1675, 2079]. Contention-based [942]. ContentPlace [47]. Context [92, 208, 360, 716, 912, 1075, 1116, 1318, 1364, 1482, 1616, 1688, 2012, 2092, 2140, 2145, 2207, 2309, 2355, 2373, 2476, 2529, 2727, 2881]. Context-aware [92, 360, 716, 1318, 1364, 2092, 2140, 2145, 2309, 2355, 2476, 2727, 2881]. Context-awareness [1482]. context-based [2373]. contexts [1145]. Contextual [2151, 2646]. Contextualized [1525]. continuity [861, 1065]. Continuous [710, 2354, 2647, 2948, 2997]. Continuum [78]. contours [35]. contract [1719, 2706]. contract-based [1719]. contracting [627]. contribution [1614, 2472, 2736]. contributors [1775]. Control [3, 21, 39, 40, 62, 105, 109, 127, 133, 253, 255, 283, 292, 293, 343, 374, 387, 390, 396, 400, 403, 527, 580, 587, 597, 629, 630, 674, 681, 685, 689, 735, 755, 804, 836, 851, 858, 868, 890, 900, 906, 927, 940, 953, 1000, 1001, 1052, 1059, 1064, 1143, 1176, 1203, 1205, 1236, 1262, 1305, 1364, 1371, 1389, 1418, 1450, 1470, 1591, 1842, 1858, 1860, 1932, 1964, 2021, 2023, 2030, 2074, 2091, 2110, 2242, 2327, 2344, 2357, 2393, 2434, 2439, 2443, 2452, 2462, 2480, 2513, 2562, 2584, 2622, 2632, 2739, 2807, 2840, 2852, 2880, 2904, 2926, 2945, 2960, 2968, 3001, 3005]. control [120, 201, 375, 515, 535, 566, 598, 770, 841, 1124, 1217, 1463, 1514, 1632, 1852, 1894, 2076, 2234, 2266, 2422, 2465, 2495, 2715, 2765, 2787, 2800, 2923, 2928, 2930, 2952, 2954]. Controlled [914, 981, 1532, 1792, 1937, 2092, 2688]. Controller [1108, 1285, 1693, 2060, 2189, 2468, 2498, 2729, 2848, 2889, 2965, 2976, 3006]. controllers [449, 558, 1787, 2186, 2556, 2994]. Controlling [456, 962]. convergecast [310, 2667]. converged [357, 363, 1591, 2318]. convergence [44, 263, 356, 392, 621, 2380]. converging [308]. conversation [667]. conversational [2081]. conversion [56]. converter [721]. converters [693, 1329]. convincing [2171]. Convolutional [2957]. cookie [166, 655]. cooperated [999]. Cooperating [551]. cooperation [396, 802, 1348, 1498, 1537, 2190, 2231, 2317, 2587]. Cooperative [49, 84, 134, 169, 280, 423, 526, 616, 719, 774, 826, 941, 951, 1096, 1104, 1115, 1167, 1339, 1343, 1350, 1393, 1432, 1444, 1542, 1547, 1550, 1596, 1639, 1779, 1795, 1804, 1830, 1894, 1947, 2019, 2064, 2088, 2119, 2219, 2262, 2313, 2485, 2565, 2588, 2596, 2624, 2804, 2808, 2815, 2972]. Coordinated [49, 212, 1340, 1915, 2035, 2216, 2696]. coordinates [381, 1074]. Coordination [276, 1837, 1846, 2303, 2684]. copying [2265]. core [246, 1054, 1120, 1229, 1280, 1567, 2190, 2583, 2733, 2820]. core/edge [246].

CoreCast [246]. CORHYS [1499].

correction [1005, 1682, 1903]. correlated [897, 1521, 1749, 1807, 2000]. correlating [410, 1769]. Correlation [304, 907, 1027, 1315, 2990]. correlations [1109]. Correspondent [894]. corresponding [721]. Corrigendum [67, 1783, 2127, 2997]. corruption [743]. corruption-resilience [743]. Cost [258, 300, 344, 404, 548, 552, 670, 704, 757, 896, 952, 1000, 1053, 1294, 1329, 1390, 1534, 1578, 1607, 1650, 1680, 1722, 2015, 2059, 2072, 2075, 2170, 2202, 2207, 2251, 2370, 2401, 2463, 2603, 2635, 2907]. Cost-aware [344]. cost-effective [1053, 1680, 2170]. Cost-efficient [404, 1607]. cost-optimal [552]. cost-saving [1650]. cost/delay [1294]. costs [831]. count [581, 2935]. countermeasures [770, 783, 1014, 1555]. Counters [1859]. Coupled [1829, 2294, 2303]. coupling [2708]. Coverage [181, 250, 302, 352, 409, 455, 647, 669, 776, 979, 1388, 1533, 1857, 1866, 1940, 1954, 1972, 2078, 2135, 2228, 2282, 2410, 2465, 2539, 2751, 2897, 2967, 3001]. coverage-aware [647]. covert [340, 448, 864]. CPS [2444]. CPU [1748]. CR [459, 843, 2305]. CR-Chord [459]. CR-VANET [2305]. Cramér [2457]. crawling [811]. Creating [1159, 1178, 1277]. credential [1263]. credibility [2177]. credit [1051, 1379, 2109]. credit-based [1051, 2109]. Crisis [821]. criteria [579, 2053, 2883]. critical [179, 352, 409, 623, 728, 827, 1187, 1708, 2487]. critical-square-grid [409]. Criticality [89]. CRNs [774, 1369]. Cross [29, 86, 88, 287, 378, 417, 546, 559, 565, 572, 588, 645, 665, 766, 795, 838, 862, 948, 1017, 1090, 1096, 1359, 1408, 1425, 1511, 1577, 1585, 1652, 1804, 1852, 1853, 1959, 2291, 2404, 2451]. Cross-AS [2404]. Cross-cultural [1408]. Cross-layer [29, 88, 287, 378, 417, 559, 565, 572, 588, 645, 665, 838, 862, 948, 1090, 1096, 1359, 1425, 1577, 1585, 1652, 1852, 1853, 2451]. cross-platform [1959]. cross-tier [1017]. crosstalk [754]. crowd [1620, 2125, 2369, 2400, 2549, 2726]. crowds [1948]. crowdsensed [2719]. crowdsensing [2177, 2395, 2403, 2581, 2844, 2920]. Crowdsourced [1465, 1617, 1621, 1622]. Crowdsourcee [2879]. crowdsourcers [1775]. Crowdsourcing [1611, 1613, 1614, 1616, 1618, 1775, 1837, 1861, 2454]. CRT [2778]. crypto [1560]. crypto-less [1560]. cryptography [2321]. cryptosystem [1005]. CS [2664]. CSI [2347]. CSL [2696]. CSMA [25, 435, 734, 767, 1003, 1352, 2201, 2205, 2263, 2284]. CSMA-based [2284]. CSMA/CA [25, 734, 767, 1003, 1352]. CTAC [770]. CubeSats [2753]. CUBIC [724]. cuckoo [2135]. cultural [1408]. culture [286]. Current [108, 1139, 1231, 2713, 2973]. curve [2654]. custodian [2557]. custom [2867]. customer [534]. customer-managed [534]. customers [1246]. customization [1286]. cut [1238, 1442]. cut-through [1238]. cuts [2340]. Cutting [700]. Cyber [909, 961, 968, 1083, 1769, 1982, 2043, 2140, 2141, 2349, 2434, 2486, 2505, 2506, 2548, 2753, 2769, 2876, 2928]. Cyber-physical [1982, 2043, 2140, 2141, 2349, 2434, 2505, 2506, 2548, 2753]. cyber-physical-systems [2928]. cybercrime [1821]. cyberdefenses [1491]. Cybersecurity [2623, 3005]. cyberspace [2716]. cycle [57, 305, 895, 1282, 2436]. cycled [289, 1810, 2447, 2525]. cycles [102, 126, 224, 1440, 2520]. Cyclic [842]. cycling [1050].

D [582, 634, 1204, 1489, 2377, 2496, 2757, 2885, 2925]. D2D [1353, 1670, 1960, 2088, 2235, 2289, 2422, 2578, 2656, 2682, 2979]. DACCER [993]. DAFNES [1765]. damage [2005]. DARE [1032]. DASH [2821]. DASH-QoS [2821]. Data [13, 47, 127, 131, 269, 552, 623, 630, 678, 723, 724, 795, 831, 832, 865, 897, 902, 911, 938, 994, 1027, 1040, 1044, 1049, 1055, 1060, 1071, 1073, 1177, 1179, 1208, 1213, 1216, 1232, 1267, 1268, 1282, 1337, 1368, 1395, 1437, 1490, 1507, 1562, 1578, 1591, 1600, 1637, 1641, 1654, 1655, 1660, 1716, 1717, 1748, 1797, 1893, 1902, 1909, 1918, 1934, 1983, 2010, 2018, 2020, 2061, 2079, 2080, 2100, 2152, 2187, 2279, 2288, 2312, 2364, 2365, 2369, 2380, 2383, 2413, 2517, 2530, 2597, 2626, 2698, 2714, 2743, 2807, 2818, 2828, 2846, 2849, 2851, 2871, 2875, 2891, 2901, 2914, 2933, 2940, 3003, 3013]. data [63, 212, 219, 298, 332, 377, 437, 446, 498, 614, 620, 673, 683, 714, 772, 893, 929, 933, 1018, 1041, 1088, 1090, 1113, 1118, 1131, 1148, 1226, 1273, 1301, 1341, 1366, 1384, 1398, 1407, 1426, 1472, 1473, 1483, 1534, 1579, 1592, 1649, 1656, 1676, 1691, 1775, 1795, 1842, 1846, 1868, 1878, 1904, 1917, 1926, 1928, 1929, 1969, 1977, 1988, 2002, 2048, 2055, 2070, 2075, 2110, 2154, 2172, 2227, 2252, 2253, 2327, 2328, 2332, 2361, 2381, 2394, 2440, 2460, 2484, 2524, 2552, 2561, 2582, 2620, 2653, 2672, 2695, 2718, 2719, 2735, 2779, 2816, 2824, 2835, 2843, 2844, 2900, 2921, 2946, 2952, 2990, 2991, 3002]. data [261, 408, 431, 528, 1032, 1321, 1839, 1934, 2040, 2069, 2083, 2143, 2177, 2179, 2211, 2306, 2586, 2607, 2628, 2643, 2884, 2955, 2973]. data-center [2061]. Data-centric [13, 552, 1893]. data-control [2952]. Data-Gathering [528]. data-plane [2394]. database [572]. datacenter [1257, 1270, 1493, 1642, 1771, 2085, 2207, 2234, 2290, 2593, 2921, 2987]. datacenters [1273, 2063, 2877]. date [2058]. DBA [2993]. DBit [1941]. DBridges [1067]. DBStream [1935]. DCB [2809]. DCF [128, 198, 566]. DDoS [203, 778, 1515, 1978, 2005, 2137, 2486, 2864, 2995]. Deadline [138, 1267, 1396, 1443, 1795, 1925, 2306, 2327, 2575, 2816]. Deadline-aware [1795]. Deadline-based [138, 1443]. Deadline-constrained [2575]. deaggregation [1506]. deal [753]. Dealing [1865]. Decapitation [260]. decays [1140]. December [2974, 2998]. decentralised [151, 209]. Decentralized [1, 67, 152, 213, 1221, 1241, 1401, 1604, 1797, 1857, 2312, 2720]. Decentralizing [2594]. decision [58, 142, 576, 633, 1064, 1443, 1594, 1631, 1824, 2203, 2323, 2407, 2456, 2883]. decision-making [633, 2456]. decision-tree [2203]. decisions [361, 1920]. Declarative [542]. decompositions [1741]. Decoupling [154, 321]. dedicated [952, 2277, 2881]. Deducing [1352]. Deep [370, 2406, 2482, 2779, 2789, 2826, 2955, 3007]. defence [530]. defending [234]. Defense [200, 260, 412, 1998, 2970]. defenses [2696]. deficit [500]. DEFIDNET [1491]. Defined [1128, 1161, 1183, 1245, 1271, 1276, 1328, 1402, 1433, 1502, 1515, 1548, 1564, 1580, 1680, 1685, 1687, 1695, 1697, 1698, 1700, 1709, 1729, 1789, 1832, 1859, 1862, 1867, 1932, 1956, 1996, 2006, 2033, 2060, 2076, 2096, 2107, 2134, 2172, 2197, 2200, 2243, 2247, 2248, 2333, 2402, 2420, 2429, 2458, 2468, 2473, 2498, 2500, 2507, 2556, 2616, 2640, 2644, 2678, 2679, 2686, 2698, 2722, 2729, 2748, 2750, 2801, 2825, 2841, 2848, 2857, 2886, 2889, 2914, 2933, 2944, 2965, 2966, 2976, 2982, 2994]. Definition [2017, 2950]. degradation [2184]. degree [1767, 2927]. degrees [731]. Delay [25, 120, 145, 226, 253, 311, 379, 418, 449, 494, 619, 888, 904, 925, 943, 949, 1101, 1130, 1134, 1143, 1217, 1260, 1294, 1457, 1552, 1578, 1580, 1624, 1683, 1892, 1950, 2031, 2034, 2045, 2079, 2080, 2201, 2205, 2239, 2384, 2436, 2450, 2460, 2568, 2611, 2640, 2711, 2852, 2863, 2923, 2993]. Delay-Aware [925, 1892, 2611]. Delay-based [120, 1143]. Delay-constrained [1624, 2450, 2460, 2568]. delay-sensitive [2852]. Delay-tolerant [418, 494, 1294, 1552, 1950, 2239]. Delayed [611]. delays [83]. delegation [1511]. deletions [632]. delivering [782]. Delivery [269, 289, 446, 452, 498, 883, 929, 994, 1035, 1055, 1276, 1349, 1395, 1494, 1577, 1659, 1753, 1912, 1927, 2027, 2136, 2192, 2218, 2272, 2547, 2624, 2626, 2644, 2814, 2827, 2923, 2946]. Demand [31, 63, 281, 1277, 1380, 1478, 1653, 1930, 1974, 2037, 2145, 2270, 2439, 2450, 2465, 2558, 2828, 2845]. Demand-oblivious [2558]. demanding [550]. demands [252, 453, 663, 1337, 1738, 1803, 1927]. demonstration [1164]. Denial [328, 783, 1792, 2492, 2696]. Denial-of-service [783]. Denial-of-sleep [2696]. DeNoise [124]. DeNoise-and-Forward [124]. dense [78, 229, 713, 833, 1541, 1758, 2119, 2175, 2368, 2441]. densely [624]. Density [126, 503, 1180, 1459, 1989, 2060, 2455, 2784, 2790, 3001]. Density-aware [3001]. density-barrier [1180]. Dependability [1548, 1589]. dependable [243, 1996]. dependencies [2341]. dependent [33, 195, 546, 1594, 1903, 2395, 2521, 2799, 2877]. Deploy [2952]. deployable [1030]. deployed [624]. Deploying [470, 2812]. deployment [75, 250, 258, 482, 903, 1133, 1154, 1361, 1461, 1484, 1584, 1786, 1848, 1889, 1972, 2026, 2141, 2371, 2410, 2619, 2638, 2823, 2845, 2925, 3015]. deployments [1161, 1455, 1826]. depth [950]. Derivation [176]. derived [273]. description [887, 2124]. Design [13, 47, 57, 87, 98, 107, 123, 210, 221, 255, 259, 261, 270, 287, 316– 318, 320, 321, 332, 348, 349, 369, 400, 404, 415, 446, 449, 478, 510, 587, 670, 674, 707, 720, 727, 751, 755, 756, 779, 849, 857, 888, 931, 1009, 1057, 1114, 1133, 1140, 1146, 1160, 1175, 1219, 1242, 1260, 1272, 1305, 1329, 1359, 1425, 1461, 1531, 1547, 1585, 1641, 1663, 1710, 1767, 1814, 1835, 1838, 1852, 1858, 1859, 1983, 1987, 1996, 2023, 2035, 2073, 2150, 2162, 2189, 2265, 2329, 2372, 2413, 2571, 2618, 2625, 2867, 2971, 3006]. Designing [454, 865, 1159, 1192, 1483, 2553]. designs [1050, 2635, 3012]. desktop [1272]. destination [739, 1169, 2285]. destructing [2553]. Detect [1178, 1317, 1634, 2142]. Detecting [348, 564, 593, 753, 848, 855, 863, 1083, 2090, 2183, 2475, 2498, 2671, 2757]. Detection [86, 99, 135, 174, 305, 340, 370, 371, 485, 496, 549, 583, 737, 752, 774, 778, 845, 853, 854, 864, 874, 878, 983, 999, 1040, 1058, 1069, 1081, 1092, 1105, 1125, 1148, 1174, 1210, 1296, 1308, 1410, 1479, 1491, 1527, 1554, 1573, 1579, 1626, 1641, 1662, 1793, 1821, 1866, 1883, 1944, 1965, 1971, 1979, 2003, 2006, 2017, 2077, 2116, 2137, 2174, 2185, 2297, 2350, 2361, 2453, 2478, 2481, 2482, 2499, 2542, 2617, 2668, 2705, 2756, 2775, 2779, 2789, 2792, 2811, 2823, 2856, 2864, 2865, 2887, 2909, 2919, 2927, 2961, 2995]. detection-resistant [340]. detector [2142]. detectors [1938]. Determination [897, 1725, 1847]. Determining [2305]. Deterministic [466, 696, 1911, 2173]. Development [472, 1196, 1305]. Deviation [1239]. Device [1480, 1552, 1593, 1931, 1984, 2162, 2178, 2317, 2318, 2326, 2394, 2411, 2552, 2715, 2738, 2739, 2784, 2834]. device-aware [1984]. Device-to-Device [2162, 2178, 2317, 2394, 2411, 2552, 2739, 2784, 2834]. devices [398, 575, 668, 1005, 1131, 1280, 1393, 1684, 2000, 2084, 2104, 2187, 2326, 2520, 2575, 2716, 2724, 2795, 2977]. dew [2738]. DHT [459, 613, 2320]. DHT-paradigm [2320]. DHTs [155]. Di-Sense [2767]. diagnosis [1347, 1525, 2432]. diagrams [2885]. difference [917, 2932]. differences [1408, 1941]. different [362, 444, 687, 1534, 2210]. Differential [48, 693, 1337, 2954]. Differentially [2598]. Differentially-private [2598]. differentiated [486, 677, 822, 1770, 2908]. differentiation [107, 244, 433, 525, 1169, 1763, 2306, 2821]. difficulty [479]. Diffusion [213, 353, 1014, 2116, 2298, 2773, 2844]. diffusion-based [353]. DiFS [1909]. digest [350]. Dilemma [802]. dimensional [1088, 1133, 1592, 1782, 2564, 2906]. Dimensional-IP [1133]. Dimensionality [1626, 2705]. dimensioning [1264, 2010, 2912]. DIMR [1645]. direct [454, 1511, 1992, 2369, 2439, 2784, 3013]. direct-sequence [1992]. direction [260, 2559]. Directional [173, 1474, 1535, 1727, 1972, 2344, 2417]. directions [2337, 2930]. directories [59]. disaggregation [2144]. disaster [773, 1628, 1736, 2236, 2439, 2698]. disaster-affected [2698]. disaster-resilient [1736]. disciplines [97]. disclosure [777, 1594]. Discontinuous [1333]. discovery [9, 158, 179, 330, 341, 648, 1132, 1173, 1256, 1465, 1530, 1587, 1593, 1820, 1931, 2100, 2369, 2485, 2706, 2719, 2750, 2784, 2805, 2832]. discrete [188, 329, 2246]. discrete-time [188, 329, 2246]. Discriminating [497]. Discriminatory [116]. disjoint [585, 732, 1645, 2121]. disk [1442, 1749, 1918]. disorder [1947]. disorder-avoidance [1947]. dispatching [1079]. dispersion [329]. disruption [1346, 1580]. disruption-tolerant [1346]. disruptive [2653]. Dissecting [849]. dissemination [47, 125, 261, 623, 694, 723, 1026, 1218, 1366, 1398, 1654, 2143, 2369, 2390, 2539, 2914]. Distance [724, 789, 2270, 2611]. Distance-vector [2270]. distinctive [2390]. distort [104]. distraction [2605]. Distributed [64, 79, 133, 193, 222, 247, 254, 273, 275, 308, 328, 330, 350, 382, 429, 540, 562, 563, 600, 648, 735, 739, 764, 917, 954, 973, 979, 1010, 1044, 1052, 1057, 1069, 1083, 1087, 1092, 1095, 1119, 1138, 1153, 1197, 1214, 1221, 1226, 1250, 1296, 1307, 1340, 1364, 1380, 1451, 1467, 1469, 1492, 1499, 1524, 1534, 1550, 1574, 1636, 1641, 1653, 1659, 1715, 1756, 1758, 1759, 1765, 1781, 1787, 1797, 1807, 1909, 1911, 1976, 1978, 2065, 2075, 2089, 2102, 2118, 2189, 2191, 2224, 2258, 2262, 2302, 2343, 2455, 2468, 2495, 2496, 2564, 2642, 2659, 2783, 2841, 2875, 2919, 2928, 2939, 2984, 2991]. Distributed [993]. Distributing [1409]. Distribution [127, 186, 197, 339, 431, 555, 642, 675, 755, 916, 1019, 1058, 1123, 1150, 1537, 1638, 1830, 1841, 1949, 1988, 2156, 2317, 2450, 2638, 2728, 2747]. Distribution-based [1058]. distributions [2433, 2927]. distrust [737]. diurnal [1148]. diverse [186, 1072, 2272]. diversified [2647]. diversity [187, 334, 375, 514, 1522, 2105, 2849, 2916]. diversity-based [2849]. divide [234, 1089]. divide-and-conquer [1089]. divide-conquer-scanning [234]. division [763, 1694, 2221]. DMM [2416]. DMMS [1756]. DNS [539, 589, 1793, 1821, 2008, 2131]. DNStamp [1211]. do [784]. DOCSIS [1755]. documents [387]. DODS [2919]. Domain [127, 246, 322, 443, 827, 841, 884, 886, 996, 1109, 1128, 1207, 1481, 1488, 1511, 1527, 1613, 1798, 1869, 1938, 2020, 2283, 2337, 2349, 2532, 2540, 2604, 2953]. domain-based [2604]. domain-specific [1613]. domains [1027, 1385, 1690]. dominated [1386]. Dominating [303, 2072, 2224, 2773]. domination [299]. Don’t [1405]. Double [102, 945, 1122, 2154, 2697, 2834]. double-link [1122]. Double-Ring [102]. double-ruling [2154]. down [6, 1255, 2567]. Downlink [372, 866, 929, 930, 932, 1024, 1417, 1603, 1664, 1759, 1770, 2165, 2543, 2654, 2658, 2882]. download [672]. downloading [574, 941, 1841, 1947, 2650]. DPI [1419, 1676]. DPillar [683]. drive [941, 1398]. drive-thru [941]. driven [10, 212, 425, 528, 1062, 1496, 1503, 1712, 2092, 2372, 2530, 2833]. driver [473, 2605]. driving [2641]. drone [2639]. drone-enabled [2639]. drop [915]. dropping [1169, 1579, 1800, 2185]. DRUID [318]. DRX [1333, 1440, 2744]. DSR [173]. DT [2272]. DT-RPL [2272]. DTN [694, 1540, 2336]. DTN-based [2336]. DTNs [2991]. dual [102, 683, 1601, 2039, 2926, 2952, 2957]. dual-factor [2957]. Dual-port [683]. dual-reinforcement-learning [1601]. duplex [948, 2344, 2904]. duplicates [2160]. durable [1539]. duration [1201, 2584]. during [1521]. duty [289, 895, 1050, 1282, 1810, 2436, 2447, 2525]. duty-cycled [289, 2447, 2525]. DVB [565]. DVB-S2 [565]. DVB-S2/ETSI [565]. DWDM [238, 756]. Dynamic [5, 8, 164, 181, 194, 222, 257, 298, 386, 388, 413, 416, 439, 525, 542, 554, 575, 603, 605, 609, 633, 659, 677, 706, 721, 725, 736, 758, 763, 780, 791, 819, 828, 906, 926, 980, 1000, 1006, 1024, 1090, 1116, 1138, 1219, 1252, 1262, 1319, 1357, 1366, 1383, 1389, 1443, 1463, 1469, 1498, 1531, 1557, 1566, 1601, 1637, 1701, 1703, 1755, 1766, 1826, 1875, 1921, 1958, 1967, 1977, 2089, 2146, 2194, 2208, 2226, 2252, 2253, 2264, 2279, 2315, 2382, 2407, 2510, 2528, 2536, 2537, 2543, 2549, 2583, 2585, 2623, 2626, 2700, 2702, 2816, 2832, 2888, 2903, 2929, 2939]. dynamic [318, 540, 1493, 1738, 1951, 2316, 2586]. dynamic-alternate [721]. Dynamic-cost-reward [1000]. dynamic-identity [5]. Dynamical [223]. dynamics [120, 201, 606, 1178]. DynMAC [1415].

e-assessment [2910]. e-health [2881, 2940]. e-healthcare [2738]. E2E [726]. E2E/ERN [726]. EA [1407]. EA-based [1407]. EAODR [2336]. EAP [197, 1387]. EAP-based [197, 1387]. early [752, 2100, 2818]. Earth [1969]. east [2765]. EasyGo [2603]. EC2 [1735]. ECC [2102, 2328]. ECC-based [2102, 2328]. ECDS [1841]. ECN [39, 421, 1975]. ECN-nonce [39]. ECN/RED [1975]. ECO [692, 2664]. ECO-ALOC [692]. Economic [40, 196, 1179, 1506, 1792, 2458]. economics [260, 815, 1848]. economy [286, 2194]. ecosystem [363, 1123, 2318]. ecosystems [2706]. EDCA [167, 296]. Edge [246, 886, 1157, 1339, 1441, 1988, 2206, 2283, 2360, 2383, 2387, 2695, 2772, 2788, 2810, 2858, 2862, 2870, 2931, 2977, 2986, 2989]. edge-as-a-service [2810]. Edge-aware [2283]. edge-computing [2870]. edge-intelligence [2387]. Editorial [2, 14, 28, 41, 53, 68, 81, 95, 103, 118, 132, 147, 148, 160, 175, 191, 206, 207, 215, 232, 251, 277, 288, 314, 327, 355, 368, 395, 419, 432, 441, 451, 469, 470, 480, 508, 524, 536, 537, 570, 602, 612, 617, 641, 651, 664, 679, 690, 698, 699, 715, 730, 745, 759, 775, 788, 794, 806, 817, 818, 844, 845, 859, 876, 892, 910, 924, 936, 947, 960, 961, 1034, 1099, 1414, 1518, 1602, 1747, 1946, 2057, 2376, 2414, 2431, 2474, 2488, 2504, 2522, 2534, 2691, 2712, 2763]. Editorial [975, 988, 1002, 1016, 1033, 1278, 1289, 1313, 1322, 1354, 1362, 1374, 1403, 1412, 1434, 1448, 1449, 1487, 1497, 1517, 1529, 1556, 1563, 1569, 1575, 1582, 1623, 1673, 1686, 1704, 1732, 1733, 1745, 1774, 1784, 1790, 1799, 1808, 1818, 1833, 1850, 1863, 1880, 1896, 1914, 1933, 1945, 1962, 1990, 2001, 2013, 2014, 2042, 2056, 2082, 2103, 2117, 2129, 2139, 2148, 2157, 2166, 2180, 2196, 2214, 2229, 2244, 2260, 2281, 2307, 2325, 2351, 2374, 2375, 2388, 2399, 2551, 2570, 2591, 2612, 2630, 2652, 2670, 2690, 2721, 2742, 2764, 2785]. Editorial [509, 1151, 1182, 2786, 2806, 2819, 2829, 2839, 2847, 2861, 2878, 2893, 2911, 2937, 2959, 2975, 2999]. EDoS [1792]. education [808]. Effect [250, 354, 1302, 2175, 2625, 2675, 2908]. Effective [244, 447, 678, 761, 824, 854, 889, 999, 1053, 1174, 1199, 1479, 1557, 1680, 1848, 1905, 1963, 2020, 2170, 2348, 2809, 2830, 2891]. effectiveness [525, 1021, 1047, 2960]. Effects [463, 734, 2495]. effici [2344]. effici-ent [2344]. efficiencies [1677]. Efficiency [705, 752, 903, 904, 1062, 1101, 1124, 1231, 1255, 1454, 1455, 1463, 1504, 1538, 1543, 1977, 2096, 2162, 2264, 2355, 2574, 2596, 2734, 2741, 2758, 2787, 2838, 2882, 2890, 2899]. Efficiency-driven [1062]. Efficient [37, 111, 129, 139, 173, 185, 211, 269, 301, 372, 454, 461, 481, 489, 514, 530, 574, 652, 681, 689, 692, 695, 703, 727, 764, 780, 782, 860, 886, 925, 939, 941, 956, 992, 1011, 1050, 1055, 1120, 1133, 1149, 1168, 1226, 1281, 1350, 1359, 1386, 1410, 1451, 1479, 1484, 1531, 1536, 1559, 1600, 1625, 1748, 1754, 1787, 1800, 1822, 1841, 1869, 1876, 1948, 1971, 2022, 2033, 2049, 2149, 2152, 2158, 2235, 2253, 2265, 2273, 2282, 2295, 2381, 2426, 2436, 2441, 2451, 2480, 2517, 2523, 2656, 2669, 2739, 2744, 2759, 2823, 2828, 2832, 2875, 2935, 2942, 2983, 2985, 2987, 3011]. efficient [5, 65, 117, 125, 155, 163, 228, 240, 268, 280, 312, 334, 386, 404, 413, 516, 553, 558, 620, 701, 704, 707, 934, 994, 999, 1015, 1022, 1026, 1061, 1066, 1103, 1115, 1132, 1264, 1287, 1291, 1293, 1300, 1365, 1382, 1396, 1398, 1461, 1462, 1471, 1483, 1560, 1607, 1649, 1707, 1758, 1760, 1819, 1867, 1870, 1872, 1926, 1928, 1980, 1988, 2031, 2073, 2091, 2108, 2125, 2162, 2187, 2208, 2213, 2222, 2252, 2262, 2330, 2391, 2406, 2424, 2437, 2531, 2539, 2563, 2579, 2674, 2687, 2707, 2732, 2740, 2761, 2774, 2780, 2805, 2815, 2827, 2842, 2843, 2898, 2918, 2946, 2947, 2951, 2971, 2977, 2989, 3002]. efficient [123, 177, 261, 406, 492, 598, 708, 811, 827, 1290, 1381, 1387, 1892, 2040, 2064, 2150, 2182, 2241, 2248, 2367, 2520, 2804]. efficiently [972]. effort [999, 1038, 2016]. efforts [1488]. EFMMRP [2150]. egoist [784]. egress [1661]. egression [1438]. eHealth [2354]. EHU [1189]. EIR [2283]. Elastic [88, 869, 906, 1475, 1549, 1586, 1667, 1674, 1737, 1906, 2138, 2149, 2625, 2681, 2993]. Electric [2515, 2601]. electro [1714]. electromagnetic [1750, 2371]. element [1966]. Elementary [1076]. elements [2460]. elephant [1697, 2179]. elephants [401, 753]. elimination [590]. Elisa^{T M} [476]. elliptical [1210]. Elman [2780]. Elsevier [42]. email [1125]. embedded [977]. Embedding [660, 898, 1171, 1607, 1633, 1692, 1738, 1741, 1742, 1768, 1924, 1928, 1977, 2062, 2188, 2252, 2592, 2662, 2932]. embryogenic [74]. emergency [1145, 1801, 2836]. emerging [365, 786, 1139, 2308, 2322, 2363]. EMGR [2330]. emission [122, 2632]. Emitter [1597]. emotion [2605]. Empirical [1125, 2298]. Employing [2460]. emulation [774, 1690, 2970]. emulator [1676]. Enable [334, 726, 1286]. Enabled [110, 685, 727, 756, 984, 1266, 1463, 1849, 2419, 2433, 2443, 2520, 2590, 2592, 2601, 2639, 2748, 2788, 2810, 2818, 2851, 2892, 2905, 2939, 2985]. Enabling [134, 149, 209, 210, 1196, 1565, 1591, 1593, 1702, 1916, 1988, 2050, 2133, 2159, 2479, 2629, 2666, 2679, 2706, 2790, 2817, 2818, 2986]. eNB [682, 1484]. encapsulation [2479]. encoding [370, 1039, 1112]. encourage [784]. encrypted [347, 3007]. encryption [483, 2329, 2425, 2426, 2548]. End [107, 164, 227, 374, 390, 546, 654, 735, 824, 1005, 1046, 1111, 1186, 1292, 1625]. End-to-end [107, 164, 227, 374, 390, 546, 735, 824, 1046, 1111, 1186, 1292, 1625]. End-user [654]. Endeavouring [1540]. Endurance [1070]. Energy [163, 211, 213, 240, 280, 300, 312, 406, 454, 484, 490, 516, 598, 668, 700, 703, 705, 707, 709, 713, 764, 780, 796, 903, 919, 925, 949, 954, 995, 1055, 1060, 1066, 1101, 1115, 1120, 1124, 1149, 1168, 1229, 1231, 1255, 1281, 1295, 1297, 1359, 1365, 1390, 1393, 1452, 1455, 1459, 1461– 1463, 1599, 1705, 1738, 1739, 1748, 1760, 1764, 1819, 1822, 1872, 1889, 1894, 1926, 2032, 2073, 2084, 2091, 2108, 2121, 2155, 2230, 2253, 2282, 2295, 2330, 2367, 2391, 2424, 2509, 2520, 2531, 2563, 2579, 2581, 2601, 2656, 2739, 2781, 2799, 2815, 2843, 2882, 2898, 2918, 2977, 3011]. energy [217, 261, 268, 290, 413, 592, 620, 670, 689, 692, 701, 702, 706, 708, 752, 860, 895, 934, 939, 1004, 1022, 1029, 1103, 1244, 1290, 1293, 1367, 1368, 1382, 1396, 1420, 1451, 1456, 1457, 1464, 1504, 1650, 1677, 1754, 1765, 1796, 1892, 2026, 2031, 2086, 2087, 2096, 2125, 2144, 2187, 2212, 2213, 2251, 2252, 2262, 2265, 2355, 2368, 2370, 2393, 2408, 2436, 2441, 2446, 2471, 2487, 2494, 2495, 2564, 2565, 2571, 2587, 2606, 2620, 2664, 2674, 2741, 2755, 2810, 2827, 2838, 2851, 2880, 2923, 2942, 2951, 3002]. Energy-Aware [484, 668, 706, 919, 939, 954, 995, 1029, 1295, 1368, 1456, 1739, 1764, 2084, 2155, 2880]. Energy-balanced [1894]. energy-based [213]. energy-critical [2487]. energy-delay [949]. energy-efficiency [1504]. Energy-Efficient [163, 211, 312, 406, 516, 598, 620, 689, 692, 708, 764, 780, 934, 1066, 1120, 1293, 1359, 1365, 1382, 1451, 1461, 1462, 1754, 1822, 1872, 1892, 1926, 2073, 2108, 2125, 2252, 2253, 2282, 2295, 2330, 2367, 2391, 2424, 2436, 2441, 2520, 2531, 2563, 2674, 2739, 2843, 2898, 2951, 3002, 3011]. energy-free [2408]. energy-harvesting [1004]. Energy-optimal [1459]. Energy-performance [713]. energy-QoE [2212]. Energy-relevant [1452]. Energy-saving [709, 1060, 2393]. Energy-sustainable [1889]. enforce [1422]. enforced [2606]. Enforcement [1973, 2033, 2208, 2226, 2594, 2917]. Enforcing [39, 2514]. engagement [1630, 2280]. engine [151, 809, 1959, 2273, 2649]. Engineering [233, 550, 557, 850, 886, 919, 954, 1232, 1300, 1314, 1339, 1566, 1835, 2106, 2253, 2275, 2540, 2825]. engines [104, 387]. English [1406]. Engset [1674]. enhance [426, 738, 840, 2095, 2901]. Enhanced [262, 301, 625, 739, 826, 916, 1233, 1287, 1370, 1757, 2536, 2544, 2600, 2808, 2899, 2957, 2972]. Enhancement [529, 615, 838, 990, 1245, 1907, 2028, 2215, 2814, 2838, 2857]. enhancements [170, 618]. Enhancing [350, 590, 1043, 1208, 1220, 1474, 1725, 1864, 1992, 2277, 2784, 2933]. eNodeB [1397]. enrichment [2871]. Ensemble [2287, 2705]. ensure [1625]. ent [2344]. Enterprise [567, 1081, 1125, 1411, 2114, 2472]. enterprise-like [567]. enterprises [964, 1620]. entries [2197]. Entropy [499, 1058, 1881, 1884, 2481]. Entropy-based [1884]. ENTRUST [2032]. entry [553]. enumerator [591]. envelope [2654]. environment [33, 84, 113, 117, 495, 624, 626, 955, 986, 1045, 1186, 1367, 1756, 2514, 2610, 2636, 2694, 2810, 2887, 2916]. environment-aware [113]. Environmental [714, 2361, 2990]. Environmental-aware [714]. environments [63, 151, 211, 408, 415, 430, 687, 891, 901, 1043, 1146, 1159, 1174, 1222, 1276, 1462, 1576, 1737, 1739, 1742, 1913, 2204, 2297, 2342, 2378, 2519, 2588, 2695, 2943, 2966]. EODL [2564]. EPC [1710]. ephemeral [152, 337]. epidemic [1877, 3010]. epidemics [1877]. epidemiological [847]. epilepsy [2432]. EPLC [2808]. EPON [196]. EPONs [1219]. equal [2300]. equal-split [2300]. equalization [1794]. Equilibria [78, 888, 1261]. equilibrium [1013, 1807]. equipment [866]. era [1515, 2002]. Erasure [1437, 1797]. ergodic [116]. Erlang [1674]. ERN [726]. Erratum [94, 697]. Error [182, 223, 403, 718, 1005, 1682, 1903, 2078, 2447]. error-control [403]. error-prone [2447]. errors [312, 504, 512]. ERSS [1381]. ESC [1560]. establishment [385, 1215, 1544, 2000, 2724]. Estimating [59, 329, 731, 1216]. Estimation [76, 83, 85, 195, 265, 311, 386, 399, 421, 491, 502, 724, 792, 800, 823, 878, 1173, 1287, 1351, 1583, 1610, 1986, 2161, 2377, 2387, 2433, 2455, 2471, 2519, 2615, 2767, 2888, 2969, 3001]. estimator [2510]. Ethernet [6, 7, 398, 445, 708, 796, 822, 860, 2213, 2380]. ETMRM [2531]. ETSI [565]. Europe [471]. European [1160]. EVaaS [2601]. evacuation [2698, 2836]. evades [864]. Evading [174]. Evaluating [295, 448, 573, 638, 666, 667, 1032, 2011]. Evaluation [47, 55, 112, 123, 129, 242, 264, 364, 504, 505, 522, 533, 566, 595, 645, 648, 756, 779, 836, 913, 931, 965, 1139, 1183, 1185, 1196, 1213, 1234, 1399, 1440, 1469, 1514, 1526, 1589, 1596, 1618, 1642, 1709, 1711, 1726, 1779, 1829, 1862, 1886, 1942, 1983, 2007, 2035, 2047, 2181, 2223, 2257, 2259, 2298, 2304, 2421, 2438, 2501, 2502, 2574, 2605, 2617, 2623, 2704, 2777, 2831, 2876, 2879, 2943, 2968, 2988, 3014]. evaluations [687]. Event [10, 268, 425, 486, 1341, 1379, 1971, 2223, 2313, 2447]. event-aware [268]. event-based [1379]. event-driven [425]. event-triggered [2447]. events [2017, 2100, 2137]. Everything [2600]. Evidence [397, 1405]. evolution [59, 247, 693, 812, 1051, 1247, 1329, 1422, 1690, 1770, 1835]. Evolutionary [74, 75, 496, 763, 888, 2333, 2509]. Evolved [1054]. evolving [1753]. ex [2323]. ex-ante [2323]. exact [225, 1009, 2463, 2702]. examine [802]. exchange [117, 494, 541, 744, 1159, 1795, 1838, 2939]. exclusive [1069]. execution [9, 462, 1025, 1959]. exist [1317]. existing [311, 977, 1054]. ExMin [1130]. expanding [587, 2339, 2820]. Expected [392, 1985]. Expediting [1811]. Experience [320, 1155, 1162, 1275, 1601, 1853, 2009, 2109, 2337, 2580]. experience-aware [1601, 1853]. experiment [1154]. Experimental [292, 397, 458, 645, 756, 836, 881, 1159, 1183, 1189, 1193, 1528, 1751, 1794, 2194, 2483, 2988]. Experimentation [1155, 1158, 1165, 1186, 2023, 2267, 2688]. experimentations [2163]. experimenting [2449]. experiments [1042, 1110, 1152, 1184, 1191, 1192, 1197, 1552, 1618, 1937]. expert [216]. explicit [2544]. exploit [514]. exploitation [967]. Exploiting [203, 620, 990, 2000, 2256, 2387, 2501, 2560, 2935]. exploration [208, 1473]. exploratory [1408]. Explore [2218]. Exploring [2599]. exponentially [2991]. exposed [767]. exposure [1750]. expression [492]. expressive [2426]. Extended [637, 885, 1392, 1440]. extension [891, 1580, 2270]. Extensions [295, 2203]. Extensive [2968, 3010]. external [1727]. Extra [800, 943]. Extracting [578, 1315]. extraction [850, 1022, 1075, 1471, 1621, 2158]. extreme [1986, 2731].

F [2654]. F-OFDM [2654]. fabrics

[1329, 1642, 1902]. face [249]. facets [2070]. facilitate [137, 1005, 1571]. facilities [1163, 1191]. facility [1159, 1160, 1189]. factor [255, 1334, 1838, 2455, 2957]. factories [1277]. factors [2992]. factory [1846]. fading [696, 899, 1432, 2783]. failover [11].

Failure [224, 305, 417, 585, 854, 1070, 1351, 1442, 1525, 1749, 1868, 2089, 2313, 2317, 2722, 2755, 2876, 2951]. failure-aware [1868]. failures [1100, 1117, 1122, 1520, 1521]. Fair [36, 176, 422, 532, 541, 640, 751, 868, 906, 937, 951, 1169, 1378, 1534, 1697, 1873, 2255, 2450]. Fairness [105, 308, 380, 383, 834, 1056, 1131, 1169, 1176, 1335, 1454, 1545, 1546, 1647, 1708, 1827, 2031, 2332, 2342, 2503, 2523]. Fairness-related [834]. Falcon [2954]. false [245, 1112, 2361, 2621]. family [971, 1930]. FANET [2963]. FANETs [1535]. Farewell [2352]. farm [328]. FaRNet [1088]. Fast [76, 100, 144, 197, 305, 390, 430, 445, 463, 492, 514, 594, 776, 853, 1088, 1290, 1301, 1395, 1474, 1520, 1694, 1714, 2182, 2249, 2356, 2709, 2737, 2755, 2962]. fast-flux [853]. Fastlane [2678]. Fastlane-ing [2678]. FatTree [1909]. Fault [428, 867, 1524, 1568, 1570, 2030, 2083, 2134, 2316, 2874, 2951, 2954, 3008]. Fault-tolerant [867, 1524, 2030, 2083, 2316]. faults [2607]. FavorQueue [1147]. fBm [546]. FDMA [1785]. feasibility [1565, 2277]. Feature [34, 956, 982– 984, 1075, 2153, 2158, 2164, 2406, 2777]. feature-based [983]. Feature-interaction [983]. Features [402, 973, 987, 2049, 2544, 2627, 2752, 2856, 2891, 2958]. FEC [223, 580, 2296]. federated [984, 1152, 1154, 1186, 1188, 1192, 2634]. Federation [209, 1163, 1194, 1650]. federations [2184]. FEDERICA [1162]. feedback [665, 1361, 1616, 1846, 1884, 2066, 2219, 2561, 2859]. feedback-based [2561]. femto [643, 1462, 1989, 2578]. femto/macro [1989]. Femtocell [624, 830, 891, 903, 1017, 1023, 1024, 1068, 1340, 1416, 1533, 1713, 1758, 1759, 1860, 1915, 2210, 2346, 2518, 2888, 2904]. femtocells [1496, 1855, 2219, 2315]. Ferry [636]. few [2195]. Fi [454, 738, 1465, 1630, 1643, 2145, 2303, 2441, 2503, 2584, 2957]. Fi/cellular [2442]. FIB [1012, 1536]. Fiber [196, 1149, 2592, 2912]. fiber-wireless [196, 1149]. fictitious [1332]. FIDC [2177]. field [258, 471, 476, 850, 1971, 2596]. FIFO [2667]. Fight [412, 889]. Fighting [1236, 1468]. File [638, 765, 2173, 2650, 2682, 2728]. file-sharing [638]. files [1019]. Filter [245, 382, 928, 1098, 2238, 2559, 2957]. Filtered [144]. filtering [370, 485, 750, 778, 835, 1978, 2479]. filters [13, 349, 595, 1235, 2562]. Finding [126, 1648, 2473]. Fine [51, 235, 852, 1567, 1660, 2378, 2597, 2776, 2807]. Fine-grained [51, 852, 1567, 1660, 2378, 2597, 2776, 2807]. fingerprint [1643, 1847, 2238, 2322]. fingerprint-based [1847]. fingerprinting [2716]. fingerprints [1843]. finish [2306]. finite [370, 632]. firefly [1794, 2943]. firewall [1740]. first [1667]. first-last [1667]. fit [1667, 2796]. fitness [2777]. FITS [1195]. Fitting [247]. FiWi [196]. fixed [626, 2099, 2993]. Fixed-Elastic [2993]. fixed-line [2099]. fixed-point [626]. flagellated [307]. flash [1948, 2137]. Flat [77, 364, 1604]. flat-name [1604]. flaws [848]. Flex [2583]. Flex-Grid [2583]. Flex-Grid/SDM [2583]. flexgrid [736]. flexibility [1887]. Flexible [123, 177, 241, 272, 317, 705, 769, 1050, 1067, 1195, 1269, 1337, 1481, 1756, 2138, 2188, 2278, 2358, 2372, 2467, 2665, 2689, 2722, 2776, 2868]. flexible-bandwidth [2467]. flexible-choice [272]. flexible-grid [705]. Flickr [611]. flip [921]. floating [1185]. flocking [900]. flooding [125, 173, 458, 547, 1708, 2436, 2539, 2671]. floodless [1067]. Flow [40, 62, 99, 133, 167, 177, 333, 380, 411, 424, 517, 555, 560, 597, 629, 650, 778, 829, 831, 870, 902, 941, 1006, 1110, 1143, 1239, 1433, 1442, 1447, 1512, 1653, 1660, 1668, 1678, 1680, 1694, 1695, 1770, 1771, 1851, 1909, 1911, 1952, 2171, 2176, 2179, 2197, 2200, 2243, 2247, 2299, 2339, 2342, 2401, 2473, 2477, 2489, 2526, 2631, 2644]. Flow-based [902, 1110]. flow-holding [2243]. Flow-level [99, 167, 380, 1694, 2489].

Flow-oriented [424]. flow-table [2247]. FLOWER [2016]. flows [348, 703, 906, 923, 1059, 1097, 1109, 1239, 1669, 1948, 2152, 2306, 2385, 2475, 2678]. fluid [10]. flux [853]. flying [2925]. FMIPv6 [871]. focused [2432]. Fog [2383, 2609, 2636, 2858, 2930, 2951]. fog-supported [2951]. folded [1176]. folded-bus [1176]. follow [1410]. followee [1410]. follower [797]. football [2779]. footprint [1643]. footprints [1939]. forbidden [2757]. forecasting [216]. Formal [841, 2433, 2717]. format [850]. formation [64, 631, 1115, 1149, 1340, 1395, 2055, 2274]. formulation [558, 2463]. formulations [57, 139, 2575, 3003]. forthcoming [2706]. forward [124, 1682, 1903]. forwarders [606]. Forwarding [418, 436, 437, 579, 659, 888, 942, 949, 994, 1067, 1200, 1201, 1458, 1579, 1697, 1716, 1798, 1919, 2023, 2025, 2044, 2069, 2110, 2195, 2392, 2536, 2649, 2813]. foster [1194]. Fountain [447]. Fourth [142]. FP7 [1160]. FPAN [2597]. FPGA [510]. Fractional [283, 1444, 1593, 2848]. Fragmentation [1073]. Fragmented [2991]. Frame [642, 738, 869, 930, 1067, 1219, 1672, 1903, 2447, 2672]. Frame-based [869, 1903]. frame-oriented [1219]. framed [2831]. framelet [434]. framelet-based [434]. frames [755]. Framework [25, 65, 84, 106, 317, 345, 389, 462, 658, 665, 706, 743, 796, 802, 843, 883, 920, 999, 1008, 1049, 1074, 1085, 1132, 1168, 1187, 1190, 1217, 1254, 1311, 1326, 1407, 1428, 1466, 1485, 1491, 1514, 1526, 1532, 1551, 1571, 1584, 1594, 1624, 1664, 1676, 1690, 1692, 1700, 1718, 1768, 1801, 1804, 1806, 1849, 1937, 1960, 1968, 2015, 2069, 2127, 2128, 2177, 2184, 2206, 2329, 2346, 2418, 2449, 2453, 2515, 2518, 2543, 2581, 2600, 2602, 2646, 2647, 2655, 2688, 2719, 2779, 2801, 2810, 2889, 2966]. frameworks [1570, 1934]. frauds [2142]. FREDI [2675]. Free [100, 126, 461, 551, 789, 938, 1101, 1129, 1205, 1256, 1296, 1644, 1777, 1831, 1921, 1970, 2154, 2408]. Free-scaling [938]. free-space [461]. freeriders [1218]. Freezing [1458]. frequency [192, 502, 763, 1069, 1481, 1531, 1593, 2323, 2619, 2970]. frequency-based [1069]. freshness [59]. friend [1221]. friendliness [587, 666]. friendly [396, 562, 587]. friends [279]. friendship [1181, 1321]. fronthaul [2993]. FSM [85]. FSO [2912]. FSO/fiber [2912]. FSTR [822]. FSTR-based [822]. FTRS [2197]. full [155, 948, 1270, 1851, 2344, 2589, 2751, 2904]. full-duplex [948, 2344, 2904]. full-text [155]. Fully [2102]. Function [162, 670, 1533, 1662, 1702, 1703, 1709, 1741, 1742, 1882, 2251, 2271, 2303, 2310, 2427, 2430, 2514, 2788, 2903, 2936, 2952, 2962]. Functional [323, 3012]. functionality [642, 1458]. Functions [1565, 1584, 2111, 2246, 2415, 2462, 2507, 2576, 2735, 2776, 2822, 2913]. Fundamental [1046, 2822]. fusion [216, 1044, 1472, 2147, 2694, 2871]. Future [84, 196, 282, 285, 315– 317, 323, 326, 495, 621, 736, 746, 920, 969, 974, 1054, 1139, 1151, 1158, 1182, 1187, 1190, 1227, 1286, 1300, 1402, 1650, 1706, 2067, 2254, 2283, 2337, 2359, 2416, 2461, 2930, 2973]. Fuzzy [384, 485, 579, 926, 1108, 1319, 1421, 1669, 2016, 2150, 2441, 2510, 2605, 2781, 2888]. Fuzzy-based [384, 2441]. FWNs [282].

G [1000, 1158, 1722, 2309, 2748, 2880, 2926, 2936, 2956, 3009]. G-enabled [2748]. G-Lab [1158]. G.9959 [2827]. gain [1130]. gains [23, 1029]. Game [73, 189, 216, 248, 259, 412, 464, 526, 940, 1013, 1048, 1149, 1209, 1304, 1317, 1444, 1501, 1576, 1590, 1631, 1759, 1776, 1855, 1865, 2021, 2055, 2193, 2204, 2210, 2215, 2360, 2367, 2373, 2418, 2509, 2518, 2529, 2634, 2732, 2749, 2857, 2879, 2904, 2961]. Game-based [1759, 2961]. game-theoretic [464, 940]. game-theoretical [1590, 2367]. games [888, 1115, 1261, 1340, 1391]. Gap [740]. Garch [226]. Garch-based [226]. Gateway [12, 20, 26, 94, 293, 2382, 2536]. gateways [712, 2226, 2666]. gathering [528, 893, 897, 1282, 2460, 2505, 2561, 2582, 2643, 2824]. Gaussian [2638]. Gbit [2798]. Gbit/s [2798]. Gbps [38, 756, 2101]. GÉANT [533]. gender [1408]. general [65, 503, 636, 802, 1058, 2329, 2837]. generalizable [1006]. Generalization [1827]. Generalized [40, 382, 499, 1058, 1692, 1994, 2801]. Generalizing [829, 1494]. Generated [104, 977, 2544]. generating [1676, 2866]. Generation [111, 142, 169, 237, 255, 283, 511, 706, 755, 786, 849, 918, 981, 1064, 1108, 1207, 1223, 1455, 1632, 1908, 2156, 2190, 2270, 2324, 2406, 2619]. Generation-Based [918]. gEnerator [1439]. Generic [1387]. Genetic [247, 853, 1985]. Genetic-based [853]. GENI [1152, 1154, 1183, 1184]. Genomics [2633]. geo [608, 2624]. geo-location [608]. geo-temporal [2624]. Geographic [65, 312, 330, 418, 949, 1442, 1521, 2065, 2330, 2428, 2555, 2603]. geographically [1521, 1749]. Geolocation [114, 543, 2132, 2457]. geometric [1528, 2162]. geometry [1805]. GeoPath [1522]. GERAN [450]. GerbilSphere [607]. GEYSERS [1164]. GHz [122, 1113, 2019]. Gibbs [1214]. gigabit [1456]. GitHub [2113]. GLARM [1813]. Global [19, 78, 266, 284, 444, 547, 813, 1159, 1194, 2171, 2733]. globally [2401]. globe [2063]. GMPLS [756]. GMPLS-enabled [756]. Gnutella [111]. go [1303]. Going [429]. Good [1540, 1609, 1940]. goodput [2223, 2385]. GoS [1433]. gossip [1078, 1218]. gossip-based [1218]. GpENI [1155]. GPS [977, 2154]. GPS-free [2154]. GPU [1039]. GPU-accelerated [1039]. grade [11, 376, 1433, 1689, 1936]. gradient [2392]. Graduating [1243]. grained [51, 135, 852, 1567, 1660, 2378, 2597, 2776, 2807].

GrainFlow [1286]. GrAnt [606]. granularity [300, 309, 704, 1632, 1720]. Graph [125, 208, 389, 428, 593, 833, 890, 978, 1063, 1508, 1821, 1924, 2249, 2336, 2340, 2826]. graph-based [389]. graphlet [591]. graphs [410, 497, 1543, 1918, 2407, 2700, 2978]. Graption [389]. GRASP [1572]. Greedy [123, 249, 606, 785, 1528, 1844, 1911, 2541, 2665]. Green [300, 699, 764, 830, 920, 926, 1122, 1391, 1425, 1450, 1451, 1454, 1460, 1464, 1486, 1533, 1683, 1720, 1781, 1842, 2018, 2080, 2251, 2267, 2278, 2315, 2335, 2370, 2371, 2582, 2763]. Greener [955, 1458]. Greening [1018, 1349, 1452]. GreenMap [2080]. grey [1944, 2964]. Grid [252, 256, 409, 507, 687, 705, 729, 875, 901, 909, 1129, 1146, 1222, 1253, 1309, 1312, 1432, 1482, 1571, 1624, 1836, 1865, 1893, 2032, 2142, 2143, 2412, 2674, 2679, 2810, 2876, 2935]. grid-based [1432, 1836, 2935]. Grid/SDM [2583]. GRiDA [764]. grids [243, 1183, 1248, 2328, 2387]. grooming [309, 704, 1288, 1572]. ground [336]. Group [31, 306, 483, 539, 633, 744, 1061, 1085, 1170, 1605, 1712, 1813, 1882, 2138, 2289, 2664, 2726, 2728]. Group-based [1813]. group-key [306]. grouping [555, 1510, 1984]. groups [127, 1438, 2289]. GROUPS-NET [2289]. growing [1724]. growth [1828, 2820]. GSA [123]. GSM [170]. guarantee [1367, 1396]. guaranteed [1018, 1283, 1600, 1956, 2993]. guaranteeing [2079]. guarantees [60, 928, 1355, 2284, 2290, 2679]. guard [1262, 1955]. Guest [148, 509, 1151, 1182, 1946]. guidance [661]. guided [3012]. guidelines [1146]. Guifi.net [1622, 1723, 1730]. Guiltiness [2913].

H.264 [505]. H.264/SVC [505]. habits [2581]. HAIR [324]. HAN [1253]. handheld [2104]. handling [2261]. handoff [58, 376, 662, 1106, 1287, 1344, 1387, 1474, 2038, 2782, 2803]. handoff-aware [662]. handoffs [442]. Handover [142, 239, 262, 360, 361, 520, 682, 1080, 1170, 1262, 1416, 1421, 1544, 1603, 1773, 1824, 1895, 1920, 1957, 2084, 2319, 2321, 2356, 2578]. handovers [362, 560, 1065]. HAPs [461]. hard [576, 2284, 2679]. Hardware [324, 751, 1242, 1859, 2101, 2477, 2613, 2770]. hardware-amenable [324]. HARQ [223, 657, 2219]. harvest [2494]. harvesting [217, 1004, 1599, 1894, 2471, 2495, 2565, 2571, 2587, 2606, 2620, 2851, 2977]. Hash [632, 1112, 1869, 2978]. Hash-routing [1869]. hashing [241]. HCCA [803]. head [2870]. heading [2417]. HealCam [2520]. healing [544, 1144, 1866, 2309]. health [1849, 2372, 2408, 2780, 2881, 2940]. healthcare [199, 1359, 1592, 2365, 2366, 2646, 2731, 2738, 2817]. heat [2298]. heat-diffusion [2298]. heavy [646, 1641, 2475]. HEER [1892]. held [1622]. helicoidal [1608]. Helix [1608]. Helm [2352]. help [246, 279]. HeNB [682]. herd [379]. heterogeneity [1198, 2560]. Heterogeneous [58, 98, 139, 142, 167, 172, 190, 332, 361, 363, 484, 545, 633, 672, 804, 920, 1048, 1186, 1191, 1207, 1283, 1315, 1358, 1382, 1385, 1433, 1461, 1464, 1484, 1495, 1498, 1505, 1627, 1658, 1684, 1807, 1844, 1948, 1963, 2068, 2219, 2291, 2310, 2318, 2370, 2371, 2378, 2385, 2395, 2407, 2410, 2442, 2476, 2514, 2641, 2658, 2704, 2711, 2727, 2741, 2743, 2809, 2842, 2853, 2907, 2914, 2990]. HetNet [1826, 2536]. HetNets [1481, 1764, 2315, 2368, 2744, 2880]. Heuristic [126, 837, 1876, 2702]. heuristics [444, 939, 2901]. hidden [198, 216, 580, 599, 767, 856, 1094, 1898, 1984, 2420, 2425, 2426]. hide [2173]. hiding [2807]. Hierarchical [6, 112, 131, 171, 178, 377, 380, 592, 747, 886, 1234, 1288, 1429, 1489, 1759, 1796, 1874, 2164, 2210, 2256, 2258, 2468, 2540, 2748, 2904]. Hierarchically [779]. hierarchies [681, 964]. hierarchy [865, 2021]. HiFIND [99]. High [10, 38, 82, 99, 237, 241, 348, 385, 746, 751, 872, 884, 1072, 1088, 1217, 1270, 1455, 1483, 1507, 1592, 1671, 1874, 1989, 2003, 2061, 2083, 2256, 2264, 2443, 2608, 2619, 2635, 2643, 2644, 2673, 2735, 2790, 2804, 2902]. high-availability [1270]. high-available [2643]. high-capacity [872]. high-density [2790]. high-dimensional [1088, 1592]. High-efficient [2804]. High-performance [746, 884, 2061, 2083, 2635]. high-precision [2608]. High-speed [10, 38, 99, 237, 751, 2256, 2644, 2735]. high-throughput [872, 2902]. high-traffic [2673]. higher [404, 2445]. higher-order [2445]. highly [1027, 2860]. highway [941, 1398, 2596]. highways [2787]. HIoTSP [2646]. HIP [364, 1625]. HIP-based [1625]. HIP/PMIP [364]. historical [462]. history [233, 586, 815]. hit [678]. HLLS [586]. HMAC [695]. HMAC-based [695]. HMM [661]. Hoc [8, 46, 64, 78, 92, 112, 113, 130, 143, 146, 271, 306, 388, 416, 423, 425, 464, 467, 477, 484, 496, 572, 588, 598, 601, 628, 653, 720, 784, 789, 805, 838, 942, 990, 1001, 1071, 1096, 1177, 1284, 1343, 1386, 1398, 1399, 1432, 1495, 1500, 1501, 1526, 1550, 1595, 1639, 1652, 1831, 1905, 1951, 1964, 1980, 1997, 2028, 2077, 2102, 2150, 2154, 2185, 2193, 2204, 2220, 2230, 2270, 2379, 2390, 2397, 2485, 2757, 2794, 2892, 2983]. holding [2243]. hole [1343, 1866, 2026, 2108, 2343, 2751]. hole-bypassing [2343]. holes [250, 2113]. Holistic [88, 261, 1935, 2708]. home [985, 986, 1172, 1779, 2456, 2505, 2865]. homed [333, 1156, 1684, 2971]. HomePlug [1172, 1538]. homes [2135, 2699]. homing [1967]. homogeneous [672]. homomorphic [2329]. Hop [15, 17, 60, 61, 254, 294, 378, 429, 437, 561, 581, 637, 696, 766, 792, 798, 918, 942, 1096, 1142, 1259, 1293, 1325, 1343, 1424, 1495, 1557, 1906, 1907, 1922, 1951, 1993, 2224, 2225, 2289, 2519, 2674, 2711, 2814, 2935]. hop-by-hop [1907]. hopping [1167, 2323, 2970]. hormone [1494]. Hose [884, 1009, 1285]. hose-based [884]. Host [386, 854, 970, 999, 1441, 2593, 2693]. Host-based [2593]. host-network [999]. hosting [1, 67]. hosts [333, 2132, 2457]. hot [383]. hot-spot [383]. hotspot [742]. hotspots [1216, 1562]. hour [299, 402]. HPC5 [111]. HPSIPT [2608]. HSR [2039]. HSTCP [449]. HTTP [166, 852, 1516, 1626, 2011, 2101, 2183, 2226, 2421, 2655]. HTTP-based [852, 2183, 2421]. hub [1446]. HUBsFLOW [2905]. Human [605, 858, 1216, 1246, 1612, 1750, 1843, 2052, 2439, 2520, 2605]. human-machine [1612]. Hunting [856]. HURBA [6]. HURP [6]. HURP/HURBA [6]. Hybrid [19, 45, 171, 371, 403, 404, 410, 418, 503, 512, 554, 575, 726, 750, 1264, 1348, 1457, 1496, 1499, 1612, 1641, 1653, 1666, 1715, 1965, 1989, 2257, 2275, 2293, 2368, 2393, 2410, 2416, 2478, 2513, 2517, 2546, 2701, 2812, 2825, 2826, 2836, 2862, 2864, 2868, 2912, 2943, 2972, 2976, 2993]. hygiene [1179]. HyPeer [272]. hyper [2777]. hyper-parameter [2777]. hyperbolic [911]. hypertextual [809]. hysteresis [338].

I-CSMA [2205]. IA [2149]. IA-RMLSA [2149]. IaaS [1269]. iBGP [186]. ICARUS [2024]. ice [1084, 2273]. ICMP [2791]. ICN [1038, 2221, 2256, 2391, 2449, 2747, 2984]. ICOD [1212]. ICP [1845]. ID [282, 320, 744, 881, 891, 1136]. ID-based [744]. ID/Locator [282, 1136]. Identification [249, 481, 733, 1646, 1677, 2331, 2746, 2985]. identified [347]. identifier [321, 974, 2409]. identifier-to-locator [321]. identifiers [321]. Identifying [23, 297, 539, 1516, 2472]. Identity [5, 87, 214, 717, 969, 970, 974, 1228, 1605, 1781, 2184, 2199, 2321]. Identity-based [974, 1605, 2321]. idle [519]. IDMS [1305]. IEEE [51, 128, 137, 164, 184, 198, 230, 262, 296, 364, 372, 376, 430, 435, 456, 513–519, 521– 523, 529, 566, 582, 634, 718, 734, 740, 793, 803, 899, 930, 992, 1056, 1103, 1259, 1262, 1377, 1480, 1489, 1573, 1637, 1780, 1829, 1920, 1953, 1984, 2523, 2615, 2696, 2852]. IEEE802.22 [1320]. IG [2705]. IG-PCA [2705]. II [1182, 1734]. III [518]. IIoT [1849]. ILP [57, 2463]. image [1022]. images [2789]. imitation [945]. Impact [91, 121, 122, 187, 331, 397, 487, 488, 512, 688, 700, 1140, 1417, 1480, 1506, 1595, 1621, 1754, 1957, 2008, 2115]. impacting [2992]. impacts [986, 1893, 2776]. Impairment [1665]. Impairment-aware [1665]. impairments [2625]. imperfect [188, 2673]. Implementation [87, 221, 270, 271, 510, 667, 674, 742, 751, 779, 836, 1114, 1155, 1160, 1190, 1217, 1859, 2130, 2337, 2502, 2631, 2680, 2745, 2828]. Implementing [320, 1159]. implications [881, 1140, 1940, 2104]. implicit [2544]. importance [2484]. Imposter [1979]. improve [108, 620, 1147, 1321, 1812, 2209, 2342]. Improved [51, 636, 752, 941, 1200, 1696, 1980, 2135, 2201, 2574, 2596, 2884, 2922, 2943]. Improvement [520, 670, 1078, 1093, 1172, 1205, 1377, 2024, 2429, 2899]. improvements [1514]. Improving [36, 114, 245, 246, 459, 513, 560, 762, 795, 1056, 1105, 1173, 1179, 1271, 1519, 1522, 1576, 1614, 1643, 1731, 1844, 2177, 2205, 2313, 2380, 2448, 2501, 2644, 2649, 2729, 2782, 2793, 2890]. IMS [295, 358, 443, 1080]. IMS-based [358, 1080]. in-advance [2521]. in-band [237]. in-building [1057]. in-depth [950]. in-home [1172, 1779]. in-network [989, 1037, 1397, 1451, 1503, 1659, 2035, 2389, 2875]. In-packet [349]. in-vehicle [2530]. inband [1826]. InCan [1397]. incast [1267, 2080, 2234]. Incentive [525, 1051, 1348, 1613, 1670, 1861, 1966, 1974, 2109, 2360, 2395, 2400, 2403, 2454, 2549, 2636]. incentive-compatible [1966]. incentives [101]. Incentivizing [284, 931, 2461]. incident [2462]. incoming [1339]. incomplete [1884, 2440]. inconsistencies [121]. inconsistency [34, 2041].

Incorporating [2221, 2749]. increase [213, 946, 1905]. increases [979]. increasing [2305]. incremental [1133, 1239, 2249, 2642]. Incrementally [911]. Independent [360, 946, 1170, 1311, 1419, 2217]. index [1059, 2241]. Index-Trie [2241]. indexed [2978]. indexing [155, 156, 731]. indicator [1360, 2730]. indicators [1525]. indirect [3013]. individual [1396]. Indoor [381, 1474, 1643, 1658, 1847, 1915, 2170, 2559, 2916, 2957]. induced [2115, 2566, 2640]. Induction [1508]. Industrial [415, 1834, 1836, 1840, 1843, 1849, 2319, 2711, 2772, 2803, 2827, 2852, 2853, 2866, 3005]. industry [1846]. inertial [1843]. infections [1769]. inference [850, 2333, 2840]. inferences [1179]. Inferring [606, 964, 1769]. infested [1910]. inflation [187]. influence [589, 667, 1324, 1372, 2217, 2883, 3013]. Information [125, 158, 179, 216, 262, 294, 319, 360, 462, 547, 586, 611, 614, 799, 810, 1034– 1038, 1041, 1042, 1045, 1077, 1121, 1247, 1315, 1342, 1375, 1429, 1494, 1513, 1530, 1559, 1612, 1644, 1663, 1696, 1869, 1912, 1950, 1951, 1999, 2015, 2023, 2025, 2058, 2093, 2124, 2126, 2137, 2147, 2224, 2233, 2242, 2254, 2255, 2268, 2276, 2304, 2348, 2368, 2390, 2501, 2505, 2539, 2541, 2547, 2557, 2569, 2723, 2766, 2934, 2986, 3012]. Information-Centric [319, 1037, 1041, 1045, 1077, 1121, 1247, 1513, 1530, 1559, 1644, 1696, 1869, 1912, 1950, 1951, 2015, 2023, 2025, 2126, 2254, 2255, 2276, 2541, 2547, 2557, 2723, 2986, 3012]. Infrastructure [15, 89, 152, 240, 269, 365, 369, 433, 442, 476, 478, 563, 1041, 1145, 1155, 1162–1164, 1175, 1622, 2068, 2090, 2311, 2589, 2610, 2640]. infrastructure-based [433, 2311]. infrastructure-less [2090, 2610]. infrastructured [1754]. infrastructures [332, 493, 623, 628, 1295, 1722, 1929, 2251, 2679]. ing [2678]. ingress [1169]. inhomogeneity [2123]. inhomogeneous [2588]. initiated [1418]. Initiation [648]. initiative [1153]. injected [678]. injection [2361]. Inner [607]. Innocence [1243]. innovation [1159]. innovative [1152, 1956, 2016, 2873]. Inoculation [1181]. insecure [2872]. inspection [347]. inspired [70, 71, 136, 495, 1805, 2141, 2276, 2605]. InstaGENI [1153]. Instant [567, 1428]. Instantaneous [1523, 1845]. instantiation [1702]. instead [2890]. instruction [2490]. instrumentation [237, 1187]. Integer [139, 162, 545, 558, 803, 3003]. Integer-multiple-spacing-based [803]. Integrated [29, 149, 362, 364, 452, 665, 756, 1159, 1266, 1291, 1438, 1463, 1543, 2291, 2536]. Integrating [376, 2743]. Integration [271, 461, 971, 1080, 1138, 1164, 1971, 2684, 3011]. Integrity [377, 1995, 2365]. intellig [2344]. intellig-ent [2344]. intelligence [658, 2048, 2387, 3009]. Intelligent [64, 130, 1188, 1212, 1640, 2278, 2393, 2486, 2530, 2566, 2616, 2622, 2648, 2663, 2688, 2734, 2748, 2787, 2995, 2996]. intensive [1655, 2995]. Inter [127, 246, 322, 386, 582, 739, 805, 886, 1064, 1087, 1109, 1128, 1179, 1470, 1488, 1527, 1668, 1798, 2020, 2176, 2283, 2631, 2683, 2921, 2953, 2987, 2991]. inter-AS [1179]. inter-cluster [2683]. inter-contact [2991]. inter-datacenter [2921, 2987]. Inter-Destination [739]. Inter-domain [127, 246, 322, 886, 1109, 1128, 1488, 1527, 1798, 2283, 2953]. Inter-flow [1668, 2176, 2631]. inter-host [386]. inter-RAT [1064]. inter-router [1470]. inter-session [582]. inter-user [1087]. inter-vehicular [805]. interacting [2934]. Interaction [964, 982–984, 1188, 2749]. interactions [34, 169, 343, 986, 1576, 2418, 2876]. interactive [650, 656, 927, 1886, 1943, 2443]. interconnected [2434, 2672]. interconnecting [1207, 2063]. interconnection [216, 683, 981, 1197, 2085]. interdependency [1222]. Interdomain [1306, 1645, 1981]. interest [1654]. interest-tree [1654]. interests [2173]. interface [110, 1958, 2084, 2905]. interfaces [30, 2326, 2357]. Interference [58, 171, 276, 460, 720, 819, 1017, 1087, 1359, 1420, 1595, 1660, 1776, 1816, 1918, 2030, 2176, 2423, 2432, 2446, 2480, 2621, 2675, 2683, 2684, 2744, 2774, 2811, 2831, 3001]. Interference-Aware [1359, 2030, 2480]. Interference-based [3001]. interfering [646]. interleaved [642]. interleaving [793]. intermittently [1201]. Internet [12, 59, 205, 280, 285, 311, 313, 315, 317, 318, 324, 326, 336, 339, 389, 428, 531, 549, 654, 658, 691, 700– 702, 733, 746, 790, 820, 833, 862, 956, 969, 973, 974, 1049, 1059, 1151, 1182, 1185, 1187, 1190, 1207, 1254, 1286, 1292, 1300, 1338, 1347, 1402, 1439, 1508, 1554, 1625, 1669, 1706, 1729, 1769, 1839, 1840, 1843, 1844, 1849, 1856, 2067, 2074, 2132, 2164, 2283, 2295, 2339, 2364, 2404, 2445, 2457, 2497, 2567, 2594, 2621, 2622, 2629, 2646, 2680, 2710, 2713, 2714, 2738, 2753, 2780, 2791, 2817, 2823, 2881, 2891, 2908, 2909, 2931, 2973]. Internet [44, 89, 203, 236, 286, 316, 442, 751, 941, 977, 978, 1109, 1158, 1215, 1277, 1339, 1423, 1488, 1834, 1835, 1910, 2038, 2257, 2361, 2363, 2428, 2492, 2524, 2666, 2775, 2793, 2805, 2850, 2934, 2950]. Internet-based [442]. Internet-of-Things [2934]. internet-of-vehicles [2497]. Internet-scale [1769, 2257]. Internet-wide [311]. Internetworking [87]. interoperability [1263]. interoperable [1263]. Interplanetary [2336]. Interplay [1118, 2973]. intersection [475]. intersession [1537]. Intertwined [1395]. Interval [85, 265, 650, 748, 2731]. Interval-based [650]. interworking [170, 657, 2409]. intra [827, 930, 2176, 2631, 2683]. intra- [2683]. intra-domain [827]. intra-flow [2176, 2631]. intra-frame [930]. intradomain [239]. intrinsic [402]. Intrinsically [385, 609]. Introducing [326, 2760]. intruders [1178]. Intrusion [99, 370, 371, 496, 874, 966, 1069, 1491, 2006, 2297, 2478, 2617, 2705, 2775, 2823, 2887, 2909]. intrusive [2615]. invalidation [2435]. investigate [1324]. Investigation [292, 458, 1627, 1954, 2909]. inward [2574]. IoMT [2870]. IoT [1165, 1823, 1842, 1968, 2021, 2065, 2163, 2318, 2353, 2366, 2373, 2382, 2387, 2425, 2444, 2609, 2621, 2629, 2660, 2695, 2703, 2706, 2712, 2713, 2715, 2716, 2718, 2736, 2743, 2749, 2766, 2772, 2778, 2781, 2783, 2789, 2795–2797, 2807, 2815, 2818, 2827, 2836, 2884, 2940, 2949, 2966, 3010]. IoT-based [2778, 2836, 2940, 2949]. IOTA [2720]. IoV [2501]. IP [24, 84, 100, 107, 135, 237, 309, 346, 347, 402, 417, 543, 553, 594, 640, 704, 726, 764, 776, 800, 828, 835, 919, 954, 1122, 1133, 1135, 1145, 1206, 1291, 1318, 1324, 1330, 1342, 1390, 1425, 1520, 1608, 1777, 1791, 1809, 1820, 1858, 1919, 1939, 1943, 2073, 2161, 2167, 2538, 2608, 2927, 2944, 2980]. IP-based [24, 84, 828]. IP-ERN [726]. IP-multicast [640]. IPD [340]. IPsec [2903]. IPTV [119, 346, 417, 562, 1150, 1283]. IPv4 [977]. IPv6 [749, 872, 996, 1330, 1854, 2356, 2827, 2832, 2850, 3015]. IR [2715]. island [2138]. island-based [2138]. isolation [135]. isotonic [2637]. ISP [610, 1302, 2095]. ISP-CDN [2095]. ISPs [2027, 2730]. Issue [15, 42, 69, 148, 315, 470, 612, 699, 746, 845, 961, 969, 982, 1034, 1151, 1182, 1363, 1460, 1611, 1946, 1991, 2112, 2353, 2712, 2996]. Issues [38, 631, 2076, 2087, 2181, 2352, 2629, 2770, 2851, 2963, 3005]. iterated [260]. iteration [420]. iterative [871, 1196]. ITU [3014]. ITU-T [3014].

Jacobi [420]. jammer

[1081, 1573, 1997, 2550]. jammers [250]. jamming [412, 1512, 1815, 1913, 1998, 2133, 2323, 2344, 2496, 2565, 3010]. jamming-aware [2344]. Japan [473]. JITeR [1890]. job [163, 1524, 2115, 2877]. Join [2502]. Joint [20, 46, 94, 359, 375, 502, 519, 545, 767, 773, 776, 861, 976, 1064, 1081, 1104, 1295, 1320, 1380, 1421, 1486, 1493, 1550, 1683, 1748, 1803, 1822, 1826, 1836, 1930, 1952, 2025, 2029, 2037, 2252, 2264, 2381, 2389, 2422, 2633, 2658, 2665, 2683, 2882, 2902, 2951]. Joint-family [1930]. Jointly [1789, 2165, 2744, 2952]. Journal [2057]. JumpFlow [1695]. Just [1890, 2855]. Just-in-time [1890, 2855].

K-GENI [1154]. Kademlia [55, 944, 1761, 1811]. Kademlia-based [55]. Kalman [2238, 2559, 2562, 2957]. kernel [792, 1840, 2294]. Kernelized [2756]. key [117, 193, 197, 306, 385, 431, 483, 681, 744, 1061, 1102, 1108, 1215, 1233, 1470, 1581, 1641, 1757, 1838, 1882, 1901, 1916, 2000, 2058, 2102, 2321, 2338, 2362, 2398, 2412, 2546, 2572, 2707, 2724, 2958, 2978]. key-node [2978]. keys [977]. kinematics [2182]. kinematics-based [2182]. Kinetics [2052]. KioskNet [270]. Kleinrock [1237]. knowledge [1297, 1621, 2101, 2154, 2472, 2685, 2934]. known [410]. KNXnet [1342]. KNXnet/IP [1342]. KPS [2978]. KREONET [1154].

L [1899]. L-MAC [1899]. Lab [1158]. Label [178, 1944]. labeled [18]. labeling [2154]. labor [1619]. laboratory [1618]. Laman [2889]. Lambda [1159]. LANC [342]. landmark [386, 2154]. landmark-labeling [2154]. language [1207, 1613]. languages [569]. LANs [3, 49, 192, 1573, 1632, 1754, 2269]. Laplacian [427]. Large [135, 202, 399, 437, 449, 502, 549, 583, 591, 613, 623, 656, 683, 809, 883, 1125, 1197, 1307, 1380, 1388, 1471, 1504, 1521, 1581, 1642, 1781, 1793, 1840, 1935, 1984, 2017, 2026, 2060, 2116, 2170, 2234, 2290, 2315, 2650, 2741, 2761, 2796, 2805, 2981]. Large-scale [135, 502, 549, 583, 613, 623, 656, 809, 883, 1125, 1197, 1307, 1380, 1388, 1504, 1521, 1781, 1793, 1840, 1935, 1984, 2017, 2026, 2060, 2116, 2234, 2290, 2315, 2650, 2741, 2761, 2796, 2805, 2981]. largest [2573]. last [1667]. Last.FM [1408]. Latency [386, 452, 1082, 1430, 1887, 1897, 1923, 1942, 2018, 2025, 2098, 2510, 2936]. latency-aware [2018]. latest [473]. law [286, 1975]. laws [1220]. Layer [17, 29, 86, 88, 287, 305, 378, 417, 511, 559, 565, 572, 588, 645, 665, 707, 795, 838, 862, 913, 948, 970, 977, 1056, 1090, 1096, 1189, 1319, 1359, 1425, 1509, 1577, 1585, 1652, 1692, 1755, 1804, 1852, 1853, 1890, 1968, 1991, 1992, 1995, 1999, 2126, 2183, 2220, 2291, 2322, 2451, 2821, 2944]. layer-2 [1189, 1968]. Layered [39, 625, 748, 793, 1400, 1699, 2467]. layered-video [793]. layers [2936]. LBAS [1199]. Lcast [1128]. LCMSC [2186]. LDMAC [2611]. LDPC [1830]. LE-MRC [2755]. leader [797]. leads [1377]. leakage [1014]. leaks [1527]. learned [1824]. Learning [64, 82, 181, 227, 359, 1144, 1252, 1319, 1332, 1421, 1601, 1631, 1634, 1640, 1699, 1801, 1840, 1852, 1865, 1884, 1899, 1934, 1938, 2100, 2162, 2231, 2269, 2271, 2405, 2406, 2606, 2650, 2731, 2775, 2779, 2789, 2795, 2864, 2887, 2906, 2920, 2922, 2931, 2932, 2955, 3007]. Learning-based [1699, 2775, 2920]. learning-enforced [2606]. leasing [439, 1369, 2194]. least [427]. LEDBAT [943]. ledger [2720]. legacy [871, 2300]. length [1637, 2201]. lengths [734]. LEO [231]. Leonard [1237]. less [632, 1560, 2090, 2610, 2678]. lessons [1824]. Let [2306]. Level [99, 167, 183, 223, 228, 229, 242, 374, 380, 489, 555, 778, 833, 879, 882, 1326, 1596, 1664, 1694, 1763, 1810, 1858, 2469, 2484, 2489, 2684, 2766]. Leveraging [738, 1321, 1542, 1906, 2318, 2841]. Libra [176]. licensed [2497]. LICITUS [1968]. life [1725, 2482]. lifecycle [847]. Lifetime [213, 258, 403, 452, 592, 653, 893, 1004, 1213, 1224, 1303, 1902, 1904, 2901, 2945]. Light [586, 666, 1440, 1716]. light-traffic [1716]. lightpath [453]. Lightweight [228, 306, 523, 1040, 1087, 1215, 1331, 1474, 1579, 1625, 1813, 1968, 2051, 2065, 2186, 2270, 2444, 2493, 2538, 2548, 2632, 2724, 2750, 2917, 2983]. like [567, 905, 1131, 1659, 1891, 2805]. likelihood [195, 1058, 2377, 2433]. limit [291]. limitation [1913, 2094]. Limitations [1073]. Limited [48, 455, 1021, 1297, 1921, 2134, 2401, 2642, 2685]. Limiting [754, 1342]. limits [1046, 2457]. line [247, 1281, 1312, 1749, 2099, 2967]. linear [162, 545, 594, 681, 1381, 2971, 3003]. Link [27, 90, 120, 224, 229, 305, 329, 334, 460, 534, 776, 792, 823, 827, 976, 1100, 1111, 1122, 1173, 1264, 1438, 1520, 1728, 1739, 1810, 1910, 2034, 2121, 2138, 2174, 2205, 2405, 2528, 2558, 2615, 2644, 2671, 2733, 2750, 2752, 2762, 2783, 2912, 2927]. link-flooding [2671]. link-level [229]. Link-state [27]. linked [2268]. links [52, 425, 461, 646, 901, 939, 946, 1113, 1301, 1586, 1816, 2447]. Linux [568]. Liquid [340]. LISP [1635]. listener [1756]. listening [2696]. literature [2114]. Live [93, 101, 154, 246, 787, 791, 807, 889, 1026, 1126, 1299, 1358, 1498, 1705, 2160, 2386, 2512, 3014]. live-streaming [1126]. lived [1211]. LMS [2296]. Load [136, 231, 255, 281, 384, 405, 579, 619, 684, 686, 697, 870, 902, 916, 1271, 1330, 1350, 1383, 1421, 1764, 1826, 1925, 2118, 2200, 2240, 2300, 2306, 2316, 2343, 2442, 2556, 2773, 2859, 2880, 2943, 2994]. Load-aware [405]. load-balanced [684, 697, 2200]. Load-balancing [619, 686, 2556]. Load-conscious [2240]. LOADng [2270]. Local [415, 776, 837, 1641, 1998, 2101, 2529, 2733, 2901]. localisation [1597]. locality [157, 291, 342]. locality-aware [157, 342]. Localization [145, 182, 312, 342, 379, 381, 427, 504, 789, 921, 1011, 1074, 1081, 1568, 1574, 1843, 1847, 1924, 2036, 2105, 2151, 2170, 2238, 2550, 2564, 2588, 2597, 2826, 2874, 2954, 2957]. Localized [1251, 2761, 2783]. Localizing [1100]. Location [126, 139, 240, 294, 450, 519, 586, 608, 716, 741, 877, 1017, 1256, 1296, 1446, 1510, 1535, 1594, 1713, 1719, 1758, 1778, 1875, 1915, 2078, 2240, 2253, 2292, 2395, 2400, 2454, 2516, 2535, 2614, 2643, 2661, 2694, 2703, 2854, 2960]. Location-aware [1713]. Location-based [716, 741, 1017, 1719, 1915]. Location-dependent [1594, 2395]. Location-free [126, 1256, 1296]. location-sensitive [2400]. locator [282, 320, 321, 881, 974, 1136, 2409]. Locator/ID [320, 881]. Locator/Identifier [2409]. LODMAC [1535]. log [1626]. logging [1606]. logic [1421, 2510, 2605]. Logical [2320]. long [195, 313, 546, 643, 724, 1770, 2611]. long-distance [2611]. long-range [195]. long-term [313]. longevity [573]. longitudinal [2094]. look [1666]. lookup [459, 553, 749, 872, 1039, 1071, 1206, 1608, 1811, 2167, 2980]. Loop [100, 631, 1777, 1835, 2703]. Loop-free [100, 1777]. loosely [156]. LoRaWAN [2717]. LOS/NLOS [2175]. loss [10, 107, 129, 580, 915, 1342, 1674, 2142, 2528, 2571, 2860]. loss-event [10]. losses [2607]. Lossy [425, 1301, 1961, 2272, 2335, 2860]. love [2012]. Low [52, 76, 126, 200, 289, 490, 709, 778, 901, 943, 1005, 1236, 1282, 1293, 1307, 1648, 1669, 1722, 1780, 1961, 2000, 2192, 2202, 2243, 2272, 2335, 2436, 2444, 2481, 2525, 2542, 2603, 2619, 2632, 2635, 2755, 2792, 2936]. low-complexity [2542]. Low-cost [1722, 2603, 2635]. low-duty-cycle [2436]. low-end [1005]. Low-energy [2755]. Low-power [52, 709, 1780, 1961, 2335]. low-rate [200, 778, 2481, 2792]. low-resource [2000, 2444]. lower [2016, 2457]. Lower-than-Best-Effort [2016]. LPM [2538]. LPWANs [2225]. LRU [2259, 2435]. LSTM [2779]. LTE [359, 682, 932, 1061, 1080, 1095, 1127, 1287, 1320, 1333, 1336, 1350, 1384, 1397, 1416, 1417, 1420, 1421, 1484, 1505, 1593, 1664, 1710, 1750, 1757, 1772, 1826, 1931, 1963, 2029, 2059, 2066, 2079, 2119, 2165, 2216, 2274, 2303, 2324, 2346, 2369, 2394, 2452, 2523, 2536, 2543, 2545, 2562, 2578, 2641, 2654, 2658, 2728, 2744, 2747, 2854, 3014]. LTE-A [1350, 1772, 1826, 1963, 2029, 2079, 2119, 2216, 2744]. LTE-Advanced [1095, 1287, 1336, 1420, 1920, 2394, 2452, 2658, 2728]. LTE-based [1127, 1416, 2324]. LTE-U [2523]. LTE-Unlicensed [2346]. LTE/Femto [2578]. LXC [2988]. Lyapunov [1478].

M [535, 1800]. M-WiMAX [535]. M/M/1 [1800]. M2C [868]. M2M [1350, 2293, 2367, 2562, 2831, 2979]. MAC [17, 49, 268, 310, 393, 407, 490, 637, 644, 689, 720, 825, 826, 1007, 1056, 1124, 1167, 1251, 1352, 1395, 1415, 1535, 1538, 1639, 1675, 1779, 1804, 1887, 1899, 1959, 2239, 2262, 2284, 2293, 2516, 2611, 2620, 2803, 2808, 2838, 2852, 2972]. MAC-layer [17]. MAC-routing [1251]. mac80211 [1514]. Machine [831, 1257, 1379, 1612, 1634, 1655, 1737, 1813, 1840, 1934, 2349, 2357, 2405, 2650, 2651, 2731, 2775, 2795, 2906, 2931]. Machine-to-Machine [2349, 2651]. machines [1705, 1740, 2617]. Macro [1017, 1462, 1989, 2059, 2518, 2916]. macro-femtocell [1017, 2518]. macro/femto [1462]. macroscopic [273, 1477]. made [1269]. MAF [2020]. MAF-SAM [2020]. MAFM [1073]. MAGMA [2004]. Maintaining [1854]. maintenance [455, 688, 1166]. major [1253]. make [1458]. Make&activate [2677]. Make&activate-before-break [2677]. Making [48, 552, 633, 1631, 1781, 2456]. malicious [459, 744, 1317, 1579, 1965, 2499]. malware [852, 999, 1123, 1769, 1970, 2004, 2345, 2856, 2906]. malware-analysis [2345]. man [1086, 2981]. man-in-the-middle [1086, 2981]. manageability [762]. managed [534, 957, 1013, 1820]. Management [29, 32, 79, 84, 105, 130, 193, 209, 210, 212, 214, 271, 281, 290, 301, 306, 400, 478, 517, 521, 569, 575, 600, 616, 627, 681, 700, 717, 781, 796, 819, 821, 877, 923, 1008, 1116, 1138, 1147, 1168, 1170, 1228, 1233, 1263, 1295, 1298, 1311, 1332, 1350, 1385, 1416, 1431, 1433, 1457, 1470, 1539, 1544, 1559, 1567, 1591, 1613, 1615, 1647, 1650, 1660, 1662, 1672, 1696, 1705, 1708, 1751, 1754– 1756, 1766, 1772, 1773, 1781, 1901, 1925, 2022, 2102, 2118, 2146, 2191, 2236, 2305, 2341, 2359, 2412, 2459, 2531, 2568, 2656, 2693, 2754, 2782, 2854, 2917, 3008]. management [45, 166, 244, 359, 449, 535, 935, 1117, 1269, 1326, 1420, 1462, 1464, 1882, 1895, 1929, 2015, 2199, 2396, 2616, 2744, 2951]. manager [2418]. Managing [54, 208, 750, 1744, 2080]. MANET [90, 1786, 1870, 2773, 2838]. MANETS [86, 213, 586, 1204, 1233, 1366, 1579, 2286, 2320, 2392, 2585, 2923]. Mantis [1445]. Many [185, 295, 1094, 1720, 1980, 2227, 2278, 2403]. Many-to-many [295, 1720, 1980, 2227, 2278]. many-to-one [185]. map [91, 2292]. mapping [321, 513, 761, 887, 930, 1073, 1446, 1924, 2080, 2377, 2404]. mappings [881, 1635]. MapReduce [1570, 2071, 2080]. MapReduce-based [2080]. maps [926]. market [54, 534, 538, 1304, 1558]. market-based [54, 534]. marketplaces [2521]. markets [365, 1619]. marking [2538, 2765]. Markov [188, 580, 1064, 1489, 1971]. Markovian [573]. MARS [2345]. MAS [821]. MAS-based [821]. mashments [1766]. mashups [1202, 1766]. Massive [1076, 1298, 1435, 2122, 2347, 2684]. massive-scale [1076]. match [2475]. matchability [2737]. matching [492, 842, 1069, 2179, 2529, 2737, 2875]. Mathematical [1057, 2751]. matrices [195, 299, 2275, 2333]. Matrix [121, 235, 491, 676, 2501, 2537]. Maturing [1161]. max [36, 133, 308, 1442, 1653]. max-flow [1442]. MaxCD [941]. Maximization [267, 440, 627, 893, 1345, 1372, 1672, 1881, 1993, 2019, 2188, 2217, 2240, 2263, 2552, 2882, 2888]. maximize [2162, 2809, 2890]. maximized [289]. Maximizing [1000, 1004, 1303, 1391, 1718, 1904, 2119, 2266, 2525, 2655, 2897]. maximum [195, 1058, 1767, 1902, 2340, 2377, 2967]. MBMS [301]. MBS [719]. MC [1082]. MC-MLAS [1082]. MDP [1627]. MDP-based [1627]. Mean [923, 1351]. Mean-variance [923]. means [481, 1059]. measure [1070]. measured [161, 1073]. Measurement [4, 38, 237, 354, 543, 596, 611, 612, 615, 712, 766, 991, 1109, 1111, 1187, 1221, 1292, 1377, 1447, 1533, 2017, 2477, 2802]. Measurement-based [4, 612, 712, 1109, 1533]. measurements [51, 135, 311, 652, 1049, 1053, 1157, 1275, 1503, 1617, 1943]. Measuring [614, 1184, 1735]. MEC [2459]. mechanical [1714]. mechanics [72, 136]. Mechanism [37, 111, 223, 386, 445, 494, 523, 527, 559, 718, 735, 888, 977, 1040, 1062, 1174, 1199, 1301, 1348, 1397, 1422, 1456, 1670, 1788, 1800, 1861, 1978, 2044, 2069, 2081, 2171, 2186, 2197, 2221, 2276, 2360, 2391, 2395, 2400, 2403, 2454, 2490, 2491, 2512, 2531, 2726, 2805]. Mechanisms [193, 525, 529, 538, 560, 597, 638, 711, 873, 2094, 2109, 2295, 2326, 2549]. Media [174, 360, 550, 1023, 1150, 1170, 1193, 1279, 1311, 1406, 1408, 2112, 2115, 2156, 2747, 2871]. media-centric [1193]. media-induced [2115]. media/video [1150]. Medical [2545, 2607, 2891]. medium [307, 689, 1418, 2344, 2452]. medium-range [307]. meet [286, 790, 1775]. meeting [1534]. meetings [2289]. meets [658]. member [1438]. membership [31, 1001, 1170]. memory [492, 1432, 2273]. memory-efficient [492, 2273]. merging [2931]. Mesh [16, 19–22, 24–27, 45, 57, 94, 164, 224, 263, 305, 369, 405, 407, 411, 442, 446, 447, 515, 571, 582, 584, 597, 601, 622, 634, 645, 646, 662, 740, 743, 837, 976, 1051, 1079, 1144, 1200, 1344, 1454, 1531, 1544, 1572, 1601, 1637, 1646, 1668, 1696, 1725, 1726, 1729, 1803, 1806, 2030, 2073, 2176, 2316, 2405, 2496, 2631, 2701, 2986]. mesh-based [446]. mesh-pull-based [447]. Message [452, 494, 636, 850, 858, 1201, 1509, 1648, 2324, 2738]. messages [425, 1093, 1470, 1964, 1995]. messaging [567, 2047]. meta [210, 444, 1154]. meta-design [210]. meta-heuristics [444]. metaheuristic [2028]. MetaRing [1176]. meter [1129]. metering [933]. Method [10, 163, 228, 381, 421, 564, 639, 1006, 1011, 1079, 1221, 1262, 1456, 1490, 1570, 1761, 1793, 1809, 1866, 1883, 1947, 2020, 2417, 2427, 2564, 2624, 2640, 2802, 2866]. MeTHODICAL [1223]. methodologies [2261]. Methodology [59, 415, 475, 1936, 2257, 2267, 2477]. Methods [371, 438, 603, 731, 782, 792, 830, 1054, 1249, 1411, 1436, 1713, 1913, 1954, 1981, 2621, 2762]. metric [1130, 1727, 1831]. Metrics [18, 220, 405, 499, 504, 965, 2137, 2580, 2623, 2637]. metro [102, 445, 822, 1665, 1711]. metro/access [1711]. metropolitan [263, 521]. MFR [636]. MI3M [1311]. mice [1697]. micro [1714, 2383]. micro-electro-mechanical [1714]. mid [1481]. mid-term [1481]. middle [1086, 2981]. middlebox [2033, 2208]. middleware [628, 867, 1494, 1949, 2051]. Midgar [1207]. migration [1199, 1705, 2266, 2848, 2863, 2883, 2988]. migrations [1740]. MIH [262, 2291]. MIMETIC [3007]. Mimic [864]. Mimics [1073]. MIMO [510, 819, 1028, 1298, 1551, 1675, 2122, 2347, 2542, 2684]. MIMO-OFDM [510]. MIMO-OFDMA [819]. min [36, 133, 308, 1442]. min-cut [1442]. mine [2826, 2928]. mines [1891]. minimal [1536]. minimization [276, 1367, 1534, 1683, 2075]. minimize [2877]. minimized [1112]. Minimizing [1084, 1507, 1771, 1957, 2079, 2370, 2371, 2883, 2967]. Minimum [619, 777, 896, 1082, 1180, 1705, 2015, 2059, 2072, 2224, 2290, 2423, 2436, 2463, 2654, 2729, 2935, 3006]. minimum-delay [2436]. Minimum-energy [1705]. Mining [1821, 1934, 1965, 2440, 2524]. mining-based [1965]. Minnie [2195]. MIPv6 [1385]. misbehaving [530]. misinformation [963]. mismatched [2022]. MISO [2494]. missing [491, 2866, 2985]. mission [179, 669, 1708]. mission-critical [179, 1708]. mission-oriented [669]. mist [2685]. Mistrustful [2173]. misuse [2478]. mitigate [1679, 1761, 2774]. mitigating [192, 465, 1086, 1087, 1490, 1792, 2671]. mitigation [1017, 1174, 1776, 1984, 2176, 2481, 2560]. Mixed [63, 998, 1411, 2987]. mixed-type [63]. mixes [1260, 1707]. mixing [17, 2720]. mixture [2455]. ML [1075, 1509, 2499]. ML- [2499]. MLAS [1082]. mMTC [2740]. mmWave [2916]. MNP [2043]. Mob [2846]. MobiCache [1542]. Mobile [92, 130, 131, 216, 239, 257, 285, 330, 361, 372, 384, 388, 423, 461, 510–512, 520, 522, 668, 669, 720, 770, 784, 826, 912, 916, 929, 950, 1071, 1157, 1180, 1201, 1216, 1225, 1228, 1245, 1256, 1262, 1283, 1284, 1296, 1330, 1368, 1373, 1386, 1393, 1479, 1513, 1562, 1567, 1580, 1587, 1595, 1648, 1677, 1760, 1783, 1831, 1879, 1940, 1946, 1951, 1955, 1979, 2044, 2049, 2070, 2077, 2087, 2102, 2107, 2146, 2190, 2191, 2207, 2319, 2321, 2359, 2369, 2383, 2400, 2460, 2489, 2548, 2549, 2559, 2581, 2657, 2682, 2692, 2726, 2768, 2794, 2858, 2899, 2920, 2925, 2977, 3013]. mobile [8, 42, 44, 46, 117, 333, 365, 464, 482, 496, 505, 572, 575, 631, 819, 830, 834, 838, 841, 866, 1011, 1066, 1085, 1363, 1365, 1367, 1383, 1443, 1492, 1504, 1542, 1615, 1616, 1629, 1710, 1717, 1744, 1751, 1775, 1778, 1796, 1812, 1837, 1861, 1923, 1948, 1956– 1958, 1988, 2012, 2028, 2043, 2045, 2051, 2066, 2084, 2125, 2154, 2177, 2185, 2279, 2280, 2283, 2304, 2326, 2327, 2340, 2356, 2385, 2403, 2419, 2424, 2454, 2455, 2485, 2490, 2552, 2643, 2660, 2843, 2850, 2862, 2931, 2973, 3007, 3011]. mobile-edge [2858]. mobile-sink [1256].

Mobility [45, 113, 178, 266, 358, 455, 523, 577, 858, 877, 979, 996, 1138, 1311, 1318, 1385, 1420, 1452, 1469, 1477, 1706, 1756, 1760, 1773, 1961, 1988, 2037, 2169, 2191, 2396, 2464, 2497, 2536, 2610, 2624, 2656, 2693, 2759, 2782, 2846, 2850, 2907, 2963]. mobility-aware [2656, 2907]. mobility-cast [2169]. Mode [421, 519, 521, 582, 634, 711, 793, 1029, 1450, 2411, 2868, 2955]. Mode-oriented [2868]. Model [40, 46, 113, 114, 178, 188, 212, 214, 257, 296, 410, 435, 450, 452, 453, 526, 600, 662, 765, 797, 841, 847, 869, 880, 907, 965, 977, 1140, 1166, 1210, 1329, 1336, 1369, 1427, 1442, 1477, 1489, 1512, 1579, 1690, 1703, 1754, 1816, 1844, 1862, 1952, 1965, 1970, 1972, 2010, 2011, 2021, 2035, 2098, 2100, 2173, 2193, 2209, 2231, 2237, 2246, 2336, 2355, 2372, 2467, 2475, 2489, 2566, 2607, 2617, 2636, 2681, 2738, 2743, 2777, 2778, 2799, 2866, 2924, 2982, 3010, 3014]. model-based [907]. Model-driven [212]. Modeling [51, 242, 343, 420, 433, 442, 457, 493, 543, 611, 615, 723, 839, 978, 991, 1029, 1057, 1073, 1107, 1548, 1586, 1651, 1662, 1886, 1908, 2009, 2034, 2070, 2213, 2261, 2280, 2304, 2435, 2535, 2547, 2668, 2688, 2812]. modelled [2927]. Modelling [566, 712, 957, 1976, 2315, 2445]. Models [34, 52, 128, 273, 313, 580, 847, 854, 966, 1674, 1683, 1749, 1937, 2007, 2217, 2223, 2766, 2963]. moderating [2115]. Modern [798, 1609]. modes [709, 1021, 1047, 1459]. Modified [2336]. modulation [380, 1475, 1906]. module [658, 862, 2294]. MoHoP [2693]. moisture [2767]. Molecular [353, 1805]. monitor [613, 1780, 1791]. Monitoring [76, 121, 177, 271, 335, 344, 350, 426, 959, 1076, 1100, 1111, 1202, 1214, 1227, 1284, 1308, 1359, 1504, 1588, 1680, 1835, 1849, 1891, 1934–1936, 1942, 2181, 2237, 2354, 2372, 2408, 2473, 2520, 2602, 2646, 2673, 2687, 2756, 2777, 2780, 2852]. monitors [870]. Monte [1011]. month [2992]. Montra [613]. MOSAIC [542]. MOSC [2427]. most [2097]. mostly [1282]. mostly-off [1282]. motifs [1324]. motion [1997, 2465]. motivate [1348]. Motivating [1614]. movement [1180, 1681, 2967]. MPLS [338, 756]. MPLS-TP [756]. MPLS-TP/PWE3 [756]. MPTCP [1657, 2860, 2981]. MQTT [2745]. MRC [2755]. MSAID [2811]. MSDG [2582]. MSP [2516]. MU [1675]. Multi [15, 21, 60, 75, 110, 133, 153, 221, 225, 264, 294, 299, 376, 405, 407, 429, 443, 468, 514, 561, 579, 696, 704, 766, 797, 798, 874, 884, 918, 949, 1003, 1020, 1044, 1048, 1079, 1096, 1108, 1120, 1141, 1156, 1163, 1200, 1229, 1232, 1259, 1268, 1280, 1293, 1339, 1343, 1360, 1416, 1424, 1495, 1521, 1532, 1557, 1606, 1642, 1658, 1684, 1707, 1806, 1809, 1817, 1906, 1911, 1951, 1966, 2010, 2030, 2053, 2101, 2120, 2225, 2253, 2265, 2285, 2337, 2425, 2446, 2463, 2484, 2540, 2547, 2566, 2674, 2681, 2739, 2772, 2797, 2814, 2844, 2848, 2854, 2906, 2910, 2915, 2926, 2928, 2939, 2944, 3008]. multi [17, 154, 183, 300, 309, 333, 408, 437, 500, 625, 640, 662, 671, 707, 757, 792, 837, 841, 941, 942, 994, 1046, 1082, 1176, 1214, 1222, 1276, 1300, 1319, 1509, 1652, 1675, 1692, 1718, 1720, 1803, 1846, 1851, 1872, 1897, 1901, 1907, 1922, 1967, 1993, 2068, 2150, 2215, 2216, 2236, 2289, 2349, 2469, 2512, 2513, 2519, 2559, 2578, 2694, 2711, 2766, 2872, 2883, 2898, 2955, 2971, 2986]. multi-access [264]. multi-access/multi-service [264]. multi-agent [1846]. multi-antenna [2446]. Multi-Anycast [994]. multi-application [1044]. Multi-AS [1339]. multi-attribute [2772]. multi-authority [2425]. Multi-Binomial [1707]. multi-carrier [1360]. multi-cell [2739]. multi-channel [110, 154, 405, 407, 468, 500, 671, 1048, 1082, 1214, 1222, 1806, 2030]. Multi-Cloud [1163, 1276, 2915, 3008]. multi-constrained [225, 757, 2463]. multi-constraint [2150]. multi-controller [2848]. multi-copying [2265]. multi-core [1120, 1229, 1280]. Multi-criteria [579, 2053, 2883]. multi-destination [2285]. multi-dimensional [2906]. Multi-Direction [2559]. Multi-Domain [443, 841, 884, 2337, 2349, 2540]. multi-flow [941, 1911]. multi-follower [797]. Multi-Gbps [2101]. multi-grade [376]. Multi-granularity [300, 309, 704, 1720]. multi-homed [333, 1156, 1684, 2971]. multi-homing [1967]. Multi-Hop [15, 17, 60, 294, 429, 437, 561, 696, 766, 792, 798, 918, 942, 1096, 1259, 1293, 1343, 1424, 1495, 1557, 1906, 1907, 1922, 1951, 1993, 2225, 2289, 2519, 2674, 2711, 2814]. Multi-hour [299]. multi-interface [110]. multi-layer [707, 1319, 1509, 1692, 2944]. multi-layered [625]. Multi-leader [797]. Multi-level [2484, 2766]. multi-location [2253]. Multi-MetaRing [1176]. Multi-model [2566]. Multi-objective [75, 153, 949, 1232, 1658, 1817, 2513, 2854]. multi-operator [2939]. Multi-Packet [110, 1003, 1141]. multi-path [837, 1200, 1532, 1897]. multi-phase [1901]. multi-plane [1809]. multi-point [2216]. Multi-portal [1079]. multi-provider [183, 1966]. multi-radio [21, 333, 405, 662, 1048, 1806, 2030]. multi-RAT [1319, 2926]. multi-rate [133, 221, 1046]. multi-request [2512]. Multi-resolution [1606, 2797]. multi-robot [2898]. multi-round [1675]. multi-scale [2955]. multi-sensor [1044, 2694]. multi-servers [1718]. multi-service [264, 2681]. Multi-sink [579, 2928]. multi-source [640, 1020, 2010, 2285, 2547]. multi-stage [1642]. multi-stream [1020]. Multi-Task [2844]. multi-tenancy [2986]. multi-tenant [1268, 1851, 2120, 2469]. multi-tier [408, 1416, 2068, 2578]. multi-timescale [1652]. Multi-Topology [1300, 1521]. multi-Tx [1803]. multi-Tx/Rx [1803]. multi-UAV [2236, 2872]. Multi-user [514, 1044, 1108, 1360, 1872, 2215]. multiadaptive [1053]. MultiCache [319]. multicarrier [1252]. Multicast [31, 39, 90, 119, 127, 133, 140, 143, 146, 153, 204, 295, 346, 394, 405, 417, 423, 488, 489, 503, 574, 640, 644, 663, 680, 757, 791, 793, 798, 820, 868, 882, 887, 896, 913, 958, 994, 1020, 1046, 1128, 1170, 1172, 1209, 1311, 1330, 1480, 1520, 1537, 1665, 1720, 1756, 1858, 1876, 1887, 2049, 2065, 2150, 2250, 2278, 2330, 2463, 2517, 2642, 2923]. multicast-based [295]. multicast-capable [394, 1876]. multicasting [685, 719, 918, 1585, 2285, 2402, 2656, 2892]. multichannel [27, 1632, 2302, 2897]. multicoloring [1816]. multicore [30]. Multidimensional [1942]. multifaceted [1191]. multifractal [2654]. multihomed [239, 1223, 2575]. Multihoming [1376]. Multihop [23, 287, 310, 341, 378, 424, 436, 529, 644, 898, 1000, 1168, 1262, 1997, 2055, 2204, 2262, 2537]. Multihybrid [1524]. Multimedia [109, 130, 220, 574, 621, 739, 791, 803, 1315, 1409, 1474, 1700, 1772, 1788, 2254, 2441, 2451, 2638]. multimodal [3007]. multipart [781]. multiparty [1076]. Multipath [168, 294, 326, 464, 585, 666, 686, 885, 1198, 1436, 1645, 1981, 2023, 2061, 2081, 2130, 2327, 2401, 2500, 2628, 2638, 2675, 2816, 2946, 2988]. Multiple [21, 84, 143, 163, 220, 239, 325, 633, 782, 796, 803, 866, 887, 896, 933, 1019, 1047, 1070, 1106, 1273, 1274, 1326, 1355, 1409, 1492, 1537, 1576, 1655, 1691, 1720, 1775, 1796, 1827, 1855, 1918, 2062, 2064, 2077, 2147, 2207, 2265, 2275, 2326, 2424, 2427, 2556, 2666, 2755, 2811, 2845, 2899, 2977]. multiple-access [2064]. multiple-level [1326]. multiple-mobile-routers-based [239]. multiplexed [131]. multiplexers [915]. multiplexing [667, 763]. multiplexing-based [763]. multipoint [1912]. Multipurpose [1706]. multirate [440, 644, 1674]. multisource [913, 2126, 2638]. Multiuser [1141]. Multivariate [912]. Multiview [2156]. mutual [523, 2362, 2546, 2706, 2707]. my [1405].

Name [322, 799, 1039, 1604, 1644, 1752, 2532]. name-based [322, 1752]. named [1040, 1044, 1490, 2069, 2628, 2846, 2849, 2875]. named-data [2849]. namespace [589]. NAN [1253]. nano [353, 2950]. nano-communication [353]. Nano-Things [2950]. nanonetworks [307]. narrowband [1281]. Nash [1013, 1547]. NAT [158, 1131, 1936]. natural [2236]. navigability [1864]. navigation [1432]. NCC [630]. NDN [2277, 2649]. Near [2493, 2593, 2639]. near-optimal [2493, 2593]. nearest [1648]. NECPPA [2437]. needed [800, 1094]. negatives [245]. negotiation [2706, 2850]. Neighbor [341, 1173, 1479, 1546, 2077, 2485, 2782]. neighborhood [649, 789]. NEMO [239, 894, 1791, 2517]. Nested [631, 1050, 1568]. NET [2289]. NetCluster [1049]. NetCore [1156]. NetFlow [331, 2645]. NetFPGA [1456]. Network [24, 30, 69, 113, 140, 169, 211, 242, 247, 263, 306, 369, 382, 397, 403, 416, 450, 452, 543, 544, 592, 621, 661, 702, 742, 786, 790, 824, 893, 918, 937, 955, 1081, 1112, 1116, 1127, 1152, 1155, 1203, 1227, 1243, 1268, 1279, 1280, 1316, 1320, 1355, 1376, 1503, 1519, 1541, 1547, 1596, 1622, 1637, 1655, 1701, 1709, 1725, 1741, 1742, 1749, 1814, 1870, 1907, 1923, 1929, 1935, 1936, 1963, 2004, 2092, 2174, 2188, 2204, 2263, 2336, 2345, 2381, 2386, 2415, 2418, 2543, 2617, 2686, 2689, 2701, 2796, 2799, 2827, 2857, 2886, 2892, 2914, 2915, 2942, 2962]. network [5, 66, 254, 259, 325, 329, 400, 415, 421, 424, 458, 502, 523, 562, 569, 590, 630, 660, 684, 697, 712, 780, 799, 831, 854, 865, 875, 882, 927, 989, 1073, 1076, 1096, 1098, 1139, 1142, 1149, 1171, 1196, 1200, 1249, 1281, 1307, 1400, 1431, 1451, 1471, 1475, 1552, 1617, 1630, 1640, 1641, 1688, 1696, 1702, 1737, 1739, 1757, 1810, 1820, 1873, 1912, 1934, 1937, 1971, 1983, 1987, 2080, 2111, 2151, 2152, 2191, 2232, 2238, 2250, 2265, 2287, 2313, 2372, 2389, 2413, 2452, 2500, 2514, 2527, 2576, 2592, 2597, 2631, 2773, 2814, 2873, 2875, 2897, 2901, 2902, 2925, 2945, 2982, 3008, 3013]. network [4, 105, 146, 151, 197, 216, 290, 299, 303, 365, 390, 404, 578, 583, 609, 680, 683, 714, 737, 773, 784, 830, 855, 887, 896, 970, 974, 977, 986, 1019, 1023, 1037, 1088, 1097, 1125, 1159, 1195, 1224, 1253, 1319, 1324, 1356, 1365, 1381, 1397, 1445, 1458, 1540, 1553, 1633, 1692, 1724, 1730, 1735, 1765–1767, 1782, 1787, 1788, 1858, 1864, 1868, 1958, 1967, 1988, 2003, 2011, 2028, 2062, 2126, 2224, 2251–2253, 2314, 2333, 2334, 2371, 2448, 2511, 2546, 2603, 2650, 2687, 2705, 2735, 2776, 2780, 2788, 2798, 2801, 2821, 2913, 2918, 2947, 2968, 2971, 2994]. network [17, 91, 145, 157, 162, 178, 194, 261, 335, 342, 344, 356, 378, 396, 408, 431, 438, 440, 448, 499, 582, 606, 607, 624, 676, 898, 921, 991, 999, 1041, 1046, 1053, 1110, 1178, 1366, 1382, 1420, 1446, 1483, 1505, 1555, 1574, 1585, 1638, 1646, 1659, 1668, 1703, 1723, 1738, 1762, 1768, 1795, 1812, 1819, 1877, 1885, 1894, 1905, 1920, 1928, 2015, 2035, 2084, 2088, 2097, 2150, 2172, 2182, 2215, 2241, 2246, 2248, 2303, 2311, 2430, 2453, 2482, 2552, 2566, 2579, 2607, 2625, 2653, 2662, 2668, 2673, 2802, 2822, 2841, 2842, 2845, 2867, 2868, 2932, 2952, 2957]. network-assisted [2011]. Network-aware [684, 697, 1365, 1655, 2527, 2842]. network-controlled [2092]. network-friendly [562]. network-layer [2126]. Network-level [1596]. network-supported [390]. network-traffic [1280]. network-wide [344]. network/interface [2084]. networked [1060, 1193, 1616, 1973, 2242, 2898, 2930]. Networking [15, 70, 71, 74, 319, 349, 363, 612, 671, 726, 786, 834, 842, 955, 989, 1034–1038, 1040, 1042, 1043, 1045, 1161, 1163, 1183, 1192, 1245, 1247, 1265, 1276, 1280, 1323, 1328, 1363, 1392, 1402, 1490, 1502, 1515, 1530, 1548, 1564, 1580, 1601, 1644, 1654, 1656, 1685, 1696, 1698, 1718, 1734, 1801, 1867, 2025, 2033, 2035, 2043, 2050, 2051, 2058, 2069, 2076, 2119, 2126, 2254, 2276, 2458, 2616, 2633, 2678, 2723, 2750, 2846, 2848, 2849, 2933]. networkings [2060]. networklets [2812]. Networks [67, 94, 107, 169, 199, 222, 267, 276, 282, 388, 504, 544, 563, 601, 603, 669, 678, 790, 826, 869, 900, 903, 912, 916, 940, 949, 964, 1044, 1071, 1168, 1295, 1297, 1314, 1343, 1390, 1391, 1427, 1433, 1437, 1467, 1495, 1523, 1595, 1610, 1709, 1713, 1727, 1729, 1783, 1797, 1805– 1807, 1815, 1838, 1869, 1881, 1915, 1950, 1951, 2049, 2054, 2057, 2072, 2093, 2136, 2146, 2149, 2155, 2193, 2268, 2309, 2312, 2335, 2350, 2353, 2377, 2379, 2472, 2480, 2498, 2503, 2509, 2516, 2517, 2529, 2691, 2733, 2757, 2763, 2919, 2936, 2944, 2961, 2979, 2996]. networks [47, 58, 111, 129, 135, 140, 238, 241, 248, 252, 287, 337, 372, 394, 412, 414, 418, 424, 427, 467, 517, 577, 584, 592, 622, 637, 644, 682, 699, 720, 736, 738, 764, 780, 796, 805, 828, 843, 884, 904, 918, 990, 995, 1000, 1001, 1021, 1029, 1048, 1064, 1077, 1100, 1153, 1231, 1284, 1411, 1424, 1428, 1522, 1552, 1637, 1660, 1700, 1701, 1708, 1721, 1822, 1871, 1911, 1986, 1987, 1997, 2010, 2023, 2077, 2092, 2218, 2239, 2240, 2282, 2323, 2373, 2405, 2416, 2436, 2496, 2528, 2541, 2557, 2577, 2599, 2611, 2635, 2747, 2831, 2941, 2983, 3014]. networks [138, 179–181, 217, 220, 224, 250, 260, 279, 283, 294, 385, 440, 456, 487, 491, 523, 526, 528, 547, 550, 633, 689, 723, 727, 768, 769, 795, 803, 888, 895, 902, 948, 950, 954, 1009, 1119, 1149, 1150, 1173, 1200, 1206, 1221, 1234, 1245, 1248, 1256, 1281, 1296, 1337, 1351, 1368, 1415, 1418, 1436, 1451, 1480, 1486, 1513, 1525, 1533, 1550, 1578, 1597, 1642, 1667, 1750, 1810, 1876, 1899, 1906, 1909, 1946, 1955, 1979, 1984, 2036, 2061, 2085, 2120, 2159, 2169, 2175, 2261, 2302, 2357, 2441, 2442, 2569, 2584, 2623, 2657, 2860, 2875, 2890, 2964, 3003]. networks [6, 29, 60, 92, 98, 110, 139, 164, 188, 196, 229, 264, 271, 274, 285, 325, 330, 341, 366, 380, 402, 407, 466, 468, 579, 597, 677, 704, 767, 813, 825, 914, 945, 957, 1013, 1015, 1056, 1074, 1096, 1220, 1232, 1267, 1271, 1349, 1353, 1386, 1416, 1432, 1441, 1453, 1454, 1567, 1587, 1591, 1647, 1658, 1695, 1697, 1759, 1777, 1809, 1817, 1820, 1836, 1841, 1888, 1930, 1970, 2006, 2080, 2102, 2176, 2190, 2206, 2231, 2247, 2264, 2275, 2288, 2298, 2321, 2330, 2390, 2402, 2415, 2502, 2507, 2547, 2631, 2674, 2682, 2683, 2692, 2704, 2739, 2797, 2818, 2825, 2851, 2854, 2907, 3004]. networks [10, 27, 65, 75, 82, 130, 131, 149, 184, 218, 239, 253, 262, 268, 273, 298, 302, 361, 405, 409, 423, 512, 513, 521, 531, 551, 571, 604, 605, 693, 732, 770, 793, 798, 804, 880, 897, 932, 962, 994, 998, 1061, 1091, 1103, 1167, 1201, 1205, 1224, 1261, 1262, 1302, 1307, 1325, 1327, 1336, 1340, 1352, 1373, 1377, 1404, 1498, 1544, 1557, 1604, 1627, 1681, 1716, 1788, 1831, 1845, 1866, 1891, 1954, 1972, 1989, 1998, 2034, 2109, 2112, 2138, 2197, 2219, 2243, 2258, 2272, 2296, 2308, 2319, 2332, 2362, 2382, 2393, 2407, 2410, 2539, 2606, 2661, 2665, 2700, 2729, 2798, 2885, 2899, 2901, 2945, 2965, 3012]. networks [21, 163, 172, 197, 208, 219, 230, 265, 280, 290, 334, 460, 516, 554, 583, 591, 634, 670, 696, 735, 758, 801, 838, 853, 890, 920, 939, 979, 992, 1047, 1051, 1052, 1070, 1079, 1166, 1181, 1226, 1259, 1293, 1331, 1347, 1364, 1370, 1397, 1476, 1489, 1492, 1501, 1524, 1532, 1589, 1639, 1649, 1683, 1728, 1753, 1755, 1758, 1789, 1796, 1830, 1860, 1875, 1889, 1901, 1908, 1927, 1932, 1965, 2012, 2022, 2030, 2044, 2045, 2106, 2108, 2110, 2134, 2147, 2156, 2210, 2253, 2290, 2291, 2355, 2359, 2417, 2461, 2468, 2489, 2510, 2582, 2624, 2658, 2659, 2698, 2722, 2788, 2794, 2888, 2904, 2929, 2970, 2989, 3010]. networks [8, 23, 26, 32, 35, 50, 61, 77, 96, 121, 137, 231, 237, 255, 377, 411, 434, 439, 445, 500, 533, 546, 572, 576, 588, 662, 706, 707, 792, 946, 963, 980, 1022, 1092, 1144, 1172, 1210, 1251, 1255, 1266, 1270, 1283, 1315, 1330, 1341, 1346, 1378, 1409, 1421, 1457, 1459, 1504, 1560, 1561, 1607, 1652, 1657, 1665, 1691, 1710, 1726, 1878, 1904, 1907, 1926, 1948, 1961, 1980, 1993, 1996, 2026, 2046, 2053, 2063, 2068, 2073, 2090, 2122, 2145, 2185, 2208, 2289, 2301, 2310, 2327, 2344, 2366, 2443, 2455, 2466, 2545, 2672, 2711, 2732, 2803, 2862, 2889, 2948, 2990, 2997]. networks [19, 42, 56, 64, 78, 97, 124, 126, 134, 143, 171, 223, 249, 310, 316, 323, 344, 357, 373, 374, 403, 444, 455, 477, 494, 496, 558, 663, 671, 673, 724, 725, 728, 821, 822, 893, 934, 942, 981, 1017, 1062, 1087, 1124, 1178, 1209, 1230, 1277, 1294, 1348, 1357, 1383, 1461, 1473, 1477, 1485, 1509, 1512, 1593, 1636, 1670, 1680, 1711, 1778, 1780, 1785, 1931, 1964, 2039, 2048, 2127, 2128, 2154, 2162, 2227, 2257, 2262, 2303, 2317, 2324, 2346, 2384, 2385, 2486, 2490, 2513, 2518, 2563, 2571, 2578, 2583, 2620, 2728, 2781, 2815, 2841, 2842, 2912, 2986]. networks [16, 38, 57, 73, 80, 104, 142, 189, 201, 300, 305, 309, 335, 375, 436, 437, 484, 486, 503, 515, 525, 567, 582, 646, 647, 653, 705, 721, 743, 789, 837, 841, 856, 878, 993, 1011, 1018, 1024, 1046, 1090, 1104, 1113, 1135, 1214, 1217, 1240, 1291, 1301, 1317, 1334, 1344, 1382, 1388, 1398, 1425, 1426, 1444, 1463, 1531, 1572, 1581, 1668, 1720, 1803, 1851, 1862, 1872, 1905, 1924, 2015, 2019, 2071, 2075, 2078, 2096, 2114, 2116, 2200, 2217, 2220, 2228, 2234, 2294, 2304, 2348, 2370, 2395, 2422, 2423, 2482, 2485, 2487, 2493, 2525, 2550, 2556, 2564, 2565, 2573, 2590, 2593, 2644, 2660, 2664, 2676, 2843, 2853, 2882, 2922, 2976, 3002]. networks [17, 22, 25, 45, 46, 102, 112, 190, 266, 281, 378, 393, 406, 425, 442, 449, 452, 464, 483, 540, 598, 600, 612, 628, 631, 699, 740, 785, 871, 898, 961, 976, 1004, 1028, 1082, 1106, 1117, 1122, 1130, 1282, 1326, 1387, 1401, 1410, 1422, 1464, 1472, 1491, 1493, 1500, 1542, 1546, 1559, 1583, 1629, 1631, 1646, 1663, 1687, 1717, 1731, 1776, 1802, 1832, 1892, 1898, 1922, 2040, 2083, 2086, 2107, 2173, 2179, 2192, 2212, 2230, 2231, 2236, 2255, 2306, 2316, 2322, 2336, 2356, 2367, 2420, 2429, 2465, 2467, 2491, 2495, 2561, 2587, 2607, 2628, 2638, 2643, 2697, 2727, 2751, 2787, 2837, 2956, 3011]. Networks [20, 193, 845, 1121, 1389, 2292, 2473, 2595, 2813, 2894]. neural [371, 2780, 2797, 2957]. neuro [1669, 2781]. neuro-fuzzy [1669, 2781]. neutrosophic [2858]. news [694]. Next [169, 237, 255, 283, 511, 706, 755, 981, 1064, 1223, 1455, 1632, 1908, 2156, 2190, 2191, 2270, 2619]. next-generation [255, 706, 1455, 1632, 1908, 2190]. NFA [492]. NFA-OBDDs [492]. NFV [1688, 1709, 2095, 2190, 2459, 2590, 2702, 2788, 2833]. NFV-enabled [2590]. NFV/SDN [2459]. NFV/SDN-based [2459]. NGN [2245]. NH [186]. NH-diverse [186]. niche [2037]. NICs [30]. NITOS [1186]. NLOS [2175]. no [619]. no-regret [619]. NoCs [2288]. Node [87, 503, 573, 653, 731, 732, 756, 897, 1032, 1117, 1520, 1739, 1889, 1898, 1965, 2089, 2105, 2613, 2638, 2823, 2885, 2978]. node-to-node [653]. nodes [240, 394, 459, 598, 1202, 1317, 1649, 1876, 2077, 2204, 2265, 2319, 2334, 2348, 2619, 2868]. nodewatcher [1724]. Noise [1595, 2238, 2322, 2640]. Noise-aware [2238]. Noise-induced [2640]. NOMA [2882]. Non [57, 182, 296, 526, 546, 578, 696, 998, 1141, 1489, 1536, 1830, 2005, 2142, 2334, 2615, 2627, 2637, 2700, 2796, 2908]. non- [2908]. Non-altering [2700]. Non-asymptotic [546]. non-binary [1830]. non-collaborative [182]. non-content [578]. Non-cooperative [526]. non-deterministic [696]. Non-intrusive [2615]. Non-isotonic [2637]. non-overlapping [1536]. non-QoS [998]. non-redundant [2796]. Non-saturated [1489]. non-saturation [296, 1141]. non-SDN [2334]. non-simple [57]. non-targets [2005]. non-technical [2142]. Non-VPN [2627]. nonbeacon [727]. nonbeacon-enabled [727]. nonce [39]. Nonconvex [725]. nonorthogonal [2977]. nonparametric [619]. Nornet [1157]. north [472, 2253]. north-south [2253]. not-via [100]. Note [2894, 2938, 3000]. notes [233]. notification [2299, 2380]. Novel [82, 192, 238, 263, 622, 685, 789, 917, 1121, 1130, 1249, 1315, 1326, 1477, 1573, 1588, 1646, 1769, 1960, 2039, 2124, 2137, 2142, 2164, 2201, 2209, 2287, 2336, 2437, 2452, 2517, 2523, 2582, 2595, 2659, 2699, 2758, 2826, 2870, 2905, 2983]. NP [409]. NP-Complete [409]. ns3 [2130]. nulling [1997, 2496]. number [923, 2599]. numbers [347]. Numerical [1336]. nutshell [760].

OAuth [971]. OAZE [562]. OBDDs [492]. Object [1022, 1753, 2259, 2948, 2997]. objective [75, 153, 949, 1232, 1658, 1744, 1817, 2513, 2854]. objectives [1638, 1763, 2987]. objects [1207, 1864, 1973, 2594, 2736]. oblivious [16, 2558, 2676]. OBS [56, 238, 768]. Observation [2052, 2333]. observing [1820]. obsolescence [387]. Obstacle [878]. obstructed [2301]. obstruction [35]. Occupancy [2144, 2237, 2969]. Occupancy-aided [2144]. ODT [1443]. OEF [1189]. OFA [921]. OFDM [510, 705, 1230, 1888, 2625, 2654]. OFDM-based [705]. OFDMA [123, 380, 665, 780, 785, 819, 998, 1068, 1209, 1325, 1353, 1670, 1759, 2124, 2888]. OFDMA-based [380, 1068, 1209]. OFELIA [1042, 1160]. off [245, 548, 707, 713, 949, 1282, 1462, 1637, 1887, 2105, 2416, 2798, 2880, 2907]. off-the-shelf [2798]. office [211]. Offline [1738, 2863]. offloading [1266, 1365, 1542, 1562, 1917, 1989, 2327, 2360, 2552, 2816, 2845, 2898, 2977, 3013]. offloading-enabled [1266]. offs [1220, 1715]. Oligopolistic [1304]. oligopoly [73]. OLSRv2 [1727]. OMA [782]. On-demand [31, 63, 1277, 2270, 2828, 2845]. on-line [247]. on-off [707]. on-the-road [1258]. one [61, 185, 378, 1882, 2394, 2403]. one-hop [61, 378]. one-to-many [2403]. one-to-one [2394]. one-way [1882]. onion [1063]. Online [106, 564, 578, 703, 813, 856, 962, 963, 1135, 1221, 1257, 1321, 1379, 1401, 1404, 1409, 1410, 1450, 1525, 1619, 1626, 1738, 1928, 1986, 2116, 2334, 2402, 2405, 2768, 2789, 2855, 2863, 2869]. onto [2080]. ontology [1361, 2699, 2910].

OPAC [2168]. OPCSC [564]. open [152, 256, 483, 1717, 2521, 2533, 2629, 2939]. open-source [2533]. OPEnflow [1042, 1160, 1161, 1174, 1189, 1314, 1679, 1693, 1755, 1787, 1851, 1862, 1925, 2022, 2098, 2120, 2477, 2982]. OPEnflow-based [1189, 1679, 1862, 2982]. openness [1571]. openstack [2041]. operated [19]. operating [1751, 2250]. operation [317, 490, 1445, 1461, 1800, 2122, 2216, 2890]. Operational [1375, 1534, 1688, 2075]. operations [688, 773, 1145, 1154, 2171]. operative [1913]. operator [396, 1701, 1812, 1963, 2310, 2415, 2939]. operator-friendly [396]. operators [216, 365, 1590, 1688, 1722, 1827, 2190]. OPERETTA [1679]. opinion [962]. OPNET [442]. OPNET-based [442]. Opportunistic [47, 88, 217, 279, 398, 406, 425, 457, 506, 588, 780, 834, 929, 998, 1007, 1056, 1097, 1353, 1373, 1499, 1509, 1607, 1670, 1717, 1949, 2043, 2046, 2048, 2049, 2051, 2088, 2110, 2159, 2169, 2209, 2218, 2219, 2347, 2348, 2373, 2392, 2471, 2552, 2581, 2603, 2624, 2732, 2734, 2846, 2942]. Opportunities [990, 1402, 1599, 1612, 1812, 1905, 2074, 2651]. opportunity [1130, 1871]. Optical [32, 38, 102, 165, 180, 263, 265, 300, 305, 394, 404, 461, 554, 663, 680, 693, 705, 707, 721, 736, 754, 820, 903, 915, 1100, 1144, 1164, 1329, 1445, 1463, 1475, 1522, 1586, 1591, 1665, 1667, 1876, 1902, 1906, 2138, 2149, 2155, 2583, 2625, 2665, 2944]. optically [2672]. Optimal [135, 225, 256, 257, 332, 337, 387, 429, 449, 474, 540, 552, 625, 637, 640, 678, 680, 708, 757, 768, 832, 906, 915, 928, 937, 976, 1009, 1013, 1028, 1030, 1057, 1091, 1095, 1097, 1104, 1119, 1150, 1274, 1294, 1369, 1396, 1443, 1450, 1457, 1459, 1491, 1512, 1516, 1537, 1629, 1648, 1653, 1663, 1672, 1752, 1768, 1812, 1814, 1830, 1888, 1927, 1972, 2023, 2071, 2168, 2172, 2300, 2310, 2323, 2333, 2341, 2410, 2450, 2493, 2526, 2576, 2593, 2619, 2639, 2661, 2795, 2843, 2880, 2898]. optimality [2637, 2793]. Optimally [890, 1393]. Optimising [2342]. optimistic [541, 921]. Optimization [62, 72, 78, 153, 204, 378, 403, 515, 519, 543, 545, 548, 606, 612, 660, 662, 693, 725, 796, 836, 861, 894, 895, 904, 948, 949, 1135, 1240, 1361, 1380, 1385, 1421, 1425, 1432, 1438, 1493, 1503, 1543, 1658, 1681, 1683, 1703, 1777, 1817, 1819, 1915, 1955, 1957, 1987, 2053, 2109, 2135, 2232, 2234, 2319, 2357, 2411, 2468, 2494, 2503, 2513, 2574, 2590, 2622, 2636, 2638, 2665, 2711, 2751, 2777, 2788, 2826, 2851, 2854, 2910, 2939, 2943, 2964, 2989]. Optimization-based [1240]. optimize [1376, 2401, 2945]. Optimized [48, 290, 333, 360, 458, 732, 801, 1011, 1112, 1510, 1700, 1741, 1789, 2215, 2314, 2326, 2541, 2564, 2664, 2837]. Optimizing [156, 345, 414, 535, 831, 871, 1260, 1638, 1750, 1853, 1895, 2246, 2702, 2964]. Optimum [139, 1031, 2471, 2599]. OQ [2288]. orchestrating [1192]. orchestration [733, 1577, 1983, 2500, 2576, 2833, 2924]. Order [1352, 1549, 2022, 2445, 2667]. order-preserving [2022]. ordered [2958]. ordering [811]. ORE [1607]. OREN [337]. organization [1714]. organizational [1688]. organized [1001, 1233, 1846]. organizing [157, 283, 1713]. ORGMA [2392]. Oriented [13, 55, 183, 210, 244, 415, 424, 628, 669, 954, 1219, 1266, 1363, 1370, 1444, 1535, 1543, 2051, 2086, 2233, 2493, 2844, 2868]. orthogonal [763]. OSPF [1135, 1809]. OSPF-based [1809]. other [176]. OTTs [2027]. Outage [1860, 2411]. outage-based [1860]. Outbound [928]. outbreak [1410]. outcomes [2113]. Outlier [1965, 2350, 2919, 2927]. outliers [2756]. outreach [547]. Outsourcing [1592, 2425, 2427, 2962]. ovarian [2891]. over-provisioning [958]. Over-The-Top [2580]. overbooking [1428]. Overcoming [968, 1072]. Overflow [275, 781, 1976, 2247]. Overhead [46, 76, 79, 223, 1293, 1578, 1648, 2243]. overhearing [341, 1905]. overlaid [2162]. overlap [2790]. overlapped [2367]. overlapping [1536]. Overlay [9, 121, 133, 149, 150, 156, 272, 319, 488, 489, 542, 791, 944, 974, 1204, 1358, 1486, 1576, 1691, 1767, 1786, 2235, 2500]. overlays [153, 616]. overload [116, 735, 2562]. overloads [991]. overprovisioning [800]. Overview [189, 1514, 1709, 1723, 2651]. OWL [2840]. own [848, 1724]. ownership [115, 1900]. Ozarow [1994].

p [436]. P.1203 [3014]. P2P [4, 55, 82, 134, 148, 154, 155, 157, 158, 203, 303, 342, 346, 351, 384, 389, 396, 397, 447, 525, 551, 563, 581, 604, 612, 638, 642, 656, 659, 717, 765, 779, 801, 889, 905, 931, 1015, 1019, 1118, 1126, 1150, 1166, 1204, 1213, 1285, 1299, 1308, 1348, 1383, 1468, 1498, 1539, 1546, 1659, 1731, 1965, 2087, 2173, 2286, 2526, 2645, 2731, 2771]. P2P-based [642, 1150]. P2P-like [1659]. P2P-TV [351, 397, 1285]. P2P-VoD [931]. p2pWeb [152]. pacing [1441]. Packet [32, 110, 131, 165, 223, 329, 348, 349, 370, 456, 555, 580, 652, 747, 750, 751, 782, 882, 932, 1003, 1007, 1054, 1073, 1141, 1144, 1169, 1177, 1200, 1331, 1396, 1436, 1453, 1471, 1503, 1578, 1579, 1800, 1874, 2022, 2079, 2153, 2185, 2203, 2209, 2288, 2341, 2470, 2538, 2545, 2608, 2735, 2814, 2827, 2835, 2923, 2946, 2969]. packet-pair [329]. packet-rate [348]. packet-switch [2288]. packet-switched [1453]. packets [781, 1239, 1247, 1301, 2009, 2831]. packing [915]. PACMAN [24]. PACOM [1356]. Padoc [2050]. page [108]. paging [519, 991]. PAIN [2730]. pair [329]. pairing [1006, 2133, 2918]. pairings [1605]. PANs [48]. paradigm [2320, 2844]. Paradigms [69, 2203]. Parallel [136, 917, 1944]. parallelism [2256]. parameter [2777]. Parameterized [1767]. parameterless [1147]. parameters [167, 187, 977, 1260, 2519]. parametric [917]. parasite [1308]. Parasitic [1356]. Pareto [548, 1013]. Pareto-optimal [1013]. Pareto-optimization [548]. parking [2233, 2508]. parks [202]. Part [1151, 1182]. Partial [56, 659, 822, 1544, 1875, 2066, 2200, 2338]. partially [2122, 2426]. participant [1837]. participants [744]. participation [659, 1614, 1921]. participatory [1335, 1345, 1615, 1974, 2397, 2566, 2636]. participatory-sensing-based [2566]. particle [40, 136, 204, 725, 1819, 2574, 2854, 2943]. partition [309, 1667]. partitioned [2314]. partitioning [238, 828, 1017, 1091, 1110, 1407, 2340, 2527]. partners [1250]. party [197, 1102, 1634, 2749]. PASCCC [1371]. passenger [2440, 2598]. Passing [2895]. passive [556, 1025, 1049, 2202, 2511, 2730]. Passpoint [1562]. password [997, 2929]. Path [18, 79, 80, 121, 187, 225, 311, 334, 420, 421, 533, 561, 595, 627, 766, 798, 801, 823, 837, 902, 919, 952, 1112, 1198, 1200, 1212, 1292, 1316, 1331, 1389, 1395, 1473, 1532, 1544, 1572, 1637, 1881, 1897, 1966, 1988, 2393, 2521, 2571, 2628, 2637, 2668, 2939]. path-loss [2571]. Path-specified [2628]. Path-vector [420, 627]. Pathbook [948]. paths [35, 235, 564, 585, 594, 972, 1791, 2039, 2121, 2626]. PathZip [1331]. pattern [310, 485]. pattern-based [485]. Patterns [241, 578, 968, 1088, 1452, 1930, 1996, 2123, 2440, 2586, 2784]. Payload [347, 874, 1134, 2332]. Payload-based [874]. PCA [2705]. PcapWT [1471]. PCE [1966]. PCI [1417, 2165]. PCIR [155]. PCN [161]. PCN-based [161]. peak [2584]. Peer [54, 93, 101, 150, 155, 157, 159, 243, 391, 446, 487, 501, 562, 572, 614, 648, 675, 684, 697, 787, 802, 880, 944, 957, 1026, 1132, 1140, 1225, 1261, 1279, 1327, 1394, 1513, 1783, 1786, 1828, 1921, 1930, 2109, 2136, 2173, 2386, 2512, 2574]. Peer-assisted [675, 957, 2136]. Peer-to-Peer [54, 93, 101, 150, 157, 243, 391, 446, 487, 501, 562, 614, 648, 684, 697, 787, 802, 880, 944, 1026, 1132, 1140, 1225, 1261, 1279, 1327, 1394, 1513, 1783, 1786, 1828, 1930, 2109, 2173, 2386, 2512]. peers [396, 1394, 2160]. penalty [1377]. People [2581, 2646]. per-bit [1286]. Per-flow [517]. per-link [2528]. per-packet [2835]. Per-station [383]. Per-user [1922]. perceive [2020]. perceived [667, 1302]. perfectly [1994]. Performance [47, 55, 61, 108, 116, 128, 131, 178, 184, 198, 264, 379, 380, 385, 463, 505, 518, 529, 533, 548, 648, 687, 711, 752, 754, 804, 840, 860, 884, 899, 913, 928, 989, 996, 1003, 1055, 1077, 1078, 1093, 1131, 1139, 1141, 1147, 1186, 1200, 1234, 1243, 1271, 1302, 1310, 1316, 1376, 1417, 1469, 1495, 1522, 1538, 1647, 1711, 1753, 1757, 1779, 1862, 1874, 1908, 1941, 1953, 1983, 2007, 2008, 2047, 2061, 2115, 2184, 2223, 2259, 2261, 2409, 2447, 2507, 2542, 2552, 2553, 2580, 2587, 2617, 2635, 2649, 2668, 2704, 2798, 2831, 2907, 2936, 2943, 2968, 2995, 3008, 3015]. performance [38, 91, 237, 450, 522, 535, 546, 566, 620, 626, 713, 734, 746, 871, 980, 1072, 1285, 1321, 1478, 1483, 1586, 1596, 1715, 1829, 1886, 1957, 1980, 2002, 2063, 2083, 2205, 2220, 2257, 2304, 2438, 2473, 2487, 2495, 2497, 2560, 2730, 2913, 2950]. performance-cost [2907]. performances [436, 634]. Peri [856]. Peri-Watchdog [856]. period [1992]. Periodic [752, 2029]. periphery [856]. permission [2830]. permittivity [2767]. Persistent [1594, 1743, 1998, 2003, 2799]. personal [2019]. Personalized [361, 1592, 2280]. personnel [2694]. perspective [440, 610, 628, 787, 952, 1144, 1399, 1561, 1650, 1762, 1856, 1931, 2609, 2633]. perspectives [654, 1143, 1587, 1744, 2254, 2714, 3013]. persuasion [2879]. perturbation [1588]. Pervasive [1594, 2044]. pervasiveness [1404]. PEV [1715]. PGRP [2555]. phase [235, 1179, 1794, 1901, 2061, 2322, 2726]. phases [734]. phenomena [1504]. phenomenon [813]. PHOABE [2425]. phone [1216]. phones [482]. Photonic [2672]. Photonic-Frame-Based [2672]. PHY [51]. Physical [307, 511, 1758, 1816, 1918, 1982, 1991, 1992, 1995, 1999, 2043, 2140, 2141, 2162, 2220, 2322, 2334, 2349, 2434, 2449, 2505, 2506, 2548, 2753, 2777, 2823, 2876, 2928]. Physical-layer [1991, 1995, 1999, 2322]. Physically [52]. Physically-based [52]. physics [136]. physics-inspired [136]. PIaaS [2233]. PiCasso [2986]. piece [2526]. piggybacked [2501]. Piggybacking [1980]. pinball [631]. ping [2791]. pipeline [2756]. Pipelined [1874, 2167]. pitfalls [472, 1072]. PKI [563]. Placement [20, 94, 693, 721, 732, 831, 915, 1316, 1454, 1493, 1600, 1655, 1736, 1737, 1868, 2018, 2060, 2159, 2168, 2251, 2258, 2468, 2541, 2563, 2695, 2729, 2788, 2869, 2965]. placements [1105, 2075]. plane [400, 1041, 1463, 1691, 1809, 2020, 2076, 2394, 2735, 2835, 2933, 2954, 2967]. planes [2618]. planetary [493]. planetary-scale [493]. PlanetLab [493, 1186]. planned [2558]. planning [192, 299, 444, 450, 995, 1417, 1445, 1561, 1658, 1839, 1927, 2165, 2182, 2316, 2342, 2641, 2668]. plateaux [1917]. Platform [542, 984, 1114, 1157, 1164, 1225, 1783, 1959, 2403, 2533, 3011]. platform-centric [2403]. platforms [30, 43, 1269, 1399, 1408, 2223, 2613]. play [1219, 1332, 1455, 2204]. playback [1618]. playout [226]. PLC [1779, 2808, 2972]. PLC/wireless [2972]. plug [2250]. plug-in [2250]. plus [2212]. PMIP [364]. PMIPv6 [1385]. PNPL [2676]. PoC [782]. point [181, 626, 1191, 1754, 2123, 2216, 2410, 2529]. points [454, 624, 2145]. poisoning [1086]. poisoning-based [1086]. policies [267, 545, 638, 841, 1409, 1674, 2011, 2226]. Policy [420, 426, 527, 569, 722, 762, 877, 985, 1059, 1422, 1524, 1955, 1973, 2033, 2208, 2327, 2425, 2426, 2470, 2471, 2514, 2606, 2807, 2917]. Policy-Based [426, 722]. policy-capable [420]. policy-constrained [1524]. polling [2761]. Pollution [391, 889, 1040, 1327, 1468, 1490, 1646, 2371]. polymorphic [593]. polynomial [1924]. POMDP [2069]. PON [925, 1455, 2031]. PONs [2993]. pool [1260]. pooling [2300]. PoP [608]. popular [639, 1419, 1841]. Popularity [346, 1140, 1753, 1828, 2037, 2544]. Popularity-based [1828]. population [1610]. port [347, 683, 704]. port-cost-efficient [704]. portable [1677]. portal [1079]. Position [504, 1084, 2559]. positioning [540, 1598, 1643, 1805]. positives [245, 1112]. post [1628]. post-disaster [1628]. Potential [258, 1038, 1231, 1349, 2131, 2715]. Potentials [930, 2383]. Power [52, 75, 146, 219, 283, 372, 375, 398, 484, 515, 517, 518, 575, 700, 704, 709, 710, 712, 738, 830, 831, 901, 930, 940, 953, 1018, 1052, 1104, 1206, 1209, 1241, 1281, 1291, 1307, 1312, 1320, 1336, 1357, 1384, 1388, 1453, 1456, 1457, 1492, 1512, 1600, 1683, 1750, 1754, 1759, 1780, 1788, 1819, 1830, 1852, 1860, 1872, 1894, 1961, 2021, 2059, 2210, 2272, 2335, 2371, 2381, 2382, 2422, 2480, 2481, 2495, 2501, 2619, 2626, 2642, 2711, 2739, 2744, 2794, 2863, 2902, 2904, 2923, 2928, 2960, 2980]. power- [704]. Power-aware [1018, 2382]. power-delay [1683]. Power-efficient [1291, 1600]. power-saving [146, 2980]. powered [217, 1764, 1888, 2382, 2446, 2804, 2902]. PPCU [2835]. PPM [1324]. PPM-based [1324]. PR [1093]. PR-SCTP [1093]. Practical [60, 83, 422, 1142, 1551, 1678, 1804, 1913, 2365, 2421, 2470, 2725, 2913, 2948, 2997]. practice [740, 1375]. Practices [1617, 1620]. Pre [370, 431, 2038, 2299, 2980]. pre-classifier [2980]. pre-congestion [2299]. pre-distribution [431]. pre-filtering [370]. Pre-scheduled [2038]. preamble [490, 599]. precise [2445, 2479]. precision [1643, 2608]. precomputed [1875]. predefined [680]. predict [1728, 2405, 2906]. predicted [2784]. Predicting [34, 580, 1985, 2113, 2544, 2650].

Prediction

[313, 520, 554, 605, 639, 665, 1566, 1682, 1716, 1827, 1840, 2037, 2100, 2145, 2211, 2379, 2584, 2624, 2644, 2752, 2759, 2948, 2992, 2997]. Prediction-based [1682, 1716, 2379, 2759]. predictive [16, 1895, 2282, 2555, 2643, 2782]. predistribution [2978]. preemptive [2697]. Preface [278]. preference [1712, 2793]. preferential [1427]. prefetching [108]. prefix [242, 313, 750, 2167]. prefix-based [750]. prefix-level [242]. prefixes [639, 1536]. Preplanned [663]. presence [198, 312, 459, 592, 711, 2133, 2183, 2523, 2820]. Present [1227]. preservation [1592, 1719, 2365]. preserve [1447]. Preserving [202, 716, 933, 1010, 1221, 1588, 1837, 1838, 1856, 2022, 2045, 2161, 2173, 2328, 2419, 2437, 2454, 2520, 2540, 2632, 2718, 2719, 2723, 2778, 2940, 2962]. Presto [1928]. prevent [2342]. Preventing [203, 2247]. prevention [1327, 2951]. previous [1538]. Price [172, 218, 418, 1181, 1768, 2908]. pricing [625, 1091, 1199, 1337, 1798, 1802, 1812, 2194]. pricing-based [1802]. primary [774, 2568, 2970]. primate [1427]. PRIME [1544]. principal [676]. principle [1241, 1334]. principles [97, 316, 670, 1983]. PRINGL [1613]. prioritization [688]. prioritized [1953, 2620, 2837]. Prioritizing [1964]. Priority [3, 1236, 1242, 1371, 1661, 1675, 1802, 2062, 2727]. Priority-aware [1802]. Priority-based [1371, 1675, 2727]. Privacy [202, 214, 716, 777, 810, 933, 973, 1010, 1014, 1208, 1221, 1409, 1423, 1435, 1466, 1588, 1592, 1719, 1837, 1856, 1861, 2045, 2143, 2151, 2161, 2173, 2199, 2233, 2240, 2267, 2328, 2419, 2437, 2454, 2506, 2516, 2520, 2535, 2594, 2632, 2661, 2693, 2703, 2708, 2712, 2718– 2720, 2723, 2749, 2778, 2840, 2940, 2960, 2962]. Privacy-aware [214, 1466, 2143, 2151, 2661]. privacy-conscious [2233]. Privacy-preserving [202, 716, 933, 1010, 1221, 1588, 1837, 1856, 2161, 2173, 2419, 2437, 2454, 2520, 2632, 2718, 2719, 2723, 2778, 2940, 2962]. Private [373, 481, 600, 1009, 1137, 1171, 1250, 2169, 2598]. PrivICN [2723]. proactive [17, 1657, 2309, 2579, 2765]. Probabilistic [85, 125, 399, 481, 547, 1012, 1431, 1595, 1868, 2072, 2133, 2237, 2539, 2732]. probabilities [275, 1674]. Probability [721, 1351, 2079, 2455, 2621, 2734]. probes [2798]. probing [560, 1108, 1769]. Problem [20, 75, 94, 181, 302, 409, 444, 599, 636, 662, 669, 693, 757, 768, 949, 1094, 1338, 1648, 1683, 1881, 1898, 2026, 2060, 2108, 2228, 2285, 3003]. problems [18, 654, 767]. procedure [1127, 1264, 1963, 2665]. procedures [360]. process [865, 965, 1064, 1101, 1811, 1970, 2654]. process-based [1064]. processing [1571, 2246, 2264, 2735, 2772, 2776, 2866]. processor [116]. producer [2846, 2984]. product [449, 1217]. productivity [415]. productivity-oriented [415]. professionals [2115]. Professor [1237]. profile [2218, 2871]. profile-casting [2218]. profiles [104, 2316]. profiling [242, 1456, 1940, 2865]. profit [627, 1031, 1391]. prognosis [472]. programmability [1269]. programmable [30, 1153]. programming [139, 162, 545, 558, 769, 1024, 1443, 2676, 2751, 2868, 3003]. progress [1854]. progressive [574, 2484, 2597]. project [621]. projections [497]. projects [471]. promise [472]. Promises [742]. Promoting [244]. prone [2447]. proof [2163, 2924]. proof-of-concept [2163]. Propagation [737, 904, 962, 1372, 1944, 1970, 2110, 2175, 2588, 2611, 2876, 3010]. proper [100]. properties [392, 691, 1346, 1608, 1986, 2228, 2428]. Proportional [107, 380, 945, 1378, 1546, 2835]. proportionality [2800]. proposal [1207, 1707, 2287]. prospects [2383]. protect [717]. protected [265, 758]. Protecting [377, 820, 1435]. protection [224, 544, 718, 776, 822, 827, 829, 952, 1355, 1515, 1708, 1861, 2206, 2411, 2703, 2749, 2840, 2917]. protective [1512]. Protocol [24, 105, 143, 197, 233, 236, 306, 385, 407, 420, 423, 435, 466, 481, 528, 584, 618, 637, 644, 677, 689, 720, 726, 795, 805, 825, 850, 879, 904, 916, 970, 1007, 1022, 1023, 1061, 1071, 1134, 1146, 1167, 1177, 1185, 1251, 1254, 1281, 1282, 1352, 1361, 1371, 1398, 1415, 1432, 1474, 1535, 1639, 1757, 1779, 1780, 1838, 1845, 1887, 1891, 1899, 1900, 1959, 2016, 2023, 2065, 2202, 2222, 2262, 2270, 2272, 2282–2284, 2293, 2298, 2338, 2343, 2354, 2373, 2386, 2398, 2479, 2516, 2555, 2577, 2579, 2620, 2674, 2693, 2701, 2750, 2758, 2808, 2827, 2905, 2935, 2948, 2972, 2997]. protocol [45, 117, 266, 320, 483, 618, 647, 648, 744, 894, 943, 1102, 1212, 1892, 2150, 2329, 2331, 2409, 2628, 2676, 2734, 2787, 2838]. Protocol-level [879]. protocol-oblivious [2676]. Protocols [38, 115, 146, 168, 247, 255, 315, 364, 393, 455, 458, 467, 490, 587, 729, 970, 996, 1003, 1025, 1051, 1078, 1107, 1196, 1297, 1306, 1399, 1418, 1473, 1526, 1726, 1895, 2108, 2209, 2239, 2321, 2379, 2572, 2881, 2968]. prototyping [596]. Provable [266]. Provably [1102]. provenance [2116]. provide [977, 1436, 1981, 2284]. Provided [281]. provider [183, 1966]. providers [700, 873, 1873, 2190]. Providing [186, 544, 680, 1373, 2516, 2693, 2778]. provision [301, 525, 1534]. Provisioning [238, 274, 325, 345, 822, 873, 958, 1270, 1665, 1855, 1963, 2066, 2155, 2904]. proxies [498, 1393]. Proximity [1279, 1479, 1960, 2043, 2050, 2369, 2728]. proximity-based [1479]. Proxy [93, 527, 1330, 2267, 2356, 2672]. pruning [702, 2558]. PSO [1370, 2422]. PsyBoG [1793]. PTP [2628]. public [498, 1233, 1274, 1581, 2063, 2321, 2728, 2844]. public-cloud [2063]. Publication [1225, 1783]. Publish [134, 674, 741, 883, 908, 1078, 1549, 1767, 2304, 2737]. publish-subscribe [1078, 1767]. publish/subscribe [134, 674, 741, 883, 908, 1549, 2737]. Publisher [2894, 2938, 3000]. publishing [2598]. pull [447, 2386, 2512]. pull-based [2512]. pulse [2294]. pulse-coupled [2294]. pure [802, 1132, 2572]. pursuit [676]. push [2386]. push-pull [2386]. Pushing [291]. PUT [1407]. puzzles [717, 1781]. PWE3 [756].

Q [1319, 1852, 2162, 2271]. Q-Learning [1319, 2162, 2271]. Q-learning-based [1852]. QoE [357, 935, 1503, 1516, 1618, 2027, 2212, 2655, 2688, 2689, 2992, 3014]. QoE-aware [2689]. QoE-centric [2027]. QoE-driven [1503]. QoI [1837]. QoI-aware [1837]. QoM [258]. QoS [88, 130, 220, 222, 238, 274, 345, 366, 369, 376, 393, 426, 513, 550, 557, 565, 621, 625, 738, 761, 838, 843, 918, 932, 940, 998, 1283, 1298, 1367, 1382, 1384, 1425, 1432, 1463, 1464, 1505, 1544, 1705, 1712, 1755, 1764, 1770, 1772, 1789, 1809, 1840, 1855, 1860, 1943, 1956, 2028, 2066, 2172, 2278, 2391, 2452, 2592, 2620, 2627, 2686, 2816, 2821, 2904, 2939, 3004]. QoS-Aware [393, 761, 838, 932, 940, 1505, 1764, 1772, 1789, 1809, 2172, 2686, 2816]. QoS-based [843, 1382]. QoS-Classifier [2627]. QoS-compliant [3004]. QoS-demanding [550]. QoS-driven [1712]. QoS-guaranteed [1283, 1956]. QoS-path [2939]. QoS-sensitive [222]. QoT [2138]. QoT-aware [2138]. QR [977]. QRM [1708]. qualitative [1709]. Quality [58, 227, 250, 488, 667, 751, 792, 795, 861, 1105, 1275, 1287, 1292, 1360, 1369, 1393, 1474, 1601, 1614, 1728, 1853, 1963, 2009, 2215, 2242, 2284, 2384, 2405, 2580, 2773, 2908, 2992]. Quantifying [336, 465, 589, 2090, 2123, 2913]. quantitative [1076, 2580]. quantization [422]. quantized [896, 2380]. Quantum [2141]. Quantum-inspired [2141]. Queen [689]. Queen-MAC [689]. queries [684, 697]. query [2772]. querying [572]. Queue [244, 275, 449, 1079, 1147, 1647, 1660, 1708, 2201, 2205, 2568, 2662]. Queue-aware [2662]. queue-length-based [2201]. queue-overflow [275]. Queueing [329, 1244, 1862, 2080, 2681, 2685, 2982]. Queues [116, 304, 1242, 1246, 1800, 1925]. Queuing [120, 281, 422, 713, 751, 1661, 2451]. QUIC [2745]. Quorum [146, 689, 1426]. Quorum-based [146, 689].

Rabin [1005]. race [631]. RAD [498].

Radiation [2230]. Radiation-constrained [2230]. Radio [21, 29, 73, 84, 188, 189, 264, 274, 283, 301, 333, 345, 359, 362, 375, 405, 526, 559, 576, 662, 677, 725, 771, 797, 819, 937, 945, 995, 1048, 1052, 1075, 1106, 1127, 1167, 1350, 1400, 1428, 1485, 1496, 1499, 1599, 1627, 1628, 1631, 1636, 1806, 1807, 1819, 1871, 1872, 1922, 1931, 1954, 2030, 2122, 2127, 2128, 2181, 2215, 2305, 2342, 2432, 2467, 2493, 2543, 2565, 2568, 2571, 2573, 2587, 2673, 2675, 2747, 2751, 2803, 2818, 2837, 2907, 2918, 2947, 2970, 2989]. radio-based [345, 2803]. radio-enabled [2818]. radios [439, 722, 1727]. Radius [581]. RAGE [591]. RAM [1748]. RAN [1710, 2469, 2510, 2614, 2741, 3009]. Random [46, 61, 125, 303, 348, 559, 1129, 1381, 1596, 1632, 1803, 1875, 1971, 2000, 2046, 2079, 2367, 2395, 2571, 2971]. randomized [528, 1119, 2665]. Randomizing [2332]. Range [33, 195, 307, 353, 546, 643, 684, 697, 789, 1681, 2153, 2305, 2799, 2872, 2897]. Range-dependent [2799]. range-free [789]. ranging [2675]. ranking [993, 2484, 2942]. ransomware [2363]. Rao [2457]. Rapid [591, 596, 618, 1639, 2868]. RAT [633, 1064, 1319, 2092, 2926]. Rate [105, 133, 161, 200, 219, 221, 230, 348, 374, 375, 448, 597, 629, 630, 778, 804, 899, 940, 1046, 1072, 1262, 1285, 1367, 1514, 1534, 1551, 1583, 1637, 1708, 1794, 1852, 1989, 2094, 2212, 2269, 2324, 2421, 2481, 2792, 2851, 3003]. rate-based [1989]. rate-diverse [1072].

Rate-sensitive [2212]. rates [896, 1918, 2888]. rating [2752]. ratio [289, 823, 1058, 2923]. ratio-maximized [289]. raw [2718]. Rayleigh [2783]. RCA [1386]. RDSR [8]. RDSR-V. [8]. Re [186, 450, 736, 1064, 2024]. re-distribution [186]. re-planning [450]. Re-routing [2024]. re-selection [1064]. REACT [202]. REActing [2492]. Reactive [1081, 2925]. reader [1677]. readers [481]. reading [1129]. Real [62, 86, 109, 269, 289, 486, 510, 557, 566, 853, 874, 916, 918, 927, 976, 1085, 1251, 1467, 1552, 1588, 1669, 1682, 1718, 1725, 1742, 1791, 1886, 1893, 1943, 2100, 2156, 2237, 2269, 2297, 2316, 2482, 2581, 2719, 2738, 2767, 2779, 2803, 2853, 2873, 2951]. real-device [1552]. real-life [1725, 2482]. Real-time [62, 86, 109, 269, 289, 486, 510, 557, 566, 853, 874, 916, 918, 927, 976, 1085, 1251, 1588, 1669, 1682, 1718, 1886, 1893, 1943, 2100, 2156, 2237, 2269, 2581, 2719, 2738, 2767, 2803, 2853, 2873, 2951]. real-world [2297]. Realising [1045]. realistic [429, 786, 1930, 2982]. reality [610, 814]. Realization [952, 2971]. reallocation [2816]. really [972]. Reasoning [722, 1235, 2231]. REATO [2492]. rebalancing [136]. rebuffering [194]. recapture [59, 1610]. receive [375, 622, 1198]. received [1643]. Receiver [13, 510, 530, 1418, 1858]. Receiver-initiated [1418]. Receiver-oriented [13]. receivers [2599]. Receiving [2896]. Reception [110, 1333, 1540, 1597]. Rechargeable [50, 1454, 1836, 2639, 2697]. Recipient [498]. Recipient-anonymous [498]. reciprocity [638]. RECODAN [2040]. recognition [1088, 2147]. recognizing [851]. recommendation [1212]. recommendations [477]. reconciliation [1999]. reconfigurable [915, 2031]. reconfiguration [438, 548, 937, 1429, 1430, 1578, 1739, 1740, 2172, 2345, 2592]. reconnection [445]. reconstruction [77]. Recoverable [2366]. recovery [223, 585, 773, 882, 1213, 1523, 1628, 2089, 2317, 2722, 2755, 2845, 2951]. recruiting [2726]. rectangles [2124]. rectification [2530]. recurrent [2797]. Recursive [318, 648, 1242]. recycling [2446]. RED [83, 1975]. Redesigning [154]. reduce [223, 581, 831]. reduced [599, 1420]. Reducing [1206, 1384, 1695, 2039, 2165, 2197, 2515, 2872, 2976]. Reduction [79, 300, 553, 1626, 1923, 2432, 2439, 2510, 2705]. redundancy [332, 1118, 1213, 1301, 1500, 1507, 1539, 1810, 2040]. redundancy-based [1301]. redundant [590, 1648, 2039, 2796]. REED [769]. reference [2165, 2600]. refinement [837]. reflection [1725]. REFRAHN [1526]. region [515, 1868]. Regional [2701]. regions [182, 2698]. regression [619, 1778]. regret [619]. Regrouping [1898]. regular [492]. regularity [864]. regularity-based [864]. regularized [427]. Regulated [789, 2799]. regulations [2461]. Reinforcement [359, 1144, 1421, 1601, 1801, 2887, 2920]. Reinforcement-learning-based [1801]. reinjection [1917]. related [475, 834, 2627]. relation [4, 2249]. relationship [923, 1820, 1877, 2580]. relationships [649]. relative [2218]. relaxed [676]. Relay [124, 516, 685, 732, 780, 798, 992, 1104, 1259, 1262, 1316, 1325, 1486, 1496, 1550, 1596, 1639, 1810, 1822, 1826, 1872, 1888, 1889, 1894, 1912, 2424, 2587, 2638, 2804, 2851, 2902, 2918, 2979]. relay-assisted [1822, 2979]. Relay-based [1912]. relay-enabled [685, 2851]. relaying [1000, 1670, 2804]. relays [1303]. Release [407]. Release-time-based [407]. relevant [1452]. Reliability [102, 822, 908, 1173, 1179, 1476, 1480, 1677, 1749, 1887, 2119, 2305, 2429, 2814]. Reliable [8, 140, 369, 644, 719, 728, 805, 882, 918, 1055, 1175, 1301, 1507, 1619, 1887, 1994, 2067, 2147, 2265, 2297, 2392, 2662, 2666, 2667, 2867, 2935]. reliant [1527]. relocation [1648]. REmedy [1679]. remote [823, 2471, 2715]. removing [980]. rendezvous [648, 2493, 2649]. rendezvous-based [2649]. renewable [1357, 1764]. reorganization [762]. repacking [529]. repair [2314]. repeatable [2163]. RePIDS [874]. replacement [832, 1490]. replay [927]. replica [1166, 2695]. Replication [61, 673, 1071, 1089, 1437, 1711, 1979, 2154, 2192]. report [1772]. reporting [1258, 1360, 2066]. reports [576]. repository [2840]. representation [2419, 2797]. representative [2097]. Reprint [809, 811, 1238]. reproducible [1192, 1196]. Reputation [112, 214, 243, 889, 1032, 1465, 1558, 1615, 1761]. Reputation-based [243, 1465]. request [63, 173, 1062, 1534, 2023, 2095, 2211, 2389, 2512, 2527]. requested [1481]. requests [898, 1291, 1977, 2590, 2812]. Required [800]. Requirements [629, 875, 986, 1253, 1358, 1416, 1482, 1558, 2210]. reroute [100, 594, 1520]. rerouting [1290, 2095]. rerouting-based [1290]. rescheduling [56]. rescue [773]. RescueNet [1801]. research [71, 471, 538, 1156, 1158, 1159, 1328, 1402, 2114, 2629, 2713]. Reservation [267, 1379, 1674, 2083, 2467, 2521, 2831]. reservations [414]. residential [1349, 1609, 2226]. ResiDI [2456]. resili [2344]. resili-ent [2344]. Resilience [7, 97, 101, 743, 761, 1105, 1399, 1526, 1607, 1685, 2532, 2729]. resiliency [99, 1390]. Resilient [96, 98, 445, 1126, 1169, 1485, 1528, 1607, 1736, 1832, 1882, 2127, 2128, 2152, 2912]. resistant [340, 632, 744, 1415]. resizing [659]. resolution [1103, 1606, 1644, 2797]. resolve [1094]. resolvers [2008]. resolving [1257]. Resource [29, 84, 131, 171, 252, 271, 301, 303, 359, 363, 440, 493, 516, 525, 529, 535, 600, 665, 668, 692, 780, 819, 992, 998, 1095, 1127, 1132, 1175, 1188, 1203, 1230, 1240, 1241, 1259, 1269, 1272, 1298, 1308, 1350, 1353, 1481, 1524, 1550, 1577, 1591, 1599, 1600, 1651, 1692, 1758, 1772, 1785, 1813, 1823, 1826, 1827, 1875, 1888, 1929, 1963, 2000, 2120, 2162, 2222, 2226, 2235, 2266, 2279, 2300, 2346, 2358, 2359, 2389, 2444, 2469, 2518, 2543, 2565, 2584, 2592, 2656, 2658, 2669, 2702, 2727, 2739, 2774, 2788, 2834, 2863, 2882, 2888, 2917, 2918, 2921, 2947, 2951, 3009]. resource-constrained [1813, 2226]. resource-guaranteed [1600]. resource-provision [525]. resources [48, 453, 1094, 1792, 1976, 2062, 2448, 2734]. response [256, 966, 2378, 2462, 2976]. Responsive [2443]. restaurant [1631]. restoration [417, 663, 2302]. Restricted [2617]. restriction [2439]. restructuring [1405]. results [18, 128]. retail [2194]. Rethinking [943, 2331]. retouched [245]. retransmission [1056, 1474, 2081]. retransmissions [2088]. Retrieval [614, 1118, 1225, 1783, 1950, 1952, 2058, 2154, 2241, 2312, 2597, 2723]. Reuse [20, 94, 946, 990, 1593]. Revealing [2070]. revenge [848]. revenue [267, 414, 2188]. reverse [850, 1007]. reverse-engineering [850]. review [477, 506, 749, 875, 1312, 2114, 2216, 2629, 2710]. reviewers [1413, 1746]. revision [2113]. revisited [2426]. Revisiting [902, 1378, 2820]. revocable [744]. revocations [337]. reward [1000, 1405, 2119]. REX [2152]. RFC [2153]. RFID [115, 202, 257, 481, 556, 777, 953, 1254, 1677, 1857, 1900, 1913, 2182, 2202, 2758, 2761, 2985]. RFID-based [202]. RFL [2874]. Rich [1766]. riders [551]. rigid [890]. ring [102, 618, 2039, 2232, 2635, 2703]. ring-based [2039, 2232]. ring-loop [2703]. rings [1169, 1545]. RIPE [2409]. RISC [2221]. rise [2363]. risk [54, 462, 627, 966, 1594, 2138, 2433, 2715]. risk-based [966]. RIVER [805]. RLNC [1381]. RMLSA [2149]. Road [1258, 1423, 1590, 2292, 2605]. Road-side [1590]. ROADM [980]. roadmap [1314, 1916]. roadside [474]. roadway [2694]. roaming [266]. robin [500]. robot [2898]. robotic [1714, 2465, 2925]. robots [387, 2692]. Robust [252, 431, 437, 489, 682, 704, 868, 1090, 1351, 1360, 1381, 1398, 1531, 1624, 1714, 1782, 1860, 2158, 2170, 2210, 2312, 2322, 2331, 2361, 2494, 2603, 2666, 2675, 2707, 2874, 2888, 2904, 2929]. robustness [1070, 1390, 1844, 1877, 2251, 2933, 2981]. role [365, 841, 2115, 2364]. role-based [841]. roles [2472]. round [500, 1675]. Route [173, 186, 187, 310, 571, 636, 894, 1385, 1527, 1831, 2273, 2639, 2641, 2677]. route-aware [310]. routed [138, 706]. Router [114, 304, 894, 1454, 1456, 1470, 2569]. routers [239, 293, 639, 692, 1059, 1203, 1229, 1508, 2256]. routes [180, 646, 691, 799]. Routing [6, 13, 27, 86, 90, 112, 139, 217, 220, 271, 274, 294, 297, 306, 420, 452, 457, 464, 467, 498, 506, 561, 584, 601, 645, 703, 736, 805, 831, 872, 884, 916, 919, 954, 976, 1020, 1038, 1130, 1146, 1149, 1177, 1235, 1306, 1399, 1432, 1469, 1475, 1486, 1495, 1522, 1523, 1526, 1550, 1567, 1580, 1600, 1601, 1642, 1690, 1700, 1752, 1777, 1809, 1836, 1852, 1881, 1890, 1906, 1951, 2020, 2028, 2093, 2111, 2176, 2218, 2270, 2330, 2336, 2373, 2389, 2392, 2401, 2509, 2517, 2521, 2531, 2537, 2555, 2557, 2577, 2638, 2674, 2686, 2701, 2755, 2796, 2849, 2924, 2935, 2942, 3015]. routing [16, 18, 26, 65, 78, 121, 145, 168, 183, 187, 204, 219, 231, 235, 249, 272, 273, 312, 322, 324, 334, 405, 413, 416, 423, 484, 572, 585, 670, 701, 721, 729, 746, 757, 762, 827, 837, 880, 886, 939, 1012, 1018, 1092, 1097, 1121, 1133, 1185, 1251, 1290, 1291, 1293, 1300, 1357, 1367, 1429, 1488, 1509, 1527, 1528, 1572, 1604, 1720, 1726, 1789, 1791, 1803, 1867, 1869, 2033, 2045, 2065, 2099, 2106, 2108, 2121, 2150, 2200, 2204, 2209, 2220, 2221, 2272, 2276–2278, 2283, 2289, 2320, 2335, 2391, 2428, 2556, 2558, 2579, 2603, 2604, 2637, 2687, 2703, 2732, 2781, 2801, 2949, 2953, 2976, 2991]. routing [8, 45, 631, 1063, 1222, 1521, 1636, 1645, 1717, 1892, 1911, 2024, 2248, 2586]. routing-based [954]. routing-delivery [2218]. routing-policy [762]. RPL [760, 2272]. RRH [2989]. RRM [545]. RS [1822]. RSS [2675]. RSS-based [2675]. RT [1664]. RTCP [739, 1305]. RTCP-based [739, 1305]. RTPT [2719]. RTSP [782]. RTSP-based [782]. RTT [176, 381]. RTT-fair [176]. RTTs [491]. RTXP [1251]. Rule [582, 769, 1694, 2323, 2616]. rule-based [582, 2323]. RuleBender [2470]. rules [945, 1011, 2195, 2479, 2842]. ruling [2154]. rumor [1970]. running [2777]. runtime [478, 769]. rural [1722]. RWA [706]. Rx [1803].

s [673, 1407, 1439, 1761, 2798, 2993]. S-BITE [1439]. S-CLONE [673]. S-Kademlia [1761]. S-PUT [1407]. S2/ETSI [565]. SAE [1757]. SAE/LTE [1757]. Safeguarding [1001]. SafePath [2598]. Safety [429, 1891, 2324, 2605, 2708, 2728, 2787]. salvaging [2209]. SAM [2020]. sample [1847, 2484, 2653]. sampled [2696]. sampler [1214]. Sampling [80, 144, 159, 177, 331, 348, 490, 599, 652, 1053, 1503, 2006, 2183, 2688]. sandstorm [335]. SAP [1975]. SAP-LAW [1975]. SARD [1694]. satellite [231, 726, 1145, 2698]. satellites [461]. SATI [269]. satisfaction [374, 2592]. Satisfy [2890]. Satisfying [1224]. saturated [1489]. saturation [296, 1141]. saving [146, 300, 398, 518, 701, 702, 709, 710, 930, 1060, 1336, 1650, 1765, 1788, 2393, 2980]. savings [290, 1231]. SC-FDMA [1785]. scalability [881, 1493, 1635, 2076, 2583, 2771]. Scalable [9, 10, 25, 124, 137, 186, 269, 311, 431, 577, 749, 852, 956, 1015, 1028, 1071, 1085, 1174, 1254, 1472, 1549, 1560, 1604, 1619, 1641, 1697, 1782, 1793, 1828, 1883, 1943, 2167, 2170, 2217, 2288, 2293, 2320, 2394, 2412, 2500, 2569, 2604, 2704, 2821, 2828, 2889]. SCALAR [1071]. scale [135, 399, 437, 493, 502, 549, 583, 613, 623, 656, 683, 809, 883, 1076, 1125, 1197, 1307, 1380, 1388, 1504, 1521, 1769, 1781, 1793, 1840, 1921, 1935, 1970, 1984, 2017, 2026, 2060, 2116, 2234, 2257, 2290, 2315, 2650, 2741, 2761, 2796, 2805, 2955, 2981]. scale-free [1921, 1970]. scales [2700]. Scaling [137, 503, 938, 1456, 1743, 2733]. scam [1123]. SCaN [2846]. SCaN-Mob [2846]. scanners [106]. scanning [234, 430, 871, 1083]. scarce [1871]. scenario [1618, 2373]. scenarios [359, 410, 786, 1070, 1207, 1628, 1826, 2600, 2926, 2968]. schedule [1050]. schedule-based [1050]. scheduled [453, 825, 2038]. scheduler [252, 422, 1378, 1481, 2452, 2860]. schedulers [754]. schedules [2661]. Scheduling [22, 50, 60, 63, 88, 123, 163, 181, 240, 243, 268, 275, 289, 308, 343, 345, 372, 406, 460, 487, 500, 517, 529, 622, 624, 646, 696, 740, 768, 771, 785, 803, 866, 869, 932, 941, 967, 976, 992, 1015, 1028, 1082, 1120, 1222, 1226, 1259, 1273, 1299, 1353, 1384, 1395, 1396, 1524, 1657, 1664, 1670, 1770, 1771, 1803, 1816, 1822, 1884, 1902, 1909, 1918, 1964, 2124, 2125, 2179, 2198, 2205, 2207, 2227, 2271, 2279, 2282, 2286, 2347, 2358, 2423, 2466, 2545, 2593, 2654, 2685, 2697, 2783, 2853, 2877, 2882, 2885, 2901, 2921, 2922, 3003]. Scheme [3, 5, 33, 37, 90, 158, 184, 202, 348, 403, 426, 486, 535, 556, 582, 636, 655, 657, 678, 681, 716, 797, 827, 996, 1017, 1022, 1069, 1087, 1103, 1121, 1215, 1219, 1287, 1293, 1327, 1331, 1367, 1381, 1382, 1385, 1451, 1464, 1511, 1558, 1573, 1581, 1588, 1603, 1608, 1636, 1667, 1675, 1720, 1802, 1812, 1813, 1823, 1840, 1841, 1974, 2040, 2061, 2066, 2099, 2154, 2162, 2164, 2174, 2218, 2249, 2287, 2292, 2296, 2328, 2349, 2358, 2362, 2426, 2437, 2451, 2501, 2523, 2546, 2548, 2578, 2582, 2595, 2604, 2608, 2611, 2616, 2636, 2703, 2707, 2724, 2732, 2759, 2827, 2875, 2919, 2929, 2942, 2970, 2983, 2994, 3002, 3009]. scheme [45, 112, 431, 1646, 1659, 2143, 2322, 2329, 2643, 2826, 2834]. schemes [60, 332, 364, 390, 406, 430, 625, 667, 787, 953, 1138, 1139, 1213, 1305, 1324, 1334, 1335, 1387, 1438, 1595, 1798, 1829, 1901, 1999, 2300, 2517, 3012].

Science [256, 812, 816, 961, 1246]. scientific [2669]. SCMA [2358]. score [2259]. ScoreForCore [1978]. scouting [604]. screaming [2160]. SCTP [11, 109, 1093]. SD [2886]. SD-WISE [2886]. SDHetN [1433]. SDM [2583]. SDMA [785]. SDMA-OFDMA [785]. SDN [1042, 1174, 1314, 1463, 1565, 1577, 1678, 1688, 1690, 1692, 1694, 1696, 1699, 1859, 1925, 1983, 2041, 2076, 2095, 2171, 2186, 2189, 2190, 2195, 2208, 2221, 2249, 2250, 2266, 2275, 2290, 2334, 2337, 2393, 2443, 2449, 2458, 2475, 2502, 2540, 2578, 2589, 2601, 2609, 2655, 2671, 2676, 2677, 2810, 2811, 2821, 2830, 2905, 2917, 2939, 2951, 2954, 3006]. SDN&NFV [2736]. SDN-based [1696, 2290, 2459, 2578, 2655, 2917]. SDN-enabled [2443, 2601, 2810, 2905, 2939]. SDN/NFV [2190]. SDNs [1689, 1709, 2098, 2835]. SDRCS [486]. SDS [2773]. SDWSNs [2531]. SE [1061]. SE-AKA [1061]. SEAHORSE [1494]. seamless [37, 262, 360, 1577, 2038, 2677]. Search [104, 387, 581, 773, 809, 1225, 1250, 1431, 1475, 1494, 1574, 1783, 1848, 2135, 2159, 2202, 2222, 2557, 2665, 2773, 2849]. search-and-routing [2849]. searching [156, 303, 2758]. second [186]. secondary [267, 538, 1038, 1922, 1976, 2215]. secondary-spectrum [538]. secrecy [1993, 2494]. secret [1994]. sector [880, 2674]. sector-based [880]. Secure [5, 119, 159, 168, 197, 266, 306, 341, 367, 382, 385, 556, 567, 571, 655, 695, 728, 894, 972, 1061, 1076, 1102, 1208, 1228, 1263, 1306, 1310, 1381, 1387, 1485, 1560, 1605, 1707, 1814, 1858, 1996, 2107, 2127, 2128, 2188, 2202, 2233, 2328, 2329, 2356, 2398, 2446, 2565, 2572, 2632, 2663, 2750, 2810, 2915, 2950, 2956, 2958, 2971, 3011]. secure-multiparty [1076]. secured [1885, 2778]. Securely [2425]. SecureMDD [1107]. Securing [446, 743, 1470, 1488, 1968]. Security [19, 112, 170, 218, 271, 569, 743, 816, 841, 909, 961, 965, 968, 969, 971, 973, 977, 1041, 1098, 1107, 1233, 1245, 1254, 1378, 1400, 1423, 1555, 1565, 1625, 1633, 1757, 1769, 1856, 1938, 1960, 1973, 1991, 1992, 1999, 2171, 2209, 2220, 2295, 2363, 2418, 2486, 2491, 2506, 2507, 2511, 2553, 2567, 2595, 2618, 2712–2714, 2717, 2769, 2770, 2778, 2857, 2858, 2963, 2966]. security-aware [1633, 2220]. security-centric [1960]. seed [487]. seen [227]. segment [1867, 2106, 2687]. segments [823]. selected [245]. Selecting [594]. selection [26, 91, 162, 222, 264, 326, 457, 571, 579, 633, 635, 801, 902, 956, 1008, 1062, 1064, 1382, 1389, 1410, 1531, 1546, 1550, 1627, 1637, 1639, 1712, 1826, 1894, 1952, 1958, 1980, 1988, 2084, 2092, 2105, 2158, 2164, 2215, 2259, 2268, 2269, 2285, 2296, 2346, 2406, 2407, 2455, 2510, 2709, 2777, 2815, 2855, 2870, 2914, 2918, 2920]. selective [1500, 2029, 2081, 2948, 2997]. Self [82, 157, 283, 345, 384, 434, 544, 701, 926, 1001, 1013, 1060, 1144, 1214, 1233, 1527, 1640, 1646, 1713, 1714, 1725, 1846, 1899, 1915, 2309, 2366, 2446, 2450, 2553]. Self-adaptive [434, 701]. Self-checking [1646]. self-coexistence [2450]. self-configuration [384, 926]. self-configuring [1060]. self-destructing [2553]. self-determination [1725]. self-energy [2446]. Self-healing [544, 1144, 2309]. self-learning [82, 1640, 1899]. self-managed [1013]. self-optimization [1915]. self-optimizing [345]. self-organization [1714]. Self-organized [1001, 1233, 1846]. Self-organizing [157, 283, 1713]. Self-Recoverable [2366]. Self-reliant [1527]. Self-tuned [1214]. selfish [598, 677, 1394, 1800]. selfishness [488]. sellers [2577]. semantic [649, 1188, 2805, 2840]. semantic-aware [2805]. semantic-based [2840]. semantics [850, 2934]. Semi [427, 1064, 2813]. semi-Markov [1064]. semi-stateless [2813]. Semi-supervised [427]. Sender [127]. seniors [2408]. Sense [2767]. sensible [2212]. Sensing [43, 48, 188, 269, 482, 486, 491, 576, 671, 771, 1055, 1106, 1145, 1167, 1335, 1345, 1485, 1541, 1615, 1842, 1926, 1972, 1974, 2048, 2064, 2125, 2127, 2128, 2140, 2369, 2397, 2400, 2408, 2549, 2566, 2596, 2603, 2618, 2664, 2726, 2767, 2824]. sensing-based [2824]. sensitive [222, 253, 893, 1897, 2212, 2395, 2400, 2852]. sensitivity [1453]. Sensor [33, 98, 139, 179, 181, 193, 199, 217, 220, 250, 385, 427, 452, 504, 528, 573, 592, 628, 637, 669, 689, 723, 769, 795, 895, 900, 912, 914, 916, 949, 1044, 1074, 1082, 1119, 1173, 1208, 1256, 1282, 1296, 1297, 1316, 1368, 1389, 1415, 1418, 1437, 1467, 1472, 1648, 1797, 1810, 1831, 1836, 1838, 1899, 1971, 1979, 1987, 2036, 2072, 2146, 2223, 2231, 2236, 2238–2240, 2261, 2265, 2268, 2282, 2292, 2298, 2302, 2323, 2330, 2350, 2365, 2372, 2377, 2436, 2441, 2491, 2509, 2528, 2561, 2579, 2638, 2674, 2692, 2697, 2763, 2814, 2827, 2897, 2919, 2942, 2964, 3011]. sensor [35, 50, 75, 77, 126, 163, 218, 240, 268, 273, 298, 302, 330, 335, 344, 377, 403, 409, 434, 455, 458, 468, 482, 579, 647, 670, 728, 732, 804, 821, 825, 878, 890, 893, 897, 934, 979, 1011, 1022, 1087, 1092, 1124, 1178, 1210, 1224, 1226, 1251, 1255, 1315, 1331, 1334, 1341, 1357, 1476, 1492, 1504, 1553, 1649, 1681, 1788, 1796, 1817, 1845, 1866, 1878, 1889, 1891, 1901, 1904, 1926, 1972, 2026, 2108, 2110, 2147, 2224, 2227, 2366, 2410, 2417, 2455, 2484, 2513, 2563, 2582, 2603, 2606, 2613, 2620, 2653, 2659, 2661, 2668, 2781, 2803, 2843, 2851, 2885, 2899, 2901, 2945, 2948, 2990, 2997]. sensor [261, 393, 431, 486, 540, 789, 1004, 1090, 1117, 1301, 1326, 1388, 1574, 1892, 1894, 1924, 2040, 2078, 2228, 2294, 2465, 2487, 2495, 2525, 2565, 2586, 2607, 2643, 2664, 2694, 2853, 2868, 2886, 2922, 3002]. Sensor-free [1831]. Sensors [1004, 1843, 2353, 2362, 2424, 2456, 2530, 2553, 2564, 2646, 2967]. sensory [2743]. sentiment [1406, 2834]. separate [1339]. Separation [246, 294, 320, 974, 2409]. Sequence [1992, 2792, 2837, 2856]. sequences [72]. Sequential [241, 1106, 2795, 2824]. serial [1472]. series [1728, 2484, 2700, 2709, 2797, 2855]. served [866]. Server [117, 158, 262, 360, 683, 1261, 1392, 1483, 1711, 1736, 1771, 1929, 2252, 2381, 2398]. server-based [1771]. server-centric [1483]. servers [152, 200, 203, 735, 860, 1120, 1246, 1679, 1718, 2183]. Service [9, 44, 58, 116, 227, 228, 244, 264, 266, 301, 328, 330, 486, 495, 525, 574, 586, 628, 648, 677, 700, 751, 765, 782, 783, 853, 858, 873, 951, 971, 1031, 1091, 1193, 1246, 1266, 1283, 1292, 1335, 1369, 1370, 1373, 1393, 1433, 1577, 1587, 1603, 1630, 1702, 1703, 1712, 1741, 1763, 1770, 1953, 1956, 1963, 2027, 2038, 2086, 2207, 2215, 2233, 2284, 2292, 2384, 2415, 2427, 2492, 2500, 2533, 2601, 2633, 2647, 2654, 2679, 2681, 2706, 2738, 2773, 2793, 2805, 2810, 2821, 2855, 2903]. service-aware [2038]. service-chaining [2415]. Service-differentiated [486, 1770]. Service-oriented [628, 1266, 1363, 1370, 2086]. Services [119, 130, 134, 183, 295, 321, 361, 384, 443, 464, 565, 621, 629, 716, 793, 820, 842, 908, 986, 1085, 1164, 1170, 1219, 1275, 1455, 1478, 1510, 1701, 1706, 1718, 1722, 1730, 1743, 1823, 1853, 2007, 2389, 2398, 2489, 2521, 2535, 2646, 2908, 2939, 3008]. SERVitES [2222]. SeRViTR [1190]. Session [295, 531, 555, 582, 648, 1665, 2645]. Session-based [531, 2645]. sessions [688]. set [303, 2224, 2773]. set-stochastic [2773]. sets [299]. setting [1609]. setup [31, 138, 295, 2903]. severe [2100]. SFC [2951]. sFlow [1174]. SGor [1063]. sGSA [785]. Shades [1811]. shaped [182]. shapelet [2709, 2855]. shapers [1583]. shaping [340, 1302]. Shapley [1823, 1864, 2086]. Shapley-value [1864]. share [532]. Shared [165, 493, 534, 564, 758, 1108, 1302, 1736, 2138, 2271, 2812, 3004]. Shared-per-wavelength [165]. shared-protected [758]. Sharing [4, 116, 152, 188, 281, 365, 367, 638, 668, 765, 797, 951, 1077, 1175, 1241, 1501, 1722, 1729, 1802, 1827, 2011, 2062, 2086, 2173, 2461, 2472, 2610]. shelf [2798]. SHMO [2408]. Short [353, 858, 1211, 1545, 2897]. Short-lived [1211]. short-range [353, 2897]. Short-term [1545]. shorter [594]. Shortest [18, 80, 180, 235, 919, 1316, 1881, 2637]. Should [1303]. shuffled [977]. shuffling [1257]. Shutdown [845]. SIC [847]. side [1590, 2554]. side-channel [2554]. sided [842]. Sign [972, 2482]. Signal [52, 77, 915, 1092, 1643, 1913, 2165]. signaling [11, 1499, 2299]. signalling [364]. signals [1643]. signature [5, 593, 1069]. signatures [185]. Signed [2978]. significant [241, 1941]. signs [2100]. similarity [595]. Simple [57, 296, 386, 618, 682, 889, 1217, 1343, 1572, 2465, 2842]. Simplification [422]. Simplifying [2676]. SimplyRep [889]. simulated [725, 1552]. Simulation [10, 52, 442, 477, 478, 1333, 1715, 1824, 2515]. simulations [917]. simulative [913]. simulators [1969]. simultaneous [1975]. Single [120, 634, 997, 1047, 1191, 1446, 1520, 1872, 1963, 2089, 2342, 2608]. single-AF-relay [1872]. single-link [120]. single-operator [1963]. single-packet [2608]. Single-stage [1446]. Single-Switched-Beam [634]. Sink [139, 579, 1256, 2563, 2928]. sinks [1443, 2227, 2643, 2899]. SINR [1481]. SIoT [790]. SIP [141, 364, 376, 735, 2486]. SIP-based [141]. SIR [3010]. SIS [1877]. site [2544]. situ [2767]. Situation [1766, 2048]. Size [244, 456, 555, 895, 1847, 2205, 2332]. Size-oriented [244]. sizes [447]. sizing [190]. sketching [1641]. skewed [620]. skill [1619]. skull [2789]. Skype [292]. SLA [873, 981, 1748]. SLA-compliant [1748]. SLA-controlled [981]. SLAs [1534]. SLDP [2750]. SLDRT [766]. Sleep [517, 521, 711, 793, 1021, 1029, 1047, 1450, 1459, 1822, 2696]. sleep-mode [793]. sleeping [2447, 2818]. SLIC [1640]. slices [2800]. slicing [1025, 2662]. sliding [348]. slope [2571]. slot [1667, 2853]. slotted [32, 1545, 2831]. Slow [401, 532, 560, 1167]. Slow-start [560]. SlowDrop [2760]. Small [98, 553, 1093, 1383, 1441, 1824, 1844, 1919, 2119, 2212, 2371, 2536, 3013]. Smart [478, 507, 628, 687, 729, 875, 901, 909, 933, 990, 1102, 1129, 1146, 1165, 1177, 1178, 1183, 1222, 1248, 1253, 1309, 1312, 1482, 1541, 1567, 1571, 1624, 1636, 1694, 1839, 1842, 1846, 1864, 1865, 1893, 1973, 2032, 2051, 2135, 2142, 2143, 2178, 2315, 2328, 2354, 2387, 2412, 2448, 2456, 2505, 2508, 2515, 2520, 2594, 2595, 2601, 2618, 2647, 2679, 2686, 2699, 2876, 2949, 2950]. smart-cities [2315]. smarter [2456]. smartphone [1345, 1372, 2091, 2132, 2355]. Smartphone-based [2132]. smartphones [2125, 2549]. SmartSantander [1165]. smoothing [194, 2490, 2660]. SmPFT [2871]. SNEAK [174]. snooze [2212]. SNR [1675]. SNVC [2054]. Social [47, 104, 149, 673, 737, 790, 813, 856, 857, 922, 963, 1181, 1221, 1321, 1335, 1345, 1401, 1404, 1406, 1408–1411, 1427, 1509, 1542, 1587, 1629, 1717, 1820, 1864, 1944, 1960, 1986, 2044, 2049, 2050, 2054, 2055, 2075, 2099, 2112, 2114– 2116, 2174, 2217, 2472, 2544, 2600, 2680, 2734, 2749, 2805, 2834, 2871, 2875, 3013]. Social-aware [47, 2055, 2875, 3013]. Social-based [2099]. social-enhanced [2600]. social-like [2805]. socialization [2348]. Socially [259, 673, 784, 1407, 1654]. Socially-aware [259, 673]. SocialVPN [149]. SODA [2966]. SoftAir [1564]. Software [692, 982, 987, 1128, 1161, 1183, 1203, 1229, 1245, 1271, 1276, 1280, 1328, 1402, 1433, 1502, 1515, 1548, 1564, 1680, 1685, 1687, 1695, 1697, 1698, 1700, 1709, 1729, 1789, 1832, 1859, 1862, 1867, 1932, 1956, 1996, 2006, 2033, 2060, 2076, 2096, 2107, 2113, 2134, 2172, 2197, 2200, 2243, 2247, 2248, 2333, 2357, 2402, 2420, 2429, 2458, 2468, 2473, 2498, 2500, 2507, 2556, 2616, 2640, 2644, 2678, 2679, 2686, 2698, 2722, 2725, 2729, 2748, 2750, 2801, 2802, 2825, 2841, 2848, 2857, 2886, 2889, 2914, 2933, 2944, 2965, 2966, 2976, 2982, 2994]. software-based [2725]. Software-Defined [1128, 1161, 1271, 1402, 1502, 1515, 1685, 1695, 1698, 1709, 1729, 1862, 1867, 1956, 2006, 2033, 2076, 2243, 2247, 2402, 2420, 2429, 2473, 2507, 2640, 2644, 2678, 2679, 2686, 2698, 2748, 2841, 2848, 2886, 2914, 2933, 2944, 2966, 2994]. Softwarization [2245, 2657]. soil [2767]. solar [2059, 2606]. Solution [123, 181, 643, 674, 757, 767, 944, 954, 1305, 1547, 1560, 1817, 2387, 2416, 2463]. solutions [1252, 1334, 1416, 1856, 2191, 2449, 2706, 2769, 2973]. solvable [2637]. solve [444]. Solving [631, 2108]. Some [233]. SON [2310]. SOR [1509]. sorted [2124]. SOTE [2825]. Souk [2052]. Sounding [1298]. Source [8, 183, 640, 855, 1020, 2010, 2285, 2533, 2547, 2661, 2703, 2804]. sources [1073, 1457]. south [2253]. Space [461, 587, 2105, 2161, 2248, 2753, 2832]. space-efficient [2248]. Spaces [1505]. spacing [803]. SpaDeS [855]. SPAIS [1646]. spam [1125]. spammers [855]. Spanish [583]. Spanning [445, 618, 796, 1430]. spare [2138]. Sparse [12, 502, 1876, 2053, 2419, 2920]. sparse-splitting [1876]. sparsity [2561]. SPARTA [2184]. Spatial [20, 94, 294, 691, 822, 946, 990, 1315, 1960, 2052, 2123]. spatially [2665]. Spatio [1974, 1987]. Spatio-temporal [1974, 1987]. Special [15, 42, 69, 148, 315, 470, 612, 699, 746, 845, 961, 969, 982, 1034, 1151, 1182, 1363, 1460, 1611, 1946, 1991, 2112, 2352, 2353, 2712, 2996]. Specific [1207, 1277, 1371, 1553, 1558, 1613, 2867]. specification [1188]. specifications [750, 983]. specified [2628]. Specifying [987]. spectra [2445]. spectral [497, 2162, 2264, 2527]. spectrally [2665]. spectrally-spatially [2665]. Spectrum [73, 188, 222, 267, 345, 439, 526, 538, 722, 725, 736, 763, 797, 843, 945, 1008, 1017, 1106, 1129, 1252, 1304, 1369, 1370, 1475, 1485, 1496, 1501, 1550, 1557, 1558, 1586, 1636, 1667, 1807, 1992, 2064, 2127, 2128, 2181, 2194, 2219, 2237, 2461, 2481, 2523, 2560, 2602, 2656, 2673, 2803, 2973]. SpectruM-Aware [1636]. Speculations [816]. SpecWatch [2602]. speed [10, 38, 82, 99, 237, 241, 751, 1072, 1455, 2256, 2644, 2735, 2843, 2955]. speedometers [354]. SPFP [2356]. sphere [607, 1604]. spike [753]. spike-detecting [753]. spilt [2502]. Spine} [1519]. SPIT [928]. splicing [533]. Split [282, 1136, 2300, 2926, 2936]. splitting [350, 1291, 1876]. spoofing [835]. spot [383]. SPR [527]. spraying [2946]. spread [1992]. spread-spectrum [1992]. spreaders [106]. spreading [1970]. Sprint [533]. SpyEye [849]. square [409]. squares [427]. SR [2081]. SRLG [1568]. SRTS [2366]. Stability [83, 90, 120, 253, 515, 1041, 1478, 1831, 2927]. Stabilizing [201]. stable [489, 842, 1798, 2179, 2282]. stack [2009]. stackable [980]. Stackelberg [797, 2360]. stacks [1361]. stadium [2779]. stage [1446, 1642, 1776]. staircase [37]. stall [1567]. stamps [2835]. stand [1125]. stand-alone [1125]. standard [1968]. Standardization [471, 1488]. standardized [1305]. standards [971, 2699]. STAR [2247]. StarLight [1159]. start [401, 532, 560]. startup [390]. starvation [229, 772, 2342]. state [27, 334, 370, 1098, 1328, 2041, 2387, 2471, 2770, 2838, 2841]. state-of-the-art [1098]. stateful [2952]. Stateless [902, 2813]. states [2888]. static [78, 341, 898, 1171]. station [383, 540, 1047, 1124, 1459, 2240, 2599, 2880]. stationary [116]. stations [198, 558, 926, 1450, 2059, 2845]. Statistical [72, 121, 227, 293, 421, 1073, 1936, 2711, 2990]. statistically [1941]. statistics [580, 1010, 2200, 2445, 2602, 2856]. status [1405]. stay [1303]. stealthy [2995]. steering [1319, 2303]. stellar [2085]. step [832, 847, 1127]. step-wise [832]. stepping [174]. stepping-stone [174]. Stochastic [547, 557, 587, 629, 824, 888, 957, 965, 1501, 2098, 2162, 2185, 2242, 2411, 2455, 2773]. stone [174]. Stopping [1123]. Storage [332, 431, 552, 1077, 1118, 1437, 1507, 1629, 1797, 2007, 2089, 2279, 2312, 2329, 2778, 2795]. stored [1213]. stores [2256]. storming [27]. story [2012]. stranger [638]. Strangers [279]. strategic [1506]. Strategies [97, 194, 222, 264, 438, 600, 685, 952, 1181, 1601, 1642, 1650, 1717, 1929, 2083, 2199, 2259, 2308, 2310, 2415]. strategy [412, 801, 1166, 1171, 1469, 1484, 2319, 2407, 2484, 2541, 2545, 2585, 2642, 2695, 2803, 2846, 2872, 2904]. stream [1020, 2752, 2927]. Streaming [8, 93, 101, 109, 174, 223, 246, 346, 446, 447, 464, 488, 505, 557, 616, 659, 718, 779, 787, 791, 860, 861, 889, 957, 1015, 1026, 1126, 1279, 1299, 1327, 1358, 1468, 1498, 1500, 1503, 1513, 1516, 1653, 1682, 1699, 1718, 1731, 1760, 1853, 1870, 1903, 1930, 2011, 2012, 2091, 2109, 2254, 2286, 2365, 2386, 2421, 2489, 2512, 2526, 2574, 2655, 2797, 3014]. streams [1567, 1641, 2101, 2341, 2719, 2733, 2762]. strength [1643]. Strengthening [1604]. stress [1765]. stress-centrality [1765]. stretch [2343]. strip [2603, 2843]. strip-based [2843]. stroke [2789]. Structural [608, 676, 1986, 2113]. structure [108, 511, 661, 1439, 1447, 2152, 2780, 2802]. Structure-Based [1439]. Structured [150, 156, 272, 880, 1786]. structures [2301, 2426, 2586]. Struggling [1343]. studies [1516, 1552]. study [4, 239, 302, 442, 493, 504, 578, 626, 713, 867, 905, 1109, 1244, 1377, 1408, 1438, 1553, 1565, 1715, 1824, 1825, 1957, 1969, 2008, 2160, 2315, 2428, 2505, 2532, 2684, 2738, 2766, 3008, 3010]. Studying [1138]. sub [186, 223]. Sub-Packet [223]. sub-second [186]. Subcarrier [1209, 1667, 2918]. subcarrier-slot [1667]. Subchannel [254, 513, 1068]. subjective [1516, 1744]. sublambda [1291]. subscribe [134, 674, 741, 883, 908, 1078, 1549, 1767, 2304, 2737]. subscription [1116]. subsidy [2461]. substrate [325, 1724, 2062]. substructures [2757]. Successful [2682]. successive [460, 2423, 2831]. Suitability [1054]. suitable [584, 1897]. sum [1375]. summaries [2097]. supervised [427, 2864]. supervisor [2889]. supplier [2207]. supplies [484]. support [58, 88, 130, 249, 376, 557, 656, 694, 751, 1208, 1298, 1544, 1594, 1685, 1692, 1756, 1770, 1778, 2048, 2391, 2521, 2577, 2846]. supported [390, 987, 2780, 2951]. Supporting [550, 799, 887, 1254, 1741, 2278, 2715, 2868]. suppression [2621]. surface [979, 1541]. Surveillance [35, 250, 1258, 2141, 2361, 2448, 2530, 2925]. Survey [43, 44, 66, 70, 74, 97, 168, 193, 199, 205, 225, 248, 357, 373, 438, 467, 468, 490, 507, 569, 601, 728, 729, 760, 813, 846, 907, 909, 959, 1098, 1136, 1230, 1255, 1275, 1284, 1323, 1347, 1399– 1401, 1416, 1436, 1476, 1488, 1502, 1554, 1555, 1587, 1591, 1598, 1599, 1615, 1650, 1656, 1668, 1671, 1684, 1773, 1785, 1786, 1815, 1878, 1901, 1929, 1961, 1981, 1982, 1999, 2036, 2068, 2076, 2096, 2136, 2189, 2199, 2228, 2254, 2295, 2350, 2379, 2384, 2390, 2413, 2511, 2524, 2567, 2609, 2613, 2692, 2706, 2713, 2714, 2746, 2769, 2817, 2852, 2921, 2944, 2965]. survey [142, 393, 787, 1036, 1117, 1685, 2430, 2775, 2956]. Survivability [97, 1797, 2308, 2311]. Survivable [96, 1122, 2420, 2614]. survival [2184]. survive [2496]. SURVIVOR [2810]. Suspicious [2006]. Sustainability [1220, 1792]. sustainable [413, 1650, 1889, 2407, 3002]. SUT [465]. SUTIL [162]. SVC [505, 2655]. SVC-based [2655]. SVDR [2604]. SVM [2922]. SVPS [2508]. Swapping [2516]. Swarm [204, 672, 725, 1143, 1819, 2574, 2826, 2854, 2910, 2943, 2964, 3009]. swarming [1131, 1321, 2771]. swarms [339, 1628]. SWIPT [2494]. SWISH [367]. Switch [186, 343, 772, 1578, 1660, 1787, 1902, 1925, 2266, 2288, 2848, 2880, 2933, 2982]. switch-based [772]. switch-off [2880]. switch-over [186]. Switched [32, 634, 1453]. switches [754, 1642, 2477]. switching [165, 263, 746, 1144, 1238, 1329, 1438, 1462, 1543, 1909, 2626]. switching-off [1462]. Sybil [717, 1761]. symbol [447, 1992]. symbolic [1025]. Synchronization [466, 739, 917, 1271, 1424, 1794, 1913, 2178, 2284, 2294, 2366]. synchronous [2792]. SYNFLOOD [1679]. syntactic [661]. system [9, 47, 55, 79, 86, 101, 150, 166, 178, 257, 270, 370, 447, 527, 534, 562, 615, 642, 684, 697, 872, 874, 889, 1000, 1001, 1030, 1068, 1141, 1199, 1213, 1228, 1250, 1287, 1405, 1494, 1596, 1619, 1634, 1644, 1680, 1791, 1835, 1842, 1846, 2020, 2024, 2189, 2250, 2305, 2358, 2408, 2432, 2456, 2478, 2486, 2508, 2511, 2530, 2532, 2601, 2605, 2618, 2622, 2632, 2635, 2645, 2681, 2684, 2734, 2748, 2753, 2780, 2789, 2809, 2830, 2865, 2917, 2949]. Systematic [311, 1083, 2524, 2623, 2710, 2746]. Systematically [1693]. Systems [4, 72, 122, 148, 203, 214, 240, 257, 397, 472, 473, 478, 501, 545, 638, 657, 665, 674, 717, 763, 771, 777, 802, 819, 953, 990, 1057, 1118, 1131, 1132, 1140, 1153, 1218, 1244, 1252, 1298, 1360, 1394, 1507, 1549, 1564, 1612, 1615, 1677, 1709, 1714, 1751, 1772, 1779, 1781, 1794, 1797, 1847, 1976, 1982, 2032, 2043, 2055, 2087, 2089, 2140, 2141, 2236, 2242, 2296, 2345, 2347, 2349, 2365, 2412, 2434, 2459, 2462, 2470, 2505, 2506, 2530, 2542, 2548, 2566, 2598, 2621, 2663, 2689, 2702, 2737, 2741, 2743, 2758, 2761, 2808, 2815, 2870, 2898, 2900, 2909, 2971, 2972, 2996, 3005]. systems [83, 141, 154, 462, 535, 866, 885, 931, 941, 982, 1426, 1438, 1484, 1543, 1555, 1588, 1861, 2059, 2124, 2125, 2454, 2512, 2619, 2771, 2928, 2930, 2985].

T [3014]. table

[297, 1206, 1678, 1695, 2171, 2197, 2247, 2477]. Tables [632, 1752, 1919]. Tabu [1475]. Tackling [2087]. tag [115, 2758, 2985]. tags [481, 556, 1857, 2202]. TAISC [1959]. take [1405]. Taking [1815]. Talent [604]. Taming [401]. tap [1994]. Target [1224, 2296, 2559, 2668, 2837]. targeted [1466, 2792]. targeting [2206]. targets [352, 2005, 2925]. Task [1187, 1338, 2125, 2271, 2768, 2844, 2898, 2922]. tasks [2279, 2400, 2566, 2692]. tat [600, 2526]. Taxonomy [71, 1136, 1717, 1785, 2379, 3012]. TCAM [748, 2134, 2153, 2248, 2980]. TCAM-based [748, 2153, 2980]. TCAM-limited [2134]. TCP [105, 116, 176, 383, 401, 463, 530, 532, 565, 568, 587, 626, 666, 724, 752, 838, 840, 1059, 1145, 1147, 1198, 1267, 1679, 1708, 1829, 1871, 1975, 1985, 2012, 2067, 2094, 2101, 2130, 2234, 2332, 2502, 2672, 2765, 2792, 2968, 2988]. TCP-friendliness [666]. TCP-friendly [587]. TCP-targeted [2792]. TCP/IP [1145]. TCPs [176]. TD [2932]. TDM [925]. TDM-PON [925]. TDMA [976, 1168, 1887, 2661]. TDMA-based [1168, 1887]. technical [538, 1619, 2142]. technical-skill [1619]. technique [489, 633, 766, 953, 1053, 1111, 1238, 1544, 1851, 1992, 2378, 2673, 2756, 2795, 2864, 2870, 2871, 2960]. Techniques [32, 200, 496, 573, 907, 1025, 1117, 1230, 1598, 1626, 1684, 1869, 1967, 2036, 2154, 2406, 2995]. Techno [196]. Techno-economic [196]. technological [1723]. technologies [7, 84, 362, 1188, 1312, 1591, 1834, 1916, 2359, 2383, 2387, 2629, 2706, 2817, 2941, 3005]. technology [286, 471, 511, 1627, 2648]. technostress [2115]. TEFIS [1191]. telecom [1701]. telecommunications [1722]. telecoms [700]. teleconference [1618]. telehaptic [3004]. telephone [12]. telephony [11, 12]. temperature [1842]. Temporal [345, 1588, 1974, 1987, 2046, 2070, 2336, 2490, 2624, 2660, 2932]. temporal-difference [2932]. temporal-spectrum [345]. tenancy [2986]. tenant [1268, 1851, 2120, 2469, 2812]. term [313, 1481, 1545, 1770]. terminal [577, 599, 767, 1094]. terminals [866]. termination [161, 2299]. Terrestrial [559]. tessellations [911]. test [1058, 1107, 1191, 1676]. Testbed [458, 742, 1042, 1110, 1152, 1154–1156, 1160, 1164, 1165, 1190, 1195, 1197, 1467, 2163, 2747]. testbeds [1151, 1159, 1182, 1186, 1188, 1192, 1194, 2449]. Testing [85, 471, 475, 1025, 1286, 1570, 1619, 1693]. Tethering [1505]. text [155]. TGBA [2726]. Thank [1413, 1746]. Thanks [2352, 2472]. their [339, 402, 1261, 1893, 2588, 2953]. theoretic [259, 329, 412, 464, 836, 940, 1048, 1317, 2215, 2373]. Theoretical [526, 913, 1590, 1776, 2367, 2518, 2857]. theory [189, 248, 740, 1052, 1375, 1584, 1855, 2113, 2137, 2193, 2631, 2732, 2858, 2875]. Things [205, 790, 862, 973, 1207, 1215, 1254, 1338, 1423, 1625, 1834, 1835, 1839, 1840, 1843, 1844, 1849, 1856, 2074, 2295, 2361, 2363, 2364, 2444, 2492, 2524, 2567, 2594, 2621, 2629, 2646, 2666, 2710, 2713, 2714, 2738, 2753, 2775, 2780, 2791, 2805, 2817, 2823, 2881, 2891, 2909, 2931, 2934, 2950, 2973]. Things-based [2361, 2738]. Things/CubeSats [2753]. thinking [2021]. Third [849, 1634, 2445]. Third-order [2445]. third-party [1634]. Thorough [2163]. threat [2003, 2433]. threats [1041, 1893, 2020, 2746]. three [197, 842, 1102, 1664, 1838, 2467, 2564, 2749]. three-dimensional [2564]. three-factor [1838]. three-layered [2467]. three-party [197, 1102, 2749]. three-sided [842]. Threshold [387, 1260, 1299, 1674, 1877, 2765]. thresholds [520, 2697]. Throughput [46, 60, 296, 383, 503, 772, 871, 872, 895, 990, 1018, 1046, 1095, 1142, 1205, 1417, 1672, 1735, 1922, 1985, 1993, 2019, 2094, 2234, 2255, 2262, 2443, 2493, 2525, 2590, 2615, 2902]. throughput-guaranteed [1018]. Throughput-optimal [1095]. throughput-overhead [46]. throwboxes [1629]. thru [941]. tick [466]. Ticket [1344, 2356]. Ticket-based [1344, 2356]. tier [408, 874, 1017, 1416, 1420, 2068, 2210, 2215, 2219, 2311, 2578, 2865]. tiered [1091, 1185]. ties [2834]. Tight [1661]. tightly [2303]. Time [50, 62, 83, 86, 109, 186, 188, 269, 289, 329, 407, 414, 466, 486, 502, 510, 528, 557, 566, 594, 629, 724, 853, 874, 893, 899, 916, 918, 927, 976, 1019, 1085, 1251, 1510, 1588, 1598, 1600, 1669, 1677, 1682, 1718, 1728, 1791, 1886, 1890, 1893, 1899, 1919, 1939, 1943, 1952, 2064, 2093, 2100, 2156, 2237, 2246, 2269, 2294, 2366, 2378, 2395, 2484, 2521, 2581, 2624, 2627, 2642, 2644, 2650, 2700, 2709, 2719, 2738, 2758, 2761, 2767, 2779, 2797, 2803, 2835, 2853, 2855, 2873, 2877, 2902, 2927, 2951, 2976, 2987]. Time-activity [1939]. time-aware [1600]. time-based [1598]. time-driven [528]. Time-efficient [2064, 2761]. time-frequency [502]. time-limited [2642]. Time-optimized [1510]. time-related [2627]. time-sensitive [893, 2395]. time-validity-constrained [2624]. time-varying [629, 899]. Timely [2286]. timeout [298]. times [116, 672, 1118, 1771, 2515, 2991]. timescale [1652]. timestamping [1211, 1248]. timestamps [1597]. timing [340, 448]. tit [600, 2526]. Tit-for-Tat [600]. tit-for-tat-based [2526]. Titan [848]. TMA [994]. TOD [1352]. TOD-MAC [1352]. today [1035]. token [1583, 2519]. token-buckets [2519]. tolerance [428, 1570, 2134, 2933]. Tolerant [145, 379, 418, 494, 867, 888, 1130, 1134, 1294, 1346, 1524, 1552, 1580, 1950, 2030, 2045, 2078, 2083, 2239, 2316, 2384, 2923]. Tolerating [1117, 1198]. tomography [2528, 2789]. tool [76, 478, 786, 987, 1193, 1445, 1471]. tool-supported [987]. Tools [477, 596, 603, 1194, 1419, 1616, 1676]. Top [106, 1255, 2567, 2580]. top-down [1255, 2567]. topic [1767]. topic-based [1767]. topological [158]. topologies [77, 102, 125, 413, 618, 1508, 2859, 2964]. Topology [21, 91, 111, 238, 598, 660, 833, 890, 934, 978, 1117, 1300, 1324, 1439, 1462, 1464, 1465, 1493, 1521, 1543, 1585, 1806, 1828, 1921, 2030, 2073, 2198, 2232, 2314, 2377, 2404, 2540, 2573, 2603, 2867, 2945, 3015]. Topology-aware [238]. Topology-Preserving [2540]. topology-transparent [2198]. TOPSIS [1958]. TOPSIS-based [1958]. Tor [879, 2554]. TorrentGuard [1123]. total [1180, 1549, 2877]. tours [1100]. TP/PWE3 [756]. TPD [1716]. traceability [1005, 2143]. traceback [1324, 2538, 2608]. Traceband [76]. traces [336, 578, 1471]. tracing [650, 1331]. trackability [1422]. TrackerDetector [1634]. trackers [1634]. Tracking [202, 548, 950, 1218, 1258, 2152, 2424, 2559, 2948, 2997, 3015]. Trade [534, 548, 713, 949, 1220, 1637, 1715, 1887, 2105, 2416, 2907]. Trade-off [548, 713, 949, 1637, 1887, 2105, 2416, 2907]. trade-offs [1220, 1715]. Tradeoff [1426, 1438, 1578, 2212, 2487]. tradeoffs [1294, 2220, 2473]. Trading [245, 534, 538, 758, 1393, 1558, 2032, 2601, 2810]. traditional [847]. Traffic [82, 129, 167, 195, 242, 253, 269, 304, 331, 336, 347, 350, 351, 383, 389, 396, 402, 413, 502, 513, 539, 554, 557, 587, 590, 671, 711, 719, 733, 766, 770, 831, 840, 851, 875, 919, 923, 927, 954, 956, 976, 1006, 1073, 1147, 1231, 1232, 1280, 1302, 1314, 1336, 1347, 1350, 1419, 1435, 1440, 1441, 1475, 1562, 1566, 1603, 1640, 1655, 1661, 1664, 1697, 1716, 1772, 1793, 1820, 1821, 1902, 1927, 1932, 1935, 1939, 1975, 2002, 2004, 2006, 2034, 2106, 2151, 2164, 2272, 2275, 2287, 2339, 2345, 2479, 2489, 2536, 2540, 2584, 2596, 2622, 2627, 2665, 2681, 2865, 2873, 3007]. traffic [16, 235, 299, 309, 310, 333, 342, 436, 478, 546, 550, 581, 583, 613, 616, 646, 650, 652, 661, 676, 707, 835, 886, 980, 1088, 1288, 1291, 1300, 1319, 1339, 1429, 1438, 1464, 1493, 1542, 1572, 1662, 1668, 1669, 1676, 1791, 1897, 1943, 2003, 2024, 2039, 2070, 2101, 2158, 2241, 2253, 2303, 2406, 2445, 2453, 2466, 2482, 2490, 2586, 2660, 2673, 2765, 2820, 2825, 2844, 2845, 2927, 2952]. traffic-aware [707, 1464, 2466]. traffic-efficient [269]. traffic-intensive [1655]. Traffic-prediction-assisted [554]. trail [57]. trails [1568]. trajectories [1216, 2598]. trajectory [994, 1443, 2955]. Trajectory-based [994]. Trans [1431]. Trans-Algorithmic [1431]. transactions [54, 2720]. Transfer [115, 326, 666, 1297, 1357, 1492, 1796, 1900, 2230, 2368, 2794]. transfers [297, 1273, 2083, 2921, 2987]. transformation [2085]. transformations [2470]. transit [885, 886, 1339, 2098]. transit-edge [886, 1339]. transition [37, 2838]. translucent [707]. Transmission [109, 219, 240, 448, 515, 782, 1003, 1141, 1352, 1436, 1681, 1782, 1812, 1871, 1894, 2040, 2048, 2066, 2081, 2126, 2213, 2239, 2305, 2348, 2446, 2451, 2471, 2484, 2495, 2500, 2501, 2625, 2628, 2666, 2682, 2799, 2818, 2844, 2882, 2940, 2960, 2979]. transmissions [599, 680, 1593, 1981, 2079, 2447]. transmit [622]. transmit/receive [622]. Transparent [180, 358, 516, 577, 992, 1144, 1883, 2149, 2198]. transparent-relay [516]. transponders [756]. Transport [38, 129, 201, 515, 943, 1480, 1523, 1720, 1897, 1956, 2016, 2067, 2138, 2593, 2628, 2946]. transportation [2515, 2530, 2566, 2598, 2601, 2663, 2734, 2748, 2955, 2996]. travel [1716, 2440]. traveling [977]. Traversal [158, 2174]. Tree [143, 204, 388, 445, 489, 553, 618, 1101, 1173, 1222, 1235, 1293, 1316, 1341, 1430, 1443, 1528, 1644, 1654, 1866, 1882, 2072, 2203, 2418, 2470, 3006]. Tree-based [204, 1222, 1293, 1341, 1528, 1644, 1866, 2470]. trees [80, 796, 1537, 1608, 1904, 2901]. trend [473]. trends [2383, 2657]. trial [2596]. triangles [386]. tribrid [2801]. Tribute [1237]. Trie [2241]. tries [2167]. triggered [2447]. triggering [438]. TRILL [1268].

TRILL-based [1268]. TRIM [358]. TRIMS [214]. triple [1219, 1455]. triple-play [1219, 1455]. TRM [1887]. TRM-MAC [1887]. trolls [737]. troubleshooting [1934, 1937]. Trust [214, 426, 563, 737, 1063, 1326, 1327, 1348, 1423, 1615, 1761, 1966, 2055, 2174, 2193, 2305, 2491, 2531, 2663, 2766]. Trust-Aware [426, 1966]. Trust-based [1327, 1348, 2055, 2491, 2663]. trusted [1211, 1511, 2718]. Trustworthy [1105, 1190, 2634]. truth [336, 2719]. Truthful [2400, 2454, 2549]. TSCH [1424, 2466]. TTL [1234]. TTL-based [1234]. Tumbler [1910]. tunable [756]. tuned [1214]. Tuning [11, 167]. tunnel [594, 1895]. tunnel-based [594]. tunneling [770]. tunnels [1239, 2928]. tutorial [1288, 3008]. TV [346, 351, 397, 1285, 1505]. TVWS [2560]. TWDM [2031]. Two [224, 235, 254, 364, 378, 463, 520, 1127, 1133, 1142, 1334, 1420, 1596, 1776, 1782, 2034, 2061, 2210, 2215, 2219, 2224, 2227, 2311, 2398, 2469, 2568, 2571, 2587, 2726, 2865]. two-dimensional [1782]. two-factor [1334]. two-hop [254, 1142, 2224]. two-level [2469]. Two-link [224]. two-phase [235, 2061, 2726]. two-relay [1596]. two-server [2398]. Two-slope [2571]. two-stage [1776]. two-step [1127]. Two-tier [1420, 2210, 2215, 2219, 2311, 2865]. two-user [2568]. Two-way [378, 463, 2034, 2587]. Tx/Rx [1803]. TXOP [433]. Type [63, 518, 1379, 1826, 2147, 2357]. typosquatting [465].

U [1604, 2523]. U-Sphere [1604]. UAV [2236, 2483, 2496, 2611, 2668, 2872]. UAV-based [2668]. UAV-to-car [2483]. ubiquitous [210, 1614, 2508, 2753, 2850]. ubiquity [1404]. UDP [1065, 1975]. ultra [122, 289, 364, 2119, 2175, 2315, 2368]. ultra-dense [2119, 2175, 2368]. ultra-low [289]. ultrasound [2545]. UMTS [170, 359, 362, 444]. un-traceability [1005]. unavailability [1480, 2557]. uncertain [1927, 2228, 2640]. uncertainties [252]. uncertainty [16, 1084, 1865, 2032, 2918]. uncooperative [2105]. Uncoordinated [2970]. Uncovering [1346]. under-ice [1084]. underground [2767]. underlay [1126, 1872, 2587, 2739]. underlay-aware [1126]. Understanding [408, 610, 849, 968, 986, 1630, 2213]. Underwater [33, 1007, 1084, 2110, 2261, 2564, 2791, 2885, 2897]. unequal [718]. unfairness [192]. unicast [394, 436, 742, 1475, 1523, 1668, 1951, 2665]. unicasting [2402]. Unified [46, 318, 743, 1385]. union [80]. units [1494, 1590]. Universal [559]. university [1197]. unknown [410, 2602]. unlicensed [222, 2219, 2346, 2497]. unlinkable [2720]. Unmanned [3002]. unpredictability [867]. unreliable [2874]. unresponsive [1508]. unstructured [303, 551, 801, 1132, 1166]. untraceability [2707]. untrusted [1250, 2173]. unveiling [241]. unwanted [733]. Up-to-date [2058]. Up/Down [6]. Update [519, 801, 832, 1694, 1919, 2249, 2357, 2477]. update-based [832]. updates [678, 2677, 2835]. updating [410]. upgradable [911]. Upgrading [394]. Uplink [516, 535, 1287, 1603, 1785, 2165, 2225, 2264, 2303, 2423, 2683]. upload [154, 931, 2575]. uploading [724, 2055]. UPS [2626]. Uranus [2800]. Urban [43, 694, 791, 1055, 1398, 1839, 1841, 2048, 2397, 2440, 2577, 2949]. URL [465, 811]. URLLC [2936]. Usable [541, 2830]. usage [493, 1408, 1452, 1695]. USB [2554]. Use [1201, 1500, 1540, 2242, 2440, 2669, 2941]. useful [1781]. User [4, 104, 117, 169, 210, 281, 374, 488, 514, 578, 654, 658, 774, 816, 866, 1010, 1037, 1044, 1087, 1108, 1137, 1188, 1208, 1302, 1310, 1360, 1364, 1368, 1478, 1510, 1534, 1676, 1872, 1922, 1988, 1989, 2075, 2092, 2097, 2173, 2210, 2211, 2215, 2280, 2340, 2472, 2490, 2497, 2523, 2544, 2568, 2636, 2724, 2727, 2784, 2842, 2907, 2970, 3013]. user-activity [2497]. User-assisted [1037, 1478]. User-aware [2340]. user-centric [1310, 1364, 1368]. user-driven [2092]. User-level [374]. user-mobility [2497]. user-oriented [210]. user-participatory [2636]. user-perceived [1302]. User-Provided [281]. users [259, 922, 1106, 1481, 1800, 1912, 1922, 1957, 2207, 2411]. Using [34, 39, 85, 104, 108, 135, 216, 220, 307, 329, 364, 371, 381, 420, 458, 520, 544, 581, 595, 626, 633, 634, 754, 795, 840, 851, 911, 918, 926, 1007, 1025, 1056, 1100, 1115, 1183, 1225, 1235, 1324, 1332, 1370, 1421, 1431, 1505, 1536, 1568, 1598, 1626, 1643, 1649, 1727, 1755, 1760, 1783, 1797, 1821, 1857, 1904, 1971, 1985, 2008, 2028, 2049, 2095, 2100, 2113, 2137, 2151, 2178, 2194, 2224, 2237, 2296, 2305, 2321, 2347, 2354, 2360, 2369, 2401, 2415, 2433, 2456, 2510, 2530, 2536, 2638, 2642, 2646, 2654, 2777, 2795, 2854, 2857, 2875, 2885, 2888, 2901, 3007, 3008, 3014]. using [52, 102, 282, 447, 515, 553, 792, 821, 827, 871, 977, 1032, 1102, 1129, 1188, 1202, 1209, 1214, 1276, 1509, 1605, 1654, 1682, 1778, 1779, 1782, 1819, 1835, 1839, 1882, 1903, 1917, 2086, 2154, 2233, 2329, 2439, 2449, 2562, 2605, 2767, 2778, 2787, 2835, 2838, 2927, 2988]. uth [2984]. utilisation [2161]. Utility [162, 440, 629, 696, 1496, 2263, 2809, 2868]. Utility-based [629, 696]. Utility-driven [1496]. utilization [776, 2266, 2523, 2702, 2795]. Utilizing [21, 439, 1926, 2219]. UTRAN [301]. UWSNs [2935].

V [472]. V-to-X [472]. V. [8]. V2I [474]. V2V [1258, 2222, 2542]. V2X [471, 473, 478, 2476, 2494, 2769]. V2X-SWIPT [2494]. validated [2605]. validation [1025, 1164, 1528, 1858, 3010]. validity [614, 1538, 2624]. value [1823, 1827, 1864, 2086, 2271]. values [2866]. VANET [2099, 2305, 2437, 2536, 2641]. VANET-LTE [2536, 2641]. VANETs [695, 791, 1212, 1886, 2193, 2233, 2313, 2508, 2555, 2642, 2708, 2759, 2782]. Variable [447, 630, 1027, 1992, 2198, 3003]. Variable-weight [2198]. variance [923, 1271]. variance-based [1271]. variants [568]. variational [2957]. varied [2968]. Variegated [1394]. various [128, 1829]. varying [629, 899]. VASE [835]. VBR [1664]. VCR [642]. VDC [1272]. VDC-Analyst [1272]. VDR [630]. vector [420, 627, 1778, 2270, 2339]. vectorizing [2735]. Vehicle [470, 471, 475, 479, 1980, 2068, 2147, 2292, 2305, 2508, 2530, 2600, 2601, 2648, 2810, 3002]. vehicle-2-x [470, 475, 479]. vehicle-as-a-service [2601]. vehicle-assisted [3002]. Vehicle-to-Everything [2600]. vehicle-to-grid [2810]. Vehicle-to-infrastructure [2068]. vehicles [1084, 2476, 2497, 2515, 2595, 2622, 2680, 2759, 2793]. Vehicular [43, 285, 425, 429, 477, 522, 577, 805, 942, 994, 1043, 1258, 1398, 1477, 1500, 1716, 1841, 1964, 1980, 2034, 2054, 2193, 2222, 2292, 2324, 2379, 2390, 2438, 2464, 2468, 2482, 2491, 2577, 2641, 2644, 2647, 2787, 2799, 2801, 2892, 2914, 2983]. verbose [421]. Verification [1272, 1740, 2041, 2279]. VerSAMI [2412]. Versatile [2412]. versus [1453, 1523]. vertex [1816, 2774]. Vertical [58, 142, 262, 362, 930, 1065, 1920]. VHub [1446]. via [100, 194, 338, 348, 350, 461, 593, 619, 661, 676, 725, 848, 1058, 1304, 1690, 1702, 1848, 1855, 2064, 2075, 2133, 2265, 2266, 2277, 2461, 2637, 2671, 2728, 2757, 2955]. Viability [2458]. vicinity [1346]. Video [8, 143, 154, 223, 292, 403, 447, 464, 505, 579, 616, 642, 718, 779, 787, 793, 860, 861, 887, 935, 1015, 1026, 1028, 1150, 1199, 1358, 1380, 1478, 1500, 1503, 1513, 1567, 1630, 1653, 1682, 1698, 1699, 1711, 1760, 1853, 1870, 1903, 1930, 2012, 2066, 2081, 2091, 2164, 2448, 2544, 2574, 2580, 2704, 3014]. video-on-demand [1380, 1478, 1653, 1930]. video-streaming [8, 464, 3014]. videos [1782, 2280, 2545]. view [154, 408, 950]. View-Upload [154]. viewer [1630]. Viewing [2109]. viewpoints [2390]. village [284]. violence [2779]. viral [1877]. Virtual [151, 325, 366, 373, 430, 600, 660, 714, 731, 814, 831, 887, 898, 996, 1009, 1079, 1091, 1171, 1175, 1195, 1202, 1238, 1249, 1257, 1272, 1364, 1405, 1446, 1589, 1607, 1633, 1655, 1692, 1705, 1737, 1738, 1740, 1742, 1768, 1812, 1832, 1904, 1905, 1927, 1928, 2062, 2071, 2073, 2077, 2080, 2168, 2188, 2251, 2253, 2334, 2415, 2449, 2462, 2527, 2576, 2592, 2604, 2659, 2788, 2798, 2800, 2913, 2932, 2952, 3008]. virtual-queue [1079]. Virtualization [66, 211, 282, 955, 1709, 1789, 1825, 1851, 1885, 2252, 2387, 2430, 2438, 2507, 2589, 2657, 2735, 2736, 2770, 2776, 2822, 2915]. Virtualization-based [2438]. Virtualized [408, 1203, 1257, 1701, 1739, 1928, 2111, 2246, 2251, 2381]. VirtualKnotter [1257]. virus [839]. ViSiBiD [2100]. VISION [621]. visual [1315, 1621, 2141]. visualization [208, 607]. visualizing [2418]. vital [2100, 2520]. VLC [2135]. VM [1493, 1600]. VMPlanner [831]. VNE [2640, 2932]. VNE-TD [2932]. VNF [2883]. VNFs [2869]. VoD [4, 615, 642, 905, 931, 957, 1031, 1140, 1199, 2087]. Voice [588, 630, 643]. voice/audio [630]. void [2261]. void-handling [2261]. VoIP [22, 141, 226, 376, 513, 667, 922, 1310]. VoIP-over-WLAN [376]. Volume [135, 927, 1237, 1471, 2127, 2997]. volumes [2003]. Voronoi [2885]. VPN [2627]. VPNs [884, 2903]. vs [346, 491, 659, 996, 1143, 1364, 1437, 1464, 1618]. VSDN [1698]. VSimRTI [478]. vulnerabilities [141]. Vulnerability [115].

W [1709]. W-SDNs [1709]. wait [2218]. waiting [2515]. Wake [599, 1899, 2818]. wake-up [1899, 2818]. wakeup [289, 2948, 2997]. Walk [2949]. Walker [303]. walks [61, 1875, 2046]. WAN [1253]. WANs [2921]. war [172]. warning [425, 2697]. Watchdog [856]. watching [4]. Water [2756, 2791]. waterflow [136]. watermarking [650]. watershed [35]. WAVE [1953, 2067]. WAVE/IEEE [1953]. wavelength [56, 138, 165, 238, 693, 706, 1329]. wavelength-routed [138, 706]. waves [1239]. way [378, 463, 1882, 1905, 2034, 2587, 2949]. WBAN [2667]. WBANs [2707, 2774, 2905]. WBSS [436]. WBSS-based [436]. WDA [328]. WDM [57, 138, 224, 309, 438, 704, 721, 758, 904, 980, 1122, 1176, 1291, 1425, 1545, 1572, 2073, 2599, 2614, 2635]. weak [302]. wearable [2187, 2362, 2646, 2724, 2777]. weather [2912]. Weaver [1616]. Web [104, 108, 152, 328, 387, 578, 799, 807– 809, 812, 815, 816, 971, 1014, 1066, 1188, 1250, 1392, 1679, 1712, 1840, 1923, 2183, 2259, 2267, 2730, 2992]. WebMon [2499]. webpage [2499]. Weight [2198, 2559, 3006]. weighted [308, 1169, 1574, 1643, 2400, 2501, 2800, 2802, 2877, 2978]. weights [260, 1135]. Welcome [2352]. welfare [1335, 1345]. west [2765]. WFQ [308]. wheel [2826]. wheels [2577]. Where [286]. while [1534]. White [1505]. Whom [1410]. Wi [454, 738, 1465, 1630, 1643, 2145, 2303, 2441, 2442, 2503, 2584, 2957]. Wi-Fi [454, 738, 1465, 1630, 1643, 2145, 2303, 2441, 2442, 2503, 2584, 2957]. Wi-Fi/cellular [2442]. Wide [122, 149, 201, 311, 344, 2063]. wide-area [149, 201, 2063]. WiFi [367, 643, 742, 1081, 1598, 1609, 1908, 2327, 2941, 2969]. wiki [150]. Wikipedia [1, 67]. WiMAX [88, 122, 123, 196, 276, 345, 426, 505, 510–512, 518, 520, 522, 535, 657, 719, 785, 1080, 1544]. WiMax/LTE [1080]. Windfall [1181]. window [120, 348, 659, 895]. WiNeMO [1956]. wire [1994]. wire-tap [1994]. wired [1436, 2413]. WIreless [15, 24, 58, 94, 113, 140, 192, 199, 248, 263, 282, 284, 287, 369, 412, 424, 427, 482, 504, 517, 557, 573, 584, 592, 601, 622, 644, 645, 669, 678, 826, 828, 832, 869, 882, 900, 903, 916, 918, 940, 949, 967, 976, 990, 1000, 1048, 1064, 1142, 1168, 1297, 1316, 1323, 1343, 1437, 1467, 1495, 1505, 1564, 1599, 1709, 1729, 1794, 1797, 1806, 1815, 1827, 1838, 1911, 1918, 1924, 1971, 1987, 2146, 2231, 2240, 2282, 2312, 2323, 2350, 2353, 2377, 2405, 2416, 2436, 2480, 2509, 2529, 2592, 2701, 2757, 2763, 2846, 2852, 2886, 2919, 2942, 2961]. wireless [49, 60, 98, 110, 139, 164, 179, 181, 217, 220, 229, 250, 264, 294, 341, 361, 366, 385, 407, 409, 440, 458, 466, 468, 521, 523, 550, 571, 597, 633, 637, 689, 712, 723, 767, 769, 770, 795, 803, 804, 825, 895, 897, 914, 948, 1074, 1096, 1108, 1149, 1173, 1200, 1205, 1220, 1245, 1248, 1256, 1296, 1368, 1415, 1418, 1454, 1557, 1573, 1601, 1627, 1658, 1750, 1754, 1810, 1836, 1884, 1888, 1899, 1912, 1946, 2036, 2176, 2238, 2261, 2265, 2298, 2302, 2308, 2321, 2330, 2372, 2413, 2441, 2447, 2456, 2606, 2613, 2631, 2674, 2692, 2814, 2890, 2902, 2945, 2964, 3003]. wireless [3, 21, 35, 65, 75, 130, 163, 172, 197, 218, 219, 230, 268, 280, 290, 298, 302, 377, 405, 434, 439, 460, 531, 545, 554, 561, 626, 670, 696, 732, 798, 890, 901, 920, 998, 1051, 1079, 1097, 1210, 1224, 1226, 1251, 1293, 1307, 1315, 1319, 1331, 1352, 1370, 1377, 1476, 1492, 1544, 1553, 1639, 1649, 1657, 1681, 1683, 1696, 1726, 1728, 1782, 1796, 1830, 1845, 1866, 1889, 1891, 1901, 1904, 1926, 1972, 1993, 1998, 2030, 2053, 2108, 2224, 2269, 2291, 2301, 2319, 2355, 2362, 2407, 2417, 2446, 2461, 2539, 2545, 2546, 2553, 2582, 2588, 2603, 2659, 2661, 2794, 2899, 2929, 2948, 2972, 2997, 3010]. wireless [16, 18, 26, 42, 52, 64, 126, 137, 143, 171, 223, 249, 310, 335, 344, 362, 403, 404, 411, 437, 455, 484, 486, 503, 515, 620, 646, 647, 653, 662, 671, 728, 743, 789, 792, 819, 821, 837, 878, 893, 934, 946, 1022, 1046, 1062, 1090, 1092, 1094, 1104, 1113, 1124, 1214, 1255, 1317, 1334, 1341, 1344, 1357, 1378, 1388, 1504, 1512, 1531, 1560, 1585, 1596, 1652, 1668, 1722, 1780, 1785, 1803, 1816, 1878, 1905, 1995, 2019, 2026, 2078, 2227, 2228, 2262, 2311, 2317, 2326, 2370, 2423, 2513, 2550, 2563, 2620, 2653, 2668, 2711, 2767, 2781, 2803, 2843, 2853, 2922, 3002]. wireless [17, 20, 22, 25, 45, 142, 182, 193, 196, 378, 393, 406, 442, 452, 540, 598, 628, 740, 898, 1004, 1028, 1082, 1117, 1282, 1326, 1387, 1389, 1472, 1574, 1581, 1632, 1646, 1731, 1802, 1892, 1894, 2040, 2086, 2150, 2230, 2236, 2316, 2322, 2561, 2579, 2619, 2638, 2643, 2697, 2727, 2804, 2827, 3011]. wireless-optical [404]. wireless-powered [1888, 2902]. wireless-sensor [2945]. wireless/optical [903]. wiretapping [1814]. Wisdom [1620]. wise [832, 1926, 2886]. within [1171, 1620]. without [27, 236, 260, 294, 347, 635, 1233, 1913, 2439, 2565, 2718]. WLAN [19, 29, 376, 383, 635, 1244, 1382, 1957, 2516]. WLANs [167, 221, 430, 433, 685, 713, 718, 836, 899, 1072, 1141, 1452, 1474, 1675, 2284, 2790, 2809]. WMNs [1089, 2342]. WOBAN [404, 1266]. Wolves [2964]. Woodpecker [2671]. work [2290]. work-conservation [2290]. workers [2726]. workflow [243, 462, 1977]. workflows [2071, 2669, 2695]. Workload [1, 67, 786, 1009, 1561, 2007]. world [98, 1383, 1844, 2195, 2297, 2753]. worlds [814]. worm [1372]. wormhole [2757]. worms [234, 593]. WRAN [2450]. writeable [1857]. WSN [258, 1361, 1395, 1443, 1848, 2396, 2663, 2766, 2827]. WSN-assisted [2766]. WSN-based [2663]. WSNs [168, 289, 352, 490, 599, 649, 1180, 1222, 1584, 1804, 1887, 2271, 2343, 2447, 2451, 2460, 2639, 2756, 2824]. WWAN [29]. Wyner [1994].

X [470, 472, 475, 479, 2404]. X-AS [2404]. XCP [36]. Xen [1203]. Xen-based [1203]. XG [2993]. XGBoost [2777]. XGBoost-based [2777]. XM [885]. XPLIT [565].

YARA [2499]. YARA-based [2499]. years [1239]. YouTube [2009].

Z [1780]. Z-Monitor [1780]. zapping [562]. Zero [6, 2154]. Zero-configuration [6]. Zero-knowledge [2154]. Zeus [848]. ZigBee [687, 727, 1341]. Zone [156, 1788]. zones [2632].

References

Urdaneta:2009:WWA

[1] Guido Urdaneta, Guillaume Pierre, and Maarten van Steen. Wikipedia workload analysis for decentralized hosting. Computer Networks (Amsterdam, Netherlands: 1999), 53(11): 1830–1845, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). See corrigendum [67].

Akyildiz:2010:E

[2] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 54(1): 1–2, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2010:ACS

[3] Sunmyeng Kim, Young-Jong Cho, and Yong K. Kim. Admission control scheme based on priority access for wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 54 (1):3–12, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2010:MBS

[4] Chunxi Li and Changjia Chen. Measurement-based study on the relation between users’ watching behavior and network sharing in P2P VoD systems. Computer Networks (Amsterdam, Netherlands: 1999), 54(1): 13–27, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jiang:2010:EDI

[5] Yixin Jiang, Haojin Zhu, Minghui Shi, Xuemin (Sherman) Shen, and Chuang Lin. An efficient dynamic-identity based signature scheme for secure network coding. Computer Networks (Amsterdam, Netherlands: 1999), 54(1): 28–40, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ibanez:2010:HHZ

[6] Guillermo Ibáñez, Alberto García-Martínez, Juan A. Carral, Pedro A. González, Arturo Azcorra, and José M. Arco. HURP/HURBA: Zeroconfiguration hierarchical Up/Down routing and bridging architecture for Ethernet backbones and campus networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(1): 41–56, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huynh:2010:RTE

[7] Minh Huynh, Stuart Goose, and Prasant Mohapatra. Resilience technologies in Ethernet. Computer Networks (Amsterdam, Netherlands: 1999), 54 (1):57–78, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Munoz:2010:RVR

[8] Jose L. Muñoz, Oscar Esparza, Mónica Aguilar, Victor Carrascal, and Jordi Forné. RDSR-V. Reliable Dynamic Source Routing for video-streaming over mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(1):79–96, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stevens:2010:AAB

[9] Tim Stevens, Tim Wauters, Chris Develder, Filip De Turck, Bart Dhoedt, and Piet Demeester. Analysis of an anycast based overlay system for scalable service discovery and execution. Computer Networks (Amsterdam, Netherlands: 1999), 54(1):97– 111, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kumar:2010:LED

[10] Suman Kumar, Seung-Jong Park, and S. Sitharama Iyengar. A loss-event driven scalable fluid simulation method for high-speed networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(1):112–132, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Eklund:2010:TSF

[11] Johan Eklund, Karl-Johan Grinnemo, Stephan Baucke, and Anna Brunstrom. Tuning SCTP failover for carrier grade telephony signaling. Computer Networks (Amsterdam, Netherlands: 1999), 54(1):133–149, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Leung:2010:STG

[12] Yiu-Wing Leung. Sparse telephone gateway for Internet telephony. Computer Networks (Amsterdam, Netherlands: 1999), 54(1):150–164, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guo:2010:ROD

[13] Deke Guo, Yuan He, and Panlong Yang. Receiver-oriented design of Bloom filters for data-centric routing. Computer Networks (Amsterdam, Netherlands: 1999), 54(1):165– 174, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBa

[14] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(1):??, January 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Boutaba:2010:SIW

[15] Raouf Boutaba, Seán Murphy, and Albert Banchs. Special issue on “Wireless Multi-Hop Networking for Infrastructure Access”. Computer Networks (Amsterdam, Netherlands: 1999), 54 (2):175–177, February 15, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wellons:2010:APO

[16] Jonathan Wellons, Liang Dai, Yuan Xue, and Yui Cui. Augmenting predictive with oblivious routing for wireless mesh networks under traffic uncertainty. Computer Networks (Amsterdam, Netherlands: 1999), 54(2): 178–195, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2010:MLP

[17] Jian Zhang, Yuanzhu Peter Chen, and Ivan Marsic. MAC-layer proactive mixing for network coding in multi-hop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (2):196–207, February 15, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ward:2010:CRL

[18] Charles B. Ward and Nathan M. Wiegand. Complexity results on labeled shortest path problems from wireless routing metrics. Computer Networks (Amsterdam, Netherlands: 1999), 54 (2):208–217, February 15, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Toubiana:2010:GSA

[19] Vincent Toubiana, Houda Labiod, Laurent Reynaud, and Yvon Gourhant. A global security architecture for operated hybrid WLAN mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(2):218– 230, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Targon:2010:JGP

[20] Valerio Targon, Brunilde Sansò, and Antonio Capone. The joint Gateway Placement and Spatial Reuse Problem in Wireless Mesh Networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(2):231–240, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). See erratum [94].

Marina:2010:TCA

[21] Mahesh K. Marina, Samir R. Das, and Anand Prabhu Subramanian. A topology control approach for utilizing multiple channels in multi-radio wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (2):241–256, February 15, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bayer:2010:AAS

[22] Nico Bayer, Bangnan Xu, Veselin Rakocevic, and Joachim Habermann. Application-aware scheduling for VoIP in Wireless Mesh Networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(2):257–277, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Radwan:2010:ICG

[23] Ayman Radwan, Hossam S. Hassanein, and Abd-Elhamid M. Taha. Identifying the capacity gains of multihop cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (2):278–290, February 15, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bernardos:2010:BIB

[24] Carlos J. Bernardos, Maria Calderon, Ignacio Soto, Ana Beatriz Solana, and Kilian Weniger. Building an IP-based community wireless mesh network: Assessment of PACMAN as an IP address autoconfiguration protocol. Computer Networks (Amsterdam, Netherlands: 1999), 54(2):291–303, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhou:2010:SDB

[25] Jiazhen Zhou and Kenneth Mitchell. A scalable delay based analytical framework for CSMA/CA wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(2):304– 318, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Papadaki:2010:GSR

[26] Katerina Papadaki and Vasilis Friderikos. Gateway selection and routing in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(2):319–329, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2010:LSR

[27] Cheolgi Kim, Young-Bae Ko, and Nitin H. Vaidya. Link-state routing without broadcast storming for multichannel mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(2):330–340, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBb

[28] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(2):??, February 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hasib:2010:CLR

[29] Abdul Hasib and Abraham O. Fapojuwo. Cross-layer radio resource management in integrated WWAN and WLAN networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (3):341–356, February 26, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ortiz:2010:NIP

[30] Andrés Ortiz, Julio Ortega, Antonio F. Díaz, and Alberto Prieto. Network interfaces for programmable NICs and multicore platforms. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):357–376, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yu:2010:ABD

[31] Shucheng Yu, Kui Ren, and Wenjing Lou. Attribute-based on-demand multicast group setup with membership anonymity. Computer Networks (Amsterdam, Netherlands: 1999), 54(3): 377–386, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rahbar:2010:ABM

[32] Akbar Ghaffar Pour Rahbar and Oliver Yang. Agile bandwidth management techniques in slotted all-optical packet switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):387–403, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Golen:2010:USA

[33] Erik F. Golen, Sumita Mishra, and Nirmala Shenoy. An underwater sensor allocation scheme for a range dependent environment. Computer Networks (Amsterdam, Netherlands: 1999), 54 (3):404–415, February 26, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Crespo:2010:PFI

[34] Rui Gustavo Crespo. Predicting feature interactions by using inconsistency models. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):416– 427, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Onur:2010:SWS

[35] Ertan Onur, Cem Ersoy, Hakan Deliç, and Lale Akarun. Surveillance with wireless sensor networks in obstruction: Breach paths as watershed contours. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):428– 441, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2010:IXA

[36] Xiaowei Yang, Yanbin Lu, and Lei Zan. Improving XCP to achieve max– min fair bandwidth allocation. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):442–461, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chand:2010:ESS

[37] Satish Chand and Hari Om. Efficient staircase scheme with seamless channel transition mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):462–474, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2010:MPI

[38] Yixin Wu, Suman Kumar, and Seung-Jong Park. Measurement and performance issues of transport protocols over 10 Gbps high-speed optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):475– 488, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kulatunga:2010:ELM

[39] Chamil Kulatunga and Gorry Fairhurst. Enforcing layered multicast congestion control using ECN-nonce. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):489–505, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Feng:2010:NEG

[40] Xiang Feng and Francis C. M. Lau. A new economic generalized particle model for flow control. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):506–524, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBc

[41] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(3):??, February 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Syrotiuk:2010:CNE

[42] Violet R. Syrotiuk and Brahim Bensaou. Computer networks (Elsevier) special issue on advances in wireless and mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(4):525–526, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2010:SUV

[43] Uichin Lee and Mario Gerla. A survey of urban vehicular sensing platforms. Computer Networks (Amsterdam, Netherlands: 1999), 54(4): 527–544, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

TalebiFard:2010:ASC

[44] Peyman TalebiFard, Terrence Wong, and Victor C. M. Leung. Access and service convergence over the mobile Internet — a survey. Computer Networks (Amsterdam, Netherlands: 1999), 54 (4):545–557, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2010:MMS

[45] Zhenxia Zhang, Richard W. Pazzi, and Azzedine Boukerche. A mobility management scheme for wireless mesh networks based on a hybrid routing protocol. Computer Networks (Amsterdam, Netherlands: 1999), 54(4): 558–572, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ye:2010:UMJ

[46] Zhenzhen Ye and Alhussein A. Abouzeid. A unified model for joint throughput-overhead analysis of random access mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(4): 573–588, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Boldrini:2010:DPE

[47] Chiara Boldrini, Marco Conti, and Andrea Passarella. Design and performance evaluation of ContentPlace, a social-aware data dissemination system for opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(4):589–604, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Misic:2010:MBL

[48] Jelena Mišić and Vojislav B. Mišić. Making the best of limited resources: Optimized differential sensing in cognitive PANs. Computer Networks (Amsterdam, Netherlands: 1999), 54(4): 605–617, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hu:2010:CCC

[49] Zhengqing Hu and Chen-Khong Tham. CCMAC: Coordinated cooperative MAC for wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 54(4):618–630, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pryyma:2010:ATS

[50] Volodymyr Pryyma, Damla Turgut, and Ladislau Bölöni. Active time scheduling for rechargeable sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(4): 631–640, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2010:IMI

[51] Jeongkeun Lee, Jiho Ryu, Sung-Ju Lee, and Ted Taekyoung Kwon. Improved modeling of IEEE 802.11a PHY through fine-grained measurements. Computer Networks (Amsterdam, Netherlands: 1999), 54(4): 641–657, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rusak:2010:PBM

[52] Tal Rusak and Philip Levis. Physicallybased models of low-power wireless links using signal power simulation. Computer Networks (Amsterdam, Netherlands: 1999), 54(4): 658–673, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBd

[53] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(4):??, March 19, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Androutsellis-Theotokis:2010:MBA

[54] Stephanos Androutsellis-Theotokis, Diomidis Spinellis, Panos Louridas, and Kostas Stroggylos. A marketbased approach to managing the risk of peer-to-peer transactions. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):675–688, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ou:2010:PEK

[55] Zhonghong Ou, Erkki Harjula, Otso Kassinen, and Mika Ylianttila. Performance evaluation of a Kademlia-based communication-oriented P2P system under churn. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):

689–705, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shan:2010:BRO

[56] D. M. Shan, K. C. Chua, and G. Mohan. On burst rescheduling in OBS networks with partial wavelength conversion capability. Computer Networks (Amsterdam, Netherlands: 1999), 54 (5):706–715, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2010:IFN

[57] Bin Wu, Kwan L. Yeung, and Pin-Han Ho. ILP formulations for non-simple p-cycle and p-trail design in WDM mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (5):716–725, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ceken:2010:IAV

[58] Celal Çeken, Serhan Yarkan, and Hüseyin Arslan. Interference aware vertical handoff decision algorithm for quality of service support in wireless heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):726–740, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anagnostopoulos:2010:EEF

[59] Ioannis Anagnostopoulos, Christos Anagnostopoulos, and Dimitrios D. Vergados. Estimating evolution of freshness in Internet cache directories under the capture–recapture methodology. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):741– 765, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gupta:2010:PSS

[60] Gagan Raj Gupta and Ness B. Shroff. Practical scheduling schemes with throughput guarantees for multi-hop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (5):766–780, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rodero-Merino:2010:PRW

[61] Luis Rodero-Merino, Antonio Fernández Anta, Luis López, and Vicent Cholvi. Performance of random walks in onehop replication networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):781–796, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dalalah:2010:RTO

[62] Doraid Dalalah. Real-time optimization flow control. Computer Networks (Amsterdam, Netherlands: 1999), 54 (5):797–810, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lei:2010:SDD

[63] Ming Lei, Susan V. Vrbsky, and Yang Xiao. Scheduling on-demand data broadcast in mixed-type request environments. Computer Networks (Amsterdam, Netherlands: 1999), 54(5): 811–825, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Torkestani:2010:IBF

[64] Javad Akbari Torkestani and Mohammad Reza Meybodi. An intelligent backbone formation algorithm for wireless ad hoc networks based on distributed learning automata. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):826–843, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2010:GFE

[65] Seungjoon Lee, Bobby Bhattacharjee, Suman Banerjee, and Bo Han. A general framework for efficient geographic routing in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):844–861, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chowdhury:2010:SNV

[66] N. M. Mosharaf Kabir Chowdhury and Raouf Boutaba. A survey of network virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 54 (5):862–876, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Urdaneta:2010:CWW

[67] Guido Urdaneta, Guillaume Pierre, and Maarten van Steen. Corrigendum to “Wikipedia workload analysis for decentralized hosting” [Computer Networks 53 (11) (2009) 1830–1845]. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):877–878, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). See [1].

Anonymous:2010:EBe

[68] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):??, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2010:SIN

[69] Eitan Altman, Tamer Başar, Emma Hart, Daniele Miorandi, Aris L. Moustakas, and Stavros Toumpis. Special issue on “New Network Paradigms”. Computer Networks (Amsterdam, Netherlands: 1999), 54(6): 879–880, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dressler:2010:SBI

[70] Falko Dressler and Ozgur B. Akan. A survey on bio-inspired networking. Computer Networks (Amsterdam, Netherlands: 1999), 54(6):881– 900, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Meisel:2010:TBI

[71] Michael Meisel, Vasileios Pappas, and Lixia Zhang. A taxonomy of biologically inspired research in computer networking. Computer Networks (Amsterdam, Netherlands: 1999), 54(6): 901–916, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kitagawa:2010:OSC

[72] Koichiro Kitagawa and Toshiyuki Tanaka. Optimization of sequences in CDMA systems: a statistical– mechanics approach. Computer Networks (Amsterdam, Netherlands: 1999), 54(6):917–924, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2010:OSA

[73] Yuedong Xu, John C. S. Lui, and Dah-Ming Chiu. On oligopoly spectrum allocation game in cognitive radio networks with capacity constraints. Computer Networks (Amsterdam, Netherlands: 1999), 54(6):925–943, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Miorandi:2010:SEE

[74] Daniele Miorandi, Lidia Yamamoto, and Francesco De Pellegrini. A survey of evolutionary and embryogenic approaches to autonomic networking. Computer Networks (Amsterdam, Netherlands: 1999), 54(6): 944–959, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Konstantinidis:2010:MOE

[75] Andreas Konstantinidis, Kun Yang, Qingfu Zhang, and Demetrios Zeinalipour-Yazti. A multi-objective evolutionary algorithm for the deployment and power assignment problem in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (6):960–976, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guerrero:2010:TFL

[76] Cesar D. Guerrero and Miguel A. Labrador. Traceband: a fast, low overhead and accurate tool for available bandwidth estimation and monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 54(6): 977–990, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nordio:2010:SRS

[77] Alessandro Nordio, Carla-Fabiana Chiasserini, and Armando Muscariello. Signal reconstruction in sensor networks with flat and clustered topologies. Computer Networks (Amsterdam, Netherlands: 1999), 54(6):991–1004, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Silva:2010:CEG

[78] Alonso Silva, Eitan Altman, Pierre Bernhard, and Mérouane Debbah. Continuum equilibria and global optimization for routing in dense static ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (6):1005–1018, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Heegaard:2010:ORD

[79] Poul E. Heegaard and Otto J. Wittner. Overhead reduction in a distributed path management system. Computer Networks (Amsterdam, Netherlands: 1999), 54(6):1019–1041, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2010:SNU

[80] Huijuan Wang and Piet Van Mieghem. Sampling networks by the union of m shortest path trees. Computer Networks (Amsterdam, Netherlands: 1999), 54(6):1042–1053, April 29, 2010.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBf

[81] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(6):??, April 29, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Keralapura:2010:NSL

[82] Ram Keralapura, Antonio Nucci, and Chen-Nee Chuah. A novel self-learning architecture for P2P traffic classification in high speed networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(7):1055–1068, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2010:BEP

[83] Lijun Wang, Lin Cai, Xinzhi Liu, and Xuemin (Sherman) Shen. Bounds estimation and practical stability of AIMD/RED systems with time delays. Computer Networks (Amsterdam, Netherlands: 1999), 54(7):1069–1082, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ong:2010:CRR

[84] Eng Hwee Ong and Jamil Y. Khan. Cooperative radio resource management framework for future IP-based multiple radio access technologies environment. Computer Networks (Amsterdam, Netherlands: 1999), 54(7): 1083–1107, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hwang:2010:TPF

[85] Iksoon Hwang and Ana Cavalli. Testing a probabilistic FSM using interval estimation. Computer Networks (Amsterdam, Netherlands: 1999), 54 (7):1108–1125, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Joseph:2010:CCL

[86] John Felix Charles Joseph, Amitabha Das, Bu-Sung Lee, and Boon-Chong Seet. CARRADS: Cross layer based adaptive real-time routing attack detection system for MANETS. Computer Networks (Amsterdam, Netherlands: 1999), 54(7):1126–1141, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Schutz:2010:DIN

[87] Simon Schütz, Henrik Abrahamsson, Bengt Ahlgren, and Marcus Brunner. Design and implementation of the Node Identity Internetworking Architecture. Computer Networks (Amsterdam, Netherlands: 1999), 54(7): 1142–1154, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lu:2010:CLE

[88] Jinchang Lu and Maode Ma. A crosslayer elastic CAC and holistic opportunistic scheduling for QoS support in WiMAX. Computer Networks (Amsterdam, Netherlands: 1999), 54(7): 1155–1168, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yan:2010:CAI

[89] Guanhua Yan, Stephan Eidenbenz, Sunil Thulasidasan, Pallab Datta, and Venkatesh Ramaswamy. Criticality analysis of Internet infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 54(7):1169–1182, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Biradar:2010:LSB

[90] Rajashekhar Biradar, Sunilkumar Manvi, and Mylara Reddy. Link stability based multicast routing scheme in MANET. Computer Networks (Amsterdam, Netherlands: 1999), 54(7): 1183–1196, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vilhar:2010:INT

[91] Andrej Vilhar, Roman Novak, and Gorazd Kandus. The impact of network topology on the performance of MAP selection algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 54(7):1197–1209, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garbinato:2010:CAB

[92] Benoît Garbinato, Adrian Holzer, and François Vessaz. Context-aware broadcasting approaches in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(7): 1210–1228, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2010:PCP

[93] Ke Xu, Ming Zhang, Jiangchuan Liu, Zhijing Qin, and Mingjiang Ye. Proxy caching for peer-to-peer live streaming. Computer Networks (Amsterdam, Netherlands: 1999), 54(7):1229–1241, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Targon:2010:EJG

[94] Valerio Targon, Brunilde Sansò, and Antonio Capone. Erratum to “The joint Gateway Placement and Spatial Reuse Problem in Wireless Mesh Networks” [Computer Networks 54 (2010) 231–240]. Computer Networks (Amsterdam, Netherlands: 1999), 54(7): 1242, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). See [20].

Anonymous:2010:EBg

[95] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(7):??, May 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Plattner:2010:RSN

[96] Bernhard Plattner, David Hutchison, and James P. G. Sterbenz. Resilient and survivable networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(8):1243–1244, June 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sterbenz:2010:RSC

[97] James P. G. Sterbenz, David Hutchison, Egemen K. Çetinkaya, Abdul Jabbar, Justin P. Rohrer, Marcus Schöller, and Paul Smith. Resilience and survivability in communication networks: Strategies, principles, and survey of disciplines. Computer Networks (Amsterdam, Netherlands: 1999), 54(8): 1245–1265, June 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guidoni:2010:DRH

[98] Daniel L. Guidoni, Raquel A. F. Mini, and Antonio A. F. Loureiro. On the design of resilient heterogeneous wireless sensor networks based on small world concepts. Computer Networks (Amsterdam, Netherlands: 1999), 54 (8):1266–1281, June 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2010:HHS

[99] Zhichun Li, Yan Gao, and Yan Chen. HiFIND: a high-speed flow-level intrusion detection approach with DoS resiliency. Computer Networks (Amsterdam, Netherlands: 1999), 54(8): 1282–1299, June 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Menth:2010:LFA

[100] Michael Menth, Matthias Hartmann, Rüdiger Martin, Tarik Čičić, and Amund Kvalbein. Loop-free alternates and not-via addresses: a proper combination for IP fast reroute? Computer Networks (Amsterdam, Netherlands: 1999), 54(8):1300–1315, June 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Park:2010:APP

[101] Kunwoo Park, Sangheon Pack, and Ted “Taekyoung” Kwon. An adaptive peer-to-peer live streaming system with incentives for resilience. Computer Networks (Amsterdam, Netherlands: 1999), 54(8):1316–1327, June 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

delRio:2010:RAD

[102] P. M. Santiago del Río, J. A. Hernández, J. Aracil, J. E. López de Vergara, J. Domżał, R. Wójcik, P. Chołda, K. Wajda, J. P. Fernández Palacios, Ó. González de Dios, and R. Duque. A reliability analysis of Double-Ring topologies with Dual Attachment using p-cycles for optical metro networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (8):1328–1341, June 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBh

[103] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(8):??, June 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Viejo:2010:USN

[104] Alexandre Viejo and Jordi Castellà-Roca. Using social networks to distort users’ profiles generated by Web search engines. Computer Networks (Amsterdam, Netherlands: 1999), 54 (9):1343–1357, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rosberg:2010:NRM

[105] Zvi Rosberg, John Matthews, and Moshe Zukerman. A network rate management protocol with TCP congestion control and fairness for all. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):1358–1374, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shi:2010:OFC

[106] Xingang Shi, Dah-Ming Chiu, and John C. S. Lui. An online framework for catching top spreaders and scanners. Computer Networks (Amsterdam, Netherlands: 1999), 54(9): 1375–1388, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Argibay-Losada:2010:NDE

[107] Pablo J. Argibay-Losada, Andrés Suárez-González, Cándido López-García, and Manuel Fernández-Veiga. A new design for end-to-end proportional loss differentiation in IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(9): 1389–1403, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Domenech:2010:UCW

[108] Josep Domenech, Jose A. Gil, Julio Sahuquillo, and Ana Pont. Using current Web page structure to improve prefetching performance. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):1404–1417, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2010:TCS

[109] Kyung-Hoe Kim, Kwang-Min Jeong, Chul-Hee Kang, and Seung-Joon Seok. A transmission control SCTP for realtime multimedia streaming. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):1418–1425, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guo:2010:CMP

[110] Ming-Fei Guo, Xinbing Wang, and Min-You Wu. On the capacity of multi-packet reception enabled multichannel multi-interface wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(9): 1426–1439, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chandra:2010:HET

[111] Joydeep Chandra, Santosh Kumar Shaw, and Niloy Ganguly. HPC5: an efficient topology generation mechanism for Gnutella networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):1440–1459, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yu:2010:RSS

[112] Yao Yu, Lei Guo, Xingwei Wang, and Cuixiang Liu. Routing security scheme based on reputation evaluation in hierarchical ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):1460–1469, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ahmed:2010:EAM

[113] Sabbir Ahmed, Gour C. Karmakar, and Joarder Kamruzzaman. An environment-aware mobility model for wireless ad hoc network. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):1470–1489, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Laki:2010:MBA

[114] S. Laki, P. Mátray, P. Hága, I. Csabai, and G. Vattay. A model based approach for improving router geolocation. Computer Networks (Amsterdam, Netherlands: 1999), 54(9): 1490–1501, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Peris-Lopez:2010:VAR

[115] Pedro Peris-Lopez, Julio C. Hernandez-Castro, Juan M. E. Tapiador, Tieyan Li, and Yingjiu Li. Vulnerability analysis of RFID protocols for tag ownership transfer. Computer Networks (Amsterdam, Netherlands: 1999), 54 (9):1502–1508, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2010:DPS

[116] Eitan Altman, Tania Jiménez, and Daniel Kofman. Discriminatory processor sharing queues with stationary ergodic service times and the performance of TCP in overload. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):1509–1519, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2010:EUA

[117] Tsu-Yang Wu and Yuh-Min Tseng. An efficient user authentication and key exchange protocol for mobile client–server environment. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):1520–1530, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBi

[118] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(9):??, June 17, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pinto:2010:SMI

[119] António Pinto and Manuel Ricardo. Secure multicast in IPTV services. Computer Networks (Amsterdam, Netherlands: 1999), 54(10):1531–1542, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2010:QDS

[120] Yueping Zhang, Yong Xiong, Steve Liu, and Dmitri Loguinov. Queuing dynamics and single-link stability of delay-based window congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 54(10):1543–1553, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Qazi:2010:IRM

[121] Sameer Qazi and Tim Moors. On the impact of routing matrix inconsistencies on statistical path monitoring in overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (10):1554–1572, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ahmed:2010:IUW

[122] Bazil Taha Ahmed, Jose Luis Masa Campos, and Jorge A. Ruiz-Cruz. Impact of Ultra Wide Band emission on WiMAX systems at 2.5 and 3.5 GHz. Computer Networks (Amsterdam, Netherlands: 1999), 54(10): 1573–1583, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zubow:2010:GSA

[123] Anatolij Zubow, Daniel Camps Mur, Xavier Perez Costa, and Paolo Favaro. Greedy scheduling algorithm (GSA) — Design and evaluation of an efficient and flexible WiMAX OFDMA scheduling solution. Computer Networks (Amsterdam, Netherlands: 1999), 54(10): 1584–1606, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sorensen:2010:SDF

[124] Jesper H. Sørensen, Rasmus Krigslund, Petar Popovski, Toshiaki Koike Akino, and Torben Larsen. Scalable DeNoiseand-Forward in bidirectional relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(10): 1607–1614, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Oikonomou:2010:PFE

[125] Konstantinos Oikonomou, Dimitrios Kogias, and Ioannis Stavrakakis. Probabilistic flooding for efficient information dissemination in random graph topologies. Computer Networks (Amsterdam, Netherlands: 1999), 54(10): 1615–1629, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sitanayah:2010:HAF

[126] Lanny Sitanayah, Amitava Datta, and Rachel Cardell-Oliver. Heuristic algorithm for finding boundary cycles in location-free low density wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(10): 1630–1645, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Islam:2010:SAD

[127] Salekul Islam and J. William Atwood. Sender access and data distribution control for inter-domain multicast groups. Computer Networks (Amsterdam, Netherlands: 1999), 54(10): 1646–1671, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kritzinger:2010:CRV

[128] P. S. Kritzinger, Henry Msiska, Tino Mundangepfupfu, Paolo Pileggi, and Andrew Symington. Comparing the results from various performance models of IEEE 802.11g DCF. Computer Networks (Amsterdam, Netherlands: 1999), 54(10):1672–1682, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Booker:2010:ETL

[129] G. Booker, A. Sprintson, E. Zechman, C. Singh, and S. Guikema. Efficient traffic loss evaluation for transport backbone networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(10):1683–1691, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khoukhi:2010:IQM

[130] Lyes Khoukhi and Soumaya Cherkaoui. Intelligent QoS management for multimedia services support in wireless mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(10):1692–1706, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ke:2010:PAH

[131] Kai-Wei Ke, Chen-Nien Tsai, and Ho-Ting Wu. Performance analysis for hierarchical resource allocation in multiplexed mobile packet data networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(10): 1707–1725, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBj

[132] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(10):??, July 1, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2010:DMM

[133] Hyang-Won Lee, Jeong woo Cho, and Song Chong. Distributed max–min flow control for multi-rate overlay multicast. Computer Networks (Amsterdam, Netherlands: 1999), 54(11): 1727–1738, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tran:2010:ECB

[134] Duc A. Tran and Cuong Pham. Enabling content-based publish/subscribe services in cooperative P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(11): 1739–1749, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Casas:2010:OVA

[135] P. Casas, S. Vaton, L. Fillatre, and I. Nikiforov. Optimal volume anomaly detection and isolation in large-scale IP networks using coarse-grained measurements. Computer Networks (Amsterdam, Netherlands: 1999), 54(11): 1750–1766, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Feng:2010:PPI

[136] Xiang Feng and Francis C. M. Lau. Parallel physics-inspired waterflow particle mechanics algorithm for load rebalancing. Computer Networks (Amsterdam, Netherlands: 1999), 54(11): 1767–1777, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Papadopoulos:2010:SIF

[137] Fragkiskos Papadopoulos. On scaling the IEEE 802.11 to facilitate scalable wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (11):1778–1791, August 2, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fawaz:2010:DBC

[138] Wissam Fawaz, Iyad Ouaiss, Ken Chen, and Harry Perros. Deadlinebased connection setup in wavelengthrouted WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(11):1792–1804, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guney:2010:EIP

[139] Evren Güney, Necati Aras, İ. Kuban Altınel, and Cem Ersoy. Efficient integer programming formulations for optimum sink location and routing in heterogeneous wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(11): 1805–1822, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chi:2010:NCB

[140] Kaikai Chi, Xiaohong Jiang, and Susumu Horiguchi. Network codingbased reliable multicast in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(11): 1823–1836, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2010:BVS

[141] Ruishan Zhang, Xinyuan Wang, Xiaohui Yang, and Xuxian Jiang. On the billing vulnerabilities of SIP-based VoIP systems. Computer Networks (Amsterdam, Netherlands: 1999), 54 (11):1837–1847, August 2, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yan:2010:SVH

[142] Xiaohuan Yan, Y. Ahmet Şekercioğlu, and Sathya Narayanan. A survey of vertical handover decision algorithms in Fourth Generation heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (11):1848–1863, August 2, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tamma:2010:TMT

[143] Bheemarjuna Reddy Tamma, Anirudh Badam, C. Siva Ram Murthy, and Ramesh R. Rao. K-Tree: a multiple tree video multicast protocol for ad hoc wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (11):1864–1884, August 2, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mai:2010:FFS

[144] Jianning Mai, Ashwin Sridharan, Hui Zang, and Chen-Nee Chuah. Fast Filtered Sampling. Computer Networks (Amsterdam, Netherlands: 1999), 54 (11):1885–1898, August 2, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lu:2010:ALA

[145] Xiaofeng Lu, Pan Hui, Don Towsley, Juahua Pu, and Zhang Xiong. Antilocalization anonymous routing for Delay Tolerant Network. Computer Networks (Amsterdam, Netherlands: 1999), 54(11):1899–1910, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kuo:2010:QBP

[146] Yu-Chen Kuo. Quorum-based powersaving multicast protocols in the asynchronous ad hoc network. Computer Networks (Amsterdam, Netherlands: 1999), 54(11):1911–1922, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBk

[147] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(11):??, August 2, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garcia-Lopez:2010:GES

[148] Pedro García-López, Michael W. Sobolewski, and Marc Sánchez-Artigas. Guest editorial for the special issue collaborative P2P systems. Computer Networks (Amsterdam, Netherlands: 1999), 54(12):1923–1925, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Juste:2010:SEW

[149] Pierre St. Juste, David Wolinsky, P. Oscar Boykin, Michael J. Covington, and Renato J. Figueiredo. SocialVPN: Enabling wide-area collaboration with integrated social and overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (12):1926–1938, August 26, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Oster:2010:BCP

[150] Gérald Oster, Rubén Mondéjar, Pascal Molli, and Sergiu Dumitriu. Building a collaborative peer-to-peer wiki system on a structured overlay. Computer Networks (Amsterdam, Netherlands: 1999), 54(12):1939–1952, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kulkarni:2010:BDN

[151] Santosh Kulkarni, Scott Douglas, and David Churchill. Badumna: a decentralised network engine for virtual environments. Computer Networks (Amsterdam, Netherlands: 1999), 54(12): 1953–1967, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sanchez-Artigas:2010:POD

[152] Marc Sànchez-Artigas, Jordi Pujol-Ahulló, Lluis Pamies-Juarez, and Pedro García-López. p2pWeb: an open, decentralized infrastructure of Web servers for sharing ephemeral Web content. Computer Networks (Amsterdam, Netherlands: 1999), 54(12):1968– 1985, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rzadca:2010:MOO

[153] Krzysztof Rzadca, Jackson Tan Teck Yong, and Anwitaman Datta. Multiobjective optimization of multicast overlays for collaborative applications. Computer Networks (Amsterdam, Netherlands: 1999), 54(12):1986– 2006, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2010:RMC

[154] Di Wu, Chao Liang, Yong Liu, and Keith W. Ross. Redesigning multichannel P2P live video systems with View-Upload Decoupling. Computer Networks (Amsterdam, Netherlands: 1999), 54(12):2007–2018, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Papapetrou:2010:PCD

[155] Odysseas Papapetrou, Wolf Siberski, and Wolfgang Nejdl. PCIR: Combining DHTs and peer clusters for efficient full-text P2P indexing. Computer Networks (Amsterdam, Netherlands: 1999), 54(12):2019–2040, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Beijar:2010:ZIO

[156] Nicklas S. Beijar. Zone indexing: Optimizing the balance between searching and indexing in a loosely structured overlay. Computer Networks (Amsterdam, Netherlands: 1999), 54(12): 2041–2055, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tirado:2010:APS

[157] Juan M. Tirado, Daniel Higuero, Florin Isaila, Jesús Carretero, and Adriana Iamnitchi. Affinity P2P: a self-organizing content-based localityaware collaborative peer-to-peer network. Computer Networks (Amsterdam, Netherlands: 1999), 54(12):2056– 2070, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cuevas:2010:CPS

[158] Rubén Cuevas, Ángel Cuevas, Albert Cabellos-Aparicio, Loránd Jakab, and Carmen Guerrero. A collaborative P2P scheme for NAT Traversal Server discovery based on topological information. Computer Networks (Amsterdam, Netherlands: 1999), 54(12): 2071–2085, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jesi:2010:SPS

[159] Gian Paolo Jesi, Alberto Montresor, and Maarten van Steen. Secure peer sampling. Computer Networks (Amsterdam, Netherlands: 1999), 54(12): 2086–2098, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBl

[160] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(12):??, August 26, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Menth:2010:PBM

[161] Michael Menth and Frank Lehrieder. PCN-based measured rate termination. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2099– 2116, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pirmez:2010:SNS

[162] Luci Pirmez, Jaime C. Carvalho, Jr., Flávia C. Delicato, Fábio Protti, Luiz F. R. C. Carmo, Paulo F. Pires, and Marcos Pirmez. SUTIL — Network selection based on utility function and integer linear programming. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2117–2136, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mao:2010:EES

[163] Xiaomao Mao, Huifang Chen, Peiliang Qiu, and Zhaoyang Zhang. Energyefficient scheduling for multiple access in wireless sensor networks: a job scheduling method. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2137–2146, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ge:2010:DEE

[164] Yu Ge, Chen-Khong Tham, Peng-Yong Kong, and Yew-Hock Ang. Dynamic end-to-end capacity in IEEE 802.16 wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2147–2165, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akar:2010:SPW

[165] N. Akar, C. Raffaelli, M. Savi, and E. Karasan. Shared-per-wavelength asynchronous optical packet switching: a comparative analysis. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2166–2181, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yue:2010:AHC

[166] Chuan Yue, Mengjun Xie, and Haining Wang. An automatic HTTP cookie management system. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2182–2198, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cano:2010:TEP

[167] C. Cano, B. Bellalta, A. Sfairopoulou, and J. Barceló. Tuning the EDCA parameters in WLANs with heterogeneous traffic: a flow-level analysis. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2199– 2214, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stavrou:2010:SSM

[168] Eliana Stavrou and Andreas Pitsillides. A survey on secure multipath routing protocols in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2215–2238, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Antoniou:2010:CUN

[169] Josephina Antoniou, Vicky Papadopoulou, Vasos Vassiliou, and Andreas Pitsillides. Cooperative user–network interactions in next generation communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (13):2239–2255, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ahmadian:2010:SEA

[170] Zahra Ahmadian, Somayeh Salimi, and Ahmad Salahi. Security enhancements against UMTS–GSM interworking attacks. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2256– 2270, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Thulasiraman:2010:IAR

[171] Preetha Thulasiraman and Xuemin (Sherman) Shen. Interference aware resource allocation for hybrid hierarchical wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (13):2271–2280, September 15, 2010.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Maille:2010:PWH

[172] Patrick Maillé and Bruno Tuffin. Price war in heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2281– 2292, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Santosa:2010:EDR

[173] Rully Adrian Santosa, Bu-Sung Lee, and Chai Kiat Yeo. Efficient DSR route request flooding with directional antennas. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):2293– 2309, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Padhye:2010:ESS

[174] Jaideep D. Padhye, Kush Kothari, Madhu Venkateshaiah, and Matthew Wright. Evading stepping-stone detection under the cloak of streaming media with SNEAK. Computer Networks (Amsterdam, Netherlands: 1999), 54 (13):2310–2325, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBm

[175] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(13):??, September 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marfia:2010:TLD

[176] Gustavo Marfia, Claudio E. Palazzi, Giovanni Pau, Mario Gerla, and Marco Roccetti. TCP Libra: Derivation, analysis, and comparison with other RTTfair TCPs. Computer Networks (Amsterdam, Netherlands: 1999), 54(14): 2327–2344, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Saxena:2010:CEC

[177] Mohit Saxena and Ramana Rao Kompella. CLAMP: Efficient class-based sampling for flexible flow monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 54(14):2345– 2356, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Berzin:2010:HML

[178] Oleg Berzin. Hierarchical Mobility Label Based Network: System model and performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 54(14):2357–2382, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Doss:2010:IDM

[179] Robin Doss, Gang Li, Vicky Mak, and Menik Tissera. Information discovery in mission-critical wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(14): 2383–2399, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Duraes:2010:CBA

[180] Gilvan M. Durães, André Soares, José R. Amazonas, and William Giozza. The choice of the best among the shortest routes in transparent optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (14):2400–2409, October 6, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Esnaashari:2010:LAB

[181] M. Esnaashari and M. R. Meybodi. A learning automata based scheduling solution to the dynamic point coverage problem in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(14): 2410–2438, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhou:2010:EAN

[182] Junyi Zhou and Jing Shi. Error analysis of non-collaborative wireless localization in circular-shaped regions. Computer Networks (Amsterdam, Netherlands: 1999), 54(14):2439–2452, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Secci:2010:LSR

[183] Stefano Secci, Jean-Louis Rougier, and Achille Pattavina. AS-level source routing for multi-provider connectionoriented services. Computer Networks (Amsterdam, Netherlands: 1999), 54 (14):2453–2467, October 6, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2010:PAB

[184] Hyewon Lee, Ilenia Tinnirello, Jeonggyun Yu, and Sunghyun Choi. A performance analysis of block ACK scheme for IEEE 802.11e networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(14):2468–2481, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2010:EMO

[185] Lei Zhang, Bo Qin, Qianhong Wu, and Futai Zhang. Efficient many-to-one authentication with certificateless aggregate signatures. Computer Networks (Amsterdam, Netherlands: 1999), 54 (14):2482–2491, October 6, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pelsser:2010:PSN

[186] Cristel Pelsser, Steve Uhlig, Tomonori Takeda, Bruno Quoitin, and Kohei Shiomoto. Providing scalable NH-diverse iBGP route re-distribution to achieve sub-second switch-over time. Computer Networks (Amsterdam, Netherlands: 1999), 54(14):2492–2505, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Muhlbauer:2010:IRP

[187] Wolfgang Mühlbauer, Steve Uhlig, Anja Feldmann, Olaf Maennel, Bruno Quoitin, and Bingjie Fu. Impact of routing parameters on route diversity and path inflation. Computer Networks (Amsterdam, Netherlands: 1999), 54 (14):2506–2518, October 6, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gelabert:2010:SSC

[188] Xavier Gelabert, Oriol Sallent, Jordi Pérez-Romero, and Ramon Agustí. Spectrum sharing in cognitive radio networks with imperfect sensing: a discrete-time Markov model. Computer Networks (Amsterdam, Netherlands: 1999), 54(14):2519–2536, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2010:GTC

[189] Beibei Wang, Yongle Wu, and K. J. Ray Liu. Game theory for cognitive radio networks: an overview. Computer Networks (Amsterdam, Netherlands: 1999), 54(14):2537–2561, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2010:AAB

[190] Yueping Zhang and Dmitri Loguinov. ABS: Adaptive buffer sizing for heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (14):2562–2574, October 6, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBn

[191] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(14):??, October 6, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bejerano:2010:NFP

[192] Yigal Bejerano, Seung-Jae Han, and Mark Smith. A novel frequency planning algorithm for mitigating unfairness in wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 54(15):2575–2590, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Simplicio:2010:SKM

[193] Marcos A. Simplício, Jr., Paulo S. L. M. Barreto, Cintia B. Margi, and Tereza C. M. B. Carvalho. A survey on key management mechanisms for distributed Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(15):2591– 2612, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2010:SAS

[194] Hui Yang, Stephen D. Patek, and Zhenyu Yan. Strategies for adaptive smoothing and rebuffering via dynamic network calculus. Computer Networks (Amsterdam, Netherlands: 1999), 54 (15):2613–2625, October 28, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Conti:2010:BML

[195] P. L. Conti, L. De Giovanni, and M. Naldi. Blind maximum likelihood estimation of traffic matrices under long-range dependent traffic. Computer Networks (Amsterdam, Netherlands: 1999), 54(15):2626–2639, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ghazisaidi:2010:TEA

[196] Navid Ghazisaidi and Martin Maier. Techno-economic analysis of EPON and WiMAX for future Fiber-Wireless (FiWi) networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (15):2640–2650, October 28, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marin-Lopez:2010:STP

[197] R. Marin-Lopez, F. Pereniguez, F. Bernal, and A. F. Gomez. Secure three-party key distribution protocol for fast network access in EAP-based wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (15):2651–2673, October 28, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hung:2010:PAI

[198] Fu-Yi Hung and Ivan Marsic. Performance analysis of the IEEE 802.11 DCF in the presence of the hidden stations. Computer Networks (Amsterdam, Netherlands: 1999), 54(15): 2674–2687, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alemdar:2010:WSN

[199] Hande Alemdar and Cem Ersoy. Wireless sensor networks for healthcare: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 54(15):2688– 2710, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Macia-Fernandez:2010:DTL

[200] Gabriel Maciá-Fernández, Rafael A. Rodríguez-Gómez, and Jesús E. Díaz-Verdejo. Defense techniques for lowrate DoS attacks against application servers. Computer Networks (Amsterdam, Netherlands: 1999), 54(15): 2711–2727, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2010:STD

[201] Qishi Wu, Nageswara S. V. Rao, Xukang Lu, and Ki-Hyeon Kwon. Stabilizing transport dynamics of control channels over wide-area networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(15):2728– 2743, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lin:2010:RRB

[202] Xiaodong Lin, Rongxing Lu, Davis Kwan, and Xuemin (Sherman) Shen. REACT: an RFID-based privacypreserving children tracking scheme for large amusement parks. Computer Networks (Amsterdam, Netherlands: 1999), 54(15):2744–2755, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sun:2010:PDA

[203] Xin Sun, Ruben Torres, and Sanjay Rao. Preventing DDoS attacks on Internet servers exploiting P2P systems. Computer Networks (Amsterdam, Netherlands: 1999), 54(15):2756– 2774, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2010:TBP

[204] Hua Wang, Xiangxu Meng, Shuai Li, and Hong Xu. A tree-based particle swarm optimization for multicast routing. Computer Networks (Amsterdam, Netherlands: 1999), 54(15): 2775–2786, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Atzori:2010:ITS

[205] Luigi Atzori, Antonio Iera, and Giacomo Morabito. The Internet of Things: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 54 (15):2787–2805, October 28, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBo

[206] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(15):??, October 28, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Feridun:2010:E

[207] Metin Feridun and Joseph Hellerstein. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 54(16): 2807–2808, November 15, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liao:2010:MNT

[208] Qi Liao, Andrew Blaich, Dirk Van-Bruggen, and Aaron Striegel. Managing networks through context: Graph visualization and exploration. Computer Networks (Amsterdam, Netherlands: 1999), 54(16):2809–2824, November 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Feeney:2010:EDM

[209] Kevin Feeney, Rob Brennan, John Keeney, Hendrik Thomas, Dave Lewis, Aidan Boran, and Declan O’Sullivan. Enabling decentralised management through federation. Computer Networks (Amsterdam, Netherlands: 1999), 54(16):2825–2839, November 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guo:2010:EUO

[210] Bin Guo, Daqing Zhang, and Michita Imai. Enabling user-oriented management for ubiquitous computing:

The meta-design approach. Computer Networks (Amsterdam, Netherlands: 1999), 54(16):2840–2855, November 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Berl:2010:NVE

[211] Andreas Berl, Nicholas Race, Johnathan Ishmael, and Hermann de Meer. Network virtualization in energyefficient office environments. Computer Networks (Amsterdam, Netherlands: 1999), 54(16):2856–2868, November 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mukherjee:2010:MDC

[212] Tridib Mukherjee, Ayan Banerjee, Georgios Varsamopoulos, and Sandeep K. S. Gupta. Model-driven coordinated management of data centers. Computer Networks (Amsterdam, Netherlands: 1999), 54(16):2869–2886, November 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bahi:2010:DEB

[213] Jacques M. Bahi and Michel Salomon. A decentralized energy-based diffusion algorithm to increase the lifetime of MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 54(16): 2887–2898, November 15, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marmol:2010:TPA

[214] Félix Gómez Mármol, Joao Girao, and Gregorio Martínez Pérez. TRIMS, a privacy-aware trust and reputation model for identity management systems. Computer Networks (Amsterdam, Netherlands: 1999), 54(16):2899– 2912, November 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBp

[215] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(16):??, November 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

LeCadre:2010:IGB

[216] Hélène Le Cadre and Mustapha Bouhtou. An interconnection game between mobile network operators: Hidden information forecasting using expert advice fusion. Computer Networks (Amsterdam, Netherlands: 1999), 54(17): 2913–2942, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Eu:2010:ORW

[217] Zhi Ang Eu, Hwee-Pink Tan, and Winston K. G. Seah. Opportunistic routing in wireless sensor networks powered by ambient energy harvesting. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):2943– 2966, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2010:PSW

[218] Jongdeog Lee, Krasimira Kapitanova, and Sang H. Son. The price of security in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):2967–2978, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Macedo:2010:TPD

[219] Daniel F. Macedo, Aldri L. dos Santos, Luiz H. A. Correia, José M. Nogueira, and Guy Pujolle. Transmission power and data rate aware routing on wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(17): 2979–2990, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cobo:2010:ABR

[220] Luis Cobo, Alejandro Quintero, and Samuel Pierre. Ant-based routing for wireless multimedia sensor networks using multiple QoS metrics. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):2991–3010, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2010:CAR

[221] Seongkwan Kim, Lochan Verma, Sunghyun Choi, and Daji Qiao. Collision-Aware Rate Adaptation in multi-rate WLANs: Design and implementation. Computer Networks (Amsterdam, Netherlands: 1999), 54(17): 3011–3030, December 3, 2010. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alyfantis:2010:DQS

[222] George Alyfantis, Stathes Hadjiefthymiades, and Lazaros Merakos. Distributed QoS-sensitive band selection strategies for dynamic spectrum access in unlicensed CDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):3031– 3048, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tsai:2010:DCB

[223] Ming-Fong Tsai, Tzu-Chi Huang, Ce-Kuen Shieh, and Kuo-Chih Chu. Dynamical combination of byte level and Sub-Packet level FEC in HARQ mechanism to reduce error recovery overhead on video streaming over wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(17): 3049–3067, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Feng:2010:TLF

[224] Taiming Feng, Long Long, Ahmed E. Kamal, and Lu Ruan. Two-link failure protection in WDM mesh networks with p-cycles. Computer Networks (Amsterdam, Netherlands: 1999), 54 (17):3068–3080, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garroppo:2010:SMC

[225] Rosario G. Garroppo, Stefano Giordano, and Luca Tavanti. A survey on multi-constrained optimal path computation: Exact and approximate algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 54(17): 3081–3107, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2010:GBA

[226] Ying Zhang, Damien Fay, Liam Kilmartin, and Andrew W. Moore. A Garch-based adaptive playout delay algorithm for VoIP. Computer Networks (Amsterdam, Netherlands: 1999), 54 (17):3108–3122, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Belzarena:2010:EEQ

[227] Pablo Belzarena and Laura Aspirot. End-to-end quality of service seen by applications: a statistical learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):3123– 3143, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Serral-Gracia:2010:ELM

[228] René Serral-Gracià, Marcelo Yannuzzi, Yann Labit, Philippe Owezarski, and Xavi Masip-Bruin. An efficient and lightweight method for Service Level Agreement assessment. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):3144–3158, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hua:2010:LLS

[229] Cunqing Hua and Rong Zheng. On link-level starvation in dense 802.11 wireless community networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):3159–3172, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lv:2010:RAI

[230] Shaohe Lv, Xiaodong Wang, and Xingming Zhou. On the rate adaptation for IEEE 802.11 wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):3173– 3186, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rao:2010:ABL

[231] Yuan Rao and Ru chuan Wang. Agentbased load balancing routing for LEO satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (17):3187–3195, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2010:EBq

[232] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 54(17):??, December 3, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bochmann:2010:SNH

[233] Gregor v. Bochmann, Dave Rayner, and Colin H. West. Some notes on the history of protocol engineering. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3197– 3209, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2010:CDA

[234] Chao Chen, Zesheng Chen, and Yubin Li. Characterizing and defending against divide-conquer-scanning worms. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3210– 3222, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Oki:2010:FTP

[235] Eiji Oki and Ayako Iwaki. Fine twophase routing over shortest paths with traffic matrix. Computer Networks (Amsterdam, Netherlands: 1999), 54 (18):3223–3231, December 20, 2010.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shue:2010:IIP

[236] Craig A. Shue and Minaxi Gupta. An Internet without the Internet protocol. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3232– 3245, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pezaros:2010:HSB

[237] Dimitrios P. Pezaros, Konstantinos Georgopoulos, and David Hutchison. High-speed, in-band performance measurement instrumentation for next generation IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3246–3263, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Belbekkouche:2010:TAW

[238] Abdeltouab Belbekkouche, Abdelhakim Hafid, Mariam Tagmouti, and Michel Gendreau. Topology-aware wavelength partitioning for DWDM OBS networks: a novel approach for absolute QoS provisioning. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3264–3279, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Labiod:2010:ASI

[239] Houda Labiod, Hai Lin, and Riccardo Nonni. Analytical study of intradomain handover in multiple-mobile-routers-based multihomed NEMO networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3280– 3294, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2010:EET

[240] Jongwook Lee, Saewoong Bahk, and Jin-Ghoo Choi. Energy efficient transmission scheduling for infrastructure sensor nodes in location systems. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3295–3308, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bu:2010:SHF

[241] Tian Bu, Jin Cao, Aiyou Chen, and Patrick P. C. Lee. Sequential hashing: a flexible approach for unveiling significant patterns in high speed networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3309– 3326, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jiang:2010:NPL

[242] Hongbo Jiang, Zihui Ge, Shudong Jin, and Jia Wang. Network prefixlevel traffic profiling: Characterizing, modeling, and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3327–3340, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rahman:2010:RBD

[243] Mustafizur Rahman, Rajiv Ranjan, and Rajkumar Buyya. Reputationbased dependable scheduling of workflow applications in Peer-to-Peer Grids. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3341– 3359, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dimitriou:2010:PES

[244] Stylianos Dimitriou and Vassilis Tsaoussidis. Promoting effective service differentiation with Size-oriented Queue Management. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3360–3372, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Donnet:2010:IRB

[245] Benoit Donnet, Bruno Baynat, and Timur Friedman. Improving retouched Bloom filter for trading off selected false positives against false negatives. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3373– 3387, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jakab:2010:CHC

[246] Loránd Jakab, Albert Cabellos-Aparicio, Thomas Silverston, Marc Solé, Florin Coras, and Jordi Domingo-Pascual. CoreCast: How core/edge separation can help improving interdomain live streaming. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3388–3401, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alouf:2010:FGA

[247] Sara Alouf, Giovanni Neglia, Iacopo Carreras, Daniele Miorandi, and Álvaro Fialho. Fitting genetic algorithms to distributed on-line evolution of network protocols. Computer Networks (Amsterdam, Netherlands:

1999), 54(18):3402–3420, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Charilas:2010:SGT

[248] Dimitris E. Charilas and Athanasios D. Panagopoulos. A survey on game theory applications in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54(18):3421– 3430, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tao:2010:GFR

[249] Shao Tao, A. L. Ananda, and Mun Choon Chan. Greedy face routing with face identification support in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 54 (18):3431–3448, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Donmez:2010:AAD

[250] Mehmet Yunus Donmez, Rabun Kosar, and Cem Ersoy. An analytical approach to the deployment quality of surveillance wireless sensor networks considering the effect of jammers and coverage holes. Computer Networks (Amsterdam, Netherlands: 1999), 54 (18):3449–3466, December 20, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2011:E

[251] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):1–2, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Batista:2011:RSG

[252] Daniel M. Batista and Nelson L. S. da Fonseca. Robust scheduler for grid networks under uncertainties of both application demands and resource availability. Computer Networks (Amsterdam, Netherlands: 1999), 55(1): 3–19, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2011:CCS

[253] Ying Li, Antonis Papachristodoulou, Mung Chiang, and A. Robert Calderbank. Congestion control and its stability in networks with delay sensitive traffic. Computer Networks (Amsterdam, Netherlands: 1999), 55(1): 20–32, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Heikkinen:2011:DSA

[254] T. Heikkinen and A. Hottinen. Distributed subchannel assignment in a two-hop network. Computer Networks (Amsterdam, Netherlands: 1999), 55 (1):33–44, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Qazi:2011:DLF

[255] Ihsan Ayyub Qazi and Taieb Znati. On the design of load factor based congestion control protocols for nextgeneration networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):45–60, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altunay:2011:ORA

[256] Mine Altunay, Sven Leyffer, Jeffrey T. Linderoth, and Zhen Xie. Optimal response to attacks on the open science grid. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):61–73, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alcaraz:2011:DSM

[257] Juan J. Alcaraz, Esteban Egea-López, Javier Vales-Alonso, and Joan García-Haro. Dynamic system model for optimal configuration of mobile RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 55(1): 74–83, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Aitsaadi:2011:APF

[258] Nadjib Aitsaadi, Nadjib Achir, Khaled Boussetta, and Guy Pujolle. Artificial potential field approach in WSN deployment: Cost, QoM, connectivity, and lifetime constraints. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):84–105, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Elias:2011:GTA

[259] Jocelyne Elias, Fabio Martignon, Konstantin Avrachenkov, and Giovanni Neglia. A game theoretic analysis of network design with sociallyaware users. Computer Networks (Amsterdam, Netherlands: 1999), 55(1): 106–118, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Domingo-Ferrer:2011:DNW

[260] Josep Domingo-Ferrer and Úrsula González-Nicolás. Decapitation of networks with and without weights and direction: The economics of iterated attack and defense. Computer Networks (Amsterdam, Netherlands: 1999), 55 (1):119–130, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhou:2011:HSN

[261] Zehua Zhou, Xiaojing Xiang, Xin Wang, and Jianping Pan. A holistic sensor network design for energy conservation and efficient data dissemination. Computer Networks (Amsterdam, Netherlands: 1999), 55(1): 131–146, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2011:EIS

[262] Younghyun Kim, Sangheon Pack, Chung Gu Kang, and Soonjun Park. An enhanced information server for seamless vertical handover in IEEE 802.21 MIH networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):147–158, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Belbekkouche:2011:WMO

[263] Abdeltouab Belbekkouche, Jihene Rezgui, and Abdelhakim Hafid. Wireless mesh and optical burst switching convergence for a novel metropolitan area network architecture. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):159–172, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gelabert:2011:PER

[264] Xavier Gelabert, Oriol Sallent, Jordi Pérez-Romero, and Ramon Agustí. Performance evaluation of radio access selection strategies in constrained multi-access/multi-service wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(1): 173–192, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mello:2011:IAE

[265] Darli A. A. Mello, Helio Waldman, and Gustavo S. Quitério. Interval availability estimation for protected connections in optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):193–204, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhou:2011:PSA

[266] Tao Zhou and Jing Xu. Provable secure authentication protocol with anonymity for roaming service in global mobility networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (1):205–213, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

AlDaoud:2011:RPR

[267] Ashraf Al Daoud, Murat Alanyali, and David Starobinski. Reservation policies for revenue maximization from secondary spectrum access in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(1): 214–224, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2011:EAM

[268] Donggook Kim, Jaesub Kim, and Kyu Ho Park. An event-aware MAC scheduling for energy efficient aggregation in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):225–240, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cho:2011:SST

[269] Kyungmin Cho, Younghyun Ju, Sungjae Jo, Yunseok Rhee, and Junehwa Song. SATI: a scalable and trafficefficient data delivery infrastructure for real-time sensing applications. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):241–263, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guo:2011:DIK

[270] S. Guo, M. Derakhshani, M. H. Falaki, U. Ismail, R. Luk, E. A. Oliver, S. Ur Rahman, A. Seth, M. A. Zaharia, and S. Keshav. Design and implementation of the KioskNet system. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):264–281, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hadjichristofi:2011:RSR

[271] George C. Hadjichristofi, Luiz A. DaSilva, Scott F. Midkiff, Unghee Lee, and Waltemar De Sousa. Routing, security, resource management, and monitoring in ad hoc networks: Implementation and integration. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):282–299, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Serbu:2011:HSO

[272] Sabina Serbu, Pascal Felber, and Peter Kropf. HyPeer: Structured overlay with flexible-choice routing. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):300–313, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ko:2011:DRA

[273] Ren-Song Ko. A distributed routing algorithm for sensor networks derived from macroscopic models. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):314–329, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

How:2011:RQP

[274] Kiam Cheng How, Maode Ma, and Yang Qin. Routing and QoS provisioning in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):330–342, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhao:2011:QOP

[275] Can Zhao and Xiaojun Lin. On the queue-overflow probabilities of a class of distributed scheduling algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 55(1): 343–355, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Antonopoulos:2011:BBC

[276] Christos P. Antonopoulos. On BS/BS coordination towards interference minimization in WiMAX networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):356–369, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2011:EBa

[277] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(1):??, January 7, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2011:P

[278] Eitan Altman, Sajal K. Das, Luciano Lenzini, and Adam Wolisz. Preface. Computer Networks (Amsterdam, Netherlands: 1999), 55(2):371– 373, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gaito:2011:SHF

[279] Sabrina Gaito, Elena Pagani, and Gian Paolo Rossi. Strangers help friends to communicate in opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(2): 374–385, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marsan:2011:EEW

[280] Marco Ajmone Marsan and Michela Meo. Energy efficient wireless Internet access with cooperative cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(2): 386–398, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Psaras:2011:DCS

[281] Ioannis Psaras and Lefteris Mamatas. On Demand Connectivity Sharing: Queuing management and load balancing for User-Provided Networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(2):399–414, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

So-In:2011:VAU

[282] Chakchai So-In, Raj Jain, Subharthi Paul, and Jianli Pan. Virtualization architecture using the ID/Locator split concept for Future Wireless Networks (FWNs). Computer Networks (Amsterdam, Netherlands: 1999), 55(2): 415–430, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dirani:2011:SON

[283] Mariana Dirani and Zwi Altman. Selforganizing networks in next generation radio access networks: Application to fractional power control. Computer Networks (Amsterdam, Netherlands: 1999), 55(2):431–438, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Biczok:2011:IGW

[284] Gergely Biczók, László Toka, András Gulyás, Tuan A. Trinh, and Attila Vidács. Incentivizing the global wireless village. Computer Networks (Amsterdam, Netherlands: 1999), 55(2): 439–456, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gerla:2011:VNF

[285] Mario Gerla and Leonard Kleinrock. Vehicular networks and the future of the mobile Internet. Computer Networks (Amsterdam, Netherlands: 1999), 55(2):457–469, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wong:2011:IAW

[286] Sulan Wong, Eitan Altman, and Julio Rojas-Mora. Internet access: Where law, economy, culture and technology meet. Computer Networks (Amsterdam, Netherlands: 1999), 55(2): 470–479, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2011:CLD

[287] Lijun Chen, Steven H. Low, and John C. Doyle. Cross-layer design in multihop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(2):480–496, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2011:EBb

[288] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(2):??, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2011:DRM

[289] Fei Yang and Isabelle Augé-Blum. Delivery ratio-maximized wakeup scheduling for ultra-low duty-cycled WSNs under real-time constraints. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):497–513, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lorincz:2011:ONM

[290] Josip Lorincz, Antonio Capone, and Dinko Begušić. Optimized network management for energy savings of wireless access networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):514–540, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

LeBlond:2011:PBL

[291] Stevens Le Blond, Arnaud Legout, and Walid Dabbous. Pushing BitTorrent locality to the limit. Computer Networks (Amsterdam, Netherlands: 1999), 55 (3):541–557, February 21, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

DeCicco:2011:SVC

[292] Luca De Cicco, Saverio Mascolo, and Vittorio Palmisano. Skype Video congestion control: an experimental investigation. Computer Networks (Amsterdam, Netherlands: 1999), 55(3): 558–571, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Barrera:2011:SAC

[293] Ivan D. Barrera, Gonzalo R. Arce, and Stephan Bohacek. Statistical approach for congestion control in gateway routers. Computer Networks (Amsterdam, Netherlands: 1999), 55(3): 572–582, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Galvez:2011:MRS

[294] Juan J. Gálvez, Pedro M. Ruiz, and Antonio F. G. Skarmeta. Multipath routing with spatial separation in wireless multi-hop networks without location information. Computer Networks (Amsterdam, Netherlands: 1999), 55 (3):583–599, February 21, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vidal:2011:EEI

[295] Ivan Vidal, Jaime Garcia-Reinoso, Ignacio Soto, and Francisco Valera. Evaluating extensions to IMS session setup for multicast-based many-to-many services. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):600– 621, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kosek-Szott:2011:SAT

[296] Katarzyna Kosek-Szott, Marek Natkaniec, and Andrzej R. Pach. A simple but accurate throughput model for IEEE 802.11 EDCA in saturation and non-saturation conditions. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):622–635, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cheng:2011:IBR

[297] Pei chun Cheng, Beichuan Zhang, Daniel Massey, and Lixia Zhang. Identifying BGP routing table transfers. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):636– 649, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kwon:2011:DTD

[298] Soonmok Kwon, Jae Hoon Ko, Jeongkyu Kim, and Cheeha Kim. Dynamic timeout for data aggregation in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):650–664, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pavon-Marino:2011:MHN

[299] P. Pavon-Marino, B. Garcia-Manrubia, and R. Aparicio-Pardo. Multi-hour network planning based on domination between sets of traffic matrices. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):665–675, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2011:ESC

[300] Xingwei Wang, Weigang Hou, Lei Guo, Jiannong Cao, and Dingde Jiang. Energy saving and cost reduction in multi-granularity green optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):676– 688, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Christophorou:2011:ERR

[301] Christophoros Christophorou, Andreas Pitsillides, and Tomas Lundborg. Enhanced radio resource management algorithms for efficient MBMS service provision in UTRAN. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):689–710, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2011:SWB

[302] Lei Li, Baoxian Zhang, Xiaojun Shen, Jun Zheng, and Zheng Yao. A study on the weak barrier coverage problem in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):711–721, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Leu:2011:RSU

[303] Jenq-Shiou Leu, Cheng-Wei Tsai, and Wei-Hsiang Lin. Resource searching in an unstructured P2P network based on Cloning Random Walker assisted by Dominating Set. Computer Networks (Amsterdam, Netherlands: 1999), 55 (3):722–733, February 21, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cai:2011:ATC

[304] Yan Cai, Patrick P. C. Lee, Weibo Gong, and Don Towsley. Analysis of traffic correlation attacks on router queues. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):734– 747, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2011:COL

[305] Bin Wu, Kwan L. Yeung, Bing Hu, and Pin-Han Ho. M^{2}-CYCLE: an optical layer algorithm for fast link failure detection in all-optical mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):748– 758, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

CastroFernandes:2011:LGK

[306] Natalia Castro Fernandes and Otto Carlos Muniz Bandeira Duarte. A lightweight group-key management protocol for secure ad-hoc-network routing. Computer Networks (Amsterdam, Netherlands: 1999), 55(3): 759–778, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gregori:2011:PCC

[307] Maria Gregori, Ignacio Llatser, Albert Cabellos-Aparicio, and Eduard Alarcón. Physical channel characterization for medium-range nanonetworks using flagellated bacteria. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):779–791, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chrysos:2011:DWS

[308] Nikolaos Chrysos and Manolis Katevenis. Distributed WFQ scheduling converging to weighted max–min fairness. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):792– 806, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2011:NMG

[309] Xingwei Wang, Weigang Hou, Lei Guo, Jiannong Cao, and Dingde Jiang. A new multi-granularity grooming algorithm based on traffic partition in IP over WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):807–821, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Theoleyre:2011:RAM

[310] Fabrice Theoleyre. A route-aware MAC for wireless multihop networks with a convergecast traffic pattern. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):822–837, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2011:SSI

[311] DK Lee, Keon Jang, Changhyun Lee, Gianluca Iannaccone, and Sue Moon. Scalable and systematic Internet-wide path and delay estimation from existing measurements. Computer Networks (Amsterdam, Netherlands: 1999), 55 (3):838–855, February 21, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Peng:2011:EEG

[312] B. Peng and A. H. Kemp. Energyefficient geographic routing in the presence of localization errors. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):856–872, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khosla:2011:PML

[313] Ravish Khosla, Sonia Fahmy, Y. Charlie Hu, and Jennifer Neville. Prediction models for long-term Internet prefix availability. Computer Networks (Amsterdam, Netherlands: 1999), 55 (3):873–889, February 21, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2011:EBc

[314] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(3):??, February 21, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Eggert:2011:SIA

[315] Lars Eggert and Tilman Wolf. Special issue on architectures and protocols for the Future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 55(4):891–892, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sanchez-Loro:2011:CFI

[316] Xavier Sanchez-Loro, José Luis Ferrer, Carles Gomez, Jordi Casademont, and Josep Paradells. Can Future Internet be based on constrained networks design principles? Computer Networks (Amsterdam, Netherlands: 1999), 55 (4):893–909, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Martin:2011:FFF

[317] Denis Martin, Lars Völker, and Martina Zitterbart. A flexible framework for Future Internet design, assessment, and operation. Computer Networks (Amsterdam, Netherlands: 1999), 55 (4):910–918, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Touch:2011:DRU

[318] Joe Touch, Ilia Baldine, Rudra Dutta, Gregory G. Finn, Bryan Ford, Scott Jordan, Dan Massey, Abraham Matta, Christos Papadopoulos, Peter Reiher, and George Rouskas. A Dynamic Recursive Unified Internet Design (DRUID). Computer Networks (Amsterdam, Netherlands: 1999), 55 (4):919–935, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Katsaros:2011:MOA

[319] Konstantinos Katsaros, George Xylomenos, and George C. Polyzos. MultiCache: an overlay architecture for information-centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 55(4):936–947, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Iannone:2011:ILI

[320] Luigi Iannone, Damien Saucez, and Olivier Bonaventure. Implementing the Locator/ID Separation Protocol: Design and experience. Computer Networks (Amsterdam, Netherlands: 1999), 55(4):948–958, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Luo:2011:DDI

[321] Hongbin Luo, Hongke Zhang, and Moshe Zukerman. Decoupling the design of identifier-to-locator mapping services from identifiers. Computer Networks (Amsterdam, Netherlands: 1999), 55(4):959–974, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rajahalme:2011:NBI

[322] Jarno Rajahalme, Mikko Särelä, Kari Visala, and Janne Riihijärvi. On namebased inter-domain routing. Computer Networks (Amsterdam, Netherlands: 1999), 55(4):975–986, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sifalakis:2011:FCF

[323] M. Sifalakis, A. Louca, G. Bouabene, M. Fry, A. Mauthe, and D. Hutchison. Functional composition in future networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(4): 987–998, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mochizuki:2011:BHH

[324] Brent Mochizuki, Firat Kiyak, Eric Keller, and Matthew Caesar. Better by a HAIR: hardware-amenable Internet routing. Computer Networks (Amsterdam, Netherlands: 1999), 55(4): 999–1010, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Houidi:2011:VNP

[325] Ines Houidi, Wajdi Louati, Walid Ben Ameur, and Djamal Zeghlache. Virtual network provisioning across multiple substrate networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(4):1011–1023, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liao:2011:IMS

[326] Jianxin Liao, Jingyu Wang, Tonghong Li, and Xiaomin Zhu. Introducing multipath selection for concurrent multipath transfer in the future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 55(4): 1024–1035, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2011:EBd

[327] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(4):??, March 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Doron:2011:WWF

[328] Ehud Doron and Avishai Wool. WDA: a Web farm Distributed Denial of Service attack attenuator. Computer Networks (Amsterdam, Netherlands: 1999), 55(5):1037–1051, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dey:2011:ENL

[329] Bikash Kumar Dey, D. Manjunath, and Supriyo Chakraborty. Estimating network link characteristics using packet-pair dispersion: a discrete-time queueing theoretic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 55(5):1052–1068, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hwang:2011:DGS

[330] Choonha Hwang, Elmurod Talipov, and Hojung Cha. Distributed geographic service discovery for mobile sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (5):1069–1082, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Carela-Espanol:2011:AIS

[331] Valentín Carela-Español, Pere Barlet-Ros, Albert Cabellos-Aparicio, and Josep Solé-Pareta. Analysis of the impact of sampling on NetFlow traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 55 (5):1083–1099, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pamies-Juarez:2011:TDO

[332] Lluis Pamies-Juarez, Pedro García-López, Marc Sánchez-Artigas, and Blas Herrera. Towards the design of optimal data redundancy schemes for heterogeneous cloud storage infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 55(5): 1100–1113, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zafeiris:2011:OTF

[333] Vassilis E. Zafeiris and E. A. Giakoumakis. Optimized traffic flow assignment in multi-homed, multi-radio mobile hosts. Computer Networks (Amsterdam, Netherlands: 1999), 55(5): 1114–1131, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Merindol:2011:EAE

[334] P. Mérindol, P. François, O. Bonaventure, S. Cateloin, and J.-J. Pansiot. An efficient algorithm to enable path diversity in link state routing networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(5): 1132–1149, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2011:NCW

[335] Pu Wang, Zhi Sun, Mehmet C. Vuran, Mznah A. Al-Rodhaan, Abdullah M. Al-Dhelaan, and Ian F. Akyildiz. On network connectivity of wireless sensor networks for sandstorm monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 55(5): 1150–1157, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dusi:2011:QAG

[336] Maurizio Dusi, Francesco Gringoli, and Luca Salgarelli. Quantifying the accuracy of the ground truth associated with Internet traffic traces. Computer Networks (Amsterdam, Netherlands: 1999), 55(5):1158–1167, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bilogrevic:2011:OOR

[337] Igor Bilogrevic, Mohammad Hossein Manshaei, Maxim Raya, and Jean-Pierre Hubaux. OREN: Optimal revocations in ephemeral networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(5):1168–1180, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akar:2011:MAB

[338] N. Akar and M. A. Toksöz. MPLS automatic bandwidth allocation via adaptive hysteresis. Computer Networks (Amsterdam, Netherlands: 1999), 55 (5):1181–1196, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hossfeld:2011:CBS

[339] Tobias Hoßfeld, Frank Lehrieder, David Hock, Simon Oechsner, Zoran Despotovic, Wolfgang Kellerer, and Maximilian Michel. Characterization of BitTorrent swarms and their distribution in the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 55(5):1197–1215, April 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Walls:2011:LDR

[340] Robert J. Walls, Kush Kothari, and Matthew Wright. Liquid: a detectionresistant covert timing channel based on IPD shaping. Computer Networks (Amsterdam, Netherlands: 1999), 55 (6):1217–1228, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003580.

Hariharan:2011:SND

[341] Srikanth Hariharan, Ness B. Shroff, and Saurabh Bagchi. Secure neighbor discovery through overhearing in static multihop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(6):1229–1241, April 25, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003671.

Zhang:2011:LLA

[342] Guoqiang Zhang, Guoqing Zhang, and Suqi Cheng. LANC: Locality-aware network coding for better P2P traffic localization. Computer Networks (Amsterdam, Netherlands: 1999), 55 (6):1242–1256, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861000366X.

Shpiner:2011:MIC

[343] Alexander Shpiner and Isaac Keslassy. Modeling the interactions of congestion control and switch scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 55(6): 1257–1275, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003646.

Talebi:2011:CAM

[344] Mohammad Sadegh Talebi, Ahmad Khonsari, Amin Mohtasham, and Ali Abbasi. Cost-aware monitoring of network-wide aggregates in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (6):1276–1290, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003634.

Lu:2011:CRB

[345] Jinchang Lu and Maode Ma. Cognitive radio-based framework and selfoptimizing temporal-spectrum block scheduling for QoS provisioning in WiMAX. Computer Networks (Amsterdam, Netherlands: 1999), 55(6): 1291–1309, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003701.

Bikfalvi:2011:PVI

[346] Alex Bikfalvi, Jaime García-Reinoso, Iván Vidal, Francisco Valera, and Arturo Azcorra. P2P vs. IP multicast: Comparing approaches to IPTV streaming based on TV channel popularity. Computer Networks (Amsterdam, Netherlands: 1999), 55(6): 1310–1325, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003877.

Alshammari:2011:CET

[347] Riyad Alshammari and A. Nur Zincir-Heywood. Can encrypted traffic be identified without port numbers, IP addresses and payload inspection? Computer Networks (Amsterdam, Netherlands: 1999), 55(6): 1326–1350, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003695.

Kudo:2011:DSW

[348] Takanori Kudo and Tetsuya Takine. Design of a sliding window scheme for detecting high packet-rate flows via random packet sampling. Computer Networks (Amsterdam, Netherlands: 1999), 55(6):1351–1363, April 25, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003865.

Rothenberg:2011:PBF

[349] Christian Esteve Rothenberg, Carlos Alberto Braz Macapuna, Maurício Ferreira Magalhães, Fábio Luciano Verdi, and Alexander Wiesmaier. In-packet Bloom filters: Design and networking applications. Computer Networks (Amsterdam, Netherlands: 1999), 55 (6):1364–1378, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003725.

Lam:2011:EDT

[350] Chi ho Lam, Wing Cheong Lau, and Onching Yue. Enhancing distributed traffic monitoring via traffic digest splitting. Computer Networks (Amsterdam, Netherlands: 1999), 55(6): 1379–1393, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003889.

Bermolen:2011:AAB

[351] Paola Bermolen, Marco Mellia, Michela Meo, Dario Rossi, and Silvio Valenti. Abacus: Accurate behavioral classification of P2P-TV traffic. Computer Networks (Amsterdam, Netherlands: 1999), 55(6):1394–1411, April 25, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003713.

Zorbas:2011:CCW

[352] Dimitrios Zorbas and Christos Douligeris. Connected coverage in WSNs based on critical targets. Computer Networks (Amsterdam, Netherlands: 1999), 55 (6):1412–1425, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003907.

Arifler:2011:CAD

[353] Dogu Arifler. Capacity analysis of a diffusion-based short-range molecular nano-communication channel. Computer Networks (Amsterdam, Netherlands: 1999), 55(6): 1426–1434, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003919.

Ramos:2011:ECA

[354] J. Ramos, P. M. Santiago del Río, J. Aracil, and J. E. López de Vergara. On the effect of concurrent applications in bandwidth measurement speedometers. Computer Networks (Amsterdam, Netherlands: 1999), 55 (6):1435–1453, April 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003683.

Anonymous:2011:EBe

[355] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(6):??, April 25, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100106X.

Rost:2011:RAN

[356] P. Rost, R. Boutaba, K. Doppler, and A. Gumaste. Recent Advances in Network Convergence. Computer Networks (Amsterdam, Netherlands: 1999), 55 (7):1455–1458, May 16, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100123X.

Stankiewicz:2011:SQA

[357] R. Stankiewicz and A. Jajszczyk. A survey of QoE assurance in converged networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(7): 1459–1473, May 16, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861100051X.

Vidal:2011:TAT

[358] Ivan Vidal, Antonio de la Oliva, Jaime Garcia-Reinoso, and Ignacio Soto. TRIM: an architecture for transparent IMS-based mobility. Computer Networks (Amsterdam, Netherlands: 1999), 55(7):1474–1486, May 16, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000326.

Vucevic:2011:RLJ

[359] Nemanja Vucević, Jordi Pérez-Romero, Oriol Sallent, and Ramon Agustí. Reinforcement learning for joint radio resource management in LTE– UMTS scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 55 (7):1487–1497, May 16, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000545.

Neves:2011:CAM

[360] Pedro Neves, João Soares, Susana Sargento, Hugo Pires, and Francisco Fontes. Context-aware media independent information server for optimized seamless handover procedures. Computer Networks (Amsterdam, Netherlands: 1999), 55(7): 1498–1519, May 16, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000715.

Kang:2011:APH

[361] Joon-Myung Kang, John Strassner, Sin seok Seo, and James Won-Ki Hong. Autonomic personalized handover decisions for mobile services in heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(7):1520–1532, May 16, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000442.

Vulic:2011:VHA

[362] Natasa Vulic, Sonia M. Heemstra de Groot, and Ignas G. M. M. Niemegeers. Vertical handovers among different wireless technologies in a UMTS radio access-based integrated architecture. Computer Networks (Amsterdam, Netherlands: 1999), 55(7): 1533–1548, May 16, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000454.

Jamalipour:2011:RCC

[363] Abbas Jamalipour, Farshad Javadi, and Kumudu S. Munasinghe. Resource competition in a converged heterogeneous networking ecosystem. Computer Networks (Amsterdam, Netherlands: 1999), 55(7):1549–1559, May 16, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000557.

Faigl:2011:ETI

[364] Zoltán Faigl, László Bokor, Pedro Miguel Neves, Khadija Daoud, and Philippe Herbelin. Evaluation of two integrated signalling schemes for the Ultra Flat Architecture using SIP, IEEE 802.21, and HIP/ PMIP protocols. Computer Networks (Amsterdam, Netherlands: 1999), 55 (7):1560–1575, May 16, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000521.

Meddour:2011:RIS

[365] Djamal-Eddine Meddour, Tinku Rasheed, and Yvon Gourhant. On the role of infrastructure sharing for mobile network operators in emerging markets. Computer Networks (Amsterdam, Netherlands: 1999), 55(7): 1576–1591, May 16, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000776.

Ishibashi:2011:QCV

[366] B. Ishibashi, N. Bouabdallah, and R. Boutaba. QoS capacity of virtual wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (7):1592–1613, May 16, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861100034X.

Leroy:2011:SSW

[367] Damien Leroy, Gregory Detal, Julien Cathalo, Mark Manulis, François Koeune, and Olivier Bonaventure. SWISH: Secure WiFi sharing. Computer Networks (Amsterdam, Netherlands: 1999), 55(7):1614–1630, May 16, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000338.

Anonymous:2011:EBf

[368] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(7):??, May 16, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001253.

Benyamina:2011:DRW

[369] D. Benyamina, A. Hafid, M. Gendreau, and J. C. Maureira. On the design of reliable wireless mesh network infrastructure with QoS constraints. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1631–1647, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003658.

Weng:2011:DPP

[370] Ning Weng, Luke Vespa, and Benfano Soewito. Deep packet pre-filtering and finite state encoding for adaptive intrusion detection system. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1648–1661, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003749.

Govindarajan:2011:IDU

[371] M. Govindarajan and R. M. Chandrasekaran. Intrusion detection using neural based hybrid classification methods. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1662–1671, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003750.

Caretti:2011:EDS

[372] M. Caretti, C. Cicconetti, D. Franceschini, L. Lenzini, D. Migliorini, E. Mingozzi, R. Rossi, and D. Sabella. Efficient downlink scheduling with power boosting in mobile IEEE 802.16 networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(8): 1672–1683, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128610003890.

Rossberg:2011:SAC

[373] Michael Rossberg and Guenter Schaefer. A survey on automatic configuration of virtual private networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(8): 1684–1699, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000053.

Roh:2011:ULS

[374] Hee-Tae Roh and Jang-Won Lee. Userlevel satisfaction aware end-to-end rate control in communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(8): 1700–1710, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000041.

Wang:2011:RAJ

[375] Jui Teng Wang. Rate adaptation with joint receive diversity and power control for cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(8): 1711–1718, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100003X.

Chen:2011:ISI

[376] Jen-Jee Chen, Ling Lee, and Yu-Chee Tseng. Integrating SIP and IEEE 802.11e to support handoff and multi-grade QoS for VoIP-over-WLAN applications. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1719–1734, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000090.

Ozdemir:2011:IPH

[377] Suat Ozdemir and Yang Xiao. Integrity protecting hierarchical concealed data aggregation for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1735–1746, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000107.

You:2011:CLO

[378] Lei You, Lianghui Ding, Ping Wu, Zhiwen Pan, Honglin Hu, Mei Song, and Junde Song. Cross-layer optimization of wireless multihop networks with one-hop two-way network coding. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1747–1769, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000211.

Dopico:2011:PAD

[379] Nelson I. Dopico, Álvaro Gutiérrez, and Santiago Zazo. Performance analysis of a delay tolerant application for herd localization. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1770–1783, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100020X.

Jdidi:2011:FLP

[380] Anis Jdidi and Tijani Chahed. Flowlevel performance of proportional fairness with hierarchical modulation in OFDMA-based networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1784–1793, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000223.

Dakkak:2011:ILM

[381] M. Dakkak, A. Nakib, B. Daachi, P. Siarry, and J. Lemoine. Indoor localization method based on RTT and AOA using coordinates clustering. Computer Networks (Amsterdam, Netherlands: 1999), 55(8): 1794–1803, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000296.

Laufer:2011:GBF

[382] Rafael P. Laufer, Pedro B. Velloso, and Otto Carlos M. B. Duarte. A generalized Bloom filter to secure distributed network applications. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1804–1819, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128610003920.

Detti:2011:PST

[383] A. Detti, N. Blefari-Melazzi, I. Habib, and A. Ordine. Per-station throughput fairness in a WLAN hot-spot with TCP traffic. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1820–1833, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000089.

Macedo:2011:FBL

[384] Daniel F. Macedo, Aldri L. dos Santos, José M. Nogueira, and Guy Pujolle.

Fuzzy-based load self-configuration in mobile P2P services. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1834–1848, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000399.

Fanian:2011:HPI

[385] Ali Fanian, Mehdi Berenjkoub, Hossein Saidi, and T. Aaron Gulliver. A high performance and intrinsically secure key establishment protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1849–1863, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000387.

Wen:2011:DLT

[386] Zhihua Wen and Michael Rabinovich. Dynamic landmark triangles: a simple and efficient mechanism for interhost latency estimation. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1864–1879, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000375.

Avrachenkov:2011:OTC

[387] Konstantin Avrachenkov, Alexander Dudin, Valentina Klimenok, Philippe Nain, and Olga Semenova. Optimal threshold control by the robots of Web search engines with obsolescence of documents. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1880–1893, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000351.

Al-Mistarihi:2011:TBD

[388] Mamoun F. Al-Mistarihi, Mohammad Al-Shurman, and Ahmad Qudaimat. Tree based dynamic address autoconfiguration in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(8): 1894–1908, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000417.

Iliofotou:2011:GGB

[389] Marios Iliofotou, Hyun chul Kim, Michalis Faloutsos, Michael Mitzenmacher, Prashanth Pappu, and George Varghese. Graption: a graph-based P2P traffic classification framework for the Internet backbone. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1909–1920, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000430.

Scharf:2011:CEE

[390] Michael Scharf. Comparison of endto-end and network-supported fast startup congestion control schemes. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1921– 1940, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000491.

Yan:2011:AAP

[391] Guanhua Yan, Duc T. Ha, and Stephan Eidenbenz. AntBot: Anti-pollution peer-to-peer botnets. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):1941–1956, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000533.

Viswanathan:2011:ECP

[392] Ramesh Viswanathan, Krishan K. Sabnani, Robert J. Holt, and Arun N. Netravali. Expected convergence properties of BGP. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1957–1981, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000508.

Yigitel:2011:QAM

[393] M. Aykut Yigitel, Ozlem Durmaz Incel, and Cem Ersoy. QoS-aware MAC protocols for wireless sensor networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):1982–2004, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000703.

Chan:2011:UUN

[394] Tony K. C. Chan, Yiu-Wing Leung, and Gaoxi Xiao. Upgrading unicast nodes to multicast-capable nodes in alloptical networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (8):2005–2021, June 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000740.

Anonymous:2011:EBg

[395] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(8):??, June 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001447.

Lee:2011:BNO

[396] HyunYong Lee, Akihiro Nakao, and JongWon Kim. BiCo: Network operator-friendly P2P traffic control through bilateral cooperation with peers. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2023–2034, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000739.

Alessandria:2011:IAN

[397] Eugenio Alessandria, Massimo Gallo, Emilio Leonardi, Marco Mellia, and Michela Meo. Impact of adverse network conditions on P2P-TV systems: Experimental evidence. Computer Networks (Amsterdam, Netherlands: 1999), 55(9):2035–2050, June 23, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000727.

Herreria-Alonso:2011:OPS

[398] Sergio Herrería-Alonso, Miguel Rodríguez-Pérez, Manuel Fernández-Veiga, and Cándido López-García. Opportunistic power saving algorithms for Ethernet devices. Computer Networks (Amsterdam, Netherlands: 1999), 55 (9):2051–2064, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000764.

Thouin:2011:LSP

[399] Frederic Thouin, Mark Coates, and Michael Rabbat. Large scale probabilistic available bandwidth estimation. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2065–2078, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000752.

Iqbal:2011:DNC

[400] Hammad Iqbal and Taieb Znati. On the design of network control and management plane. Computer Networks (Amsterdam, Netherlands: 1999), 55 (9):2079–2091, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000405.

Ha:2011:TEN

[401] Sangtae Ha and Injong Rhee. Taming the elephants: New TCP slow start. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2092–2110, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000363.

Garcia-Dorado:2011:CBH

[402] José Luis García-Dorado, José Alberto Hernández, Javier Aracil, Jorge E. López de Vergara, and Sergio Lopez-Buedo. Characterization of the busyhour traffic of IP networks based on their intrinsic features. Computer Networks (Amsterdam, Netherlands: 1999), 55(9):2111–2125, June 23, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000806.

Tan:2011:NLO

[403] Chong Tan, Junni Zou, Min Wang, and Ruifeng Zhang. Network lifetime optimization for wireless video sensor networks with network coding/ARQ hybrid adaptive error-control scheme. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2126–2137, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100079X.

Reaz:2011:CED

[404] Abu (Sayeem) Reaz, Vishwanath Ramamurthi, Massimo Tornatore, Suman Sarkar, Dipak Ghosal, and Biswanath Mukherjee. Cost-efficient design for higher capacity hybrid wirelessoptical broadband access network (WOBAN). Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2138–2149, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000788.

Li:2011:LAM

[405] Fangmin Li, Yilin Fang, Fei Hu, and Xinhua Liu. Load-aware multicast routing metrics in multi-radio multi-channel wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2150–2167, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000934.

Yoon:2011:EEO

[406] Sung-Guk Yoon, Changhee Joo, and Saewoong Bahk. Energy-efficient opportunistic scheduling schemes in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (9):2168–2175, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000922.

Jeng:2011:RTB

[407] Andy An-Kai Jeng, Rong-Hong Jan, Chi-Yu Li, and Chien Chen. Releasetime-based multi-channel MAC protocol for wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(9):2176– 2195, June 23, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000909.

Zhang:2011:UDC

[408] Yueping Zhang, Ao-Jan Su, and Guofei Jiang. Understanding data center network architectures in virtualized environments: a view from multitier applications. Computer Networks (Amsterdam, Netherlands: 1999), 55 (9):2196–2208, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000946.

Ke:2011:CSG

[409] Wei-Chieh Ke, Bing-Hong Liu, and Ming-Jer Tsai. The critical-square-grid coverage problem in wireless sensor networks is NP-Complete. Computer Networks (Amsterdam, Netherlands: 1999), 55(9):2209–2220, June 23, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000971.

Ahmadinejad:2011:HMC

[410] Seyed Hossein Ahmadinejad, Saeed Jalili, and Mahdi Abadi. A hybrid model for correlating alerts of known and unknown attack scenarios and updating attack graphs. Computer Networks (Amsterdam, Netherlands: 1999), 55(9):2221–2240, June 23, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000983.

Ramamurthi:2011:CCF

[411] Vishwanath Ramamurthi, Abu (Sayeem) Reaz, Dipak Ghosal, Sudhir Dixit, and Biswanath Mukherjee. Channel, capacity, and flow assignment in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(9):2241–2258, June 23, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001009.

Chen:2011:FJJ

[412] Lin Chen and Jean Leneutre. Fight jamming with jamming — a game theoretic analysis of jamming attack in wireless networks and defense strategy. Computer Networks (Amsterdam, Netherlands: 1999), 55 (9):2259–2270, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611000995.

Kist:2011:DTS

[413] Alexander A. Kist and Abdelnour Aldraho. Dynamic topologies for sustainable and energy efficient traffic routing. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2271–2288, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001010.

Belzarena:2011:ORB

[414] Pablo Belzarena, Fernando Paganini, and Andrés Ferragut. Optimizing revenue for bandwidth auctions over networks with time reservations. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2289–2302, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001022.

Estepa:2011:POM

[415] Rafael Estepa, Antonio Estepa, and Thiago Cupertino. A productivityoriented methodology for local area network design in industrial environments. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2303–2314, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001046.

Paillassa:2011:NAD

[416] Beatrice Paillassa, Cholatip Yawut, and Riadh Dhaou. Network awareness and dynamic routing: The ad hoc network case. Computer Networks (Amsterdam, Netherlands: 1999), 55 (9):2315–2328, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001034.

Yuksel:2011:CLF

[417] M. Yuksel, K. K. Ramakrishnan, R. D. Doverspike, R. K. Sinha, G. Li, K. N. Oikonomou, and D. Wang. Crosslayer failure restoration of IP multicast with applications to IPTV. Computer Networks (Amsterdam, Netherlands: 1999), 55(9):2329–2351, June 23, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001186.

Boc:2011:PHG

[418] Mathias Boc, Anne Fladenmuller, Marcelo Dias de Amorim, Laura Galluccio, and Sergio Palazzo. Price:

Hybrid geographic and co-based forwarding in delay-tolerant networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(9): 2352–2360, June 23, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001174.

Anonymous:2011:EBh

[419] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(9):??, June 23, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001678.

Carl:2011:MPC

[420] Glenn Carl and George Kesidis. Modeling a policy-capable path-vector routing protocol using Jacobi iteration over a path algebra. Computer Networks (Amsterdam, Netherlands: 1999), 55 (10):2361–2379, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001162.

Diana:2011:EVM

[421] Rémi Diana and Emmanuel Lochin. ECN verbose mode: a statistical method for network path congestion estimation. Computer Networks (Amsterdam, Netherlands: 1999), 55(10): 2380–2391, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001216.

Dwekat:2011:PFQ

[422] Z. Dwekat and G. N. Rouskas. A practical fair queuing scheduler: Simplification through quantization. Computer Networks (Amsterdam, Netherlands: 1999), 55(10): 2392–2406, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001320.

Lee:2011:CMR

[423] I-Ta Lee, Guann-Long Chiou, and Shun-Ren Yang. A cooperative multicast routing protocol for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (10):2407–2424, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001332.

Chi:2011:FON

[424] Kaikai Chi, Xiaohong Jiang, Baoliu Ye, and Yanjun Li. Flow-oriented network coding architecture for multihop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (10):2425–2442, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001356.

Li:2011:OBE

[425] Ming Li, Kai Zeng, and Wenjing Lou. Opportunistic broadcast of eventdriven warning messages in Vehicular Ad Hoc Networks with lossy links. Computer Networks (Amsterdam, Netherlands: 1999), 55(10):

2443–2464, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001344.

Safa:2011:PBT

[426] Haidar Safa and Farah Abu Shahla. A policy-based trust-aware adaptive monitoring scheme to enhance WiMax QoS. Computer Networks (Amsterdam, Netherlands: 1999), 55(10): 2465–2480, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001368.

Chen:2011:SSL

[427] Jiming Chen, Chengqun Wang, Youxian Sun, and Xuemin (Sherman) Shen. Semi-supervised Laplacian regularized least squares algorithm for localization in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(10): 2481–2491, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001393.

Deng:2011:FTI

[428] Wenping Deng, Merkouris Karaliopoulos, Wolfgang Mühlbauer, Peidong Zhu, Xicheng Lu, and Bernhard Plattner. k-Fault tolerance of the Internet AS graph. Computer Networks (Amsterdam, Netherlands: 1999), 55 (10):2492–2503, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001381.

Amoroso:2011:GRO

[429] Alessandro Amoroso, Gustavo Marfia, and Marco Roccetti. Going realistic and optimal: a distributed multihop broadcast algorithm for vehicular safety. Computer Networks (Amsterdam, Netherlands: 1999), 55(10): 2504–2519, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100140X.

Jin:2011:FSS

[430] Sunggeun Jin, Munhwan Choi, Lei Wang, and Sunghyun Choi. Fast scanning schemes for IEEE 802.11 WLANs in virtual AP environments. Computer Networks (Amsterdam, Netherlands: 1999), 55(10): 2520–2533, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100137X.

Zeng:2011:SRK

[431] Rongfei Zeng, Yixin Jiang, Chuang Lin, Yanfei Fan, and Xuemin (Sherman) Shen. A scalable and robust key pre-distribution scheme with network coding for sensor data storage. Computer Networks (Amsterdam, Netherlands: 1999), 55(10): 2534–2544, July 14, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001228.

Anonymous:2011:EBi

[432] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(10):??, July 14, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002088.

Min:2011:MAT

[433] Geyong Min, Jia Hu, and Mike E. Woodward. Modeling and analysis of TXOP differentiation in infrastructurebased WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 55 (11):2545–2557, August 4, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001526.

ODonovan:2011:SAF

[434] Tony O’Donovan, Utz Roedig, Jonathan Benson, and Cormac J. Sreenan. Self-adaptive framelet-based communication for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(11): 2558–2575, August 4, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001538.

Shu:2011:NAM

[435] Feng Shu and Taka Sakurai. A new analytical model for the IEEE 802.15.4 CSMA–CA protocol. Computer Networks (Amsterdam, Netherlands: 1999), 55(11):2576–2591, August 4, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611001587.

Wang:2011:PFM

[436] Shie-Yuan Wang, Chih-Che Lin, Wei-Jyun Hong, and Kuang-Che Liu. On the performances of forwarding multihop unicast traffic in WBSS-based 802.11(p)/1609 networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(11): 2592–2607, August 4, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001575.

Xing:2011:TRM

[437] Fei Xing and Wenye Wang. Toward robust multi-hop data forwarding in large scale wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(11): 2608–2621, August 4, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001514.

Wu:2011:SWN

[438] Jing Wu. A survey of WDM network reconfiguration: Strategies and triggering methods. Computer Networks (Amsterdam, Netherlands: 1999), 55 (11):2622–2645, August 4, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001794.

Murawski:2011:UDS

[439] R. Murawski and E. Ekici. Utilizing dynamic spectrum leasing for cognitive radios in 802.11-based wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(11): 2646–2657, August 4, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001782.

Ferragut:2011:RAM

[440] Andrés Ferragut and Fernando Paganini. Resource allocation over multirate wireless networks: a Network Utility Maximization perspective. Computer Networks (Amsterdam, Netherlands: 1999), 55(11): 2658–2674, August 4, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001824.

Anonymous:2011:EBj

[441] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(11):??, August 4, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611002222.

Zhao:2011:OBM

[442] Weiyi Zhao and Jiang Xie. OPNETbased modeling and simulation study on handoffs in Internet-based infrastructure wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(12):2675– 2688, August 25, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001502.

Islam:2011:MDA

[443] Salekul Islam and Jean-Charles Grégoire. Multi-domain authentication for IMS services. Computer Networks (Amsterdam, Netherlands: 1999), 55(12): 2689–2704, August 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001423.

St-Hilaire:2011:CDM

[444] Marc St-Hilaire and Shangyun Liu. Comparison of different meta-heuristics to solve the global planning problem of UMTS networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (12):2705–2716, August 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001630.

Qiu:2011:FST

[445] Jian Qiu, Yong Liu, Gurusamy Mohan, and Kee Chaing Chua. Fast spanning tree reconnection mechanism for resilient Metro Ethernet networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(12): 2717–2729, August 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001551.

Seibert:2011:DSD

[446] Jeff Seibert, Xin Sun, Cristina Nita-Rotaru, and Sanjay Rao. A design for securing data delivery in mesh-based peer-to-peer streaming. Computer Networks (Amsterdam, Netherlands: 1999), 55(12):

2730–2745, August 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001836.

Oh:2011:EMP

[447] Hyung Rai Oh, Dapeng Oliver Wu, and Hwangjun Song. An effective meshpull-based P2P video streaming system using Fountain codes with variable symbol sizes. Computer Networks (Amsterdam, Netherlands: 1999), 55 (12):2746–2759, August 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100154X.

Zi:2011:ETR

[448] Xiaochao Zi, Lihong Yao, Xinghao Jiang, Li Pan, and Jianhua Li. Evaluating the transmission rate of covert timing channels in a network. Computer Networks (Amsterdam, Netherlands: 1999), 55(12): 2760–2771, August 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100185X.

deSanti:2011:DOA

[449] Juliana de Santi and Nelson L. S. da Fonseca. Design of optimal Active Queue Management controllers for HSTCP in large bandwidth-delay product networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (12):2772–2790, August 25, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001964.

Toril:2011:NPM

[450] Matías Toril, Volker Wille, S. Luna-Ramírez, and K. Jarvinen. Network performance model for location area re-planning in GERAN. Computer Networks (Amsterdam, Netherlands: 1999), 55(12):2791– 2802, August 25, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001885.

Anonymous:2011:EBk

[451] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(12):??, August 25, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611002404.

Aioffi:2011:BMD

[452] Wagner Moro Aioffi, Cristiano Arbex Valle, Geraldo R. Mateus, and Alexandre Salles da Cunha. Balancing message delivery latency and network lifetime through an integrated model for clustering and routing in Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (13):2803–2820, September 15, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001952.

Chen:2011:NMA

[453] Ying Chen, Arunita Jaekel, and Ataul Bari. A new model for allocating resources to scheduled lightpath demands. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2821–2837, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001897.

Camps-Mur:2011:DEE

[454] Daniel Camps-Mur, Xavier Pérez-Costa, and Sebastià Sallent-Ribes. Designing energy efficient access points with Wi-Fi Direct. Computer Networks (Amsterdam, Netherlands: 1999), 55 (13):2838–2855, September 15, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002179.

Sahoo:2011:LMC

[455] Prasan Kumar Sahoo and Jang-Ping Sheu. Limited mobility coverage and connectivity maintenance protocols for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2856–2872, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002167.

Ciccarese:2011:ACP

[456] Giovanni Ciccarese, Mario De Blasi, Pierluigi Marra, Cosimo Palazzo, and Luigi Patrono. An algorithm for controlling packet size in IEEE 802.16e networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2873–2885, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002155.

Darehshoorzadeh:2011:MCC

[457] Amir Darehshoorzadeh, Llorenç Cerdà-Alabern, and Vicent Pla. Modeling and comparison of candidate selection algorithms in opportunistic routing. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2886–2898, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002064.

Jacobsson:2011:EIO

[458] Martin Jacobsson, Cheng Guo, and Ignas Niemegeers. An experimental investigation of optimized flooding protocols using a wireless sensor network testbed. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2899–2913, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002003.

Nechaev:2011:CCI

[459] B. Nechaev, D. Korzun, and A. Gurtov. CR-Chord: Improving lookup availability in the presence of malicious DHT nodes. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2914–2928, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001927.

Lv:2011:LSW

[460] Shaohe Lv, Weihua Zhuang, Xiaodong Wang, and Xingming Zhou. Link scheduling in wireless networks with successive interference cancellation. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2929–2941, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002039.

Korcak:2011:EIH

[461] Ömer Korçak and Fatih Alagöz. Efficient integration of HAPs and mobile satellites via free-space optical links. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2942–2953, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001940.

Wickboldt:2011:FRA

[462] Juliano Araujo Wickboldt, Luís Armando Bianchin, Roben Castagna Lunardi, Lisandro Zambenedetti Granville, Luciano Paschoal Gaspary, and Claudio Bartolini. A framework for risk assessment based on analysis of historical information of workflow execution in IT systems. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2954–2975, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002015.

Ge:2011:PET

[463] Fei Ge, Sammy Chan, Lachlan L. H. Andrew, Fan Li, Liansheng Tan, and Moshe Zukerman. Performance effects of two-way FAST TCP. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 2976–2984, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002180.

Igartua:2011:GTM

[464] Mónica Aguilar Igartua, Luis J. de la Cruz Llopis, Víctor Carrascal Frías, and Emilio Sanvicente Gargallo. A game-theoretic multipath routing for video-streaming services over Mobile Ad Hoc Networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (13):2985–3000, September 15, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002027.

Banerjee:2011:SQM

[465] Anirban Banerjee, Md Sazzadur Rahman, and Michalis Faloutsos. SUT: Quantifying and mitigating URL typosquatting. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 3001–3014, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001976.

Gotzhein:2011:BBS

[466] R. Gotzhein and T. Kuhn. Black Burst Synchronization (BBS) — a protocol for deterministic tick and time synchronization in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 3015–3031, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001812.

Boukerche:2011:RPA

[467] Azzedine Boukerche, Begumhan Turgut, Nevin Aydin, Mohammad Z. Ahmad, Ladislau Bölöni, and Damla Turgut. Routing protocols in ad hoc networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 55(13): 3032–3080, September 15, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001654.

Incel:2011:SMC

[468] Ozlem Durmaz Incel. A survey on multi-channel communication in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(13):3081–3099, September 15, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001873.

Anonymous:2011:EBl

[469] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(13):??, September 15, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611002647.

Herrtwich:2011:ESI

[470] R. G. Herrtwich and I. Radusch. Editorial for the special issue: Deploying vehicle-2-x communication. Computer Networks (Amsterdam, Netherlands: 1999), 55(14):3101– 3102, October 6, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002386.

Weiss:2011:VCE

[471] Christian Weiß. V2X communication in Europe — From research projects towards standardization and field testing of vehicle communication technology. Computer Networks (Amsterdam, Netherlands: 1999), 55(14): 3103–3119, October 6, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001198.

Misener:2011:DVX

[472] James A. Misener, Subir Biswas, and Greg Larson. Development of V-to-X systems in North America: The promise, the pitfalls and the prognosis. Computer Networks (Amsterdam, Netherlands: 1999), 55(14): 3120–3133, October 6, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001319.

Fukushima:2011:LTV

[473] Masao Fukushima. The latest trend of V2X driver assistance systems in Japan. Computer Networks (Amsterdam, Netherlands: 1999), 55(14):

3134–3141, October 6, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001150.

Vales-Alonso:2011:OCR

[474] J. Vales-Alonso, F. Vicente-Carrasco, and J. J. Alcaraz. Optimal configuration of roadside beacons in V2I communications. Computer Networks (Amsterdam, Netherlands: 1999), 55 (14):3142–3153, October 6, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001617.

Roglinger:2011:MTI

[475] Sebastian Röglinger. A methodology for testing intersection related Vehicle-2-X applications. Computer Networks (Amsterdam, Netherlands: 1999), 55 (14):3154–3168, October 6, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002192.

Schweiger:2011:ECI

[476] Benno Schweiger, Christian Raubitschek, Bernard Bäker, and Johann Schlichter. Elisa^{T M} — Car to infrastructure communication in the field. Computer Networks (Amsterdam, Netherlands: 1999), 55(14):3169–3178, October 6, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611001915.

Stanica:2011:SVA

[477] Razvan Stanica, Emmanuel Chaput, and André-Luc Beylot. Simulation of vehicular ad-hoc networks: Challenges, review of tools and recommendations. Computer Networks (Amsterdam, Netherlands: 1999), 55(14): 3179–3188, October 6, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001629.

Schunemann:2011:VSR

[478] Björn Schünemann. V2X simulation runtime infrastructure VSim-RTI: an assessment tool to design smart traffic management systems. Computer Networks (Amsterdam, Netherlands: 1999), 55(14): 3189–3198, October 6, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001605.

Troncoso:2011:DAA

[479] Carmela Troncoso, Enrique Costa-Montenegro, Claudia Diaz, and Stefan Schiffner. On the difficulty of achieving anonymity for Vehicle-2-X communication. Computer Networks (Amsterdam, Netherlands: 1999), 55(14): 3199–3210, October 6, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001599.

Anonymous:2011:EBm

[480] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(14):??, October 6, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611002878.

Trujillo-Rasua:2011:EPC

[481] Rolando Trujillo-Rasua and Agusti Solanas. Efficient probabilistic communication protocol for the private identification of RFID tags by means of collaborative readers. Computer Networks (Amsterdam, Netherlands: 1999), 55 (15):3211–3223, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001800.

Ruan:2011:WSD

[482] Zheng Ruan, Edith C.-H. Ngai, and Jiangchuan Liu. Wireless sensor deployment for collaborative sensing with mobile phones. Computer Networks (Amsterdam, Netherlands: 1999), 55 (15):3224–3245, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002295.

Zhang:2011:AGK

[483] Lei Zhang, Qianhong Wu, Bo Qin, Josep Domingo-Ferrer, and Úrsula González-Nicolás. Asymmetric group key agreement protocol for open networks and its application to broadcast encryption. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3246–3255, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002283.

Vazifehdan:2011:EAR

[484] Javad Vazifehdan, R. Venkatesha Prasad, Ertan Onur, and Ignas Niemegeers. Energy-aware routing algorithms for wireless ad hoc networks with heterogeneous power supplies. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3256–3274, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002271.

Wang:2011:FPB

[485] Kuochen Wang, Chun-Ying Huang, Shang-Jyh Lin, and Ying-Dar Lin. A fuzzy pattern-based filtering algorithm for botnet detection. Computer Networks (Amsterdam, Netherlands: 1999), 55(15):3275–3286, October 27, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611002040.

Xue:2011:SSD

[486] Yuyan Xue, Byrav Ramamurthy, and Mehmet C. Vuran. SDRCS: a servicedifferentiated real-time communication scheme for event sensing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (15):3287–3302, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002301.

Esposito:2011:ISS

[487] Flavio Esposito, Ibrahim Matta, Debajyoti Bera, and Pietro Michiardi. On the impact of seed scheduling in peerto-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (15):3303–3317, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002325.

Li:2011:IUS

[488] Dan Li, Jianping Wu, Yong Cui, Jiangchuan Liu, and Ke Xu. Impact of user selfishness in construction action on the streaming quality of overlay multicast. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3318–3331, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002313.

Alkubeily:2011:NAL

[489] M. Alkubeily, H. Bettahar, and A. Bouabdallah. A new Application-Level Multicast technique for stable, robust and efficient overlay tree construction. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3332–3350, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001939.

Cano:2011:LEO

[490] C. Cano, B. Bellalta, A. Sfairopoulou, and M. Oliver. Low energy operation in WSNs: a survey of preamble sampling MAC protocols. Computer Networks (Amsterdam, Netherlands: 1999), 55 (15):3351–3363, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002349.

Dong:2011:EMR

[491] Ziqian Dong, Santhanakrishnan Anand, and Rajarathnam Chandramouli. Estimation of missing RTTs in computer networks: Matrix completion vs compressed sensing. Computer Networks (Amsterdam, Netherlands: 1999), 55 (15):3364–3375, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002490.

Yang:2011:FME

[492] Liu Yang, Rezwana Karim, Vinod Ganapathy, and Randy Smith. Fast, memory-efficient regular expression matching with NFA-OBDDs. Computer Networks (Amsterdam, Netherlands: 1999), 55(15):3376–3393, October 27, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611002489.

Leon:2011:MRU

[493] Xavier León, Tuan Anh Trinh, and Leandro Navarro. Modeling resource usage in planetary-scale shared infrastructures: PlanetLab’s case study. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3394–3407, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002532.

Talipov:2011:CCB

[494] Elmurod Talipov and Hojung Cha. Communication capacity-based message exchange mechanism for delaytolerant networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (15):3408–3422, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002520.

Balasubramaniam:2011:BIF

[495] Sasitharan Balasubramaniam, Dmitri Botvich, Raymond Carroll, Julien Mineraud, Tadashi Nakano, Tatsuya Suda, and William Donnelly. Biologically inspired future service environment. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3423–3440, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002507.

Sen:2011:ECT

[496] Sevil Sen and John A. Clark. Evolutionary computation techniques for intrusion detection in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3441–3457, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002477.

Fay:2011:DGT

[497] Damien Fay, Hamed Haddadi, Steve Uhlig, Liam Kilmartin, Andrew W. Moore, Jérôme Kunegis, and Marios Iliofotou. Discriminating graphs through spectral projections. Computer Networks (Amsterdam, Netherlands: 1999), 55(15):3458–3468, October 27, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611002374.

Kim:2011:RRA

[498] Hyoungshick Kim and Jaehoon Jeong. RAD: Recipient-anonymous data delivery based on public routing proxies. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3469–3484, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002568.

Tellenbach:2011:ANA

[499] Bernhard Tellenbach, Martin Burkhart, Dominik Schatzmann, David Gugelmann, and Didier Sornette. Accurate network anomaly classification with generalized entropy metrics. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3485–3502, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002556.

Nikolova:2011:BDR

[500] Dessislava Nikolova and Chris Blondia. Bonded deficit round robin scheduling for multi-channel networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3503–3516, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002544.

Ciccarelli:2011:CPP

[501] Gianluca Ciccarelli and Renato Lo Cigno. Collusion in peer-to-peer systems. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3517–3532, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002581.

Jiang:2011:JTF

[502] Dingde Jiang, Zhengzheng Xu, Zhenhua Chen, Yang Han, and Hongwei Xu. Joint time-frequency sparse estimation of large-scale network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3533–3547, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861100257X.

Wang:2011:MTS

[503] Cheng Wang, Changjun Jiang, Xiang-Yang Li, and Yunhao Liu. On multicast throughput scaling of hybrid wireless networks with general node density. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3548–3561, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002593.

Aksu:2011:SLM

[504] Hidayet Aksu, Demet Aksoy, and Ibrahim Korpeoglu. A study of localization metrics: Evaluation of position errors in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3562–3577, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002350.

Migliorini:2011:PEH

[505] Daniele Migliorini, Enzo Mingozzi, and Carlo Vallati. Performance evaluation of H.264/SVC video streaming over mobile WiMAX. Computer Networks (Amsterdam, Netherlands: 1999), 55 (15):3578–3591, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002623.

Hsu:2011:ORR

[506] Che-Jung Hsu, Huey-Ing Liu, and Winston K. G. Seah. Opportunistic routing — a review and the challenges ahead. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3592–3603, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002337.

Wang:2011:SCA

[507] Wenye Wang, Yi Xu, and Mohit Khanna. A survey on the communication architectures in smart grid. Computer Networks (Amsterdam, Netherlands: 1999), 55(15): 3604–3629, October 27, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861100260X.

Anonymous:2011:EBn

[508] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(15):??, October 27, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611003082.

Tugcu:2011:GE

[509] Tuna Tugcu. Guest Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 55(16):3631–3633, November 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002982.

Font-Bach:2011:RTM

[510] O. Font-Bach, N. Bartzoudis, A. Pascual-Iserte, and D. López Bueno. A realtime MIMO-OFDM mobile WiMAX receiver: Architecture, design and FPGA implementation. Computer Networks (Amsterdam, Netherlands: 1999), 55(16):3634–3647, November 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611000910.

Cho:2011:PLS

[511] HanGyu Cho, Taeyoung Kim, Yu-Tao Hsieh, and Jong-Kae Fwu. Physical layer structure of next generation mobile WiMAX technology. Computer Networks (Amsterdam, Netherlands: 1999), 55(16):

3648–3658, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001988.

Kunst:2011:IHE

[512] Rafael Kunst, Cristiano Bonato Both, Lisandro Zambenedetti Granville, and Juergen Rochol. On the impact of hybrid errors on mobile WiMAX networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3659–3671, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001769.

Kaarthick:2011:SAM

[513] Balakrishnan Kaarthick, N. Nagarajan, E. Raguvaran, and G. Saimethun. Subchannel allocation and mapping algorithms for improving the QoS of VoIP traffic in IEEE 802.16e networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3672–3679, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002362.

Cicconetti:2011:FEA

[514] C. Cicconetti, L. Lenzini, A. Lodi, S. Martello, E. Mingozzi, and M. Monaci. A fast and efficient algorithm to exploit multiuser diversity in IEEE 802.16 BandAMC. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3680–3693, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002854.

Vejarano:2011:SRA

[515] G. Vejarano, D. Wang, and J. Mc-Nair. Stability region adaptation using transmission power control for transport capacity optimization in IEEE 802.16 wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3694–3704, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001770.

Liang:2011:EEU

[516] Jia-Ming Liang, You-Chiun Wang, Jen-Jee Chen, Jui-Hsiang Liu, and Yu-Chee Tseng. Energy-efficient uplink resource allocation for IEEE 802.16j transparent-relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(16):3705– 3720, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002209.

Chen:2011:PFS

[517] Jen-Jee Chen, Shih-Lin Wu, Shiou-Wen Wang, and Yu-Chee Tseng. Perflow sleep scheduling for power management in IEEE 802.16 wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3721–3733, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100096X.

Jin:2011:PAP

[518] Shunfu Jin and Wuyi Yue. Performance analysis for power saving class type III of IEEE 802.16 in WiMAX. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3734–3743, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100199X.

Kim:2011:JPA

[519] Ronny Yongho Kim, Ritesh Kumar Kalle, and Debabrata Das. Joint paging area and location update optimization for IEEE 802.16m idle mode. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3744–3758, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001903.

Becvar:2011:IHP

[520] Zdenek Becvar, Pavel Mach, and Boris Simak. Improvement of handover prediction in mobile WiMAX by using two thresholds. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3759–3773, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001411.

Jin:2011:ASM

[521] Sunggeun Jin, Xi Chen, Daji Qiao, and Sunghyun Choi. Adaptive sleep mode management in IEEE 802.16m wireless metropolitan area networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3774–3783, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000958.

Fernandez-Carames:2011:MWV

[522] T. M. Fernández-Caramés, M. González-López, and L. Castedo. Mobile WiMAX for vehicular applications: Performance evaluation and comparison against IEEE 802.11p/a. Computer Networks (Amsterdam, Netherlands: 1999), 55(16):3784– 3795, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611000892.

Chuang:2011:LMA

[523] Ming-Chin Chuang and Jeng-Farn Lee. A lightweight mutual authentication mechanism for network mobility in IEEE 802.16e wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(16): 3796–3809, November 10, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002052.

Anonymous:2011:EBo

[524] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(16):??, November 10, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611003239.

Wang:2011:ESD

[525] Yufeng Wang, Akihiro Nakao, Athanasios V. Vasilakos, and Jianhua Ma. On the effectiveness of service differentiation based resourceprovision incentive mechanisms in dynamic and autonomous P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(17): 3811–3831, December 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002611.

Elias:2011:NCS

[526] Jocelyne Elias, Fabio Martignon, Antonio Capone, and Eitan Altman. Noncooperative spectrum access in cognitive radio networks: a game theoretical model. Computer Networks (Amsterdam, Netherlands: 1999), 55(17): 3832–3846, December 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002829.

Sou:2011:SPM

[527] Sok-Ian Sou and Chuan-Sheng Lin. SPR proxy mechanism for 3GPP Policy and Charging Control System. Computer Networks (Amsterdam, Netherlands: 1999), 55(17): 3847–3862, December 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002957.

Galmes:2011:RDG

[528] Sebastià Galmés and Ramon Puigjaner. Randomized Data-Gathering protocol for time-driven sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(17): 3863–3885, December 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002970.

Kao:2011:PER

[529] Chien-Chi Kao, Shun-Ren Yang, and Tung-Lin Tsai. Performance enhancement of repacking and borrowing mechanisms for IEEE 802.16j multihop resource scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 55 (17):3886–3903, December 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002842.

Chan:2011:EDA

[530] Aldar C.-F. Chan. Efficient defence against misbehaving TCP receiver DoS attacks. Computer Networks (Amsterdam, Netherlands: 1999), 55(17): 3904–3914, December 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003008.

Lee:2011:SBC

[531] Seongjin Lee, Jongwoo Song, Soohan Ahn, and Youjip Won. Sessionbased classification of Internet applications in 3G wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 55(17):3915– 3931, December 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003069.

Yu:2011:TSS

[532] InKwan Yu and Richard Newman. TCP slow start with fair share of bandwidth. Computer Networks (Amsterdam, Netherlands: 1999), 55(17): 3932–3946, December 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003057.

Page:2011:PEP

[533] Christopher Page, Emad Guirguis, and Mina Guirguis. Performance evaluation of path splicing on the GÉANT and the Sprint networks. Computer Networks (Amsterdam, Netherlands: 1999), 55 (17):3947–3958, December 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003045.

Londono:2011:TCC

[534] Jorge Londoño, Azer Bestavros, and Nikolaos Laoutaris. Trade & Cap: a customer-managed, market-based system for trading bandwidth allowances at a shared link. Computer Networks (Amsterdam, Netherlands: 1999), 55 (17):3959–3974, December 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001861.

Theodoridis:2011:CRM

[535] Georgios Theodoridis and Fotini-Niovi Pavlidou. A combined resource management and admission control scheme for optimizing uplink performance of M-WiMAX systems. Computer Networks (Amsterdam, Netherlands: 1999), 55(17): 3975–3986, December 1, 2011. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611001642.

Anonymous:2011:EBp

[536] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 55(17):??, December 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611003458.

Akyildiz:2012:E

[537] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 1–2, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611004221.

Yoon:2012:ARS

[538] Hyenyoung Yoon, Junseok Hwang, and Martin B. H. Weiss. An analytic research on secondary-spectrum trading mechanisms based on technical and market changes. Computer Networks (Amsterdam, Netherlands: 1999), 56 (1):3–19, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611001848.

Choi:2012:IBC

[539] Hyunsang Choi and Heejo Lee. Identifying botnets by capturing group activities in DNS traffic. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):20–33, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002787.

Zadeh:2012:DOD

[540] P. D. Hossein Zadeh, C. Schlegel, and M. H. MacGregor. Distributed optimal dynamic base station positioning in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):34–49, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611002969.

Kupcu:2012:UOF

[541] Alptekin Küpçü and Anna Lysyanskaya. Usable optimistic fair exchange. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 50–63, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100301X.

Mao:2012:MDP

[542] Yun Mao, Boon Thau Loo, Zachary Ives, and Jonathan M. Smith. MO-SAIC: Declarative platform for dynamic overlay composition. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):64–84, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003021.

Dong:2012:NMB

[543] Ziqian Dong, Rohan D. W. Perera, Rajarathnam Chandramouli, and K. P. Subbalakshmi. Network measurement based modeling and optimization for IP geolocation. Computer Networks (Amsterdam, Netherlands: 1999), 56 (1):85–98, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003173.

Aly:2012:NPC

[544] Salah A. Aly, Ahmed E. Kamal, and Osameh M. Al-Kofahi. Network protection codes: Providing self-healing in autonomic networks using network coding. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 99–111, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003203.

Lucas-Estan:2012:ILP

[545] M. C. Lucas-Estañ, J. Gozalvez, and J. Sanchez-Soriano. Integer linear programming optimization of joint RRM policies for heterogeneous wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 112–126, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003197.

Rizk:2012:NAE

[546] Amr Rizk and Markus Fidler. Nonasymptotic end-to-end performance bounds for networks with long range dependent fBm cross traffic. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):127– 141, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003185.

Crisostomo:2012:PFS

[547] Sérgio Crisóstomo, Udo Schilcher, Christian Bettstetter, and João Barros. Probabilistic flooding in stochastic networks: Analysis of global information outreach. Computer Networks (Amsterdam, Netherlands: 1999), 56 (1):142–156, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003215.

Razavi:2012:PCT

[548] Sara Modarres Razavi, Di Yuan, Fredrik Gunnarsson, and Johan Moe. Performance and cost tradeoff in Tracking Area reconfiguration: a Pareto-optimization approach. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 157–168, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003331.

Gamer:2012:CAB

[549] Thomas Gamer. Collaborative anomaly-based detection of large-scale Internet attacks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (1):169–185, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003318.

Cordeschi:2012:TEW

[550] Nicola Cordeschi, Valentina Polli, and Enzo Baccarelli. Traffic Engineering for wireless connectionless access networks supporting QoS-demanding media applications. Computer Networks (Amsterdam, Netherlands: 1999), 56 (1):186–197, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861100332X.

Joung:2012:CFR

[551] Yuh-Jzer Joung, Terry Hui-Ye Chiu, and Shy-Min Chen. Cooperating with free riders in unstructured P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 198–212, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003367.

Joung:2012:MDC

[552] Yuh-Jzer Joung, Shih-Hsiang Huang, and Shi-Hang Lin. Making datacentric storage adaptive and costoptimal. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):

213–230, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003343.

Park:2012:EIA

[553] Hyuntae Park, Hyejeong Hong, and Sungho Kang. An efficient IP address lookup algorithm based on a small balanced tree using entry reduction. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 231–243, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003379.

Mirahmadi:2012:TPA

[554] Maysam Mirahmadi and Abdallah Shami. Traffic-prediction-assisted dynamic bandwidth assignment for hybrid optical wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):244–259, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611003392.

Lu:2012:SLF

[555] Chun-Nan Lu, Chun-Ying Huang, Ying-Dar Lin, and Yuan-Cheng Lai. Session level flow classification by packet size distribution and session grouping. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 260–272, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003410.

Moessner:2012:SAS

[556] M. Moessner and Gul N. Khan. Secure authentication scheme for passive C1G2 RFID tags. Computer Networks (Amsterdam, Netherlands: 1999), 56 (1):273–286, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003422.

Baccarelli:2012:QST

[557] Enzo Baccarelli, Nicola Cordeschi, and Tatiana Patriarca. QoS Stochastic Traffic Engineering for the wireless support of real-time streaming applications. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 287–302, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003525.

Toril:2012:EIP

[558] M. Toril, P. Guerrero-García, S. Luna-Ramírez, and V. Wille. An efficient integer programming formulation for the assignment of base stations to controllers in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 303–314, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003549.

Yun:2012:CLA

[559] Ji-Hoon Yun. Cross-layer analysis of the random access mechanism in Universal Terrestrial Radio Access. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 315–328, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003537.

Li:2012:ISS

[560] Dagang Li, Emmanuel Van Lil, and Antoine Van de Capelle. Improving Slow-start based probing mechanisms for flow adaptation after handovers. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 329–344, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003033.

Kilavuz:2012:PAM

[561] Mustafa O. Kilavuz and Murat Yuksel. Path approximation for multihop wireless routing under applicationbased accuracy constraints. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):345–364, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611003550.

Chen:2012:ONF

[562] Yiping Chen, Erwan Le Merrer, Zhe Li, Yaning Liu, and Gwendal Simon. OAZE: a network-friendly distributed zapping system for peer-to-peer IPTV. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 365–377, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003355.

Avramidis:2012:CPD

[563] Agapios Avramidis, Panayiotis Kotzanikolaou, Christos Douligeris, and Mike Burmester. Chord-PKI: a distributed trust infrastructure based on P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 378–398, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003574.

Bai:2012:COO

[564] Huali Bai and Ming Chen. CCIPCA-OPCSC: an online method for detecting shared congestion paths. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):399–411, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611003586.

Alins:2012:XCL

[565] Juanjo Alins, Jorge Mata-Diaz, Jose L. Muñoz, Elizabeth Rendón-Morales, and Oscar Esparza. XPLIT: a cross-layer architecture for TCP services over DVB-S2/ETSI QoS BSM. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 412–434, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003380.

Tian:2012:MPE

[566] Guosong Tian and Yu-Chu Tian. Modelling and performance evaluation of the IEEE 802.11 DCF for real-time control. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 435–447, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003598.

Xie:2012:SIM

[567] Mengjun Xie, Zhenyu Wu, and Haining Wang. Secure instant messaging in enterprise-like networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(1): 448–461, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003409.

Callegari:2012:BAT

[568] C. Callegari, S. Giordano, M. Pagano, and T. Pepe. Behavior analysis of TCP Linux variants. Computer Networks (Amsterdam, Netherlands: 1999), 56 (1):462–476, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003677.

Han:2012:SPL

[569] Weili Han and Chang Lei. A survey on policy languages in network and security management. Computer Networks (Amsterdam, Netherlands: 1999), 56 (1):477–489, January 12, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003562.

Anonymous:2012:EBa

[570] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(1):??, January 12, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611004257.

Khan:2012:SRS

[571] Shafiullah Khan, Nabil Ali Alrajeh, and Kok-Keong Loo. Secure route selection in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 491–503, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611002519.

Papapetrou:2012:CLR

[572] Evangelos Papapetrou, Panos Vassiliadis, Efthymia Rova, and Apostolos Zarras. Cross-layer routing for peer database querying over mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):504–520, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003707.

Bruneo:2012:EWS

[573] Dario Bruneo, Salvatore Distefano, Francesco Longo, Antonio Puliafito, and Marco Scarpa. Evaluating wireless sensor node longevity through Markovian techniques. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):521–532, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003690.

Yetgin:2012:EPD

[574] Zeki Yetgın and Turgay Çelık. Efficient progressive downloading over multimedia broadcast multicast service. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 533–547, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003689.

Shih:2012:AHD

[575] Hung-Cheng Shih and Kuochen Wang. An adaptive hybrid dynamic power management algorithm for mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 548–565, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003719.

Rifa-Pous:2012:AHD

[576] Helena Rifà-Pous and Carles Garrigues. Authenticating hard decision sensing reports in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):566–576, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003720.

Carneiro:2012:TST

[577] Gustavo Carneiro, Pedro Fortuna, Jaime Dias, and Manuel Ricardo. Transparent and scalable terminal mobility for vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(2):577–597, February 2, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611003732.

Macia-Fernandez:2012:EUW

[578] Gabriel Maciá-Fernández, Yong Wang, Rafael A. Rodri’guez-Gómez, and Aleksandar Kuzmanovic. Extracting user Web browsing patterns from non-content network traces: The online advertising case study. Computer Networks (Amsterdam, Netherlands: 1999), 56(2):598–614, February 2, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611003781.

Isik:2012:MSL

[579] Sinan Isik, Mehmet Yunus Donmez, and Cem Ersoy. Multi-sink load balanced forwarding with a multi-criteria fuzzy sink selection for video sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 615–627, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003768.

Silveira:2012:PPL

[580] Fernando Silveira and Edmundo de Souza e Silva. Predicting packet loss statistics with hidden Markov models for FEC control. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):628–641, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003756.

Pereira:2012:ASR

[581] Ricardo Lopes Pereira, Teresa Vazão, and Rodrigo Rodrigues. Adaptive Search Radius — using hop count to reduce P2P traffic. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):642–660, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100380X.

Wang:2012:RBI

[582] Shie-Yuan Wang, Chih-Che Lin, and Yu-Chi Chang. A rule-based intersession network coding scheme over IEEE 802.16(d) mesh CDS-mode networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 661–685, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003793.

Mata:2012:DTC

[583] Felipe Mata, José Luis García-Dorado, and Javier Aracil. Detection of traffic changes in large-scale backbone networks: The case of the Spanish academic network. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):686–702, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003835.

Campista:2012:RPS

[584] Miguel Elias M. Campista, Luís Henrique M. K. Costa, and Otto Carlos M. B. Duarte. A routing protocol suitable for backhaul access in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):703–718, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003823.

Lee:2012:CDP

[585] Yong Oh Lee and A. L. Narasimha Reddy. Constructing disjoint paths for failure recovery and multipath routing. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 719–730, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003811.

Yang:2012:HHI

[586] Xinyu Yang, Xiaojing Fan, Wei Yu, Xinwen Fu, and Shusen Yang. HLLS: a History information based Light Location Service for MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 56(2):731–744, February 2, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611003847.

Feng:2012:STF

[587] Jie Feng and Lisong Xu. Stochastic TCP friendliness: Expanding the design space of TCP-friendly traffic control protocols. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):745–761, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003860.

Obeidat:2012:CLO

[588] Suhaib A. Obeidat, Abraham N. Aldaco, and Violet R. Syrotiuk. Cross-layer opportunistic adaptation for voice over ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 762–779, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003859.

Deccio:2012:QDN

[589] Casey Deccio, Jeff Sedayao, Krishna Kant, and Prasant Mohapatra. Quantifying DNS namespace influence. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 780–794, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003926.

Halepovic:2012:ERN

[590] Emir Halepovic, Carey Williamson, and Majid Ghaderi. Enhancing redundant network traffic elimination. Computer Networks (Amsterdam, Netherlands: 1999), 56(2):

795–809, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003914.

Marcus:2012:RRG

[591] D. Marcus and Y. Shavitt. RAGE — a rapid graphlet enumerator for large networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):810–819, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003902.

Cheng:2012:NLB

[592] Bo-Chao Cheng, Guo-Tan Liao, Ryh-Yuh Tseng, and Ping-Hai Hsu. Network lifetime bounds for hierarchical wireless sensor networks in the presence of energy constraints. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 820–831, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003884.

Bayoglu:2012:GBS

[593] Burak Bayoglu and Ibrahim Sogukpinar. Graph based signature classes for detecting polymorphic worms via content analysis. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):832–844, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100394X.

Xu:2012:SSA

[594] Mingwei Xu, Yuan Yang, and Qi Li. Selecting shorter alternate paths for tunnel-based IP Fast ReRoute in linear time. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 845–857, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003938.

Donnet:2012:PSE

[595] Benoit Donnet, Bamba Gueye, and Mohamed Ali Kaafar. Path similarity evaluation using Bloom filters. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 858–869, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003872.

Ziviani:2012:RPA

[596] Artur Ziviani, Thiago B. Cardozo, and Antônio Tadeu A. Gomes. Rapid prototyping of active measurement tools. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 870–883, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004002.

Jamshaid:2012:MCF

[597] Kamran Jamshaid, Paul A. S. Ward, and Martin Karsten. Mechanisms for centralized flow rate control in 802.11-based wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(2):

884–901, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004014.

Zarifzadeh:2012:EET

[598] Sajjad Zarifzadeh, Nasser Yazdani, and Amir Nayyeri. Energy-efficient topology control in wireless ad hoc networks with selfish nodes. Computer Networks (Amsterdam, Netherlands: 1999), 56 (2):902–914, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004026.

Cano:2012:WAT

[599] C. Cano, B. Bellalta, and M. Oliver. Wake up after transmissions and reduced channel contention to alleviate the hidden terminal problem in preamble sampling WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 56(2):915–926, February 2, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611004051.

Robert:2012:DRM

[600] Jean-Marc Robert, Hadi Otrok, Ahmad Nahar Quttoum, and Rihab Boukhris. A distributed resource management model for Virtual Private Networks: Tit-for-Tat strategies. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 927–939, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004075.

Alotaibi:2012:SRA

[601] Eiman Alotaibi and Biswanath Mukherjee. A survey on routing algorithms for wireless Ad-Hoc and mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(2): 940–965, February 2, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100377X.

Anonymous:2012:EBb

[602] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(2):??, February 2, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612000084.

Alvarez-Hamelin:2012:CDN

[603] José Ignacio Alvarez-Hamelin, Éric Fleury, Alessandro Vespignani, and Artur Ziviani. Complex dynamic networks: Tools and methods. Computer Networks (Amsterdam, Netherlands: 1999), 56(3):967–969, February 23, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611004233.

Koenigstein:2012:TSP

[604] N. Koenigstein and Y. Shavitt. Talent scouting in P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(3):970–982, February 23, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128611003963.

Kim:2012:CPD

[605] Hyoungshick Kim, John Tang, Ross Anderson, and Cecilia Mascolo. Centrality prediction in dynamic human contact networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (3):983–996, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003975.

Vendramin:2012:GIB

[606] Ana Cristina Kochem Vendramin, Anelise Munaretto, Myriam Regattieri Delgado, and Aline Carneiro Viana. GrAnt: Inferring best forwarders from complex networks’ dynamics through a greedy Ant Colony Optimization. Computer Networks (Amsterdam, Netherlands: 1999), 56 (3):997–1015, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611004099.

Shelley:2012:GIS

[607] David S. Shelley and Mehmet Hadi Gunes. GerbilSphere: Inner sphere network visualization. Computer Networks (Amsterdam, Netherlands: 1999), 56 (3):1016–1028, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611003987.

Feldman:2012:SAP

[608] Dima Feldman, Yuval Shavitt, and Noa Zilberman. A structural approach for PoP geo-location. Computer Networks (Amsterdam, Netherlands: 1999), 56 (3):1029–1040, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004191.

Mitra:2012:IDN

[609] Bivas Mitra, Lionel Tabourier, and Camille Roth. Intrinsically dynamic network communities. Computer Networks (Amsterdam, Netherlands: 1999), 56(3):1041–1053, February 23, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611003999.

Varvello:2012:UBR

[610] Matteo Varvello, Moritz Steiner, and Koen Laevens. Understanding BitTorrent: a reality check from the ISP’s perspective. Computer Networks (Amsterdam, Netherlands: 1999), 56(3): 1054–1065, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004488.

Cha:2012:DIC

[611] Meeyoung Cha, Fabrício Benevenuto, Yong-Yeol Ahn, and Krishna P. Gummadi. Delayed information cascades in Flickr: Measurement, analysis, and modeling. Computer Networks (Amsterdam, Netherlands: 1999), 56(3): 1066–1076, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003951.

Fu:2012:ECN

[612] Xiaoming Fu, Yang Chen, Guy Leduc, and Laurent Mathy. Editorial for Computer Networks special issue on “Measurement-based optimization of P2P networking and applications”. Computer Networks (Amsterdam, Netherlands: 1999), 56(3): 1077–1079, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000217.

Memon:2012:MLS

[613] Ghulam Memon, Reza Rejaie, Yang Guo, and Daniel Stutzbach. Montra: a large-scale DHT traffic monitor. Computer Networks (Amsterdam, Netherlands: 1999), 56(3): 1080–1091, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100404X.

Koenigstein:2012:MVP

[614] Noam Koenigstein, Yuval Shavitt, Ela Weinsberg, and Udi Weinsberg. Measuring the validity of peer-to-peer data for information retrieval applications. Computer Networks (Amsterdam, Netherlands: 1999), 56(3): 1092–1102, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004038.

Ma:2012:MME

[615] Zhen Ma, Ke Xu, Jiangchuan Liu, and Haiyang Wang. Measurement, modeling and enhancement of BitTorrentbased VoD system. Computer Networks (Amsterdam, Netherlands: 1999), 56 (3):1103–1117, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100421X.

Pussep:2012:CTM

[616] Konstantin Pussep, Frank Lehrieder, Christian Gross, Simon Oechsner, Markus Guenther, and Sebastian Meyer. Cooperative traffic management for video streaming overlays. Computer Networks (Amsterdam, Netherlands: 1999), 56(3): 1118–1130, February 23, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004087.

Anonymous:2012:EBc

[617] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(3):??, February 23, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612000291.

Marchese:2012:SPE

[618] M. Marchese and M. Mongelli. Simple protocol enhancements of Rapid Spanning Tree Protocol over ring topologies. Computer Networks (Amsterdam, Netherlands: 1999), 56(4): 1131–1151, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003744.

Larroca:2012:MDL

[619] Federico Larroca and Jean-Louis Rougier. Minimum delay loadbalancing via nonparametric regression and no-regret algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):1152– 1166, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004117.

Shin:2012:ESA

[620] Haw-Yun Shin. Exploiting skewed access and energy-efficient algorithm to improve the performance of wireless data broadcasting. Computer Networks (Amsterdam, Netherlands: 1999), 56 (4):1167–1182, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004105.

Perez:2012:NCQ

[621] Luis Pérez, Luis Velasco, Juan Rodríguez, Pedro Capelastegui, Guillem Hernández-Sola, Lorena Calavia, Antonio Marqués, Borja Iribarne, Amador Pozo, and Antoine De Poorter. Network convergence and QoS for future multimedia services in the VI-SION project. Computer Networks (Amsterdam, Netherlands: 1999), 56 (4):1183–1199, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004154.

Chin:2012:NSA

[622] Kwan-Wu Chin, Sieteng Soh, and Chen Meng. Novel scheduling algorithms for concurrent transmit/receive wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (4):1200–1214, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004130.

Cinque:2012:DDL

[623] Marcello Cinque, Catello Di Martino, and Christian Esposito. On data dissemination for large-scale complex critical infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):1215–1235, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128611004129.

Yun:2012:CBS

[624] Jeongkyun Yun, Sung-Guk Yoon, Jin-Ghoo Choi, and Saewoong Bahk. Contention based scheduling for femtocell access points in a densely deployed network environment. Computer Networks (Amsterdam, Netherlands: 1999), 56 (4):1236–1248, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004178.

Vemu:2012:OML

[625] Koteswara Rao Vemu, Shalabh Bhatnagar, and N. Hemachandra. Optimal multi-layered congestion based pricing schemes for enhanced QoS. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):1249– 1262, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100418X.

Moltchanov:2012:STP

[626] Dmitri Moltchanov. A study of TCP performance in wireless environment using fixed-point approximation. Computer Networks (Amsterdam, Netherlands: 1999), 56(4): 1263–1285, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004063.

Muthuswamy:2012:PVC

[627] Praveen K. Muthuswamy, Aparna Gupta, Murat Yuksel, and Koushik Kar. Path-vector contracting: Profit maximization and risk management. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):1286– 1302, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611003896.

Familiar:2012:BSO

[628] Miguel S. Familiar, José F. Martínez, Iván Corredor, and Carlos García-Rubio. Building service-oriented Smart Infrastructures over Wireless Ad Hoc Sensor Networks: a middleware perspective. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):

1303–1328, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004208.

Kim:2012:SUB

[629] Byung-Gook Kim and Jang-Won Lee. Stochastic utility-based flow control algorithm for services with time-varying rate requirements. Computer Networks (Amsterdam, Netherlands: 1999), 56 (4):1329–1342, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004397.

Cohen:2012:VDR

[630] Aaron E. Cohen, Jian-Hung Lin, and Keshab K. Parhi. Variable data rate (VDR) network congestion control (NCC) applied to voice/audio communication. Computer Networks (Amsterdam, Netherlands: 1999), 56(4): 1343–1356, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004427.

Yousaf:2012:SPR

[631] Faqir Zarrar Yousaf and Christian Wietfeld. Solving pinball routing, race condition and loop formation issues in nested mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(4): 1357–1375, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861100449X.

Kanizo:2012:HTF

[632] Yossi Kanizo, David Hay, and Isaac Keslassy. Hash tables with finite buckets are less resistant to deletions. Computer Networks (Amsterdam, Netherlands: 1999), 56(4): 1376–1389, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611004476.

Falowo:2012:DRS

[633] Olabisi E. Falowo and H. Anthony Chan. Dynamic RAT selection for multiple calls in heterogeneous wireless networks using group decisionmaking technique. Computer Networks (Amsterdam, Netherlands: 1999), 56 (4):1390–1401, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611004518.

Lin:2012:PID

[634] Chih-Che Lin, Shie-Yuan Wang, and Teng-Wei Hsu. On the performances of IEEE 802.16(d) mesh CDS-mode networks using Single-Switched-Beam Antennas. Computer Networks (Amsterdam, Netherlands: 1999), 56(4): 1402–1423, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128611004506.

Leith:2012:WCS

[635] D. J. Leith, P. Clifford, V. Badarla, and D. Malone. WLAN channel selection without communication. Computer Networks (Amsterdam, Netherlands: 1999), 56(4): 1424–1441, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000023.

Wang:2012:GMF

[636] Ting Wang and Chor Ping Low. The general Message Ferry Route (MFR^{∗}) problem and the An-Improved-Route (AIR) scheme. Computer Networks (Amsterdam, Netherlands: 1999), 56 (4):1442–1457, March 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000060.

Cho:2012:OHE

[637] Kyong-Tak Cho and Saewoong Bahk. Optimal Hop Extended MAC protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):1458– 1469, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000059.

Li:2012:ESP

[638] Yunzhao Li, Don Gruenbacher, and Caterina Scoglio. Evaluating stranger policies in P2P file-sharing systems with reciprocity mechanisms. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):1470– 1485, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000047.

Zhang:2012:CPP

[639] Wei Zhang, Jun Bi, Jianping Wu, and Baobao Zhang. Catching popular prefixes at AS border routers with a prediction based method. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):1486– 1502, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000175.

Rahimi:2012:FOM

[640] M. Reza Rahimi, Abdul Bais, and Nima Sarshar. On fair and optimal multi-source IP-multicast. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):1503–1524, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128612000199.

Anonymous:2012:EBd

[641] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(4):??, March 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000503.

Chang:2012:IVF

[642] Ching-Lung Chang and Sih-Ping Huang. The interleaved video frame distribution for P2P-based VoD system with VCR functionality. Computer Networks (Amsterdam, Netherlands: 1999), 56(5):

1525–1537, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000035.

Fitzpatrick:2012:VCC

[643] J. Fitzpatrick. Voice call capacity analysis of long range WiFi as a femto backhaul solution. Computer Networks (Amsterdam, Netherlands: 1999), 56 (5):1538–1553, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000205.

Campolo:2012:MMP

[644] Claudia Campolo, Claudio Casetti, Carla-Fabiana Chiasserini, and Antonella Molinaro. A multirate MAC protocol for reliable multicast in multihop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(5): 1554–1567, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000229.

Bezahaf:2012:EEC

[645] Mehdi Bezahaf, Luigi Iannone, Marcelo Dias de Amorim, and Serge Fdida. An experimental evaluation of cross-layer routing in a wireless mesh backbone. Computer Networks (Amsterdam, Netherlands: 1999), 56(5): 1568–1583, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000187.

Vieira:2012:SLH

[646] Fabio R. J. Vieira, José F. de Rezende, Valmir C. Barbosa, and Serge Fdida. Scheduling links for heavy traffic on interfering routes in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(5): 1584–1598, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000357.

Wang:2012:CAC

[647] Bang Wang, Hock Beng Lim, and Di Ma. A coverage-aware clustering protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(5): 1599–1611, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000400.

Maenpaa:2012:PER

[648] Jouni Mäenpää. Performance evaluation of Recursive Distributed Rendezvous based service discovery for Peer-to-Peer Session Initiation Protocol. Computer Networks (Amsterdam, Netherlands: 1999), 56(5): 1612–1626, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000394.

Rocha:2012:WCB

[649] Atslands R. Rocha, Luci Pirmez, Flávia C. Delicato, Érico Lemos, Igor Santos, Danielo G. Gomes, and José Neuman de Souza. WSNs clustering based on semantic neighborhood relationships. Computer Networks (Amsterdam, Netherlands: 1999), 56(5): 1627–1645, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000382.

Pyun:2012:IBF

[650] Young June Pyun, Younghee Park, Douglas S. Reeves, Xinyuan Wang, and Peng Ning. Interval-based flow watermarking for tracing interactive traffic. Computer Networks (Amsterdam, Netherlands: 1999), 56(5): 1646–1665, March 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000412.

Anonymous:2012:EBe

[651] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(5):??, March 30, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000692.

Raspall:2012:EPS

[652] Frederic Raspall. Efficient packet sampling for accurate traffic measurements. Computer Networks (Amsterdam, Netherlands: 1999), 56(6): 1667–1684, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004142.

Vazifehdan:2012:LNN

[653] Javad Vazifehdan, R. Venkatesha Prasad, and Ignas Niemegeers. On the lifetime of node-to-node communication in wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(6): 1685–1709, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861100452X.

Lee:2012:EUP

[654] Sihyung Lee and Hyong S. Kim. Enduser perspectives of Internet connectivity problems. Computer Networks (Amsterdam, Netherlands: 1999), 56 (6):1710–1722, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000230.

Liu:2012:SCS

[655] Alex X. Liu, Jason M. Kovacs, and Mohamed G. Gouda. A secure cookie scheme. Computer Networks (Amsterdam, Netherlands: 1999), 56(6): 1723–1730, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000370.

Hu:2012:PCS

[656] Yi Hu, Laxmi N. Bhuyan, and Min Feng. P2P consistency support for large-scale interactive applications. Computer Networks (Amsterdam, Netherlands: 1999), 56(6): 1731–1744, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000369.

Baek:2012:AAH

[657] Joo-Young Baek and Young-Joo Suh. An adaptive ARQ-HARQ interworking scheme in WiMAX systems. Computer Networks (Amsterdam, Netherlands: 1999), 56(6):1745–1762, April 19, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000448.

Lu:2012:WAI

[658] Yu Lu, Mehul Motani, and Wai-Choong Wong. When Ambient Intelligence meets the Internet: User Module framework and its applications. Computer Networks (Amsterdam, Netherlands: 1999), 56(6): 1763–1781, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000436.

Ouyang:2012:PFV

[659] Zhipeng Ouyang, Lisong Xu, Byrav Ramamurthy, and Negede Yossef. Partial forwarding vs. partial participation for dynamic window resizing in P2P streaming. Computer Networks (Amsterdam, Netherlands: 1999), 56 (6):1782–1796, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000424.

Cheng:2012:VNE

[660] Xiang Cheng, Sen Su, Zhongbao Zhang, Kai Shuang, Fangchun Yang, Yan Luo, and Jie Wang. Virtual network embedding through topology awareness and optimization. Computer Networks (Amsterdam, Netherlands: 1999), 56(6):1797–1813, April 19, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000461.

Yin:2012:NTC

[661] ChengGuo Yin, ShuangQing Li, and Qi Li. Network traffic classification via HMM under the guidance of syntactic structure. Computer Networks (Amsterdam, Netherlands: 1999), 56 (6):1814–1825, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200045X.

Rezgui:2012:OMH

[662] Jihene Rezgui, Abdelhakim Hafid, Racha Ben Ali, and Michel Gendreau. Optimization model for handoffaware channel assignment problem for multi-radio wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(6): 1826–1846, April 19, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200059X.

Soproni:2012:PRM

[663] Péter Soproni and Tibor Cinkler. Preplanned restoration of multicast demands in optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(6):1847–1861, April 19, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000588.

Anonymous:2012:EBf

[664] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(6):??, April 19, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000849.

Awal:2012:ICL

[665] Mohammad Abdul Awal, Lila Boukhatem, and Lin Chen. An integrated crosslayer framework of adaptive FEedback REsource allocation and Prediction for OFDMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):1863–1875, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000606.

Aydin:2012:ETF

[666] Ilknur Aydin, Janardhan Iyengar, Phillip Conrad, Chien-Chung Shen, and Paul Amer. Evaluating TCPfriendliness in light of Concurrent Multipath Transfer. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):1876–1892, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000242.

Saldana:2012:EIM

[667] Jose Saldana, Julián Fernández-Navajas, José Ruiz-Mas, Jenifer Murillo, Eduardo Viruete Navarro, and José I. Aznar. Evaluating the influence of multiplexing schemes and buffer implementation on perceived VoIP conversation quality. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):1893–1919, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000618.

Furthmuller:2012:EAR

[668] Jochen Furthmüller and Oliver P. Waldhorst. Energy-aware resource sharing with mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 56(7):1920–1934, May 3, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000655.

Ammari:2012:PCM

[669] Habib M. Ammari. On the problem of k-coverage in mission-oriented mobile wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(7):1935–1950, May 3, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000667.

Liu:2012:DPI

[670] Anfeng Liu, Ju Ren, Xu Li, Zhigang Chen, and Xuemin (Sherman) Shen. Design principles and improvement of cost function based energy aware routing algorithms for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):1951–1967, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000643.

Tamma:2012:TSC

[671] Bheemarjuna Reddy Tamma, B. S. Manoj, and Ramesh R. Rao. Traffic sensing and characterization in multichannel wireless networks for cognitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):1968–1982, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128611004166.

Murai:2012:HDT

[672] Fabricio Murai, Antonio A. de A. Rocha, Daniel R. Figueiredo, and Edmundo A. de Souza e Silva. Heterogeneous download times in a homogeneous Bit-Torrent swarm. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):1983–2000, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000679.

Tran:2012:CSA

[673] Duc A. Tran, Khanh Nguyen, and Cuong Pham. S-CLONE: Sociallyaware data replication for social networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(7): 2001–2013, May 3, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000746.

Ion:2012:DIC

[674] Mihaela Ion, Giovanni Russello, and Bruno Crispo. Design and implementation of a confidentiality and access control solution for publish/ subscribe systems. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):2014–2037, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000771.

Michiardi:2012:PAC

[675] Pietro Michiardi, Damiano Carra, Francesco Albanese, and Azer Bestavros. Peer-assisted content distribution on a budget. Computer Networks (Amsterdam, Netherlands: 1999), 56(7): 2038–2048, May 3, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000758.

Wang:2012:SAN

[676] Zhe Wang, Kai Hu, Ke Xu, Baolin Yin, and Xiaowen Dong. Structural analysis of network traffic matrix via relaxed principal component pursuit. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):2049–2067, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000813.

How:2012:ADS

[677] Kiam Cheng How, Maode Ma, and Yang Qin. An altruistic differentiated service protocol in dynamic cognitive radio networks against selfish behaviors. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):2068–2079, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000783.

Akon:2012:BEH

[678] Mursalin Akon, Mohammad Towhidul Islam, Xuemin (Sherman) Shen, and Ajit Singh. A bandwidth and effective hit optimal cache scheme for wireless data access networks with client injected updates. Computer Networks (Amsterdam, Netherlands: 1999), 56 (7):2080–2095, May 3, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000795.

Anonymous:2012:EBg

[679] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(7):??, May 3, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001090.

Reinhard:2012:OCO

[680] Vincent Reinhard, Johanne Cohen, Joanna Tomasik, Dominique Barth, and Marc-Antoine Weisser. Optimal configuration of an optical network providing predefined multicast transmissions. Computer Networks (Amsterdam, Netherlands: 1999), 56 (8):2097–2106, May 24, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200062X.

Hassen:2012:EKM

[681] Hani Ragab Hassen, Hatem Bettahar, Abdalmadjid Bouadbdallah, and Yacine Challal. An efficient key management scheme for content access control for linear hierarchies. Computer Networks (Amsterdam, Netherlands: 1999), 56(8):2107–2118, May 24, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000631.

Cao:2012:SRH

[682] Jin Cao, Hui Li, Maode Ma, Yueyu Zhang, and Chengzhe Lai. A simple and robust handover authentication between HeNB and eNB in LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (8):2119–2131, May 24, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200076X.

Liao:2012:DDP

[683] Yong Liao, Jiangtao Yin, Dong Yin, and Lixin Gao. DPillar: Dual-port server interconnection network for large scale data centers. Computer Networks (Amsterdam, Netherlands: 1999), 56 (8):2132–2147, May 24, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000801.

Joung:2012:BNA

[684] Yuh-Jzer Joung, Wing-Tat Wong, Hsiao-Mei Huang, and Yi-Fang Chou. Building a network-aware and loadbalanced peer-to-peer system for range queries. Computer Networks (Amsterdam, Netherlands: 1999), 56(8): 2148–2167, May 24, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000825. See

erratum [697].

Chin:2012:NAC

[685] Kwan-Wu Chin and Shinan Li. Novel association control strategies for multicasting in relay-enabled WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 56(8): 2168–2178, May 24, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000916.

Makela:2012:CLB

[686] Antti Mäkelä, Sebastian Siikavirta, and Jukka Manner. Comparison of load-balancing approaches for multipath connectivity. Computer Networks (Amsterdam, Netherlands: 1999), 56 (8):2179–2195, May 24, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200093X.

Bilgin:2012:PEZ

[687] B. E. Bilgin and V. C. Gungor. Performance evaluations of ZigBee in different smart grid environments. Computer Networks (Amsterdam, Netherlands: 1999), 56(8):2196–2205, May 24, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000941.

Lee:2012:IAB

[688] Sihyung Lee, Kyriaki Levanti, and Hyong S. Kim. Impact analysis of BGP sessions for prioritization of maintenance operations. Computer Networks (Amsterdam, Netherlands: 1999), 56 (8):2206–2220, May 24, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000953.

Ekbatanifard:2012:QMQ

[689] GholamHossein Ekbatanifard, Reza Monsefi, Mohammad H. Yaghmaee M., and Seyed Amin Hosseini S. Queen-MAC: a quorum-based energy-efficient medium access control protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(8):2221–2236, May 24, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000977.

Anonymous:2012:EBh

[690] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(8):??, May 24, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200134X.

Matray:2012:SPI

[691] Péter Mátray, Péter Hága, Sándor Laki, Gábor Vattay, and István Csabai. On the spatial properties of Internet routes. Computer Networks (Amsterdam, Netherlands: 1999), 56(9): 2237–2248, June 14, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000989.

Fragni:2012:EAE

[692] Carlo Fragni and Luís Henrique Maciel Kosmalski Costa. ECO-ALOC: Energy-efficient resource allocation for cluster-based software routers. Computer Networks (Amsterdam, Netherlands: 1999), 56(9):2249– 2261, June 14, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001004.

Lezama:2012:DEO

[693] Fernando Lezama, Gerardo Castañón, and Ana Maria Sarmiento. Differential evolution optimization applied to the wavelength converters placement problem in all optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(9):2262–2275, June 14, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001181.

Pham:2012:DSN

[694] Tuan-Minh Pham and Serge Fdida. DTN support for news dissemination in an urban area. Computer Networks (Amsterdam, Netherlands: 1999), 56 (9):2276–2291, June 14, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861200117X.

Hu:2012:EHB

[695] Changhui Hu, Tat Wing Chim, S. M. Yiu, Lucas C. K. Hui, and Victor O. K. Li. Efficient HMAC-based secure communication for VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 56(9):2292–2303, June 14, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001314.

Liu:2012:UBS

[696] Zheng Liu, Maode Ma, and Jufeng Dai. Utility-based scheduling in wireless multi-hop networks over non-deterministic fading channels. Computer Networks (Amsterdam, Netherlands: 1999), 56(9):2304– 2315, June 14, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001193.

Joung:2012:ESN

[697] Yuh-Jzer Joung, Wing-Tat Wong, Hsiao-Mei Huang, and Yi-Fang Chou. Erratum to “Building a network-aware and load-balanced peer-to-peer system for range queries”, COMNET, 56(8) 2012, 2148–2167. Computer Networks (Amsterdam, Netherlands: 1999), 56(9):2316, June 14, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001697. See

[684].

Anonymous:2012:EBi

[698] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(9):??, June 14, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001636.

Capone:2012:ECN

[699] Antonio Capone, Dan Kilper, and Zhisheng Niu. Editorial for Computer Networks special issue on “green communication networks”. Computer Networks (Amsterdam, Netherlands: 1999), 56(10):2317–2319, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001843.

Bolla:2012:CEB

[700] Raffaele Bolla, Roberto Bruschi, Alessandro Carrega, Franco Davoli, Diego Suino, Constantinos Vassilakis, and Anastasios Zafeiropoulos. Cutting the energy bills of Internet Service Providers and telecoms through power management: an impact analysis. Computer Networks (Amsterdam, Netherlands: 1999), 56(10): 2320–2342, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001326.

Kim:2012:ACB

[701] Young-Min Kim, Eun-Jung Lee, Hea-Sook Park, Jun-Kyun Choi, and Hong-Shik Park. Ant colony based selfadaptive energy saving routing for energy efficient Internet. Computer Networks (Amsterdam, Netherlands: 1999), 56(10):2343–2354, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001235.

Cuomo:2012:NPE

[702] Francesca Cuomo, Antonio Cianfrani, Marco Polverini, and Daniele Mangione. Network pruning for energy saving in the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 56 (10):2355–2367, July 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001028.

Avallone:2012:EEO

[703] Stefano Avallone and Giorgio Ventre. Energy efficient online routing of flows with additive constraints. Computer Networks (Amsterdam, Netherlands: 1999), 56(10):2368–2382, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001041.

Hou:2012:MGR

[704] Weigang Hou, Lei Guo, Xuetao Wei, and Xiaoxue Gong. Multi-granularity and robust grooming in power- and port-cost-efficient IP over WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(10): 2383–2399, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001223.

Vizcaino:2012:EEA

[705] Jorge López Vizcaíno, Yabin Ye, and Idelfonso Tafur Monroy. Energy efficiency analysis for flexible-grid OFDMbased optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(10):2400–2419, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001053.

Ricciardi:2012:EAD

[706] Sergio Ricciardi, Francesco Palmieri, Ugo Fiore, Davide Careglio, Germán Santos-Boada, and Josep Solé-Pareta. An energy-aware dynamic RWA framework for next-generation wavelengthrouted networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (10):2420–2442, July 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001156.

Rizzelli:2012:EET

[707] Giuseppe Rizzelli, Annalisa Morea, Massimo Tornatore, and Olivier Rival. Energy efficient Traffic-Aware design of on-off Multi-Layer translucent optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (10):2443–2455, July 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001065.

Herreria-Alonso:2012:OCE

[708] Sergio Herrería-Alonso, Miguel Rodri’guez-Pérez, Manuel Fernández-Veiga, and Cándido López-García. Optimal configuration of Energy-Efficient Ethernet. Computer Networks (Amsterdam, Netherlands: 1999), 56(10): 2456–2467, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612000990.

Wolkerstorfer:2012:ESL

[709] Martin Wolkerstorfer, Driton Statovci, and Tomas Nordström. Energysaving by low-power modes in ADSL2. Computer Networks (Amsterdam, Netherlands: 1999), 56(10):2468– 2480, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001144.

Mancuso:2012:APS

[710] Vincenzo Mancuso and Sara Alouf. Analysis of power saving with continuous connectivity. Computer Networks (Amsterdam, Netherlands: 1999), 56 (10):2481–2493, July 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861200103X.

DeTurck:2012:PAS

[711] Koen De Turck, Stijn De Vuyst, Dieter Fiems, Sabine Wittevrongel, and Herwig Bruneel. Performance analysis of sleep mode mechanisms in the presence of bidirectional traffic. Computer Networks (Amsterdam, Netherlands: 1999), 56(10):2494–2505, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001077.

Gomez:2012:MBM

[712] Karina Gomez, Dejene Boru, Roberto Riggio, Tinku Rasheed, Daniele Miorandi, and Fabrizio Granelli. Measurement-based modelling of power consumption at wireless access network gateways. Computer Networks (Amsterdam, Netherlands: 1999), 56 (10):2506–2521, July 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001442.

daSilva:2012:EPT

[713] Ana Paula Couto da Silva, Michela Meo, and Marco Ajmone Marsan. Energy-performance trade-off in dense WLANs: a queuing study. Computer Networks (Amsterdam, Netherlands: 1999), 56(10):2522–2537, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001168.

Nguyen:2012:EAV

[714] Kim Khoa Nguyen, Mohamed Cheriet, Mathieu Lemay, Victor Reijs, Andrew Mackarel, and Alin Pastrama.

Environmental-aware virtual data center network. Computer Networks (Amsterdam, Netherlands: 1999), 56 (10):2538–2550, July 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001016.

Anonymous:2012:EBj

[715] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(10):??, July 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001867.

Pingley:2012:CAS

[716] Aniket Pingley, Wei Yu, Nan Zhang, Xinwen Fu, and Wei Zhao. A context-aware scheme for privacypreserving location-based services. Computer Networks (Amsterdam, Netherlands: 1999), 56(11):2551– 2568, July 31, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001211.

Cordeiro:2012:IMB

[717] Weverton Luis da Costa Cordeiro, Flávio Roberto Santos, Gustavo Huff Mauch, Marinho Pilla Barcelos, and Luciano Paschoal Gaspary. Identity management based on adaptive puzzles to protect P2P systems from Sybil attacks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (11):2569–2589, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001417.

Lin:2012:UEP

[718] Cheng-Han Lin, Yu-Chi Wang, Ce-Kuen Shieh, and Wen-Shyang Hwang. An unequal error protection mechanism for video streaming over IEEE 802.11e WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 56 (11):2590–2599, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001430.

Elrabiei:2012:RCM

[719] Sara Moftah Elrabiei and Mohamed Hadi Habaebi. Reliable cooperative multicasting for MBS WiMAX traffic. Computer Networks (Amsterdam, Netherlands: 1999), 56(11): 2600–2613, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001466.

Chen:2012:IAM

[720] Ching-Wen Chen, Chuan-Chi Weng, and Po-Yueh Chen. An interference avoidance MAC protocol design in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(11):2614– 2634, July 31, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001454.

Wang:2012:PBD

[721] Sheng-Wei Wang. Probability based dynamic-alternate routing and the corresponding converter placement algorithm in all-optical WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(11): 2635–2648, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001491.

Bahrak:2012:SAP

[722] Behnam Bahrak, Amol Deshpande, and Jung-Min ‘Jerry’ Park. Spectrum access policy reasoning for policy-based cognitive radios. Computer Networks (Amsterdam, Netherlands: 1999), 56 (11):2649–2663, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200148X.

Dong:2012:BDD

[723] Wei Dong, Chun Chen, Xue Liu, Guodong Teng, Jiajun Bu, and Yunhao Liu. Bulk data dissemination in wireless sensor networks: Modeling and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 56 (11):2664–2676, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001478.

Tomita:2012:DUT

[724] Nobuyoshi Tomita and Shahrokh Valaee. Data uploading time estimation for CUBIC TCP in long distance networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (11):2677–2689, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001508.

Tang:2012:NDS

[725] Meiqin Tang, Chengnian Long, Xinping Guan, and Xinjiang Wei. Nonconvex dynamic spectrum allocation for cognitive radio networks via particle swarm optimization and simulated annealing. Computer Networks (Amsterdam, Netherlands: 1999), 56 (11):2690–2699, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001521.

Pacheco:2012:IEA

[726] Dino Martin Lopez Pacheco, Tuan Tran Thai, Emmanuel Lochin, and Fabrice Arnal. An IP-ERN architecture to enable hybrid E2E/ERN protocol and application to satellite networking. Computer Networks (Amsterdam, Netherlands: 1999), 56(11): 2700–2713, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200151X.

Fernandez-Lopez:2012:TDE

[727] Helena Fernández-López, José A. Afonso, J. H. Correia, and Ricardo Simoes. Towards the design of efficient nonbeacon-enabled ZigBee networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(11): 2714–2725, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001533.

Schaffer:2012:SRC

[728] Péter Schaffer, Károly Farkas, Ádám Horváth, Tamás Holczer, and Levente Buttyán. Secure and reliable clustering in wireless sensor networks: a critical survey. Computer Networks (Amsterdam, Netherlands: 1999), 56 (11):2726–2741, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861200120X.

Saputro:2012:SRP

[729] Nico Saputro, Kemal Akkaya, and Suleyman Uludag. A survey of routing protocols for smart grid communications. Computer Networks (Amsterdam, Netherlands: 1999), 56(11): 2742–2771, July 31, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001429.

Anonymous:2012:EBk

[730] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(11):??, July 31, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002095.

Wang:2012:VIB

[731] Pinghui Wang, Xiaohong Guan, Don Towsley, and Jing Tao. Virtual indexing based methods for estimating node connection degrees. Computer Networks (Amsterdam, Netherlands: 1999), 56(12):2773– 2787, August 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001272.

Lee:2012:ORN

[732] Sookyoung Lee and Mohamed Younis. Optimized relay node placement for connecting disjoint wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(12): 2788–2804, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001594.

Feitosa:2012:OAU

[733] Eduardo Feitosa, Eduardo Souto, and Djamel H. Sadok. An orchestration approach for unwanted Internet traffic identification. Computer Networks (Amsterdam, Netherlands: 1999), 56 (12):2805–2831, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001582.

Rashwand:2012:EAP

[734] Saeed Rashwand and Jelena Misić. Effects of access phases lengths on performance of IEEE 802.15.6 CSMA/ CA. Computer Networks (Amsterdam, Netherlands: 1999), 56(12): 2832–2846, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001740.

Liao:2012:DEE

[735] Jianxin Liao, Jinzhu Wang, Tonghong Li, Jing Wang, Jingyu Wang, and Xiaomin Zhu. A distributed end-to-end overload control mechanism for networks of SIP servers. Computer Networks (Amsterdam, Netherlands: 1999), 56(12):2847– 2868, August 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001752.

Castro:2012:DRS

[736] Alberto Castro, Luis Velasco, Marc Ruiz, Miroslaw Klinkowski, Juan Pedro Fernández-Palacios, and Davide Careglio. Dynamic routing and spectrum (re)allocation in future flexgrid optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (12):2869–2883, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001764.

Ortega:2012:PTD

[737] F. Javier Ortega, José A. Troyano, Fermín L. Cruz, Carlos G. Vallejo, and Fernando Enríquez. Propagation of trust and distrust for the detection of trolls in a social network. Computer Networks (Amsterdam, Netherlands: 1999), 56(12): 2884–2895, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200179X.

Camps-Mur:2012:LFA

[738] Daniel Camps-Mur, Manil Dev Gomony, Xavier Pérez-Costa, and Sebastià Sallent-Ribes. Leveraging 802.11n frame aggregation to enhance QoS and power consumption in Wi-Fi networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(12): 2896–2911, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200182X.

Montagud:2012:EAR

[739] Mario Montagud and Fernando Boronat. Enhanced adaptive RTCP-based Inter-Destination Multimedia Synchronization approach for distributed applications. Computer Networks (Amsterdam, Netherlands: 1999), 56(12): 2912–2933, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001818.

Yoo:2012:BSI

[740] Jae-Yong Yoo, Cigdem Sengul, Ruben Merz, and JongWon Kim. Backpressure scheduling in IEEE 802.11 wireless mesh networks: Gap between theory and practice. Computer Networks (Amsterdam, Netherlands: 1999), 56 (12):2934–2948, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001971.

Holzer:2012:AAL

[741] Adrian Holzer, Patrick Eugster, and Benoît Garbinato. ALPS — Adaptive Location-based Publish/ Subscribe. Computer Networks (Amsterdam, Netherlands: 1999), 56(12): 2949–2962, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001958.

Chieochan:2012:NCU

[742] S. Chieochan and E. Hossain. Network coding for unicast in a WiFi hotspot: Promises, challenges, and testbed implementation. Computer Networks (Amsterdam, Netherlands: 1999), 56 (12):2963–2980, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001946.

Wang:2012:SWM

[743] Ze Wang, Maode Ma, and Jigang Wu. Securing wireless mesh networks in a unified security framework with corruption-resilience. Computer Networks (Amsterdam, Netherlands: 1999), 56(12):2981–2993, August 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128612001983.

Wu:2012:RIB

[744] Tsu-Yang Wu, Yuh-Min Tseng, and Tung-Tso Tsai. A revocable ID-based authenticated group key exchange protocol with resistant to malicious participants. Computer Networks (Amsterdam, Netherlands: 1999), 56(12):

2994–3006, August 16, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001995.

Anonymous:2012:EBl

[745] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(12):??, August 16, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861200223X.

Pavon-Marino:2012:SIC

[746] Pablo Pavon-Marino, Mohammed Atiquzzaman, and Joan Garcia-Haro. Special issue on “Challenges in high-performance switching and routing in the Future Internet”. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3007–3009, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002502.

Lim:2012:NHP

[747] Hyesook Lim, Soohyun Lee, and Earl E. Swartzlander, Jr. A new hierarchical packet classification algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3010–3022, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001545.

Bremler-Barr:2012:LIC

[748] Anat Bremler-Barr, David Hay, and Danny Hendler. Layered interval codes for TCAM-based classification. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3023–3039, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001788.

Smiljanic:2012:CRS

[749] Aleksandra Smiljanić and Zoran Cica. A comparative review of scalable lookup algorithms for IPv6. Computer Networks (Amsterdam, Netherlands: 1999), 56(13):3040– 3054, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001806.

Neji:2012:PBA

[750] Nizar Ben Neji and Adel Bouhoula. A prefix-based approach for managing hybrid specifications in complex packet filtering. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3055–3064, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002034.

Sanli:2012:HDI

[751] Mustafa Sanli, Ece Güran Schmidt, and Hasan Cengiz Güran. Hardware design and implementation of packet fair queuing algorithms for the quality of service support in the high-speed Internet. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3065–3075, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001557.

Francini:2012:PED

[752] Andrea Francini. Periodic early detection for improved TCP performance and energy efficiency. Computer Networks (Amsterdam, Netherlands: 1999), 56(13):3076–3086, September 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001570.

Divakaran:2012:SDA

[753] Dinil Mon Divakaran. A spikedetecting AQM to deal with elephants. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3087–3098, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001776.

Hermida:2012:ABO

[754] D. Fernández Hermida, M. Rodelgo Lacruz, A. Bianco, D. Cuda, G. Gavilanes Castillo, C. López Bravo, and F. J. González Castaño. AWGbased optical switches performance using crosstalk limiting schedulers. Computer Networks (Amsterdam, Netherlands: 1999), 56(13):3099– 3109, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001831.

Cuda:2012:DCN

[755] Davide Cuda, Paolo Giaccone, and Massimo Montalto. Design and control of next generation distribution frames. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3110–3122, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002010.

Vilalta:2012:GEM

[756] Ricard Vilalta, Raul Muñoz, Ramon Casellas, Ricardo Martinez, and Javier Vílchez. GMPLS-enabled MPLS-TP/PWE3 node with integrated 10 Gbps tunable DWDM transponders: design and experimental evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3123–3135, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001569.

Molnar:2012:COS

[757] Miklós Molnár, Alia Bellabas, and Samer Lahoud. The cost optimal solution of the multi-constrained multicast routing problem. Computer Networks (Amsterdam, Netherlands: 1999), 56 (13):3136–3149, September 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612001600.

Lucerna:2012:TAA

[758] Diego Lucerna, Massimo Tornatore, Biswanath Mukherjee, and Achille Pattavina. Trading availability among shared-protected dynamic connections in WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(13): 3150–3162, September 5, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612001612.

Anonymous:2012:EBm

[759] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(13):??, September 5, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612002526.

Gaddour:2012:RNS

[760] Olfa Gaddour and Anis Koubâa. RPL in a nutshell: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 56(14):3163–3178, September 28, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002423.

Zhang:2012:AER

[761] Xian Zhang, Xiuzhong Chen, and Chris Phillips. Achieving effective resilience for QoS-aware application mapping. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3179–3191, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002411.

Lee:2012:IMT

[762] Sihyung Lee, Tina Wong, and Hyong S. Kim. Improving manageability through reorganization of routing-policy configurations. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3192–3205, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200240X.

Ahmadi:2012:EAO

[763] H. Ahmadi and Y. H. Chew. Evolutionary algorithms for orthogonal frequency division multiplexingbased dynamic spectrum access systems. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3206–3218, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002381.

Bianzino:2012:GGD

[764] Aruna Prem Bianzino, Luca Chiaraviglio, Marco Mellia, and Jean-Louis Rougier. GRiDA: GReen Distributed Algorithm for energy-efficient IP backbone networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (14):3219–3232, September 28, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002344.

Mondal:2012:PCN

[765] Amit Mondal, Ionut Trestian, Zhen Qin, and Aleksandar Kuzmanovic. P2P as a CDN: a new service model for file sharing. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3233–3246, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002290.

Hu:2012:SMT

[766] Zhiguo Hu, Dalu Zhang, Anqi Zhu, Zhiwei Chen, and Hualei Zhou. SLDRT: a measurement technique for available bandwidth on multihop path with bursty cross traffic. Computer Networks (Amsterdam, Netherlands: 1999), 56(14):3247– 3260, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002289.

Huang:2012:JSH

[767] Caishi Huang, Chin-Tau Lea, and Albert Kai-Sun Wong. A joint solution for the hidden and exposed terminal problems in CSMA/CA wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3261–3273, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002174.

Figueiredo:2012:OAB

[768] Gustavo B. Figueiredo, Eduardo Candido Xavier, and Nelson L. S. da Fonseca. Optimal algorithms for the batch scheduling problem in OBS networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3274–3286, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002149.

Fei:2012:RFR

[769] Xiang Fei and Evan Magill. REED: Flexible rule based programming of wireless sensor networks at runtime. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3287–3299, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002071.

Khalil:2012:CCT

[770] Issa Khalil, Mamoun Awad, and Abdallah Khreishah. CTAC: Control traffic tunneling attacks’ countermeasures in mobile wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3300–3317, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200206X.

Jeon:2012:ASS

[771] Wha Sook Jeon and Dong Geun Jeong. Adaptive sensing scheduling for cognitive radio systems. Computer Networks (Amsterdam, Netherlands: 1999), 56 (14):3318–3332, September 28, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002058.

Shpiner:2012:SBA

[772] Alexander Shpiner, Isaac Keslassy, Gabi Bracha, Eyal Dagan, Ofer Iny, and Eyal Soha. A switchbased approach to throughput collapse and starvation in data centers. Computer Networks (Amsterdam, Netherlands: 1999), 56(14):3333– 3346, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002046.

Narayanan:2012:JND

[773] Ram G. Lakshmi Narayanan and Oliver C. Ibe. A joint network for disaster recovery and search and rescue operations. Computer Networks (Amsterdam, Netherlands: 1999), 56(14): 3347–3373, September 28, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002022.

Leon:2012:CDP

[774] Olga León, Juan Hernández-Serrano, and Miguel Soriano. Cooperative detection of primary user emulation attacks in CRNs. Computer Networks (Amsterdam, Netherlands: 1999), 56 (14):3374–3384, September 28, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200196X.

Anonymous:2012:EBn

[775] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(14):??, September 28, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861200271X.

Tseng:2012:JCL

[776] Po-Kai Tseng and Wei-Ho Chung. Joint coverage and link utilization for fast IP local protection. Computer Networks (Amsterdam, Netherlands: 1999), 56(15):3385– 3400, October 15, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002435.

Doss:2012:MDA

[777] Robin Doss, Wanlei Zhou, Saravanan Sundaresan, Shui Yu, and Longxiang Gao. A minimum disclosure approach to authentication and privacy in RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 56(15): 3401–3416, October 15, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002447.

Zhang:2012:FLD

[778] Changwang Zhang, Zhiping Cai, Weifeng Chen, Xiapu Luo, and Jianping Yin. Flow level detection and filtering of low-rate DDoS. Computer Networks (Amsterdam, Netherlands: 1999), 56(15): 3417–3431, October 15, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002496.

Guo:2012:HCP

[779] Yang Guo, Chao Liang, and Yong Liu. Hierarchically Clustered P2P Video Streaming: Design, implementation, and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 56 (15):3432–3445, October 15, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002484.

Chen:2012:EED

[780] Yu Chen and Xuming Fang. Energyefficient dynamic resource allocation with opportunistic network coding in OFDMA relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(15):3446– 3455, October 15, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002460.

Mansour:2012:OMM

[781] Yishay Mansour, Boaz Patt-Shamir, and Dror Rawitz. Overflow management with multipart packets. Computer Networks (Amsterdam, Netherlands: 1999), 56(15):3456– 3467, October 15, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002459.

Bang:2012:ERB

[782] Jiwoong Bang and Daewon Kim. Efficient RTSP-based multiple buffering and packet transmission methods for delivering OMA PoC Box service. Computer Networks (Amsterdam, Netherlands: 1999), 56(15): 3468–3478, October 15, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002617.

Lehmann:2012:DSA

[783] Matheus Brenner Lehmann, Flávio Roberto Santos, Luciano Paschoal Gaspary, and Marinho Pilla Barcellos. Denial-of-service attacks and countermeasures on BitTorrent. Computer Networks (Amsterdam, Netherlands: 1999), 56(15): 3479–3498, October 15, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002605.

Krzesinski:2012:HDW

[784] A. E. Krzesinski, G. Latouche, and P. G. Taylor. How do we encourage an egoist to act socially in an ad hoc mobile network? Computer Networks (Amsterdam, Netherlands: 1999), 56 (15):3499–3510, October 15, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002630.

Zubow:2012:SSO

[785] Anatolij Zubow, Johannes Marotzke, Daniel Camps-Mur, and Xavier Perez Costa. sGSA: a SDMA-OFDMA greedy scheduling algorithm for WiMAX networks. Computer Networks (Amsterdam, Netherlands: 1999), 56 (15):3511–3530, October 15, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002642.

Botta:2012:TGR

[786] Alessio Botta, Alberto Dainotti, and Antonio Pescapé. A tool for the generation of realistic network workload for emerging networking scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 56(15): 3531–3547, October 15, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612000928.

Zhang:2012:SPP

[787] Xiangyang Zhang and Hossam Hassanein. A survey of peer-to-peer live video streaming schemes — an algorithmic perspective. Computer Networks (Amsterdam, Netherlands: 1999), 56(15):3548–3579, October 15, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612002393.

Anonymous:2012:EBo

[788] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(15):??, October 15, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861200299X.

Wu:2012:NRF

[789] Guang Wu, Shu Wang, Bang Wang, Yan Dong, and Shu Yan. A novel range-free localization based on regulated neighborhood distance for wireless ad hoc and sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(16): 3581–3593, November 14, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002629.

Atzori:2012:SIT

[790] Luigi Atzori, Antonio Iera, Giacomo Morabito, and Michele Nitti. The Social Internet of Things (SIoT) — When social networks meet the Internet of Things: Concept, architecture and network characterization. Computer Networks (Amsterdam, Netherlands: 1999), 56(16): 3594–3608, November 14, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002654.

Hsieh:2012:DOM

[791] Yi-Ling Hsieh and Kuochen Wang. Dynamic overlay multicast for live multimedia streaming in urban VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 56(16):3609– 3628, November 14, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861200268X.

Palma:2012:LQE

[792] David Palma, Helder Araujo, and Marilia Curado. Link quality estimation in wireless multi-hop networks using Kernel based methods. Computer Networks (Amsterdam, Netherlands: 1999), 56(16): 3629–3638, November 14, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002691.

Kao:2012:SMI

[793] Chien-Chi Kao, Shun-Ren Yang, and Hsin-Chen Chen. A sleep-mode interleaving algorithm for layered-video multicast services in IEEE 802.16e networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(16): 3639–3654, November 14, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002770.

Anonymous:2012:EBp

[794] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(16):??, November 14, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612003180.

Coen-Porisini:2012:IDQ

[795] Alberto Coen-Porisini and Sabrina Sicari. Improving data quality using a cross layer protocol in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(17): 3655–3665, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002836.

Capone:2012:OFE

[796] Antonio Capone, Daniele Corti, Luca Gianoli, and Brunilde Sansó. An optimization framework for the energy management of carrier Ethernet networks with Multiple Spanning Trees. Computer Networks (Amsterdam, Netherlands: 1999), 56(17): 3666–3681, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002848.

Kim:2012:MLM

[797] Sungwook Kim. Multi-leader multifollower Stackelberg model for cognitive radio spectrum sharing scheme. Computer Networks (Amsterdam, Netherlands: 1999), 56(17):3682– 3692, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002940.

Kuo:2012:MHM

[798] Wen-Hsing Kuo and Ting-Yang Lin. Multi-hop multicast path construction in modern wireless relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(17): 3693–3704, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612002952.

Detti:2012:SWI

[799] A. Detti, M. Pomposini, N. Blefari-Melazzi, and S. Salsano. Supporting the Web with an information centric network that routes by name. Computer Networks (Amsterdam, Netherlands: 1999), 56(17): 3705–3722, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002964.

Yuksel:2012:REC

[800] M. Yuksel, K. K. Ramakrishnan, S. Kalyanaraman, J. D. Houle, and R. Sadhvani. Required extra capacity: a comparative estimation of overprovisioning needed for a classless IP backbone. Computer Networks (Amsterdam, Netherlands: 1999), 56(17): 3723–3743, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002976.

Meng:2012:OSU

[801] Xianfu Meng, Yanli Wang, and Yalin Ding. An optimized strategy for update path selection in unstructured P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(17): 3744–3755, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003106.

Palomar:2012:PDG

[802] Esther Palomar, Almudena Alcaide, Arturo Ribagorda, and Yan Zhang.

The Peer’s Dilemma: a general framework to examine cooperation in pure peer-to-peer systems. Computer Networks (Amsterdam, Netherlands: 1999), 56(17):3756–3766, November 30, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003118.

Feng:2012:IMS

[803] Li Feng and Jianqing Li. Integermultiple-spacing-based scheduling for multimedia applications in IEEE 802.11e HCCA wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(17):3767– 3782, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200312X.

Jin:2012:RCH

[804] Jiong Jin, Marimuthu Palaniswami, and Bhaskar Krishnamachari. Rate control for heterogeneous wireless sensor networks: Characterization, algorithms and performance. Computer Networks (Amsterdam, Netherlands: 1999), 56(17):3783–3794, November 30, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003131.

Bernsen:2012:RRI

[805] James Bernsen and D. Manivannan. RIVER: a reliable inter-vehicular routing protocol for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 56(17): 3795–3807, November 30, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003167.

Anonymous:2012:EBq

[806] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(17):??, November 30, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612003404.

Husemann:2012:WWL

[807] Dirk Husemann and Harry Rudin. The WEB we live in. Computer Networks (Amsterdam, Netherlands: 1999), 56 (18):3809–3810, December 17, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003672.

Allison:2012:WE

[808] Colin Allison, Alan Miller, Iain Oliver, Rosa Michaelson, and Thanassis Tiropanis. The Web in education. Computer Networks (Amsterdam, Netherlands: 1999), 56(18): 3811–3824, December 17, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003489.

Brin:2012:RAL

[809] Sergey Brin and Lawrence Page. Reprint of: The anatomy of a largescale hypertextual Web search engine. Computer Networks (Amsterdam, Netherlands: 1999), 56(18): 3825–3833, December 17, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003611.

Camenisch:2012:IP

[810] Jan Camenisch. Information privacy?! Computer Networks (Amsterdam, Netherlands: 1999), 56(18): 3834–3848, December 17, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003660.

Cho:2012:REC

[811] Junghoo Cho, Hector Garcia-Molina, and Lawrence Page. Reprint of: Efficient crawling through URL ordering. Computer Networks (Amsterdam, Netherlands: 1999), 56(18): 3849–3858, December 17, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861200360X.

Hall:2012:WEW

[812] Wendy Hall and Thanassis Tiropanis. Web evolution and Web Science. Computer Networks (Amsterdam, Netherlands: 1999), 56(18): 3859–3865, December 17, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003581.

Heidemann:2012:OSN

[813] Julia Heidemann, Mathias Klier, and Florian Probst. Online social networks: a survey of a global phenomenon. Computer Networks (Amsterdam, Netherlands: 1999), 56(18): 3866–3878, December 17, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003088.

Hughes:2012:VWA

[814] Ian Hughes. Virtual worlds, augmented reality, blended reality. Computer Networks (Amsterdam, Netherlands: 1999), 56(18):3879–3885, December 17, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003477.

Odlyzko:2012:WHE

[815] Andrew Odlyzko. Web history and economics. Computer Networks (Amsterdam, Netherlands: 1999), 56(18): 3886–3890, December 17, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003659.

Sandhu:2012:SSW

[816] Ravi Sandhu. Speculations on the science of Web user security. Computer Networks (Amsterdam, Netherlands: 1999), 56(18): 3891–3895, December 17, 2012. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003647.

Anonymous:2012:EBr

[817] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 56(18):??, December 17, 2012. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612003842.

Akyildiz:2013:CE

[818] Ian F. Akyildiz and Harry Rudin. COMNET editorial for 2012. Computer Networks (Amsterdam, Netherlands: 1999), 57(1):1–2, January 16, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004185.

Papathanasiou:2013:DRR

[819] Christos Papathanasiou, Nikos Dimitriou, and Leandros Tassiulas. Dynamic radio resource and interference management for MIMO-OFDMA mobile broadband wireless access systems. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 3–16, January 16, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002824.

Long:2013:PMS

[820] Long Long and Ahmed E. Kamal. Protecting multicast services in optical Internet backbones. Computer Networks (Amsterdam, Netherlands: 1999), 57 (1):17–28, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003064.

Sardouk:2013:CMU

[821] Ahmad Sardouk, Majdi Mansouri, Leila Merghem-Boulahia, Dominique Gaiti, and Rana Rahim-Amoud. Crisis management using MAS-based wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (1):29–45, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200309X.

Shan:2013:PSP

[822] Dong Mei Shan, Kee Chaing Chua, Gurusamy Mohan, and Jian Qiu. Partial spatial protection for provisioning differentiated reliability in FSTR-based Metro Ethernet networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(1):46–60, January 16, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003155.

Nam:2013:EAB

[823] Seung Yeob Nam, Seong Joon Kim, Sihyung Lee, and Hyong S. Kim. Estimation of the available bandwidth ratio of a remote link or path segments. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 61–77, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003143.

Angrishi:2013:EES

[824] Kishore Angrishi. An end-to-end stochastic network calculus with effective bandwidth and effective capacity. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 78–84, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003258.

Jang:2013:ASM

[825] Beakcheol Jang, Jun Bum Lim, and Mihail L. Sichitiu. An asynchronous scheduled MAC protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (1):85–98, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003246.

An:2013:ECC

[826] Dognhyeok An, Honguk Woo, Hyunsoo Yoon, and Ikjun Yeom. Enhanced cooperative communication MAC for mobile wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(1):99–116, January 16, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003234.

Xu:2013:ECP

[827] Mingwei Xu, Meijia Hou, Dan Wang, and Jiahai Yang. An efficient critical protection scheme for intradomain routing using link characteristics. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 117–133, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003283.

Barkoosaraei:2013:DPI

[828] A. Mirsayar Barkoosaraei and A. Hamid Aghvami. Dynamic partitioning of IP-based wireless access networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 134–146, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003271.

Eshete:2013:GCF

[829] Addisu Eshete and Yuming Jiang. Generalizing the CHOKe flow protection. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 147–161, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200326X.

Mukherjee:2013:FBG

[830] Anwesha Mukherjee, Srimoyee Bhattacherjee, Sucheta Pal, and Debashis De. Femtocell based green power consumption methods for mobile network. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 162–178, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003295.

Fang:2013:VOV

[831] Weiwei Fang, Xiangmin Liang, Shengxin Li, Luca Chiaraviglio, and Naixue Xiong. VMPlanner: Optimizing virtual machine placement and traffic flow routing to reduce network power costs in cloud data centers. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 179–196, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003301.

Chen:2013:UBS

[832] Hui Chen, Yang Xiao, and Susan V. Vrbsky. An update-based step-wise optimal cache replacement for wireless data access. Computer Networks (Amsterdam, Netherlands: 1999), 57 (1):197–212, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003374.

Gregori:2013:DCI

[833] Enrico Gregori, Luciano Lenzini, and Chiara Orsini. k-Dense communities in the Internet AS-level topology graph. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 213–227, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003362.

Mtibaa:2013:FRC

[834] Abderrahmen Mtibaa and Khaled A. Harras. Fairness-related challenges in mobile opportunistic networking. Computer Networks (Amsterdam, Netherlands: 1999), 57(1):228– 242, January 16, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003350.

Yao:2013:VFI

[835] Guang Yao, Jun Bi, and Peiyao Xiao. VASE: Filtering IP spoofing traffic with agility. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 243–257, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003349.

Serrano:2013:CTO

[836] Pablo Serrano, Paul Patras, Andrea Mannocci, Vincenzo Mancuso, and Albert Banchs. Control theoretic optimization of 802.11 WLANs: Implementation and experimental evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 258–272, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003337.

Vieira:2013:LHR

[837] Fabio R. J. Vieira, José F. de Rezende, Valmir C. Barbosa, and Serge Fdida. Local heuristic for the refinement of multi-path routing in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 273–285, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003325.

Mbarushimana:2013:CLT

[838] C. Mbarushimana, A. Shahrabi, and T. Buggy. A cross-layer TCP enhancement in QoS-aware mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (1):286–301, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003386.

Amador:2013:MCV

[839] Julia Amador and Jesus R. Artalejo. Modeling computer virus with the BSDE approach. Computer Networks (Amsterdam, Netherlands: 1999), 57 (1):302–316, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003453.

Mahmoodi:2013:UTA

[840] Toktam Mahmoodi, Vasilis Friderikos, and Hamid Aghvami. Using traffic asymmetry to enhance TCP performance. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 317–329, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003465.

Unal:2013:FRB

[841] D. Unal and M. U. Caglayan. A formal role-based access control model for security policies in multi-domain mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (1):330–350, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003490.

Cui:2013:CSM

[842] Lin Cui and Weijia Jia. Cyclic stable matching for three-sided networking services. Computer Networks (Amsterdam, Netherlands: 1999), 57(1): 351–363, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003520.

Canberk:2013:AQB

[843] Berk Canberk. An adaptive and QoSbased spectrum awareness framework for CR networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (1):364–373, January 16, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003519.

Anonymous:2013:EBa

[844] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(1):??, January 16, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612004203.

Salles:2013:ECN

[845] Ronaldo Salles, Guofei Gu, and Morton Swimmer. Editorial for Computer Networks special issue on “Botnet Activity: Analysis, Detection and Shutdown”. Computer Networks (Amsterdam, Netherlands: 1999), 57(2):

375–377, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003076.

Silva:2013:BS

[846] Sérgio S. C. Silva, Rodrigo M. P. Silva, Raquel C. G. Pinto, and Ronaldo M. Salles. Botnets: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 57(2): 378–403, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003568.

Khosroshahy:2013:SBL

[847] Masood Khosroshahy, Mustafa K. Mehmet Ali, and Dongyu Qiu. The SIC botnet lifecycle model: a step beyond traditional epidemiological models. Computer Networks (Amsterdam, Netherlands: 1999), 57(2):404– 421, February 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003313.

Riccardi:2013:TRD

[848] Marco Riccardi, Roberto Di Pietro, Marta Palanques, and Jorge Aguilà Vila. Titans’ revenge: Detecting Zeus via its own flaws. Computer Networks (Amsterdam, Netherlands: 1999), 57 (2):422–435, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003556.

Sood:2013:DSU

[849] Aditya K. Sood, Richard J. Enbody, and Rohit Bansal. Dissecting SpyEye — Understanding the design of third generation botnets. Computer Networks (Amsterdam, Netherlands: 1999), 57(2): 436–450, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002666.

Caballero:2013:APR

[850] Juan Caballero and Dawn Song. Automatic protocol reverse-engineering: Message format extraction and field semantics inference. Computer Networks (Amsterdam, Netherlands: 1999), 57 (2):451–474, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002927.

Dietrich:2013:CCR

[851] Christian J. Dietrich, Christian Rossow, and Norbert Pohlmann. CoCoSpot: Clustering and recognizing botnet command and control channels using traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 57(2):475– 486, February 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002472.

Perdisci:2013:SFG

[852] Roberto Perdisci, Davide Ariu, and Giorgio Giacinto. Scalable fine-grained behavioral clustering of HTTP-based malware. Computer Networks (Amsterdam, Netherlands: 1999), 57(2): 487–500, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002678.

Lin:2013:GBR

[853] Hui-Tang Lin, Ying-You Lin, and Jui-Wei Chiang. Genetic-based realtime fast-flux service networks detection. Computer Networks (Amsterdam, Netherlands: 1999), 57(2): 501–513, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002915.

Huang:2013:EBH

[854] Chun-Ying Huang. Effective bot host detection based on network failure models. Computer Networks (Amsterdam, Netherlands: 1999), 57(2): 514–525, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002939.

Las-Casas:2013:SDS

[855] Pedro H. B. Las-Casas, Dorgival Guedes, Jussara M. Almeida, Artur Ziviani, and Humberto T. Marques-Neto. SpaDeS: Detecting spammers at the source network. Computer Networks (Amsterdam, Netherlands: 1999), 57 (2):526–539, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002897.

Yan:2013:PWH

[856] Guanhua Yan. Peri-Watchdog: Hunting for hidden botnets in the periphery of online social networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(2):540–555, February 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612002903.

Boshmaf:2013:DAS

[857] Yazan Boshmaf, Ildar Muslukhov, Konstantin Beznosov, and Matei Ripeanu. Design and analysis of a social botnet. Computer Networks (Amsterdam, Netherlands: 1999), 57(2): 556–578, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002150.

Hua:2013:BCC

[858] Jingyu Hua and Kouichi Sakurai. Botnet command and control based on Short Message Service and human mobility. Computer Networks (Amsterdam, Netherlands: 1999), 57(2): 579–597, February 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612002162.

Anonymous:2013:EBb

[859] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(2):??, February 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128613000236.

delaOliva:2013:PAE

[860] Antonio de la Oliva, Tito R. Vargas Hernández, Juan Carlos Guerri, José Alberto Hernández, and Pedro Reviriego. Performance analysis of Energy Efficient Ethernet on video streaming servers. Computer Networks (Amsterdam, Netherlands: 1999), 57 (3):599–608, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003507.

Tan:2013:JOC

[861] Evan Tan and Chun Tung Chou. Joint optimization of continuity and quality for streaming video. Computer Networks (Amsterdam, Netherlands: 1999), 57(3):609–621, February 26, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128612003532.

Han:2013:CLC

[862] Chong Han, Josep Miquel Jornet, Etimad Fadel, and Ian F. Akyildiz. A cross-layer communication module for the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 57(3): 622–633, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200357X.

Chu:2013:BBD

[863] Zi Chu, Steven Gianvecchio, Aaron Koehl, Haining Wang, and Sushil Jajodia. Blog or block: Detecting blog bots through behavioral biometrics. Computer Networks (Amsterdam, Netherlands: 1999), 57(3): 634–646, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003593.

Kothari:2013:MAC

[864] Kush Kothari and Matthew Wright. Mimic: an active covert channel that evades regularity-based detection. Computer Networks (Amsterdam, Netherlands: 1999), 57(3): 647–657, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003623.

Kamiyama:2013:DDC

[865] Noriaki Kamiyama. Designing data center network by analytic hierarchy process. Computer Networks (Amsterdam, Netherlands: 1999), 57(3): 658–667, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003635.

Wang:2013:DSU

[866] Yu Wang, Hari Krishna Garg, and Mehul Motani. Downlink scheduling for user equipment served by multiple mobile terminals in cellular systems. Computer Networks (Amsterdam, Netherlands: 1999), 57(3):

668–681, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003684.

Dumitras:2013:SUF

[867] Tudor Dumitras and Priya Narasimhan. A study of unpredictability in faulttolerant middleware. Computer Networks (Amsterdam, Netherlands: 1999), 57(3):682–698, February 26, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128612003696.

Lucas:2013:RFM

[868] Vincent Lucas, Jean-Jacques Pansiot, Dominique Grad, and Benoît Hilt. Robust and fair Multicast Congestion Control (M2C). Computer Networks (Amsterdam, Netherlands: 1999), 57 (3):699–724, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003702.

Baek:2013:ECM

[869] Joo-Young Baek, Jeong-Yoon Lee, and Young-Joo Suh. An elastic compensation model for frame-based scheduling algorithms in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(3): 725–740, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003714.

Kamiyama:2013:ALB

[870] Noriaki Kamiyama, Tatsuya Mori, and Ryoichi Kawahara. Autonomic load balancing of flow monitors. Computer Networks (Amsterdam, Netherlands: 1999), 57(3): 741–761, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003726.

Yousaf:2013:OTP

[871] Faqir Zarrar Yousaf and Christian Wietfeld. Optimizing throughput performance of FMIPv6 over legacy 802.11 networks using iterative scanning. Computer Networks (Amsterdam, Netherlands: 1999), 57(3): 762–781, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003738.

Hsiao:2013:HTH

[872] Yi-Mao Hsiao, Yuan-Sun Chu, Jeng-Farn Lee, and Jinn-Shyan Wang. A high-throughput and high-capacity IPv6 routing lookup system. Computer Networks (Amsterdam, Netherlands: 1999), 57(3):782–794, February 26, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861200374X.

Casalicchio:2013:MSP

[873] Emiliano Casalicchio and Luca Silvestri. Mechanisms for SLA provisioning in cloud-based service providers. Computer Networks (Amsterdam, Netherlands: 1999), 57(3):795–810, February 26, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128612003763.

Jamdagni:2013:RMT

[874] Aruna Jamdagni, Zhiyuan Tan, Xiangjian He, Priyadarsi Nanda, and Ren Ping Liu. RePIDS: a multi tier Real-time Payload-based Intrusion Detection System. Computer Networks (Amsterdam, Netherlands: 1999), 57 (3):811–824, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003544.

Khan:2013:CRA

[875] Reduan H. Khan and Jamil Y. Khan. A comprehensive review of the application characteristics and traffic requirements of a smart grid communications network. Computer Networks (Amsterdam, Netherlands: 1999), 57 (3):825–845, February 26, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003751.

Anonymous:2013:EBc

[876] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(3):??, February 26, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128613000406.

Liou:2013:MMC

[877] Ren-Huang Liou and Yi-Bing Lin. Mobility management with the centralbased location area policy. Computer Networks (Amsterdam, Netherlands: 1999), 57(4):847–857, March 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003775.

Wang:2013:ODE

[878] Wei-Tong Wang and Kuo-Feng Ssu. Obstacle detection and estimation in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(4):858–868, March 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003787.

Ling:2013:PLA

[879] Zhen Ling, Junzhou Luo, Wei Yu, Xinwen Fu, Weijia Jia, and Wei Zhao. Protocol-level attacks against Tor. Computer Networks (Amsterdam, Netherlands: 1999), 57(4): 869–886, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003799.

Lee:2013:ASB

[880] Jiunn-Jye Lee, Hann-Huei Chiou, Chia-Chang Hsu, and Chin-Laung Lei. An adaptive sector-based routing model over structured peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (4):887–896, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003805.

Kim:2013:CLI

[881] Juhoon Kim, Luigi Iannone, and Anja Feldmann. Caching Locator/ID mappings: an experimental scalability analysis and its implications. Computer Networks (Amsterdam, Netherlands: 1999), 57(4):897–909, March 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003817.

Chi:2013:BLP

[882] Kaikai Chi, Xiaohong Jiang, Yi hua Zhu, Jing Wang, and Yanjun Li. Block-level packet recovery with network coding for wireless reliable multicast. Computer Networks (Amsterdam, Netherlands: 1999), 57(4): 910–923, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003829.

Diallo:2013:CBP

[883] Mohamed Diallo, Vasilis Sourlas, Paris Flegkas, Serge Fdida, and Leandros Tassiulas. A content-based publish/ subscribe framework for large-scale content delivery. Computer Networks (Amsterdam, Netherlands: 1999), 57 (4):924–943, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612003921.

Chen:2013:HPR

[884] Xiuzhong Chen, Marc De Leenheer, Rui Wang, Chaitanya S. K. Vadrevu, Lei Shi, Jie Zhang, and Biswanath Mukherjee. High-performance routing for hose-based VPNs in multidomain backbone networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(4):944–953, March 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003933.

Camacho:2013:BXB

[885] Jose M. Camacho, Alberto García-Martínez, Marcelo Bagnulo, and Francisco Valera. BGP-XM: BGP eXtended Multipath for transit Autonomous Systems. Computer Networks (Amsterdam, Netherlands: 1999), 57(4): 954–975, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003957.

Secci:2013:EID

[886] Stefano Secci, Kunpen Liu, and Bijan Jabbari. Efficient inter-domain traffic engineering with transit-edge hierarchical routing. Computer Networks (Amsterdam, Netherlands: 1999), 57 (4):976–989, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003969.

Miao:2013:MVN

[887] Yuting Miao, Qiang Yang, Chunming Wu, Ming Jiang, and Jinzhou Chen.

Multicast virtual network mapping for supporting multiple description codingbased video applications. Computer Networks (Amsterdam, Netherlands: 1999), 57(4):990–1002, March 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003970.

El-Azouzi:2013:EFG

[888] Rachid El-Azouzi, Francesco De Pellegrini, Habib B. A. Sidi, and Vijay Kamble. Evolutionary forwarding games in delay tolerant networks: Equilibria, mechanism design and stochastic approximation. Computer Networks (Amsterdam, Netherlands: 1999), 57 (4):1003–1018, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003982.

BorgesVieira:2013:SSE

[889] Alex Borges Vieira, Rafael Barra de Almeida, Jussara Marques de Almeida, and Sérgio Vale Aguiar Campos. SimplyRep: a simple and effective reputation system to fight pollution in P2P live streaming. Computer Networks (Amsterdam, Netherlands: 1999), 57 (4):1019–1036, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612003994.

Luo:2013:TCB

[890] Xiaoyuan Luo, Yanlin Yan, Shaobao Li, and Xinping Guan. Topology control based on optimally rigid graph in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(4): 1037–1047, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004008.

Kwon:2013:CIE

[891] Sungoh Kwon and Neung-Hyung Lee. Cell ID extension in femtocell environments. Computer Networks (Amsterdam, Netherlands: 1999), 57(4): 1048–1062, March 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861200401X.

Anonymous:2013:EBd

[892] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(4):??, March 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000509.

Shan:2013:NLM

[893] Feng Shan, Weifa Liang, Jun Luo, and Xiaojun Shen. Network lifetime maximization for time-sensitive data gathering in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(5): 1063–1077, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004045.

Bauer:2013:SCR

[894] Christian Bauer. A Secure Correspondent Router Protocol for NEMO Route Optimization. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1078–1100, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612004057.

Donmez:2013:CAC

[895] Mehmet Yunus Donmez, Sinan Isik, and Cem Ersoy. Combined analysis of contention window size and duty cycle for throughput and energy optimization in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1101–1112, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612004069.

Raayatpanah:2013:MCM

[896] M. A. Raayatpanah, H. Salehi Fathabadi, B. H. Khalaj, and S. Khodayifar. Minimum cost multiple multicast network coding with quantized rates. Computer Networks (Amsterdam, Netherlands: 1999), 57(5): 1113–1123, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612004070.

Karasabun:2013:AND

[897] Efe Karasabun, Ibrahim Korpeoglu, and Cevdet Aykanat. Active node determination for correlated data gathering in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(5): 1124–1138, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004082.

Yun:2013:EVN

[898] Donggyu Yun, Jungseul Ok, Bongjhin Shin, Soobum Park, and Yung Yi. Embedding of virtual network requests over static wireless multihop networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(5): 1139–1152, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004094.

Yun:2013:PAI

[899] Ji-Hoon Yun. Performance analysis of IEEE 802.11 WLANs with rate adaptation in time-varying fading channels. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1153–1166, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004100.

Antoniou:2013:CCW

[900] Pavlos Antoniou, Andreas Pitsillides, Tim Blackwell, Andries Engelbrecht, and Loizos Michael. Congestion control in wireless sensor networks based on bird flocking behavior. Computer Networks (Amsterdam, Netherlands: 1999), 57(5):1167–1191, April 7, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612004112.

Kilic:2013:ALP

[901] Necati Kilic and V. Cagri Gungor. Analysis of low power wireless links in smart grid environments. Computer Networks (Amsterdam, Netherlands: 1999), 57(5):1192–1203, April 7, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612004136.

Detal:2013:RFB

[902] Gregory Detal, Christoph Paasch, Simon van der Linden, Pascal Mérindol, Gildas Avoine, and Olivier Bonaventure. Revisiting flow-based load balancing: Stateless path selection in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1204–1216, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861200415X.

Aleksic:2013:EEF

[903] Slavisa Aleksic, Margot Deruyck, Willem Vereecken, Wout Joseph, Mario Pickavet, and Luc Martens. Energy efficiency of femtocell deployment in combined wireless/optical access networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1217–1233, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612004173.

Baziana:2013:APE

[904] P. A. Baziana and I. E. Pountourakis. An access protocol for efficiency optimization in WDM networks: a propagation delay and collisions avoidance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1234–1252, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004318.

DAcunto:2013:BLP

[905] Lucia D’Acunto, Nitin Chiluka, Tamás Vinkó, and Henk Sips. BitTorrentlike P2P approaches for VoD: a comparative study. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1253–1276, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000029.

Casagrande:2013:FOD

[906] Daniele Casagrande, Pier Luca Montessoro, and Franco Blanchini. Fair and optimal dynamic admission control of elastic flows. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1277–1288, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004148.

Salah:2013:MBS

[907] Saeed Salah, Gabriel Maciá-Fernández, and Jesús E. Díaz-Verdejo. A modelbased survey of alert correlation techniques. Computer Networks (Amsterdam, Netherlands: 1999), 57(5):

1289–1317, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128612004124.

Esposito:2013:RPS

[908] Christian Esposito, Domenico Cotroneo, and Stefano Russo. On reliability in publish/subscribe services. Computer Networks (Amsterdam, Netherlands: 1999), 57(5): 1318–1343, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861200432X.

Wang:2013:CSS

[909] Wenye Wang and Zhuo Lu. Cyber security in the Smart Grid: Survey and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 57 (5):1344–1371, April 7, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000042.

Anonymous:2013:EBe

[910] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(5):??, April 7, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000625.

Csernai:2013:IUD

[911] Márton Csernai, András Gulyás, Attila Körösi, Balázs Sonkoly, and Gergely Biczók. Incrementally upgradable data center architecture using hyperbolic tessellations. Computer Networks (Amsterdam, Netherlands: 1999), 57 (6):1373–1393, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004033.

Anagnostopoulos:2013:MCC

[912] Christos Anagnostopoulos and Stathes Hadjiefthymiades. Multivariate context collection in mobile sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(6): 1394–1407, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000078.

Ke:2013:PEM

[913] Kai-Wei Ke and Chia-Hui Huang. Performance evaluation of multisource Application Layer Multicast (ALM): Theoretical and simulative aspects. Computer Networks (Amsterdam, Netherlands: 1999), 57(6): 1408–1424, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000091.

Hunkeler:2013:CCC

[914] Urs Hunkeler, Clemens Lombriser, Hong Linh Truong, and Beat Weiss. A case for centrally controlled wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (6):1425–1442, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000108.

Borraz-Sanchez:2013:OPR

[915] Conrado Borraz-Sánchez and Diego Klabjan. Optimal placement of reconfigurable optical add/drop multiplexers with packing, blocking, and signal loss. Computer Networks (Amsterdam, Netherlands: 1999), 57(6): 1443–1458, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000121.

Ahmed:2013:ERT

[916] Adel Ali Ahmed. An enhanced realtime routing protocol with load distribution for mobile wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(6): 1459–1473, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000133.

Fan:2013:NCS

[917] Linjun Fan, Yunxiang Ling, Tao Wang, Xiaomin Zhu, and Xiaoyong Tang. Novel clock synchronization algorithm of parametric difference for parallel and distributed simulations. Computer Networks (Amsterdam, Netherlands: 1999), 57(6):1474–1487, April 22, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000145.

Benfattoum:2013:QRT

[918] Youghourta Benfattoum, Steven Martin, and Khaldoun Al Agha. QoS for real-time reliable multicasting in wireless multi-hop networks using a Generation-Based Network Coding. Computer Networks (Amsterdam, Netherlands: 1999), 57(6): 1488–1502, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000157.

Amaldi:2013:EAI

[919] E. Amaldi, A. Capone, and L. G. Gianoli. Energy-aware IP traffic engineering with shortest path routing. Computer Networks (Amsterdam, Netherlands: 1999), 57(6): 1503–1517, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000169.

Mahapatra:2013:GFF

[920] Rajarshi Mahapatra, Antonio De Domenico, Rohit Gupta, and Emilio Calvanese Strinati. Green framework for future heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(6): 1518–1528, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000170.

Wang:2013:OOA

[921] Xiaoping Wang, Yunhao Liu, Zheng Yang, Kai Lu, and Jun Luo. OFA: an optimistic approach to conquer flip ambiguity in network localization. Computer Networks (Amsterdam, Netherlands: 1999), 57(6): 1529–1544, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000182.

Chiappetta:2013:ABA

[922] S. Chiappetta, C. Mazzariello, R. Presta, and S. P. Romano. An anomaly-based approach to the analysis of the social behavior of VoIP users. Computer Networks (Amsterdam, Netherlands: 1999), 57(6):1545–1559, April 22, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000194.

Kawahara:2013:MVR

[923] Ryoichi Kawahara, Tetsuya Takine, Tatsuya Mori, Noriaki Kamiyama, and Keisuke Ishibashi. Mean-variance relationship of the number of flows in traffic aggregation and its application to traffic management. Computer Networks (Amsterdam, Netherlands: 1999), 57 (6):1560–1576, April 22, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000200.

Anonymous:2013:EBf

[924] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(6):??, April 22, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000923.

Newaz:2013:ADA

[925] S. H. Shah Newaz, Ángel Cuevas, Gyu Myoung Lee, Noël Crespi, and Jun Kyun Choi. Adaptive delay-aware energy efficient TDM-PON. Computer Networks (Amsterdam, Netherlands: 1999), 57(7):1577–1596, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300011X.

Facchini:2013:DGS

[926] Christian Facchini, Oliver Holland, Fabrizio Granelli, Nelson L. S. da Fonseca, and Hamid Aghvami. Dynamic green self-configuration of 3G base stations using fuzzy cognitive maps. Computer Networks (Amsterdam, Netherlands: 1999), 57(7): 1597–1610, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000212.

Chu:2013:RTV

[927] Weibo Chu, Xiaohong Guan, Zhongmin Cai, and Lixin Gao. Real-time volume control for interactive network traffic replay. Computer Networks (Amsterdam, Netherlands: 1999), 57 (7):1611–1629, May 8, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000303.

Jung:2013:OSF

[928] Tobias Jung, Sylvain Martin, Mohamed Nassar, Damien Ernst, and Guy Leduc. Outbound SPIT filter with optimal performance guarantees. Computer Networks (Amsterdam, Netherlands: 1999), 57(7):1630–1643, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000315.

Lee:2013:ODD

[929] Jongwook Lee, Jin-Ghoo Choi, and Saewoong Bahk. Opportunistic downlink data delivery for mobile collaborative communities. Computer Networks (Amsterdam, Netherlands: 1999), 57 (7):1644–1655, May 8, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000327.

Lagkas:2013:AIF

[930] T. D. Lagkas, P. Sarigiannidis, and M. Louta. On analyzing the intraframe power saving potentials of the IEEE 802.16e downlink vertical mapping. Computer Networks (Amsterdam, Netherlands: 1999), 57(7): 1656–1673, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000339.

Wu:2013:IUC

[931] Weijie Wu, John C. S. Lui, and Richard T. B. Ma. On incentivizing upload capacity in P2P-VoD systems: Design, analysis and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 57(7):1674–1688, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000340.

Lai:2013:QAD

[932] Wei Kuang Lai and Chang-Lung Tang. QoS-aware downlink packet scheduling for LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (7):1689–1698, May 8, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000352.

Rottondi:2013:PPS

[933] Cristina Rottondi, Giacomo Verticale, and Antonio Capone. Privacypreserving smart metering with multiple data consumers. Computer Networks (Amsterdam, Netherlands: 1999), 57(7):1699–1713, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000364.

Torkestani:2013:EET

[934] Javad Akbari Torkestani. An energyefficient topology construction algorithm for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(7): 1714–1725, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000479.

Vakili:2013:QMV

[935] Ahmad Vakili and Jean-Charles Grégoire. QoE management for video conferencing applications. Computer Networks (Amsterdam, Netherlands: 1999), 57(7):1726–1738, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000558.

Anonymous:2013:EBg

[936] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(7):??, May 8, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001254.

Arkoulis:2013:OFC

[937] Stamatios Arkoulis, Evangelos Anifantis, Vasileios Karyotis, Symeon Papavassiliou, and Nikolaos Mitrou. On the optimal, fair and channelaware cognitive radio network reconfiguration. Computer Networks (Amsterdam, Netherlands: 1999), 57 (8):1739–1757, June 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000583.

Gyarmati:2013:FSY

[938] László Gyarmati, András Gulyás, Balázs Sonkoly, Tuan A. Trinh, and Gergely Biczók. Free-scaling your data center. Computer Networks (Amsterdam, Netherlands: 1999), 57 (8):1758–1773, June 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000674.

Lin:2013:EHE

[939] Gongqi Lin, Sieteng Soh, Kwan-Wu Chin, and Mihai Lazarescu. Efficient heuristics for energy-aware routing in networks with bundled links. Computer Networks (Amsterdam, Netherlands: 1999), 57(8):1774–1788, June 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000686.

Bacci:2013:QAG

[940] Giacomo Bacci and Marco Luise. QoSaware game-theoretic rate & power control for CDMA wireless communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (8):1789–1804, June 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000698.

Yang:2013:MEM

[941] Shengbo Yang, Chai Kiat Yeo, and Bu Sung Lee. MaxCD: Efficient multiflow scheduling and cooperative downloading for improved highway drivethru Internet systems. Computer Networks (Amsterdam, Netherlands: 1999), 57(8):1805–1820, June 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000741.

Rondinone:2013:CBF

[942] Michele Rondinone and Javier Gozalvez. Contention-based forwarding with multi-hop connectivity awareness in vehicular ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(8):1821–1837, June 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000753.

Carofiglio:2013:RLE

[943] Giovanna Carofiglio, Luca Muscariello, Dario Rossi, Claudio Testa, and Silvio Valenti. Rethinking the Low Extra Delay Background Transport (LED-BAT) Protocol. Computer Networks (Amsterdam, Netherlands: 1999), 57 (8):1838–1852, June 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000765.

Czirkos:2013:SBK

[944] Zoltán Czirkos and Gábor Hosszú. Solution for the broadcasting in the Kademlia peer-to-peer overlay. Computer Networks (Amsterdam, Netherlands: 1999), 57(8): 1853–1862, June 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000777.

Iellamo:2013:PDI

[945] Stefano Iellamo, Lin Chen, and Marceau Coupechoux. Proportional and double imitation rules for spectrum access in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(8): 1863–1879, June 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000789.

Pires:2013:ILN

[946] Alexandre A. Pires and José F. de Rezende. Independent links: a new approach to increase spatial reuse in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (8):1880–1893, June 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000844.

Anonymous:2013:EBh

[947] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(8):??, June 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001345.

Fang:2013:PCL

[948] Xi Fang, Dejun Yang, and Guoliang Xue. Pathbook: Cross-layer optimization for full-duplex wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(9): 1895–1912, June 19, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000571.

Ammari:2013:EDT

[949] Habib M. Ammari. On the energydelay trade-off in geographic forwarding in always-on wireless sensor networks: a multi-objective optimization problem. Computer Networks (Amsterdam, Netherlands: 1999), 57 (9):1913–1935, June 19, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000832.

Ficek:2013:ATM

[950] Michal Ficek, Tomás Pop, and Lukás Kencl. Active tracking in mobile networks: an in-depth view. Computer Networks (Amsterdam, Netherlands: 1999), 57(9):1936–1954, June 19, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613000996.

Militano:2013:FCC

[951] L. Militano, A. Iera, and F. Scarcello. A fair cooperative contentsharing service. Computer Networks (Amsterdam, Netherlands: 1999), 57 (9):1955–1973, June 19, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861300100X.

Babarczi:2013:RSD

[952] Péter Babarczi, Gergely Biczók, Harald Øverby, János Tapolcai, and Péter Soproni. Realization strategies of dedicated path protection: a bandwidth cost perspective. Computer Networks (Amsterdam, Netherlands: 1999), 57 (9):1974–1990, June 19, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001011.

Alsalih:2013:PCT

[953] Waleed Alsalih, Kashif Ali, and Hossam Hassanein. A power control technique for anti-collision schemes in RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 57 (9):1991–2003, June 19, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001023.

Coiro:2013:EAT

[954] A. Coiro, M. Listanti, A. Valenti, and F. Matera. Energy-aware traffic engineering: a routing-based distributed solution for connection-oriented IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(9): 2004–2020, June 19, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001035.

Botero:2013:GNN

[955] Juan Felipe Botero and Xavier Hesselbach. Greener networking in a network virtualization environment. Computer Networks (Amsterdam, Netherlands: 1999), 57(9):2021–2039, June 19, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001151.

Fahad:2013:TES

[956] Adil Fahad, Zahir Tari, Ibrahim Khalil, Ibrahim Habib, and Hussein Alnuweiri. Toward an efficient and scalable feature selection approach for Internet traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 57 (9):2040–2057, June 19, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001163.

Gramatikov:2013:SMP

[957] S. Gramatikov, F. Jaureguizar, J. Cabrera, and N. García. Stochastic modelling of peer-assisted VoD streaming in managed networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(9):2058–2074, June 19, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001175.

Logota:2013:AMC

[958] Evariste Logota, Carlos Campos, Susana Sargento, and Augusto Neto. Advanced multicast class-based bandwidth over-provisioning. Computer Networks (Amsterdam, Netherlands: 1999), 57(9):2075–2092, June 19, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001217.

Aceto:2013:CMS

[959] Giuseppe Aceto, Alessio Botta, Walter de Donato, and Antonio Pescapè. Cloud monitoring: a survey. Computer Networks (Amsterdam, Netherlands:

1999), 57(9):2093–2115, June 19, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001084.

Anonymous:2013:EBi

[960] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(9):??, June 19, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001461.

Eidenbenz:2013:ECN

[961] Stephan Eidenbenz, Madhav Marathe, and Arunabha Sen. Editorial for Computer Networks special issue on “Towards a Science of Cyber Security”. Computer Networks (Amsterdam, Netherlands: 1999), 57(10): 2119–2120, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001515.

Kuhlman:2013:COP

[962] Chris J. Kuhlman, V. S. Anil Kumar, and S. S. Ravi. Controlling opinion propagation in online networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(10): 2121–2132, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000893.

Nguyen:2013:AMC

[963] Nam P. Nguyen, Guanhua Yan, and My T. Thai. Analysis of misinformation containment in online social networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (10):2133–2146, July 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001126.

Wang:2013:ACI

[964] Yi Wang, Marios Iliofotou, Michalis Faloutsos, and Bin Wu. Analyzing Communication Interaction Networks (CINs) in enterprises and inferring hierarchies. Computer Networks (Amsterdam, Netherlands: 1999), 57 (10):2147–2158, July 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001187.

Almasizadeh:2013:SMA

[965] Jaafar Almasizadeh and Mohammad Abdollahi Azgomi. A stochastic model of attack process for the evaluation of security metrics. Computer Networks (Amsterdam, Netherlands: 1999), 57(10):2159–2180, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000856.

Caskurlu:2013:AMR

[966] Bugra Caskurlu, Ashish Gehani, Cemal Cagatay Bilgin, and K. Subramani. Analytical models for risk-based intrusion response. Computer Networks (Amsterdam, Netherlands: 1999), 57 (10):2181–2192, July 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000881.

Ben-Porat:2013:ECB

[967] Udi Ben-Porat, Anat Bremler-Barr, and Hanoch Levy. On the exploitation of CDF based wireless scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 57(10): 2193–2205, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000376.

Julisch:2013:UOC

[968] Klaus Julisch. Understanding and overcoming cyber security anti-patterns. Computer Networks (Amsterdam, Netherlands: 1999), 57(10):2206– 2211, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000388.

Anonymous:2013:SIS

[969] Anonymous. Special issue on “Security and identity architecture for the future Internet”. Computer Networks (Amsterdam, Netherlands: 1999), 57 (10):2215–2217, July 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001795.

Leva:2013:ABN

[970] Tapio Levä, Miika Komu, Ari Keränen, and Sakari Luukkainen. Adoption barriers of network layer protocols: the case of host identity protocol. Computer Networks (Amsterdam, Netherlands: 1999), 57(10): 2218–2232, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000480.

Torroglosa-Garcia:2013:IOW

[971] Elena Torroglosa-García, Antonio D. Pérez-Morales, Pedro Martinez-Julia, and Diego R. Lopez. Integration of the OAuth and Web Service family security standards. Computer Networks (Amsterdam, Netherlands: 1999), 57 (10):2233–2249, July 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001138.

Xiang:2013:SWY

[972] Yang Xiang, Xingang Shi, Jianping Wu, Zhiliang Wang, and Xia Yin. Sign what you really care about — Secure BGP AS-paths efficiently. Computer Networks (Amsterdam, Netherlands: 1999), 57(10):2250–2265, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000030.

Roman:2013:FCS

[973] Rodrigo Roman, Jianying Zhou, and Javier Lopez. On the features and challenges of security and privacy in distributed Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 57(10):2266–2279, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000054.

Martinez-Julia:2013:BSI

[974] Pedro Martinez-Julia and Antonio F. Skarmeta. Beyond the separation of identifier and locator: Building an identity-based overlay network architecture for the Future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 57(10): 2280–2300, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000066.

Anonymous:2013:EBj

[975] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(10):??, July 5, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001540.

Cappanera:2013:OJR

[976] Paola Cappanera, Luciano Lenzini, Alessandro Lori, Giovanni Stea, and Gigliola Vaglini. Optimal joint routing and link scheduling for real-time traffic in TDMA Wireless Mesh Networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(11): 2301–2312, August 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300008X.

Sengupta:2013:APN

[977] Soham Sengupta. An approach to provide a network layer security model with QR code generated with shuffled GPS parameters as embedded keys traveling over Internet using existing IPv4 mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 57 (11):2313–2330, August 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300114X.

Tarissan:2013:TBG

[978] Fabien Tarissan, Bruno Quoitin, Pascal Mérindol, Benoit Donnet, Jean-Jacques Pansiot, and Matthieu Latapy. Towards a bipartite graph modeling of the Internet topology. Computer Networks (Amsterdam, Netherlands: 1999), 57(11): 2331–2347, August 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001199.

Liu:2013:MIS

[979] Xiao-Yang Liu, Kai-Liang Wu, Yanmin Zhu, Linghe Kong, and Min-You Wu. Mobility increases the surface coverage of distributed sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(11): 2348–2363, August 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001205.

Nooruzzaman:2013:CRP

[980] Md. Nooruzzaman, Osanori Koyama, and Yutaka Katsuyama. Congestion removing performance of stackable ROADM in WDM networks under dynamic traffic. Computer Networks (Amsterdam, Netherlands: 1999), 57 (11):2364–2373, August 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001230.

Stojanovic:2013:SCI

[981] Mirjana D. Stojanovic, Aleksandra M. Kostic-Ljubisavljevic, and Vesna M. Radonjic-Djogatovic. SLA-controlled interconnection charging in next generation networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (11):2374–2394, August 5, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001321.

Turner:2013:SIF

[982] Kenneth J. Turner and Evan H. Magill. Special issue on feature interaction in communications and software systems. Computer Networks (Amsterdam, Netherlands: 1999), 57(12): 2395–2398, August 20, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001527.

Apel:2013:FID

[983] Sven Apel, Alexander von Rhein, Thomas Thüm, and Christian Kästner. Feature-interaction detection based on feature-based specifications. Computer Networks (Amsterdam, Netherlands: 1999), 57(12):2399–2409, August 20, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128613001102.

Kolberg:2013:FIF

[984] M. Kolberg, J. F. Buford, K. Dhara, X. Wu, and V. Krishnaswamy. Feature interaction in a federated communications-enabled collaboration platform. Computer Networks (Amsterdam, Netherlands: 1999), 57(12): 2410–2428, August 20, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861300087X.

Maternaghan:2013:PCH

[985] Claire Maternaghan and Kenneth J. Turner. Policy conflicts in home automation. Computer Networks (Amsterdam, Netherlands: 1999), 57(12): 2429–2441, August 20, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001114.

Nakamura:2013:CIR

[986] Masahide Nakamura, Kousuke Ikegami, and Shinsuke Matsumoto. Considering impacts and requirements for better understanding of environment interactions in home network services. Computer Networks (Amsterdam, Netherlands: 1999), 57(12): 2442–2453, August 20, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001096.

Tun:2013:SSF

[987] Thein Than Tun, Robin Laney, Yijun Yu, and Bashar Nuseibeh. Specifying software features for composition: a tool-supported approach.

Computer Networks (Amsterdam, Netherlands: 1999), 57(12):2454– 2464, August 20, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613000868.

Anonymous:2013:EBk

[988] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(12):??, August 20, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128613002041.

Kim:2013:PAN

[989] Yusung Kim and Ikjun Yeom. Performance analysis of in-network caching for content-centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2465–2482, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300090X.

Chang:2013:TEE

[990] Chao-Tsun Chang, Chih-Yung Chang, Tzu-Lin Wang, and Yun-Jung Lu. Throughput enhancement by exploiting spatial reuse opportunities with smart antenna systems in wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (13):2483–2498, September 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001229.

Serror:2013:MMP

[991] Jérémy Serror, Hui Zang, and Jean C. Bolot. Measurement and modeling of paging channel overloads on a cellular network. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2499–2513, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001308.

Lin:2013:EAR

[992] Hui-Tang Lin, Ying-You Lin, and Chung-Jui Sun. Efficient and adaptive resource scheduling in IEEE 802.16j transparent relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(13):2514– 2535, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861300131X.

Wehmuth:2013:DDA

[993] Klaus Wehmuth and Artur Ziviani. DACCER: Distributed Assessment of the Closeness CEntrality Ranking in complex networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (13):2536–2548, September 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001412.

Jeong:2013:TTB

[994] Jaehoon (Paul) Jeong, Tian He, and David H. C. Du. TMA: Trajectorybased Multi-Anycast forwarding for efficient multicast data delivery in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 57 (13):2549–2563, September 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001424.

Boiardi:2013:RPE

[995] Silvia Boiardi, Antonio Capone, and Brunilde Sansó. Radio planning of energy-aware cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2564–2577, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001436.

Tuncer:2013:PAV

[996] Hasan Tuncer, Andres Kwasinski, and Nirmala Shenoy. Performance analysis of Virtual Mobility Domain scheme vs. IPv6 mobility protocols. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2578–2596, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001448.

Acar:2013:SPA

[997] Tolga Acar, Mira Belenkiy, and Alptekin Küpçü. Single password authentication. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2597–2614, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001667.

Lee:2013:ORA

[998] Jeng-Wei Lee, I-Hsun Chuang, Win-Bin Huang, and Yau-Hwang Kuo. An opportunistic resource allocation approach for mixed QoS and non-QoS connections in OFDMA wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2615–2627, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001709.

Shin:2013:ENH

[999] Seungwon Shin, Zhaoyan Xu, and Guofei Gu. EFFORT: a new hostnetwork cooperated framework for efficient and effective bot malware detection. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2628–2642, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001783.

Chang:2013:DCR

[1000] Ben-Jye Chang, Ying-Hsin Liang, and Yu-Hsien Lee. Dynamic-cost-reward connection admission control for maximizing system reward in 4G wireless multihop relaying networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2643–2655, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001825.

CastroFernandes:2013:SAH

[1001] Natalia Castro Fernandes, Marcelo Duffles Donato Moreira, and Otto Carlos Muniz Bandeira Duarte. Safeguarding ad hoc networks with a self-organized membership control system. Computer Networks (Amsterdam, Netherlands: 1999), 57(13): 2656–2674, September 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001886.

Anonymous:2013:EBl

[1002] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(13):??, September 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128613002296.

Bellalta:2013:PAC

[1003] B. Bellalta, A. Faridi, D. Staehle, J. Barcelo, A. Vinel, and M. Oliver. Performance analysis of CSMA/CA protocols with multi-packet transmission. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2675–2688, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001710.

Zhang:2013:CAM

[1004] Pengfei Zhang, Gaoxi Xiao, and Hwee-Pink Tan. Clustering algorithms for maximizing the lifetime of wireless sensor networks with energy-harvesting sensors. Computer Networks (Amsterdam, Netherlands: 1999), 57(14):

2689–2704, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001874.

Chien:2013:CRC

[1005] Hung-Yu Chien. Combining Rabin cryptosystem and error correction codes to facilitate anonymous authentication with un-traceability for lowend devices. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2705–2717, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001898.

Camacho:2013:GDF

[1006] José Camacho, Pablo Padilla, Pedro García-Teodoro, and Jesús Díaz-Verdejo. A generalizable dynamic flow pairing method for traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2718–2732, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001904.

Ng:2013:UAM

[1007] Hai-Heng Ng, Wee-Seng Soh, and Mehul Motani. An underwater acoustic MAC protocol using reverse opportunistic packet appending. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2733–2751, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001916.

Bouali:2013:CMF

[1008] F. Bouali, O. Sallent, J. Pérez-Romero, and R. Agustí. A cognitive management framework for spectrum selection. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2752–2765, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001928.

Diarrassouba:2013:HWB

[1009] Ibrahima Diarrassouba, Ali Lourimi, Ali Ridha Mahjoub, and Habib Youssef. Hose workload based exact algorithm for the optimal design of virtual private networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(14):2766– 2774, October 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300193X.

Tschorsch:2013:APP

[1010] Florian Tschorsch and Björn Scheuermann. An algorithm for privacypreserving distributed user statistics. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2775–2787, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001941.

Wang:2013:ELM

[1011] Ze Wang, Yunlong Wang, Maode Ma, and Jigang Wu. Efficient localization for mobile sensor networks based on constraint rules optimized Monte Carlo method. Computer Networks (Amsterdam, Netherlands: 1999), 57 (14):2788–2801, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001953.

Rossini:2013:FAT

[1012] Giuseppe Rossini, Dario Rossi, Christophe Betoule, Remi Clavier, and Gilles Thouenon. FIB Aplasia through probabilistic routing and autoforwarding. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2802–2816, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001965.

Gasior:2013:PON

[1013] Dariusz Gasior and Maciej Drwal. Pareto-optimal Nash equilibrium in capacity allocation game for self-managed networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2817–2832, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613001977.

Malandrino:2013:PLW

[1014] Delfina Malandrino and Vittorio Scarano. Privacy leakage on the Web: Diffusion and countermeasures. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2833–2855, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001989.

Hua:2013:ESA

[1015] Kai-Lung Hua, Ge-Ming Chiu, Hsing-Kuo Pao, and Yi-Chi Cheng. An efficient scheduling algorithm for scalable video streaming over P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(14): 2856–2868, October 4, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613001990.

Anonymous:2013:EBm

[1016] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(14):??, October 4, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128613002673.

Ryoo:2013:LBS

[1017] Sunheui Ryoo, Changhee Joo, and Saewoong Bahk. Location-based spectrum allocation and partitioning scheme for cross-tier interference mitigation in macro-femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 2869–2879, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128612004021.

Xu:2013:GDC

[1018] Mingwei Xu, Yunfei Shang, Dan Li, and Xin Wang. Greening data center networks with throughputguaranteed power-aware routing. Computer Networks (Amsterdam, Netherlands: 1999), 57(15):2880–2899, October 29, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128612004161.

Meng:2013:ADT

[1019] Xiang Meng, Pui-Sze Tsang, and King-Shan Lui. Analysis of distribution time of multiple files in a P2P network. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 2900–2915, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002004.

Chen:2013:MSM

[1020] Yuh-Rong Chen, Sridhar Radhakrishnan, Sudarshan Dhall, and Suleyman Karabuk. On multi-stream multisource multicast routing. Computer Networks (Amsterdam, Netherlands: 1999), 57(15):2916–2930, October 29, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128613002016.

Chiaraviglio:2013:SME

[1021] Luca Chiaraviglio, Antonio Cianfrani, Esther Le Rouzic, and Marco Polverini. Sleep modes effectiveness in backbone networks with limited configurations. Computer Networks (Amsterdam, Netherlands: 1999), 57(15):

2931–2948, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002028.

Pham:2013:OES

[1022] Duc Minh Pham and Syed Mahfuzul Aziz. Object extraction scheme and protocol for energy efficient image communication over wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 2949–2960, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002144.

Meerja:2013:MAP

[1023] Khalim Amjad Meerja, Pin-Han Ho, Bin Wu, and Hsiang-Fu Yu. Media access protocol for a coexisting cognitive femtocell network. Computer Networks (Amsterdam, Netherlands: 1999), 57 (15):2961–2975, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002168.

Xiang:2013:DPA

[1024] Xudong Xiang, Jianxiong Wan, Chuang Lin, and Xin Chen. A dynamic programming approximation for downlink channel allocation in cognitive femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 2976–2991, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861300217X.

Mouttappa:2013:UPT

[1025] Pramila Mouttappa, Stephane Maag, and Ana Cavalli. Using passive testing based on symbolic execution and slicing techniques: Application to the validation of communication protocols. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 2992–3008, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002181.

Hu:2013:TEV

[1026] Chao Hu, Ming Chen, and Changyou Xing. Towards efficient video chunk dissemination in peer-to-peer live streaming. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 3009–3024, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002193.

Tosi:2013:DCB

[1027] Stefania Tosi, Sara Casolari, and Michele Colajanni. Data clustering based on correlation analysis applied to highly variable domains. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 3025–3038, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861300220X.

Zhou:2013:OAC

[1028] Chao Zhou, Xinggong Zhang, and Zongming Guo. Optimal adaptive channel scheduling for scalable video broadcasting over MIMO wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 3039–3050, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002260.

Chiaraviglio:2013:MSM

[1029] Luca Chiaraviglio, Delia Ciullo, Marco Mellia, and Michela Meo. Modeling sleep mode gains in energy-aware networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 3051–3066, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002272.

Habak:2013:ODB

[1030] Karim Habak, Moustafa Youssef, and Khaled A. Harras. An optimal deployable bandwidth aggregation system. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 3067–3080, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300234X.

Kamiyama:2013:OPA

[1031] Noriaki Kamiyama, Ryoichi Kawahara, and Haruhisa Hasegawa. Optimum profit allocation in coalitional VoD service. Computer Networks (Amsterdam, Netherlands: 1999), 57(15):

3081–3097, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002351.

Sicari:2013:DED

[1032] Sabrina Sicari, Alberto Coen-Porisini, and Roberto Riggio. DARE: evaluating Data Accuracy using node REputation. Computer Networks (Amsterdam, Netherlands: 1999), 57(15): 3098–3111, October 29, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002363.

Anonymous:2013:EBn

[1033] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 57(15):??, October 29, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128613002909.

Detti:2013:ESI

[1034] Andrea Detti, Diego Perino, Mario Gerla, and Yanghee Choi. Editorial: Special issue on Information Centric Networking. Computer Networks (Amsterdam, Netherlands: 1999), 57(16): 3113–3115, November 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003307.

Carofiglio:2013:CDT

[1035] G. Carofiglio, G. Morabito, L. Muscariello, I. Solis, and M. Varvello.

From content delivery today to information centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 57(16):3116–3127, November 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002156.

Zhang:2013:CIC

[1036] Guoqiang Zhang, Yang Li, and Tao Lin. Caching in information centric networking: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 57(16): 3128–3141, November 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002235.

Lee:2013:UAN

[1037] HyunYong Lee and Akihiro Nakao.

User-assisted in-network caching in information-centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 57(16):3142–3153, November 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002247.

Eum:2013:PBR

[1038] S. Eum, K. Nakauchi, M. Murata, Y. Shoji, and N. Nishinaga. Potential based routing as a secondary best-effort routing for Information Centric Networking (ICN). Computer Networks (Amsterdam, Netherlands: 1999), 57 (16):3154–3164, November 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002739.

Wang:2013:GAN

[1039] Yi Wang, Huichen Dai, Ting Zhang, Wei Meng, Jindou Fan, and Bin Liu. GPU-accelerated name lookup with component encoding. Computer Networks (Amsterdam, Netherlands: 1999), 57(16):3165–3177, November 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002223.

Conti:2013:LMD

[1040] Mauro Conti, Paolo Gasti, and Marco Teoli. A lightweight mechanism for detection of cache pollution attacks in Named Data Networking. Computer Networks (Amsterdam, Netherlands: 1999), 57(16): 3178–3191, November 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002818.

Wahlisch:2013:BDP

[1041] Matthias Wählisch, Thomas C. Schmidt, and Markus Vahlenkamp. Backscatter from the data plane — Threats to stability and security in information-centric network infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 57(16): 3192–3206, November 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002259.

Salsano:2013:ICN

[1042] S. Salsano, N. Blefari-Melazzi, A. Detti, G. Morabito, and L. Veltri. Information centric networking over SDN and OpenFlow: Architectural aspects and experiments on the OFE-LIA testbed. Computer Networks (Amsterdam, Netherlands: 1999), 57(16): 3207–3221, November 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002727.

Amadeo:2013:ECC

[1043] Marica Amadeo, Claudia Campolo, and Antonella Molinaro. Enhancing content-centric networking for vehicular environments. Computer Networks (Amsterdam, Netherlands: 1999), 57 (16):3222–3234, November 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002211.

Bandara:2013:DMU

[1044] H. M. N. Dilum Bandara and Anura P.

Jayasumana. Distributed, multi-user, multi-application, and multi-sensor data fusion over named data networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(16): 3235–3248, November 13, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002806.

Tagger:2013:RAE

[1045] Ben Tagger, Dirk Trossen, Alexandros Kostopoulos, Stuart Porter, and George Parisis. Realising an application environment for informationcentric networking. Computer Networks (Amsterdam, Netherlands: 1999), 57(16):3249–3266, November 13, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300279X.

Vieira:2013:FLE

[1046] Luiz Filipe M. Vieira, Mario Gerla, and Archan Misra. Fundamental limits on end-to-end throughput of network coding in multi-rate and multicast wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3267–3275, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002375.

Marsan:2013:ESM

[1047] Marco Ajmone Marsan, Luca Chiaraviglio, Delia Ciullo, and Michela Meo. On the effectiveness of single and multiple base station sleep modes in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3276–3290, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002429.

Chen:2013:CAH

[1048] Xiaohui Chen, Jing Xu, Wei Yuan, Wei Liu, and Wenqing Cheng. Channel assignment in heterogeneous multi-radio multi-channel wireless networks: a game theoretic approach. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3291–3299, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002430.

Baralis:2013:NCB

[1049] Elena Baralis, Andrea Bianco, Tania Cerquitelli, Luca Chiaraviglio, and Marco Mellia. NetCluster: a clustering-based framework to analyze Internet passive measurements data. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3300–3315, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002454.

Carrano:2013:NBD

[1050] Ricardo C. Carrano, Diego Passos, Luiz C. S. Magalhães, and Célio V. N. Albuquerque. Nested block designs: Flexible and efficient schedule-based asynchronous duty cycling. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3316–3326, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002466.

Liu:2013:CEC

[1051] Honghuing Liu, Patrick P. C. Lee, and John C. S. Lui. On the credit evolution of credit-based incentive protocols in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3327–3343, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002478.

Matsui:2013:DPA

[1052] Genki Matsui, Takuji Tachibana, Yukinori Nakamura, and Kenji Sugimoto. Distributed power adjustment based on control theory for cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3344–3356, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861300248X.

Silva:2013:MST

[1053] João Marco C. Silva, Paulo Carvalho, and Solange Rito Lima. A multiadaptive sampling technique for cost-effective network measurements. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3357–3369, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002491.

Faigl:2013:SAE

[1054] Zoltán Faigl, László Bokor, Jani Pellikka, and Andrei Gurtov. Suitability analysis of existing and new authentication methods for future 3GPP Evolved Packet Core. Computer Networks (Amsterdam, Netherlands: 1999), 57(17):3370– 3388, December 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002508.

Borgia:2013:EER

[1055] Eleonora Borgia, Giuseppe Anastasi, and Marco Conti. Energy efficient and reliable data delivery in urban sensing applications: a performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3389–3409, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300251X.

Hagelstein:2013:IFI

[1056] B. Hagelstein, M. Abolhasan, D. Franklin, and F. Safaei. Improving fairness in IEEE 802.11 networks using MAC layer opportunistic retransmission. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3410–3427, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002521.

Chen:2013:MMO

[1057] Lei Chen and Di Yuan. Mathematical modeling for optimal design of in-building distributed antenna systems. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3428–3445, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002533.

Coluccia:2013:DBA

[1058] A. Coluccia, A. D’Alconzo, and F. Ricciato. Distribution-based anomaly detection via generalized likelihood ratio test: a general Maximum Entropy approach. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3446–3462, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002545.

Avrachenkov:2013:CCT

[1059] K. Avrachenkov, U. Ayesta, J. Doncel, and P. Jacko. Congestion control of TCP flows in Internet routers by means of index policy. Computer Networks (Amsterdam, Netherlands: 1999), 57 (17):3463–3478, December 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002557.

Cordeschi:2013:ESS

[1060] Nicola Cordeschi, Mohammad Shojafar, and Enzo Baccarelli. Energysaving self-configuring networked data centers. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3479–3491, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002569.

Lai:2013:SAS

[1061] Chengzhe Lai, Hui Li, Rongxing Lu, and Xuemin (Sherman) Shen. SE-AKA: a secure and efficient group authentication and key agreement protocol for LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3492–3510, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002570.

Shieh:2013:EDS

[1062] Ce-Kuen Shieh, Chia-Yu Yu, Cheng-Han Lin, and J. Morris Chang. Efficiency-driven selection of bandwidth request mechanism in broadband wireless access networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17):3511–3521, December 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128613002582.

Zhou:2013:STG

[1063] Peng Zhou, Xiapu Luo, Ang Chen, and Rocky K. C. Chang. SGor: Trust graph based onion routing. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3522–3544, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002594.

Carvalho:2013:SMD

[1064] Glaucio H. S. Carvalho, Isaac Woungang, Alagan Anpalagan, Rodolfo W. L. Coutinho, and João C. W. A. Costa. A semi-Markov decision process-based joint call admission control for inter-RAT cell reselection in next generation wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3545–3562, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002600.

Ha:2013:UCV

[1065] Jihun Ha, Jiung Yu, Byungjo Kim, and Hyogon Kim. On UDP continuity over vertical handovers. Computer Networks (Amsterdam, Netherlands: 1999), 57 (17):3563–3580, December 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002612.

Wang:2013:EEM

[1066] Le Wang and Jukka Manner. Energyefficient mobile web in a bundle. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3581–3600, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002624.

Varis:2013:DFF

[1067] Nuutti Varis, Jukka Manner, Mikko Särelä, and Timo Kiravuo. DBridges: Flexible floodless frame forwarding. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3601–3616, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002636.

Kim:2013:SAO

[1068] Byung-Gook Kim, Jeong-Ahn Kwon, and Jang-Won Lee. Subchannel allocation for the OFDMA-based femtocell system. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3617–3629, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002648.

Meng:2013:TAC

[1069] Yuxin Meng, Wenjuan Li, and Lam-For Kwok. Towards adaptive character frequency-based exclusive signature matching scheme and its applications in distributed intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3630–3640, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300265X.

Manzano:2013:ENR

[1070] M. Manzano, E. Calle, V. Torres-Padrosa, J. Segovia, and D. Harle. Endurance: a new robustness measure for complex networks under multiple failure scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 57 (17):3641–3653, December 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002740.

Atsan:2013:SSD

[1071] Emre Atsan and Öznur Özkasap.

SCALAR: Scalable data lookup and replication protocol for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3654–3672, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002752.

Zubeldia:2013:OPP

[1072] Martín Zubeldía, Andrés Ferragut, and Fernando Paganini. Overcoming performance pitfalls in rate-diverse high speed WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 57 (17):3673–3685, December 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002764.

Fras:2013:LMA

[1073] Matjaz Fras, Joze Mohorko, and Zarko Cucej. Limitations of a Mapping Algorithm with Fragmentation Mimics (MAFM) when modeling statistical data sources based on measured packet network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 57 (17):3686–3700, December 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002776.

Hou:2013:FBB

[1074] Cuiqin Hou, Yibin Hou, and Zhangqin Huang. A framework based on barycentric coordinates for localization in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17):3701– 3712, December 9, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002788.

Bolea:2013:MAC

[1075] Liliana Bolea, Jordi Pérez-Romero, and Ramón Agustí. ML aided context feature extraction for cognitive radio. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3713–3727, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300282X.

Iacovazzi:2013:ESM

[1076] A. Iacovazzi, A. D’Alconzo, F. Ricciato, and M. Burkhart. Elementary secure-multiparty computation for massive-scale collaborative network monitoring: a quantitative assessment. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3728–3742, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002831.

Carofiglio:2013:PBS

[1077] Giovanna Carofiglio, Massimo Gallo, and Luca Muscariello. On the performance of bandwidth and storage sharing in information-centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(17): 3743–3758, December 9, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002843.

Anadiotis:2013:PIG

[1078] Angelos-Christos G. Anadiotis, Charalampos Z. Patrikakis, and Iakovos S. Venieris. On the performance improvement of gossip protocols for content-based publish-subscribe through caching. Computer Networks (Amsterdam, Netherlands: 1999), 57 (18):3759–3772, December 24, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003071.

Lin:2013:MPA

[1079] Peng Lin, Alan Kai-Hau Yeung, and Angus Kin-Yeung Wong. Multi-portal association based dispatching and a virtual-queue method in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3773–3789, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003113.

Akkari:2013:IBI

[1080] Nadine Akkari. An IMS-based integration architecture for WiMax/LTE handover. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3790–3798, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003101.

Cai:2013:JRJ

[1081] Yifeng Cai, Konstantinos Pelechrinis, Xin Wang, Prashant Krishnamurthy, and Yijun Mo. Joint reactive jammer detection and localization in an enterprise WiFi network. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3799–3811, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003095.

Ghods:2013:MMM

[1082] Fatemeh Ghods, Hamed Yousefi, Ali Mohammad Afshin Hemmatyar, and Ali Movaghar. MC-MLAS: Multichannel Minimum Latency Aggregation Scheduling in Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3812–3825, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003125.

Bou-Harb:2013:SAD

[1083] Elias Bou-Harb, Mourad Debbabi, and Chadi Assi. A systematic approach for detecting and clustering distributed cyber scanning. Computer Networks (Amsterdam, Netherlands: 1999), 57 (18):3826–3839, December 24, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003137.

Chen:2013:MPU

[1084] Baozhi Chen and Dario Pompili.

Minimizing position uncertainty for under-ice autonomous underwater vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3840–3854, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003149.

Naor:2013:SFM

[1085] Zohar Naor and Sajal K. Das. A scalable framework for mobile real-time group communication services. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3855–3865, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003150.

Nam:2013:CAM

[1086] Seung Yeob Nam, Sirojiddin Djuraev, and Minho Park. Collaborative approach to mitigating ARP poisoning-based Man-in-the-Middle attacks. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3866–3884, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003162.

Sun:2013:LDS

[1087] Wen Sun, Yu Ge, and Wai-Choong Wong. A lightweight distributed scheme for mitigating inter-user interference in body sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(18):3885– 3896, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003393.

Paredes-Oliva:2013:FFR

[1088] Ignasi Paredes-Oliva, Pere Barlet-Ros, and Xenofontas Dimitropoulos. FaRNet: Fast recognition of highdimensional patterns from big network traffic data. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3897–3913, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003435.

Al-Arnaout:2013:DCA

[1089] Zakwan Al-Arnaout, Qiang Fu, and Marcus Frean. A divide-and-conquer approach for content replication in WMNs. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3914–3928, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003423.

Wu:2013:RDD

[1090] Weigang Wu, Jiannong Cao, Hejun Wu, and Jingjing Li. Robust and dynamic data aggregation in wireless sensor networks: a cross-layer approach. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3929–3940, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003411.

Lee:2013:OPC

[1091] Seung-Ho Lee, Han-You Jeong, and Seung-Woo Seo. Optimal pricing and capacity partitioning for tiered access service in virtual networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3941–3956, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300340X.

Qiu:2013:DRS

[1092] Wanzhi Qiu and Efstratios Skafidas. Distributed routing for signal detection in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3957–3966, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003459.

Rajiullah:2013:PAI

[1093] Mohammad Rajiullah, Reine Lundin, Anna Brunstrom, and Stefan Lindskog. Performance analysis and improvement of PR-SCTP for small messages. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3967–3986, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003447.

Wu:2013:HMW

[1094] Zhenhua Wu and Yu Hen Hu. How many wireless resources are needed to resolve the hidden terminal problem? Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 3987–3996, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003472.

Accongiagioco:2013:TOR

[1095] Giovanni Accongiagioco, Matteo Andreozzi, Daniele Migliorini, and Giovanni Stea. Throughput-optimal resource allocation in LTE-Advanced with distributed antennas. Computer Networks (Amsterdam, Netherlands: 1999), 57(18):3997–4009, December 24, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003496.

Gokturk:2013:CLM

[1096] M. Sarper Gokturk, Ozgur Gurbuz, and Elza Erkip. A cross-layer multi-hop cooperative network architecture for wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 4010–4029, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003587.

Mehmood:2013:OOR

[1097] Tahir Mehmood, Lavy Libman, Hooman Reisi Dehkordi, and Sanjay K. Jha. Optimal opportunistic routing and network coding for bidirectional wireless flows. Computer Networks (Amsterdam, Netherlands: 1999), 57 (18):4030–4046, December 24, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003605.

Geravand:2013:BFA

[1098] Shahabeddin Geravand and Mahmood Ahmadi. Bloom filter applications in network security: a state-of-the-art survey. Computer Networks (Amsterdam, Netherlands: 1999), 57(18): 4047–4064, December 24, 2013. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003083.

Akyildiz:2014:E

[1099] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 58 (??):1, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004258.

Cho:2014:LLF

[1100] Sangman Cho and Srinivasan Ramasubramanian. Localizing link failures in all-optical networks using monitoring tours. Computer Networks (Amsterdam, Netherlands: 1999), 58 (??):2–12, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002442.

Block:2014:BPA

[1101] R. Block, G. T. Peeters, and B. Van Houdt. A branching process approach to compute the delay and energy efficiency of tree algorithms with free access. Computer Networks (Amsterdam, Netherlands: 1999), 58 (??):13–28, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002880.

Yang:2014:PST

[1102] Haomin Yang, Yaoxue Zhang, Yuezhi Zhou, Xiaoming Fu, Hao Liu, and Athanasios V. Vasilakos. Provably secure three-party authenticated key agreement protocol using smart cards. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 29–38, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002867.

Jung:2014:ACR

[1103] Kyoung-Hak Jung, Hyo-Ryun Lee, Wan-Seon Lim, and Young-Joo Suh. An adaptive collision resolution scheme for energy efficient communication in IEEE 802.15.4 networks. Computer Networks (Amsterdam, Netherlands: 1999), 58(??):39–57, January 15, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002855.

Uddin:2014:JOA

[1104] Mohammad Faisal Uddin, Chadi Assi, and Ali Ghrayeb. Joint optimal AF relay assignment and power allocation in wireless cooperative networks. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 58–69, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002971.

Perez:2014:TPI

[1105] Manuel Gil Pérez, Juan E. Tapiador, John A. Clark, Gregorio Martínez Pérez, and Antonio F. Skarmeta Gómez. Trustworthy placements: Improving quality and resilience in collaborative attack detection. Computer Networks (Amsterdam, Netherlands: 1999), 58(??):70–86, January 15, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003009.

Zhang:2014:SSB

[1106] Wenjie Zhang and Chai Kiat Yeo. Sequential sensing based spectrum handoff in cognitive radio networks with multiple users. Computer Networks (Amsterdam, Netherlands: 1999), 58 (??):87–98, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002995.

Katkalov:2014:MTC

[1107] Kuzman Katkalov, Nina Moebius, Kurt Stenzel, Marian Borek, and Wolfgang Reif. Modeling test cases for security protocols with SecureMDD. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 99–111, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613002983.

Dou:2014:MUW

[1108] Lihua Dou, Yunchuan Wei, and Jun Ni. Multi-user wireless channel probing for shared key generation with a fuzzy controller. Computer Networks (Amsterdam, Netherlands: 1999), 58 (??):112–126, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003010.

Wu:2014:MBS

[1109] Xiaofei Wu, Xin Wang, Ke Yu, and Frank Y. Li. A measurementbased study on the correlations of inter-domain Internet application flows. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 127–140, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003034.

Yao:2014:FBP

[1110] Wei-Min Yao and Sonia Fahmy. Flowbased partitioning of network testbed experiments. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 141–157, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300306X.

Kim:2014:EEM

[1111] Jin Cheol Kim and Younghee Lee.

An end-to-end measurement and monitoring technique for the bottleneck link capacity and its available bandwidth. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 158–179, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003058.

Carrea:2014:OHN

[1112] Laura Carrea, Alexei Vernitski, and Martin Reed. Optimized hash for network path encoding with minimized false positives. Computer Networks (Amsterdam, Netherlands: 1999), 58 (??):180–191, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003046.

Vardhan:2014:GWL

[1113] Hars Vardhan, Seong-Ryong Ryu, Bhaskar Banerjee, and Ravi Prakash. 60 GHz wireless links in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 58 (??):192–205, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003460.

Sherr:2014:DIA

[1114] Micah Sherr, Harjot Gill, Taher Aquil Saeed, Andrew Mao, William R. Marczak, Saravana Soundararajan, Wenchao Zhou, Boon Thau Loo, and Matt Blaze. The design and implementation of the A^{3} application-aware anonymity platform. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 206–227, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003599.

Lai:2014:EEC

[1115] Hung-Quoc Lai, Yan Chen, and K. J. Ray Liu. Energy efficient cooperative communications using coalition formation games. Computer Networks (Amsterdam, Netherlands: 1999), 58 (??):228–238, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003630.

Shawky:2014:NAD

[1116] A. Shawky, R. Olsen, J. Pedersen, and H. Schwefel. Network aware dynamic context subscription management. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 239–253, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003629.

Younis:2014:TMT

[1117] Mohamed Younis, Izzet F. Senturk, Kemal Akkaya, Sookyoung Lee, and Fatih Senel. Topology management techniques for tolerating node failures in wireless sensor networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 58(??): 254–283, January 15, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613002879.

Pamies-Juarez:2014:IBD

[1118] Lluis Pamies-Juarez, Marc Sanchez-Artigas, Pedro García-López, Rubén Mondéjar, and Rahma Chaabouni.

On the interplay between data redundancy and retrieval times in P2P storage systems. Computer Networks (Amsterdam, Netherlands: 1999), 59 (??):1–16, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004088.

Dabirmoghaddam:2014:ORC

[1119] Ali Dabirmoghaddam, Majid Ghaderi, and Carey Williamson. On the optimal randomized clustering in distributed sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 59 (??):17–32, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004118.

Asghari:2014:EES

[1120] Naser M. Asghari, M. Mandjes, and Anwar Walid. Energy-efficient scheduling in multi-core servers. Computer Networks (Amsterdam, Netherlands: 1999), 59(??):33–43, February 11, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861300412X.

Sourlas:2014:NCA

[1121] Vasilis Sourlas, Paris Flegkas, and Leandros Tassiulas. A novel cache aware routing scheme for Information-Centric Networks. Computer Networks (Amsterdam, Netherlands: 1999), 59 (??):44–61, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004039.

Zhang:2014:SGI

[1122] Xiaoning Zhang, Haoran Wang, and Zian Zhang. Survivable green IP over WDM networks against doublelink failures. Computer Networks (Amsterdam, Netherlands: 1999), 59 (??):62–76, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004143.

Cuevas:2014:TSS

[1123] Rubén Cuevas, Michal Kryczka, Roberto González, Angel Cuevas, and Arturo Azcorra. TorrentGuard: Stopping scam and malware distribution in the BitTorrent ecosystem. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 77–90, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004106.

Toklu:2014:BME

[1124] Sinan Toklu and O. Ayhan Erdem.

BSC-MAC: Energy efficiency in wireless sensor networks with base station control. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 91–100, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004155.

Ouyang:2014:LSE

[1125] Tu Ouyang, Soumya Ray, Mark Allman, and Michael Rabinovich. A large-scale empirical analysis of email spam detection through network characteristics in a stand-alone enterprise. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 101–121, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003484.

Fischer:2014:RUA

[1126] Mathias Fischer, Sascha Grau, Giang Nguyen, and Guenter Schaefer. Resilient and underlay-aware P2P livestreaming. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 122–136, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003617.

Asheralieva:2014:TSR

[1127] Alia Asheralieva and Kaushik Mahata. A two-step resource allocation procedure for LTE-based cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 137–152, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861300368X.

Coras:2014:LSD

[1128] Florin Coras, Jordi Domingo-Pascual, Fabio Maino, Dino Farinacci, and Albert Cabellos-Aparicio. Lcast:

Software-defined inter-domain multicast. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 153–170, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003708.

Yaacoub:2014:AMR

[1129] Elias Yaacoub and Adnan Abu-Dayya.

Automatic meter reading in the smart grid using contention based random access over the free cellular spectrum. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 171–183, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003691.

Jeong:2014:ERM

[1130] Jaeseong Jeong, Kyunghan Lee, Yung Yi, Injong Rhee, and Song Chong. ExMin: a routing metric for novel opportunity gain in Delay Tolerant Networks. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 184–196, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003782.

Liu:2014:PFB

[1131] Yangyang Liu, Le Chang, and Jianping Pan. On the performance and fairness of BitTorrent-like data swarming systems with NAT devices. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 197–212, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003770.

Mirtaheri:2014:ERD

[1132] Seyedeh Leili Mirtaheri and Mohsen Sharifi. An efficient resource discovery framework for pure unstructured peerto-peer systems. Computer Networks (Amsterdam, Netherlands: 1999), 59 (??):213–226, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003769.

Xu:2014:ETD

[1133] Mingwei Xu, Shu Yang, Dan Wang, and Jianping Wu. Efficient Two Dimensional-IP routing: an incremental deployment design. Computer Networks (Amsterdam, Netherlands: 1999), 59(??):227–243, February 11, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128613003757.

Papastergiou:2014:DTP

[1134] Giorgos Papastergiou, Ioannis Alexiadis, Scott Burleigh, and Vassilis Tsaoussidis. Delay Tolerant Payload Conditioning protocol. Computer Networks (Amsterdam, Netherlands: 1999), 59(??): 244–263, February 11, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003745.

Vallet:2014:OOW

[1135] Josselin Vallet and Olivier Brun. Online OSPF weights optimization in IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 60 (??):1–12, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004179.

Ramirez:2014:STI

[1136] W. Ramirez, X. Masip-Bruin, M. Yannuzzi, R. Serral-Gracia, A. Martinez, and M. S. Siddiqui. A survey and taxonomy of ID/Locator Split Architectures. Computer Networks (Amsterdam, Netherlands: 1999), 60(??): 13–33, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300409X.

Jia:2014:UBP

[1137] Adele L. Jia, Xiaowei Chen, Xiaowen Chu, Johan A. Pouwelse, and Dick H. J. Epema. User behaviors in private BitTorrent communities. Computer Networks (Amsterdam, Netherlands: 1999), 60(??): 34–45, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004131.

Condeixa:2014:SID

[1138] Tiago Condeixa and Susana Sargento.

Studying the integration of distributed and dynamic schemes in the mobility management. Computer Networks (Amsterdam, Netherlands: 1999), 60 (??):46–59, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004167.

Faigl:2014:PEC

[1139] Zoltán Faigl, Jani Pellikka, László Bokor, and Andrei Gurtov. Performance evaluation of current and emerging authentication schemes for future 3GPP network architectures. Computer Networks (Amsterdam, Netherlands: 1999), 60(??):60–74, February 26, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128613004180.

Chen:2014:PDP

[1140] Fei Chen, Haitao Li, and Jiangchuan Liu. Popularity decays in peer-to-peer VoD systems: Impact, model, and design implications. Computer Networks (Amsterdam, Netherlands: 1999), 60 (??):75–87, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004313.

Bellalta:2014:PAM

[1141] Boris Bellalta, Azadeh Faridi, Jaume Barcelo, Vanesa Daza, and Miquel Oliver. Performance analysis of a Multiuser Multi-Packet Transmission system for WLANs in non-saturation conditions. Computer Networks (Amsterdam, Netherlands: 1999), 60(??): 88–100, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004350.

Chi:2014:PTA

[1142] Kaikai Chi, Yi hua Zhu, Xiaohong Jiang, and Xianzhong Tian. Practical throughput analysis for two-hop wireless network coding. Computer Networks (Amsterdam, Netherlands: 1999), 60(??):101– 114, February 26, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004283.

Testa:2014:DBC

[1143] Claudio Testa and Dario Rossi.

Delay-based congestion control: Flow vs. BitTorrent swarm perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 60(??):115– 128, February 26, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004295.

Razo-Zapata:2014:SHT

[1144] Iván S. Razo-Zapata, Gerardo Castañón, and Carlos Mex-Perera. Self-healing in transparent optical packet switching mesh networks: a reinforcement learning perspective. Computer Networks (Amsterdam, Netherlands: 1999), 60 (??):129–146, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003733.

Gotta:2014:TIS

[1145] A. Gotta, M. Luglio, and C. Roseti. A TCP/IP satellite infrastructure for sensing operations in emergency contexts. Computer Networks (Amsterdam, Netherlands: 1999), 60(??): 147–159, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003721.

Temel:2014:RPD

[1146] Samil Temel, Vehbi Çagri Gungor, and Taskin Koçak. Routing protocol design guidelines for smart grid environments. Computer Networks (Amsterdam, Netherlands: 1999), 60(??): 160–170, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003927.

Anelli:2014:FPA

[1147] Pascal Anelli, Rémi Diana, and Emmanuel Lochin. FavorQueue: a parameterless active queue management to improve TCP traffic performance. Computer Networks (Amsterdam, Netherlands: 1999), 60(??): 171–186, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003915.

Mata:2014:ADD

[1148] Felipe Mata, Piotr Zuraniewski, Michel Mandjes, and Marco Mellia. Anomaly detection in diurnal data. Computer Networks (Amsterdam, Netherlands: 1999), 60(??):

187–200, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003940.

Coimbra:2014:EER

[1149] J. Coimbra, G. Schütz, and N. Correia. Energy efficient routing algorithm for fiber-wireless access networks: a network formation game approach. Computer Networks (Amsterdam, Netherlands: 1999), 60(??): 201–216, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003976.

Cui:2014:OMV

[1150] Hao Cui, Xiao Su, and Weijia Shang.

On optimal media/video distribution in closed P2P-based IPTV networks. Computer Networks (Amsterdam, Netherlands: 1999), 60(??): 217–232, February 26, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003861.

Sterbenz:2014:SIFa

[1151] James P. G. Sterbenz, David Hutchison, Paul Müller, and Chip Elliott. Special issue on Future Internet Testbeds — Part I: Guest Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 61(??): 1–4, March 14, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000504.

Berman:2014:GFT

[1152] Mark Berman, Jeffrey S. Chase, Lawrence Landweber, Akihiro Nakao, Max Ott, Dipankar Raychaudhuri, Robert Ricci, and Ivan Seskar. GENI: a federated testbed for innovative network experiments. Computer Networks (Amsterdam, Netherlands: 1999), 61 (??):5–23, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004507.

Bastin:2014:IIA

[1153] Nicholas Bastin, Andy Bavier, Jessica Blaine, Jim Chen, Narayan Krishnan, Joe Mambretti, Rick McGeer, Rob Ricci, and Nicki Watts. The InstaGENI initiative: an architecture for distributed systems and advanced programmable networks. Computer Networks (Amsterdam, Netherlands: 1999), 61(??):24–38, March 14, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004477.

Kim:2014:KGT

[1154] Dongkyun Kim, Joobum Kim, Gicheol Wang, Jin-Hyung Park, and Seung-Hae Kim. K-GENI testbed deployment and federated meta operations experiment over GENI and KREONET. Computer Networks (Amsterdam, Netherlands: 1999), 61(??):39–50, March 14, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004052.

Medhi:2014:GTN

[1155] Deep Medhi, Byrav Ramamurthy, Caterina Scoglio, Justin P. Rohrer, Egemen K. Çetinkaya, Ramkumar Cherukuri, Xuan Liu, Pragatheeswaran Angu, Andy Bavier, Cort Buffington, and James P. G. Sterbenz. The GpENI testbed: Network infrastructure, implementation experience, and experimentation. Computer Networks (Amsterdam, Netherlands: 1999), 61(??):51–74, March 14, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004404.

Gran:2014:COM

[1156] Ernst Gunnar Gran, Thomas Dreibholz, and Amund Kvalbein. N or NetCore — a multi-homed research testbed. Computer Networks (Amsterdam, Netherlands: 1999), 61 (??):75–87, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004489.

Kvalbein:2014:NEP

[1157] Amund Kvalbein, Dziugas Baltrunas, Kristian Evensen, Jie Xiang, Ahmed Elmokashfi, and Simone Ferlin-Oliveira. The Nornet Edge platform for mobile broadband measurements. Computer Networks (Amsterdam, Netherlands: 1999), 61(??): 88–101, March 14, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004490.

Schwerdel:2014:FIR

[1158] Dennis Schwerdel, Bernd Reuther, Thomas Zinner, Paul Müller, and Phouc Tran-Gia. Future Internet research and experimentation: the G-Lab approach. Computer Networks (Amsterdam, Netherlands: 1999), 61 (??):102–117, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004362.

Mambretti:2014:CEI

[1159] Joe Mambretti, Jim Chen, and Fei Yeh. Creating environments for innovation: Designing and implementing advanced experimental network research testbeds based on the Global Lambda Integrated Facility and the StarLight Exchange. Computer Networks (Amsterdam, Netherlands: 1999), 61(??): 118–131, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004374.

Sune:2014:DIO

[1160] M. Suñé, L. Bergesio, H. Woesner, T. Rothe, A. Köpsel, D. Colle, B. Puype, D. Simeonidou, R. Nejabati, M. Channegowda, M. Kind, T. Dietz, A. Autenrieth, V. Kotronis, E. Salvadori, S. Salsano, M. Körner, and S. Sharma. Design and implementation of the OFELIA FP7 facility: the European OpenFlow testbed. Computer Networks (Amsterdam, Netherlands: 1999), 61(??):132–150, March 14, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004301.

Kobayashi:2014:MOS

[1161] Masayoshi Kobayashi, Srini Seetharaman, Guru Parulkar, Guido Appenzeller, Joseph Little, Johan van Reijendam, Paul Weissmann, and Nick McKeown. Maturing of OpenFlow and Software-defined Networking through deployments. Computer Networks (Amsterdam, Netherlands: 1999), 61 (??):151–175, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300371X.

Campanella:2014:FIE

[1162] M. Campanella and F. Farina. The FEDERICA infrastructure and experience. Computer Networks (Amsterdam, Netherlands: 1999), 61(??): 176–183, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004428.

Jofre:2014:FBM

[1163] Jordi Jofre, Celia Velayos, Giada Landi, Michal Giertych, Alastair C. Hume, Gareth Francis, and Albert Vico Oton. Federation of the BonFIRE multi-cloud infrastructure with networking facilities. Computer Networks (Amsterdam, Netherlands: 1999), 61 (??):184–196, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003952.

Belter:2014:GOT

[1164] Bartosz Belter, Juan Rodriguez Martinez, José Ignacio Aznar, Jordi Ferrer Riera, Luis M. Contreras, Monika Antoniak-Lewandowska, Matteo Biancani, Jens Buysse, Chris Develder, Yuri Demchenko, Pasquale Donadio, Dimitra Simeonidou, Reza Nejabati, Shuping Peng, Lukasz Drzewiecki, Eduard Escalona, Joan Antoni Garcia Espin, Steluta Gheorghiu, Mattijs Ghijsen, Jakub Gutkowski, Giada Landi, et al. The GEYSERS optical testbed: a platform for the integration, validation and demonstration of cloud-based infrastructure services. Computer Networks (Amsterdam, Netherlands: 1999), 61 (??):197–216, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004441.

Sanchez:2014:SIE

[1165] Luis Sanchez, Luis Muñoz, Jose Antonio Galache, Pablo Sotres, Juan R. Santana, Veronica Gutierrez, Rajiv Ramdhany, Alex Gluhak, Srdjan Krco, Evangelos Theodoridis, and Dennis Pfisterer. SmartSantander: IoT experimentation over a smart city testbed. Computer Networks (Amsterdam, Netherlands: 1999), 61 (??):217–238, March 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004337.

Meng:2014:ACM

[1166] Xianfu Meng and Changyuan Zhang. An ant colony model based replica consistency maintenance strategy in unstructured P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 62(??):1–11, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000024.

Lee:2014:SHB

[1167] Yoh han Lee and Daeyoung Kim.

Slow hopping based cooperative sensing MAC protocol for cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):12–28, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000061.

Amokrane:2014:EEM

[1168] Ahmed Amokrane, Rami Langar, Raouf Boutaba, and Guy Pujolle. Energy efficient management framework for multihop TDMA-based wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):29–42, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400005X.

Yilmaz:2014:ADD

[1169] Mete Yilmaz and Nirwan Ansari.

Achieving destination differentiation in ingress aggregated fairness for resilient packet rings by weighted destination based fair dropping. Computer Networks (Amsterdam, Netherlands: 1999), 62(??):43–54, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000048.

Guimaraes:2014:MGM

[1170] Carlos Guimarães, Daniel Corujo, Antonio de la Oliva, Yoshihiro Ohba, and Rui L. Aguiar. Multicast group membership management in media independent handover services. Computer Networks (Amsterdam, Netherlands: 1999), 62(??):55–68, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000176.

Fajjari:2014:NVN

[1171] Ilhem Fajjari, Nadjib Aitsaadi, Michal Pióro, and Guy Pujolle. A new virtual network static embedding strategy within the Cloud’s private backbone network. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):69–88, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861400022X.

Pinero:2014:AIM

[1172] P. J. Piñero, J. A. Cortés, J. Malgosa, F. J. Cañete, P. Manzanares, and L. Díez. Analysis and improvement of multicast communications in HomePlug AV-based in-home networks. Computer Networks (Amsterdam, Netherlands: 1999), 62(??): 89–100, April 7, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000218.

Dezfouli:2014:IBR

[1173] Behnam Dezfouli, Marjan Radi, Shukor Abd Razak, Kamin Whitehouse, Kamalrulnizam Abu Bakar, and Tan Hwee-Pink. Improving broadcast reliability for neighbor discovery, link estimation and collection tree construction in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):101–121, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000206.

Giotis:2014:COS

[1174] K. Giotis, C. Argyropoulos, G. Androulidakis, D. Kalogeras, and V. Maglaris. Combining OpenFlow and sFlow for an effective and scalable anomaly detection and mitigation mechanism on SDN environments. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):122–136, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004003.

Di:2014:DRV

[1175] Hao Di, Vishal Anand, Hongfang Yu, Lemin Li, Gang Sun, and Dan Liao. Design of reliable virtual infrastructure with resource sharing. Computer Networks (Amsterdam, Netherlands: 1999), 62(??):137–151, April 7, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300399X.

Bianco:2014:MMF

[1176] Andrea Bianco, Davide Cuda, and Jorge M. Finochietto. Multi-MetaRing fairness control in a WDM foldedbus architecture. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):152–161, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003988.

Amin:2014:SDP

[1177] Saman Hameed Amin, H. S. Al-Raweshidy, and Rafed Sabbar Abbas. Smart data packet ad hoc routing protocol. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):162–181, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004015.

Tian:2014:DSI

[1178] Jie Tian, Guiling Wang, Tan Yan, and Wensheng Zhang. Detect smart intruders in sensor networks by creating network dynamics. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):182–196, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004064.

Gregori:2014:IRI

[1179] Enrico Gregori, Alessandro Improta, Luciano Lenzini, Lorenzo Rossi, and Luca Sani. Improving the reliability of inter-AS economic inferences through a hygiene phase on BGP data. Computer Networks (Amsterdam, Netherlands: 1999), 62(??):197–207, April 7, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004040.

Cheng:2014:DBC

[1180] Chien-Fu Cheng, Ting-Ying Wu, and Hsien-Chun Liao. A densitybarrier construction algorithm with minimum total movement in mobile WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 62(??): 208–220, April 7, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004027.

Meier:2014:WPF

[1181] Dominic Meier, Yvonne Anne Pignolet, Stefan Schmid, and Roger Wattenhofer. On the Windfall and price of friendship: Inoculation strategies on social networks. Computer Networks (Amsterdam, Netherlands: 1999), 62 (??):221–236, April 7, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004076.

Sterbenz:2014:SIFb

[1182] James P. G. Sterbenz, David Hutchison, Paul Müller, and Chip Elliott. Special issue on Future Internet Testbeds — Part II: Guest Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 63 (??):1–4, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861400098X.

Sydney:2014:UGE

[1183] Ali Sydney, David S. Ochs, Caterina Scoglio, Don Gruenbacher, and Ruth Miller. Using GENI for experimental evaluation of Software Defined Networking in smart grids. Computer Networks (Amsterdam, Netherlands: 1999), 63(??):5–16, April 22, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004349.

Griffioen:2014:MEG

[1184] James Griffioen, Zongming Fei, Hussamuddin Nasir, Xiongqi Wu, Jeremy Reed, and Charles Carpenter. Measuring experiments in GENI. Computer Networks (Amsterdam, Netherlands: 1999), 63(??):17–32, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004325.

Nozaki:2014:ETR

[1185] Yoshihiro Nozaki, Parth Bakshi, Hasan Tuncer, and Nirmala Shenoy. Evaluation of tiered routing protocol in floating cloud tiered Internet architecture. Computer Networks (Amsterdam, Netherlands: 1999), 63(??): 33–47, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613003939.

Keranidis:2014:EEE

[1186] Stratos Keranidis, Dimitris Giatsios, Thanasis Korakis, Iordanis Koutsopoulos, Leandros Tassiulas, Thierry Rakotoarivelo, Max Ott, and Thierry Parmentelat. Experimentation on endto-end performance aware algorithms in the federated environment of the heterogeneous PlanetLab and NI-TOS testbeds. Computer Networks (Amsterdam, Netherlands: 1999), 63 (??):48–67, April 22, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004398.

Mehani:2014:IFC

[1187] Olivier Mehani, Guillaume Jourjon, Thierry Rakotoarivelo, and Max Ott. An instrumentation framework for the critical task of measurement collection in the future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 63(??):68–83, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000085.

Tosic:2014:RSI

[1188] Milorad Tosic and Ivan Seskar. Resource specification and intelligent user interaction for federated testbeds using Semantic Web technologies. Computer Networks (Amsterdam, Netherlands: 1999), 63(??):84–100, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000231.

Matias:2014:EOO

[1189] Jon Matias, Alaitz Mendiola, Nerea Toledo, Borja Tornero, and Eduardo Jacob. The EHU–OEF: an OpenFlow-based Layer-2 experimental facility. Computer Networks (Amsterdam, Netherlands: 1999), 63(??): 101–127, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613003964.

Ata:2014:SFI

[1190] Shingo Ata, Dijiang Huang, Xuan Liu, Akira Wada, Tianyi Xing, Parikshit Juluri, Chun-Jen Chung, Yasuhiro Sato, and Deep Medhi. SeRViTR: a framework, implementation, and a testbed for a trustworthy future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 63(??): 128–146, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004416.

Yannuzzi:2014:TSA

[1191] M. Yannuzzi, M. S. Siddiqui, A. Sällström, B. Pickering, R. Serral-Gracià, A. Martínez, W. Chen, S. Taylor, F. Benbadis, J. Leguay, E. Borrelli, I. Ormaetxea, K. Campowsky, G. Giammatteo, G. Aristomenopoulos, S. Papavassiliou, T. Kuczynski, S. Zielinski, J. M. Seigneur, C. Ballester Lafuente, J. Johansson, X. Masip-Bruin, M. Caria, J. R. Ribeiro Junior, E. Salageanu, and J. Latanicki. TEFIS: a single access point for conducting multifaceted experiments on heterogeneous test facilities. Computer Networks (Amsterdam, Netherlands: 1999), 63 (??):147–172, April 22, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861300443X.

Rakotoarivelo:2014:DOR

[1192] Thierry Rakotoarivelo, Guillaume Jourjon, and Max Ott. Designing and orchestrating reproducible experiments on federated networking testbeds. Computer Networks (Amsterdam, Netherlands: 1999), 63 (??):173–187, April 22, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004465.

Han:2014:ESC

[1193] Sang Woo Han, Namgon Kim, and JongWon Kim. An experimental service composition tool for media-centric networked applications. Computer Networks (Amsterdam, Netherlands: 1999), 63(??):188–204, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004386.

Auge:2014:TFG

[1194] Jordan Augé, Thierry Parmentelat, Nicolas Turro, Sandrine Avakian, Loïc Baron, Mohamed Amine Larabi, Mohammed Yasin Rahman, Timur Friedman, and Serge Fdida. Tools to foster a global federation of testbeds. Computer Networks (Amsterdam, Netherlands: 1999), 63(??):205–220, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128613004519.

Moraes:2014:FFV

[1195] Igor M. Moraes, Diogo M. F. Mattos, Lyno Henrique G. Ferraz, Miguel Elias M. Campista, Marcelo G. Rubinstein, Luís Henrique M. K. Costa, Marcelo D. de Amorim, Pedro B. Velloso, Otto Carlos M. B. Duarte, and Guy Pujolle. FITS: a flexible virtual network testbed architecture. Computer Networks (Amsterdam, Netherlands: 1999), 63(??): 221–237, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000036.

Kim:2014:EID

[1196] Young-Hwan Kim, Alina Quereilhac, Mohamed Amine Larabi, Julien Tribino, Thierry Parmentelat, Thierry Turletti, and Walid Dabbous. Enabling iterative development and reproducible evaluation of network protocols. Computer Networks (Amsterdam, Netherlands: 1999), 63(??): 238–250, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000073.

Dulinski:2014:UTL

[1197] Zbigniew Duliński, Kamil Palkowski, and Piotr Cholda. A university testbed for large-scale interconnection experiments on distributed applications. Computer Networks (Amsterdam, Netherlands: 1999), 63(??): 251–264, April 22, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128613004453.

Li:2014:TPH

[1198] Ming Li, Andrey Lukyanenko, Sasu Tarkoma, Yong Cui, and Antti Ylä-Jääski. Tolerating path heterogeneity in multipath TCP with bounded receive buffers. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):1–14, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000425.

Cong:2014:LEP

[1199] Xin Cong, Kai Shuang, Sen Su, FangChun Yang, and LingLing Zi. LBAS: an effective pricing mechanism towards video migration in cloudassisted VoD system. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):15–25, May 8, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000413.

DiStasi:2014:CMP

[1200] Giovanni Di Stasi, Jonas Karlsson, Stefano Avallone, Roberto Canonico, Andreas Kassler, and Anna Brunstrom. Combining multi-path forwarding and packet aggregation for improved network performance in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):26–37, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000346.

Kim:2014:UCD

[1201] Sun-Hyun Kim, Yeonsik Jeong, and Seung-Jae Han. Use of contact duration for message forwarding in intermittently connected mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):38–54, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000358.

Rendon:2014:MVN

[1202] Oscar Mauricio Caicedo Rendon, Carlos Raniery Paula dos Santos, Arthur Selle Jacobs, and Lisandro Zambenedetti Granville. Monitoring Virtual Nodes using mashups. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):55–70, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128614000498.

Couto:2014:NRC

[1203] Rodrigo S. Couto, Miguel Elias M. Campista, and Luís Henrique M. K. Costa. Network resource control for Xen-based virtualized software routers. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):71–88, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000450.

Abid:2014:POM

[1204] S. A. Abid, Mazliza Othman, and Nadir Shah. 3D P2P overlay over MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):89–111, May 8, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000486.

Kang:2014:AFC

[1205] Daeho Kang, Sangkyu Park, Changhee Joo, and Saewoong Bahk. Addressfree contention in wireless access networks with common control channel for throughput improvement. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):112–124, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000516.

Coiro:2014:RPC

[1206] Angelo Coiro, Luca Chiaraviglio, Antonio Cianfrani, Marco Listanti, and Marco Polverini. Reducing power consumption in backbone IP networks through table lookup bypass. Computer Networks (Amsterdam, Netherlands: 1999), 64(??): 125–142, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000474.

Garcia:2014:MGH

[1207] Cristian González García, B. Cristina Pelayo G-Bustelo, Jordán Pascual Espada, and Guillermo Cueva-Fernandez. Midgar:

Generation of heterogeneous objects interconnecting applications. A Domain Specific Language proposal for Internet of Things scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):143–158, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000528.

Astorga:2014:ESA

[1208] Jasone Astorga, Eduardo Jacob, Nerea Toledo, and Juanjo Unzilla. Enhancing secure access to sensor data with user privacy support. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):159–179, May 8, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000449.

Tan:2014:SPA

[1209] C. K. Tan, T. C. Chuah, and S. W. Tan. Subcarrier and power allocation for OFDMA-based multicast cellular networks using a coalitional game. Computer Networks (Amsterdam, Netherlands: 1999), 64(??): 180–194, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000668.

Moshtaghi:2014:AEA

[1210] Masud Moshtaghi, Christopher Leckie, Shanika Karunasekera, and Sutharshan Rajasegarar. An adaptive elliptical anomaly detection model for wireless sensor networks. Computer Networks (Amsterdam, Netherlands:

1999), 64(??):195–207, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000462.

Neumann:2014:DSL

[1211] Christoph Neumann, Olivier Heen, and Stéphane Onno. DNStamp: Shortlived trusted timestamping. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):208–224, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000784.

Younes:2014:IPR

[1212] Maram Bani Younes, Azzedine Boukerche, and Graciela Rom’an-Alonso. An intelligent path recommendation protocol (ICOD) for VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):225– 242, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400067X.

Dandoush:2014:LAD

[1213] Abdulhalim Dandoush, Sara Alouf, and Philippe Nain. Lifetime and availability of data stored on a P2P system: Evaluation of redundancy and recovery schemes. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):243–260, May 8, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000772.

Wang:2014:STD

[1214] Yufei Wang, Rong Zheng, and Qixin Wang. Self-tuned distributed monitoring of multi-channel wireless networks using Gibbs sampler. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):261–272, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000681.

Saied:2014:LCK

[1215] Yosra Ben Saied, Alexis Olivereau, Djamal Zeghlache, and Maryline Laurent. Lightweight collaborative key establishment scheme for the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):273–295, May 8, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000437.

Hoteit:2014:EHT

[1216] Sahar Hoteit, Stefano Secci, Stanislav Sobolevsky, Carlo Ratti, and Guy Pujolle. Estimating human trajectories and hotspots through mobile phone data. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):296–307, May 8, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000656.

Wang:2014:ACC

[1217] Min Wang, Junfeng Wang, and Sunyoung Han. Adaptive congestion control framework and a simple implementation on high bandwidth-delay product networks. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):308–321, May 8, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000875.

Guerraoui:2014:TFG

[1218] Rachid Guerraoui, Kévin Huguenin, Anne-Marie Kermarrec, Maxime Monod, Swagatika Prusty, and Aline Roumy. Tracking freeriders in gossip-based content dissemination systems. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):322–338, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000899.

Lin:2014:DAF

[1219] Hui-Tang Lin, Chia-Lin Lai, and Chin-Lien Liu. Design and analysis of a frame-oriented dynamic bandwidth allocation scheme for triple-play services over EPONs. Computer Networks (Amsterdam, Netherlands: 1999), 64 (??):339–352, May 8, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000905.

Gitzenis:2014:EWN

[1220] S. Gitzenis, G. S. Paschos, and L. Tassiulas. Enhancing wireless networks with caching: Asymptotic laws, sustainability & trade-offs. Computer Networks (Amsterdam, Netherlands: 1999), 64(??):353–368, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000802.

Fu:2014:CDP

[1221] Yongquan Fu, Yijie Wang, and Wei Peng. CommonFinder: a decentralized and privacy-preserving common-friend measurement method for the distributed online social networks. Computer Networks (Amsterdam, Netherlands: 1999), 64(??): 369–389, May 8, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000863.

Yigit:2014:IBM

[1222] Melike Yigit, Ozlem Durmaz Incel, and Vehbi Cagri Gungor. On the interdependency between multichannel scheduling and tree-based routing for WSNs in smart grid environments. Computer Networks (Amsterdam, Netherlands: 1999), 65 (??):1–20, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000917.

Sousa:2014:MTN

[1223] Bruno Sousa, Kostas Pentikousis, and Marilia Curado. MeTHODI-CAL: Towards the next generation of multihomed applications. Computer Networks (Amsterdam, Netherlands: 1999), 65(??):21–40, June 2, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000930.

Lee:2014:STN

[1224] Keontaek Lee, Hak-Jin Kim, Sunju Park, and Seungjae Han. Satisfying the target network lifetime in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 65(??):41–55, June 2, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000942.

Lombera:2014:PPP

[1225] Isaí Michel Lombera, Louise E. Moser, P. Michael Melliar-Smith, and Yung-Ting Chuang. Peer-to-peer publication, search and retrieval using the Android mobile platform. Computer Networks (Amsterdam, Netherlands: 1999), 65(??):56–72, June 2, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000954. See

corrigendum [1783].

Liu:2014:EDD

[1226] Bing-Hong Liu and Jyun-Yu Jhang.

Efficient distributed data scheduling algorithm for data aggregation in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 65(??):73–83, June 2, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000978.

Lee:2014:NMP

[1227] Sihyung Lee, Kyriaki Levanti, and Hyong S. Kim. Network monitoring: Present and future. Computer Networks (Amsterdam, Netherlands: 1999), 65(??):84–98, June 2, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400111X.

Khalil:2014:CIM

[1228] Issa Khalil, Abdallah Khreishah, and Muhammad Azeem. Consolidated Identity Management System for secure mobile cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 65(??): 99–110, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001194.

Bolla:2014:EAM

[1229] Raffaele Bolla, Roberto Bruschi, and Paolo Lago. Energy adaptation in multi-core software routers. Computer Networks (Amsterdam, Netherlands: 1999), 65(??):111–128, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000966.

Shams:2014:SRA

[1230] Farshad Shams, Giacomo Bacci, and Marco Luise. A survey on resource allocation techniques in OFDM(A) networks. Computer Networks (Amsterdam, Netherlands: 1999), 65(??): 129–150, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001212.

Bonetto:2014:EEA

[1231] Edoardo Bonetto, Alessandro Finamore, Marco Mellia, and Riccardo Fiandra. Energy efficiency in access and aggregation networks: From current traffic to potential savings. Computer Networks (Amsterdam, Netherlands: 1999), 65(??): 151–166, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001121.

Ho:2014:MOT

[1232] Trong-Viet Ho, Yves Deville, and Olivier Bonaventure. Multi-objective traffic engineering for data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 65(??): 167–182, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001224.

Maity:2014:SOP

[1233] Soumyadev Maity and R. C. Hansdah. Self-organized public key management in MANETs with enhanced security and without certificatechains. Computer Networks (Amsterdam, Netherlands: 1999), 65(??): 183–211, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001200.

Fofack:2014:PEH

[1234] Nicaise Choungmo Fofack, Philippe Nain, Giovanni Neglia, and Don Towsley. Performance evaluation of hierarchical TTL-based cache networks. Computer Networks (Amsterdam, Netherlands: 1999), 65(??): 212–231, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001108.

Battaglia:2014:RBR

[1235] Caroline Battaglia, Véronique Bruyère, Olivier Gauwin, Cristel Pelsser, and Bruno Quoitin. Reasoning on BGP routing filters using tree automata. Computer Networks (Amsterdam, Netherlands: 1999), 65(??): 232–254, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400019X.

Gong:2014:FBC

[1236] Y. Gong, D. Rossi, C. Testa, S. Valenti, and M. D. Täht. Fighting the bufferbloat: On the coexistence of AQM and low priority congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 65(??): 255–267, June 2, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000188.

Silvester:2014:PLK

[1237] John Silvester and Parviz Kermani. Professor Leonard Kleinrock — Tribute Volume. Computer Networks (Amsterdam, Netherlands: 1999), 66 (??):1–3, June 19, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001467.

Kermani:2014:RSC

[1238] Parviz Kermani and Leonard Kleinrock. Reprint of “Virtual cutthrough: a new computer communication switching technique”. Computer Networks (Amsterdam, Netherlands: 1999), 66(??):4–17, June 19, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001510.

Fratta:2014:FDY

[1239] Luigi Fratta, Mario Gerla, and Leonard Kleinrock. Flow Deviation: 40 years of incremental flows for packets, waves, cars and tunnels. Computer Networks (Amsterdam, Netherlands: 1999), 66 (??):18–31, June 19, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001443.

Tychogiorgos:2014:OBR

[1240] G. Tychogiorgos and K. K. Leung.

Optimization-based resource allocation in communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 66(??):32–45, June 19, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001170.

Yemini:2014:BPP

[1241] Yechiam Yemini. A balance of power principle for decentralized resource sharing. Computer Networks (Amsterdam, Netherlands: 1999), 66 (??):46–51, June 19, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001157.

Afek:2014:RDH

[1242] Y. Afek, A. Bremler-Barr, and L. Schiff. Recursive design of hardware priority queues. Computer Networks (Amsterdam, Netherlands: 1999), 66 (??):52–67, June 19, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001145.

Ben-Porat:2014:CNP

[1243] Udi Ben-Porat, Anat Bremler-Barr, and Hanoch Levy. Computer and network performance: Graduating from the “Age of Innocence”. Computer Networks (Amsterdam, Netherlands: 1999), 66(??):68–81, June 19, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001236.

Marsan:2014:QSS

[1244] Marco Ajmone Marsan and Michela Meo. Queueing systems to study the energy consumption of a campus WLAN. Computer Networks (Amsterdam, Netherlands: 1999), 66 (??):82–93, June 19, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001169.

Ding:2014:SDN

[1245] Aaron Yi Ding, Jon Crowcroft, Sasu Tarkoma, and Hannu Flinck. Software defined networking for security enhancement in wireless mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 66 (??):94–101, June 19, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001133.

Takagi:2014:CSS

[1246] Hideaki Takagi. From computer science to service science: Queues with human customers and servers. Computer Networks (Amsterdam, Netherlands: 1999), 66(??):102–111, June 19, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001182.

Kurose:2014:ICN

[1247] Jim Kurose. Information-centric networking: the evolution from circuits to packets to content. Computer Networks (Amsterdam, Netherlands: 1999), 66 (??):112–120, June 19, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001455.

Ferrari:2014:AAT

[1248] Paolo Ferrari, Emiliano Sisinni, Alessandra Flammini, and Alessandro Depari. Adding accurate timestamping capability to wireless networks for smart grids. Computer Networks (Amsterdam, Netherlands:

1999), 67(??):1–13, July 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001091.

Hammad:2014:NMV

[1249] Ali Hammad, Reza Nejabati, and Dimitra Simeonidou. Novel methods for virtual network composition. Computer Networks (Amsterdam, Netherlands: 1999), 67(??):14–25, July 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400125X.

Romero-Tris:2014:DSP

[1250] Cristina Romero-Tris, Jordi Castellà-Roca, and Alexandre Viejo. Distributed system for private Web search with untrusted partners. Computer Networks (Amsterdam, Netherlands: 1999), 67(??):26–42, July 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001261.

Mouradian:2014:RLR

[1251] Alexandre Mouradian, Isabelle Augé-Blum, and Fabrice Valois. RTXP: a localized real-time MAC-routing protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 67 (??):43–59, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001248.

Ahmadi:2014:LSA

[1252] H. Ahmadi, Y. H. Chew, N. Reyhani, C. C. Chai, and L. A. DaSilva. Learning solutions for auction-based dynamic spectrum access in multicarrier systems. Computer Networks (Amsterdam, Netherlands: 1999), 67 (??):60–73, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001406.

Kuzlu:2014:CNR

[1253] Murat Kuzlu, Manisa Pipattanasomporn, and Saifur Rahman. Communication network requirements for major smart grid applications in HAN, NAN and WAN. Computer Networks (Amsterdam, Netherlands: 1999), 67 (??):74–88, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001431.

Ray:2014:SRS

[1254] Biplob R. Ray, Jemal Abawajy, and Morshed Chowdhury. Scalable RFID security framework and protocol supporting Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 67(??):89–103, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001273.

Rault:2014:EEW

[1255] Tifenn Rault, Abdelmadjid Bouabdallah, and Yacine Challal. Energy efficiency in wireless sensor networks: a top-down survey. Computer Networks (Amsterdam, Netherlands: 1999), 67 (??):104–122, July 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001418.

Chu:2014:SDL

[1256] Wei-Cheng Chu and Kuo-Feng Ssu. Sink discovery in location-free and mobile-sink wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 67(??):123– 140, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400142X.

Zou:2014:VOV

[1257] Shihong Zou, Xitao Wen, Kai Chen, Shan Huang, Yan Chen, Yongqiang Liu, Yong Xia, and Chengchen Hu. VirtualKnotter: Online virtual machine shuffling for congestion resolving in virtualized datacenter. Computer Networks (Amsterdam, Netherlands: 1999), 67(??):141–153, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400139X.

Chen:2014:SRV

[1258] Lien-Wu Chen, Yu-Chee Tseng, and Kun-Ze Syue. Surveillance on-the-road: Vehicular tracking and reporting by V2V communications. Computer Networks (Amsterdam, Netherlands: 1999), 67(??):154–163, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001480.

Liu:2014:MHR

[1259] I-Hsien Liu, Chuan-Gang Liu, Chien-Tung Lu, Yi-Tsen Kuo, and Jung-Shian Li. A multi-hop resource scheduling algorithm for IEEE 802.16j relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 67 (??):164–179, July 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001479.

Rebollo-Monedero:2014:ODP

[1260] David Rebollo-Monedero, Javier Parra-Arnau, Jordi Forné, and Claudia Diaz. Optimizing the design parameters of threshold pool mixes for anonymity and delay. Computer Networks (Amsterdam, Netherlands: 1999), 67(??): 180–200, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001522.

Koutsopoulos:2014:CSG

[1261] Iordanis Koutsopoulos, Leandros Tassiulas, and Lazaros Gkatzikis. Clientserver games and their equilibria in peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 67(??):201–218, July 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001509.

Lakshmi:2014:HRB

[1262] L. Rajya Lakshmi. Handover rate based dynamic guard bandwidth allocation method for call admission control in IEEE 802.16j Mobile Multihop Relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 67 (??):219–234, July 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614001546.

Benson:2014:BIS

[1263] Glenn Benson, Shiu-Kai Chin, Sean Croston, Karthick Jayaraman, and Susan Older. Banking on interoperability: Secure, interoperable credential management. Computer Networks (Amsterdam, Netherlands: 1999), 67 (??):235–251, July 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614001376.

Schmidt:2014:HPE

[1264] R. de O. Schmidt, R. Sadre, A. Sperotto, H. van den Berg, and A. Pras. A hybrid procedure for efficient link dimensioning. Computer Networks (Amsterdam, Netherlands: 1999), 67 (??):252–269, July 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614001534.

Boutaba:2014:CNC

[1265] Raouf Boutaba, Noura Limam, Stefano Secci, and Tarik Taleb. Cloud networking and communications. Computer Networks (Amsterdam, Netherlands: 1999), 68(??):1–4, August 5, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002072.

Reaz:2014:CIW

[1266] Abu (Sayeem) Reaz, Vishwanath Ramamurthi, Massimo Tornatore, and Biswanath Mukherjee. Cloudintegrated WOBAN: an offloadingenabled architecture for serviceoriented access networks. Computer Networks (Amsterdam, Netherlands: 1999), 68(??):5–19, August 5, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000796.

Hwang:2014:DIA

[1267] Jaehyun Hwang, Joon Yoo, and Nakjung Choi. Deadline and Incast Aware TCP for cloud data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 68 (??):20–34, August 5, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000401.

Amamou:2014:TBM

[1268] Ahmed Amamou, Kamel Haddadou, and Guy Pujolle. A TRILLbased multi-tenant data center network. Computer Networks (Amsterdam, Netherlands: 1999), 68(??): 35–53, August 5, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000851.

Wickboldt:2014:RMI

[1269] Juliano Araujo Wickboldt, Rafael Pereira Esteves, Márcio Barbosa de Carvalho, and Lisandro Zambenedetti Granville. Resource management in IaaS cloud platforms made flexible through programmability. Computer Networks (Amsterdam, Netherlands: 1999), 68 (??):54–70, August 5, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400084X.

Ni:2014:PHA

[1270] Wenda Ni, Changcheng Huang, and Jing Wu. Provisioning high-availability datacenter networks for full bandwidth communication. Computer Networks (Amsterdam, Netherlands: 1999), 68 (??):71–94, August 5, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000929.

Guo:2014:IPL

[1271] Zehua Guo, Mu Su, Yang Xu, Zhemin Duan, Luo Wang, Shufeng Hui, and H. Jonathan Chao. Improving the performance of load balancing in software-defined networks through load variance-based synchronization. Computer Networks (Amsterdam, Netherlands: 1999), 68(??): 95–109, August 5, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000814.

Calyam:2014:VAD

[1272] Prasad Calyam, Sudharsan Rajagopalan, Sripriya Seetharam, Arunprasath Selvadhurai, Khaled Salah, and Rajiv Ramnath. VDC-Analyst: Design and verification of virtual desktop cloud resource allocations. Computer Networks (Amsterdam, Netherlands: 1999), 68(??):110– 122, August 5, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000887.

Wang:2014:MBD

[1273] Yiwen Wang, Sen Su, Alex X. Liu, and Zhongbao Zhang. Multiple bulk data transfers scheduling among datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 68(??): 123–137, August 5, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000838.

Woo:2014:OAA

[1274] Simon S. Woo and Jelena Mirkovic.

Optimal application allocation on multiple public clouds. Computer Networks (Amsterdam, Netherlands: 1999), 68 (??):138–148, August 5, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614000371.

Casas:2014:QEC

[1275] Pedro Casas and Raimund Schatz.

Quality of experience in Cloud services: Survey and measurements. Computer Networks (Amsterdam, Netherlands: 1999), 68(??):149–165, August 5, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000383.

Paul:2014:ADM

[1276] Subharthi Paul, Raj Jain, Mohammed Samaka, and Jianli Pan. Application delivery in multi-cloud environments using software defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 68(??): 166–186, August 5, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000826.

Strijkers:2014:IFC

[1277] Rudolf Strijkers, Marc X. Makkes, Cees de Laat, and Robert Meijer. Internet factories: Creating application-specific networks on-demand. Computer Networks (Amsterdam, Netherlands: 1999), 68(??):187–198, August 5, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614000395.

Anonymous:2014:EBa

[1278] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 68(??):ifc, August 5, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128614002096.

Jin:2014:NAPa

[1279] Xin Jin and Yu-Kwong Kwok. Network aware peer-to-peer media streaming: Capacity or proximity? Computer Networks (Amsterdam, Netherlands: 1999), 69(??):1–18, August 20, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001492.

Buh:2014:ANT

[1280] Tomaz Buh, Roman Trobec, and Andrej Ciglic. Adaptive network-traffic balancing on multi-core software networking devices. Computer Networks (Amsterdam, Netherlands: 1999), 69 (??):19–34, August 20, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001686.

Dvorak:2014:EEN

[1281] Jan Dvorák, Jirí Novák, and Petr Kocourek. Energy efficient network protocol architecture for narrowband power line communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 69(??): 35–50, August 20, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001650.

Bober:2014:BLD

[1282] Wojciech Bober and Chris J. Bleakley. BailighPulse: a low duty cycle data gathering protocol for mostly-off Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands:

1999), 69(??):51–65, August 20, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001716.

Park:2014:QGM

[1283] Soohong Park, Jaehoon (Paul) Jeong, and Choong Seon Hong. QoSguaranteed Mobile IPTV service in heterogeneous access networks. Computer Networks (Amsterdam, Netherlands: 1999), 69(??):66–81, August 20, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001844.

Battat:2014:MMA

[1284] Nadia Battat, Hamida Seba, and Hamamache Kheddouci. Monitoring in mobile ad hoc networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 69(??): 82–100, August 20, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001662.

Traverso:2014:PCH

[1285] S. Traverso, C. Kiraly, E. Leonardi, and M. Mellia. A performance comparison of hose rate controller approaches for P2P-TV applications. Computer Networks (Amsterdam, Netherlands: 1999), 69(??): 101–120, August 20, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001558.

Liu:2014:GET

[1286] Zhongjin Liu, Yong Li, Bo Cui, Li Su, Depeng Jin, and Lieguang Zeng. Grain-Flow: Enable testing for future Internet architectures by per-bit customization. Computer Networks (Amsterdam, Netherlands: 1999), 69(??): 121–132, August 20, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001832.

Kim:2014:EHS

[1287] Ji-Su Kim, Jae-Hyun Kim, and Sunghyun Cho. Enhanced handoff scheme based on efficient uplink quality estimation in LTE-Advanced system. Computer Networks (Amsterdam, Netherlands: 1999), 69(??): 133–146, August 20, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001868.

Wang:2014:HTG

[1288] Hui Wang and George N. Rouskas.

Hierarchical traffic grooming: a tutorial. Computer Networks (Amsterdam, Netherlands: 1999), 69(??): 147–156, August 20, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001820.

Anonymous:2014:EBb

[1289] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 69(??):ifc, August 20, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128614002291.

Yang:2014:TFR

[1290] Yuan Yang, Mingwei Xu, and Qi Li. Towards fast rerouting-based energy efficient routing. Computer Networks (Amsterdam, Netherlands: 1999), 70 (??):1–15, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614001674.

Wu:2014:PEI

[1291] Gaofeng Wu and Gurusamy Mohan. Power-efficient integrated routing of sublambda connection requests with traffic splitting in IP over WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 16–29, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614001704.

Abdelkefi:2014:ASQ

[1292] Atef Abdelkefi, Yuming Jiang, Bjarne Emil Helvik, Gergely Biczók, and Alexandru Calu. Assessing the service quality of an Internet path through end-to-end measurement. Computer Networks (Amsterdam, Netherlands: 1999), 70 (??):30–44, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614001698.

Liang:2014:LOT

[1293] Liang Liang, Yu Ge, Gang Feng, Wei Ni, and Aung Aung Phyo Wai. A low overhead tree-based energy-efficient routing scheme for multi-hop wireless body area networks. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 45–58, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001881.

Tasiopoulos:2014:OAC

[1294] Argyrios G. Tasiopoulos, Christos Tsiaras, and Stavros Toumpis. Optimal and achievable cost/delay tradeoffs in delay-tolerant networks. Computer Networks (Amsterdam, Netherlands: 1999), 70(??):59–74, September 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001960.

Addis:2014:EAJ

[1295] Bernardetta Addis, Danilo Ardagna, Antonio Capone, and Giuliana Carello. Energy-aware joint management of networks and Cloud infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 70(??):75–95, September 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001649.

Chu:2014:LFB

[1296] Wei-Cheng Chu and Kuo-Feng Ssu.

Location-free boundary detection in mobile wireless sensor networks with a distributed approach. Computer Networks (Amsterdam, Netherlands:

1999), 70(??):96–112, September 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128614001959.

Angelopoulos:2014:WET

[1297] Constantinos Marios Angelopoulos, Sotiris Nikoletseas, and Theofanis P. Raptis. Wireless energy transfer in sensor networks with adaptive, limited knowledge protocols. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 113–141, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614001972.

Lee:2014:SRM

[1298] Taeseop Lee, Sangkyu Park, Hyung-Sin Kim, and Saewoong Bahk. Sounding resource management for QoS support in massive MIMO systems. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 142–153, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614001856.

Li:2014:TBS

[1299] Chunxi Li, Changjia Chen, Yong Liu, and Baoxian Zhang. Threshold bipolar scheduling for P2P live streaming. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 154–169, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002060.

Mirzamany:2014:ETE

[1300] Esmat Mirzamany, Aboubaker Lasebae, and Orhan Gemikonakli. An efficient traffic engineering based on multi-topology routing for future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 170–178, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001996.

Wu:2014:CPL

[1301] Celimuge Wu, Yusheng Ji, Juan Xu, Satoshi Ohzahata, and Toshihiko Kato. Coded packets over lossy links: a redundancy-based mechanism for reliable and fast data collection in sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 179–191, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002059.

Kim:2014:EIT

[1302] Kyeong Soo Kim. The effect of ISP traffic shaping on user-perceived performance in broadband shared access networks. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 192–209, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002242.

Terlecky:2014:SSS

[1303] Peter Terlecky, Brian Phelan, Amotz Bar-Noy, Theodore Brown, and Dror Rawitz. Should I stay or should I go? Maximizing lifetime with relays. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 210–224, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002229.

La:2014:OSA

[1304] Quang Duy La, Yong Huat Chew, and Boon-Hee Soong. Oligopolistic spectrum allocation game via market competition under spectrum broker. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 225–239, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002047.

Montagud:2014:DDA

[1305] Mario Montagud, Fernando Boronat, Hans Stokking, and Pablo Cesar. Design, development and assessment of control schemes for IDMS in a standardized RTCP-based solution. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 240–259, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002278.

Goldberg:2014:HSS

[1306] Sharon Goldberg, Michael Schapira, Pete Hummon, and Jennifer Rexford.

How secure are secure interdomain routing protocols? Computer Networks (Amsterdam, Netherlands: 1999), 70 (??):260–287, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614001984.

Kim:2014:DNC

[1307] Hyung-Sin Kim, Jae-Seok Bang, and Yong-Hwan Lee. Distributed network configuration in large-scale low power wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 70 (??):288–301, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002217.

Rodriguez-Gomez:2014:RMD

[1308] Rafael A. Rodríguez-Gómez, Gabriel Maciá-Fernández, Pedro García-Teodoro, Moritz Steiner, and Davide Balzarotti. Resource monitoring for the detection of parasite P2P botnets. Computer Networks (Amsterdam, Netherlands: 1999), 70(??):302– 311, September 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002230.

Yigit:2014:CCS

[1309] Melike Yigit, V. Cagri Gungor, and Selcuk Baktir. Cloud computing for Smart Grid applications. Computer Networks (Amsterdam, Netherlands: 1999), 70 (??):312–329, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002382.

Frangoudis:2014:PSU

[1310] Pantelis A. Frangoudis and George C. Polyzos. On the performance of secure user-centric VoIP communication. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 330–344, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002400.

Figueiredo:2014:MFM

[1311] Sérgio Figueiredo, Carlos Guimarães, Daniel Corujo, and Rui L. Aguiar. MI3M: a framework for media independent multicast mobility management. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 345–365, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002370.

Yigit:2014:PLC

[1312] Melike Yigit, V. Cagri Gungor, Gurkan Tuna, Maria Rangoussi, and Etimad Fadel. Power line communication technologies for smart grid applications: a review of advances and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 70(??): 366–383, September 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002369.

Anonymous:2014:EBc

[1313] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 70(??):ifc, September 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128614002515.

Akyildiz:2014:RTE

[1314] Ian F. Akyildiz, Ahyoung Lee, Pu Wang, Min Luo, and Wu Chou. A roadmap for traffic engineering in SDN–OpenFlow networks. Computer Networks (Amsterdam, Netherlands: 1999), 71(??):1–30, October 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002254.

Mowafi:2014:NAE

[1315] Moad Y. Mowafi, Fahed H. Awad, and Walid A. Aljoby. A novel approach for extracting spatial correlation of visual information in heterogeneous wireless multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 71(??):31–47, October 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002412.

Bhattacharya:2014:SPT

[1316] Abhijit Bhattacharya and Anurag Kumar. A shortest path tree based algorithm for relay placement in a wireless sensor network and its performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 71 (??):48–62, October 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002424.

Wang:2014:GTA

[1317] Wenjing Wang, Mainak Chatterjee, Kevin Kwiat, and Qing Li. A game theoretic approach to detect and coexist with malicious nodes in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 71 (??):63–83, October 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002394.

Condeixa:2014:CAA

[1318] Tiago Condeixa and Susana Sargento. Context-aware adaptive IP mobility anchoring. Computer Networks (Amsterdam, Netherlands: 1999), 71 (??):84–99, October 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002448.

Munoz:2014:DTS

[1319] P. Muñoz, D. Laselva, R. Barco, and P. Mogensen. Dynamic traffic steering based on fuzzy Q-Learning approach in a multi-RAT multi-layer wireless network. Computer Networks (Amsterdam, Netherlands: 1999), 71 (??):100–116, October 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002266.

Asheralieva:2014:JPB

[1320] Alia Asheralieva and Kaushik Mahata. Joint power and bandwidth allocation in IEEE802.22 based cognitive LTE network. Computer Networks (Amsterdam, Netherlands: 1999), 71 (??):117–129, October 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400245X.

Zhang:2014:LOS

[1321] Honggang Zhang, Benyuan Liu, Bin Nie, Zhiyong Xu, Xiayin Weng, and Chao Yu. Leveraging online social friendship to improve data swarming performance. Computer Networks (Amsterdam, Netherlands: 1999), 71 (??):130–143, October 4, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002436.

Anonymous:2014:EBd

[1322] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 71(??):ifc, October 4, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128614002709.

Amadeo:2014:CCW

[1323] Marica Amadeo, Claudia Campolo, Antonella Molinaro, and Giuseppe Ruggeri. Content-centric wireless networking: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 72 (??):1–13, October 29, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002497.

Kiremire:2014:UNM

[1324] Ankunda R. Kiremire, Matthias R. Brust, and Vir V. Phoha. Using network motifs to investigate the influence of network topology on PPMbased IP traceback schemes. Computer Networks (Amsterdam, Netherlands: 1999), 72(??):14–32, October 29, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861400259X.

Jeong:2014:HCB

[1325] Dong Geun Jeong, Jeong Ae Han, and Wha Sook Jeon. Hop capacity balancing in OFDMA relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 72(??): 33–44, October 29, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002485.

Zhang:2014:NML

[1326] Bo Zhang, Zhenhua Huang, and Yang Xiang. A novel multiple-level trust management framework for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 72 (??):45–61, October 29, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002461.

Kang:2014:TBP

[1327] Xin Kang and Yongdong Wu. A trust-based pollution attack prevention scheme in peer-to-peer streaming networks. Computer Networks (Amsterdam, Netherlands: 1999), 72 (??):62–73, October 29, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002667.

Jammal:2014:SDN

[1328] Manar Jammal, Taranpreet Singh, Abdallah Shami, Rasool Asal, and Yiming Li. Software defined networking: State of the art and research challenges. Computer Networks (Amsterdam, Netherlands: 1999), 72 (??):74–98, October 29, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002588.

Savi:2014:CEM

[1329] Michele Savi, Harald Øverby, Norvald Stol, and Carla Raffaelli. Cost evolution model to design optical switching fabrics with wavelength converters. Computer Networks (Amsterdam, Netherlands: 1999), 72(??): 99–112, October 29, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002631.

Nguyen:2014:CIM

[1330] Tien-Thinh Nguyen and Christian Bonnet. Considerations of IP multicast for load balancing in Proxy Mobile IPv6 networks. Computer Networks (Amsterdam, Netherlands: 1999), 72 (??):113–126, October 29, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400262X.

Lu:2014:PLS

[1331] Xiaopei Lu, Dezun Dong, Xiangke Liao, Shanshan Li, and Xiaodong Liu. PathZip: a lightweight scheme for tracing packet path in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 73 (??):1–14, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002813.

Kim:2014:LBB

[1332] Sungwook Kim. Learning based bandwidth management algorithms by using bargaining and fictitious play approaches. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 15–21, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002783.

Stea:2014:CSA

[1333] Giovanni Stea and Antonio Virdis. A comprehensive simulation analysis of LTE Discontinuous Reception (DRX). Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 22–40, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002680.

Wang:2014:ATF

[1334] Ding Wang and Ping Wang. On the anonymity of two-factor authentication schemes for wireless sensor networks: Attacks, principle and solutions. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 41–57, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002643.

Tham:2014:FSW

[1335] Chen-Khong Tham and Tie Luo. Fairness and social welfare in service allocation schemes for participatory sensing. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 58–71, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002679.

Jin:2014:NAPb

[1336] Sunggeun Jin and Daji Qiao. Numerical analysis of the power saving with a bursty traffic model in LTE-Advanced networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 72–83, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002795.

Divakaran:2014:BAD

[1337] Dinil Mon Divakaran, Mohan Gurusamy, and Mathumitha Sellamuthu.

Bandwidth allocation with differential pricing for flexible demands in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 73 (??):84–97, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002837.

Colistra:2014:PTA

[1338] Giuseppe Colistra, Virginia Pilloni, and Luigi Atzori. The problem of task allocation in the Internet of Things and the consensus-based approach. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 98–111, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002655.

Zhang:2014:MCI

[1339] Yaodong Zhang, Yue Wang, Dan Pei, and Jian Yuan. Multi-AS cooperative incoming traffic engineering in a transit-edge separate Internet. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 112–127, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002771.

Khan:2014:DGC

[1340] Furqan Hameed Khan and Young-June Choi. Distributed games for coordinated coalition formation in femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 128–141, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002862.

Pan:2014:EDC

[1341] Meng-Shiuan Pan, Ping-Lin Liu, and Yen-Pei Lin. Event data collection in ZigBee tree-based wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 142–153, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002874.

Cavalieri:2014:LLI

[1342] Salvatore Cavalieri and Ferdinando Chiacchio. Limiting the loss of information in KNXnet/IP on congestion conditions. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 154–172, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002928.

Baadache:2014:SAS

[1343] Abderrahmane Baadache and Ali Belmehdi. Struggling against simple and cooperative black hole attacks in multi-hop wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 173–184, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002801.

Xu:2014:TBH

[1344] Li Xu, Yuan He, Xiaofeng Chen, and Xinyi Huang. Ticket-based handoff authentication for wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 185–194, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002886.

Liu:2014:SWM

[1345] Tong Liu and Yanmin Zhu. Social welfare maximization in participatory smartphone sensing. Computer Networks (Amsterdam, Netherlands: 1999), 73(??):195–209, November 14, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128614002941.

Phe-Neau:2014:UVP

[1346] Tiphaine Phe-Neau, Marcelo Dias de Amorim, and Vania Conan. Uncovering vicinity properties in disruptiontolerant networks. Computer Networks (Amsterdam, Netherlands: 1999), 73 (??):210–223, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002825.

Marnerides:2014:TAD

[1347] A. K. Marnerides, A. Schaeffer-Filho, and A. Mauthe. Traffic anomaly diagnosis in Internet backbone networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 224–243, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002850.

Tian:2014:TBI

[1348] Chunqi Tian, Baijian Yang, Jidong Zhong, and Xiaojian Liu. Trustbased incentive mechanism to motivate cooperation in hybrid P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 244–255, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002904.

Garroppo:2014:GPC

[1349] Rosario G. Garroppo, Gianfranco Nencioni, Luca Tavanti, and Bernard Gendron. The greening potential of content delivery in residential community networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 256–267, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003041.

Hsu:2014:ECA

[1350] Yi-Huai Hsu, Kuochen Wang, and Yu-Chee Tseng. Efficient cooperative access class barring with load balancing and traffic adaptive radio resource management for M2M communications over LTE-A. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 268–281, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614002898.

Coluccia:2014:REM

[1351] A. Coluccia, F. Ricciato, and P. Romirer-Maierhofer. Robust estimation of mean failure probability in access networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 282–301, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003090.

Lee:2014:TOD

[1352] Youngsoo Lee and Chong-Ho Choi. Transmission Order Deducing MAC (TOD-MAC) protocol for CSMA/CA wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 73 (??):302–318, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002849.

Han:2014:ORS

[1353] Min-Hong Han, Byung-Gook Kim, and Jang-Won Lee. Opportunistic resource scheduling for D2D communication in OFDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 73(??): 319–334, November 14, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614002916.

Anonymous:2014:EBe

[1354] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 73(??):ifc, November 14, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128614003338.

Kuperman:2014:NPM

[1355] Greg Kuperman, Eytan Modiano, and Aradhana Narula-Tam. Network protection with multiple availability guarantees. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part A) (??):1–12, December 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400293X.

Lv:2014:PPA

[1356] Jianming Lv, Tieying Zhang, Zhenhua Li, and Xueqi Cheng. PACOM: Parasitic anonymous communication in the BitTorrent network. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part A)(??):13– 33, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400303X.

Shi:2014:DRA

[1357] Lei Shi, Jianghong Han, Dong Han, Xu Ding, and Zhenchun Wei. The dynamic routing algorithm for renewable wireless sensor networks with wireless power transfer. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part A)(??):34–52, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003089.

Ren:2014:OLV

[1358] Dongni Ren, Wang Kit Wong, and S.- H. Gary Chan. Overlay live video streaming with heterogeneous bitrate requirements. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part A)(??):53–63, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003077.

Awad:2014:IAE

[1359] Alaa Awad, Amr Mohamed, Amr A. El-Sherif, and Omar A. Nasr. Interferenceaware energy-efficient cross-layer design for healthcare monitoring applications. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part A) (??):64–77, December 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003119.

Ahmad:2014:RCQ

[1360] Ayaz Ahmad, Naveed Ul Hassan, and Nadir Shah. Robust channel quality indicator reporting for multicarrier and multi-user systems. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part A)(??):78– 88, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003211.

Ansari:2014:COC

[1361] Junaid Ansari, Elena Meshkova, Wasif Masood, Arham Muslim, Janne Riihijärvi, and Petri Mähönen. CON-Fab: Ontology and component based optimization of WSN protocol stacks with deployment feedback. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part A)(??):89– 108, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003065.

Anonymous:2014:EBf

[1362] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part A)(??): ifc, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003491.

Wu:2014:SIM

[1363] Yulei Wu, Fei Hu, and Xinheng Wang.

Special issue on mobile computing for content/service-oriented networking architecture. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B)(??):1–3, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003752.

Matos:2014:CAC

[1364] Ricardo Matos, Carlos Marques, and Susana Sargento. Context-aware control of user-centric virtual networks: Centralized vs distributed approaches. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B)(??):4–21, December 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003272.

Magurawalage:2014:EEN

[1365] Chathura M. Sarathchandra Magurawalage, Kun Yang, Liang Hu, and Jianming Zhang. Energy-efficient and network-aware offloading algorithm for mobile cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B)(??):22– 33, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003193.

Vingelmann:2014:AAD

[1366] Péter Vingelmann, Janus Heide, Morten Videbæk Pedersen, Qi Zhang, and Frank H. P. Fitzek. All-to-all data dissemination with network coding in dynamic MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B)(??):34– 47, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003144.

Wang:2014:ECM

[1367] Gaocai Wang, Ying Peng, Peng Feng, and Nao Wang. An energy consumption minimization routing scheme based on rate adaptation with QoS guarantee for the mobile environment. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B) (??):48–57, December 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400317X.

Dong:2014:MAB

[1368] Mianxiong Dong, Kaoru Ota, Laurence T. Yang, Shan Chang, Hongzi Zhu, and Zhenyu Zhou. Mobile agentbased energy-aware and user-centric data collection in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B) (??):58–70, December 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003156.

Wei:2014:AMO

[1369] Guiyi Wei, Yang Xiang, Min Ji, and Ping Zhu. An analytical model for optimal spectrum leasing under constraints of quality of service in CRNs. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B) (??):71–80, December 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003120.

Liu:2014:SOS

[1370] Quan Liu, Hongwei Niu, Wenjun Xu, and Duzhong Zhang. A serviceoriented spectrum allocation algorithm using enhanced PSO for cognitive wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B)(??):81–91, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400320X.

Jan:2014:PPB

[1371] Mian Ahmad Jan, Priyadarsi Nanda, Xiangjian He, and Ren Ping Liu. PASCCC: Priority-based applicationspecific congestion control clustering protocol. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B)(??):92–102, December 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003168.

Peng:2014:CSW

[1372] Sancheng Peng, Min Wu, Guojun Wang, and Shui Yu. Containing smartphone worm propagation with an influence maximization algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B)(??):103– 113, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003132.

Kaushik:2014:PSA

[1373] Bhanu Kaushik, Honggang Zhang, Xinyu Yang, Xinwen Fu, Benyuan Liu, and Jie Wang. Providing service assurance in mobile opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B) (??):114–140, December 9, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003181.

Anonymous:2014:EBg

[1374] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 74 (part B)(??): ifc, December 9, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003776.

Ciftcioglu:2014:OIC

[1375] Ertugrul N. Ciftcioglu, Antonios Michaloliakos, Aylin Yener, Konstantinos Psounis, Thomas F. La Porta, and Ramesh Govindan. Operational information content sum capacity: From theory to practice. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):1–17, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003053.

Capela:2014:MNC

[1376] Nelson Capela and Susana Sargento. Multihoming and network coding: a new approach to optimize the network performance. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):18– 36, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003107.

Kuptsov:2014:HPL

[1377] Dmitriy Kuptsov, Boris Nechaev, Andrey Lukyanenko, and Andrei Gurtov. How penalty leads to improvement: a measurement study of wireless backoff in IEEE 802.11 networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A) (??):37–57, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003223.

Park:2014:RSP

[1378] Hanjin Park, Yung Yi, and Yongdae Kim. Revisiting security of proportional fair scheduler in wireless cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):58–74, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003314.

Sou:2014:BCR

[1379] Sok-Ian Sou and Dung-Ru Tsai. Bulk credit reservation in eventbased Machine Type Communications for 3GPP online charging. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):75– 85, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003442.

Ren:2014:DJO

[1380] Dongni Ren, S.-H. Gary Chan, Guangyu Shi, and Hongbo Zhang. Distributed joint optimization for large-scale video-on-demand. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):86– 98, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003478.

Noura:2014:ERE

[1381] Hassan Noura, Steven Martin, Khaldoun Al Agha, and Khaled Chahine. ERSS–RLNC: Efficient and robust secure scheme for random linear network coding. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A) (??):99–112, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003466.

Tuysuz:2014:EEQ

[1382] Mehmet Fatih Tuysuz. An energyefficient QoS-based network selection scheme over heterogeneous WLAN–3G networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):113–133, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003454.

Shen:2014:ADL

[1383] Xiang-Jun Shen, Lu Liu, Zheng-Jun Zha, Pei-Ying Gu, Zhong-Qiu Jiang, Ji-Ming Chen, and John Panneerselvam. Achieving dynamic load balancing through mobile agents in small world P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):134– 148, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400187X.

Tung:2014:RPC

[1384] Li-Ping Tung, Ying-Dar Lin, Yu-Hsien Kuo, Yuan-Cheng Lai, and Krishna M.

Sivalingam. Reducing power consumption in LTE data scheduling with the constraints of channel condition and QoS. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A) (??):149–159, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003570.

Jia:2014:UMP

[1385] Wen-Kang Jia. A unified MIPv6 and PMIPv6 route optimization scheme for heterogeneous mobility management domains. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A) (??):160–176, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003661.

Hassan:2014:REC

[1386] Dina S. M. Hassan, Hossam M. A. Fahmy, and Ayman M. Bahaa-ElDin. RCA: Efficient connected dominated clustering algorithm for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):177–191, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400365X.

Zhang:2014:GCS

[1387] Yinghui Zhang, Xiaofeng Chen, Jin Li, and Hui Li. Generic construction for secure and efficient handoff authentication schemes in EAP-based wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):192–211, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003648.

Wang:2014:CCP

[1388] Hui Wang, H. Eduardo Roman, Liyong Yuan, Yongfeng Huang, and Rongli Wang. Connectivity, coverage and power consumption in largescale wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):212– 225, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003636.

Sergiou:2014:CCW

[1389] Charalambos Sergiou, Vasos Vassiliou, and Aristodemos Paphitis. Congestion control in Wireless Sensor Networks through dynamic alternative path selection. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A) (??):226–238, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003624.

Addis:2014:ECR

[1390] B. Addis, A. Capone, G. Carello, L. G.

Gianoli, and B. Sansò. On the energy cost of robustness and resiliency in IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):239–259, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003594.

Anglano:2014:MPG

[1391] Cosimo Anglano, Marco Guazzone, and Matteo Sereno. Maximizing profit in green cellular networks through collaborative games. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):260– 275, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003582.

Qiao:2014:CEW

[1392] Xiuquan Qiao, Guoshun Nan, Wei Tan, Lei Guo, Junliang Chen, Wei Quan, and Yukai Tu. CCNxTomcat: an extended web server for Content-Centric Networking. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):276–296, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003697.

Bianzino:2014:CPO

[1393] Aruna Prem Bianzino, Mikael Asplund, Ekhiotz Jon Vergara, and Simin Nadjm-Tehrani. Cooperative proxies: Optimally trading energy and quality of service in mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A) (??):297–312, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003685.

Jin:2014:VCP

[1394] Xin Jin, Yu-Kwong Kwok, and Jian Deng. Variegated competing peer-to-peer systems with selfish peers. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):313– 330, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003727.

Bagaa:2014:IPF

[1395] Miloud Bagaa, Mohamed Younis, Abdelouahid Derhab, and Nadjib Badache. Intertwined path formation and MAC scheduling for fast delivery of aggregated data in WSN. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):331– 350, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003739.

Shan:2014:OEE

[1396] Feng Shan, Junzhou Luo, and Xiaojun Shen. Optimal energy efficient packet scheduling with arbitrary individual deadline guarantee. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):351– 366, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003843.

Ming:2014:INC

[1397] Zhongxing Ming, Mingwei Xu, and Dan Wang. InCan: In-network cache assisted eNodeB caching mechanism in 4G LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):367– 380, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003831.

Villas:2014:DER

[1398] Leandro Aparecido Villas, Azzedine Boukerche, Guilherme Maia, Richard Werner Pazzi, and Antonio A. F. Loureiro. DRIVE: an efficient and robust data dissemination protocol for highway and urban vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):381–394, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003673.

Friginal:2014:SEP

[1399] Jesús Friginal, David de Andrés, Juan-Carlos Ruiz, and Miquel Martínez. A survey of evaluation platforms for ad hoc routing protocols: a resilience perspective. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A) (??):395–413, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003430.

Hlavacek:2014:LAC

[1400] Deanna Hlavacek and J. Morris Chang. A layered approach to cognitive radio network security: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):414– 436, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003569.

Paul:2014:SDO

[1401] Thomas Paul, Antonino Famulari, and Thorsten Strufe. A survey on decentralized Online Social Networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??):437– 452, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003600.

Hakiri:2014:SDN

[1402] Akram Hakiri, Aniruddha Gokhale, Pascal Berthou, Douglas C. Schmidt, and Thierry Gayraud. Softwaredefined networking: Challenges and research opportunities for Future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A) (??):453–471, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003703.

Anonymous:2014:EBh

[1403] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part A)(??): ifc, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004046.

Klier:2014:CPU

[1404] Julia Klier, Mathias Klier, and Rolf T. Wigand. The connectedness, pervasiveness and ubiquity of online social networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B) (??):473–476, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614004332.

Mutter:2014:DTM

[1405] Tobias Mutter and Dennis Kundisch. Don’t take away my status! — evidence from the restructuring of a virtual reward system. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B)(??):477– 490, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003247.

Yan:2014:BAC

[1406] Gongjun Yan, Wu He, Jiancheng Shen, and Chuanyi Tang. A bilingual approach for conducting Chinese and English social media sentiment analysis. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B)(??): 491–503, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003235.

Tran:2014:PEB

[1407] Duc A. Tran and Ting Zhang. S-PUT: an EA-based framework for socially aware data partitioning. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B)(??):504– 518, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003296.

Putzke:2014:CCG

[1408] Johannes Putzke, Kai Fischbach, Detlef Schoder, and Peter A. Gloor. Cross-cultural gender differences in the adoption and usage of social media platforms — an exploratory study of Last.FM. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B)(??):519–530, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003302.

Patsakis:2014:DPP

[1409] Constantinos Patsakis, Athanasios Zigomitros, Achilleas Papageorgiou, and Edgar Galván-López. Distributing privacy policies over multimedia content across multiple online social networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B) (??):531–543, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003259.

Zhao:2014:WFE

[1410] Junzhou Zhao, John C. S. Lui, Don Towsley, and Xiaohong Guan. Whom to follow: Efficient followee selection for cascading outbreak detection on online social networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B)(??):544– 559, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003260.

Behrendt:2014:MMA

[1411] Sebastian Behrendt, Alexander Richter, and Matthias Trier. Mixed methods analysis of enterprise social networks. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B) (??):560–577, December 24, 2014. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003284.

Anonymous:2014:EBi

[1412] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 75 (part B)(??): ifc, December 24, 2014. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614004277.

Anonymous:2015:TYR

[1413] Anonymous. Thank you reviewers. Computer Networks (Amsterdam, Netherlands: 1999), 76(??):i– xxxi, January 15, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614004575.

Anonymous:2015:E

[1414] Anonymous. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 76(??):iii–iv, January 15, 2015.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400454X.

Correia:2015:DRM

[1415] Luiz H. A. Correia, Thanh-Dien Tran, Vasco N. S. S. Pereira, João C. Giacomin, and Jorge M. Sá Silva. DynMAC: a resistant MAC protocol to coexistence in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 76 (??):1–16, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003740.

Gyo:2015:SHM

[1416] Győzö Gódor, Zoltán Jakó, Ádám Knapp, and Sándor Imre. A survey of handover management in LTEbased multi-tier femtocell networks: Requirements, challenges and solutions. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 17–41, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003715.

Acedo-Hernandez:2015:AIP

[1417] R. Acedo-Hernández, M. Toril, S. Luna-Ramírez, I. de la Bandera, and N. Faour. Analysis of the impact of PCI planning on downlink throughput performance in LTE. Computer Networks (Amsterdam, Netherlands: 1999), 76 (??):42–54, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003855.

Fafoutis:2015:RIM

[1418] Xenofon Fafoutis, Alessio Di Mauro, Madava D. Vithanage, and Nicola Dragoni. Receiver-initiated medium access control protocols for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 76 (??):55–74, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003910.

Bujlow:2015:ICP

[1419] Tomasz Bujlow, Valentín Carela-Español, and Pere Barlet-Ros. Independent comparison of popular DPI tools for traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 76(??):75–89, January 15, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614003909.

Xenakis:2015:AMM

[1420] Dionysis Xenakis, Nikos Passas, Lazaros Merakos, and Christos Verikoukis. Advanced mobility management for reduced interference and energy consumption in the two-tier LTE-Advanced network. Computer Networks (Amsterdam, Netherlands: 1999), 76(??):90–111, January 15, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128614003867.

Munoz:2015:LBH

[1421] P. Muñoz, R. Barco, and I. de la Bandera. Load balancing and handover joint optimization in LTE networks using Fuzzy Logic and Reinforcement Learning. Computer Networks (Amsterdam, Netherlands: 1999), 76 (??):112–125, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003892.

Mangili:2015:CAM

[1422] Michele Mangili, Fabio Martignon, and Stefano Paraboschi. A cacheaware mechanism to enforce confidentiality, trackability and access policy evolution in Content-Centric Networks. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 126–145, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003995.

Sicari:2015:SPT

[1423] S. Sicari, A. Rizzardi, L. A. Grieco, and A. Coen-Porisini. Security, privacy and trust in Internet of Things: the road ahead. Computer Networks (Amsterdam, Netherlands: 1999), 76 (??):146–164, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003971.

Chang:2015:ASM

[1424] Tengfei Chang, Thomas Watteyne, Kris Pister, and Qin Wang. Adaptive synchronization in multi-hop TSCH networks. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 165–176, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003922.

Wu:2015:CLO

[1425] Yuansheng Wu, Bing Guo, Yan Shen, Jihe Wang, and Xiaobin Liu. A Cross-Layer Optimization and Design approach under QoS constraints for green IP over WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 177–190, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003879.

Wang:2015:TAC

[1426] Xu Wang, Hailong Sun, Ting Deng, and Jinpeng Huai. On the tradeoff of availability and consistency for quorum systems in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 191–206, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003958.

Akbas:2015:PAM

[1427] Mustafa Ilhan Akbas, Matthias R. Brust, Damla Turgut, and Carlos H. C. Ribeiro. A preferential attachment model for primate social networks. Computer Networks (Amsterdam, Netherlands: 1999), 76(??):

207–226, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003983.

Cavdar:2015:IOF

[1428] Tugrul Çavdar, Erkan Güler, and Zhaleh Sadreddini. Instant overbooking framework for cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 227–241, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004009.

Ohsita:2015:ATI

[1429] Yuichi Ohsita, Takashi Miyamura, Shin’ichi Arakawa, Shohei Kamamura, Daisaku Shimazaki, Kohei Shiomoto, Atsushi Hiramatsu, and Masayuki Murata. Aggregation of traffic information for hierarchical routing reconfiguration. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 242–258, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004149.

Fornasa:2015:BLS

[1430] Martino Fornasa, Michele Stecca, Massimo Maresca, and Pierpaolo Baglietto. Bounded latency spanning tree reconfiguration. Computer Networks (Amsterdam, Netherlands: 1999), 76 (??):259–274, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003934.

Gonen:2015:ANM

[1431] Bilal Gonen, Gurhan Gunduz, and Murat Yuksel. Automated network management and configuration using Probabilistic Trans-Algorithmic Search. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 275–293, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004022.

Gao:2015:GBC

[1432] Chao Gao, Guorong Zhao, Jianhua Lu, and Shuang Pan. A gridbased cooperative QoS routing protocol with fading memory optimization for navigation carrier ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 294–316, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004356.

Erel:2015:GSG

[1433] Müge Erel, Zemre Arslan, Yusuf Özçevik, and Berk Canberk. Grade of Service (GoS) based adaptive flow management for Software Defined Heterogeneous Networks (SD-HetN). Computer Networks (Amsterdam, Netherlands: 1999), 76(??): 317–330, January 15, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004010.

Anonymous:2015:EBa

[1434] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 76(??):ifc, January 15, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128614004496.

Iacovazzi:2015:PTP

[1435] Alfonso Iacovazzi and Andrea Baiocchi.

Protecting traffic privacy for massive aggregated traffic. Computer Networks (Amsterdam, Netherlands: 1999), 77 (??):1–17, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400437X.

Domzal:2015:SMP

[1436] Jerzy Domzal, Zbigniew Duliński, Miroslaw Kantor, Jacek Rzasa, Rafal Stankiewicz, Krzysztof Wajda, and Robert Wójcik. A survey on methods to provide multipath transmission in wired packet networks. Computer Networks (Amsterdam, Netherlands: 1999), 77(??):18–41, February 11, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128614004435.

Albano:2015:RVE

[1437] Michele Albano and Stefano Chessa.

Replication vs erasure coding in data centric storage for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 42–55, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004368.

Wang:2015:TSA

[1438] Dexiang Wang. Tradeoff study among traffic egression schemes and member allocation optimization for link aggregation groups in integrated switching systems. Computer Networks (Amsterdam, Netherlands: 1999), 77 (??):56–72, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004447.

Accongiagioco:2015:BSB

[1439] Giovanni Accongiagioco, Enrico Gregori, and Luciano Lenzini. S-BITE: a Structure-Based Internet Topology gEnerator. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 73–89, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400396X.

Fowler:2015:AEE

[1440] Scott Fowler, Ahmed Omar Shahidullah, Mohammed Osman, Johan M. Karlsson, and Di Yuan. Analytical evaluation of extended DRX with additional active cycles for light traffic. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 90–102, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004381.

Gharakheili:2015:CEH

[1441] Hassan Habibi Gharakheili, Arun Vishwanath, and Vijay Sivaraman. Comparing edge and host traffic pacing in small buffer networks. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 103–116, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004393.

Neumayer:2015:GMF

[1442] Sebastian Neumayer, Alon Efrat, and Eytan Modiano. Geographic max-flow and min-cut under a circular disk failure model. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 117–127, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003880.

Tashtarian:2015:OOD

[1443] Farzad Tashtarian, M. H. Yaghmaee Moghaddam, Khosrow Sohraby, and Sohrab Effati. ODT: Optimal deadline-based trajectory for mobile sinks in WSN: a decision tree and dynamic programming approach. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 128–143, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004459.

Vo:2015:CGA

[1444] Phuong Luu Vo, Duc Ngoc Minh Dang, Sungwon Lee, Choong Seon Hong, and Quan Le-Trung. A coalitional game approach for fractional cooperative caching in content-oriented networks. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 144–152, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004587.

Kretsis:2015:MCB

[1445] Aristotelis Kretsis, Panagiotis Kokkinos, Kostas Christodoulopoulos, Theodora Varvarigou, and Emmanouel (Manos) Varvarigos. Mantis: Cloud-based optical network planning and operation tool. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 153–168, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004253.

Shanbhag:2015:VSS

[1446] Shashank Shanbhag, Arun Reddy Kandoor, Cong Wang, Ramgopal Mettu, and Tilman Wolf. VHub: Single-stage virtual network mapping through hub location. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 169–180, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004605.

Lee:2015:FMA

[1447] Myungjin Lee, Mohammad Hajjat, Ramana Rao Kompella, and Sanjay G. Rao. A flow measurement architecture to preserve application structure. Computer Networks (Amsterdam, Netherlands: 1999), 77(??): 181–195, February 11, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003946.

Anonymous:2015:EBb

[1448] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 77(??):ifc, February 11, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128615000122.

Anonymous:2015:EBc

[1449] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 78(??):ibc, February 26, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861500033X.

Combes:2015:OOC

[1450] Richard Combes, Salah Eddine Elayoubi, Arshad Ali, Louai Saker, and Tijani Chahed. Optimal online control for sleep mode in green base stations. Computer Networks (Amsterdam, Netherlands: 1999), 78(??): ibc, February 26, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004216.

Fang:2015:EED

[1451] Chao Fang, F. Richard Yu, Tao Huang, Jiang Liu, and Yunjie Liu. An energy-efficient distributed in-network caching scheme for green contentcentric networks. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004186.

Ganji:2015:GCW

[1452] Fatemeh Ganji, Lukasz Budzisz, Fikru G. Debele, Nanfang Li, Michela Meo, Marco Ricca, Yi Zhang, and Adam Wolisz. Greening campus WLANs: Energy-relevant usage and mobility patterns. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004241.

Heddeghem:2015:PCS

[1453] Ward Van Heddeghem, Filip Idzikowski, Francesco Musumeci, Achille Pattavina, Bart Lannoo, Didier Colle, and Mario Pickavet. A power consumption sensitivity analysis of circuitswitched versus packet-switched backbone networks. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004150.

Huan:2015:RRP

[1454] Xiaoli Huan, Bang Wang, Yijun Mo, and Laurence T. Yang. Rechargeable router placement based on efficiency and fairness in green wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004411.

Lambert:2015:EEA

[1455] S. Lambert, B. Lannoo, A. Dixit, D. Colle, M. Pickavet, J. Montalvo, J. A. Torrijos, and P. Vetter. Energy efficiency analysis of high speed tripleplay services in next-generation PON deployments. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004472.

Ngoc:2015:NPP

[1456] Nam Pham Ngoc, Thanh Nguyen Huu, Trong Vu Quang, Vu Tran Hoang, Huong Truong Thu, Phuoc Tran-Gia, and Christian Schwartz. A new power profiling method and power scaling mechanism for energy-aware NetFPGA gigabit router. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004423.

Peng:2015:OPM

[1457] J. Peng, P. Hong, and K. Xue. Optimal power management under delay constraint in cellular networks with hybrid energy sources. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614004174.

Polverini:2015:FFF

[1458] Marco Polverini, Antonio Cianfrani, Angelo Coiro, Marco Listanti, and Roberto Bruschi. Freezing forwarding functionality to make the network greener. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861400440X.

Rengarajan:2015:EOB

[1459] Balaji Rengarajan, Gianluca Rizzo, and Marco Ajmone Marsan. Energyoptimal base station density in cellular access networks with sleep modes. Computer Networks (Amsterdam, Netherlands: 1999), 78(??): ibc, February 26, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614004228.

Serrano:2015:SIG

[1460] Pablo Serrano, Xavier Costa-Pérez, Jinsong Wu, and Ken Christensen. Special issue: Green communications. Computer Networks (Amsterdam, Netherlands: 1999), 78(??): ibc, February 26, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000390.

Son:2015:EED

[1461] Kyuho Son, Eunsung Oh, and Bhaskar Krishnamachari. Energy-efficient design of heterogeneous cellular networks from deployment to operation. Computer Networks (Amsterdam, Netherlands: 1999), 78(??): ibc, February 26, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400423X.

Suarez:2015:EEB

[1462] Luis Suárez, Loutfi Nuaymi, and Jean-Marie Bonnin. Energy-efficient BS switching-off and cell topology management for macro/femto environments. Computer Networks (Amsterdam, Netherlands: 1999), 78(??): ibc, February 26, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004162.

Wang:2015:EEQ

[1463] Jiayuan Wang, Xin Chen, Chris Phillips, and Ying Yan. Energy efficiency with QoS control in dynamic optical networks with SDN enabled integrated control plane. Computer Networks (Amsterdam, Netherlands: 1999), 78(??):ibc, February 26, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004198.

Yigitel:2015:QVE

[1464] M. Aykut Yigitel, Ozlem Durmaz Incel, and Cem Ersoy. QoS vs. energy: a traffic-aware topology management scheme for green heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 78 (??):ibc, February 26, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004204.

Frangoudis:2015:RBC

[1465] Pantelis A. Frangoudis and George C. Polyzos. Reputation-based crowdsourced Wi-Fi topology discovery. Computer Networks (Amsterdam, Netherlands: 1999), 79(??):1–16, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128614004666.

Wang:2015:PAF

[1466] Wei Wang, Linlin Yang, Yanjiao Chen, and Qian Zhang. A privacyaware framework for targeted advertising. Computer Networks (Amsterdam, Netherlands: 1999), 79(??): 17–29, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861400471X.

Al-Nakhala:2015:DAW

[1467] Noor Al-Nakhala, Ryan Riley, and Tarek Elfouly. Distributed algorithms in wireless sensor networks: an approach for applying binary consensus in a real testbed. Computer Networks (Amsterdam, Netherlands: 1999), 79 (??):30–38, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004654.

Tauhiduzzaman:2015:FPA

[1468] Md. Tauhiduzzaman and Mea Wang.

Fighting pollution attacks in P2P streaming. Computer Networks (Amsterdam, Netherlands: 1999), 79(??): 39–52, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004630.

Jeon:2015:CPE

[1469] Seil Jeon, Namhi Kang, Daniel Corujo, and Rui L. Aguiar. Comprehensive performance evaluation of distributed and dynamic mobility routing strategy. Computer Networks (Amsterdam, Netherlands: 1999), 79 (??):53–67, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000043.

Somanatha:2015:RAK

[1470] Revathi Bangalore Somanatha and J. William Atwood. Router authentication, key management, and adjacency management for securing interrouter control messages. Computer Networks (Amsterdam, Netherlands: 1999), 79(??):68–90, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004691.

Kim:2015:PEP

[1471] Young-Hwan Kim, Roberto Konow, Diego Dujovne, Thierry Turletti, Walid Dabbous, and Gonzalo Navarro. PcapWT: an efficient packet extraction tool for large volume network traces. Computer Networks (Amsterdam, Netherlands: 1999), 79(??): 91–102, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004617.

Mostefaoui:2015:SAS

[1472] A. Mostefaoui, A. Boukerche, M. A. Merzoug, and M. Melkemi. A scalable approach for serial data fusion in Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 79(??):103–119, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004629.

Rojas:2015:APB

[1473] Elisa Rojas, Guillermo Ibañez, Jose Manuel Gimenez-Guzman, Juan A. Carral, Alberto Garcia-Martinez, Isaias Martinez-Yelmo, and Jose Manuel Arco. All-path bridging: Path exploration protocols for data center and campus networks. Computer Networks (Amsterdam, Netherlands: 1999), 79 (??):120–132, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000055.

Han:2015:FDH

[1474] Sangyup Han, Myungchul Kim, Ben Lee, and Sungwon Kang. Fast Directional Handoff and lightweight retransmission protocol for enhancing multimedia quality in indoor WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 79(??): 133–147, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000031.

Goscien:2015:TSA

[1475] Róza Goścień, Krzysztof Walkowiak, and Miroslaw Klinkowski. Tabu search algorithm for routing, modulation and spectrum allocation in elastic optical network with anycast and unicast traffic. Computer Networks (Amsterdam, Netherlands: 1999), 79 (??):148–165, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004460.

Mahmood:2015:RWS

[1476] Muhammad Adeel Mahmood, Winston K. G. Seah, and Ian Welch. Reliability in wireless sensor networks: a survey and challenges ahead. Computer Networks (Amsterdam, Netherlands: 1999), 79(??):166–187, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004708.

Silva:2015:NMM

[1477] Fabrício A. Silva, Azzedine Boukerche, Thais R. M. B. Silva, Linnyer B. Ruiz, and Antonio A. F. Loureiro. A novel macroscopic mobility model for vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 79 (??):188–202, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000067.

Romero:2015:LSP

[1478] Pablo Romero, Franco Robledo, Pablo Rodríguez-Bocca, and Claudia Rostagnol. Lyapunov stability and performance of user-assisted Video-on-Demand services. Computer Networks (Amsterdam, Netherlands: 1999), 79 (??):203–215, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500002X.

Bostanipour:2015:EEN

[1479] Behnaz Bostanipour and Benoît Garbinato. Effective and efficient neighbor detection for proximitybased mobile applications. Computer Networks (Amsterdam, Netherlands: 1999), 79(??):216–235, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614004678.

Daldoul:2015:IDU

[1480] Yousri Daldoul, Djamal-Eddine Meddour, Toufik Ahmed, and Raouf Boutaba. Impact of device unavailability on the reliability of multicast transport in IEEE 802.11 networks. Computer Networks (Amsterdam, Netherlands: 1999), 79(??):

236–246, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000195.

Pachon:2015:FMT

[1481] Álvaro Pachón, Andrés Navarro, and Ubaldo García-Palomares. A flexible mid-term frequency domain scheduler for resource allocation in Het-Nets based on the SINR requested by users. Computer Networks (Amsterdam, Netherlands: 1999), 79(??): 247–262, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000079.

Donohoe:2015:CAS

[1482] Michael Donohoe, Brendan Jennings, and Sasitharan Balasubramaniam. Context-awareness and the smart grid: Requirements and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 79(??): 263–282, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000109.

Wang:2015:DEH

[1483] Ting Wang, Zhiyang Su, Yu Xia, Jogesh Muppala, and Mounir Hamdi. Designing efficient high performance server-centric data center network architecture. Computer Networks (Amsterdam, Netherlands: 1999), 79(??): 283–296, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000092.

Wang:2015:EED

[1484] You-Chiun Wang and Chien-An Chuang. Efficient eNB deployment strategy for heterogeneous cells in 4G LTE systems. Computer Networks (Amsterdam, Netherlands: 1999), 79 (??):297–312, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000249.

Soto:2015:FRS

[1485] Julio Soto and Michele Nogueira. A framework for resilient and secure spectrum sensing on cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 79 (??):313–322, March 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000225. See

corrigendum [2127].

Ekici:2015:JOR

[1486] Fatma Ekici and Didem Gözüpek. Joint overlay routing and relay assignment for green networks. Computer Networks (Amsterdam, Netherlands: 1999), 79(??):323–344, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000080.

Anonymous:2015:EBd

[1487] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 79(??):ifc, March 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128615000523.

Siddiqui:2015:SRE

[1488] M. S. Siddiqui, D. Montero, R. Serral-Gracià, X. Masip-Bruin, and M. Yannuzzi. A survey on the recent efforts of the Internet Standardization Body for securing inter-domain routing. Computer Networks (Amsterdam, Netherlands: 1999), 80 (??):1–26, April 7, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000286.

Martorell:2015:NSI

[1489] Gabriel Martorell, Guillem Femenias, and Felip Riera-Palou. Non-saturated IEEE 802.11 networks. A hierarchical 3D Markov model. Computer Networks (Amsterdam, Netherlands: 1999), 80(??):27–50, April 7, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000419.

Karami:2015:ABC

[1490] Amin Karami and Manel Guerrero-Zapata. An ANFIS-based cache replacement method for mitigating cache pollution attacks in Named Data Networking. Computer Networks (Amsterdam, Netherlands: 1999), 80 (??):51–65, April 7, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000316.

Pastrana:2015:DFO

[1491] Sergio Pastrana, Juan E. Tapiador, Agustin Orfila, and Pedro Peris-Lopez. DEFIDNET: a framework for optimal allocation of cyberdefenses in Intrusion Detection Networks. Computer Networks (Amsterdam, Netherlands: 1999), 80(??):66–88, April 7, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000237.

Madhja:2015:DWP

[1492] Adelina Madhja, Sotiris Nikoletseas, and Theofanis P. Raptis. Distributed wireless power transfer in sensor networks with multiple Mobile Chargers. Computer Networks (Amsterdam, Netherlands: 1999), 80(??): 89–108, April 7, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000298.

Zhao:2015:JVP

[1493] Yangming Zhao, Yifan Huang, Kai Chen, Minlan Yu, Sheng Wang, and DongSheng Li. Joint VM placement and topology optimization for traffic scalability in dynamic datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 80 (??):109–123, April 7, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861400468X.

Sobe:2015:SGA

[1494] Anita Sobe, Wilfried Elmenreich, Tibor Szkaliczki, and Laszlo Böszörmenyi.

SEAHORSE: Generalizing an artificial hormone system algorithm to a middleware for search and delivery of information units. Computer Networks (Amsterdam, Netherlands: 1999), 80 (??):124–142, April 7, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000262.

Azari:2015:PAA

[1495] Amin Azari, Jalil S. Harsini, and Farshad Lahouti. Performance analysis of ad-hoc routing in heterogeneous clustered multi-hop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 80(??): 143–154, April 7, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000274.

Yi:2015:UDR

[1496] Leng-Gan Yi and Yi-Min Lu. Utilitydriven relay for hybrid access femtocells based on cognitive radio spectrum auction. Computer Networks (Amsterdam, Netherlands: 1999), 80 (??):155–166, April 7, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000250.

Anonymous:2015:EBe

[1497] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 80(??):ifc, April 7, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000845.

Jin:2015:CBC

[1498] Xin Jin and Yu-Kwong Kwok. Coercion builds cooperation in dynamic and heterogeneous P2P live streaming networks. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):1–18, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000456.

Marinho:2015:CHS

[1499] José Marinho and Edmundo Monteiro. CORHYS: Hybrid signaling for opportunistic distributed cognitive radio. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):19–42, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000304.

Rezende:2015:SUR

[1500] Cristiano Rezende, Azzedine Boukerche, Mohammed Almulla, and Antonio A. F. Loureiro. The selective use of redundancy for video streaming over Vehicular Ad Hoc Networks. Computer Networks (Amsterdam, Netherlands: 1999), 81(??):43–62, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128614004642.

Malanchini:2015:SCS

[1501] Ilaria Malanchini, Steven Weber, and Matteo Cesana. Stochastic characterization of the spectrum sharing game in ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):63–78, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000213.

Farhady:2015:SDN

[1502] Hamid Farhady, HyunYong Lee, and Akihiro Nakao. Software-Defined Networking: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):79–95, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000614.

Bouten:2015:QDN

[1503] Niels Bouten, Ricardo de O. Schmidt, Jeroen Famaey, Steven Latré, Aiko Pras, and Filip De Turck. QoE-driven in-network optimization for Adaptive Video Streaming based on packet sampling measurements. Computer Networks (Amsterdam, Netherlands: 1999), 81(??):96–115, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000468.

Park:2015:LSM

[1504] Soochang Park, Seung-Woo Hong, Euisin Lee, Sang-Ha Kim, and Noel Crespi. Large-scale mobile phenomena monitoring with energy-efficiency in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 81(??):116–135, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000407.

Aldabbagh:2015:QAT

[1505] Ghadah Aldabbagh, Sheikh Tahir Bakhsh, Nadine Akkari, Sabeen Tahir, Haleh Tabrizi, and John Cioffi. QoS-Aware Tethering in a Heterogeneous Wireless Network using LTE and TV White Spaces. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):136–146, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000444.

Lutu:2015:AEI

[1506] Andra Lutu, Marcelo Bagnulo, Cristel Pelsser, Kenjiro Cho, and Rade Stanojevic. An analysis of the economic impact of strategic deaggregation. Computer Networks (Amsterdam, Netherlands: 1999), 81(??): 147–163, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500047X.

Huang:2015:MDR

[1507] Zhen Huang, Jinbang Chen, Yisong Lin, Pengfei You, and Yuxing Peng. Minimizing data redundancy for high reliable cloud storage systems. Computer Networks (Amsterdam, Netherlands: 1999), 81(??): 164–177, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000602.

Kardes:2015:GBI

[1508] Hakan Kardes, Mehmet Hadi Gunes, and Kamil Sarac. Graph Based Induction of unresponsive routers in Internet topologies. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):178–200, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000596.

Socievole:2015:MSM

[1509] A. Socievole, E. Yoneki, F. De Rango, and J. Crowcroft. ML–SOR: Message routing using multi-layer social networks in opportunistic communications. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):201–219, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000638.

Anagnostopoulos:2015:TOU

[1510] Christos Anagnostopoulos, Stathes Hadjiefthymiades, and Kostas Kolomvatsos. Time-optimized user grouping in Location Based Services. Computer Networks (Amsterdam, Netherlands: 1999), 81(??):220–244, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000651.

Yang:2015:DBC

[1511] Li Yang, Jianfeng Ma, Wenjing Lou, and Qi Jiang. A delegation based cross trusted domain direct anonymous attestation scheme. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):245–257, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000717.

Sarma:2015:OPA

[1512] Siddhartha Sarma and Joy Kuri. Optimal power allocation for protective jamming in wireless networks: a flow based model. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):258–271, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500050X.

Detti:2015:MPP

[1513] Andrea Detti, Bruno Ricci, and Nicola Blefari-Melazzi. Mobile peer-to-peer video streaming over informationcentric networks. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):272–288, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000663.

Yin:2015:RCM

[1514] Wei Yin, Peizhao Hu, and Jadwiga Indulska. Rate control in the mac80211 framework: Overview, evaluation and improvements. Computer Networks (Amsterdam, Netherlands: 1999), 81 (??):289–307, April 22, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000675.

Wang:2015:DAP

[1515] Bing Wang, Yao Zheng, Wenjing Lou, and Y. Thomas Hou. DDoS attack protection in the era of cloud computing and Software-Defined Networking. Computer Networks (Amsterdam, Netherlands: 1999), 81(??): 308–319, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000742.

Hossfeld:2015:IQO

[1516] Tobias Hoßfeld, Michael Seufert, Christian Sieber, Thomas Zinner, and Phuoc Tran-Gia. Identifying QoE optimal adaptation of HTTP adaptive streaming based on subjective studies. Computer Networks (Amsterdam, Netherlands: 1999), 81(??): 320–332, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000626.

Anonymous:2015:EBf

[1517] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 81(??):ifc, April 22, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500105X.

Tavernier:2015:E

[1518] Wouter Tavernier, Deborah Frincke, Achim Autenrieth, and Didier Colle. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 82 (??):1–3, May 8, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001206.

Alashaikh:2015:SCI

[1519] Abdulaziz Alashaikh, Teresa Gomes, and David Tipper. The Spine concept for improving network availability. Computer Networks (Amsterdam, Netherlands: 1999), 82 (??):4–19, May 8, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000687.

Sundarrajan:2015:FRS

[1520] Aditya Sundarrajan and Srinivasan Ramasubramanian. Fast reroute from single link and single node failures for IP multicast. Computer Networks (Amsterdam, Netherlands: 1999), 82 (??):20–33, May 8, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000481.

Gardner:2015:GMT

[1521] M. Todd Gardner, Rebecca May, Cory Beard, and Deep Medhi. A Geographic Multi-Topology Routing approach and its benefits during large-scale geographically correlated failures. Computer Networks (Amsterdam, Netherlands: 1999), 82(??):34–49, May 8, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000729.

Cheng:2015:AGD

[1522] Yufei Cheng, M. Todd Gardner, Junyan Li, Rebecca May, Deep Medhi, and James P. G. Sterbenz. Analysing GeoPath diversity and improving routing performance in optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 82 (??):50–67, May 8, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000699.

Babarczi:2015:IRU

[1523] Péter Babarczi, Alija Pasić, János Tapolcai, Felicián Németh, and Bence Ladóczki. Instantaneous recovery of unicast connections in transport networks: Routing versus coding. Computer Networks (Amsterdam, Netherlands: 1999), 82(??): 68–80, May 8, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000493.

Moon:2015:MJS

[1524] Yong-Hyuk Moon and Chan-Hyun Youn. Multihybrid job scheduling for fault-tolerant distributed computing in policy-constrained resource networks. Computer Networks (Amsterdam, Netherlands: 1999), 82 (??):81–95, May 8, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500078X.

Fortes:2015:CIO

[1525] Sergio Fortes, Raquel Barco, Alejandro Aguilar-García, and Pablo Muñoz. Contextualized indicators for online failure diagnosis in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 82 (??):96–113, May 8, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000791.

Friginal:2015:RRE

[1526] Jesús Friginal, David de Andrés, Juan-Carlos Ruiz, and Miquel Martínez. REFRAHN: a Resilience Evaluation Framework for Ad Hoc Routing Protocols. Computer Networks (Amsterdam, Netherlands: 1999), 82(??): 114–134, May 8, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000808.

Siddiqui:2015:SRD

[1527] M. S. Siddiqui, D. Montero, R. Serral-Gracià, and M. Yannuzzi. Selfreliant detection of route leaks in interdomain routing. Computer Networks (Amsterdam, Netherlands: 1999), 82 (??):135–155, May 8, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000778.

Sahhaf:2015:EVR

[1528] Sahel Sahhaf, Wouter Tavernier, Didier Colle, Mario Pickavet, and Piet Demeester. Experimental validation of resilient tree-based greedy geometric routing. Computer Networks (Amsterdam, Netherlands: 1999), 82 (??):156–171, May 8, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000705.

Anonymous:2015:EBg

[1529] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 82(??):ifc, May 8, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500122X.

Lee:2015:CDI

[1530] Munyoung Lee, Junghwan Song, Kideok Cho, Sangheon Pack, Ted ‘Taekyoung’ Kwon, Jussi Kangasharju, and Yanghee Choi. Content discovery for information-centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 83(??): 1–14, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128614003612.

Vallati:2015:EDW

[1531] Carlo Vallati and Enzo Mingozzi. Efficient design of wireless mesh networks with robust dynamic frequency selection capability. Computer Networks (Amsterdam, Netherlands: 1999), 83 (??):15–29, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000201.

Luo:2015:AMP

[1532] Min Luo, Yulong Zeng, Jianfei Li, and Wu Chou. An adaptive multipath computation framework for centrally controlled networks. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):30–44, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000432.

Dudnikova:2015:MBC

[1533] Anna Dudnikova, Daniela Panno, and Antonio Mastrosimone. Measurementbased coverage function for green femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 83 (??):45–58, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000730.

Xu:2015:OCM

[1534] Zichuan Xu and Weifa Liang. Operational cost minimization of distributed data centers through the provision of fair request rate allocations while meeting different user SLAs. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):59–75, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000766.

Temel:2015:LLO

[1535] Samil Temel and Ilker Bekmezci. LOD-MAC: Location Oriented Directional MAC protocol for FANETs. Computer Networks (Amsterdam, Netherlands:

1999), 83(??):76–84, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500081X.

Liu:2015:EFC

[1536] Yaoqing Liu, Vince Lehman, and Lan Wang. Efficient FIB caching using minimal non-overlapping prefixes. Computer Networks (Amsterdam, Netherlands: 1999), 83(??): 85–99, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000894.

Zheng:2015:CDM

[1537] Xiaoying Zheng, Chunglae Cho, and Ye Xia. Content distribution by multiple multicast trees and intersession cooperation: Optimal algorithms and approximations. Computer Networks (Amsterdam, Netherlands: 1999), 83 (??):100–117, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000900.

Cano:2015:EVP

[1538] C. Cano and D. Malone. On efficiency and validity of previous Homeplug MAC performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):118–135, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000912.

DellAmico:2015:ARM

[1539] Matteo Dell’Amico, Pietro Michiardi, Laszlo Toka, and Pasquale Cataldi. Adaptive redundancy management for durable P2P backup. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):136–148, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000924.

Sanchez-Carmona:2015:EGB

[1540] Adrián Sánchez-Carmona, Sergi Robles, and Carlos Borrego. Endeavouring to be in the good books. Awarding DTN network use for acknowledging the reception of bundles. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):149–166, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000936.

Ali:2015:SSN

[1541] Farha N. Ali, Yvon Feaster, Jiannan Zhai, and Jason O. Hallstrom. The smart surface network: a bus-based approach to dense sensing. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):167–183, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500095X.

Zhang:2015:MCT

[1542] Sheng Zhang, Jie Wu, Zhuzhong Qian, and Sanglu Lu. MobiCache: Cellular traffic offloading leveraging cooperative caching in mobile social networks. Computer Networks (Amsterdam, Netherlands: 1999), 83(??): 184–198, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000973.

Wang:2015:BEO

[1543] Dexiang Wang. Bandwidth-efficiency-oriented topology optimization for integrated switching systems based on circulant graphs. Computer Networks (Amsterdam, Netherlands: 1999), 83 (??):199–216, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000985.

Lakshmi:2015:PPP

[1544] L. Rajya Lakshmi, Vinay J. Ribeiro, and B. N. Jain. PRIME: a partial path establishment based handover management technique for QoS support in WiMAX based wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 83 (??):217–234, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000997.

Bianco:2015:STF

[1545] Andrea Bianco, Davide Cuda, and Jorge M. Finochietto. Short-term fairness in slotted WDM rings. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):235–248, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001000.

Zubeldia:2015:NSP

[1546] Martín Zubeldía, Andrés Ferragut, and Fernando Paganini. Neighbor selection for proportional fairness in P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 83 (??):249–264, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001036.

Avrachenkov:2015:CND

[1547] Konstantin Avrachenkov, Jocelyne Elias, Fabio Martignon, Giovanni Neglia, and Leon Petrosyan. Cooperative network design: a Nash bargaining solution approach. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):265–279, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001127.

Longo:2015:DMS

[1548] Francesco Longo, Salvatore Distefano, Dario Bruneo, and Marco Scarpa. Dependability modeling of Software Defined Networking. Computer Networks (Amsterdam, Netherlands: 1999), 83 (??):280–296, June 4, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001139.

Ma:2015:SET

[1549] Xingkong Ma, Yijie Wang, Xiaoqiang Pei, and Fangliang Xu. Scalable and elastic total order in content-based publish/subscribe systems. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):297–314, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001140.

Ding:2015:DRA

[1550] Lei Ding, Tommaso Melodia, Stella N.

Batalama, and John D. Matyjas. Distributed resource allocation in cognitive and cooperative ad hoc networks through joint routing, relay selection and spectrum allocation. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):315–331, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000754.

Deek:2015:PFM

[1551] Lara Deek, Eduard Garcia-Villegas, Elizabeth Belding, Sung-Ju Lee, and Kevin Almeroth. A practical framework for 802.11 MIMO rate adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 83(??): 332–348, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001024.

Chen:2015:DTN

[1552] Yuanzhu Chen, Xu Liu, Jiafen Liu, Walter Taylor, and Jason H. Moore. Delay-tolerant networks and network coding: Comparative studies on simulated and real-device experiments. Computer Networks (Amsterdam, Netherlands: 1999), 83(??): 349–362, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001152.

Nadeem:2015:ASS

[1553] Adnan Nadeem, Muhammad Azhar Hussain, Obaidullah Owais, Abdul Salam, Sarwat Iqbal, and Kamran Ahsan. Application specific study, analysis and classification of body area wireless sensor network applications. Computer Networks (Amsterdam, Netherlands: 1999), 83(??): 363–380, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000821.

Aceto:2015:ICD

[1554] Giuseppe Aceto and Antonio Pescapé. Internet Censorship detection: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 83(??): 381–421, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615000948.

Talooki:2015:SCC

[1555] Vahid Nazari Talooki, Riccardo Bassoli, Daniel E. Lucani, Jonathan Rodriguez, Frank H. P. Fitzek, Hugo Marques, and Rahim Tafazolli. Security concerns and countermeasures in network coding based communication systems: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):422–445, June 4, 2015.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615000961.

Anonymous:2015:EBh

[1556] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 83(??):ifc, June 4, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001395.

Jiang:2015:EDS

[1557] Dingde Jiang, Yuanting Wang, Chunping Yao, and Yang Han. An effective dynamic spectrum access algorithm for multi-hop cognitive wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 84 (??):1–16, June 19, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001164.

Hassan:2015:CST

[1558] Md. Rakib Hassan, Gour Karmakar, Joarder Kamruzzaman, and Bala Srinivasan. A comprehensive spectrum trading scheme based on market competition, reputation and buyer specific requirements. Computer Networks (Amsterdam, Netherlands: 1999), 84 (??):17–31, June 19, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001176.

Saha:2015:ECA

[1559] Sumanta Saha, Andrey Lukyanenko, and Antti Ylä-Jääski. Efficient cache availability management in Information-Centric Networks. Computer Networks (Amsterdam, Netherlands: 1999), 84(??):32–45, June 19, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128615001188.

Pietro:2015:EES

[1560] Roberto Di Pietro and Gabriele Oligeri. ESC: an efficient, scalable, and cryptoless solution to secure wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 84(??): 46–63, June 19, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500119X.

Ran:2015:CNP

[1561] Chen Ran, Shaowei Wang, and Chonggang Wang. Cellular networks planning: a workload balancing perspective. Computer Networks (Amsterdam, Netherlands: 1999), 84(??): 64–75, June 19, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001346.

Hoteit:2015:MDT

[1562] Sahar Hoteit, Stefano Secci, Guy Pujolle, Adam Wolisz, Cezary Ziemlicki, and Zbigniew Smoreda. Mobile data traffic offloading over Passpoint hotspots. Computer Networks (Amsterdam, Netherlands: 1999), 84 (??):76–93, June 19, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500136X.

Anonymous:2015:EBi

[1563] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 84(??):ifc, June 19, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001504.

Akyildiz:2015:SSD

[1564] Ian F. Akyildiz, Pu Wang, and Shih-Chun Lin. SoftAir: a software defined networking architecture for 5G wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 85 (??):1–18, July 5, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001632.

Yoon:2015:ESF

[1565] Changhoon Yoon, Taejune Park, Seungsoo Lee, Heedo Kang, Seungwon Shin, and Zonghua Zhang. Enabling security functions with SDN: a feasibility study. Computer Networks (Amsterdam, Netherlands: 1999), 85 (??):19–35, July 5, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001619.

Otoshi:2015:TPD

[1566] Tatsuya Otoshi, Yuichi Ohsita, Masayuki Murata, Yousuke Takahashi, Keisuke Ishibashi, and Kohei Shiomoto. Traffic prediction for dynamic traffic engineering. Computer Networks (Amsterdam, Netherlands: 1999), 85 (??):36–50, July 5, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001565.

He:2015:SRF

[1567] Jun He and Wei Song. Smart routing: Fine-grained stall management of video streams in mobile core networks. Computer Networks (Amsterdam, Netherlands: 1999), 85 (??):51–62, July 5, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001553.

Ali:2015:SFL

[1568] Mohammed L. Ali, Pin-Han Ho, and János Tapolcai. SRLG fault localization using nested m-trails. Computer Networks (Amsterdam, Netherlands: 1999), 85(??):63–79, July 5, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001371.

Anonymous:2015:EBj

[1569] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 85(??):ifc, July 5, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001784.

Marynowski:2015:MTF

[1570] João Eugenio Marynowski, Altair Olivo Santin, and Andrey Ricardo Pimentel. Method for testing the fault tolerance of MapReduce frameworks. Computer Networks (Amsterdam, Netherlands: 1999), 86(??):1–13, July 5, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001358.

Gunes:2015:BPF

[1571] Mehmet Hadi Gunes, Murat Yuksel, and Hayreddin Ceker. A blind processing framework to facilitate openness in smart grid communications. Computer Networks (Amsterdam, Netherlands: 1999), 86(??):14–26, July 5, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001607.

Wu:2015:GTG

[1572] Xinyun Wu, Tao Ye, Qi Guo, and Zhipeng Lü. GRASP for traffic grooming and routing with simple path constraints in WDM mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 86 (??):27–39, July 5, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001590.

Garcia-Villegas:2015:NCJ

[1573] Eduard Garcia-Villegas, Muhammad Shahwaiz Afaqui, and Elena Lopez-Aguilera. A novel cheater and jammer detection scheme for IEEE 802.11-based wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 86(??):40–56, July 5, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001589.

Yao:2015:DWS

[1574] Yingbiao Yao and Nanlan Jiang. Distributed wireless sensor network localization based on weighted search. Computer Networks (Amsterdam, Netherlands: 1999), 86(??):57–75, July 5, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128615001577.

Anonymous:2015:EBk

[1575] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 86(??):ifc, July 5, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001875.

Liao:2015:CGA

[1576] Jianxin Liao, Ziteng Cui, Jingyu Wang, Tonghong Li, Qi Qi, and Jing Wang. A coalitional game approach on improving interactions in multiple overlay environments. Computer Networks (Amsterdam, Netherlands: 1999), 87(??):1–15, July 20, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001620.

Cerroni:2015:CLR

[1577] Walter Cerroni, Molka Gharbaoui, Barbara Martini, Aldo Campi, Piero Castoldi, and Franco Callegati. Crosslayer resource orchestration for cloud service delivery: a seamless SDN approach. Computer Networks (Amsterdam, Netherlands: 1999), 87(??): 16–32, July 20, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001644.

Fu:2015:SCP

[1578] Shu Fu, Bin Wu, Xiaohong Jiang, Achille Pattavina, Hong Wen, and Hongfang Yu. Switch cost and packet delay tradeoff in data center networks with switch reconfiguration overhead. Computer Networks (Amsterdam, Netherlands: 1999), 87 (??):33–43, July 20, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500170X.

Sanchez-Casado:2015:MDF

[1579] Leovigildo Sánchez-Casado, Gabriel Maciá-Fernández, Pedro García-Teodoro, and Roberto Magán-Carrión. A model of data forwarding in MANETs for lightweight detection of malicious packet dropping. Computer Networks (Amsterdam, Netherlands: 1999), 87 (??):44–58, July 20, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001723.

Borrego:2015:MCB

[1580] Carlos Borrego, Sergi Robles, Angela Fabregues, and Adrián Sánchez-Carmona. A mobile code bundle extension for application-defined routing in delay and disruption tolerant networking. Computer Networks (Amsterdam, Netherlands: 1999), 87 (??):59–77, July 20, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001942.

Yao:2015:CAB

[1581] Lin Yao, Jing Deng, Jie Wang, and Guowei Wu. A-CACHE: an anchorbased public key caching scheme in large wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 87(??):78–88, July 20, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001851.

Anonymous:2015:EBl

[1582] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 87(??):ifc, July 20, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002078.

Zhang:2015:CTR

[1583] Ertong Zhang and Lisong Xu. Capacity and token rate estimation for networks with token bucket shapers. Computer Networks (Amsterdam, Netherlands: 1999), 88(??):1–11, September 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001735.

Senouci:2015:WDF

[1584] Mustapha Reda Senouci, Abdelhamid Mellouk, Latifa Oukhellou, and Amar Aissani. WSNs deployment framework based on the theory of belief functions. Computer Networks (Amsterdam, Netherlands: 1999), 88 (??):12–26, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001747.

Vien:2015:CLT

[1585] Quoc-Tuan Vien, Wanqing Tu, Huan X.

Nguyen, and Ramona Trestian. Crosslayer topology design for network coding based wireless multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 27–39, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500198X.

Vaezi:2015:APM

[1586] Kaveh Vaezi and Nail Akar. Analytical performance modeling of elastic optical links with aligned spectrum allocation. Computer Networks (Amsterdam, Netherlands: 1999), 88 (??):40–50, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001978.

Girolami:2015:SDM

[1587] Michele Girolami, Stefano Chessa, and Antonio Caruso. On service discovery in mobile social networks: Survey and perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 88 (??):51–71, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001991.

Yang:2015:NTP

[1588] Xinyu Yang, Xuebin Ren, Shusen Yang, and Julie McCann. A novel temporal perturbation based privacypreserving scheme for real-time monitoring systems. Computer Networks (Amsterdam, Netherlands: 1999), 88 (??):72–88, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002005.

Lira:2015:AAD

[1589] Victor Lira, Eduardo Tavares, and Paulo Maciel. An automated approach to dependability evaluation of virtual networks. Computer Networks (Amsterdam, Netherlands: 1999), 88 (??):89–102, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001760.

Fux:2015:RSU

[1590] Vladimir Fux, Patrick Maillé, and Matteo Cesana. Road-side units operators in competition: a game-theoretical approach. Computer Networks (Amsterdam, Netherlands: 1999), 88(??):

103–120, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002017.

Hou:2015:RMC

[1591] Weigang Hou, Lei Guo, Yejun Liu, Cunqian Yu, and Yue Zong. Resource management and control in converged optical data center networks: Survey and enabling technologies. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 121–135, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001711.

Wang:2015:OHD

[1592] Wei Wang, Lei Chen, and Qian Zhang.

Outsourcing high-dimensional healthcare data to cloud with personalized privacy preservation. Computer Networks (Amsterdam, Netherlands: 1999), 88(??):136–148, September 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128615002157.

Tsolkas:2015:EDD

[1593] Dimitris Tsolkas, Nikos Passas, and Lazaros Merakos. Enabling device discovery transmissions in LTE networks with fractional frequency reuse. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 149–160, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002030.

Premarathne:2015:LDD

[1594] Uthpala Subodhani Premarathne, Ibrahim Khalil, and Mohammed Atiquzzaman. Location-dependent disclosure risk based decision support framework for persistent authentication in pervasive computing applications. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 161–177, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615001954.

Adarbah:2015:INI

[1595] Haitham Y. Adarbah, Shakeel Ahmad, and Alistair Duffy. Impact of noise and interference on probabilistic broadcast schemes in mobile ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 178–186, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002145.

Papadimitriou:2015:NLP

[1596] Georgios Papadimitriou, Nikolaos Pappas, Apostolos Traganitis, and Vangelis Angelakis. Network-level performance evaluation of a two-relay cooperative random access wireless system. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 187–201, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002054.

Facchi:2015:ELR

[1597] Nicolò Facchi, Francesco Gringoli, Fabio Ricciato, and Andrea Toma. Emitter localisation from reception timestamps in asynchronous networks. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 202–217, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001966.

Makki:2015:SWP

[1598] Ahmed Makki, Abubakr Siddig, Mohamed Saad, and Chris Bleakley. Survey of WiFi positioning using timebased techniques. Computer Networks (Amsterdam, Netherlands: 1999), 88 (??):218–233, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002169.

Ahmed:2015:SEH

[1599] Imran Ahmed, M. Majid Butt, Constantinos Psomas, Amr Mohamed, Ioannis Krikidis, and Mohsen Guizani. Survey on energy harvesting wireless communications: Challenges and opportunities for radio resource allocation. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 234–248, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002029.

Dalvandi:2015:PER

[1600] Aissan Dalvandi, Mohan Gurusamy, and Kee Chaing Chua. Power-efficient resource-guaranteed VM placement and routing for time-aware data center applications. Computer Networks (Amsterdam, Netherlands: 1999), 88 (??):249–268, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002182.

Coutinho:2015:DDR

[1601] Nuno Coutinho, Ricardo Matos, Carlos Marques, André Reis, Susana Sargento, Jacob Chakareski, and Andreas Kassler. Dynamic dual-reinforcement-learning routing strategies for quality of experience-aware wireless mesh networking. Computer Networks (Amsterdam, Netherlands: 1999), 88(??): 269–285, September 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002170.

Anonymous:2015:EBm

[1602] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):ifc, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002601.

Kim:2015:AHS

[1603] Ji-Su Kim, Jin-Ki Kim, and Jae-Hyun Kim. Advanced handover scheme considering downlink and uplink service traffic in asymmetric channel. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):1–13, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002212.

Kos:2015:USS

[1604] Jernej Kos, Mahdi Aiash, Jonathan Loo, and Denis Trçek. U-Sphere: Strengthening scalable flat-name routing for decentralized networks. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):14–31, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002297.

Rossi:2015:IBS

[1605] Francesco Rossi and Giovanni Schmid.

Identity-based secure group communications using pairings. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):32–43, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002303.

Fu:2015:MRA

[1606] Bo Fu and Yang Xiao. A multiresolution accountable logging and its applications. Computer Networks (Amsterdam, Netherlands: 1999), 89(??): 44–58, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128615002042.

Oliveira:2015:ORE

[1607] Rodrigo R. Oliveira, Daniel S. Marcon, Leonardo R. Bays, Miguel C.

Neves, Luciano P. Gaspary, Deep Medhi, and Marinho P. Barcellos. Opportunistic Resilience Embedding (ORE): Toward cost-efficient resilient virtual networks. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):59–77, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002339.

Rojas-Cessa:2015:HIL

[1608] Roberto Rojas-Cessa, Taweesak Kijkanjanarat, Wara Wangchai, Krutika Patil, and Narathip Thirapittayatakul. Helix: IP lookup scheme based on helicoidal properties of binary trees. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):78–89, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002418.

Hiland-Jrgensen:2015:GBW

[1609] Toke Høiland-Jørgensen, Per Hurtig, and Anna Brunstrom. The good, the bad and the WiFi: Modern AQMs in a residential setting. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):90–106, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002479.

Accettura:2015:CRA

[1610] Nicola Accettura, Giovanni Neglia, and Luigi Alfredo Grieco. The Capture– Recapture approach for population estimation in computer networks. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):107–122, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500242X.

Hossfeld:2015:SIC

[1611] Tobias Hoßfeld, Phuoc Tran-Gia, and Maja Vukovic. Special issue on crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 90 (??):1–4, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001759.

Demartini:2015:HHM

[1612] Gianluca Demartini. Hybrid humanmachine information systems: Challenges and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 90(??):5–13, October 29, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002194.

Scekic:2015:PDS

[1613] Ognjen Scekic, Hong-Linh Truong, and Schahram Dustdar. PRINGL — a domain-specific language for incentive management in crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 90(??): 14–33, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002200.

Goncalves:2015:MPI

[1614] Jorge Goncalves, Simo Hosio, Jakob Rogstadius, Evangelos Karapanos, and Vassilis Kostakos. Motivating participation and improving quality of contribution in ubiquitous crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 90(??): 34–48, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002224.

Mousa:2015:TMR

[1615] Hayam Mousa, Sonia Ben Mokhtar, Omar Hasan, Osama Younes, Mohiy Hadhoud, and Lionel Brunie. Trust management and reputation systems in mobile participatory sensing applications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 90 (??):49–73, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002340.

Sasao:2015:CWA

[1616] Tomoyo Sasao, Shin’ichi Konomi, Masatoshi Arikawa, and Hideyuki Fujita. Context Weaver: Awareness and feedback in networked mobile crowdsourcing tools. Computer Networks (Amsterdam, Netherlands: 1999), 90 (??):74–84, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002273.

Hirth:2015:CNM

[1617] Matthias Hirth, Tobias Hoßfeld, Marco Mellia, Christian Schwartz, and Frank Lehrieder. Crowdsourced network measurements: Benefits and best practices. Computer Networks (Amsterdam, Netherlands: 1999), 90(??): 85–98, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002236.

Volk:2015:CVL

[1618] Thomas Volk, Christian Keimel, Michael Moosmeier, and Klaus Diepold. Crowdsourcing vs. laboratory experiments — QoE evaluation of binaural playback in a teleconference scenario. Computer Networks (Amsterdam, Netherlands: 1999), 90(??): 99–109, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002261.

Christoforaki:2015:SSR

[1619] Maria Christoforaki and Panagiotis G.

Ipeirotis. A system for scalable and reliable technical-skill testing in online labor markets. Computer Networks (Amsterdam, Netherlands: 1999), 90 (??):110–120, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500225X.

Hosseini:2015:WCW

[1620] Mahmood Hosseini, Jack Moore, Malik Almaliki, Alimohammad Shahri, Keith Phalp, and Raian Ali. Wisdom of the Crowd within enterprises: Practices and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 90 (??):121–132, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002248.

Oosterman:2015:IKE

[1621] Jasper Oosterman, Jie Yang, Alessandro Bozzon, Lora Aroyo, and Geert-Jan Houben. On the impact of knowledge extraction and aggregation on crowdsourced annotation of visual artworks. Computer Networks (Amsterdam, Netherlands: 1999), 90(??): 133–149, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002315.

Baig:2015:GNC

[1622] Roger Baig, Ramon Roca, Felix Freitag, and Leandro Navarro. guifi.net, a crowdsourced network infrastructure held in common. Computer Networks (Amsterdam, Netherlands: 1999), 90 (??):150–165, October 29, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002327.

Anonymous:2015:EBn

[1623] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 90(??):ifc, October 29, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128615003333.

Kim:2015:ARD

[1624] Kangho Kim, Hwantae Kim, Jongtack Jung, and Hwangnam Kim. AFAR: a robust and delay-constrained communication framework for smart grid applications. Computer Networks (Amsterdam, Netherlands: 1999), 91 (??):1–25, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002509.

Sahraoui:2015:EHB

[1625] Somia Sahraoui and Azeddine Bilami.

Efficient HIP-based approach to ensure lightweight end-to-end security in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 91 (??):26–45, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002558.

Juvonen:2015:OAD

[1626] Antti Juvonen, Tuomo Sipola, and Timo Hämäläinen. Online anomaly detection using dimensionality reduction techniques for HTTP log analysis. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 46–56, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002650.

Khloussy:2015:IMB

[1627] Elissar Khloussy, Xavier Gelabert, and Yuming Jiang. Investigation on MDPbased radio access technology selection in heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 57–67, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002583.

Trotta:2015:CRP

[1628] Angelo Trotta, Marco Di Felice, Luca Bedogni, Luciano Bononi, and Fabio Panzieri. Connectivity recovery in post-disaster scenarios through Cognitive Radio swarms. Computer Networks (Amsterdam, Netherlands: 1999), 91 (??):68–89, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002510.

Fan:2015:OSA

[1629] Bo Fan, Supeng Leng, Kun Yang, and Yan Zhang. Optimal storage allocation on throwboxes in Mobile Social Networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 90–100, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002753.

Chen:2015:UVE

[1630] Yanjiao Chen, Qihong Chen, Fan Zhang, Qian Zhang, Kaishun Wu, Ruochen Huang, and Liang Zhou.

Understanding viewer engagement of video service in Wi-Fi network. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 101–116, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002662.

Zhang:2015:CRG

[1631] Biling Zhang, Yan Chen, Chih-Yu Wang, and K. J. Ray Liu. A Chinese restaurant game for learning and decision making in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 117–134, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002704.

Yin:2015:AGC

[1632] Jiechen Yin, Yuming Mao, Supeng Leng, Yuming Jiang, and Muhammad Asad Khan. Access granularity control of multichannel random access in next-generation wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 135–150, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002686.

Liu:2015:TSA

[1633] Shuhao Liu, Zhiping Cai, Hong Xu, and Ming Xu. Towards security-aware virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):

151–163, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002741.

Wu:2015:TSD

[1634] Qianru Wu, Qixu Liu, Yuqing Zhang, and Guanxing Wen. TrackerDetector: a system to detect thirdparty trackers through machine learning. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 164–173, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002728.

Coras:2015:SLM

[1635] Florin Coras, Jordi Domingo-Pascual, and Albert Cabellos-Aparicio. On the scalability of LISP mappings caches. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 174–183, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002698.

Saleem:2015:SSA

[1636] Yasir Saleem, Kok-Lim Alvin Yau, Hafizal Mohamad, Nordin Ramli, and Mubashir Husain Rehmani. SMART: a SpectruM-Aware ClusteR-based rouTing scheme for distributed cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 91 (??):196–224, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002790.

Chakraborty:2015:DRP

[1637] Sandip Chakraborty and Sukumar Nandi. Data rate, path length and network contention trade-off in IEEE 802.11s mesh networks: a dynamic data rate selection approach. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 225–243, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002868.

Zheng:2015:ONO

[1638] Xiaoying Zheng and Ye Xia. Optimizing network objectives in collaborative content distribution. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):244–261, November 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861500273X.

Liu:2015:CMP

[1639] Kai Liu, Xiaoying Chang, Feng Liu, Xin Wang, and Athanasios V. Vasilakos. A cooperative MAC protocol with rapid relay selection for wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 91 (??):262–282, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002807.

Divakaran:2015:SSL

[1640] Dinil Mon Divakaran, Le Su, Yung Siang Liau, and Vrizlynn L. L. Thing. SLIC: Self-Learning Intelligent Classifier for network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 91 (??):283–297, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002819.

Huang:2015:HLD

[1641] Qun Huang and Patrick P. C. Lee. A hybrid local and distributed sketching design for accurate and scalable heavy key detection in network data streams. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 298–315, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002856.

Chrysos:2015:LSB

[1642] N. Chrysos, F. Neeser, M. Gusat, C. Minkenberg, W. Denzel, C. Basso, M. Rudquist, K. Valk, and B. Vanderpool. Large switches or blocking multi-stage networks? An evaluation of routing strategies for datacenter fabrics. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 316–328, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002893.

Leu:2015:IIP

[1643] Jenq-Shiou Leu, Min-Chieh Yu, and Hung-Jie Tzeng. Improving indoor positioning precision by using received signal strength fingerprint and footprint based on weighted ambient Wi-Fi signals. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 329–340, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002923.

Louati:2015:BFT

[1644] Wassef Louati, Walid Ben-Ameur, and Djamal Zeghlache. A bottleneckfree tree-based name resolution system for Information-Centric Networking. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 341–355, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002844.

Yin:2015:DDI

[1645] Xia Yin, Dan Wu, Zhiliang Wang, Xingang Shi, and Jianping Wu. DIMR: Disjoint Interdomain Multipath Routing. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 356–375, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002881.

Zhu:2015:SNS

[1646] Donghai Zhu, Xinyu Yang, and Wei Yu. SPAIS: a novel Self-checking Pollution Attackers Identification Scheme in network coding-based wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):

376–389, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002832.

Hong:2015:FAP

[1647] Gongbing Hong, James Martin, and James M. Westall. On fairness and application performance of active queue management in broadband cable networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 390–406, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002789.

Cheng:2015:MSR

[1648] Chien-Fu Cheng, Chih-Wei Huang, and Lung-Hao Li. Mobile sensor relocation problem: Finding the optimal (nearest) redundant sensor with low message overhead. Computer Networks (Amsterdam, Netherlands: 1999), 91 (??):407–424, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002820.

Milyeykovski:2015:UCN

[1649] Vitaly Milyeykovski, Michael Segal, and Vladimir Katz. Using central nodes for efficient data collection in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 425–437, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003011.

Giacobbe:2015:TEM

[1650] Maurizio Giacobbe, Antonio Celesti, Maria Fazio, Massimo Villari, and Antonio Puliafito. Towards energy management in cloud federation: a survey in the perspective of future sustainable and cost-saving strategies. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 438–452, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002911.

Vakilinia:2015:MRA

[1651] Shahin Vakilinia, Mustafa Mehmet Ali, and Dongyu Qiu. Modeling of the resource allocation in cloud computing centers. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 453–470, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500290X.

Pham:2015:MTC

[1652] Quoc-Viet Pham, Hoang-Linh To, and Won-Joo Hwang. A multi-timescale cross-layer approach for wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 471–482, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002674.

Tian:2015:OBA

[1653] Chen Tian, Jingdong Sun, Weimin Wu, and Yan Luo. Optimal bandwidth allocation for hybrid Video-on-Demand streaming with a distributed max flow algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 483–494, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003114.

Xia:2015:DDU

[1654] Feng Xia, Qiuyuan Yang, Jie Li, Jiannong Cao, Li Liu, and Ahmedin Mohammed Ahmed. Data dissemination using interest-tree in socially aware networking. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 495–507, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003096.

Ilkhechi:2015:NAV

[1655] Amir Rahimzadeh Ilkhechi, Ibrahim Korpeoglu, and Özgür Ulusoy. Networkaware virtual machine placement in cloud data centers with multiple traffic-intensive components. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):508–527, November 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003023.

Wang:2015:SDC

[1656] Bin Wang, Zhengwei Qi, Ruhui Ma, Haibing Guan, and Athanasios V.

Vasilakos. A survey on data center networking for cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 528–547, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500300X.

Oh:2015:CBP

[1657] Bong-Hwan Oh and Jaiyong Lee.

Constraint-based proactive scheduling for MPTCP in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 548–563, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003047.

Goudos:2015:MOA

[1658] Sotirios K. Goudos, David Plets, Ning Liu, Luc Martens, and Wout Joseph. A multi-objective approach to indoor wireless heterogeneous networks planning based on biogeography-based optimization. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 564–576, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002972.

Zhang:2015:DNC

[1659] Xu Zhang, Ning Wang, Vassilios G.

Vassilakis, and Michael P. Howarth. A distributed in-network caching scheme for P2P-like content chunk delivery. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):

577–592, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002959.

Chen:2015:AFI

[1660] Guo Chen, Youjian Zhao, and Dan Pei. Alleviating flow interference in data center networks through fine-grained switch queue management. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 593–613, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002935.

Kawahara:2015:TAP

[1661] Jun Kawahara, Koji M. Kobayashi, and Tomotaka Maeda. Tight analysis of priority queuing for egress traffic. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 614–624, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003035.

Sou:2015:MAB

[1662] Sok-Ian Sou and Hung-Yang Hsieh. Modeling application-based charging management with traffic detection function in 3GPP. Computer Networks (Amsterdam, Netherlands: 1999), 91 (??):625–637, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002777.

Mangili:2015:ODI

[1663] Michele Mangili, Fabio Martignon, and Antonio Capone. Optimal design of Information Centric Networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 638–653, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003138.

Kumar:2015:TLL

[1664] Satish Kumar, Arnab Sarkar, Santhosh Sriram, and Arijit Sur. A three level LTE downlink scheduling framework for RT VBR traffic. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 654–674, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500287X.

Panayiotou:2015:IAM

[1665] T. Panayiotou, G. Ellinas, N. Antoniades, and A. Hadjiantonis. Impairmentaware multicast session provisioning in metro optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):675–688, November 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861500314X.

Ahmadi:2015:NLH

[1666] Mahmood Ahmadi and Sara Mehdizadeh Khalifani. A new look at hybrid Aloha: an analytical approach. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):

689–699, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002947.

Fadini:2015:SSP

[1667] Waya Fadini, Bijoy Chand Chatterjee, and Eiji Oki. A subcarrier-slot partition scheme with first-last fit spectrum allocation for elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 700–711, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003102.

Xie:2015:SIF

[1668] L. F. Xie, Peter H. J. Chong, Ivan W. H. Ho, and Y. L. Guan. A survey of inter-flow network coding in wireless mesh networks with unicast traffic. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 738–751, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003060.

Rizzi:2015:LCR

[1669] Antonello Rizzi, Alfonso Iacovazzi, Andrea Baiocchi, and Silvia Colabrese. A low complexity real-time Internet traffic flows neuro-fuzzy classifier. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 752–771, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003217.

Song:2015:OSI

[1670] Jee Hun Song, Hee-Tae Roh, and Jang-Won Lee. Opportunistic scheduling and incentive mechanism for OFDMA networks with D2D relaying. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 772–787, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003072.

Cardona:2015:BHA

[1671] J. Camilo Cardona, Pierre François, Bruno Decraene, John Scudder, Adam Simpson, and Keyur Patel. Bringing high availability to BGP: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 91(??): 788–803, November 14, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003151.

Kawahara:2015:OBM

[1672] Jun Kawahara and Koji M. Kobayashi.

Optimal buffer management for 2-frame throughput maximization. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):804–820, November 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003084.

Anonymous:2015:EBo

[1673] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):ifc, November 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128615003692.

Moscholios:2015:CPE

[1674] Ioannis D. Moscholios, Michael D. Logothetis, John S. Vardakas, and Anthony C. Boucouvalas. Congestion probabilities of elastic and adaptive calls in Erlang–Engset multirate loss models under the threshold and bandwidth reservation policies. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1) (??):1–23, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003205.

Nam:2015:SMP

[1675] Changwon Nam and Saewoong Bahk. Δ SNR–MAC: a priority-based multiround contention scheme for MU– MIMO WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1)(??):24–40, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003126.

Megyesi:2015:UBB

[1676] Péter Megyesi, Géza Szabó, and Sándor Molnár. User behavior based traffic emulator: a framework for generating test data for DPI tools. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1)(??):41–54, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003461.

Lin:2015:BTE

[1677] Xiao-Hui Lin, Yu Tan, Yu-Kwong Kwok, Hui Wang, Mingjun Dai, Bin Chen, and Gong Chao Su. Balancing time and energy efficiencies with identification reliability constraint for portable reader in mobile RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1) (??):55–71, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500328X.

Luo:2015:PFT

[1678] Shouxi Luo, Hongfang Yu, and Lemin Li. Practical flow table aggregation in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1)(??):72–88, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003278.

Fichera:2015:OOB

[1679] Silvia Fichera, Laura Galluccio, Salvatore C. Grancagnolo, Giacomo Morabito, and Sergio Palazzo. OP-ERETTA: an OPEnflow-based REmedy to mitigate TCP SYNFLOOD Attacks against web servers. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1)(??):89– 100, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002984.

Su:2015:CCE

[1680] Zhiyang Su, Ting Wang, Yu Xia, and Mounir Hamdi. CeMon: a costeffective flow monitoring system in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1)(??):101– 115, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003291.

Kim:2015:OAT

[1681] Hyunbum Kim and Jorge A. Cobb. Optimization algorithms for transmission range and actor movement in wireless sensor and actor networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1) (??):116–133, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003308.

Weng:2015:RTV

[1682] Yung-Tsung Weng, Chi-Huang Shih, Chun-I Kuo, and Ce-Kuen Shieh. Realtime video streaming using predictionbased forward error correction. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1)(??):134– 147, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500331X.

Moety:2015:OMJ

[1683] Farah Moety, Samer Lahoud, Bernard Cousin, and Kinda Khawam. Optimization models for the joint powerdelay minimization problem in green wireless access networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1)(??):148– 167, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003527.

Habak:2015:BAT

[1684] Karim Habak, Khaled A. Harras, and Moustafa Youssef. Bandwidth aggregation techniques in heterogeneous multi-homed devices: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1) (??):168–188, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002996.

daSilva:2015:RSS

[1685] Anderson Santos da Silva, Paul Smith, Andreas Mauthe, and Alberto Schaeffer-Filho. Resilience support in software-defined networking: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1) (??):189–207, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003229.

Anonymous:2015:EBp

[1686] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 1)(??): ifc, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003953.

Leon-Garcia:2015:SDN

[1687] Alberto Leon-Garcia, Peter Ashwood-Smith, and Yashar Ganjali. Software Defined Networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):209– 210, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003552.

Contreras:2015:OOB

[1688] Luis M. Contreras, Paul Doolan, Håkon Lønsethagen, and Diego R. López. Operational, organizational and business challenges for network operators in the context of SDN and NFV. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2) (??):211–217, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002492.

Rizvi:2015:TCG

[1689] Syed Naveed Rizvi, Daniel Raumer, Florian Wohlfart, and Georg Carle. Towards carrier grade SDNs. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):218– 226, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003497.

Kotronis:2015:RCA

[1690] Vasileios Kotronis, Adrian Gämperli, and Xenofontas Dimitropoulos. Routing centralization across domains via SDN: a model and emulation framework for BGP evolution. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):227– 239, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002480.

Nakamura:2015:CDP

[1691] Ryo Nakamura, Kouji Okada, Yuji Sekiya, and Hiroshi Esaki. A common data plane for multiple overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2) (??):240–250, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003515.

Nguyen:2015:GRA

[1692] Huu Thanh Nguyen, Anh Vu Vu, Duc Lam Nguyen, Van Huynh Nguyen, Manh Nam Tran, Quynh Thu Ngo, Thu-Huong Truong, Tai Hung Nguyen, and Thomas Magedanz. A generalized resource allocation framework in support of multi-layer virtual network embedding based on SDN. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2) (??):251–269, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003783.

Peresini:2015:STO

[1693] Peter Peresíni, Maciej Kuźniar, Marco Canini, Daniele Venzano, Dejan Kostić, and Jennifer Rexford. Systematically testing OpenFlow controller applications. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2) (??):270–286, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002546.

Li:2015:SSA

[1694] Qing Li, Kun Zhao, Yong Jiang, Mingwei Xu, and Shu-Tao Xia. SARD: a Smart Approach of Rule Division for fast flow-level consistent update in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2) (??):287–299, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500345X.

Guo:2015:JRF

[1695] Zehua Guo, Yang Xu, Marco Cello, Junjie Zhang, Zicheng Wang, Mingjian Liu, and H. Jonathan Chao. JumpFlow: Reducing flow table usage in softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):300–315, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128615003503.

Kim:2015:ICM

[1696] Won-Suk Kim, Sang-Hwa Chung, and Jae-Won Moon. Improved content management for informationcentric networking in SDN-based wireless mesh network. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):316– 329, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002534.

Hegde:2015:SFF

[1697] Saumya Hegde, Shashidhar G. Koolagudi, and Swapan Bhattacharya. Scalable and fair forwarding of elephant and mice traffic in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2) (??):330–340, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003242.

Owens:2015:VSD

[1698] Harold Owens II and Arjan Durresi.

Video over Software-Defined Networking (VSDN). Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):341–356, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003199.

Uzakgider:2015:LBA

[1699] Tuba Uzakgider, Cihat Cetinkaya, and Muge Sayit. Learning-based approach for layered adaptive video streaming over SDN. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):357–368, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003473.

Caraguay:2015:FOM

[1700] Ángel Leonardo Valdivieso Caraguay, Jesús Antonio Puente Fernández, and Luis Javier García Villalba. Framework for optimized multimedia routing over software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):369– 379, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003230.

Cerrato:2015:TDV

[1701] Ivano Cerrato, Alex Palesandro, Fulvio Risso, Marc Suñé, Vinicio Vercellone, and Hagen Woesner. Toward dynamic virtualized network services in telecom operator networks. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):380– 395, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003485.

Cheng:2015:ENF

[1702] Guozhen Cheng, Hongchang Chen, Hongchao Hu, Zhiming Wang, and Julong Lan. Enabling network function combination via service chain instantiation. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)

(??):396–407, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003254.

Wang:2015:DFC

[1703] Peng Wang, Julong Lan, Xiaohui Zhang, Yuxiang Hu, and Shuqiao Chen. Dynamic function composition for network service chain: Model and optimization. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??):408–418, December 9, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003266.

Anonymous:2015:EBq

[1704] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 92 (part 2)(??): ifc, December 9, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004235.

Baccarelli:2015:MEB

[1705] Enzo Baccarelli, Danilo Amendola, and Nicola Cordeschi. Minimumenergy bandwidth management for QoS live migration of virtual machines. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1) (??):1–22, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003655.

Gonzalez-Horta:2015:MMS

[1706] Francisco A. Gonzalez-Horta, Pedro Mejia-Alvarez, and Eldamira Buenfil-Alpuche. Multipurpose mobility services for the Future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):23– 40, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003539.

Madani:2015:MBM

[1707] Shaahin Madani and Ibrahim Khalil. Multi-binomial mixes: a proposal for secure and efficient anonymous communication. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):41–53, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003679.

Casoni:2015:QQR

[1708] Maurizio Casoni, Carlo Augusto Grazia, Martin Klapez, and Natale Patriciello. QRM: a queue rate management for fairness and TCP flooding protection in mission-critical networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1) (??):54–65, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003837.

Akyildiz:2015:WSD

[1709] Ian F. Akyildiz, Shih-Chun Lin, and Pu Wang. Wireless software-defined networks (W-SDNs) and network function virtualization (NFV) for 5G cellular systems: an overview and qualitative evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):66–79, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003862.

Ren:2015:DAC

[1710] Shoushou Ren, Tao Lin, Wei An, Guoqiang Zhang, Dalei Wu, Laxmi Narayan Bhuyan, and Zhen Xu. Design and analysis of collaborative EPC and RAN caching for LTE mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1) (??):80–95, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003850.

Savi:2015:PEV

[1711] Marco Savi, Roberto Fratini, Giacomo Verticale, and Massimo Tornatore. Performance evaluation of video server replication in metro/access networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1) (??):96–110, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003813.

Wang:2015:QDS

[1712] Hei-Chia Wang, Wei-Pin Chiu, and Suei-Chih Wu. QoS-driven selection of web service considering group preference. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1) (??):111–124, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003874.

Aguilar-Garcia:2015:LAS

[1713] A. Aguilar-Garcia, S. Fortes, Mariano Molina-García, Jaime Calle-Sánchez, José I. Alonso, Aaron Garrido, Alfonso Fernández-Durán, and R. Barco. Location-aware self-organizing methods in femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):125– 140, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003849.

Lakhlef:2015:FRS

[1714] Hicham Lakhlef and Julien Bourgeois. Fast and robust self-organization for micro-electro-mechanical robotic systems. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1) (??):141–152, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003059.

Mansour:2015:CSS

[1715] Samah Mansour, Intissar Harrabi, Martin Maier, and Géza Joós. Cosimulation study of performance tradeoffs between centralised, distributed, and hybrid adaptive PEV charging algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):153–165, December 24, 2015.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002960.

Jeong:2015:TTP

[1716] Jaehoon Paul Jeong, Jinyong Kim, Taehwan Hwang, Fulong Xu, Shuo Guo, Yu Jason Gu, Qing Cao, Ming Liu, and Tian He. TPD: Travel Prediction-based Data Forwarding for light-traffic vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):166– 182, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003898.

Zhu:2015:DRS

[1717] Konglin Zhu, Wenzhong Li, Xiaoming Fu, and Lin Zhang. Data routing strategies in opportunistic mobile social networks: Taxonomy and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):183–198, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003916.

Wang:2015:MRT

[1718] Tian Wang, Yiqiao Cai, Weijia Jia, Sheng Wen, Guojun Wang, Hui Tian, Yonghong Chen, and Bineng Zhong. Maximizing real-time streaming services based on a multiservers networking framework. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):199– 212, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003928.

Wang:2015:PPL

[1719] Wei Wang, Linlin Yang, and Qian Zhang. Privacy preservation in location-based advertising: a contractbased approach. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):213–224, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S138912861500393X.

Wang:2015:MMM

[1720] Xingwei Wang, Dapeng Qu, Min Huang, Keqin Li, Sajal K. Das, Jinhong Zhang, and Ruiyun Yu. Multiple many-to-many multicast routing scheme in green multigranularity transport networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):225– 242, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003886.

Braem:2015:CN

[1721] Bart Braem, Felix Freitag, and Roger Baig Viñas. Community networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2) (??):243–244, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003564.

Simo-Reigadas:2015:SLC

[1722] Javier Simo-Reigadas, Esteban Municio, Eduardo Morgado, Eva M. Castro, Andres Martinez, Luis F. Solorzano, and Ignacio Prieto-Egido. Sharing low-cost wireless infrastructures with telecommunications operators to bring 3G services to rural communities. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2)(??): 245–259, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003163.

Vega:2015:TOG

[1723] Davide Vega, Roger Baig, Llorenç Cerdà-Alabern, Esunly Medina, Roc Meseguer, and Leandro Navarro. A technological overview of the guifi.net community network. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2)(??):260– 278, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003436.

Kos:2015:NSG

[1724] Jernej Kos, Mitar Milutinović, and Luka Cehovin. nodewatcher: a substrate for growing your own community network. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2) (??):279–296, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003400.

Centelles:2015:ERS

[1725] Roger Pueyo Centelles, Victor Oncins, and Axel Neumann. Enhancing reflection and self-determination in a reallife community mesh network. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2)(??):297– 307, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003825.

Neumann:2015:EMR

[1726] Axel Neumann, Ester López, and Leandro Navarro. Evaluation of mesh routing protocols for wireless community networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2)(??):308–323, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615002522.

Barz:2015:OCN

[1727] Christoph Barz, Christoph Fuchs, Jonathan Kirchhoff, Julia Niewiejska, and Henning Rogge. OLSRv2 for community networks: Using Directional Airtime Metric with external radios. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2) (??):324–341, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003424.

Millan:2015:TSA

[1728] Pere Millan, Carlos Molina, Esunly Medina, Davide Vega, Roc Meseguer, Bart Braem, and Chris Blondia. Time series analysis to predict link quality of wireless community networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2)(??):342– 358, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003412.

Abujoda:2015:SDW

[1729] Ahmed Abujoda, David Dietrich, Panagiotis Papadimitriou, and Arjuna Sathiaseelan. Software-defined wireless mesh networks for Internet access sharing. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2) (??):359–372, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003187.

Selimi:2015:CSG

[1730] Mennan Selimi, Amin M. Khan, Emmanouil Dimogerontakis, Felix Freitag, and Roger Pueyo Centelles. Cloud services in the Guifi.net community network. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2) (??):373–388, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003175.

Baldesi:2015:IPS

[1731] Luca Baldesi, Leonardo Maccari, and Renato Lo Cigno. Improving P2P streaming in Wireless Community Networks. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2)

(??):389–403, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003448.

Anonymous:2015:EBr

[1732] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??): ifc, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004338.

Anonymous:2015:EBs

[1733] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 2)(??): ifc, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004399.

Boutaba:2015:CNC

[1734] Raouf Boutaba, Nelson Fonseca, Dzmitry Kliazovich, and Noura Limam. Cloud networking and communications II. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):405–407, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003540.

Persico:2015:MNT

[1735] Valerio Persico, Pietro Marchetta, Alessio Botta, and Antonio Pescapè. Measuring network throughput in the cloud: the case of Amazon EC2. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):408– 422, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500362X.

Couto:2015:SPS

[1736] Rodrigo S. Couto, Stefano Secci, Miguel Elias M. Campista, and Luís Henrique M. K. Costa. Server placement with shared backups for disaster-resilient clouds. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):423– 434, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003643.

Kavvadia:2015:EVM

[1737] Eleni Kavvadia, Spyros Sagiadinos, Konstantinos Oikonomou, Giorgos Tsioutsiouliklis, and Sonia Aïssa. Elastic virtual machine placement in cloud computing network environments. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3) (??):435–447, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003631.

Zhang:2015:EAV

[1738] Zhongbao Zhang, Sen Su, Junchi Zhang, Kai Shuang, and Peng Xu. Energy aware virtual network embedding with dynamic demands: Online and offline. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3) (??):448–459, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003618.

Ghazisaeedi:2015:EAN

[1739] Ebrahim Ghazisaeedi and Changcheng Huang. Energy-aware node and link reconfiguration for virtualized network environments. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):460–479, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128615003771.

Jarraya:2015:VFR

[1740] Yosr Jarraya, Arash Eghtesadi, Sahba Sadri, Mourad Debbabi, and Makan Pourzandi. Verification of firewall reconfiguration for virtual machines migrations in the cloud. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):480– 491, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500376X.

Sahhaf:2015:NSC

[1741] Sahel Sahhaf, Wouter Tavernier, Matthias Rost, Stefan Schmid, Didier Colle, Mario Pickavet, and Piet Demeester. Network service chaining with optimized network function embedding supporting service decompositions. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3) (??):492–505, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500359X.

Bellavista:2015:VNF

[1742] Paolo Bellavista, Franco Callegati, Walter Cerroni, Chiara Contoli, Antonio Corradi, Luca Foschini, Alessandro Pernafini, and Giuliano Santandrea. Virtual network function embedding in real cloud environments. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):506– 517, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003588.

Lin:2015:SPC

[1743] Wenjie Lin, Puneet Sharma, Sarbajit Chatterjee, Deepti Sharma, David Lee, Subu Iyer, and Ajay Gupta. Scaling persistent connections for cloud services. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3) (??):518–530, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615003576.

Marotta:2015:MMC

[1744] Marcelo Antonio Marotta, Leonardo Roveda Faganello, Matias Artur Klafke Schimuneck, Lisandro Zambenedetti Granville, Juergen Rochol, and Cristiano Bonato Both. Managing mobile cloud computing considering objective and subjective perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):531– 542, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003667.

Anonymous:2015:EBt

[1745] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??): ifc, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004569.

Anonymous:2016:TYR

[1746] Anonymous. Thank you reviewers. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): I–IX, January 15, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500496X.

Akyildiz:2016:CE

[1747] Ian F. Akyildiz and Harry Rudin. Comnet Editorial for 2015. Computer Networks (Amsterdam, Netherlands: 1999), 94(??):iii–iv, January 15, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004958.

Castro:2016:JCR

[1748] Pedro H. P. Castro, Vívian L. Barreto, Sand Luz Corrêa, Lisandro Zambenedetti Granville, and Kleber Vieira Cardoso. A joint CPU–RAM energy efficient and SLA-compliant approach for cloud data centers. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):1–13, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004752.

Neumayer:2016:NRU

[1749] Sebastian Neumayer and Eytan Modiano. Network reliability under geographically correlated line and disk failure models. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):14–28, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004740.

Deruyck:2016:OLW

[1750] Margot Deruyck, Emmeric Tanghe, David Plets, Luc Martens, and Wout Joseph. Optimizing LTE wireless access networks towards power consumption and electromagnetic exposure of human beings. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):29–40, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004727.

Sanchez:2016:EAC

[1751] M. Isabel Sanchez, Antonio de la Oliva, and Carlos J. Bernardos. Experimental analysis of connectivity management in mobile operating systems. Computer Networks (Amsterdam, Netherlands: 1999), 94(??):41–61, January 15, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004545.

Dai:2016:CCO

[1752] Huichen Dai and Bin Liu. CON-SERT: Constructing optimal namebased routing tables. Computer Networks (Amsterdam, Netherlands: 1999), 94(??):62–79, January 15, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004533.

Mangili:2016:PAC

[1753] Michele Mangili, Fabio Martignon, and Antonio Capone. Performance analysis of Content-Centric and Content-Delivery networks with evolving object popularity. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):80–98, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004521.

Garroppo:2016:IAP

[1754] Rosario G. Garroppo, Gianfranco Nencioni, Gregorio Procissi, and Luca Tavanti. The impact of the access point power model on the energyefficient management of infrastructured wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):99–111, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500451X.

Mendiola:2016:ADQ

[1755] Alaitz Mendiola, Victor Fuentes, Jon Matias, Jasone Astorga, Nerea Toledo, Eduardo Jacob, and Maider Huarte. An architecture for dynamic QoS management at Layer 2 for DOC-SIS access networks using Open-Flow. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 112–128, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004508.

Nguyen:2016:DFA

[1756] Tien-Thinh Nguyen and Christian Bonnet. DMMS: a flexible architecture for multicast listener support in a distributed mobility management environment. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 129–144, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500448X.

Degefa:2016:PSE

[1757] Fikadu B. Degefa, Donghoon Lee, Jiye Kim, Younsung Choi, and Dongho Won. Performance and security enhanced authentication and key agreement protocol for SAE/LTE network. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 145–163, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004211.

Mishra:2016:ELA

[1758] Sudeepta Mishra and C. Siva Ram Murthy. An efficient location aware distributed physical resource block assignment for dense closed access femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):164–175, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500420X.

Han:2016:DHG

[1759] Song Han, Xinbin Li, Zhixin Liu, and Xinping Guan. Distributed hierarchical game-based algorithm for downlink power allocation in OFDMA femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):176–188, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004181.

Kolios:2016:EEM

[1760] Panayiotis Kolios, Katerina Papadaki, and Vasilis Friderikos. Energy efficient mobile video streaming using mobility. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 189–204, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500417X.

Pecori:2016:KTR

[1761] Riccardo Pecori. S-Kademlia: a trust and reputation method to mitigate a Sybil attack in Kademlia. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 205–218, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004168.

Zhang:2016:CCN

[1762] Guoqiang Zhang and Ziqu Xu.

Combing CCN with network coding: an architectural perspective. Computer Networks (Amsterdam, Netherlands: 1999), 94(??):219– 230, January 15, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004144.

Ding:2016:ABC

[1763] Jianbing Ding, Zhenjie Zhang, Richard T. B. Ma, and Yin Yang. Auctionbased cloud service differentiation with service level objectives. Computer Networks (Amsterdam, Netherlands: 1999), 94(??):231–249, January 15, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128615004132.

Han:2016:EAQ

[1764] Qiaoni Han, Bo Yang, Cailian Chen, and Xinping Guan. Energy-aware and QoS-aware load balancing for HetNets powered by renewable energy. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 250–262, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004120.

Patota:2016:DDA

[1765] Federico Patota, Luca Chiaraviglio, Francesco Bella, Vincenzo Deriu, Silvia Fortunato, and Francesca Cuomo. DAFNES: a distributed algorithm for network energy saving based on stresscentrality. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 263–284, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004119.

Rendon:2016:RDM

[1766] Oscar Mauricio Caicedo Rendon, Felipe Estrada-Solano, Vinicius Guimarães, Liane Margarida Rockenbach Tarouco, and Lisandro Zambenedetti Granville. Rich dynamic mashments: an approach for network management based on mashups and situation management. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 285–306, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004107.

Onus:2016:PMA

[1767] Melih Onus and Andréa W. Richa. Parameterized maximum and average degree approximation in topicbased publish-subscribe overlay network design. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 307–317, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004090.

Wang:2016:BPF

[1768] Yang Wang, Qian Hu, and Xiaojun Cao. A branch-and-price framework for optimal virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 318–326, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004089.

Bou-Harb:2016:NCS

[1769] Elias Bou-Harb, Mourad Debbabi, and Chadi Assi. A novel cyber security capability: Inferring Internet-scale infections by correlating malware and probing activities. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):327–343, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004077.

Wang:2016:SDD

[1770] You-Chiun Wang and Song-Yun Hsieh.

Service-differentiated downlink flow scheduling to support QoS in long term evolution. Computer Networks (Amsterdam, Netherlands: 1999), 94 (??):344–359, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004053.

Zhang:2016:MDF

[1771] Jie Zhang, Dafang Zhang, Kun Huang, and Zheng Qin. Minimizing datacenter flow completion times with server-based flow scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 360–374, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500403X.

Liu:2016:QAR

[1772] Tzu-Chin Liu, Kuochen Wang, Chia-Yu Ku, and Yi-Huai Hsu. QoSaware resource management for multimedia traffic report systems over LTE-A. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 375–389, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004028.

Ferretti:2016:SHM

[1773] Stefano Ferretti, Vittorio Ghini, and Fabio Panzieri. A survey on handover management in mobility architectures. Computer Networks (Amsterdam, Netherlands: 1999), 94(??): 390–413, January 15, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004491.

Anonymous:2016:EBa

[1774] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 94(??):ifc, January 15, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128615004892.

Peng:2016:WDC

[1775] Jia Peng, Yanmin Zhu, Wei Shu, and Min-You Wu. When data contributors meet multiple crowdsourcers: Bilateral competition in mobile crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 95(??): 1–14, February 11, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004764.

Meharouech:2016:TSG

[1776] Amira Meharouech, Jocelyne Elias, and Ahmed Mehaoua. A two-stage game theoretical approach for interference mitigation in Body-to-Body Networks. Computer Networks (Amsterdam, Netherlands: 1999), 95(??): 15–34, February 11, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004776.

Hartmann:2016:ROI

[1777] Matthias Hartmann, David Hock, and Michael Menth. Routing optimization for IP networks with loop-free alternates. Computer Networks (Amsterdam, Netherlands: 1999), 95(??): 35–50, February 11, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004156.

Timoteo:2016:AUS

[1778] Robson D. A. Timoteo, Lizandro N. Silva, Daniel C. Cunha, and George D. C. Cavalcanti. An approach using support vector regression for mobile location in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 95(??): 51–61, February 11, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500479X.

Oliveira:2016:PEH

[1779] Roberto M. Oliveira, Michelle S. P. Facina, Moises V. Ribeiro, and Alex B. Vieira. Performance evaluation of in-home broadband PLC systems using a cooperative MAC protocol. Computer Networks (Amsterdam, Netherlands: 1999), 95(??): 62–76, February 11, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004806.

Tennina:2016:ZMP

[1780] Stefano Tennina, Olfa Gaddour, Anis Koubâa, Fernando Royo, Mário Alves, and Mohamed Abid. Z-Monitor: a protocol analyzer for IEEE 802.15.4-based low-power wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 95 (??):77–96, February 11, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004788.

Cordeiro:2016:MPG

[1781] Weverton Luis da Costa Cordeiro, Flávio Roberto Santos, Marinho Pilla Barcelos, Luciano Paschoal Gaspary, Hanna Kavalionak, Alessio Guerrieri, and Alberto Montresor. Making puzzles green and useful for adaptive identity management in large-scale distributed systems. Computer Networks (Amsterdam, Netherlands: 1999), 95 (??):97–114, February 11, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004818.

Ostovari:2016:RWT

[1782] Pouya Ostovari and Jie Wu. Robust wireless transmission of scalable coded videos using two-dimensional network coding. Computer Networks (Amsterdam, Netherlands: 1999), 95(??): 115–126, February 11, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004715.

Lombera:2016:CPP

[1783] I. Michel Lombera, L. E. Moser, P. M.

Melliar-Smith, and Y. T. Chuang. Corrigendum to “Peer-to-Peer Publication, Search and Retrieval Using the Android Mobile Platform” [Computer Networks 65(2014) 56–72]. Computer Networks (Amsterdam, Netherlands: 1999), 95(??):127, February 11, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004041. See

[1225].

Anonymous:2016:EBb

[1784] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 95(??):ifc, February 11, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128616000116.

Tsiropoulou:2016:URA

[1785] Eirini Eleni Tsiropoulou, Aggelos Kapoukakis, and Symeon Papavassiliou. Uplink resource allocation in SC-FDMA wireless networks: a survey and taxonomy. Computer Networks (Amsterdam, Netherlands: 1999), 96 (??):1–28, February 26, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861500482X.

Mojamed:2016:SPP

[1786] Mohammad Al Mojamed and Mario Kolberg. Structured Peer-to-Peer overlay deployment on MANET: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 96(??): 29–47, February 26, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004831.

Nagano:2016:ESC

[1787] Junichi Nagano and Norihiko Shinomiya. Efficient switch clustering for distributed controllers of OpenFlow network with biconnectivity. Computer Networks (Amsterdam, Netherlands: 1999), 96 (??):48–57, February 26, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615003904.

Lee:2016:PSM

[1788] Kyu-Hwan Lee, Jae-Hyun Kim, and Sunghyun Cho. Power saving mechanism with network coding in the bottleneck zone of multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 96 (??):58–68, February 26, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002285.

Lin:2016:JOQ

[1789] Shih-Chun Lin, Pu Wang, and Min Luo. Jointly optimized QoS-aware virtualization and routing in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 96 (??):69–78, February 26, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500256X.

Anonymous:2016:EBc

[1790] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 96(??):ifc, February 26, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616300056.

Raspall:2016:BNS

[1791] Frederic Raspall. Building Nemo, a system to monitor IP routing and traffic paths in real time. Computer Networks (Amsterdam, Netherlands: 1999), 97(??):1–30, March 14, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004879.

Baig:2016:CAC

[1792] Zubair A. Baig, Sadiq M. Sait, and Farid Binbeshr. Controlled access to cloud resources for mitigating Economic Denial of Sustainability (EDoS) attacks. Computer Networks (Amsterdam, Netherlands: 1999), 97(??): 31–47, March 14, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616000050.

Kwon:2016:PSB

[1793] Jonghoon Kwon, Jehyun Lee, Heejo Lee, and Adrian Perrig. PsyBoG: a scalable botnet detection method for large-scale DNS traffic. Computer Networks (Amsterdam, Netherlands: 1999), 97(??):48–73, March 14, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004843.

Brandner:2016:FSP

[1794] G. Brandner, U. Schilcher, and C. Bettstetter. Firefly synchronization with phase rate equalization and its experimental analysis in wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 97 (??):74–87, March 14, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616000049.

Sui:2016:DAC

[1795] Yang Sui, Xiumin Wang, Jin Wang, Lusheng Wang, and Saihang Hou. Deadline-aware cooperative data exchange with network coding. Computer Networks (Amsterdam, Netherlands: 1999), 97(??):88–97, March 14, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616000098.

Madhja:2016:HCW

[1796] Adelina Madhja, Sotiris Nikoletseas, and Theofanis P. Raptis. Hierarchical, collaborative wireless energy transfer in sensor networks with multiple Mobile Chargers. Computer Networks (Amsterdam, Netherlands: 1999), 97 (??):98–112, March 14, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616000219.

Al-Awami:2016:DDS

[1797] Louai Al-Awami and Hossam S. Hassanein. Distributed data storage systems for data survivability in wireless sensor networks using decentralized erasure Cod’s. Computer Networks (Amsterdam, Netherlands: 1999), 97 (??):113–127, March 14, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616000220.

Bhardwaj:2016:CSP

[1798] Onkar Bhardwaj, Elliot Anshelevich, and Koushik Kar. Coalitionally stable pricing schemes for inter-domain forwarding. Computer Networks (Amsterdam, Netherlands: 1999), 97(??): 128–146, March 14, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616000025.

Anonymous:2016:EBd

[1799] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 97(??):ifc, March 14, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616300299.

Gai:2016:PDM

[1800] Yi Gai, Hua Liu, and Bhaskar Krishnamachari. A packet dropping mechanism for efficient operation of M/M/1 queues with selfish users. Computer Networks (Amsterdam, Netherlands: 1999), 98(??):1–13, April 7, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004855.

Lee:2016:RRL

[1801] Eun Kyung Lee, Hariharasudhan Viswanathan, and Dario Pompili. RescueNet: Reinforcement-learning-based communication framework for emergency networking. Computer Networks (Amsterdam, Netherlands: 1999), 98 (??):14–28, April 7, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300032.

Yi:2016:PAP

[1802] Changyan Yi, Zhen Zhao, Jun Cai, Ricardo Lobato de Faria, and Gong (Michael) Zhang. Priorityaware pricing-based capacity sharing scheme for beyond-wireless body area networks. Computer Networks (Amsterdam, Netherlands: 1999), 98 (??):29–43, April 7, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300020.

Wang:2016:JRS

[1803] Luyao Wang, Kwan-Wu Chin, and Sieteng Soh. Joint routing and scheduling in multi-Tx/Rx wireless mesh networks with random demands. Computer Networks (Amsterdam, Netherlands: 1999), 98(??):44–56, April 7, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300123.

Gokturk:2016:PCL

[1804] M. Sarper Gokturk, Ozgur Gurbuz, and Murat Erman. A practical cross layer cooperative MAC framework for WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 98 (??):57–71, April 7, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300135.

Akbas:2016:MGI

[1805] Mustafa Ilhan Akbas, Gürkan Solmaz, and Damla Turgut. Molecular geometry inspired positioning for aerial networks. Computer Networks (Amsterdam, Netherlands: 1999), 98 (??):72–88, April 7, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300172.

Avonts:2016:FCT

[1806] Jeroen Avonts and Chris Blondia. A framework to compare topology algorithms in multi-channel multi-radio wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 98(??):89–108, April 7, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300184.

Amjad:2016:CHS

[1807] Muhammad Faisal Amjad, Mainak Chatterjee, and Cliff C. Zou. Coexistence in heterogeneous spectrum through distributed correlated equilibrium in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 98(??): 109–122, April 7, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300160.

Anonymous:2016:EBe

[1808] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 98(??):ifc, April 7, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300470.

Jaron:2016:QAM

[1809] Alexandre Jaron, Andrej Mihailovic, and A. H. Aghvami. QoS-aware multi-plane routing method for OSPFbased IP access networks. Computer Networks (Amsterdam, Netherlands: 1999), 99(??):1–14, April 22, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300214.

Ding:2016:CLB

[1810] Xu Ding, Xinjiang Sun, Cheng Huang, and Xiaobei Wu. Cluster-level based link redundancy with network coding in duty cycled relay wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 99 (??):15–36, April 22, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300196.

Einziger:2016:SEK

[1811] Gil Einziger, Roy Friedman, and Yoav Kantor. Shades: Expediting Kademlia’s lookup process. Computer Networks (Amsterdam, Netherlands: 1999), 99(??):37–50, April 22, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300147.

Xiao:2016:OPS

[1812] Xun Xiao, Rui Zhang, Jianping Wang, Chunming Qiao, and Kejie Lu. An optimal pricing scheme to improve transmission opportunities for a mobile virtual network operator. Computer Networks (Amsterdam, Netherlands: 1999), 99(??):51–65, April 22, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300202.

Lai:2016:GGB

[1813] Chengzhe Lai, Rongxing Lu, Dong Zheng, Hui Li, and Xuemin (Sherman) Shen. GLARM: Group-based lightweight authentication scheme for resource-constrained machine to machine communications. Computer Networks (Amsterdam, Netherlands: 1999), 99(??):66–81, April 22, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300238.

Chang:2016:ODS

[1814] Xiangmao Chang, Jin Wang, Jianping Wang, Kejie Lu, and Yi Zhuang. On the optimal design of secure network coding against wiretapping attack. Computer Networks (Amsterdam, Netherlands: 1999), 99(??): 82–98, April 22, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004995.

Al-Mefleh:2016:TAJ

[1815] Haithem Al-Mefleh and Osameh Al-Kofahi. Taking advantage of jamming in wireless networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 99(??): 99–124, April 22, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300275.

Vieira:2016:SWL

[1816] Fabio R. J. Vieira, José F. de Rezende, and Valmir C. Barbosa. Scheduling wireless links by vertex multicoloring in the physical interference model. Computer Networks (Amsterdam, Netherlands: 1999), 99(??): 125–133, April 22, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630024X.

Iqbal:2016:MOO

[1817] M. Iqbal, M. Naeem, A. Anpalagan, N. N. Qadri, and M. Imran. Multi-objective optimization in sensor networks: Optimization classification, applications and solution approaches. Computer Networks (Amsterdam, Netherlands: 1999), 99(??): 134–161, April 22, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300159.

Anonymous:2016:EBf

[1818] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 99(??):ifc, April 22, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300627.

Tang:2016:EEP

[1819] Meiqin Tang and Yalin Xin. Energy efficient power allocation in cognitive radio network using coevolution chaotic particle swarm optimization. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):1–11, May 8, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300263.

Jakalan:2016:SRD

[1820] Ahmad Jakalan, Jian Gong, Qi Su, Xiaoyan Hu, and Abdeldime M. S. Abdelgder. Social relationship discovery of IP addresses in the managed IP networks by observing traffic at network boundary. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):12–27, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300342.

Berger:2016:MAD

[1821] Andreas Berger, Alessandro D’Alconzo, Wilfried N. Gansterer, and Antonio Pescapé. Mining agile DNS traffic using graph analysis for cybercrime detection. Computer Networks (Amsterdam, Netherlands: 1999), 100 (??):28–44, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300251.

Chen:2016:EEJ

[1822] Hongbin Chen, Qiong Zhang, and Feng Zhao. Energy-efficient joint BS and RS sleep scheduling in relayassisted cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):45–54, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630038X.

Kim:2016:ASV

[1823] Sungwook Kim. Asymptotic Shapley value based resource allocation scheme for IoT services. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):55–63, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300433.

Xenakis:2016:HDS

[1824] Dionysis Xenakis, Nikos Passas, Lazaros Merakos, and Christos Verikoukis. Handover decision for small cells: Algorithms, lessons learned and simulation study. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):64–74, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004065.

Haque:2016:ACV

[1825] Syed Arefinul Haque, Salekul Islam, Md. Jahidul Islam, and Jean-Charles Grégoire. An architecture for client virtualization: a case study. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):75–89, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300421.

Eguizabal:2016:JDR

[1826] Miguel Eguizábal and Ángela Hernández.

Joint dynamic resource allocation and load balancing-cell selection in LTE-A HetNet scenarios based on Type 1 inband relay deployments. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):90–109, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300408.

Malanchini:2016:WRS

[1827] Ilaria Malanchini, Stefan Valentin, and Osman Aydin. Wireless resource sharing for multiple operators: Generalization, fairness, and the value of prediction. Computer Networks (Amsterdam, Netherlands: 1999), 100 (??):110–123, May 8, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300366.

Gunduz:2016:PBS

[1828] Gurhan Gunduz and Murat Yuksel. Popularity-based scalable peer-to-peer topology growth. Computer Networks (Amsterdam, Netherlands: 1999), 100 (??):124–140, May 8, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300391.

Sharon:2016:CIT

[1829] Oran Sharon and Yaron Alpert. Coupled IEEE 802.11ac and TCP performance evaluation in various aggregation schemes and Access Categories. Computer Networks (Amsterdam, Netherlands: 1999), 100 (??):141–156, May 8, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300457.

Melo:2016:OPD

[1830] Bruna L. R. Melo, Daniel C. Cunha, and Cecilio Pimentel. Optimal power distribution in non-binary LDPC code-based cooperative wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 100 (??):157–165, May 8, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300445.

Kok:2016:SFR

[1831] Gin-Xian Kok, Chee-Onn Chow, Yi-Han Xu, and Hiroshi Ishii. Sensorfree route stability metric for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 100 (??):166–178, May 8, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300536.

Gomes:2016:BAA

[1832] Rafael L. Gomes, Luiz F. Bittencourt, Edmundo R. M. Madeira, Eduardo Cerqueira, and Mario Gerla. Bandwidth-aware allocation of resilient Virtual Software Defined Networks. Computer Networks (Amsterdam, Netherlands: 1999), 100 (??):179–194, May 8, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300524.

Anonymous:2016:EBg

[1833] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):ifc, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300809.

Zhang:2016:ITA

[1834] Daqiang Zhang, Jiafu Wan, Ching-Hsien (Robert) Hsu, and Ammar Rayes. Industrial technologies and applications for the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):1–4, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630041X.

Xia:2016:CLD

[1835] Min Xia, Teng Li, Yunfei Zhang, and Clarence W. de Silva. Closed-loop design evolution of engineering system using condition monitoring through Internet of Things and cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):5–18, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615005034.

Han:2016:GBJ

[1836] Guangjie Han, Aihua Qian, Jinfang Jiang, Ning Sun, and Li Liu. A gridbased joint routing and charging algorithm for industrial wireless rechargeable sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 101(??):19–28, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615005010.

Zhang:2016:PPQ

[1837] Bo Zhang, Chi Harold Liu, Jianyu Lu, Zheng Song, Ziyu Ren, Jian Ma, and Wendong Wang. Privacy-preserving QoI-aware participant coordination for mobile crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 101(??):29–41, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616000074.

Amin:2016:DAP

[1838] Ruhul Amin, Sk Hafizul Islam, G. P.

Biswas, Muhammad Khurram Khan, Lu Leng, and Neeraj Kumar. Design of an anonymity-preserving three-factor authenticated key exchange protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 101(??):42–62, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616000207.

Rathore:2016:UPB

[1839] M. Mazhar Rathore, Awais Ahmad, Anand Paul, and Seungmin Rho. Urban planning and building smart cities based on the Internet of Things using Big Data analytics. Computer Networks (Amsterdam, Netherlands: 1999), 101(??):63–80, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616000086.

Luo:2016:LSW

[1840] Xiong Luo, Ji Liu, Dandan Zhang, and Xiaohui Chang. A large-scale web QoS prediction scheme for the Industrial Internet of Things based on a kernel machine learning algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 101(??):81–89, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616000189.

Huang:2016:EEC

[1841] Wei Huang and Liangmin Wang. ECDS: Efficient collaborative downloading scheme for popular content distribution in urban vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):90–103, June 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300226.

Liu:2016:GDC

[1842] Qiang Liu, Yujun Ma, Musaed Alhussein, Yin Zhang, and Limei Peng. Green data center with IoT sensing and cloud-assisted smart temperature control system. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):104–112, June 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615004739.

Lin:2016:HLB

[1843] Kai Lin, Wenjian Wang, Yuanguo Bi, Meikang Qiu, and Mohammad Mehedi Hassan. Human localization based on inertial sensors and fingerprints in the Industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):113–126, June 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615004193.

Qiu:2016:GMS

[1844] Tie Qiu, Diansong Luo, Feng Xia, Nakema Deonauth, Weisheng Si, and Amr Tolba. A greedy model with small world for improving the robustness of heterogeneous Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):127–143, June 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861500506X.

Kong:2016:IIC

[1845] Linghe Kong, Qiao Xiang, Xue Liu, Xiao-Yang Liu, Xiaofeng Gao, Guihai Chen, and Min-You Wu. ICP: Instantaneous clustering protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 101(??):144–157, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616000062.

Wang:2016:TSF

[1846] Shiyong Wang, Jiafu Wan, Daqiang Zhang, Di Li, and Chunhua Zhang. Towards smart factory for industry 4.0: a self-organized multi-agent system with big data based feedback and coordination. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):158–168, June 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615005046.

Kanaris:2016:SSD

[1847] Loizos Kanaris, Akis Kokkinis, Giancarlo Fortino, Antonio Liotta, and Stavros Stavrou. Sample Size Determination Algorithm for fingerprint-based indoor localization systems. Computer Networks (Amsterdam, Netherlands: 1999), 101(??):169–177, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128615005022.

Tsai:2016:EWD

[1848] Chun-Wei Tsai. An effective WSN deployment algorithm via search economics. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):178–191, June 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616000190.

Hossain:2016:CAI

[1849] M. Shamim Hossain and Ghulam Muhammad. Cloud-assisted Industrial Internet of Things (IIoT) — enabled framework for health monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 101 (??):192–202, June 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300019.

Anonymous:2016:EBh

[1850] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 101(??):ifc, June 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301256.

Yamanaka:2016:TFF

[1851] Hiroaki Yamanaka, Eiji Kawai, and Shinji Shimojo. A technique for full flow virtualization of multitenant OpenFlow networks. Computer Networks (Amsterdam, Netherlands: 1999), 102(??):1–19, June 19, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300585.

Wang:2016:RPC

[1852] Ke Wang, Teck Yoong Chai, and Wai-Choong Wong. Routing, power control and rate adaptation: a Q-learning-based cross-layer design. Computer Networks (Amsterdam, Netherlands: 1999), 102(??):20–37, June 19, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300573.

Qadir:2016:QEA

[1853] Qahhar Muhammad Qadir, Alexander A. Kist, and Zhongwei Zhang. A quality of experience-aware crosslayer architecture for optimizing video streaming services. Computer Networks (Amsterdam, Netherlands: 1999), 102 (??):38–49, June 19, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300597.

Nikkhah:2016:MPI

[1854] Mehdi Nikkhah. Maintaining the progress of IPv6 adoption. Computer Networks (Amsterdam, Netherlands: 1999), 102(??):50–69, June 19, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630055X.

Sankar:2016:QPM

[1855] V. Udaya Sankar and Vinod Sharma. QoS provisioning for multiple Femtocells via game theory. Computer Networks (Amsterdam, Netherlands: 1999), 102(??):70–82, June 19, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300548.

Malina:2016:PSP

[1856] Lukas Malina, Jan Hajny, Radek Fujdiak, and Jiri Hosek. On perspective of security and privacypreserving solutions in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 102 (??):83–95, June 19, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300779.

Jedda:2016:DRC

[1857] Ahmed Jedda, Mazen G. Khair, and Hussein T. Mouftah. Decentralized RFID coverage algorithms using writeable tags. Computer Networks (Amsterdam, Netherlands: 1999), 102 (??):96–108, June 19, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300676.

Li:2016:SRA

[1858] Bing Li and J. William Atwood. Secure receiver access control for IP multicast at the network level: Design and validation. Computer Networks (Amsterdam, Netherlands: 1999), 102 (??):109–128, June 19, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300706.

Zhao:2016:DIS

[1859] Tao Zhao, Tao Li, Biao Han, Zhigang Sun, and Jinfeng Huang. Design and implementation of Software Defined Hardware Counters for SDN. Computer Networks (Amsterdam, Netherlands: 1999), 102(??):129–144, June 19, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300688.

Liu:2016:RPC

[1860] Zhixin Liu, Peng Zhang, Xinping Guan, and Hongjiu Yang. Robust power control for femtocell networks under outage-based QoS constraints. Computer Networks (Amsterdam, Netherlands: 1999), 102 (??):145–156, June 19, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300354.

Wang:2016:IMP

[1861] Yingjie Wang, Zhipeng Cai, Guisheng Yin, Yang Gao, Xiangrong Tong, and Guanying Wu. An incentive mechanism with privacy protection in mobile crowdsourcing systems. Computer Networks (Amsterdam, Netherlands: 1999), 102(??):157–171, June 19, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300883.

Xiong:2016:PEO

[1862] Bing Xiong, Kun Yang, Jinyuan Zhao, Wei Li, and Keqin Li. Performance evaluation of OpenFlow-based software-defined networks based on queueing model. Computer Networks (Amsterdam, Netherlands: 1999), 102 (??):172–185, June 19, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630069X.

Anonymous:2016:EBi

[1863] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 102(??):ifc, June 19, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616301414.

Militano:2016:ENS

[1864] L. Militano, M. Nitti, L. Atzori, and A. Iera. Enhancing the navigability in a social network of smart objects: a Shapley-value based approach. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):1–14, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300743.

LeCadre:2016:DUS

[1865] Hélène Le Cadre and Jean-Sébastien Bedo. Dealing with uncertainty in the smart grid: a learning game approach. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):15–32, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300603.

Li:2016:TBC

[1866] Wei Li and Yuwei Wu. Tree-based coverage hole detection and healing method in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):33–43, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301037.

Lee:2016:ERA

[1867] Ming-Chieh Lee and Jang-Ping Sheu.

An efficient routing algorithm based on segment routing in software-defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):44–55, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300871.

Ma:2016:PRF

[1868] Lisheng Ma, Xiaohong Jiang, Bin Wu, Achille Pattavina, and Norio Shiratori. Probabilistic region failureaware data center network and content placement. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):56–66, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300755.

Sourlas:2016:EHR

[1869] Vasilis Sourlas, Ioannis Psaras, Lorenzo Saino, and George Pavlou. Efficient hash-routing and domain clustering techniques for information-centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):67–83, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300998.

Rhaiem:2016:NCB

[1870] Olfa Ben Rhaiem, Lamia Chaari Fourati, and Wessam Ajib. Network coding-based approach for efficient video streaming over MANET. Computer Networks (Amsterdam, Netherlands: 1999), 103(??):84– 100, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301001.

Byun:2016:TST

[1871] Sang-Seon Byun. TCP over scarce transmission opportunity in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):101–114, July 5, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300731.

Wang:2016:EEP

[1872] Shiguo Wang, Rukhsana Ruby, Victor C. M. Leung, and Zhiqiang Yao. Energy-efficient power allocation for multi-user single-AF-relay underlay cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 103(??):115–128, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301049.

Hoteit:2016:FNC

[1873] Sahar Hoteit, Mahmoud El Chamie, Damien Saucez, and Stefano Secci. On fair network cache allocation to content providers. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):129–142, July 5, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301050.

Erdem:2016:PHA

[1874] Oguzhan Erdem. Pipelined hierarchical architecture for high performance packet classification. Computer Networks (Amsterdam, Netherlands: 1999), 103(??):143–164, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301074.

Millan:2016:RLB

[1875] Víctor M. López Millán, Vicent Cholvi, Antonio Fernández Anta, and Luis López. Resource location based on precomputed partial random walks in dynamic networks. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):165–180, July 5, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301062.

Constantinou:2016:HAE

[1876] Costas K. Constantinou and Georgios Ellinas. Heuristic algorithms for efficient allocation of multicastcapable nodes in sparse-splitting optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):181–195, July 5, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300561.

Socievole:2016:ANR

[1877] A. Socievole, F. De Rango, C. Scoglio, and P. Van Mieghem. Assessing network robustness under SIS epidemics: the relationship between epidemic threshold and viral conductance. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):196–206, July 5, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301141.

Parmar:2016:CDA

[1878] Keyur Parmar and Devesh C. Jinwala. Concealed data aggregation in wireless sensor networks: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 103 (??):207–227, July 5, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301116.

Luo:2016:CMB

[1879] Xiapu Luo, Haocheng Zhou, Le Yu, Lei Xue, and Yi Xie. Characterizing mobile ^{∗}-box applications. Computer Networks (Amsterdam, Netherlands: 1999), 103(??):228–239, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300962.

Anonymous:2016:EBj

[1880] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 103(??):ifc, July 5, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301542.

Chellappan:2016:CEM

[1881] Vanniyarajan Chellappan, Krishna M. Sivalingam, and Kamala Krithivasan. A centrality entropy maximization problem in shortest path routing networks. Computer Networks (Amsterdam, Netherlands: 1999), 104 (??):1–15, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630113X.

Sun:2016:TCA

[1882] Yanming Sun, Min Chen, Abel Bacchus, and Xiaodong Lin. Towards collusion-attack-resilient group key management using one-way function tree. Computer Networks (Amsterdam, Netherlands: 1999), 104(??): 16–26, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301128.

Joldzic:2016:TSA

[1883] Ognjen Joldzic, Zoran Djuric, and Pavle Vuletic. A transparent and scalable anomaly-based DoS detection method. Computer Networks (Amsterdam, Netherlands: 1999), 104 (??):27–42, July 20, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301347.

Karaca:2016:EBA

[1884] Mehmet Karaca, Ozgur Ercetin, and Tansu Alpcan. Entropy-based active learning for wireless scheduling with incomplete channel feedback. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):43–54, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301220.

Wang:2016:TSN

[1885] Yang Wang, Phanvu Chau, and Fuyu Chen. Towards a secured network virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 104 (??):55–65, July 20, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301219.

Said:2016:MIR

[1886] Adel Mounir Said, Michel Marot, Ashraf William Ibrahim, and Hossam Afifi. Modeling interactive realtime applications in VANETs with performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):66–78, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301207.

Bhatia:2016:TMT

[1887] Ashutosh Bhatia and R. C. Hansdah. TRM-MAC: a TDMA-based reliable multicast MAC protocol for WSNs with flexibility to trade-off between latency and reliability. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):79–93, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301153.

Huang:2016:ORA

[1888] Gaofei Huang and Wanqing Tu. Optimal resource allocation in wirelesspowered OFDM relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):94– 107, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301372.

Mehajabin:2016:ESR

[1889] Nusrat Mehajabin, Md. Abdur Razzaque, Mohammad Mehedi Hassan, Ahmad Almogren, and Atif Alamri. Energy-sustainable relay node deployment in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 104 (??):108–121, July 20, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301517.

Bessani:2016:JJT

[1890] Alysson Bessani, Nuno F. Neves, Paulo Veríssimo, Wagner Dantas, Alexandre Fonseca, Rui Silva, Pedro Luz, and Miguel Correia. JITeR: Just-in-time application-layer routing. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):122–136, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301475.

Kumari:2016:APW

[1891] Shipra Kumari and Hari Om. Authentication protocol for wireless sensor networks applications like safety monitoring in coal mines. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):137–154, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301384.

Yi:2016:HDA

[1892] D. Yi and H. Yang. HEER — a delay-aware and energy-efficient routing protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 104 (??):155–173, July 20, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301190.

Wei:2016:DCT

[1893] Mingkui Wei and Wenye Wang. Datacentric threats and their impacts to real-time communications in smart grid. Computer Networks (Amsterdam, Netherlands: 1999), 104(??): 174–188, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301359.

Zhang:2016:EBC

[1894] Deyu Zhang, Zhigang Chen, Haibo Zhou, Long Chen, and Xuemin (Sherman) Shen. Energy-balanced cooperative transmission based on relay selection and power control in energy harvesting wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 104 (??):189–197, July 20, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301505.

Yousaf:2016:OTM

[1895] Faqir Zarrar Yousaf, Christian Wietfeld, and Sahibzada Ali Mahmud. Optimizing tunnel management in predictive handover protocols. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):198–212, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301463.

Anonymous:2016:EBk

[1896] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):ifc, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301682.

Yedugundla:2016:MPT

[1897] Kiran Yedugundla, Simone Ferlin, Thomas Dreibholz, Özgü Alay, Nicolas Kuhn, Per Hurtig, and Anna Brunstrom. Is multi-path transport suitable for latency sensitive traffic? Computer Networks (Amsterdam, Netherlands: 1999), 105(??):1–21, August 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301396.

Yoon:2016:RAA

[1898] Sung-Guk Yoon, Jeong-O Seo, and Saewoong Bahk. Regrouping algorithm to alleviate the hidden node problem in 802.11ah networks. Computer Networks (Amsterdam, Netherlands: 1999), 105(??):22–32, August 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301487.

Dinh:2016:MWT

[1899] Thanh Dinh, Younghan Kim, Tao Gu, and Athanasios V. Vasilakos. L-MAC: a wake-up time self-learning MAC protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):33–46, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301529.

Cao:2016:ROT

[1900] Tianjie Cao, Xiuqing Chen, Robin Doss, Jingxuan Zhai, Lucas J. Wise, and Qiang Zhao. RFID ownership transfer protocol based on cloud. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):47–59, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301621.

Messai:2016:SKM

[1901] Mohamed-Lamine Messai and Hamida Seba. A survey of key management schemes in multi-phase wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):60–74, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301360.

Bianco:2016:STM

[1902] Andrea Bianco, Paolo Giaccone, and Marco Ricca. Scheduling traffic for maximum switch lifetime in optical data center fabrics. Computer Networks (Amsterdam, Netherlands: 1999), 105(??):75–88, August 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301232.

Shih:2016:FBF

[1903] Chi-Huang Shih, Chun-I Kuo, and Yeh-Kai Chou. Frame-based forward error correction using contentdependent coding for video streaming applications. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):89–98, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630161X.

Nguyen:2016:MLD

[1904] Ngoc-Tu Nguyen, Bing-Hong Liu, Van-Trung Pham, and Yi-Sheng Luo. On maximizing the lifetime for data aggregation in wireless sensor networks using virtual data aggregation trees. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):99–110, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630175X.

Xie:2016:VOE

[1905] Ling Fu Xie, Peter Han Joo Chong, Ivan Wang-Hei Ho, and Henry C. B. Chan. Virtual overhearing: an effective way to increase network coding opportunities in wireless ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):111–123, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301761.

Costa:2016:LAM

[1906] Lucas R. Costa, Guilherme N. Ramos, and André C. Drummond. Leveraging adaptive modulation with multihop routing in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):124–137, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301785.

Pahlevani:2016:NCH

[1907] Peyman Pahlevani, Hana Khamfroush, Daniel E. Lucani, Morten V. Pedersen, and Frank H. P. Fitzek. Network coding for hop-by-hop communication enhancement in multi-hop networks. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):138–149, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301499.

Michaloliakos:2016:PMN

[1908] Antonios Michaloliakos, Ryan Rogalin, Yonglong Zhang, Konstantinos Psounis, and Giuseppe Caire. Performance modeling of next-generation WiFi networks. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):150–165, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301773.

Cui:2016:DDF

[1909] Wenzhi Cui, Ye Yu, and Chen Qian. DiFS: Distributed Flow Scheduling for adaptive switching in FatTree data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):166–179, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301839.

Zhang:2016:TAL

[1910] Yao Zhang, Xiaoyou Wang, Adrian Perrig, and Zhiming Zheng. Tumbler: Adaptable link access in the botsinfested Internet. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):180–193, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301852.

Chen:2016:DGC

[1911] Jing Chen, Kun He, Quan Yuan, Ruiying Du, Lina Wang, and Jie Wu. Distributed Greedy Coding-aware Deterministic Routing for multi-flow in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):194–206, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301803.

Frangoudis:2016:RBM

[1912] Pantelis A. Frangoudis, George C.

Polyzos, and Gerardo Rubino. Relaybased multipoint content delivery for wireless users in an information-centric network. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):207–223, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301840.

Hu:2016:PLC

[1913] Qiao Hu, Lavinia Mihaela Dinca, Anjia Yang, and Gerhard Hancke. Practical limitation of co-operative RFID jamming methods in environments without accurate signal synchronization. Computer Networks (Amsterdam, Netherlands: 1999), 105 (??):224–236, August 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630189X.

Anonymous:2016:EBl

[1914] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 105(??):ifc, August 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616302195.

Aguilar-Garcia:2016:CLB

[1915] A. Aguilar-Garcia, R. Barco, and S. Fortes. Coordinated location-based self-optimization for indoor femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 106 (??):1–16, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301906.

Akyildiz:2016:RKE

[1916] Ian F. Akyildiz, Shuai Nie, Shih-Chun Lin, and Manoj Chandrasekaran. 5G roadmap: 10 key enabling technologies. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 17–48, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301918.

Rebecchi:2016:CPC

[1917] Filippo Rebecchi, Marcelo Dias de Amorim, and Vania Conan. Circumventing plateaux in cellular data offloading using adaptive content reinjection. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 49–63, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301931.

Goussevskaia:2016:WSM

[1918] Olga Goussevskaia, Luiz F. M. Vieira, and Marcos A. M. Vieira. Wireless scheduling with multiple data rates: From physical interference to disk graphs. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 64–76, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301967.

Zhang:2016:CIF

[1919] Yuanyuan Zhang, Mingwei Xu, Ning Wang, Jun Li, Penghan Chen, and Fei Liang. Compressing IP forwarding tables with small bounded update time. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 77–90, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301955.

Xenakis:2016:AAV

[1920] Dionysis Xenakis, Nikos Passas, Lazaros Merakos, and Christos Verikoukis. ANDSF-assisted vertical handover decisions in the IEEE 802.11/LTE-Advanced network. Computer Networks (Amsterdam, Netherlands: 1999), 106(??):91–108, September 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616301888.

Lu:2016:TLS

[1921] Xiaoyan Lu, Eyuphan Bulut, and Boleslaw Szymanski. Towards limited scale-free topology with dynamic peer participation. Computer Networks (Amsterdam, Netherlands: 1999), 106 (??):109–121, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301992.

Zheng:2016:PUT

[1922] Jun Zheng, Peng Yang, Jingjing Luo, Qiuming Liu, and Li Yu. Peruser throughput analysis for secondary users in multi-hop cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 122–133, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300378.

Sehati:2016:NAL

[1923] Ali Sehati and Majid Ghaderi. Network assisted latency reduction for mobile web browsing. Computer Networks (Amsterdam, Netherlands: 1999), 106 (??):134–150, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302067.

Xu:2016:WSN

[1924] Hao Xu, Huafei Sun, Yongqiang Cheng, and Hao Liu. Wireless sensor networks localization based on graph embedding with polynomial mapping. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 151–160, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302110.

Boero:2016:BLB

[1925] L. Boero, M. Cello, C. Garibotto, M. Marchese, and M. Mongelli.

BeaQoS: Load balancing and deadline management of queues in an Open-Flow SDN switch. Computer Networks (Amsterdam, Netherlands: 1999), 106 (??):161–170, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302055.

Nguyen:2016:CEE

[1926] Minh Tuan Nguyen, Keith A. Teague, and Nazanin Rahnavard. CCS: Energy-efficient data collection in clustered wireless sensor networks utilizing block-wise compressive sensing. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 171–185, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302092.

Mangili:2016:OPV

[1927] Michele Mangili, Jocelyne Elias, Fabio Martignon, and Antonio Capone. Optimal planning of virtual content delivery networks under uncertain traffic demands. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 186–195, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302158.

Wang:2016:PTE

[1928] Ting Wang and Mounir Hamdi. Presto:

Towards efficient online virtual network embedding in virtualized cloud data centers. Computer Networks (Amsterdam, Netherlands: 1999), 106(??):

196–208, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302171.

Tso:2016:NSR

[1929] Fung Po Tso, Simon Jouet, and Dimitrios P. Pezaros. Network and server resource management strategies for data centre infrastructures: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 209–225, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302298.

Hwang:2016:JFA

[1930] Kyung-Wook Hwang, Vijay Gopalakrishnan, Rittwik Jana, Seungjoon Lee, Vishal Misra, K. K. Ramakrishnan, and Dan Rubenstein. Jointfamily: Adaptive bitrate video-on-demand streaming over peer-to-peer networks with realistic abandonment patterns. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 226–244, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301876.

Tsolkas:2016:DDL

[1931] Dimitris Tsolkas, Nikos Passas, and Lazaros Merakos. Device discovery in LTE networks: a radio access perspective. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 245–259, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302262.

Lin:2016:CTB

[1932] Shih-Chun Lin, Pu Wang, and Min Luo. Control traffic balancing in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 106(??): 260–271, September 4, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615002571.

Anonymous:2016:EBm

[1933] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 106(??):ifc, September 4, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616302572.

DAlconzo:2016:MLD

[1934] Alessandro D’Alconzo, Pere Barlet-Ros, Kensuke Fukuda, and David Choffnes. Machine learning, data mining and Big Data frameworks for network monitoring and troubleshooting. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??):1–4, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302122.

Baer:2016:DHA

[1935] Arian Baer, Pedro Casas, Alessandro D’Alconzo, Pierdomenico Fiadino, Lukasz Golab, Marco Mellia, and Erich Schikuta. DBStream: a holistic approach to large-scale network traffic monitoring and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??):5–19, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301189.

Bocchi:2016:SNM

[1936] Enrico Bocchi, Ali Safari Khatouni, Stefano Traverso, Alessandro Finamore, Maurizio Munafò, Marco Mellia, and Dario Rossi. Statistical network monitoring: Methodology and application to carrier-grade NAT. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??):20–35, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301980.

Espinet:2016:FMC

[1937] François Espinet, Diana Joumblatt, and Dario Rossi. Framework, models and controlled experiments for network troubleshooting. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??):36–54, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301815.

Grill:2016:LCA

[1938] Martin Grill and Tomás Pevný. Learning combination of anomaly detectors for security domain. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??):55– 63, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301669.

Iglesias:2016:TAF

[1939] Félix Iglesias and Tanja Zseby. Timeactivity footprints in IP traffic. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??):64– 75, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300767.

Lutu:2016:GBI

[1940] Andra Lutu, Yuba Raj Siwakoti, Özgü Alay, Dziugas Baltrunas, and Ahmed Elmokashfi. The good, the bad and the implications of profiling mobile broadband coverage. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??):76–93, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301979.

Akhtar:2016:DAS

[1941] Zahaib Akhtar, Alefiya Hussain, Ethan Katz-Bassett, and Ramesh Govindan. DBit: Assessing statistically significant differences in CDN performance. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??):94–103, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301657.

Tomanek:2016:MCL

[1942] Ondrej Tomanek, Pavol Mulinka, and Lukas Kencl. Multidimensional cloud latency monitoring and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1) (??):104–120, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630192X.

Middleton:2016:SCQ

[1943] Stuart E. Middleton and Stefano Modafferi. Scalable classification of QoS for real-time interactive applications from IP traffic measurements. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1) (??):121–132, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301165.

Zhang:2016:SCD

[1944] Qishan Zhang, Qirong Qiu, Wenzhong Guo, Kun Guo, and Naixue Xiong. A social community detection algorithm based on parallel grey label propagation. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1) (??):133–143, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301827.

Anonymous:2016:EBn

[1945] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 1)(??): ifc, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302900.

Curado:2016:GES

[1946] Marilia Curado, Ivan Ganchev, Andreas Kassler, and Yevgeni Koucheryavy. Guest editorial: Special issue on mobile wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):145– 147, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302304.

Wen:2016:ADA

[1947] Xiuxiu Wen, Guangsheng Feng, Huiqiang Wang, Hongwu Lv, and Junyu Lin. An adaptive disorderavoidance cooperative downloading method. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2) (??):148–162, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301025.

Moura:2016:EAM

[1948] Jose Moura and Christopher Edwards.

Efficient access of mobile flows to heterogeneous networks under flash crowds. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2) (??):163–177, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301086.

Helgason:2016:MOC

[1949] Ólafur Helgason, Sylvia T. Kouyoumdjieva, Ljubica Pajević, Emre A. Yavuz, and Gunnar Karlsson. A middleware for opportunistic content distribution. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2) (??):178–193, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301797.

Anastasiades:2016:ICC

[1950] Carlos Anastasiades, Tobias Schmid, Jürg Weber, and Torsten Braun. Information-centric content retrieval for delay-tolerant networks. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):194– 207, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300718.

Anastasiades:2016:DUI

[1951] Carlos Anastasiades, Jürg Weber, and Torsten Braun. Dynamic Unicast: Information-centric multi-hop routing for mobile ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):208– 219, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630072X.

Zhioua:2016:JAT

[1952] G. E. M. Zhioua, J. Zhang, H. Labiod, N. Tabbane, and S. Tabbane. A joint active time and flow selection model for cellular content retrieval through ITS. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2) (??):220–232, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630086X.

Zhou:2016:PAP

[1953] Peng Zhou, Yanheng Liu, Jian Wang, Weiwen Deng, and Heekuck Oh. Performance analysis of prioritized broadcast service in WAVE/IEEE 802.11p. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2) (??):233–245, October 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301104.

Kora:2016:ARC

[1954] Ahmed D. Kora, Brice A. Elono Ongbwa, Jean-Pierre Cances, and Vahid Meghdadi. Accurate radio coverage assessment methods investigation for 3G/4G networks. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):246– 257, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300986.

Dudin:2016:AOG

[1955] Alexander Dudin, Chesoong Kim, Sergey Dudin, and Olga Dudina. Analysis and optimization of Guard Channel Policy with buffering in cellular mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):258–269, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301013.

Silva:2016:ISD

[1956] Felipe S. Dantas Silva, Augusto Venâncio Neto, Douglas Maciel, José Castillo-Lema, Flávio Silva, Pedro Frosi, and Eduardo Cerqueira. An innovative software-defined WiNeMO architecture for advanced QoS-guaranteed mobile service transport. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):270–291, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301177.

Zola:2016:MIH

[1957] E. Zola and A. J. Kassler. Minimizing the impact of the handover for mobile users in WLAN: a study on performance optimization. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):292– 303, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300780.

Senouci:2016:TBD

[1958] Mohamed Abdelkrim Senouci, M. Sajid Mushtaq, Said Hoceini, and Abdelhamid Mellouk. TOPSIS-based dynamic approach for mobile network interface selection. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):304–314, October 9, 2016.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301098.

Jooris:2016:TCP

[1959] Bart Jooris, Jan Bauwens, Peter Ruckebusch, Peter De Valck, Christophe Van Praet, Ingrid Moerman, and Eli De Poorter. TAISC: a cross-platform MAC protocol compiler and execution engine. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):315– 326, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300974.

Ometov:2016:NSC

[1960] Aleksandr Ometov, Antonino Orsino, Leonardo Militano, Giuseppe Araniti, Dmitri Moltchanov, and Sergey Andreev. A novel security-centric framework for D2D connectivity based on spatial and social proximity. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):327– 338, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300858.

Oliveira:2016:LPL

[1961] Afonso Oliveira and Teresa Vazão.

Low-power and lossy networks under mobility: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??):339–352, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300895.

Anonymous:2016:EBo

[1962] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 107 (part 2)(??): ifc, October 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630305X.

Asheralieva:2016:ERB

[1963] Alia Asheralieva and Yoshikazu Miyanaga. Effective resource block allocation procedure for quality of service provisioning in a single-operator heterogeneous LTE-A network. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):1–14, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302419.

Taherkhani:2016:PSM

[1964] Nasrin Taherkhani and Samuel Pierre. Prioritizing and scheduling messages for congestion control in vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):15–28, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302079.

Meng:2016:OMB

[1965] Xianfu Meng and Shuang Ren. An outlier mining-based malicious node detection model for hybrid P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):29–39, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302365.

Gharbaoui:2016:ICT

[1966] Molka Gharbaoui, Barbara Martini, Carol. J. Fung, Francesco Paolucci, Alessio Giorgetti, and Piero Castoldi. An incentive-compatible and trustaware multi-provider path computation element (PCE). Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):40–54, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302407.

Pereira:2016:WNC

[1967] Carlos Pereira, Ana Aguiar, and Daniel E. Lucani. When are network coding based dynamic multihoming techniques beneficial? Computer Networks (Amsterdam, Netherlands: 1999), 108(??):55–65, October 24, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616302456.

Sciancalepore:2016:LLS

[1968] S. Sciancalepore, G. Piro, E. Vogli, G. Boggia, L. A. Grieco, and G. Cavone. LICITUS: a lightweight and standard compatible framework for securing layer-2 communications in the IoT. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):66–77, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302468.

Sharkh:2016:BCE

[1969] Mohamed Abu Sharkh, Ali Kanso, Abdallah Shami, and Peter Öhlén. Building a cloud on Earth: a study of cloud computing data center simulators. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):78–96, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630216X.

Hosseini:2016:MMP

[1970] Soodeh Hosseini and Mohammad Abdollahi Azgomi. A model for malware propagation in scale-free networks based on rumor spreading process. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 97–107, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302523.

Chen:2016:IMR

[1971] Xianda Chen, Kyung Tae Kim, and Hee Yong Youn. Integration of Markov random field with Markov chain for efficient event detection using wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):108–119, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302328.

Li:2016:ADW

[1972] Feng Li, Jun Luo, Shiqing Xin, and Ying He. Autonomous deployment of wireless sensor networks for optimal coverage with directional sensing model. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 120–132, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302493.

Sicari:2016:SPE

[1973] Sabrina Sicari, Alessandra Rizzardi, Daniele Miorandi, Cinzia Cappiello, and Alberto Coen-Porisini. Security policy enforcement for networked smart objects. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 133–147, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302663.

Tham:2016:STI

[1974] Chen-Khong Tham and Wen Sun. A spatio-temporal incentive scheme with consumer demand awareness for participatory sensing. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):148–159, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302481.

Bujari:2016:AER

[1975] Armir Bujari, Andrea Marin, Claudio E. Palazzi, and Sabina Rossi. Analysis of ECN/RED and SAP-LAW with simultaneous TCP and UDP traffic. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 160–170, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302687.

Glabowski:2016:MOS

[1976] Mariusz Glabowski, Adam Kaliszan, and Maciej Stasiak. Modelling overflow systems with distributed secondary resources. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 171–183, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302675.

Truong-Huu:2016:DEW

[1977] Tram Truong-Huu, Mohan Gurusamy, and Vishal Girisagar. Dynamic embedding of workflow requests for bandwidth efficiency in data centers. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 184–198, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302705.

Kalkan:2016:DFM

[1978] Kübra Kalkan and Fatih Alagöz. A distributed filtering mechanism against DDoS attacks: ScoreFor-Core. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 199–209, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302754.

Dimitriou:2016:IDR

[1979] Tassos Dimitriou, Ebrahim A. Alrashed, Mehmet Hakan Karaata, and Ali Hamdan. Imposter detection for replication attacks in mobile sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 210–222, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302717.

Patra:2016:PAM

[1980] Moumita Patra and C. Siva Ram Murthy. Piggybacking assisted manyto-many communication with efficient vehicle selection for improved performance in vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 223–232, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302559.

Wojcik:2016:SMP

[1981] Robert Wójcik, Jerzy Domzal, Zbigniew Duliński, Grzegorz Rzym, Andrzej Kamisiński, Piotr Gawlowicz, Piotr Jurkiewicz, Jacek Rzasa, Rafal Stankiewicz, and Krzysztof Wajda. A survey on methods to provide interdomain multipath transmissions. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 233–259, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630281X.

Chaari:2016:CPS

[1982] Rihab Chaâri, Fatma Ellouze, Anis Koubâa, Basit Qureshi, Nuno Pereira, Habib Youssef, and Eduardo Tovar. Cyber-physical systems clouds: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 260–278, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302699.

Gharbaoui:2016:CNO

[1983] M. Gharbaoui, B. Martini, D. Adami, S. Giordano, and P. Castoldi. Cloud and network orchestration in SDN data centers: Design principles and performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):279–295, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302821.

Damayanti:2016:CCM

[1984] Wayan Damayanti, Sanghyun Kim, and Ji-Hoon Yun. Collision chain mitigation and hidden device-aware grouping in large-scale IEEE 802.11ah networks. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 296–306, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302882.

Benet:2016:PET

[1985] Cristian Hernandez Benet, Andreas Kassler, and Enrica Zola. Predicting expected TCP throughput using genetic algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):307–322, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302808.

Cem:2016:ESP

[1986] Emrah Cem and Kamil Sarac. Estimation of structural properties of online social networks at the extreme. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 323–344, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302857.

Chen:2016:CNC

[1987] Siguang Chen, Chuanxin Zhao, Meng Wu, Zhixin Sun, Haijun Zhang, and Victor C. M. Leung. Compressive network coding for wireless sensor networks: Spatio-temporal coding and optimization design. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):345–356, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302973.

Plachy:2016:PSE

[1988] Jan Plachy, Zdenek Becvar, and Pavel Mach. Path selection enabling user mobility and efficient distribution of data for computation at the edge of mobile network. Computer Networks (Amsterdam, Netherlands: 1999), 108 (??):357–370, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302870.

Khalifah:2016:HFM

[1989] Afraa Khalifah, Nadine Akkari, Ghadah Aldabbagh, and Nikos Dimitriou. Hybrid femto/macro rate-based offloading for high user density networks. Computer Networks (Amsterdam, Netherlands: 1999), 108(??): 371–380, October 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302535.

Anonymous:2016:EBp

[1990] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 108(??):ifc, October 24, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616303243.

Hancke:2016:SIR

[1991] Gerhard Hancke, Aikaterini Mitrokotsa, Reihaneh Safavi-Naini, and Damien Sauveron. Special issue on recent advances in physical-layer security. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):1–3, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302547.

Quyen:2016:CDS

[1992] Nguyen Xuan Quyen, Trung Q. Duong, Nguyen-Son Vo, Qingqing Xie, and Lei Shu. Chaotic direct-sequence spreadspectrum with variable symbol period: a technique for enhancing physical layer security. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):4–12, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302031.

Nardelli:2016:TMM

[1993] Pedro H. J. Nardelli, Hirley Alves, Carlos H. M. de Lima, and Matti Latvaaho. Throughput maximization in multi-hop wireless networks under a secrecy constraint. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):13–20, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302018.

Aliberti:2016:RPS

[1994] Giulio Aliberti, Roberto Di Pietro, and Stefano Guarino. Reliable and perfectly secret communication over the generalized Ozarow–Wyner’s wire-tap channel. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1) (??):21–30, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302146.

Tippenhauer:2016:PLI

[1995] Nils Ole Tippenhauer, Kasper Bonne Rasmussen, and Srdjan Capkun.

Physical-layer integrity for wireless messages. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):31–38, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302043.

Petroulakis:2016:PDS

[1996] Nikolaos E. Petroulakis, George Spanoudakis, and Ioannis G. Askoxylakis. Patterns for the design of secure and dependable software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):39–49, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302080.

Bhunia:2016:ABN

[1997] Suman Bhunia, Vahid Behzadan, Paulo Alexandre Regis, and Shamik Sengupta. Adaptive beam nulling in multihop ad hoc networks against a jammer in motion. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):50–66, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302109.

Lee:2016:PJW

[1998] Il-Gu Lee and Myungchul Kim. Persistent jamming in wireless local area networks: Attack and defense. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):67– 83, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630202X.

Huth:2016:IRS

[1999] Christopher Huth, René Guillaume, Thomas Strohm, Paul Duplys, Irin Ann Samuel, and Tim Güneysu. Information reconciliation schemes in physicallayer security: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):84– 104, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301864.

Zenger:2016:AKE

[2000] Christian T. Zenger, Mario Pietersz, Jan Zimmer, Jan-Felix Posielek, Thorben Lenze, and Christof Paar. Authenticated key establishment for low-resource devices exploiting correlated random channels. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??):105– 123, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301943.

Anonymous:2016:EBq

[2001] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 1)(??): ifc, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303449.

Meo:2016:TPB

[2002] Michela Meo and Sabine Wittevrongel. Traffic and performance in the big data era. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2) (??):125–126, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630233X.

Marchetti:2016:AHV

[2003] Mirco Marchetti, Fabio Pierazzi, Michele Colajanni, and Alessandro Guido. Analysis of high volumes of network traffic for Advanced Persistent Threat detection. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??):127– 141, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616301633.

Bocchi:2016:MNB

[2004] Enrico Bocchi, Luigi Grimaudo, Marco Mellia, Elena Baralis, Sabyasachi Saha, Stanislav Miskovic, Gaspar Modelo-Howard, and Sung-Ju Lee. MAGMA network behavior classifier for malware traffic. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2) (??):142–156, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300949.

Somani:2016:DAC

[2005] Gaurav Somani, Manoj Singh Gaur, Dheeraj Sanghi, and Mauro Conti.

DDoS attacks in cloud computing: Collateral damage to non-targets. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??):157– 171, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300901.

Ha:2016:STS

[2006] Taejin Ha, Sunghwan Kim, Namwon An, Jargalsaikhan Narantuya, Chiwook Jeong, JongWon Kim, and Hyuk Lim. Suspicious traffic sampling for intrusion detection in software-defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??):172–182, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616301645.

Goncalves:2016:WMP

[2007] Glauber D. Gonçalves, Idilio Drago, Alex B. Vieira, Ana Paula Couto da Silva, Jussara M. Almeida, and Marco Mellia. Workload models and performance evaluation of cloud storage services. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??):183–199, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300950.

Hours:2016:SID

[2008] Hadrien Hours, Ernst Biersack, Patrick Loiseau, Alessandro Finamore, and Marco Mellia. A study of the impact of DNS resolvers on CDN performance using a causal approach. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??):200– 210, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302006.

Wamser:2016:MYS

[2009] Florian Wamser, Pedro Casas, Michael Seufert, Christian Moldovan, Phuoc Tran-Gia, and Tobias Hossfeld. Modeling the YouTube stack: From packets to quality of experience. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??):211– 224, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300925.

Bonald:2016:MSM

[2010] T. Bonald and C. Comte. The multisource model for dimensioning data networks. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??):225–233, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616300913.

Kleinrouweler:2016:MES

[2011] Jan Willem Kleinrouweler, Sergio Cabrero, Rob van der Mei, and Pablo Cesar. A model for evaluating sharing policies for network-assisted HTTP adaptive streaming. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??):234– 245, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616300937.

Metzger:2016:TVS

[2012] Florian Metzger, Eirini Liotou, Christian Moldovan, and Tobias Hoßfeld. TCP video streaming and mobile networks: Not a love story, but better with context. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2) (??):246–256, November 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302134.

Anonymous:2016:EBr

[2013] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??): ifc, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303553.

Anonymous:2016:EBs

[2014] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 109 (part 2)(??): ifc, November 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303553.

Wang:2016:MCC

[2015] Jin Wang, Jing Ren, Kejie Lu, Jianping Wang, Shucheng Liu, and Cedric Westphal. A minimum cost cache management framework for informationcentric networks with network coding. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 1–17, December 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302444.

Trang:2016:FIF

[2016] Si Quoc Viet Trang and Emmanuel Lochin. FLOWER, an innovative Fuzzy Lower-than-Best-Effort transport protocol. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):18–30, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302985.

Chen:2016:MLS

[2017] Meng Chen, Mingwei Xu, Qing Li, and Yuan Yang. Measurement of largescale BGP events: Definition, detection, and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):31–45, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303152.

Fan:2016:GLA

[2018] Yuqi Fan, Hongli Ding, Lusheng Wang, and Xiaojing Yuan. Green latencyaware data placement in data centers. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 46–57, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303139.

Wang:2016:TMC

[2019] Yu Wang, Mehul Motani, Hari Krishna Garg, Xin Kang, and Qian Chen. Throughput maximization for cooperative 60 GHz wireless personal area networks. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):58–68, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303000.

Guo:2016:MSE

[2020] Yi Guo, Haixin Duan, Jikun Chen, and Fu Miao. MAF-SAM: an effective method to perceive data plane threats of inter domain routing system. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 69–78, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303164.

Kim:2016:CHT

[2021] Sungwook Kim. Cognitive hierarchy thinking based behavioral game model for IoT power control algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):79–90, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630319X.

Mao:2016:EMP

[2022] Jianbiao Mao, Biao Han, Zhigang Sun, Xicheng Lu, and Ziwen Zhang. Efficient mismatched packet buffer management with packet order-preserving for Open-Flow networks. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):91–103, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303140.

Carofiglio:2016:OMC

[2023] Giovanna Carofiglio, Massimo Gallo, and Luca Muscariello. Optimal multipath congestion control and request forwarding in information-centric networks: Protocol design and experimentation. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 104–117, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303024.

deSouza:2016:IIT

[2024] Allan M. de Souza, R. S. Yokoyama, Azzedine Boukerche, Guilherme Maia, Eduardo Cerqueira, Antonio A. F. Loureiro, and Leandro Aparecido Villas. ICARUS: Improvement of traffic Condition through an Alerting and Rerouting System. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):118–132, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303012.

Carofiglio:2016:JFC

[2025] Giovanna Carofiglio, Léonce Mekinda, and Luca Muscariello. Joint forwarding and caching with latency awareness in information-centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 133–153, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303176.

Ramos:2016:DLS

[2026] Heitor S. Ramos, Azzedine Boukerche, Alyson L. C. Oliveira, Alejandro C. Frery, Eduardo M. R. Oliveira, and Antonio A. F. Loureiro. On the deployment of large-scale wireless sensor networks considering the energy hole problem. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):154–167, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303036.

Ahmad:2016:QCS

[2027] Arslan Ahmad, Alessandro Floris, and Luigi Atzori. QoE-centric service delivery: a collaborative approach among OTTs and ISPs. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):168–179, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303218.

Mandhare:2016:QRE

[2028] V. V. Mandhare, V. R. Thool, and R. R. Manthalkar. QoS routing enhancement using metaheuristic approach in mobile ad-hoc network. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 180–191, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630322X.

Narman:2016:JSP

[2029] Husnu S. Narman, Mohammed Atiquzzaman, Mehdi Rahmani-andebili, and Haiying Shen. Joint and selective periodic component carrier assignment for LTE-A. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 192–205, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303310.

Maleki:2016:FTI

[2030] Esmaeil Nik Maleki and Ghasem Mirjalily. Fault-tolerant interferenceaware topology control in multiradio multi-channel wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 206–222, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302791.

Valcarenghi:2016:DFR

[2031] Luca Valcarenghi, Koteswararao Kondepu, and Piero Castoldi. Delay fairness in reconfigurable and energy efficient TWDM PON. Computer Networks (Amsterdam, Netherlands: 1999), 110(??):223–231, December 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616302845.

Misra:2016:EET

[2032] Sudip Misra, Samaresh Bera, Tamoghna Ojha, Hussein T. Mouftah, and Alagan Anpalagan. EN-TRUST: Energy trading under uncertainty in smart grid systems. Computer Networks (Amsterdam, Netherlands: 1999), 110(??):232– 242, December 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303206.

Li:2016:ERM

[2033] Xin Li, Haotian Wu, Don Gruenbacher, Caterina Scoglio, and Tricha Anjali. Efficient routing for middlebox policy enforcement in software-defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 243–252, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303346.

Lee:2016:TWT

[2034] Jinho Lee, Jaehoon (Paul) Jeong, and David H. C. Du. Two-way traffic link delay modeling in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 253–265, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303358.

Xu:2016:DEC

[2035] Yuemei Xu, Song Ci, Yang Li, Tao Lin, and Gang Li. Design and evaluation of coordinated in-network caching model for content centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 266–283, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630336X.

Chowdhury:2016:ALT

[2036] Tashnim J. S. Chowdhury, Colin Elkin, Vijay Devabhaktuni, Danda B. Rawat, and Jared Oluoch. Advances on localization techniques for wireless sensor networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 284–305, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303383.

Vasilakos:2016:AND

[2037] Xenofon Vasilakos, Vasilios A. Siris, and George C. Polyzos. Addressing niche demand based on joint mobility prediction and content popularity caching. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 306–323, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303334.

Xie:2016:PSH

[2038] Kun Xie, Jiannong Cao, Xin Wang, and Jigang Wen. Pre-scheduled handoff for service-aware and seamless Internet access. Computer Networks (Amsterdam, Netherlands: 1999), 110 (??):324–337, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303371.

Tien:2016:NRB

[2039] Nguyen Xuan Tien, Semog Kim, and Jong Myung Rhee. A novel ringbased dual paths approach for reducing redundant traffic in HSR networks. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 338–350, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303413.

Ye:2016:RER

[2040] Run Ye, Azzedine Boukerche, Houjun Wang, Xiaojia Zhou, and Bin Yan. RECODAN: an efficient redundancy coding-based data transmission scheme for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 351–363, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303425.

Xu:2016:SSI

[2041] Yang Xu, Yong Liu, Rahul Singh, and Shu Tao. SDN state inconsistency verification in openstack. Computer Networks (Amsterdam, Netherlands: 1999), 110(??): 364–376, December 9, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303127.

Anonymous:2016:EBt

[2042] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 110(??):ifc, December 9, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128616303759.

Socievole:2016:CPS

[2043] A. Socievole, A. Ziviani, F. De Rango, A. V. Vasilakos, and E. Yoneki. Cyber-physical systems for Mobile Opportunistic Networking in Proximity (MNP). Computer Networks (Amsterdam, Netherlands: 1999), 111 (??):1–5, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302997.

Machado:2016:PFM

[2044] Kassio Machado, Azzedine Boukerche, Pedro O. S. Vaz de Melo, Eduardo Cerqueira, and Antonio A. F. Loureiro. Pervasive forwarding mechanism for mobile social networks. Computer Networks (Amsterdam, Netherlands: 1999), 111 (??):6–16, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616302730.

Miao:2016:PPR

[2045] Jingwei Miao, Omar Hasan, Sonia Ben Mokhtar, Lionel Brunie, and Ammar Hasan. 4PR: Privacy preserving routing in mobile delay tolerant networks. Computer Networks (Amsterdam, Netherlands: 1999), 111(??): 17–28, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630247X.

Ramiro:2016:CAT

[2046] Victor Ramiro, Emmanuel Lochin, and Patrick Sénac. Characterization and applications of temporal random walks on opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 111(??): 29–44, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302390.

Hernandez-Orallo:2016:AEP

[2047] Enrique Hernández-Orallo, Marina Murillo-Arcila, Carlos T. Calafate, Juan Carlos Cano, J. Alberto Conejero, and Pietro Manzoni. Analytical evaluation of the performance of contact-based messaging applications. Computer Networks (Amsterdam, Netherlands: 1999), 111(??): 45–54, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302353.

Rolim:2016:SAC

[2048] Carlos Oberdan Rolim, Anubis G. Rossetto, Valderi R. Q. Leithardt, Guilherme A. Borges, Cláudio F. R. Geyer, Tatiana F. M. dos Santos, and Adriano M. Souza. Situation awareness and computational intelligence in opportunistic networks to support the data transmission of urban sensing applications. Computer Networks (Amsterdam, Netherlands: 1999), 111(??): 55–70, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302420.

Chen:2016:EMA

[2049] Xiao Chen, Charles Shang, Britney Wong, Wenzhong Li, and Suho Oh. Efficient multicast algorithms in opportunistic mobile social networks using community and social features. Computer Networks (Amsterdam, Netherlands: 1999), 111(??): 71–81, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302341.

Holzer:2016:PES

[2050] Adrian Holzer, Sven Reber, Jonny Quarta, Jorge Mazuze, and Denis Gillet. Padoc: Enabling social networking in proximity. Computer Networks (Amsterdam, Netherlands: 1999), 111 (??):82–92, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630250X.

Turkes:2016:CLO

[2051] Okan Turkes, Hans Scholten, and Paul J. M. Havinga. Cocoon: a lightweight opportunistic networking middleware for community-oriented smart mobile applications. Computer Networks (Amsterdam, Netherlands: 1999), 111(??): 93–108, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302742.

Killijian:2016:SSO

[2052] Marc-Olivier Killijian, Roberto Pasqua, Matthieu Roy, Gilles Trédan, and Christophe Zanon. Souk: Spatial Observation of Human Kinetics. Computer Networks (Amsterdam, Netherlands: 1999), 111(??): 109–119, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302511.

Ojog:2016:MCO

[2053] Cristian-Octavian Ojog, Radu-Corneliu Marin, Radu-Ioan Ciobanu, and Ciprian Dobre. Multi-criteria optimization of wireless connectivity over sparse networks. Computer Networks (Amsterdam, Netherlands: 1999), 111 (??):120–128, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302389.

Oliveira:2016:SSN

[2054] Thiago Rodrigues Oliveira, Cristiano M. Silva, Daniel F. Macedo, and José Marcos S. Nogueira. SNVC: Social Networks for Vehicular Certification. Computer Networks (Amsterdam, Netherlands: 1999), 111(??): 129–140, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302778.

Militano:2016:TBS

[2055] L. Militano, A. Orsino, G. Araniti, M. Nitti, L. Atzori, and A. Iera. Trustbased and social-aware coalition formation game for multihop data uploading in 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 111 (??):141–151, December 24, 2016. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302432.

Anonymous:2016:EBu

[2056] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 111(??):ifc, December 24, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616303917.

Anonymous:2017:ECN

[2057] Anonymous. Editorial for Computer Networks Journal 2016. Computer Networks (Amsterdam, Netherlands: 1999), 112(??):iii, January 15, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304364.

Mauri:2017:DKR

[2058] Giulia Mauri and Giacomo Verticale.

Up-to-date key retrieval for information centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 112(??):1–11, January 15, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303668.

Zhang:2017:MCS

[2059] Yi Zhang, Michela Meo, Raffaella Gerboni, and Marco Ajmone Marsan. Minimum cost solar power systems for LTE macro base stations. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):12–23, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303401.

Liao:2017:DCB

[2060] Jianxin Liao, Haifeng Sun, Jingyu Wang, Qi Qi, Kai Li, and Tonghong Li. Density cluster based approach for controller placement problem in large-scale software defined networkings. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 24–35, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303620.

Ferraz:2017:HPT

[2061] Lyno Henrique G. Ferraz, Rafael Laufer, Diogo M. F. Mattos, Otto Carlos M. B. Duarte, and Guy Pujolle. A high-performance Two-Phase Multipath scheme for data-center networks. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):36–51, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303188.

Ogino:2017:VNE

[2062] Nagao Ogino, Takeshi Kitahara, Shin’ichi Arakawa, and Masayuki Murata. Virtual network embedding with multiple priority classes sharing substrate resources. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):52–66, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303395.

Persico:2017:PWA

[2063] Valerio Persico, Alessio Botta, Pietro Marchetta, Antonio Montieri, and Antonio Pescapé. On the performance of the wide-area networks interconnecting public-cloud datacenters around the globe. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):67–83, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630353X.

Zheng:2017:TEC

[2064] Meng Zheng, Chi Xu, Wei Liang, Haibin Yu, and Lin Chen. Timeefficient cooperative spectrum sensing via analog computation over multipleaccess channel. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):84–94, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303632.

Pan:2017:LDG

[2065] Meng-Shiuan Pan and Shu-Wei Yang.

A lightweight and distributed geographic multicast routing protocol for IoT applications. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):95–107, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303711.

Salman:2017:PFR

[2066] Mustafa Ismael Salman, Muntadher Qasim Abdulhasan, Chee Kyun Ng, Nor Kamariah Noordin, Borhanuddin Mohd Ali, and Aduwati Sali. A partial feedback reporting scheme for LTE mobile video transmission with QoS provisioning. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):108–121, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302869.

Abdelsalam:2017:TWN

[2067] Ahmed Abdelsalam, Michele Luglio, Cesare Roseti, and Francesco Zampognaro. TCP Wave: a new reliable transport approach for future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 122–143, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303681.

Ndashimye:2017:VIC

[2068] Emmanuel Ndashimye, Sayan K. Ray, Nurul I. Sarkar, and Jairo A. Gutiérrez. Vehicle-to-infrastructure communication over multi-tier heterogeneous networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):144–166, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303826.

Yao:2017:PFF

[2069] Jinfa Yao, Baoqun Yin, Xiaobin Tan, and Xiaofeng Jiang. A POMDP framework for forwarding mechanism in named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 167–175, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303723.

Oliveira:2017:MDT

[2070] Eduardo Mucelli Rezende Oliveira, Aline Carneiro Viana, K. P. Naveen, and Carlos Sarraute. Mobile data traffic modeling: Revealing temporal facets. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 176–193, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303644.

Xu:2017:OCV

[2071] Cong Xu, Jiahai Yang, Kevin Yin, and Hui Yu. Optimal construction of virtual networks for Cloud-based MapReduce workflows. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):194–207, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630367X.

Adasme:2017:MCD

[2072] Pablo Adasme, Rafael Andrade, and Abdel Lisser. Minimum cost dominating tree sensor networks under probabilistic constraints. Computer Networks (Amsterdam, Netherlands: 1999), 112(??):208–222, January 15, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616303875.

Ren:2017:EEV

[2073] Cheng Ren, Sheng Wang, Jing Ren, Weizhong Qian, Xiaoning Zhang, and Jie Duan. Energy-efficient virtual topology design in IP over WDM mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 223–236, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304108.

Ouaddah:2017:ACI

[2074] Aafaf Ouaddah, Hajar Mousannif, Anas Abou Elkalam, and Abdellah Ait Ouahman. Access control in the Internet of Things: Big challenges and new opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):237–262, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303735.

Xia:2017:OCM

[2075] Qiufen Xia, Weifa Liang, and Zichuan Xu. The operational cost minimization in distributed clouds via communityaware user data placements of social networks. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 263–278, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303887.

Karakus:2017:SCP

[2076] Murat Karakus and Arjan Durresi. A survey: Control plane scalability issues and approaches in Software-Defined Networking (SDN). Computer Networks (Amsterdam, Netherlands: 1999), 112(??):279– 293, January 15, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630411X.

Bostanipour:2017:NDA

[2077] Behnaz Bostanipour and Benoît Garbinato. A neighbor detection algorithm based on multiple virtual mobile nodes for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 294–313, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630370X.

Wang:2017:ALE

[2078] Zhibo Wang, Honglong Chen, Qing Cao, Hairong Qi, Zhi Wang, and Qian Wang. Achieving location error tolerant barrier coverage for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 112 (??):314–328, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630408X.

Chang:2017:MCC

[2079] Ben-Jye Chang and Gunag-Jie Jhang.

Minimizing contention collision probability and guaranteeing packet delay for cloud big data transmissions in 4G LTE-A packet random access. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 329–344, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303656.

Ghazisaeedi:2017:GGM

[2080] Ebrahim Ghazisaeedi and Changcheng Huang. GreenMap: Green mapping of MapReduce-based virtual networks onto a data center network and managing incast queueing delay. Computer Networks (Amsterdam, Netherlands: 1999), 112(??): 345–359, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304091.

Liu:2017:CSS

[2081] Shaowei Liu, Weimin Lei, Wei Zhang, and Yunchong Guan. CMT– SR: a selective retransmission based concurrent multipath transmission mechanism for conversational video. Computer Networks (Amsterdam, Netherlands: 1999), 112(??):360– 371, January 15, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304121.

Anonymous:2017:EBa

[2082] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 112(??):ifc, January 15, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128616304315.

Zuo:2017:FTB

[2083] Liudong Zuo, Michelle M. Zhu, Chase Q. Wu, and Jason Zurawski. Fault-tolerant bandwidth reservation strategies for data transfers in highperformance networks. Computer Networks (Amsterdam, Netherlands: 1999), 113(??):1–16, February 11, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128616303693.

Tuysuz:2017:EAN

[2084] Mehmet Fatih Tuysuz and Murat Uçan. Energy-aware network/ interface selection and handover application for Android-based mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 17–28, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304157.

Erickson:2017:STI

[2085] Alejandro Erickson, Iain A. Stewart, Javier Navaridas, and Abbas E. Kiasari. The stellar transformation: From interconnection networks to datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 113 (??):29–45, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304169.

Yoro:2017:SOS

[2086] Wilfried Yoro, Tijani Chahed, Mamdouh El-Tabach, Taoufik En-Najjary, and Azeddine Gati. Service-oriented sharing of energy in wireless access networks using Shapley value. Computer Networks (Amsterdam, Netherlands: 1999), 113(??):46– 57, February 11, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630384X.

Deltouzos:2017:TEB

[2087] Konstantinos Deltouzos and Spyros Denazis. Tackling energy and battery issues in mobile P2P VoD systems. Computer Networks (Amsterdam, Netherlands: 1999), 113(??):

58–71, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304170.

Yan:2017:ONC

[2088] Yan Yan, Baoxian Zhang, and Cheng Li. Opportunistic network coding based cooperative retransmissions in D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 72–83, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304194.

Itani:2017:DSN

[2089] M. Itani, S. Sharafeddine, and I. Elkabani. Dynamic single node failure recovery in distributed storage systems. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 84–93, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304200.

Panos:2017:AQD

[2090] Christoforos Panos, Christoforos Ntantogian, Stefanos Malliaros, and Christos Xenakis. Analyzing, quantifying, and detecting the blackhole attack in infrastructure-less networks. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 94–110, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304212.

Leu:2017:EES

[2091] Jenq-Shiou Leu, Min-Chieh Yu, Chun-Yao Liu, Alrezza Pradanta Bagus Budiarsa, and Vincent Utomo. Energy efficient streaming for smartphone by video adaptation and backlight control. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 111–123, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304224.

Barmpounakis:2017:CAU

[2092] Sokratis Barmpounakis, Alexandros Kaloxylos, Panagiotis Spapis, and Nancy Alonistioti. Context-aware, user-driven, network-controlled RAT selection for 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 113(??):124–147, February 11, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128616304236.

Banerjee:2017:CTR

[2093] Bitan Banerjee, Anand Seetharam, Amitava Mukherjee, and Mrinal Kanti Naskar. Characteristic time routing in information centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 148–158, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304248.

Araujo:2017:RLM

[2094] João Taveira Araujo, Raul Landa, Richard G. Clegg, George Pavlou, and Kensuke Fukuda. On rate limitation mechanisms for TCP throughput: a longitudinal analysis. Computer Networks (Amsterdam, Netherlands: 1999), 113(??):159–175, February 11, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128616304182.

Lai:2017:USN

[2095] Jeffrey Lai, Qiang Fu, and Tim Moors. Using SDN and NFV to enhance request rerouting in ISP-CDN collaborations. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 176–187, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630425X.

Tuysuz:2017:SEE

[2096] Mehmet Fatih Tuysuz, Zekiye Kubra Ankarali, and Didem Gözüpek. A survey on energy efficiency in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 113 (??):188–204, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304273.

Stein:2017:MRS

[2097] Joshua Stein, Han Hee Song, Mario Baldi, and Jun Li. On the most representative summaries of network user activities. Computer Networks (Amsterdam, Netherlands: 1999), 113(??):

205–217, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304297.

Javed:2017:SMT

[2098] Uzzam Javed, Azeem Iqbal, Saad Saleh, Syed Ali Haider, and Muhammad U. Ilyas. A stochastic model for transit latency in OpenFlow SDNs. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 218–229, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304406.

Shi:2017:SBR

[2099] Junling Shi, Xingwei Wang, Min Huang, Keqin Li, and Sajal K. Das. Social-based routing scheme for fixedline VANET. Computer Networks (Amsterdam, Netherlands: 1999), 113 (??):230–243, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304418.

Forkan:2017:VLM

[2100] Abdur Rahim Mohammad Forkan, Ibrahim Khalil, and Mohammed Atiquzzaman. ViSiBiD: a learning model for early discovery and real-time prediction of severe clinical events using vital signs as big data. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 244–257, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304431.

Vega:2017:MGH

[2101] Carlos Vega, Paula Roquero, and Javier Aracil. Multi-Gbps HTTP traffic analysis in commodity hardware based on local knowledge of TCP streams. Computer Networks (Amsterdam, Netherlands: 1999), 113(??): 258–268, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300014.

Gharib:2017:FDE

[2102] Mohammed Gharib, Zahra Moradlou, Mohammed Ali Doostari, and Ali Movaghar. Fully distributed ECCbased key management for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 113 (??):269–283, February 11, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861630442X.

Anonymous:2017:EBb

[2103] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 113(??):ifc, February 11, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128617300117.

Wei:2017:CBH

[2104] Xuetao Wei, Nicholas C. Valler, Harsha V. Madhyastha, Iulian Neamtiu, and Michalis Faloutsos. Characterizing the behavior of handheld devices and its implications. Computer Networks (Amsterdam, Netherlands: 1999), 114 (??):1–12, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300038.

Gazzah:2017:TBN

[2105] Leïla Gazzah and Leïla Najjar. Tradeoff between node selection and space diversity for accurate uncooperative localization. Computer Networks (Amsterdam, Netherlands: 1999), 114 (??):13–22, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730004X.

Moreno:2017:TES

[2106] Eduardo Moreno, Alejandra Beghelli, and Filippo Cugini. Traffic engineering in segment routing networks. Computer Networks (Amsterdam, Netherlands: 1999), 114(??): 23–31, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300063.

Liyanage:2017:SCC

[2107] Madhusanka Liyanage, An Braeken, Anca Delia Jurcut, Mika Ylianttila, and Andrei Gurtov. Secure communication channel architecture for Software Defined Mobile Networks. Computer Networks (Amsterdam, Netherlands: 1999), 114(??): 32–50, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300075.

Mohemed:2017:EER

[2108] Reem E. Mohemed, Ahmed I. Saleh, Maher Abdelrazzak, and Ahmed S. Samra. Energy-efficient routing protocols for solving energy hole problem in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 114(??): 51–66, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304261.

Kang:2017:VEO

[2109] Xin Kang and Jing Yang. Viewing experience optimization for peerto-peer streaming networks with credit-based incentive mechanisms. Computer Networks (Amsterdam, Netherlands: 1999), 114(??):67– 79, February 26, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300051.

Liu:2017:PCD

[2110] Linfeng Liu, Ping Wang, and Ran Wang. Propagation control of data forwarding in opportunistic underwater sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 114 (??):80–94, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861730018X.

Khebbache:2017:VNF

[2111] Selma Khebbache, Makhlouf Hadji, and Djamal Zeghlache. Virtualized network functions chaining and routing algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 114 (??):95–110, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300087.

Klier:2017:SIS

[2112] Julia Klier, Mathias Klier, Alexander Richter, and Rolf T. Wigand. Special issue on social media networks in business. Computer Networks (Amsterdam, Netherlands: 1999), 114(??): 111–113, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303528.

Li:2017:PSR

[2113] Libo Li, Frank Goethals, Bart Baesens, and Monique Snoeck. Predicting software revision outcomes on GitHub using structural holes theory. Computer Networks (Amsterdam, Netherlands: 1999), 114(??): 114–124, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302766.

Wehner:2017:ESN

[2114] Benjamin Wehner, Christian Ritter, and Susanne Leist. Enterprise social networks: a literature review and research agenda. Computer Networks (Amsterdam, Netherlands: 1999), 114 (??):125–142, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302833.

Brooks:2017:SMI

[2115] Stoney Brooks and Christopher Califf. Social media-induced technostress: Its impact on the job performance of it professionals and the moderating role of job characteristics. Computer Networks (Amsterdam, Netherlands: 1999), 114(??): 143–153, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616302729.

Wang:2017:TLS

[2116] Haishuai Wang, Jia Wu, Shirui Pan, Peng Zhang, and Ling Chen. Towards large-scale social networks with online diffusion provenance detection. Computer Networks (Amsterdam, Netherlands: 1999), 114(??): 154–166, February 26, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861630278X.

Anonymous:2017:EBc

[2117] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 114(??):ifc, February 26, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617300312.

Sinha:2017:DLM

[2118] Abhishek Sinha, Pradeepkumar Mani, Jie Liu, Ashley Flavel, and Dave Maltz. Distributed load management algorithms in anycast-based CDNs. Computer Networks (Amsterdam, Netherlands: 1999), 115 (??):1–15, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300282.

Chang:2017:ACC

[2119] Ben-Jye Chang and Syuan-Hong Liou.

Adaptive cooperative communication for maximizing reliability and reward in ultra-dense small cells LTE-A toward 5G cellular networking. Computer Networks (Amsterdam, Netherlands: 1999), 115(??):16–28, March 14, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300269.

DOro:2017:ABR

[2120] Salvatore D’Oro, Laura Galluccio, Panayotis Mertikopoulos, Giacomo Morabito, and Sergio Palazzo. Auctionbased resource allocation in Open-Flow multi-tenant networks. Computer Networks (Amsterdam, Netherlands: 1999), 115(??):29–41, March 14, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300099.

Wang:2017:EAR

[2121] Rui Wang, Suixiang Gao, Wenguo Yang, and Zhipeng Jiang. Energy aware routing with link disjoint backup paths. Computer Networks (Amsterdam, Netherlands: 1999), 115 (??):42–53, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300270.

Park:2017:MMO

[2122] Sangkyu Park, Hyunjoong Lee, Chan-Byoung Chae, and Saewoong Bahk. Massive MIMO operation in partially centralized cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 115 (??):54–64, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300257.

Schilcher:2017:QIS

[2123] Udo Schilcher, Günther Brandner, and Christian Bettstetter. Quantifying inhomogeneity of spatial point patterns. Computer Networks (Amsterdam, Netherlands: 1999), 115 (??):65–81, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304443.

Tsai:2017:NDA

[2124] Tsung-Yu Tsai, Yi-Hsueh Tsai, Zsehong Tsai, and Shiann-Tsong Sheu. A Novel Description approach based on sorted rectangles for scheduling information bearing in OFDMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 115 (??):82–99, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300294.

Wang:2017:TEE

[2125] Jing Wang, Jian Tang, Guoliang Xue, and Dejun Yang. Towards energyefficient task scheduling on smartphones in mobile crowd sensing systems. Computer Networks (Amsterdam, Netherlands: 1999), 115(??): 100–109, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304145.

Liu:2017:ICN

[2126] Wai-Xi Liu, Shun-Zheng Yu, Guang Tan, and Jun Cai. Information-centric networking with built-in network coding to achieve multisource transmission at network-layer. Computer Networks (Amsterdam, Netherlands: 1999), 115 (??):110–128, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128615001656.

Soto:2017:CFR

[2127] Julio Soto and Michele Nogueira. Corrigendum to “A framework for resilient and secure spectrum sensing on cognitive radio networks” [Computer Networks volume 79 (2015) 313–322]. Computer Networks (Amsterdam, Netherlands: 1999), 115 (??):129, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/

                                        article/pii/S1389128617300233. See

[1485].

Soto:2017:FRS

[2128] Julio Soto and Michele Nogueira. A framework for resilient and secure spectrum sensing on cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 115 (??):130–138, March 14, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300245.

Anonymous:2017:EBd

[2129] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 115(??):ifc, March 14, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S138912861730049X.

Coudron:2017:IMT

[2130] Matthieu Coudron and Stefano Secci. An implementation of multipath TCP in ns3. Computer Networks (Amsterdam, Netherlands: 1999), 116 (??):1–11, April 7, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300403.

MacFarland:2017:BBB

[2131] Douglas C. MacFarland, Craig A. Shue, and Andrew J. Kalafut. The best bang for the byte: Characterizing the potential of DNS amplification attacks. Computer Networks (Amsterdam, Netherlands: 1999), 116 (??):12–21, April 7, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300452.

Ciavarrini:2017:SBG

[2132] Gloria Ciavarrini, Valerio Luconi, and Alessio Vecchio. Smartphone-based geolocation of Internet hosts. Computer Networks (Amsterdam, Netherlands: 1999), 116(??):22–32, April 7, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300440.

DiPietro:2017:EBC

[2133] Roberto Di Pietro and Gabriele Oligeri.

Enabling broadcast communications in presence of jamming via probabilistic pairing. Computer Networks (Amsterdam, Netherlands: 1999), 116 (??):33–46, April 7, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300543.

Mohan:2017:FTT

[2134] Purnima Murali Mohan, Tram Truong-Huu, and Mohan Gurusamy. Fault tolerance in TCAM-limited software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 116 (??):47–62, April 7, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300476.

Sun:2017:COV

[2135] Geng Sun, Yanheng Liu, Ming Yang, Aimin Wang, Shuang Liang, and Ying Zhang. Coverage optimization of VLC in smart homes based on improved cuckoo search algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 116(??):63–78, April 7, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300580.

Anjum:2017:SPA

[2136] Nasreen Anjum, Dmytro Karamshuk, Mohammad Shikh-Bahaei, and Nishanth Sastry. Survey on peer-assisted content delivery networks. Computer Networks (Amsterdam, Netherlands: 1999), 116(??):79–95, April 7, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300464.

Behal:2017:DDA

[2137] Sunny Behal and Krishan Kumar. Detection of DDoS attacks and flash events using novel information theory metrics. Computer Networks (Amsterdam, Netherlands: 1999), 116 (??):96–110, April 7, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300592.

Koubaa:2017:QAE

[2138] Mohamed Koubàa, Maroua Bakri, Ammar Bouallègue, and Maurice Gagnaire. QoT-aware elastic bandwidth allocation and spare capacity assignment in flexible island-based optical transport networks under shared risk link group constraints. Computer Networks (Amsterdam, Netherlands: 1999), 116 (??):111–140, April 7, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300439.

Anonymous:2017:EBe

[2139] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 116(??):ifc, April 7, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730066X.

Malekian:2017:CPS

[2140] Reza Malekian, Kui Wu, Gianluca Reali, Ning Ye, and Kevin Curran. Cyberphysical systems and context-aware sensing and computing. Computer Networks (Amsterdam, Netherlands: 1999), 117(??):1–4, April 22, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300567.

Kuo:2017:QIA

[2141] Shu-Yu Kuo, Yao-Hsin Chou, and Chi-Yuan Chen. Quantum-inspired algorithm for cyber-physical visual surveillance deployment systems. Computer Networks (Amsterdam, Netherlands: 1999), 117(??):5–18, April 22, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616303899.

Han:2017:NDD

[2142] Wenlin Han and Yang Xiao. A novel detector to detect colluded non-technical loss frauds in smart grid. Computer Networks (Amsterdam, Netherlands: 1999), 117(??):19–31, April 22, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303516.

Zhang:2017:PAD

[2143] Xu Zhang, Mi Wen, Kejie Lu, and Jingsheng Lei. A privacy-aware data dissemination scheme for smart grid with abnormal data traceability. Computer Networks (Amsterdam, Netherlands: 1999), 117(??):32–41, April 22, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616303838.

Tang:2017:OAE

[2144] Guoming Tang, Zhen Ling, Fengyong Li, Daquan Tang, and Jiuyang Tang. Occupancy-aided energy disaggregation. Computer Networks (Amsterdam, Netherlands: 1999), 117 (??):42–51, April 22, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128616304133.

Rodriguez-Lozano:2017:CAP

[2145] David Rodriguez-Lozano, Juan A. Gomez-Pulido, Jose M. Lanza-Gutierrez, Arturo Duran-Dominguez, and Ramon A. Fernandez-Diaz. Contextaware prediction of access points demand in Wi-Fi networks. Computer Networks (Amsterdam, Netherlands: 1999), 117(??):52–61, April 22, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300026.

Abuarqoub:2017:DCM

[2146] Abdelrahman Abuarqoub, Mohammad Hammoudeh, Bamidele Adebisi, Sohail Jabbar, Ahcène Bounceur, and Hashem Al-Bashar. Dynamic clustering and management of mobile wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 117 (??):62–75, April 22, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300397.

Li:2017:RVT

[2147] Fuliang Li and Zhihan Lv. Reliable vehicle type recognition based on information fusion in multiple sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 117 (??):76–84, April 22, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300579.

Anonymous:2017:EBf

[2148] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 117(??):ifc, April 22, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617300944.

Fontinele:2017:EIR

[2149] Alexandre Fontinele, Iallen Santos, Juarez Nolêto Neto, Divanilson R. Campelo, and André Soares. An efficient IA-RMLSA algorithm for transparent elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 118 (??):1–14, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300634.

Yadav:2017:EDE

[2150] Ajay Kumar Yadav, Santosh Kumar Das, and Sachin Tripathi. EFMMRP: Design of efficient fuzzy based multiconstraint multicast routing protocol for wireless ad-hoc network. Computer Networks (Amsterdam, Netherlands: 1999), 118(??):15–23, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300610.

Das:2017:PAC

[2151] Aveek K. Das, Parth H. Pathak, Chen-Nee Chuah, and Prasant Mohapatra. Privacy-aware contextual localization using network traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 118(??): 24–36, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300555.

Divakaran:2017:RRE

[2152] Dinil Mon Divakaran, Li Ling Ko, Le Su, and Vrizlynn L. L. Thing. REX:

Resilient and efficient data structure for tracking network flows. Computer Networks (Amsterdam, Netherlands: 1999), 118(??):37–53, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300646.

Sun:2017:RRF

[2153] Penghao Sun, Julong Lan, Peng Wang, and Teng Ma. RFC: Range feature code for TCAM-based packet classification. Computer Networks (Amsterdam, Netherlands: 1999), 118 (??):54–61, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300609.

Tang:2017:ZKG

[2154] Yao-Jen Tang, Jian-Jhih Kuo, and Ming-Jer Tsai. Zero-knowledge GPSfree data replication and retrieval scheme in mobile ad hoc networks using double-ruling and landmark-labeling techniques. Computer Networks (Amsterdam, Netherlands: 1999), 118 (??):62–77, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300749.

Yang:2017:EAP

[2155] Song Yang, Philipp Wieder, Ramin Yahyapour, and Xiaoming Fu. Energyaware provisioning in optical cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 118 (??):78–95, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300762.

Marques:2017:MRT

[2156] Hugo Marques, Hélio Silva, Evariste Logota, Jonathan Rodriguez, Seiamak Vahid, and Rahim Tafazolli. Multiview real-time media distribution for next generation networks. Computer Networks (Amsterdam, Netherlands: 1999), 118(??):96–124, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300622.

Anonymous:2017:EBg

[2157] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 118(??):ifc, May 8, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301093.

Shi:2017:ERF

[2158] Hongtao Shi, Hongping Li, Dan Zhang, Chaqiu Cheng, and Wei Wu. Efficient and robust feature extraction and selection for traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 119(??):1–16, June 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300786.

Domingues:2017:EOS

[2159] Guilherme Domingues, Edmundo de Souza e Silva, Rosa M. M. Leão, Daniel S. Menasché, and Don Towsley. Enabling opportunistic search and placement in cache networks. Computer Networks (Amsterdam, Netherlands: 1999), 119 (??):17–34, June 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300737.

Ruckert:2017:CSP

[2160] Julius Rückert, Julian Wulfheide, Tamara Knierim, and David Hausheer. On clubs, screaming peers, and duplicates: a comprehensive study of BitTorrent Live. Computer Networks (Amsterdam, Netherlands: 1999), 119 (??):35–55, June 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300750.

Zander:2017:CPP

[2161] Sebastian Zander, Lachlan L. H. Andrew, and Grenville Armitage. Collaborative and privacy-preserving estimation of IP address space utilisation. Computer Networks (Amsterdam, Netherlands: 1999), 119 (??):56–70, June 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300798.

Swain:2017:DSG

[2162] Siba Narayan Swain, Rahul Thakur, and C. Siva Ram Murthy. Design and stochastic geometric analysis of an efficient Q-Learning based physical resource block allocation scheme to maximize the spectral efficiency of Device-to-Device overlaid cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 119 (??):71–85, June 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301056.

Papadopoulos:2017:TIT

[2163] Georgios Z. Papadopoulos, Antoine Gallais, Guillaume Schreiner, Emery Jou, and Thomas Noel. Thorough IoT testbed characterization: From proof-of-concept to repeatable experimentations. Computer Networks (Amsterdam, Netherlands: 1999), 119 (??):86–101, June 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617300804.

Dong:2017:NFS

[2164] Yu ning Dong, Jia jie Zhao, and Jiong Jin. Novel feature selection and classification of Internet video traffic based on a hierarchical scheme. Computer Networks (Amsterdam, Netherlands: 1999), 119(??):102–111, June 4, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301184.

Acedo-Hernandez:2017:PPA

[2165] R. Acedo-Hernández, M. Toril, S. Luna-Ramírez, and C. Úbeda. A PCI planning algorithm for jointly reducing reference signal collisions in LTE uplink and downlink. Computer Networks (Amsterdam, Netherlands: 1999), 119 (??):112–123, June 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301214.

Anonymous:2017:EBh

[2166] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 119(??):ifc, June 4, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301329.

Wu:2017:SPI

[2167] Yi Wu, Ge Nong, and Mounir Hamdi.

Scalable pipelined IP lookup with prefix tries. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):1–11, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301160.

Ibn-Khedher:2017:OOP

[2168] Hatem Ibn-Khedher, Emad Abd-Elrahman, Ahmed E. Kamal, and Hossam Afifi. OPAC: an optimal placement algorithm for virtual CDN. Computer Networks (Amsterdam, Netherlands: 1999), 120(??):12–27, June 19, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301391.

Costantino:2017:PMC

[2169] Gianpiero Costantino, Rajib Maiti, Fabio Martinelli, and Paolo Santi. Private mobility-cast for opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):28–42, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301512.

Guan:2017:RCE

[2170] Tong Guan, Le Fang, Wen Dong, Dimitrios Koutsonikolas, Geoffrey Challen, and Chunming Qiao. Robust, costeffective and scalable localization in large indoor areas. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):43–55, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301494.

Qiu:2017:GFT

[2171] Xiaofeng Qiu, Kai Zhang, and Qiuzheng Ren. Global Flow Table: a convincing mechanism for security operations in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):56–70, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301251.

Tajiki:2017:OQA

[2172] Mohammad Mahdi Tajiki, Behzad Akbari, and Nader Mokari. Optimal Qos-aware network reconfiguration in software defined cloud data centers. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):71–86, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301263.

daSilva:2017:MPD

[2173] Pedro Moreira da Silva, Jaime Dias, and Manuel Ricardo. Mistrustful P2P: Deterministic privacy-preserving P2P file sharing model to hide user content interests in untrusted peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):87–104, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301287.

Bo:2017:TTT

[2174] Zhang Bo, Zhang Huan, Li Meizi, Zhao Qin, and Huang Jifeng. Trust Traversal: a trust link detection scheme in social network. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):105–125, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301500.

Galiotto:2017:ENP

[2175] Carlo Galiotto, Nuno K. Pratas, Linda Doyle, and Nicola Marchetti. Effect of LOS/NLOS propagation on 5G ultradense networks. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):126–140, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301536.

Houaidia:2017:IFI

[2176] Chiraz Houaidia, Hanen Idoudi, Adrien Van Den Bossche, Leila Azouz Saidane, and Thierry Val. Inter-flow and intraflow interference mitigation routing in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 120(??):141–156, June 19, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301202.

Zhou:2017:FFI

[2177] Tongqing Zhou, Zhiping Cai, Kui Wu, Yueyue Chen, and Ming Xu. FIDC: a framework for improving data credibility in mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 120(??):157–169, June 19, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301445.

Li:2017:CSU

[2178] Wenzhong Li, Xiao Chen, and Sanglu Lu. Content synchronization using device-to-device communication in smart cities. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):170–185, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301548.

Zhang:2017:SMB

[2179] Yuxiang Zhang, Lin Cui, and Yuan Zhang. A stable matching based elephant flow scheduling algorithm in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 120 (??):186–197, June 19, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301603.

Anonymous:2017:EBi

[2180] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 120(??):ifc, June 19, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617301846.

Gavrilovska:2017:RSE

[2181] Liljana Gavrilovska, Pero Latkoski, Vladimir Atanasovski, Ramjee Prasad, Albena Mihovska, Octavian Fratu, and Pavlos Lazaridis. Radio Spectrum: Evaluation approaches, coexistence issues and monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):1–12, July 5, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301561.

Zhang:2017:EFK

[2182] Tao Zhang and Jing Liu. An efficient and fast kinematics-based algorithm for RFID network planning. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):13–24, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301585.

Jazi:2017:DHB

[2183] Hossein Hadian Jazi, Hugo Gonzalez, Natalia Stakhanova, and Ali A. Ghorbani. Detecting HTTP-based application layer DoS attacks on web servers in the presence of sampling. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):25–36, July 5, 2017.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301172.

Macedo:2017:SSP

[2184] Ricardo Macedo, Leonardo Melniski, Aldri Santos, Yacine Ghamri-Doudane, and Michele Nogueira. SPARTA: a survival performance degradation framework for identity federations. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):37–52, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301524.

Rmayti:2017:SAP

[2185] Mohammad Rmayti, Rida Khatoun, Youcef Begriche, Lyes Khoukhi, and Dominique Gaiti. A stochastic approach for packet dropping attacks detection in mobile Ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 121 (??):53–64, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301615.

Chen:2017:LLC

[2186] Ming Chen, Ke Ding, Jie Hao, Chao Hu, Gaogang Xie, Changyou Xing, and Bing Chen. LCMSC: a lightweight collaborative mechanism for SDN controllers. Computer Networks (Amsterdam, Netherlands: 1999), 121 (??):65–75, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301652.

Sung:2017:DBE

[2187] Jung-Woong Sung and Seung-Jae Han. Data bundling for energy efficient communication of wearable devices. Computer Networks (Amsterdam, Netherlands: 1999), 121 (??):76–88, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730141X.

Besiktas:2017:SVN

[2188] Cihangir Besiktas, Didem Gözüpek, Aydin Ulas, and Erhan Lokman. Secure virtual network embedding with flexible bandwidth-based revenue maximization. Computer Networks (Amsterdam, Netherlands: 1999), 121 (??):89–99, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301457.

Oktian:2017:DSC

[2189] Yustus Eko Oktian, SangGon Lee, HoonJae Lee, and JunHuy Lam. Distributed SDN controller system: a survey on design choice. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):100–111, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301706.

Gazit:2017:MOC

[2190] Nir Gazit, Francesco Malandrino, and David Hay. Mobile operators and content providers in next-generation SDN/ NFV core networks: Between cooperation and competition. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):112–123, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301627.

Cominardi:2017:DMM

[2191] Luca Cominardi, Fabio Giust, Carlos J. Bernardos, and Antonio De La Oliva. Distributed mobility management solutions for next mobile network architectures. Computer Networks (Amsterdam, Netherlands: 1999), 121 (??):124–136, July 5, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301421.

Gkatzikis:2017:LCC

[2192] Lazaros Gkatzikis, Vasilis Sourlas, Carlo Fischione, and Iordanis Koutsopoulos. Low complexity content replication through clustering in Content-Delivery Networks. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):137–151, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301743.

Mehdi:2017:GTB

[2193] Muhammad Mohsin Mehdi, Imran Raza, and Syed Asad Hussain. A game theory based trust model for Vehicular Ad hoc Networks (VANETs). Computer Networks (Amsterdam, Netherlands: 1999), 121 (??):152–172, July 5, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301573.

Gazda:2017:DSL

[2194] Juraj Gazda, Gabriel Bugár, Marcel Volosin, Peter Drotár, Denis Horváth, and Vladimír Gazda. Dynamic spectrum leasing and retail pricing using an experimental economy. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):173–184, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301718.

Rifai:2017:MSW

[2195] Myriana Rifai, Nicolas Huin, Christelle Caillouet, Frederic Giroire, Joanna Moulierac, Dino Lopez Pacheco, and Guillaume Urvoy-Keller. Minnie: an SDN world with few compressed forwarding rules. Computer Networks (Amsterdam, Netherlands: 1999), 121 (??):185–207, July 5, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301433.

Anonymous:2017:EBj

[2196] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 121(??):ifc, July 5, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302104.

Leng:2017:FMR

[2197] Bing Leng, Liusheng Huang, Chunming Qiao, Hongli Xu, and Xinglong Wang. FTRS: a mechanism for reducing flow table entries in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 122(??):1–15, July 20, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301470.

Lutz:2017:VWT

[2198] Jonathan Lutz, Charles J. Colbourn, and Violet R. Syrotiuk. Variableweight topology-transparent scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 122(??): 16–28, July 20, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301640.

Werner:2017:CIM

[2199] Jorge Werner, Carla Merkle Westphall, and Carlos Becker Westphall. Cloud identity management: a survey on privacy strategies. Computer Networks (Amsterdam, Netherlands: 1999), 122(??):29–42, July 20, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301664.

Xu:2017:PFS

[2200] Hongli Xu, Xiang-Yang Li, Liusheng Huang, Yang Du, and Zichun Liu. Partial flow statistics collection for load-balanced routing in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):43–55, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301639.

Xue:2017:NQL

[2201] Dongyue Xue, Eylem Ekici, Rania Ibrahim, and Moustafa Youssef. A novel queue-length-based CSMA algorithm with improved delay characteristics. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):56–69, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301676.

Sundaresan:2017:SSP

[2202] Saravanan Sundaresan, Robin Doss, Selwyn Piramuthu, and Wanlei Zhou. A secure search protocol for low cost passive RFID tags. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):70–82, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730155X.

Stimpfling:2017:EDT

[2203] Thibaut Stimpfling, Normand Bélanger, Omar Cherkaoui, André Béliveau, Ludovic Béliveau, and Yvon Savaria. Extensions to decision-tree based packet classification algorithms to address new classification paradigms. Computer Networks (Amsterdam, Netherlands: 1999), 122(??):83–95, July 20, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301482.

Morawski:2017:NNP

[2204] Michal Morawski and Przemyslaw Ignaciuk. Network nodes play a game — a routing alternative in multihop adhoc environments. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):96–104, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301469.

Wang:2017:IQS

[2205] Yi Wang and Ye Xia. Improving the queue size and delay performance with the I-CSMA link scheduling algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):105–119, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301408.

Hiran:2017:CFP

[2206] Rahul Hiran, Niklas Carlsson, and Nahid Shahmehri. Collaborative framework for protection against attacks targeting BGP and edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):120–137, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301792.

Chunlin:2017:MCB

[2207] Li Chunlin, Yan Xin, Zhang Yang, and Luo Youlong. Multiple context based service scheduling for balancing cost and benefits of mobile users and cloud datacenter supplier in mobile cloud. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):138–152, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730169X.

Pinheiro:2017:EAD

[2208] Antônio J. Pinheiro, Ethel B. Gondim, and Divanilson R. Campelo. An efficient architecture for dynamic middlebox policy enforcement in SDN networks. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):153–162, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301925.

Salehi:2017:NPS

[2209] Mahmood Salehi and Azzedine Boukerche. A novel packet salvaging model to improve the security of opportunistic routing protocols. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):163–178, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301597.

Liu:2017:RPA

[2210] Zhixin Liu, Shiyi Li, Kai Ma, Xinping Guan, and Xinbin Li. Robust power allocation based on hierarchical game with consideration of different user requirements in two-tier femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):179–190, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301937.

Yoon:2017:ADC

[2211] Min Sang Yoon, Ahmed E. Kamal, and Zhengyuan Zhu. Adaptive data center activation with user request prediction. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):191–204, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301780.

Yi:2017:RSC

[2212] Leng-Gan Yi. Rate-sensitive CoMP beamforming plus sensible snooze for energy-QoE tradeoff in cognitive small cell networks. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):205–216, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301688.

Meng:2017:MUB

[2213] Jinli Meng, Fengyuan Ren, and Chuang Lin. Modeling and understanding burst transmission for energy efficient Ethernet. Computer Networks (Amsterdam, Netherlands: 1999), 122 (??):217–230, July 20, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301196.

Anonymous:2017:EBk

[2214] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 122(??):ifc, July 20, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730227X.

Roy:2017:OSU

[2215] Asmita Roy, Sadip Midya, Koushik Majumder, Santanu Phadikar, and Anurag Dasgupta. Optimized secondary user selection for quality of service enhancement of two-tier multi-user Cognitive Radio Network: a game theoretic approach. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):1–18, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301949.

Qamar:2017:CRC

[2216] Faizan Qamar, Kaharudin Bin Dimyati, Mhd Nour Hindia, Kamarul Ariffin Bin Noordin, and Ahmed M. Al-Samman. A comprehensive review on coordinated multi-point operation for LTE-A. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):19–37, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301950.

Wu:2017:SIB

[2217] Peng Wu and Li Pan. Scalable influence blocking maximization in social networks under competitive independent cascade models. Computer Networks (Amsterdam, Netherlands: 1999), 123(??):38–50, August 4, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301962.

Borrego:2017:EWC

[2218] Carlos Borrego, Adrián Sánchez-Carmona, Zhiyuan Li, and Sergi Robles. Explore and wait: a composite routing-delivery scheme for relative profile-casting in opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):51–63, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302013.

Karar:2017:OSA

[2219] Sandip Karar and Abhirup Das Barman. Opportunistic spectrum access for cooperative unlicensed femtocells in two-tier heterogeneous networks utilizing HARQ feedback. Computer Networks (Amsterdam, Netherlands: 1999), 123(??):64–76, August 4, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301986.

Xu:2017:PLS

[2220] Yang Xu, Jia Liu, Yulong Shen, Xiaohong Jiang, and Norio Shiratori. Physical layer security-aware routing and performance tradeoffs in ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):77–87, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302049.

Lv:2017:RIR

[2221] Jianhui Lv, Xingwei Wang, Min Huang, Junling Shi, Keqin Li, and Jie Li. RISC: ICN routing mechanism incorporating SDN and community division. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):88–103, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302025.

Meneguette:2017:SES

[2222] Rodolfo I. Meneguette and Azzedine Boukerche. SERVitES: an efficient search and allocation resource protocol based on V2V communication for vehicular cloud. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):104–118, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302062.

Donatiello:2017:MPE

[2223] Lorenzo Donatiello and Gustavo Marfia. Models and performance evaluation of event goodput in sensor platforms. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):119–136, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302074.

Mohanty:2017:DCM

[2224] Jasaswi Prasad Mohanty, Chittaranjan Mandal, and Chris Reade. Distributed construction of minimum Connected Dominating Set in wireless sensor network using two-hop information. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):137–152, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302177.

Barrachina-Munoz:2017:MHC

[2225] Sergio Barrachina-Muñoz, Boris Bellalta, Toni Adame, and Albert Bel. Multi-hop communication in the uplink for LPWANs. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):153–168, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302207.

Bonafiglia:2017:EDH

[2226] Roberto Bonafiglia, Amedeo Sapio, Mario Baldi, Fulvio Risso, and Paolo C. Pomi. Enforcement of dynamic HTTP policies on resource-constrained residential gateways. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):169–183, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302086.

Saginbekov:2017:MMD

[2227] Sain Saginbekov and Arshad Jhumka. Many-to-many data aggregation scheduling in wireless sensor networks with two sinks. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):184–199, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302232.

Wang:2017:CPU

[2228] Yang Wang, Shuang Wu, Zhiyin Chen, Xiaofeng Gao, and Guihai Chen. Coverage problem with uncertain properties in wireless sensor networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 123 (??):200–232, August 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301998.

Anonymous:2017:EBl

[2229] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 123(??):ifc, August 4, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617302785.

Nikoletseas:2017:RCA

[2230] Sotiris Nikoletseas, Theofanis P. Raptis, and Christoforos Raptopoulos. Radiation-constrained algorithms for Wireless Energy Transfer in Ad hoc Networks. Computer Networks (Amsterdam, Netherlands: 1999), 124 (??):1–10, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302323.

Gheisari:2017:NRL

[2231] S. Gheisari and M. R. Meybodi. A new reasoning and learning model for Cognitive Wireless Sensor Networks based on Bayesian networks and learning automata cooperation. Computer Networks (Amsterdam, Netherlands: 1999), 124(??): 11–26, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302426.

Ezran:2017:AOR

[2232] Philippe Ezran, Yoram Haddad, and Mérouane Debbah. Availability optimization in a ring-based network topology. Computer Networks (Amsterdam, Netherlands: 1999), 124(??): 27–32, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302050.

Safi:2017:PCO

[2233] Qamas Gul Khan Safi, Senlin Luo, Chao Wei, Limin Pan, and Qianrou Chen. PIaaS: Cloud-oriented secure and privacy-conscious parking information as a service using VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 124(??): 33–45, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861730244X.

Xu:2017:TOT

[2234] Lei Xu, Ke Xu, Yong Jiang, Fengyuan Ren, and Haiyang Wang. Throughput optimization of TCP incast congestion control in large-scale datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 124 (??):46–60, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302475.

Yang:2017:ERA

[2235] Zi-Yang Yang and Yaw-Wen Kuo.

Efficient resource allocation algorithm for overlay D2D communication. Computer Networks (Amsterdam, Netherlands: 1999), 124(??): 61–71, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302451.

Erdelj:2017:WSN

[2236] Milan Erdelj, Michal Król, and Enrico Natalizio. Wireless Sensor Networks and Multi-UAV systems for natural disaster management. Computer Networks (Amsterdam, Netherlands: 1999), 124(??): 72–86, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302220.

Manesh:2017:RTS

[2237] Mohsen Riahi Manesh, Sririam Subramaniam, Hector Reyes, and Naima Kaabouch. Real-time spectrum occupancy monitoring using a probabilistic model. Computer Networks (Amsterdam, Netherlands: 1999), 124 (??):87–96, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302463.

Fang:2017:NAF

[2238] Xuming Fang, Lei Nan, Zonghua Jiang, and Lijun Chen. Noiseaware fingerprint localization algorithm for wireless sensor network based on adaptive fingerprint Kalman filter. Computer Networks (Amsterdam, Netherlands: 1999), 124(??): 97–107, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302608.

Boukerche:2017:MTP

[2239] Azzedine Boukerche and Xiaoli Zhou. MAC transmission protocols for delaytolerant sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 124(??):108–125, September 4, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128617302566.

Baroutis:2017:LCM

[2240] Nikolaos Baroutis and Mohamed Younis. Load-conscious maximization of base-station location privacy in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 124(??):126– 139, September 4, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302657.

Xie:2017:ITE

[2241] Gaogang Xie, Jingxiu Su, Xin Wang, Taihua He, Guangxing Zhang, Steve Uhlig, and Kave Salamatian. Index-Trie: Efficient archival and retrieval of network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 124 (??):140–156, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302542.

Olsen:2017:UIQ

[2242] Rasmus L. Olsen, Jacob Theilgaard Madsen, Jakob G. Rasmussen, and Hans-Peter Schwefel. On the use of information quality in stochastic networked control systems. Computer Networks (Amsterdam, Netherlands: 1999), 124(??): 157–169, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302505.

Jia:2017:LOF

[2243] Xuya Jia, Qing Li, Yong Jiang, Zehua Guo, and Jie Sun. A low overhead flowholding algorithm in software-defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 124(??): 170–180, September 4, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302530.

Anonymous:2017:EBm

[2244] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 124(??):ifc, September 4, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617302992.

Hossfeld:2017:SCN

[2245] Tobias Hoßfeld, S.-H. Gary Chan, Brian L. Mark, and Andreas Timm-Giel. Softwarization and caching in NGN. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):1–3, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302670.

Zinner:2017:DTM

[2246] Thomas Zinner, Stefan Geissler, Stanislav Lange, Steffen Gebert, Michael Seufert, and Phuoc Tran-Gia. A discrete-time model for optimizing the processing time of virtualized network functions. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):4–14, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301809.

Guo:2017:SPF

[2247] Zehua Guo, Ruoyan Liu, Yang Xu, Andrey Gushchin, Anwar Walid, and H. Jonathan Chao. STAR: Preventing flow-table overflow in softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):15–25, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301779.

Zhang:2017:TSE

[2248] Sai Qian Zhang, Qi Zhang, Ali Tizghadam, Byungchul Park, Hadi Bannazadeh, Raouf Boutaba, and Alberto Leon-Garcia. TCAM spaceefficient routing in a software defined network. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):26–40, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302645.

Li:2017:FIU

[2249] Qing Li, Lei Wang, Yong Jiang, Mingwei Xu, and Jianping Wu. A fast and incremental update scheme for SDN based on a relation graph. Computer Networks (Amsterdam, Netherlands: 1999), 125(??):41–52, October 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301974.

Gao:2017:CAM

[2250] Qiang Gao, Weiqin Tong, Samina Kausar, Lei Huang, Chao Shen, and Shenan Zheng. Congestion-aware multicast plug-in for an SDN network operating system. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):53–63, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301810.

Marotta:2017:ECR

[2251] Antonio Marotta, Fabio D’Andreagiovanni, Andreas Kassler, and Enrica Zola.

On the energy cost of robustness for green virtual network function placement in 5G virtualized infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):64–75, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301767.

Nam:2017:JNE

[2252] Tran Manh Nam, Nguyen Huu Thanh, Hoang Trung Hieu, Nguyen Tien Manh, Nguyen Van Huynh, and Hoang Duong Tuan. Joint network embedding and server consolidation for energy-efficient dynamic data center virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):76–89, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302517.

Maswood:2017:EED

[2253] Mirza Mohd Shahriar Maswood, Chris Develder, Edmundo Madeira, and Deep Medhi. Energy-efficient dynamic virtual network traffic engineering for north-south traffic in multi-location data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):90–102, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301731.

Majeed:2017:MSI

[2254] Muhammad Faran Majeed, Syed Hassan Ahmed, Siraj Muhammad, Houbing Song, and Danda B. Rawat. Multimedia streaming in information-centric networking: a survey and future perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):103–121, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302414.

Bonald:2017:FTA

[2255] Thomas Bonald, Léonce Mekinda, and Luca Muscariello. Fair throughput allocation in Information-Centric Networks. Computer Networks (Amsterdam, Netherlands: 1999), 125(??): 122–131, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302190.

Mansilha:2017:EPH

[2256] Rodrigo B. Mansilha, Marinho P. Barcellos, Emilio Leonardi, and Dario Rossi. Exploiting parallelism in hierarchical content stores for high-speed ICN routers. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):132–145, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730172X.

Tortelli:2017:HMP

[2257] M. Tortelli, D. Rossi, and E. Leonardi.

A hybrid methodology for the performance evaluation of Internet-scale cache networks. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):146–159, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301305.

Josilo:2017:DAC

[2258] Sladana Josilo, Valentino Pacifici, and György Dán. Distributed algorithms for content placement in hierarchical cache networks. Computer Networks (Amsterdam, Netherlands: 1999), 125 (??):160–171, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302335.

Hasslinger:2017:PEN

[2259] Gerhard Hasslinger, Konstantinos Ntougias, Frank Hasslinger, and Oliver Hohlfeld. Performance evaluation for new web caching strategies combining LRU with score based object selection. Computer Networks (Amsterdam, Netherlands: 1999), 125(??): 172–186, October 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301755.

Anonymous:2017:EBn

[2260] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 125(??):ifc, October 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128617303274.

Coutinho:2017:PMA

[2261] Rodolfo W. L. Coutinho, Azzedine Boukerche, Luiz F. M. Vieira, and Antonio A. F. Loureiro. Performance modeling and analysis of void-handling methodologies in underwater wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):1–14, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302712.

Shamna:2017:ETE

[2262] HR Shamna and Jacob Lillykutty.

An energy and throughput efficient distributed cooperative MAC protocol for multihop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):15–30, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302682.

Wang:2017:NCB

[2263] Tao Wang, Baoxian Zhang, Zheng Yao, and Hussein T. Mouftah. Network coding based adaptive CSMA for network utility maximization. Computer Networks (Amsterdam, Netherlands: 1999), 126(??):31– 43, October 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730261X.

Hajisami:2017:DJP

[2264] Abolfazl Hajisami and Dario Pompili.

Dynamic joint processing: Achieving high spectral efficiency in uplink 5G cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):44–56, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302700.

Ekmen:2017:REE

[2265] Merve Ekmen and Aysegül Altin-Kayhan. Reliable and energy efficient wireless sensor network design via conditional multi-copying for multiple central nodes. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):57–68, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302724.

Ye:2017:MSC

[2266] Xing Ye, Guozhen Cheng, and Xingguo Luo. Maximizing SDN control resource utilization via switch migration. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):69–80, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302669.

dAmbrosio:2017:PPG

[2267] Salvatore d’Ambrosio, Salvatore de Pasquale, Gerardo Iannone, Delfina Malandrino, Alberto Negro, Giovanni Patimo, Vittorio Scarano, Raffaele Spinelli, and Rocco Zaccagnino. Privacy as a proxy for Green Web browsing: Methodology and experimentation. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):81–99, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302761.

Aggarwal:2017:SSL

[2268] Charu C. Aggarwal, Amotz Bar-Noy, and Simon Shamoun. On sensor selection in linked information networks. Computer Networks (Amsterdam, Netherlands: 1999), 126(??): 100–113, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302256.

Luvisotto:2017:LAR

[2269] Michele Luvisotto, Federico Tramarin, and Stefano Vitturi. A learning algorithm for rate selection in real-time wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):114–124, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730275X.

Clausen:2017:LDA

[2270] Thomas Clausen, Jiazi Yi, and Ulrich Herberg. Lightweight Ondemand Ad hoc Distance-vector Routing — Next Generation (LOADng): Protocol, extension, and applicability. Computer Networks (Amsterdam, Netherlands: 1999), 126(??): 125–140, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302694.

Wei:2017:TSA

[2271] Zhenchun Wei, Yan Zhang, Xiangwei Xu, Lei Shi, and Lin Feng. A task scheduling algorithm based on Q-learning and shared value function for WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 126(??): 141–149, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302487.

Kim:2017:DRD

[2272] Hyung-Sin Kim, Hosoo Cho, Hongchan Kim, and Saewoong Bahk. DT-RPL: Diverse bidirectional traffic delivery through RPL routing protocol in low power and lossy networks. Computer Networks (Amsterdam, Netherlands: 1999), 126(??): 150–161, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302748.

Gregori:2017:IME

[2273] Enrico Gregori, Barbara Guidi, Alessandro Improta, and Luca Sani. ICE: a memory-efficient BGP route collecting engine. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):162–173, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302906.

Casetti:2017:AFC

[2274] Claudio Casetti, Carla-Fabiana Chiasserini, Francesco Malandrino, and Carlo Borgiattino. Area formation and content assignment for LTE broadcasting. Computer Networks (Amsterdam, Netherlands: 1999), 126(??): 174–186, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302876.

Guo:2017:TEH

[2275] Yingya Guo, Zhiliang Wang, Xia Yin, Xingang Shi, and Jianping Wu. Traffic engineering in hybrid SDN networks with multiple traffic matrices. Computer Networks (Amsterdam, Netherlands: 1999), 126(??): 187–199, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730289X.

Lv:2017:AII

[2276] Jianhui Lv, Xingwei Wang, Kexin Ren, Min Huang, and Keqin Li. ACOinspired Information-Centric Networking routing mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 126(??):200–217, October 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617302852.

Kim:2017:ENF

[2277] Dohyung Kim and Younghoon Kim.

Enhancing NDN feasibility via dedicated routing and caching. Computer Networks (Amsterdam, Netherlands: 1999), 126(??):218–228, October 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861730292X.

Wang:2017:GIR

[2278] Xingwei Wang, Jinhong Zhang, Min Huang, and Shengxiang Yang. A green intelligent routing algorithm supporting flexible QoS for many-to-many multicast. Computer Networks (Amsterdam, Netherlands: 1999), 126(??): 229–245, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302918.

Xu:2017:DVT

[2279] Guangwei Xu, Yanke Bai, Qiao Pan, Qiubo Huang, and Yanbin Yang. Data verification tasks scheduling based on dynamic resource allocation in mobile big data storage. Computer Networks (Amsterdam, Netherlands: 1999), 126 (??):246–255, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302979.

Yang:2017:PUE

[2280] Lin Yang, Mingxuan Yuan, Yanjiao Chen, Wei Wang, Qian Zhang, and Jia Zeng. Personalized user engagement modeling for mobile videos. Computer Networks (Amsterdam, Netherlands: 1999), 126(??): 256–267, October 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302931.

Anonymous:2017:EBo

[2281] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 126(??):ifc, October 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617303481.

Chenait:2017:EEC

[2282] Manel Chenait, Bahia Zebbane, Chafika Benzaid, and Nadjib Badache. Energy-efficient coverage protocol based on stable and predictive scheduling in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):1–12, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302967.

Mukherjee:2017:EEA

[2283] Shreyasee Mukherjee, Shravan Sriram, Tam Vu, and Dipankar Raychaudhuri. EIR: Edge-aware inter-domain routing protocol for the future mobile Internet. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):13–30, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302943.

Kuo:2017:CBM

[2284] Yaw-Wen Kuo and Jane-Hwa Huang.

A CSMA-based MAC protocol for WLANs with automatic synchronization capability to provide hard quality of service guarantees. Computer Networks (Amsterdam, Netherlands: 1999), 127(??):31–42, November 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617302888.

Guo:2017:SSP

[2285] Deke Guo, Xiaoqiang Teng, Zhiyao Hu, Junjie Xie, and Bangbang Ren. Source selection problem in multisource multi-destination multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 43–55, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302955.

Hu:2017:TSA

[2286] Chia-Cheng Hu, Chin-Feng Lai, Ji-Gong Hou, and Yueh-Min Huang. Timely scheduling algorithm for P2P streaming over MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 127(??):56–67, November 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617303080.

Gomez:2017:ENT

[2287] Santiago Egea Gómez, Belén Carro Martínez, Antonio J. Sánchez-Esguevillas, and Luis Hernández Callejo. Ensemble network traffic classification: Algorithm comparison and novel ensemble scheme proposal. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):68–80, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303079.

Hassen:2017:SPS

[2288] Fadoua Hassen and Lotfi Mhamdi.

A scalable packet-switch architecture based on OQ NoCs for data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):81–93, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303109.

Nunes:2017:GNG

[2289] Ivan O. Nunes, Clayson Celes, Pedro O. S. Vaz de Melo, and Antonio A. F. Loureiro. GROUPS-NET: Group meetings aware routing in multi-hop D2D networks. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):94–108, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303237.

Marcon:2017:AMB

[2290] Daniel S. Marcon, Fabrício M. Mazzola, and Marinho P. Barcellos. Achieving minimum bandwidth guarantees and work-conservation in largescale, SDN-based datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 109–125, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303195.

Mansouri:2017:CLA

[2291] Wahida Mansouri, Faouzi Zarai, Kais Mnif, and Lotfi Kamoun. Cross layer architecture with integrated MIH in heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 126–137, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303183.

Yim:2017:VLS

[2292] Yongbin Yim, Hyunchong Cho, Sang-Ha Kim, Euisin Lee, and Mario Gerla. Vehicle location service scheme based on road map in Vehicular Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 138–150, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303213.

Ghazvini:2017:SHM

[2293] Farshad Keyvan Ghazvini, Mustafa Mehmet-Ali, and Mahmoud Doughan. Scalable hybrid MAC protocol for M2M communications. Computer Networks (Amsterdam, Netherlands: 1999), 127(??):151–160, November 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128617303134.

Wang:2017:KMP

[2294] Yongqiang Wang, Krishna Mosalakanti, Felipe Núñez, Socrates Deligeorges, and Francis J. Doyle III. A kernel module for pulse-coupled time synchronization of sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 161–172, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303158.

Hellaoui:2017:EEM

[2295] Hamed Hellaoui, Mouloud Koudil, and Abdelmadjid Bouabdallah. Energyefficient mechanisms in security of the Internet of Things: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 173–189, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303146.

Lee:2017:TBS

[2296] Kyu-Hwan Lee, Dong-Hyuk Jang, Sung-Jae Lee, and Jae-Ryong Cha. Target BER selection scheme in LMS networks using AL-FEC systems. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 190–199, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303122.

Viegas:2017:TRA

[2297] Eduardo K. Viegas, Altair O. Santin, and Luiz S. Oliveira. Toward a reliable anomaly-based intrusion detection in real-world environments. Computer Networks (Amsterdam, Netherlands: 1999), 127(??):200– 216, November 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303225.

Ghosh:2017:EEH

[2298] Pradipta Ghosh, He Ren, Reza Banirazi, Bhaskar Krishnamachari, and Edmond Jonckheere. Empirical evaluation of the heat-diffusion collection protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 217–232, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303250.

Wetzels:2017:FTS

[2299] Frank Wetzels, Hans van den Berg, Joost Bosman, and Rob van der Mei. Flow termination signaling in the centralized pre-congestion notification architecture. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 233–242, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303201.

Nemeth:2017:ORP

[2300] Krisztián Németh, Attila Kőrösi, and Gábor Rétvári. Optimal resource pooling over legacy equal-split load balancing schemes. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):243–265, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303249.

Neto:2017:CBS

[2301] Manassés Ferreira Neto, Olga Goussevskaia, and Vinícius Fernandes dos Santos. Connectivity with backbone structures in obstructed wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 266–281, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303419.

Chouikhi:2017:DCR

[2302] Samira Chouikhi, Inès El Korbi, Yacine Ghamri-Doudane, and Leila Azouz Saidane. Distributed connectivity restoration in multichannel wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):282–295, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303171.

Santhappan:2017:NCF

[2303] Thomas Valerrian Pasca Santhappan, Sumanta Patro, Bheemarjuna Reddy Tamma, and Antony Franklin A. Network Coordination Function for uplink traffic steering in tightly coupled LTE Wi-Fi networks. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):296–316, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303420.

Ventrella:2017:PSM

[2304] Agnese V. Ventrella, Giuseppe Piro, and L. Alfredo Grieco. Publish– subscribe in mobile information centric networks: Modeling and performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):317–339, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303377.

Goli-Bidgoli:2017:DVR

[2305] Salman Goli-Bidgoli and Naser Movahhedinia. Determining vehicles’ radio transmission range for increasing cognitive radio VANET (CR-VANET) reliability using a trust management system. Computer Networks (Amsterdam, Netherlands: 1999), 127(??): 340–351, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303067.

Zhang:2017:MLM

[2306] Han Zhang, Xingang Shi, Yingya Guo, Zhiliang Wang, and Xia Yin. More load, more differentiation — let more flows finish before deadline in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 127 (??):352–367, November 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861730333X.

Anonymous:2017:EBp

[2307] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 127(??):ifc, November 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617303705.

Kamal:2017:SSE

[2308] Ahmed E. Kamal, Muhammad Imran, Hsiao-Hwa Chen, and Athanasios V. Vasilakos. Survivability strategies for emerging wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 128(??):1–4, December 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303110.

Asghar:2017:TPC

[2309] Muhammad Zeeshan Asghar, Paavo Nieminen, Seppo Hämäläinen, Tapani Ristaniemi, Muhammad Ali Imran, and Timo Hämäläinen. Towards proactive context-aware self-healing for 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):5–13, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301895.

Rejeb:2017:CSF

[2310] Sonia Ben Rejeb, Nidal Nasser, Samer Mansour, and Massa Boujlbane. Centralized SON function for operator optimal strategies in heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):14–27, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302621.

Xie:2017:SAT

[2311] Lang Xie, Poul E. Heegaard, and Yuming Jiang. Survivability analysis of a two-tier infrastructure-based wireless network. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):28–40, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301226.

Al-Awami:2017:RDD

[2312] Louai Al-Awami and Hossam S. Hassanein. Robust decentralized data storage and retrieval for wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):41–50, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300427.

Das:2017:NCC

[2313] Bhaskar Das and Jalal Almhana. A new cooperative communication algorithm for improving connectivity in the event of network failure in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):51–62, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301275.

Lalouani:2017:ORP

[2314] Wassila Lalouani, Mohamed Younis, and Nadjib Badache. Optimized repair of a partitioned network topology. Computer Networks (Amsterdam, Netherlands: 1999), 128(??): 63–77, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300415.

Ever:2017:MGH

[2315] Enver Ever, Fadi M. Al-Turjman, Hadi Zahmatkesh, and Mustafa Riza. Modelling green HetNets in dynamic ultra large-scale applications: a case-study for femtocells in smartcities. Computer Networks (Amsterdam, Netherlands: 1999), 128(??): 78–93, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730107X.

eddineHammami:2017:FTD

[2316] Seif eddine Hammami and Hossam Afifi. Fault-tolerant dynamic planning for wireless mesh networks based on real load profiles. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):94–107, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302529.

Sharafeddine:2017:FRW

[2317] Sanaa Sharafeddine, Karim Jahed, Omar Farhat, and Zaher Dawy. Failure recovery in wireless content distribution networks with device-to-device cooperation. Computer Networks (Amsterdam, Netherlands: 1999), 128(??): 108–122, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301901.

Galinina:2017:LHD

[2318] Olga Galinina, Sergey Andreev, Mikhail Komarov, and Svetlana Maltseva. Leveraging heterogeneous device connectivity in a converged 5G-IoT ecosystem. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):123–132, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301822.

Li:2017:CAH

[2319] Di Li, Xiaomin Li, and Jiafu Wan. A cloud-assisted handover optimization strategy for mobile nodes in industrial wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):133–141, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302359.

Tahir:2017:LCD

[2320] A. Tahir, S. A. Abid, and Nadir Shah. Logical clusters in a DHTparadigm for scalable routing in MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 128(??): 142–153, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302499.

He:2017:AHA

[2321] Debiao He, Sherali Zeadally, Libing Wu, and Huaqun Wang. Analysis of handover authentication protocols for mobile wireless networks using identity-based public key cryptography. Computer Networks (Amsterdam, Netherlands: 1999), 128(??): 154–163, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128616304285.

Zhao:2017:RAS

[2322] Caidan Zhao, Minmin Huang, Lianfen Huang, Xiaojiang Du, and Mohsen Guizani. A robust authentication scheme based on physical-layer phase noise fingerprint for emerging wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 128(??): 164–171, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302347.

Bhavathankar:2017:ODR

[2323] Prasenjit Bhavathankar, Subhadeep Sarkar, and Sudip Misra. Optimal decision rule-based ex-ante frequency hopping for jamming avoidance in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 128 (??):172–185, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617300774.

Soleimani:2017:SMG

[2324] Hossein Soleimani, Thomas Begin, and Azzedine Boukerche. Safety message generation rate adaptation in LTE-based vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 128(??): 186–196, December 9, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301913.

Anonymous:2017:EBq

[2325] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 128(??):ifc, December 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128617303833.

Sharafeddine:2017:ODC

[2326] Sanaa Sharafeddine, Karim Jahed, and Marwan Fawaz. Optimized device centric aggregation mechanisms for mobile devices with multiple wireless interfaces. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):1–16, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303432.

Peng:2017:MMD

[2327] Yi-Ting Peng, Sok-Ian Sou, Meng-Hsun Tsai, and Chuan-Sheng Lin. Multipath mobile data offloading of deadline assurance with policy and charging control in cellular/WiFi networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)

(??):17–27, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303365.

Vahedi:2017:SEB

[2328] Erfaneh Vahedi, Majid Bayat, Mohammad Reza Pakravan, and Mohammad Reza Aref. A secure ECCbased privacy preserving data aggregation scheme for smart grids. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):28– 36, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303353.

Zhang:2017:GFD

[2329] Jian Zhang, Yang Yang, Yanjiao Chen, Jing Chen, and Qian Zhang. A general framework to design secure cloud storage protocol using homomorphic encryption scheme. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):37–50, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303328.

Huang:2017:EEE

[2330] Haojun Huang, Junbao Zhang, Xu Zhang, Benshun Yi, Qilin Fan, and Feng Li. EMGR: Energy-efficient multicast geographic routing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):51–63, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861730316X.

Wang:2017:RRA

[2331] YiPeng Wang, Xiaochun Yun, Yongzheng Zhang, Liwei Chen, and Tianning Zang. Rethinking robust and accurate application protocol identification. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):64–78, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303572.

Lee:2017:RTP

[2332] Soojeon Lee, Dongman Lee, Myungjin Lee, Hyungsoo Jung, and Byoung-Sun Lee. Randomizing TCP payload size for TCP fairness in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):79–92, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303584.

Malboubi:2017:SDN

[2333] Mehdi Malboubi, Yanlei Gong, Zijun Yang, Xiong Wang, Chen-Nee Chuah, and Puneet Sharma. Software defined network inference with evolutionary optimal observation matrices. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):93–104, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303444.

Osgouei:2017:OAN

[2334] Amin Ghalami Osgouei, Amir Khorsandi Koohanestani, Hossein Saidi, and Ali Fanian. Online assignment of non-SDN virtual network nodes to a physical SDN. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):105–116, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303559.

Santos:2017:CCB

[2335] Bruno P. Santos, Luiz F. M. Vieira, and Marcos A. M. Vieira. CGR: Centrality-based green routing for lowpower and lossy networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):117– 128, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303602.

Alaoui:2017:ENR

[2336] Sara El Alaoui and Byrav Ramamurthy. EAODR: a novel routing algorithm based on the Modified Temporal Graph network model for DTN-based Interplanetary Networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):129–141, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303638.

Katsalis:2017:IEM

[2337] K. Katsalis, B. Rofoee, G. Landi, J. F. Riera, K. Kousias, M. Anastasopoulos, L. Kiraly, A. Tzanakaki, and T. Korakis. Implementation experience in multi-domain SDN: Challenges, consolidation and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):142–158, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303560.

Ermis:2017:KAP

[2338] Orhan Ermis, Serif Bahtiyar, Emin Anarim, and M. Ufuk Çaglayan. A key agreement protocol with partial backward confidentiality. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):159– 177, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303596.

Ding:2017:ITC

[2339] Lei Ding, Jun Liu, Tao Qin, and Haifei Li. Internet traffic classification based on expanding vector of flow. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):178–192, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303663.

Niu:2017:UAP

[2340] Jianwei Niu, Shihao Wang, Wei Niu, and Mohammed Atiquzzaman. Useraware partitioning algorithm for mobile cloud computing based on maximum graph cuts. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):193–206, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303626.

Scalosub:2017:TOB

[2341] Gabriel Scalosub. Towards optimal buffer management for streams with packet dependencies. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):207– 214, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303651.

Qu:2017:OCA

[2342] Ying Qu, Bryan Ng, and Michael Homer. Optimising channel assignment to prevent flow starvation and improve fairness for planning single radio WMNs in built environments. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):215–231, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303675.

Nguyen:2017:DHB

[2343] Phi Le Nguyen, Yusheng Ji, Zhi Liu, Huy Vu, and Khanh-Van Nguyen. Distributed hole-bypassing protocol in WSNs with constant stretch and load balancing. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):232–250, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303687.

Qi:2017:ERE

[2344] Ji Qi, Fei Hu, Xin Li, Koushik A. M., and Sunil Kumar. 3-ent (resili-ent, intellig-ent, and effici-ent) medium access control for full-duplex, jamming-aware, directional airborne networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):251–260, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303456.

Ceron:2017:MTC

[2345] João Marcelo Ceron, Cíntia Borges Margi, and Lisandro Zambenedetti Granville. MARS: From traffic containment to network reconfiguration in malware-analysis systems. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):261– 272, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861730378X.

Thakur:2017:RAC

[2346] Rahul Thakur, Vijeth J. Kotagi, and C. Siva Ram Murthy. Resource allocation and cell selection framework for LTE-Unlicensed femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):273–283, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303791.

Lee:2017:OSA

[2347] Hyunjoong Lee, Sangkyu Park, and Saewoong Bahk. An opportunistic scheduling algorithm using aged CSI in massive MIMO systems. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):284– 296, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730364X.

Wu:2017:EIT

[2348] Jia Wu, Zhigang Chen, and Ming Zhao. Effective information transmission based on socialization nodes in opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??):297– 305, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303808.

Qiu:2017:AAS

[2349] Yue Qiu, Maode Ma, and Shuo Chen.

An anonymous authentication scheme for multi-domain machine-to-machine communication in cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1) (??):306–318, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730381X.

Ayadi:2017:ODA

[2350] Aya Ayadi, Oussama Ghorbel, Abdulfattah M. Obeid, and Mohamed Abid. Outlier detection approaches for wireless sensor networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??): 319–333, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303900.

Anonymous:2017:EBr

[2351] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 1)(??): ifc, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303961.

Rudin:2017:CHW

[2352] Harry Rudin. Change at the Helm: Welcome, farewell, advice, thanks: Special issues. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??):iii, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304103.

Rodrigues:2017:SIW

[2353] J. J. P. C. Rodrigues, S. Zeadally, N. Kumar, and G. Han. Special issue on 5G wireless networks for IoT and body sensors. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):335–339, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304036.

Lloret:2017:APS

[2354] Jaime Lloret, Lorena Parra, Miran Taha, and Jesus Tomás. An architecture and protocol for smart continuous eHealth monitoring using 5G. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??): 340–351, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302189.

Marin:2017:CAC

[2355] Radu-Corneliu Marin, Radu-Ioan Ciobanu, Ciprian Dobre, Constandinos X. Mavromoustakis, and George Mastorakis. A context-aware collaborative model for smartphone energy efficiency over 5G wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):352–362, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302554.

You:2017:STB

[2356] Ilsun You and Jong-Hyouk Lee. SPFP:

Ticket-based secure handover for fast proxy mobile IPv6 in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):363–372, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302001.

Fortuna:2017:SIC

[2357] Carolina Fortuna, Adnan Bekan, Tomaz Javornik, Gregor Cerar, and Mihael Mohorcic. Software interfaces for control, optimization and update of 5G machine type communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):373–383, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302578.

Han:2017:FRS

[2358] Shuai Han, Cheng Guo, Weixiao Meng, and Cheng Li. A flexible resource scheduling scheme for an adaptive SCMA system. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??):384– 391, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303614.

Li:2017:RMF

[2359] Yujie Li, Zhibin Gao, Lianfen Huang, Xiaojiang Du, and Mohsen Guizani. Resource management for future mobile networks: Architecture and technologies. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):392–398, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301299.

Liu:2017:IMC

[2360] Yang Liu, Changqiao Xu, Yufeng Zhan, Zhixin Liu, Jianfeng Guan, and Hongke Zhang. Incentive mechanism for computation offloading using edge computing: a Stackelberg game approach. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):399–409, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301068.

Yang:2017:RDF

[2361] Lijun Yang, Chao Ding, Meng Wu, and Kun Wang. Robust detection of false data injection attacks for data aggregation in an Internet of Things-based environmental surveillance. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):410–428, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302372.

Li:2017:AMA

[2362] Xiong Li, Maged Hamada Ibrahim, Saru Kumari, Arun Kumar Sangaiah, Vidushi Gupta, and Kim-Kwang Raymond Choo. Anonymous mutual authentication and key agreement scheme for wearable sensors in wireless body area networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??):429– 443, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617301044.

Yaqoob:2017:RRE

[2363] Ibrar Yaqoob, Ejaz Ahmed, Muhammad Habib ur Rehman, Abdelmuttlib Ibrahim Abdalla Ahmed, Mohammed Ali Al-garadi, Muhammad Imran, and Mohsen Guizani. The rise of ransomware and emerging security challenges in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):444–458, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303468.

Ahmed:2017:RBD

[2364] Ejaz Ahmed, Ibrar Yaqoob, Ibrahim Abaker Targio Hashem, Imran Khan, Abdelmuttlib Ibrahim Abdalla Ahmed, Muhammad Imran, and Athanasios V. Vasilakos. The role of big data analytics in Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??):459– 471, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302591.

Chen:2017:PIP

[2365] Chi-Yuan Chen, Hsin-Min Wu, Lei Wang, and Chia-Mu Yu. Practical integrity preservation for data streaming in cloud-assisted healthcare sensor systems. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):472–480, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302438.

Qiu:2017:SSR

[2366] Tie Qiu, Xize Liu, Min Han, Mingchu Li, and Yushuang Zhang. SRTS: a Self-Recoverable Time Synchronization for sensor networks of healthcare IoT. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??): 481–492, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302037.

Zhou:2017:EEG

[2367] Zhenyu Zhou, Junhao Feng, Yunjian Jia, Shahid Mumtaz, Kazi Mohammed Saidul Huq, Jonathan Rodriguez, and Di Zhang. Energyefficient game-theoretical random access for M2M communications in overlapped cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??):493– 501, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302633.

Zhao:2017:HIE

[2368] Long Zhao, Kan Zheng, and Periklis Chatzimisios. Hybrid information and energy transfer in ultra-dense Het-Nets. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):502–509, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730124X.

DeBenedetto:2017:PDD

[2369] Jacopo De Benedetto, Paolo Bellavista, and Luca Foschini. Proximity discovery and data dissemination for mobile crowd sensing using LTE direct. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??): 510–521, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303092.

Wang:2017:MEC

[2370] Bang Wang, Qiang Yang, Laurence T. Yang, and Chunsheng Zhu. On minimizing energy consumption cost in green heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):522–535, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617301238.

Lin:2017:MEP

[2371] Chun-Cheng Lin, Ching-Tsorng Tsai, Der-Jiunn Deng, I-Hsin Tsai, and Shun-Yu Jhong. Minimizing electromagnetic pollution and power consumption in green heterogeneous small cell network deployment. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??):536– 547, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617302244.

Harbouche:2017:MDF

[2372] Ahmed Harbouche, Noureddine Djedi, Mohammed Erradi, Jalel Ben-Othman, and Abdellatif Kobbane. Model driven flexible design of a wireless body sensor network for health monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??):

548–571, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730258X.

Borah:2017:GTC

[2373] Satya J. Borah, Sanjay Kumar Dhurandher, Isaac Woungang, and Vinesh Kumar. A game theoretic context-based routing protocol for opportunistic networks in an IoT scenario. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2) (??):572–584, December 24, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617302864.

Anonymous:2017:EBs

[2374] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 129 (part 2)(??): ifc, December 24, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861730405X.

Anonymous:2018:EBa

[2375] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 130(??):ii, January 15, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617304346.

Akyildiz:2018:EC

[2376] Ian F. Akyildiz, Harry Rudin, and Burkhard Stiller. Editorial for COM-NET 2017. Computer Networks (Amsterdam, Netherlands: 1999), 130 (??):iii, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304371.

Gunathillake:2018:TMA

[2377] Ashanie Gunathillake, Andrey V. Savkin, and Anura P. Jayasumana. Topology mapping algorithm for 2D and 3D wireless sensor networks based on maximum likelihood estimation. Computer Networks (Amsterdam, Netherlands: 1999), 130 (??):1–15, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303948.

Hafsaoui:2018:FGR

[2378] A. Hafsaoui, A. Dandoush, G. Urvoy-Keller, M. Siekkinen, and D. Collange. A fine-grained response time analysis technique in heterogeneous environments. Computer Networks (Amsterdam, Netherlands: 1999), 130 (??):16–33, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304140.

Abdel-Halim:2018:PBP

[2379] Islam Tharwat Abdel-Halim and Hossam Mahmoud Ahmed Fahmy. Prediction-based protocols for vehicular ad hoc networks: Survey and taxonomy. Computer Networks (Amsterdam, Netherlands: 1999), 130 (??):34–50, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303924.

Shu:2018:AIC

[2380] Ran Shu, Fengyuan Ren, Jiao Zhang, Tong Zhang, and Chuang Lin. Analysing and improving convergence of quantized congestion notification in data center Ethernet. Computer Networks (Amsterdam, Netherlands: 1999), 130(??): 51–64, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304024.

Marotta:2018:JPE

[2381] Antonio Marotta, Stefano Avallone, and Andreas Kassler. A joint power efficient server and network consolidation approach for virtualized data centers. Computer Networks (Amsterdam, Netherlands: 1999), 130 (??):65–80, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304012.

Karthikeya:2018:PAG

[2382] Surabhi Abhimithra Karthikeya, Revathy Narayanan, and Siva Ram Murthy C. Power-aware gateway connectivity in battery-powered dynamic IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 130 (??):81–93, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303936.

Bilal:2018:PTP

[2383] Kashif Bilal, Osman Khalid, Aiman Erbad, and Samee U. Khan. Potentials, trends, and prospects in edge technologies: Fog, cloudlet, mobile edge, and micro data centers. Computer Networks (Amsterdam, Netherlands: 1999), 130(??): 94–120, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617303778.

Roy:2018:QSD

[2384] Animesh Roy, Tamaghna Acharya, and Sipra DasBit. Quality of service in delay tolerant networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 130(??): 121–133, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617304188.

Seetharam:2018:GFH

[2385] Anand Seetharam and Arti Ramesh. On the goodput of flows in heterogeneous mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 130(??):134–144, January 15, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128617304164.

Ayatollahi:2018:PPN

[2386] Hoda Ayatollahi, Mohammad Khansari, and Hamid R. Rabiee. A push-pull network coding protocol for live peer-to-peer streaming. Computer Networks (Amsterdam, Netherlands: 1999), 130 (??):145–155, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304152.

Meloni:2018:CBI

[2387] A. Meloni, P. A. Pegoraro, L. Atzori, A. Benigni, and S. Sulis. Cloudbased IoT solution for state estimation in smart grids: Exploiting virtualization and edge-intelligence technologies. Computer Networks (Amsterdam, Netherlands: 1999), 130(??): 156–165, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617303912.

Anonymous:2018:EBb

[2388] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 131(??):ii, February 11, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1389128618300185.

Chu:2018:JCR

[2389] Weibo Chu, Mostafa Dehghan, John C. S. Lui, Don Towsley, and Zhi-Li Zhang. Joint cache resource allocation and request routing for in-network caching services. Computer Networks (Amsterdam, Netherlands: 1999), 131 (??):1–14, February 11, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1389128617304176.

Ghebleh:2018:CCI

[2390] Reza Ghebleh. A comparative classification of information dissemination approaches in vehicular ad hoc networks from distinctive viewpoints: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 131 (??):15–37, February 11, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617304218.

Wang:2018:EEI

[2391] Xingwei Wang, Jianhui Lv, Min Huang, Keqin Li, Jie Li, and Kexin Ren. Energy-efficient ICN routing mechanism with QoS support. Computer Networks (Amsterdam, Netherlands: 1999), 131(??): 38–51, February 11, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617304206.

Kang:2018:ORO

[2392] Daeho Kang, Hyung-Sin Kim, Changhee Joo, and Saewoong Bahk. ORGMA: Reliable opportunistic routing with gradient forwarding for MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 131(??):52– 64, February 11, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S138912861730419X.

Jia:2018:IPC

[2393] Xuya Jia, Yong Jiang, Zehua Guo, Gengbiao Shen, and Lei Wang. Intelligent path control for energysaving in hybrid SDN networks. Computer Networks (Amsterdam, Netherlands: 1999), 131(??):65– 76, February 11, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S138912861730422X.

Nardini:2018:SDP

[2394] Giovanni Nardini, Giovanni Stea, and Antonio Virdis. A scalable data-plane architecture for one-to-one deviceto-device communications in LTE-Advanced. Computer Networks (Amsterdam, Netherlands: 1999), 131 (??):77–95, February 11, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1389128617304243.

Wang:2018:HIM

[2395] Zhibo Wang, Ran Tan, Jiahui Hu, Jing Zhao, Qian Wang, Feng Xia, and Xiaoguang Niu. Heterogeneous incentive mechanism for time-sensitive and location-dependent crowdsensing networks with random arrivals. Computer Networks (Amsterdam, Netherlands: 1999), 131(??): 96–109, February 11, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1389128617304280.

Wang:2018:MMW

[2396] Xiaonan Wang, Zhengxiong Dou, Dong Wang, and Qi Sun. Mobility management for 6LoWPAN WSN. Computer Networks (Amsterdam, Netherlands: 1999), 131(??): 110–128, February 11, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617304231.

Castro-Jul:2018:CAU

[2397] Fátima Castro-Jul, Rebeca P. Díaz-Redondo, and Ana Fernández-Vilas. Collaboratively assessing urban alerts in ad hoc participatory sensing. Computer Networks (Amsterdam, Netherlands: 1999), 131(??):129– 143, February 11, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1389128617304267.

Chattaraj:2018:NTS

[2398] Durbadal Chattaraj, Monalisa Sarma, and Ashok Kumar Das. A new twoserver authentication and key agreement protocol for accessing secure cloud services. Computer Networks (Amsterdam, Netherlands: 1999), 131 (??):144–164, February 11, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128617304255.

Anonymous:2018:EBc

[2399] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 132(??):ii, February 26, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618300616.

Cai:2018:TIMa

[2400] Hui Cai, Yanmin Zhu, and Zhenni Feng. A truthful incentive mechanism for mobile crowd sensing with location-sensitive weighted tasks. Computer Networks (Amsterdam, Netherlands: 1999), 132(??):1–14, February 26, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128617304309.

He:2018:UFC

[2401] Junxiao He, Oliver Yang, and Yifeng Zhou. Using flow cost to globally allocate and optimize limited bandwidth in multipath routing. Computer Networks (Amsterdam, Netherlands: 1999), 132 (??):15–25, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304279.

Huang:2018:OUM

[2402] Meitian Huang, Weifa Liang, Zichuan Xu, Wenzheng Xu, Song Guo, and Yinlong Xu. Online unicasting and multicasting in software-defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 132(??): 26–39, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304292.

Zhan:2018:IMP

[2403] Yufeng Zhan, Yuanqing Xia, and Jinhui Zhang. Incentive mechanism in platform-centric mobile crowdsensing: a one-to-many bargaining approach. Computer Networks (Amsterdam, Netherlands: 1999), 132 (??):40–52, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304401.

Nur:2018:CXI

[2404] Abdullah Yasin Nur and Mehmet Engin Tozal. Cross-AS (X-AS) Internet topology mapping. Computer Networks (Amsterdam, Netherlands: 1999), 132 (??):53–67, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300124.

Bote-Lorenzo:2018:OML

[2405] Miguel L. Bote-Lorenzo, Eduardo Gómez-Sánchez, Carlos Mediavilla-Pastor, and Juan I. Asensio-Pérez. Online machine learning algorithms to predict link quality in community wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 132(??):68–80, February 26, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618300069.

Shi:2018:EFG

[2406] Hongtao Shi, Hongping Li, Dan Zhang, Chaqiu Cheng, and Xuanxuan Cao. An efficient feature generation approach based on deep learning and feature selection techniques for traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 132 (??):81–98, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300082.

Kobayashi:2018:TSH

[2407] Hideo Kobayashi, Eiichi Kameda, Yoshiaki Terashima, and Norihiko Shinomiya. Towards sustainable heterogeneous wireless networks: a decision strategy for AP selection with dynamic graphs. Computer Networks (Amsterdam, Netherlands: 1999), 132 (??):99–107, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300094.

Xiao:2018:SSH

[2408] Fu Xiao, Qianwen Miao, Xiaohui Xie, Lijuan Sun, and Ruchuan Wang. SHMO: a seniors health monitoring system based on energy-free sensing. Computer Networks (Amsterdam, Netherlands: 1999), 132(??): 108–117, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300033.

Li:2018:ALI

[2409] Yue Li and Luigi Iannone. Assessing Locator/Identifier Separation Protocol interworking performance through RIPE Atlas. Computer Networks (Amsterdam, Netherlands: 1999), 132(??): 118–128, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128617304413.

Karatas:2018:ODH

[2410] Mumtaz Karatas. Optimal deployment of heterogeneous sensor networks for a hybrid point and barrier coverage application. Computer Networks (Amsterdam, Netherlands: 1999), 132 (??):129–144, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300021.

Thieu:2018:OPC

[2411] Quang-Tuan Thieu and Hung-Yun Hsieh. Outage protection for cellularmode users in device-to-device communications through stochastic optimization. Computer Networks (Amsterdam, Netherlands: 1999), 132(??): 145–160, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300070.

Benmalek:2018:VVS

[2412] Mourad Benmalek, Yacine Challal, Abdelouahid Derhab, and Abdelmadjid Bouabdallah. VerSAMI: Versatile and scalable key management for Smart Grid AMI systems. Computer Networks (Amsterdam, Netherlands: 1999), 132(??): 161–179, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300112.

Hadi:2018:BDA

[2413] Mohammed S. Hadi, Ahmed Q. Lawey, Taisir E. H. El-Gorashi, and Jaafar M. H. Elmirghani. Big data analytics for wireless and wired network design: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 132 (??):180–199, February 26, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300239.

Anonymous:2018:EBd

[2414] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 133(??):ii, March 14, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618300926.

Gupta:2018:SCS

[2415] Abhishek Gupta, M. Farhan Habib, Uttam Mandal, Pulak Chowdhury, Massimo Tornatore, and Biswanath Mukherjee. On service-chaining strategies using Virtual Network Functions in operator networks. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):1–16, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300379.

Carmona-Murillo:2018:HDS

[2416] J. Carmona-Murillo, V. Friderikos, and J. L. González-Sánchez. A hybrid DMM solution and trade-off analysis for future wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 133(??):17–32, March 14, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300392.

Li:2018:HAM

[2417] Wei Li, Changxin Huang, Chao Xiao, and Songchen Han. A heading adjustment method in wireless directional sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):33–41, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300410.

Arghavani:2018:AMG

[2418] Abbas Arghavani, Mahdi Arghavani, Mahmood Ahmadi, and Paul Crane. Attacker-Manager Game Tree (AMGT): A new framework for visualizing and analysing the interactions between attacker and network security manager. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):42–58, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300148.

Shen:2018:PPS

[2419] Yiran Shen, Chengwen Luo, Dan Yin, Hongkai Wen, Rus Daniela, and Wen Hu. Privacy-preserving sparse representation classification in cloudenabled mobile applications. Computer Networks (Amsterdam, Netherlands: 1999), 133(??):59–72, March 14, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300483.

Borylo:2018:SAH

[2420] Piotr Borylo, Jerzy Domzal, and Robert Wójcik. Survivable automatic hidden bypasses in Software-Defined Networks. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):73–89, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830032X.

Ayad:2018:PER

[2421] Ibrahim Ayad, Youngbin Im, Eric Keller, and Sangtae Ha. A practical evaluation of rate adaptation algorithms in HTTP-based adaptive streaming. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):90–103, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300288.

Xu:2018:JCA

[2422] Jun Xu, Chengcheng Guo, and Hao Zhang. Joint channel allocation and power control based on PSO for cellular networks with D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):104–119, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300264.

Xu:2018:CMU

[2423] Chaonong Xu, Kaichi Ma, and Yongjun Xu. Complexity of minimum uplink scheduling in backbone-assisted successive interference cancellation-based wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):120–129, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300215.

Wang:2018:EER

[2424] Tian Wang, Jiandian Zeng, MD Zakirul Alam Bhuiyan, Yonghong Chen, Yiqiao Cao, Hui Tian, and Mande Xie. Energy-efficient relay tracking with multiple mobile camera sensors. Computer Networks (Amsterdam, Netherlands: 1999), 133(??): 130–140, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300045.

Belguith:2018:PSO

[2425] Sana Belguith, Nesrine Kaaniche, Maryline Laurent, Abderrazak Jemai, and Rabah Attia. PHOABE: Securely outsourcing multi-authority attribute based encryption with policy hidden for cloud assisted IoT. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):141–156, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300495.

Cui:2018:EEC

[2426] Hui Cui, Robert H. Deng, Junzuo Lai, Xun Yi, and Surya Nepal. An efficient and expressive ciphertext-policy attribute-based encryption scheme with partially hidden access structures, revisited. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):157–165, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830046X.

Chen:2018:MMA

[2427] Huan Chen, Xiong Wang, Yangming Zhao, Tongyu Song, Yang Wang, Shizhong Xu, and Lemin Li. MOSC: a method to assign the outsourcing of service function chain across multiple clouds. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):166–182, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830029X.

Wang:2018:SGP

[2428] Jessie Hui Wang and Changqing An.

A study on geographic properties of Internet routing. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):183–194, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300422.

Moazzeni:2018:RIS

[2429] Shadi Moazzeni, Mohammad Reza Khayyambashi, Naser Movahhedinia, and Franco Callegati. On reliability improvement of Software-Defined Networks. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):195–211, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300343.

Yi:2018:CSN

[2430] Bo Yi, Xingwei Wang, Keqin Li, Sajal K. Das, and Min Huang. A comprehensive survey of Network Function Virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 133 (??):212–262, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300306.

Anonymous:2018:EBe

[2431] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 134(??):ii, April 7, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301063.

Martinez-Chavez:2018:CRS

[2432] Sergio M. Martínez-Chavez, Mario E. Rivero-Angeles, Laura I. Garay-Jimenez, and Elena F. Ruiz-Ledesma. Cognitive radio system for interference reduction in BANETs focused on epilepsy diagnosis. Computer Networks (Amsterdam, Netherlands: 1999), 134(??):1–22, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300100.

Stergiopoulos:2018:UFD

[2433] G. Stergiopoulos, D. Gritzalis, and V. Kouktzoglou. Using formal distributions for threat likelihood estimation in cloud-enabled IT risk assessment. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):23–45, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300446.

Lopez:2018:ACC

[2434] Javier Lopez and Juan E. Rubio. Access control for cyber-physical systems interconnected to the cloud. Computer Networks (Amsterdam, Netherlands: 1999), 134(??):46–54, April 7, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300501.

Detti:2018:MLC

[2435] Andrea Detti, Lorenzo Bracciale, Pierpaolo Loreti, and Nicola Blefari Melazzi. Modeling LRU cache with invalidation. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):55–65, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300380.

Cheng:2018:TMD

[2436] Long Cheng, Jianwei Niu, Chengwen Luo, Lei Shu, Linghe Kong, Zhiwei Zhao, and Yu Gu. Towards minimum-delay and energy-efficient flooding in low-duty-cycle wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):66–77, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300136.

Pournaghi:2018:NNE

[2437] Seyed Morteza Pournaghi, Behnam Zahednejad, Majid Bayat, and Yaghoub Farjami. NECPPA: A novel and efficient conditional privacy-preserving authentication scheme for VANET. Computer Networks (Amsterdam, Netherlands: 1999), 134(??):78– 92, April 7, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300227.

Malandrino:2018:VBE

[2438] Francesco Malandrino, Carla-Fabiana Chiasserini, and Claudio Casetti. Virtualization-based evaluation of backhaul performance in vehicular applications. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):93–104, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300276.

Satoh:2018:RCD

[2439] Daisuke Satoh, Yuji Takano, Ryunosuke Sudo, and Takemi Mochida. Reduction of communication demand under disaster congestion using control to change human communication behavior without direct restriction. Computer Networks (Amsterdam, Netherlands: 1999), 134(??): 105–115, April 7, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300641.

Weng:2018:MUP

[2440] Xiaoxiong Weng, Yongxin Liu, Houbing Song, Shushen Yao, and Pengfei Zhang. Mining urban passengers’ travel patterns from incomplete data with use cases. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):116–126, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300574.

Collotta:2018:FBA

[2441] Mario Collotta, Giovanni Pau, and Daniel G. Costa. A fuzzy-based approach for energy-efficient Wi-Fi communications in dense wireless multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 134(??):127–139, April 7, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300409.

Demirci:2018:CBA

[2442] Ilhan Demirci and Ömer Korçak. Cell breathing algorithms for load balancing in Wi-Fi/cellular heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):140–151, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300549.

Naman:2018:RHT

[2443] Aous Thabit Naman, Yu Wang, Hassan Habibi Gharakheili, Vijay Sivaraman, and David Taubman. Responsive high throughput congestion control for interactive applications over SDNenabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):152–166, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300525.

Feng:2018:ALA

[2444] Wei Feng, Yu Qin, Shijun Zhao, and Dengguo Feng. AAoT: Lightweight attestation and authentication of low-resource things in IoT and CPS. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):167–182, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300471.

Marnerides:2018:ITC

[2445] A. K. Marnerides, D. P. Pezaros, and D. Hutchison. Internet traffic characterisation: Third-order statistics and higher-order spectra for precise traffic modelling. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):183–201, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300677.

Li:2018:STM

[2446] Quanzhong Li and Sai Zhao. Secure transmission for multi-antenna wireless powered communication with cochannel interference and self-energy recycling. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):202–214, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300719.

Guntupalli:2018:PFT

[2447] Lakshmikanth Guntupalli, Jorge Martinez-Bauset, and Frank Y. Li. Performance of frame transmissions and event-triggered sleeping in dutycycled WSNs with error-prone wireless links. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):215–227, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300562.

Kunst:2018:INR

[2448] Rafael Kunst, Leandro Avila, Edison Pignaton, Sergio Bampi, and Juergen Rochol. Improving network resources allocation in smart cities video surveillance. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):228–244, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300513.

Siracusano:2018:FEI

[2449] G. Siracusano, S. Salsano, P. L. Ventre, A. Detti, O. Rashed, and N. Blefari-Melazzi. A framework for experimenting ICN over SDN solutions using physical and virtual testbeds. Computer Networks (Amsterdam, Netherlands: 1999), 134(??):245–259, April 7, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300367.

Zhao:2018:ODD

[2450] Yanxiao Zhao, MD Nashid Anjum, Lina Pu, Guodong Wang, and Yu Luo. Optimal on demand delay-constrained fair distribution for self-coexistence WRAN. Computer Networks (Amsterdam, Netherlands: 1999), 134 (??):260–271, April 7, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300689.

Bennis:2018:EQS

[2451] Ismail Bennis, Hacène Fouchal, Kandaraj Piamrat, and Marwane Ayaida. Efficient queuing scheme through cross-layer approach for multimedia transmission over WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 134(??):272–282, April 7, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300550.

Chaudhuri:2018:NQA

[2452] Saptarshi Chaudhuri, Irfan Baig, and Debabrata Das. A novel QoS aware medium access control scheduler for LTE-advanced network. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):1–14, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300355.

Xia:2018:BFN

[2453] Hui Xia, Bin Fang, Matthew Roughan, Kenjiro Cho, and Paul Tune. A BasisEvolution framework for network traffic anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):15–31, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300331.

Wang:2018:TIM

[2454] Yingjie Wang, Zhipeng Cai, Xiangrong Tong, Yang Gao, and Guisheng Yin. Truthful incentive mechanism with location privacy-preserving for mobile crowdsourcing systems. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):32–43, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300756.

Mozaffari:2018:PDE

[2455] Mohiyeddin Mozaffari, Behrouz Safarinejadian, and Mokhtar Shasadeghi. Probability density estimation in sensor networks based on distributed mixture of factor analyzers, mobile agents and stochastic sensor selection. Computer Networks (Amsterdam, Netherlands: 1999), 135(??): 44–53, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830077X.

Filho:2018:RTS

[2456] Geraldo P. R. Filho, Leandro A. Villas, Heitor Freitas, Alan Valejo, Daniel L. Guidoni, and Jó Ueyama. ResiDI: Towards a smarter smart home system for decision-making using wireless sensors and actuators. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):54–69, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300768.

Ciavarrini:2018:GIH

[2457] Gloria Ciavarrini, Maria S. Greco, and Alessio Vecchio. Geolocation of Internet hosts: Accuracy limits through Cramér–Rao lower bound. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):70–80, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300744.

Karakus:2018:EVS

[2458] Murat Karakus and Arjan Durresi. Economic viability of Software Defined Networking (SDN). Computer Networks (Amsterdam, Netherlands: 1999), 135(??):81–95, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300823.

Schiller:2018:CMN

[2459] E. Schiller, N. Nikaein, E. Kalogeiton, M. Gasparyan, and T. Braun. CDS-MEC: NFV/SDN-based application management for MEC in 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):96–107, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830080X.

Konstantopoulos:2018:EME

[2460] Charalampos Konstantopoulos, Nikolaos Vathis, Grammati Pantziou, and Damianos Gavalas. Employing mobile elements for delay-constrained data gathering in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):108–131, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300732.

Merwaday:2018:ISS

[2461] Arvind Merwaday, Murat Yuksel, Thomas Quint, Ismail Güvenç, Walid Saad, and Naim Kapucu. Incentivizing spectrum sharing via subsidy regulations for future wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):132–146, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300781.

Piedrahita:2018:VIR

[2462] Andrés F. Murillo Piedrahita, Vikram Gaur, Jairo Giraldo, Alvaro A. Cardenas, and Sandra Julieta Rueda. Virtual incident response functions in control systems. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):147–159, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300434.

Khallef:2018:IFE

[2463] Walid Khallef, Sylvain Durand, and Miklós Molnár. ILP formulation of the exact solution of multi-constrained minimum cost multicast. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):160–170, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300835.

Boukerche:2018:VCC

[2464] Azzedine Boukerche and Robson E. De Grande. Vehicular cloud computing: Architectures, applications, and mobility. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):171–189, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300057.

Zhao:2018:CDS

[2465] Haitao Zhao, Lingchu Mao, and Jibo Wei. Coverage on demand: A simple motion control algorithm for autonomous robotic sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):190–200, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618300720.

Rekik:2018:ATA

[2466] Sana Rekik, Nouha Baccour, Mohamed Jmaiel, Khalil Drira, and Luigi Alfredo Grieco. Autonomous and traffic-aware scheduling for TSCH networks. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):201–212, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300963.

Falcao:2018:FBM

[2467] Marcos R. M. Falcão, Andson M.

Balieiro, and Kelvin L. Dias. A flexible-bandwidth model with channel reservation and channel aggregation for three-layered Cognitive Radio Networks. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):213–225, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300690.

Liyanage:2018:CPO

[2468] Kushan Sudheera Kalupahana Liyanage, Maode Ma, and Peter Han Joo Chong. Controller placement optimization in hierarchical distributed software defined vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):226–239, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300951.

Morcos:2018:TLA

[2469] Mira Morcos, Tijani Chahed, Lin Chen, Jocelyne Elias, and Fabio Martignon.

A two-level auction for resource allocation in multi-tenant C-RAN. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):240–252, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300707.

Hager:2018:RTB

[2470] Sven Hager, Patrik John, Stefan Dietzel, and Björn Scheuermann. RuleBender: Tree-based policy transformations for practical packet classification systems. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):253–265, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300860.

Zhang:2018:OTP

[2471] Heng Zhang and Wei Xing Zheng. Optimum transmission policy for remote state estimation with opportunistic energy harvesting. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):266–274, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300884.

Cetto:2018:TSI

[2472] Alexandra Cetto, Mathias Klier, Alexander Richter, and Jan Felix Zolitschka. “Thanks for sharing” — identifying users’ roles based on knowledge contribution in enterprise social networks. Computer Networks (Amsterdam, Netherlands: 1999), 135 (??):275–288, April 22, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300793.

Suarez-Varela:2018:FMS

[2473] José Suárez-Varela and Pere Barlet-Ros. Flow monitoring in Software-Defined Networks: Finding the accuracy/performance tradeoffs. Computer Networks (Amsterdam, Netherlands: 1999), 135(??):289–301, April 22, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300872.

Anonymous:2018:EBf

[2474] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 136(??):ii, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301403.

Afek:2018:DHF

[2475] Yehuda Afek, Anat Bremler-Barr, Shir Landau Feibish, and Liron Schiff. Detecting heavy flows in the SDN match and action model. Computer Networks (Amsterdam, Netherlands: 1999), 136(??):1–12, May 8, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300859.

Sepulcre:2018:CAH

[2476] Miguel Sepulcre and Javier Gozalvez. Context-aware heterogeneous V2X communications for connected vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 136 (??):13–21, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300975.

Kuzniar:2018:MMA

[2477] Maciej Kuźniar, Peter Peresíni, Dejan Kostić, and Marco Canini. Methodology, measurement and analysis of flow table update characteristics in hardware OpenFlow switches. Computer Networks (Amsterdam, Netherlands: 1999), 136(??):22–36, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300811.

Hajisalem:2018:HID

[2478] Vajiheh Hajisalem and Shahram Babaie. A hybrid intrusion detection system based on ABC-AFS algorithm for misuse and anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 136 (??):37–50, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301014.

Cerrato:2018:EPT

[2479] Ivano Cerrato and Fulvio Risso. Enabling precise traffic filtering based on protocol encapsulation rules. Computer Networks (Amsterdam, Netherlands: 1999), 136(??):51–67, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301002.

Yao:2018:EIA

[2480] Junmei Yao, Wei Lou, Chao Yang, and Kaishun Wu. Efficient interferenceaware power control for wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 136 (??):68–79, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300847.

Chen:2018:PSE

[2481] Zhaomin Chen, Chai Kiat Yeo, Bu Sung Lee, and Chiew Tong Lau. Power spectrum entropy based detection and mitigation of low-rate DoS attacks. Computer Networks (Amsterdam, Netherlands: 1999), 136 (??):80–94, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301026.

Yang:2018:DDN

[2482] Tingting Yang, Xiang Long, Arun Kumar Sangaiah, Zhigao Zheng, and Chao Tong. Deep detection network for real-life traffic sign in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 136 (??):95–104, May 8, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300999.

Hadiwardoyo:2018:ECU

[2483] Seilendria A. Hadiwardoyo, Enrique Hernández-Orallo, Carlos T. Calafate, Juan Carlos Cano, and Pietro Manzoni. Experimental characterization of UAVto-car communications. Computer Networks (Amsterdam, Netherlands: 1999), 136(??):105–118, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301105.

Li:2018:MLS

[2484] Ming Li, Joseph Reeves, and Carlos Moreno. Multi-level sample importance ranking based progressive transmission strategy for time series body sensor data. Computer Networks (Amsterdam, Netherlands: 1999), 136 (??):119–127, May 8, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301099.

Yang:2018:CCN

[2485] Siqian Yang, Cheng Wang, and Changjun Jiang. Centron: Cooperative neighbor discovery in mobile adhoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 136 (??):128–136, May 8, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301117.

Semerci:2018:ICS

[2486] Murat Semerci, Ali Taylan Cemgil, and Bülent Sankur. An intelligent cyber security system against DDoS attacks in SIP networks. Computer Networks (Amsterdam, Netherlands: 1999), 136(??):137–154, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300987.

Xu:2018:AAB

[2487] Lijie Xu, Geng Yang, Jia Xu, Lei Wang, and Haipeng Dai. Achieving adaptive broadcasting performance tradeoff for energy-critical sensor networks: a bottom-up approach. Computer Networks (Amsterdam, Netherlands: 1999), 136(??):155–170, May 8, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301154.

Anonymous:2018:EBg

[2488] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 137(??):ii, June 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301658.

Lin:2018:FLT

[2489] Yu-Ting Lin, Thomas Bonald, and Salah Eddine Elayoubi. Flow-level traffic model for adaptive streaming services in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 137(??):1–16, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300318.

Shinkuma:2018:UIM

[2490] Ryoichi Shinkuma, Yusuke Tanaka, Yoshinobu Yamada, Eiji Takahashi, and Takeo Onishi. User instruction mechanism for temporal traffic smoothing in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 137(??):17–26, June 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301166.

Javed:2018:TBS

[2491] Muhammad Awais Javed, Sherali Zeadally, and Zara Hamid. Trust-based security adaptation mechanism for Vehicular Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 137(??):27–36, June 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830118X.

Sicari:2018:RRD

[2492] Sabrina Sicari, Alessandra Rizzardi, Daniele Miorandi, and Alberto Coen-Porisini. REATO: REActing TO Denial of Service attacks in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):37–48, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301348.

Xin:2018:TOL

[2493] ChunSheng Xin, Sharif Ullah, Min Song, Zhao Wu, Qiong Gu, and Huanqing Cui. Throughput oriented lightweight near-optimal rendezvous algorithm for cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):49–60, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301178.

Zhu:2018:REH

[2494] Zhengyu Zhu, Zhongyong Wang, Zheng Chu, Di Zhang, and Byonghyo Shim. Robust energy harvest balancing optimization with V2X-SWIPT over MISO secrecy channel. Computer Networks (Amsterdam, Netherlands: 1999), 137(??):61–68, June 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301336.

Zareei:2018:EAD

[2495] Mahdi Zareei, Cesar Vargas-Rosales, Rafaela Villalpando-Hernandez, Leyre Azpilicueta, Mohammad Hossein Anisi, and Mubashir Husain Rehmani. The effects of an Adaptive and Distributed Transmission Power Control on the performance of energy harvesting sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):69–82, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301300.

Bhunia:2018:DAB

[2496] Suman Bhunia, Paulo Alexandre Regis, and Shamik Sengupta. Distributed adaptive beam nulling to survive against jamming in 3D UAV mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):83–97, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301191.

Rawat:2018:PCI

[2497] Danda B. Rawat, Reham Alsabet, Chandra Bajracharya, and Min Song. On the performance of cognitive internet-of-vehicles with unlicensed user-mobility and licensed useractivity. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):98–106, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301294.

Anand:2018:DCC

[2498] N. Anand, Sarath Babu, and B. S. Manoj. On detecting compromised controller in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):107–118, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830135X.

Kim:2018:WMY

[2499] Sungjin Kim, Jinkook Kim, Seokwoo Nam, and Dohoon Kim. WebMon: ML- and YARA-based malicious webpage detection. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):119–131, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301142.

Guan:2018:SOS

[2500] Yunchong Guan, Weimin Lei, Wei Zhang, Shaowei Liu, and Hao Li. Scalable orchestration of software defined service overlay network for multipath transmission. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):132–146, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301130.

Sadiq:2018:TPA

[2501] Ali Safa Sadiq, Suleman Khan, Kayhan Zrar Ghafoor, Mohsen Guizani, and Seyedali Mirjalili. Transmission power adaption scheme for improving IoV awareness exploiting: evaluation weighted matrix based on piggybacked information. Computer Networks (Amsterdam, Netherlands: 1999), 137 (??):147–159, June 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301324.

Guo:2018:JST

[2502] Wei Guo, V. Mahendran, and Sridhar Radhakrishnan. Join and spilt TCP for SDN networks: Architecture, implementation, and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 137(??):160–172, June 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301361.

Amer:2018:AOB

[2503] Mohammed Amer, Anthony Busson, and Isabelle Guérin Lassous. Association optimization based on access fairness for Wi-Fi networks. Computer Networks (Amsterdam, Netherlands: 1999), 137(??):173–188, June 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301129.

Anonymous:2018:EBh

[2504] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 138(??):ii, June 19, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301828.

Do:2018:CPS

[2505] Quang Do, Ben Martini, and Kim-Kwang Raymond Choo. Cyberphysical systems information gathering: A smart home case study. Computer Networks (Amsterdam, Netherlands: 1999), 138(??):1–12, June 19, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618301440.

Alcaraz:2018:SPC

[2506] Cristina Alcaraz, Xinyi Huang, and Erich Rome. Security and privacy in cloud-assisted cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):13–14, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301464.

Hausheer:2018:SPS

[2507] David Hausheer, Oliver Hohlfeld, Stefan Schmid, and Guofei Gu. Security and performance of softwaredefined networks and functions virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):15–17, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301452.

Safi:2018:SCB

[2508] Qamas Gul Khan Safi, Senlin Luo, Limin Pan, Wangtong Liu, Rasheed Hussain, and Safdar H. Bouk. SVPS: Cloud-based smart vehicle parking system over ubiquitous VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):18–30, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301531.

Attiah:2018:ERG

[2509] Afraa Attiah, Muhammad Faisal Amjad, Mainak Chatterjee, and Cliff Zou. An evolutionary routing game for energy balance in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):31–43, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830152X.

Mahmood:2018:LRD

[2510] Ali M. Mahmood, Adil Al-Yasiri, and Omar Y. Alani. Latency reduction by dynamic channel estimator selection in C-RAN networks using fuzzy logic. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):44–56, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301476.

Liu:2018:SAP

[2511] Yantao Liu and Yasser Morgan. Security against passive attacks on network coding system — a survey. Computer Networks (Amsterdam, Netherlands: 1999), 138(??):57–76, June 19, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830121X.

Zhang:2018:TMR

[2512] Jianwei Zhang, Xinchang Zhang, and Chunling Yang. Towards the multi-request mechanism in pullbased peer-to-peer live streaming systems. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):77–89, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301518.

Singh:2018:CCW

[2513] Karishma Singh, Karan Singh, Le Hoang Son, and Ahmed Aziz. Congestion control in wireless sensor networks by hybrid multi-objective optimization algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):90–107, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301439.

Cui:2018:ENP

[2514] Lin Cui, Fung Po Tso, and Weijia Jia. Enforcing network policy in heterogeneous network function box environment. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):108–118, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830149X.

Garcia-Magarino:2018:AAB

[2515] Iván García-Magariño, Guillermo Palacios-Navarro, Raquel Lacuesta, and Jaime Lloret. ABSCEV: An agentbased simulation framework about smart transportation for reducing waiting times in charging electric vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):119–135, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301282.

Arana:2018:MPL

[2516] O. Arana, F. Garcia, J. Gomez, and V. Rangel. MSP: Providing location privacy in WLAN networks with a MAC swapping protocol. Computer Networks (Amsterdam, Netherlands: 1999), 138(??):136–148, June 19, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301506.

Gao:2018:NNE

[2517] Xiaofeng Gao, Tao Chen, Zongchen Chen, and Guihai Chen. NEMO: Novel and efficient multicast routing schemes for hybrid data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):149–163, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301488.

Rohoden:2018:GTF

[2518] Katty Rohoden, Rebeca Estrada, Hadi Otrok, and Zbigniew Dziong. Game theoretical framework for clustering and resource allocation in macro-femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 138(??):164–176, June 19, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301555.

Ramos:2018:EPT

[2519] Javier Ramos, David Muelas, Jorge E. López de Vergara, and Javier Aracil. Estimation of the parameters of token-buckets in multi-hop environments. Computer Networks (Amsterdam, Netherlands: 1999), 138 (??):177–191, June 19, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301580.

Yang:2018:HEE

[2520] Qing Yang, Yiran Shen, Fengyuan Yang, Jianpei Zhang, Wanli Xue, and Hongkai Wen. HealCam: Energyefficient and privacy-preserving human vital cycles monitoring on cameraenabled smart devices. Computer Networks (Amsterdam, Netherlands: 1999), 138(??):192–200, June 19, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301543.

Bhat:2018:TDR

[2521] Shireesh Bhat, George N. Rouskas, and Iyad Katib. On time dependent routing algorithms for open marketplaces of path services with support for inadvance path reservation. Computer Networks (Amsterdam, Netherlands: 1999), 138(??):201–212, June 19, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301610.

Anonymous:2018:EBi

[2522] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 139(??):ii, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302160.

Baswade:2018:NCS

[2523] Anand M. Baswade, Touheed Anwar Atif, Bheemarjuna Reddy Tamma, and Antony Franklin. A novel coexistence scheme for IEEE 802.11 for user fairness and efficient spectrum utilization in the presence of LTE-U. Computer Networks (Amsterdam, Netherlands: 1999), 139(??):1–18, July 5, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301567.

Shadroo:2018:SSB

[2524] Shabnam Shadroo and Amir Masoud Rahmani. Systematic survey of big data and data mining in Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 139 (??):19–47, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301579.

Xu:2018:MTL

[2525] Dan Xu, Wenli Jiao, Zhuang Yin, Junjie Huang, Yao Peng, Xiaojiang Chen, Dingyi Fang, and Zhanyong Tang. Maximizing throughput for low duty-cycled sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 139(??):48–59, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301312.

Sasabe:2018:AOP

[2526] Masahiro Sasabe. Analysis of optimal piece flow in tit-for-tat-based P2P streaming. Computer Networks (Amsterdam, Netherlands: 1999), 139 (??):60–69, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301592.

Gao:2018:SCA

[2527] Lingnan Gao and George N. Rouskas.

A spectral clustering approach to network-aware virtual request partitioning. Computer Networks (Amsterdam, Netherlands: 1999), 139 (??):70–80, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301609.

Cao:2018:APL

[2528] Chenhong Cao, Yi Gao, Wei Dong, and Jiajun Bu. Accurate per-link loss tomography in dynamic sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 139 (??):81–91, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301683.

Hassine:2018:APB

[2529] Kawther Hassine, Mounir Frikha, and Tijani Chahed. Access point backhaul capacity aggregation as a matching game in the context of wireless local area networks. Computer Networks (Amsterdam, Netherlands: 1999), 139(??):92–108, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301737.

Khalid:2018:ADD

[2530] Aaqib Khalid, Tariq Umer, Muhammad Khalil Afzal, Sheraz Anjum, Adnan Sohail, and Hafiz Muhammad Asif. Autonomous data driven surveillance and rectification system using in-vehicle sensors for intelligent transportation systems (ITS). Computer Networks (Amsterdam, Netherlands: 1999), 139(??):109–118, July 5, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301713.

Wang:2018:EEE

[2531] Rui Wang, Zhiyong Zhang, Zhiwei Zhang, and Zhiping Jia. ETMRM: An energy-efficient trust management and routing mechanism for SD-WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 139 (??):119–135, July 5, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301725.

Krohnke:2018:RDN

[2532] Lars Kröhnke, Jelte Jansen, and Harald Vranken. Resilience of the Domain Name System: A case study of the .nl-domain. Computer Networks (Amsterdam, Netherlands: 1999), 139 (??):136–150, July 5, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301889.

Cerrato:2018:CCO

[2533] Ivano Cerrato, Fulvio Risso, Roberto Bonafiglia, Kostas Pentikousis, Gergely Pongrácz, and Hagen Woesner. COM-POSER: A compact open-source service platform. Computer Networks (Amsterdam, Netherlands: 1999), 139 (??):151–174, July 5, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301750.

Anonymous:2018:EBj

[2534] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 140(??):ii, July 20, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302901.

Biswas:2018:MPA

[2535] Pratima Biswas and Ashok Singh Sairam. Modeling privacy approaches for location based services. Computer Networks (Amsterdam, Netherlands: 1999), 140(??):1–14, July 20, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301890.

Chekkouri:2018:NIV

[2536] Ahmed Salim Chekkouri, Abdellatif Ezzouhairi, and Samuel Pierre. A new integrated VANET-LTE — a architecture for enhanced mobility in small cells HetNet using dynamic gateway and traffic forwarding. Computer Networks (Amsterdam, Netherlands: 1999), 140(??):15–27, July 20, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301762.

Peres:2018:MMA

[2537] Bruna Peres, Bruno P. Santos, Otavio A. de O. Souza, Olga Goussevskaia, Marcos A. M. Vieira, Luiz F. M. Vieira, and Antonio A. F.

Loureiro. Matrix: Multihop address allocation and dynamic any-to-any routing for 6LoWPAN. Computer Networks (Amsterdam, Netherlands: 1999), 140(??):28–40, July 20, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301774.

Patel:2018:LLA

[2538] Hasmukh Patel and Devesh C. Jinwala. LPM: A lightweight authenticated packet marking approach for IP traceback. Computer Networks (Amsterdam, Netherlands: 1999), 140 (??):41–50, July 20, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301786.

Koufoudakis:2018:PFC

[2539] George Koufoudakis, Konstantinos Oikonomou, Konstantinos Giannakis, and Sonia Aïssa. Probabilistic flooding coverage analysis for efficient information dissemination in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 140 (??):51–61, July 20, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830197X.

Hua:2018:TPT

[2540] Jingyu Hua, Laiping Zhao, Suohao Zhang, Yangyang Liu, Xin Ge, and Sheng Zhong. Topologypreserving traffic engineering for hierarchical multi-domain SDN. Computer Networks (Amsterdam, Netherlands:

1999), 140(??):62–77, July 20, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301749.

Banerjee:2018:GCO

[2541] Bitan Banerjee, Adita Kulkarni, and Anand Seetharam. Greedy caching: An optimized content placement strategy for information-centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 140 (??):78–91, July 20, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301907.

Li:2018:PAL

[2542] Guoquan Li, Ying Zhou, Tong Bai, Jinzhao Lin, Yu Pang, Wei Wu, Sadia Din, and Gwanggil Jeon. Performance analysis for low-complexity detection of MIMO V2V communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 140 (??):92–100, July 20, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830210X.

Lyazidi:2018:DRA

[2543] Mohammed Yazid Lyazidi, Nadjib Aitsaadi, and Rami Langar. A dynamic resource allocation framework in LTE downlink for cloud-radio access network. Computer Networks (Amsterdam, Netherlands: 1999), 140 (??):101–111, July 20, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830207X.

Jia:2018:PIE

[2544] Adele Lu Jia, Siqi Shen, Dongsheng Li, and Shengling Chen. Predicting the implicit and the explicit video popularity in a user generated content site with enhanced social features. Computer Networks (Amsterdam, Netherlands: 1999), 140(??):112–125, July 20, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301944.

Nasralla:2018:CAP

[2545] Moustafa M. Nasralla, Manzoor Razaak, Ikram U. Rehman, and Maria G. Martini. Content-aware packet scheduling strategy for medical ultrasound videos over LTE wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 140 (??):126–137, July 20, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618302342.

Koya:2018:AHM

[2546] Aneesh M. Koya and Deepthi P. P. Anonymous hybrid mutual authentication and key agreement scheme for wireless body area network. Computer Networks (Amsterdam, Netherlands: 1999), 140(??):138–151, July 20, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618302044.

Hashemi:2018:AMM

[2547] Seyyed Naser Seyyed Hashemi and Ali Bohlooli. Analytical modeling of multisource content delivery in informationcentric networks. Computer Networks (Amsterdam, Netherlands: 1999), 140 (??):152–162, July 20, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302056.

He:2018:LAB

[2548] Qian He, Ning Zhang, Yongzhuang Wei, and Yan Zhang. Lightweight attribute based encryption scheme for mobile cloud assisted cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 140 (??):163–173, July 20, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300458.

Cai:2018:TIMb

[2549] Hui Cai, Yanmin Zhu, Zhenni Feng, Hongzi Zhu, Jiadi Yu, and Jian Cao. Truthful incentive mechanisms for mobile crowd sensing with dynamic smartphones. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):ii, August 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302433.

Wang:2018:AJL

[2550] Tongxiang Wang, Xianglin Wei, Jianhua Fan, and Tao Liang. Adaptive jammer localization in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):1–16, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301920.

Anonymous:2018:EBk

[2551] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 141(??):1–234, August 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618304250.

Tan:2018:PMN

[2552] A. Serdar Tan and Engin Zeydan. Performance maximization of network assisted mobile data offloading with opportunistic device-to-device communications. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):17–30, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302214.

Li:2018:DSD

[2553] Yu Li, Xiaotian Wang, Dae Wook Kim, Junjie Zhang, and Rui Dai. Designing self-destructing wireless sensors with security and performance assurance. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):31–43, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302299.

Yang:2018:USC

[2554] Qing Yang, Paolo Gasti, Kiran Balagani, Yantao Li, and Gang Zhou. USB side-channel attack on Tor. Computer Networks (Amsterdam, Netherlands: 1999), 141(??):44–56, August 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302640.

Karimi:2018:PPG

[2555] Ramin Karimi and Saeed Shokrollahi.

PGRP: Predictive geographic routing protocol for VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 141(??):57–66, August 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302597.

Wang:2018:LBR

[2556] Haibo Wang, Hongli Xu, Liusheng Huang, Jianxin Wang, and Xuwei Yang. Load-balancing routing in software defined networks with multiple controllers. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):67–81, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302238.

Banerjee:2018:CSR

[2557] Anubhab Banerjee, Bitan Banerjee, Anand Seetharam, and Chintha Tellambura. Content search and routing under custodian unavailability in information-centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):82–91, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618302391.

Yang:2018:DOR

[2558] Yuan Yang and Mingwei Xu. Demandoblivious routing with planned link pruning. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):92–101, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301932.

Zhu:2018:MTI

[2559] Dali Zhu, Bobai Zhao, and Siye Wang. Mobile target indoor tracking based on Multi-Direction Weight Position Kalman Filter. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):102–114, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303013.

Shin:2018:TAP

[2560] Kyubo Shin, Hyoil Kim, Se Young Chun, and Donghoon Shin. TVWS assisted performance anomaly mitigation exploiting spectrum heterogeneity. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):115–127, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830269X.

Lv:2018:SFB

[2561] Cuicui Lv, Qiang Wang, Wenjie Yan, and Jia Li. A sparsity feedback-based data gathering algorithm for Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):128–144, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303128.

Tavana:2018:CCA

[2562] Morteza Tavana, Ali Rahmati, and Vahid Shah-Mansouri. Congestion control with adaptive access class barring for LTE M2M overload using Kalman filters. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):145–156, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618300537.

Tsoumanis:2018:EES

[2563] Georgios Tsoumanis, Konstantinos Oikonomou, George Koufoudakis, and Sonia Aïssa. Energy-efficient sink placement in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):157–165, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303244.

Wang:2018:EEO

[2564] Zhuo Wang, Xiaoning Feng, Guangjie Han, Yancheng Sui, and Hongde Qin. EODL: Energy Optimized Distributed Localization Method in threedimensional underwater acoustic sensors networks. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):166–178, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303190.

Xu:2018:SRA

[2565] Chi Xu, Chunhe Song, Peng Zeng, and Haibin Yu. Secure resource allocation for energy harvesting cognitive radio sensor networks without and with cooperative jamming. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):179–188, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303219.

Shi:2018:MMI

[2566] Heyuan Shi, Xibin Zhao, Hai Wan, Huihui Wang, Jian Dong, Kun Tang, and Anfeng Liu. Multi-model induced network for participatorysensing-based classification tasks in intelligent and connected transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):199–221, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303268.

Kouicem:2018:ITS

[2567] Djamel Eddine Kouicem, Abdelmadjid Bouabdallah, and Hicham Lakhlef. Internet of Things security: a topdown survey. Computer Networks (Amsterdam, Netherlands: 1999), 141 (??):222–233, August 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301208.

Mehr:2018:QMT

[2568] Kamal Adli Mehr, Javad Musevi Niya, and Nail Akar. Queue management for two-user cognitive radio with delayconstrained primary user. Computer Networks (Amsterdam, Netherlands: 1999), 142(??):ii, September 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303232.

Detti:2018:CBS

[2569] Andrea Detti, Lorenzo Bracciale, Pierpaolo Loreti, Giulio Rossi, and Nicola Blefari Melazzi. A clusterbased scalable router for information centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 142 (??):1–12, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303499.

Anonymous:2018:EBl

[2570] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 142(??):1–290, September 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618305498.

Sabuj:2018:TSP

[2571] Saifur Rahman Sabuj and Masanori Hamamura. Two-slope path-loss design of energy harvesting in random cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 142(??):13–23, September 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618303177.

Akdogan:2018:SKA

[2572] Dilara Akdogan, Duygu Karaoglan Altop, Laleh Eskandarian, and Albert Levi. Secure key agreement protocols: Pure biometrics and cancelable biometrics. Computer Networks (Amsterdam, Netherlands: 1999), 142 (??):24–32, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303372.

Yadav:2018:ALC

[2573] Ram Narayan Yadav and Rajiv Misra. Approximating the largest connected topology in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 142 (??):33–48, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303530.

Golchi:2018:EIP

[2574] Mahya Mohammadi Golchi and Homayun Motameni. Evaluation of the improved particle swarm optimization algorithm efficiency inward peer to peer video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 142 (??):49–63, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830344X.

Khatouni:2018:DCC

[2575] Ali Safari Khatouni, Marco Ajmone Marsan, Marco Mellia, and Reza Rejaie. Deadline-constrained content upload from multihomed devices: Formulations and algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 64–75, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303633.

Gao:2018:OOV

[2576] Meihui Gao, Bernardetta Addis, Mathieu Bouet, and Stefano Secci. Optimal orchestration of virtual network functions. Computer Networks (Amsterdam, Netherlands: 1999), 142 (??):76–92, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303578.

Bhoi:2018:ARP

[2577] Sourav Kumar Bhoi, Deepak Puthal, Pabitra Mohan Khilar, Joel J. P. C. Rodrigues, Sanjaya Kumar Panda, and Laurence T. Yang. Adaptive routing protocol for urban vehicular networks to support sellers and buyers on wheels. Computer Networks (Amsterdam, Netherlands: 1999), 142(??):

93–107, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303165.

Sadik:2018:SBH

[2578] Malak Sadik, Nadine Akkari, and Ghadah Aldabbagh. SDN-based handover scheme for multi-tier LTE/ Femto and D2D networks. Computer Networks (Amsterdam, Netherlands: 1999), 142(??):108–127, September 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618303475.

Mohamed:2018:EEC

[2579] Reem E. Mohamed, Walid R. Ghanem, Abeer T. Khalil, Mohamed Elhoseny, Muhammad Sajjad, and Mohamed A. Mohamed. Energy efficient collaborative proactive routing protocol for Wireless Sensor Network. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 128–141, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303748.

Li:2018:QRB

[2580] Weiwei Li, Petros Spachos, Mark Chignell, Alberto Leon-Garcia, Leon Zucherman, and Jie Jiang. A quantitative relationship between application performance metrics and quality of experience for over-the-top video. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 142–153, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302664.

Bradai:2018:RTE

[2581] Salma Bradai, Sofien Khemakhem, and Mohamed Jmaiel. Real-time and energy aware opportunistic mobile crowdsensing framework based on people’s connectivity habits. Computer Networks (Amsterdam, Netherlands: 1999), 142(??):154–167, September 4, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618303785.

Li:2018:MNG

[2582] Zhetao Li, YuXin Liu, Ming Ma, Anfeng Liu, Xiaozhi Zhang, and Gungming Luo. MSDG: a novel green data gathering scheme for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 179–193, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303736.

Rumipamba-Zambrano:2018:SDF

[2583] Rubén Rumipamba-Zambrano, Jordi Perelló, Joan M. Gené, and Salvatore Spadaro. On the scalability of dynamic Flex-Grid/SDM optical core networks. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 194–207, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303712.

Gamage:2018:ARA

[2584] Samoda Gamage, Jamil Y. Khan, and Duy T. Ngo. Adaptive resource allocation with traffic peak duration prediction and admission control for cognitive Wi-Fi networks. Computer Networks (Amsterdam, Netherlands: 1999), 142 (??):208–222, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304122.

Naz:2018:DCS

[2585] Sheneela Naz, Rao Naveed Bin Rais, Peer Azmat Shah, Sadaf Yasmin, Amir Qayyum, Seungmin Rho, and Yunyoung Nam. A dynamic caching strategy for CCN-based MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 223–239, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303207.

Zhao:2018:CRS

[2586] Wenbo Zhao, Xueyan Tang, and Luping Xu. Constructing routing structures for sensor data collection with dynamic traffic patterns. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 240–252, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303992.

Zeng:2018:PAU

[2587] Fanzi Zeng, Jisheng Xu, Yongfeng Li, Kang Li, and Lei Jiao. Performance analysis of underlay twoway relay cooperation in cognitive radio networks with energy harvesting. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 253–271, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303189.

Li:2018:RAC

[2588] Shenghong Li, Wei Ni, Chang Kyung Sung, and Mark Hedley. Recent advances on cooperative wireless localization and their application in inhomogeneous propagation environments. Computer Networks (Amsterdam, Netherlands: 1999), 142(??): 272–290, September 4, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304067.

Li:2018:TFV

[2589] Junfeng Li, Dan Li, Yirong Yu, Yukai Huang, Jing Zhu, and Jinkun Geng. Towards full virtualization of SDN infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 143(??):ii, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303852.

Xu:2018:TOA

[2590] Zichuan Xu, Weifa Liang, Alex Galis, Yu Ma, Qiufen Xia, and Wenzheng Xu. Throughput optimization for admitting NFV-enabled requests in cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):1–14, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303980.

Anonymous:2018:EBm

[2591] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 143(??):1–288, October 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618307837.

Han:2018:QSA

[2592] Pengchao Han, Yejun Liu, and Lei Guo. QoS satisfaction aware and network reconfiguration enabled resource allocation for virtual network embedding in fiber–wireless access network. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):15–29, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830450X.

Xie:2018:HBS

[2593] Weibin Xie, Fang Wang, Dan Feng, Lingling Zhang, Tingwei Zhu, and Qingyu Shi. Host-based scheduling: Achieving near-optimal transport for datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 143(??):30–48, October 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304523.

Sagirlar:2018:DPE

[2594] Gokhan Sagirlar, Barbara Carminati, and Elena Ferrari. Decentralizing privacy enforcement for Internet of Things smart objects. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):49–61, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305322.

Wang:2018:NSS

[2595] Junchao Wang, Kaining Han, Anastasios Alexandridis, Zeljko Zilic, Yu Pang, Wei Wu, Sadia Din, and Gwanggil Jeon. A novel security scheme for Body Area Networks compatible with smart vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 143(??):62–73, October 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304924.

Ferreira:2018:CSI

[2596] Joaquim Ferreira, Muhammad Alam, Bruno Fernandes, Luis Silva, João Almeida, Lara Moura, Rui Costa, Giovanni Iovino, and Elena Cordiviola. Cooperative sensing for improved traffic efficiency: the highway field trial. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):74–81, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304894.

Chen:2018:FFG

[2597] Sijia Chen, Bin Song, Jie Guo, Yanling Zhang, Xiaojiang Du, and Mohsen Guizani. FPAN: Fine-grained and progressive attention localization network for data retrieval. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):82–97, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304948.

Al-Hussaeni:2018:SDP

[2598] Khalil Al-Hussaeni, Benjamin C. M. Fung, Farkhund Iqbal, Gaby G. Dagher, and Eun G. Park. SafePath: Differentially-private publishing of passenger trajectories in transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):98–111, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304882.

Baziana:2018:EON

[2599] P. A. Baziana. Exploring the optimum number of receivers per station in WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):112–125, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305206.

Campolo:2018:RFS

[2600] Claudia Campolo, Antonella Molinaro, and Antonio Iera. A reference framework for social-enhanced Vehicle-to-Everything communications in 5G scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):126–139, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830495X.

Aujla:2018:EEV

[2601] Gagangeet Singh Aujla, Anish Jindal, and Neeraj Kumar. EVaaS: Electric vehicle-as-a-service for energy trading in SDN-enabled smart transportation system. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):140–152, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304900.

Li:2018:SFA

[2602] Ming Li, Dejun Yang, Jian Lin, Ming Li, and Jian Tang. SpecWatch: a framework for adversarial spectrum monitoring with unknown statistics. Computer Networks (Amsterdam, Netherlands: 1999), 143(??): 153–165, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305255.

Liu:2018:ELC

[2603] Chen Liu, Dingyi Fang, Yue Hu, Shensheng Tang, Dan Xu, Wen Cui, Xiaojiang Chen, Baoying Liu, Guangquan Xu, and Hao Chen. EasyGo: Low-cost and robust geographic opportunistic sensing routing in a strip topology wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):166–175, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304857.

Li:2018:SSV

[2604] Jie Li, Xingwei Wang, and Min Huang. SVDR: a scalable virtual domain-based routing scheme for CCN. Computer Networks (Amsterdam, Netherlands: 1999), 143(??): 176–190, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305036.

Riaz:2018:VFL

[2605] Faisal Riaz, Sania Khadim, Rabia Rauf, Mudassar Ahmad, Sohail Jabbar, and Junaid Chaudhry. A validated fuzzy logic inspired driver distraction evaluation system for road safety using artificial human driver emotion. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):191–205, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303591.

Khiati:2018:ALE

[2606] Mustapha Khiati and Djamel Djenouri. Adaptive learning-enforced broadcast policy for solar energy harvesting wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 143(??):206– 220, October 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305218.

Zhang:2018:BNM

[2607] Haibin Zhang, Jiajia Liu, and Ai-Chun Pang. A Bayesian network model for data losses and faults in medical body sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):221–246, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304936.

Murugesan:2018:HHP

[2608] Vijayalakshmi Murugesan, Mercy Shalinie Selvaraj, and Ming-Hour Yang. HP-SIPT: a high-precision single-packet IP traceback scheme. Computer Networks (Amsterdam, Netherlands: 1999), 143 (??):263–274, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305073.

Salman:2018:ISS

[2609] Ola Salman, Imad Elhajj, Ali Chehab, and Ayman Kayssi. IoT survey: an SDN and fog computing perspective. Computer Networks (Amsterdam, Netherlands: 1999), 143(??): 275–288, October 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305395.

Strayer:2018:CSM

[2610] Tim Strayer, Samuel Nelson, Amando Caro, Joud Khoury, Bryan Tedesco, Olivia DeRosa, Carsten Clark, Kolia Sadeghi, Michael Matthews, Jake Kurzer, Philip Lundrigan, Vikas Kawadia, Dorene Ryder, Keith Gremban, and Wayne Phoel. Content sharing with mobility in an infrastructureless environment. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):ii, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305632.

Chen:2018:LPD

[2611] Xi Chen, Chuanhe Huang, Xiying Fan, Di Liu, and Peng Li. LD-MAC: a propagation delay-aware MAC scheme for long-distance UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):1–16, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305905.

Anonymous:2018:EBn

[2612] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 144(??):1–296, October 24, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618308697.

Karray:2018:CSW

[2613] Fatma Karray, Mohamed W. Jmal, Alberto Garcia-Ortiz, Mohamed Abid, and Abdulfattah M. Obeid. A comprehensive survey on wireless sensor node hardware platforms. Computer Networks (Amsterdam, Netherlands: 1999), 144(??):17– 39, October 24, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302202.

Khorsandi:2018:BLA

[2614] Bahare M. Khorsandi and Carla Raffaelli. BBU location algorithms for survivable 5G C-RAN over WDM. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):40–52, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305991.

Han:2018:NIE

[2615] Jung-Han Han and Seung-Jae Han.

Non-intrusive estimation of available throughput for IEEE 802.11 link. Computer Networks (Amsterdam, Netherlands: 1999), 144(??): 53–63, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305760.

Wang:2018:IRM

[2616] Lei Wang, Qing Li, Richard Sinnott, Yong Jiang, and Jianping Wu. An intelligent rule management scheme for Software Defined Networking. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):64–76, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306248.

Aldwairi:2018:EPR

[2617] Tamer Aldwairi, Dilina Perera, and Mark A. Novotny. An evaluation of the performance of Restricted Boltzmann Machines as a model for anomaly network intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):77–88, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306005.

Habibzadeh:2018:SCS

[2618] Hadi Habibzadeh, Tolga Soyata, Burak Kantarci, Azzedine Boukerche, and Cem Kaptan. Sensing, communication and security planes: a new challenge for a smart city system design. Computer Networks (Amsterdam, Netherlands: 1999), 144(??): 89–110, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306534.

Zolotukhin:2018:ODL

[2619] Mikhail Zolotukhin, Alexander Sayenko, and Timo Hämäläinen. On optimal deployment of low power nodes for high frequency next generation wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 144(??): 111–119, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306376.

Sarang:2018:QMP

[2620] Sohail Sarang, Micheal Drieberg, Azlan Awang, and Rizwan Ahmad. A QoS MAC protocol for prioritized data in energy harvesting wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 144(??): 120–140, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305723.

Promsuk:2018:PFA

[2621] Natthanan Promsuk, Attaphongse Taparugssanagorn, and Johanna Vartiainen. Probability of false alarm based interference suppression methods in Internet of Things (IoT) systems. Computer Networks (Amsterdam, Netherlands: 1999), 144(??): 141–153, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307308.

Kumar:2018:ACO

[2622] Priyan Malarvizhi Kumar, Usha Devi G, Gunasekaran Manogaran, Revathi Sundarasekar, Naveen Chilamkurti, and Ramachandran Varatharajan. Ant colony optimization algorithm with Internet of Vehicles for intelligent traffic control system. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):163–200, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304845.

Enoch:2018:SEC

[2623] Simon Yusuf Enoch, Mengmeng Ge, Jin B. Hong, Hani Alzaid, and Dong Seong Kim. A systematic evaluation of cybersecurity metrics for dynamic networks. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):201–215, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306285.

Lo:2018:GTM

[2624] Chun-Chih Lo and Yau-Hwang Kuo. A geo-temporal mobility prediction method for cooperative time-validity-constrained content delivery in opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):216–229, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830700X.

Behera:2018:ETI

[2625] Sadananda Behera, Jithin George, and Goutam Das. Effect of transmission impairments in CO-OFDM based Elastic Optical Network design. Computer Networks (Amsterdam, Netherlands: 1999), 144(??): 230–241, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307291.

AL-Hazemi:2018:DAP

[2626] Fawaz AL-Hazemi, Yuyang Peng, Chan-Hyun Youn, Josip Lorincz, Chao Li, Guo Song, and Raouf Boutaba. Dynamic allocation of power delivery paths in consolidated data centers based on adaptive UPS switching. Computer Networks (Amsterdam, Netherlands: 1999), 144(??): 242–253, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307138.

Caicedo-Munoz:2018:QCV

[2627] Julian Andres Caicedo-Muñoz, Agapito Ledezma Espino, Juan Carlos Corrales, and Alvaro Rendón. QoS-Classifier for VPN and Non-VPN traffic based on timerelated features. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):254–270, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307321.

Ye:2018:PPS

[2628] Yuhang Ye, Brian Lee, Ronan Flynn, Niall Murray, Guiming Fang, Jianwen Cao, and Yuansong Qiao. PTP: Pathspecified transport protocol for concurrent multipath transmission in named data networks. Computer Networks (Amsterdam, Netherlands: 1999), 144 (??):271–279, October 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306686.

Colakovic:2018:ITI

[2629] Alem Colaković and Mesud Hadzialić. Internet of Things (IoT): a review of enabling technologies, challenges, and open research issues. Computer Networks (Amsterdam, Netherlands: 1999), 144(??):280– 296, October 24, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305243.

Anonymous:2018:EBo

[2630] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 145(??):ii, November 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618310016.

Hansen:2018:BIF

[2631] Jonas Hansen, Jeppe Krigslund, Daniel E. Lucani, Peyman Pahlevani, and Frank H. P. Fitzek. Bridging inter-flow and intra-flow network coding in wireless mesh networks: From theory to implementation. Computer Networks (Amsterdam, Netherlands: 1999), 145(??):1–12, November 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305188.

Angles-Tafalla:2018:SPP

[2632] Carles Anglès-Tafalla, Jordi Castellà-Roca, Macià Mut-Puigserver, M. Magdalena Payeras-Capellà, and Alexandre Viejo. Secure and privacypreserving lightweight access control system for low emission zones.

Computer Networks (Amsterdam, Netherlands: 1999), 145(??):13– 26, November 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307655.

Reali:2018:GSJ

[2633] G. Reali, M. Femminella, E. Nunzi, and D. Valocchi. Genomics as a service: a joint computing and networking perspective. Computer Networks (Amsterdam, Netherlands: 1999), 145 (??):27–51, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307205.

Abusitta:2018:TFC

[2634] Adel Abusitta, Martine Bellaiche, and Michel Dagenais. On trustworthy federated clouds: a coalitional game approach. Computer Networks (Amsterdam, Netherlands: 1999), 145 (??):52–63, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307606.

Chao:2018:ABW

[2635] I-Fen Chao and Chain-Hung Lee.

AWG-based WDM ring networks: High-performance and low-cost system designs. Computer Networks (Amsterdam, Netherlands: 1999), 145 (??):64–75, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308168.

Kim:2018:UIM

[2636] Won-Suk Kim and Sang-Hwa Chung. User incentive model and its optimization scheme in user-participatory fog computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 145(??):76–88, November 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618307990.

Saad:2018:NIR

[2637] Mohamed Saad. Non-isotonic routing metrics solvable to optimality via shortest path. Computer Networks (Amsterdam, Netherlands: 1999), 145 (??):89–95, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308338.

Awad:2018:ORN

[2638] Fahed H. Awad. Optimization of relay node deployment for multisource multipath routing in Wireless Multimedia Sensor Networks using Gaussian distribution. Computer Networks (Amsterdam, Netherlands: 1999), 145 (??):96–106, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308429.

Wu:2018:NOB

[2639] Tao Wu, Panlong Yang, Haipeng Dai, Ping Li, and Xunpeng Rao. Near optimal bounded route association for drone-enabled rechargeable WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 145(??):

107–117, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304912.

Inoue:2018:NIV

[2640] Koki Inoue, Shin’ichi Arakawa, Satoshi Imai, Toru Katagiri, and Masayuki Murata. Noise-induced VNE method for software-defined infrastructure with uncertain delay behaviors. Computer Networks (Amsterdam, Netherlands: 1999), 145(??):118–127, November 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618308600.

Ahmad:2018:VLB

[2641] Iftikhar Ahmad, Rafidah Md Noor, Ismail Ahmedy, Syed Adeel Ali Shah, Ibrar Yaqoob, Ejaz Ahmed, and Muhammad Imran. VANET-LTE based heterogeneous vehicular clustering for driving assistance and route planning applications. Computer Networks (Amsterdam, Netherlands: 1999), 145(??): 128–140, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308417.

Bousbaa:2018:DTL

[2642] Fatima Zohra Bousbaa, Nasreddine Lagraa, Chaker Abdelaziz Kerrache, Fen Zhou, Mohamed Bachir Yagoubi, and Rasheed Hussain. A distributed time-limited multicast algorithm for VANETs using incremental power strategy. Computer Networks (Amsterdam, Netherlands: 1999), 145(??): 141–155, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303724.

Zhu:2018:HAL

[2643] Chuan Zhu, Kangning Quan, Guangjie Han, and Joel J. P. C. Rodrigues. A high-available and location predictive data gathering scheme with mobile sinks for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 145(??): 156–164, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308442.

Yan:2018:IFD

[2644] Xiaoyun Yan, Ping Dong, Xiaojiang Du, Tao Zheng, Jianan Sun, and Mohsen Guizani. Improving flow delivery with link available time prediction in software-defined high-speed vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 145 (??):165–174, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308454.

Wang:2018:BSB

[2645] Chun-Yu Wang, Chi-Lung Ou, Yu-En Zhang, Feng-Min Cho, Pin-Hao Chen, Jyh-Biau Chang, and Ce-Kuen Shieh. BotCluster: a session-based P2P botnet clustering system on Net-Flow. Computer Networks (Amsterdam, Netherlands: 1999), 145(??):

175–189, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308351.

Khowaja:2018:CAB

[2646] Sunder Ali Khowaja, Aria Ghora Prabono, Feri Setiawan, Bernardo Nugroho Yahya, and Seok-Lyong Lee. Contextual activity based Healthcare Internet of Things, Services, and People (HIoTSP): an architectural framework for healthcare monitoring using wearable sensors. Computer Networks (Amsterdam, Netherlands: 1999), 145 (??):190–206, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308594.

AlRidhawi:2018:CDV

[2647] Ismaeel Al Ridhawi, Moayad Aloqaily, Burak Kantarci, Yaser Jararweh, and Hussein T. Mouftah. A continuous diversified vehicular cloud service availability framework for smart cities. Computer Networks (Amsterdam, Netherlands: 1999), 145(??): 207–218, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308430.

Singh:2018:BBB

[2648] Madhusudan Singh and Shiho Kim.

Branch based blockchain technology in intelligent vehicle. Computer Networks (Amsterdam, Netherlands: 1999), 145 (??):219–231, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308399.

Chen:2018:INF

[2649] Xin Chen, Guoqiang Zhang, Qian Gao, and Huajun Cui. Improving NDN forwarding engine performance by rendezvous-based caching and forwarding. Computer Networks (Amsterdam, Netherlands: 1999), 145(??): 232–242, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308302.

Samba:2018:PFD

[2650] Alassane Samba, Yann Busnel, Alberto Blanc, Philippe Dooze, and Gwendal Simon. Predicting file downloading time in cellular network: Largescale analysis of machine learning approaches. Computer Networks (Amsterdam, Netherlands: 1999), 145(??): 243–254, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308545.

Amodu:2018:MMC

[2651] Oluwatosin Ahmed Amodu and Mohamed Othman. Machine-to-machine communication: an overview of opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 145(??): 255–276, November 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830851X.

Anonymous:2018:EBp

[2652] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 146(??):ii, December 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618311101.

Vasconcelos:2018:DSA

[2653] Israel L. C. Vasconcelos, Ivan C.

Martins, Carlos M. S. Figueiredo, and Andre L. L. Aquino. A data sample algorithm applied to wireless sensor network with disruptive connections. Computer Networks (Amsterdam, Netherlands: 1999), 146 (??):1–11, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303621.

Abrahao:2018:SAL

[2654] Diego Cruz Abrahão and Flávio Henrique Teles Vieira. Scheduling algorithm for the LTE downlink with F-OFDM using adaptive multifractal envelope process and minimum service curve. Computer Networks (Amsterdam, Netherlands: 1999), 146 (??):12–32, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308806.

Tashtarian:2018:VSB

[2655] Farzad Tashtarian, Alireza Erfanian, and Amir Varasteh. S2VC: an SDNbased framework for maximizing QoE in SVC-based HTTP adaptive streaming. Computer Networks (Amsterdam, Netherlands: 1999), 146(??): 33–46, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830879X.

Ghahfarokhi:2018:ESE

[2656] Behrouz Shahgholi Ghahfarokhi, Maryam Azadmanesh, and Samaneh Khadem Khorasani. Energy and spectrum efficient mobility-aware resource management for D2D multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 146 (??):47–64, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309198.

Condoluci:2018:SVM

[2657] Massimo Condoluci and Toktam Mahmoodi. Softwarization and virtualization in 5G mobile networks: Benefits, trends and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 146 (??):65–84, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302500.

Liu:2018:JDR

[2658] Jain-Shing Liu. Joint downlink resource allocation in LTE-Advanced heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 146(??):85–103, December 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618309101.

Luo:2018:NDA

[2659] Chuanwen Luo, Jiguo Yu, Deying Li, Honglong Chen, Yi Hong, and Lina Ni. A novel distributed algorithm for constructing virtual backbones in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 146 (??):104–114, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309265.

Yamada:2018:TTS

[2660] Yoshinobu Yamada, Ryoichi Shinkuma, Takanori Iwai, Takeo Onishi, Takahiro Nobukiyo, and Kozo Satoda. Temporal traffic smoothing for IoT traffic in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 146 (??):115–124, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308466.

Kirton:2018:TOS

[2661] Jack Kirton, Matthew Bradbury, and Arshad Jhumka. Towards optimal source location privacy-aware TDMA schedules in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 146(??): 125–137, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308958.

Tang:2018:QAR

[2662] Lun Tang, Guofan Zhao, Chenmeng Wang, Peipei Zhao, and Qianbin Chen. Queue-aware reliable embedding algorithm for 5G network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 146(??): 138–150, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309204.

Gaber:2018:TBS

[2663] Tarek Gaber, Sarah Abdelwahab, Mohamed Elhoseny, and Aboul Ella Hassanien. Trust-based secure clustering in WSN-based intelligent transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 146 (??):151–158, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309253.

Yang:2018:ECE

[2664] Hao Yang and Xiwei Wang. ECO CS: Energy consumption optimized compressive sensing in group sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 146(??): 159–166, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304869.

Lechowicz:2018:GRA

[2665] Piotr Lechowicz, Krzysztof Walkowiak, and Miroslaw Klinkowski. Greedy randomized adaptive search procedure for joint optimization of unicast and anycast traffic in spectrallyspatially flexible optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 146(??):167– 182, December 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308995.

Xu:2018:ERR

[2666] Dan Xu, Wenli Jiao, Zhuang Yin, Bin Wu, Yao Peng, Xiaojiang Chen, Feng Chen, and Dingyi Fang. Enabling robust and reliable transmission in Internet of Things with multiple gateways. Computer Networks (Amsterdam, Netherlands: 1999), 146 (??):183–199, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309708.

Bu:2018:FOR

[2667] Gewu Bu and Maria Potop-Butucaru.

FIFO order reliable convergecast in WBAN. Computer Networks (Amsterdam, Netherlands: 1999), 146(??): 200–216, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309630.

Sun:2018:PMA

[2668] Peng Sun and Azzedine Boukerche.

Performance modeling and analysis of a UAV path planning and target detection in a UAV-based wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 146(??):

217–231, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309782.

Almiani:2018:ERU

[2669] Khaled Almi’ani, Young Choon Lee, and Bernard Mans. On efficient resource use for scientific workflows in clouds. Computer Networks (Amsterdam, Netherlands: 1999), 146(??): 232–242, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303384.

Anonymous:2018:EBq

[2670] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 147(??):ii, December 24, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618312362.

Wang:2018:WDM

[2671] Lei Wang, Qing Li, Yong Jiang, Xuya Jia, and Jianping Wu. Woodpecker: Detecting and mitigating link-flooding attacks via SDN. Computer Networks (Amsterdam, Netherlands: 1999), 147 (??):1–13, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309721.

Ra:2018:PFB

[2672] Yongwook Ra, June-Koo Kevin Rhee, Junseong Bang, and Yunjoo Kim. Photonic-frame-based TCP proxy architecture in optically interconnected data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 14–26, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302408.

Thakur:2018:AHT

[2673] Prabhat Thakur, Alok Kumar, S. Pandit, G. Singh, and S. N. Satashia. Analysis of high-traffic cognitive radio network with imperfect spectrum monitoring technique. Computer Networks (Amsterdam, Netherlands: 1999), 147 (??):27–37, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310065.

Huang:2018:ASG

[2674] Jianhua Huang, Danwei Ruan, and Weiqiang Meng. An annulus sector grid aided energy-efficient multihop routing protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 147 (??):38–48, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309861.

Zhao:2018:FRR

[2675] Yu Zhao, Yunhuai Liu, Tingting Yu, Tian He, and Chen Qian. FREDI: Robust RSS-based ranging with multipath effect and radio interference. Computer Networks (Amsterdam, Netherlands: 1999), 147(??):

49–63, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618310223.

Wang:2018:PSP

[2676] Xiaodong Wang, Ye Tian, Min Zhao, Mingzheng Li, Lei Mei, and Xinming Zhang. PNPL: Simplifying programming for protocol-oblivious SDN networks. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 64–80, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830968X.

Delaet:2018:MBB

[2677] Sylvie Delaët, Shlomi Dolev, Daniel Khankin, and Shimrit Tzur-David. Make&activate-before-break for seamless SDN route updates. Computer Networks (Amsterdam, Netherlands: 1999), 147(??):81–97, December 24, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618310326.

Bu:2018:FIM

[2678] Kai Bu, Kaiwen Zhu, Yi Zheng, Yuanyuan Yang, Yutian Yang, and Linfeng Cheng. Fastlane-ing more flows with less bandwidth for softwaredefined networking. Computer Networks (Amsterdam, Netherlands: 1999), 147(??):98–111, December 24, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618306182.

Dorsch:2018:EHS

[2679] Nils Dorsch, Fabian Kurtz, and Christian Wietfeld. Enabling hard service guarantees in Software-Defined Smart Grid infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 147(??):112–131, December 24, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618310417.

Atzori:2018:TIS

[2680] Luigi Atzori, Alessandro Floris, Roberto Girau, Michele Nitti, and Giovanni Pau. Towards the implementation of the Social Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 132–145, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310041.

Hanczewski:2018:QMM

[2681] Slawomir Hanczewski, Maciej Stasiak, and Joanna Weissenberg. Queueing model of a multi-service system with elastic and adaptive traffic. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 146–161, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309836.

Jarray:2018:SFT

[2682] Chedia Jarray and Anastasios Giovanidis. Successful file transmission in mobile D2D networks with caches. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 162–179, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618301993.

Hajisami:2018:CBJ

[2683] Abolfazl Hajisami and Dario Pompili. Cloud-BSS: Joint intra- and inter-cluster interference cancellation in uplink 5G cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 180–190, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310399.

Grassi:2018:MMI

[2684] A. Grassi, G. Piro, G. Boggia, M. Kurras, W. Zirwas, R. SivaSiva Ganesan, K. Pedersen, and L. Thiele. Massive MIMO interference coordination for 5G broadband access: Integration and system level study. Computer Networks (Amsterdam, Netherlands: 1999), 147 (??):191–203, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310818.

Cohen:2018:QMB

[2685] Itamar Cohen and Gabriel Scalosub. Queueing in the mist: Buffering and scheduling with limited knowledge. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 204–220, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311010.

Attia:2018:QAS

[2686] Maroua Ben Attia, Kim-Khoa Nguyen, and Mohamed Cheriet. QoS-aware software-defined routing in smart community network. Computer Networks (Amsterdam, Netherlands: 1999), 147 (??):221–235, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311046.

Li:2018:BEN

[2687] Xiaoqian Li and Kwan L. Yeung.

Bandwidth-efficient network monitoring algorithms based on segment routing. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 236–245, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310752.

Khokhar:2018:ISF

[2688] Muhammad Jawad Khokhar, Nawfal Abbassi Saber, Thierry Spetebroot, and Chadi Barakat. An intelligent sampling framework for controlled experimentation and QoE modeling. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 246–261, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303700.

Akyildiz:2018:AFN

[2689] I. F. Akyildiz, A. Kak, E. Khorov, A. Krasilov, and A. Kureev. AR-BAT: a flexible network architecture for QoE-aware communications in 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 147(??): 262–279, December 24, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618311228.

Anonymous:2019:EBa

[2690] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 148(??):ii, January 15, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128618314233.

Akyildiz:2019:CNC

[2691] Ian F. Akyildiz, Harry Rudin, and Burkhard Stiller. Computer Networks (COMNET) editorial for 2018. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):iii–iv, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618314270.

Huang:2019:MRW

[2692] Hailong Huang, Andrey V. Savkin, Ming Ding, and Chao Huang. Mobile robots in wireless sensor networks: a survey on tasks. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):1–19, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830255X.

Park:2019:MPP

[2693] Jaehyun Park, Jaehee Ha, Brent Byunghoon Kang, and Myungchul Kim. Mo-HoP: a protocol providing for both mobility management and host privacy. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 20–33, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311022.

Yuan:2019:APL

[2694] Yazhou Yuan, Xiaoqin Sun, Zhixin Liu, Yuefeng Li, and Xinping Guan. Approach of personnel location in roadway environment based on multisensor fusion and activity classification. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):34–45, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303542.

Shao:2019:DRP

[2695] Yanling Shao, Chunlin Li, and Hengliang Tang. A data replica placement strategy for IoT workflows in collaborative edge and cloud environments. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):46–59, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311460.

Krentz:2019:DSD

[2696] Konrad-Felix Krentz and Christoph Meinel. Denial-of-sleep defenses for IEEE 802.15.4 coordinated sampled listening (CSL). Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):60–71, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305954.

Lin:2019:DWT

[2697] Chi Lin, Yu Sun, Kai Wang, Zhunyue Chen, Bo Xu, and Guowei Wu. Double warning thresholds for preemptive charging scheduling in Wireless Rechargeable Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):72–87, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618311708.

Lourenco:2019:DED

[2698] Rafael B. R. Lourenço, Gustavo B. Figueiredo, Massimo Tornatore, and Biswanath Mukherjee. Data evacuation from data centers in disaster-affected regions through software-defined satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):88–100, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618311472.

Ning:2019:NOC

[2699] Huansheng Ning, Feifei Shi, Tao Zhu, Qingjuan Li, and Liming Chen.

A novel ontology consistent with acknowledged standards in smart homes. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 101–107, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311964.

Leo:2019:NAT

[2700] Yannick Léo, Christophe Crespelle, and Eric Fleury. Non-altering time scales for aggregation of dynamic networks into series of graphs. Computer Networks (Amsterdam, Netherlands: 1999), 148(??):108– 119, January 15, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861831199X.

Chai:2019:RCA

[2701] Yuan Chai and Xiao-Jun Zeng. Regional condition-aware hybrid routing protocol for hybrid wireless mesh network. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 120–128, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312076.

Nguyen:2019:ORU

[2702] Thi-Minh Nguyen, Michel Minoux, and Serge Fdida. Optimizing resource utilization in NFV dynamic systems: New exact and heuristic approaches. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 129–141, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618302287.

Wang:2019:SLP

[2703] Hao Wang, Guangjie Han, Lina Zhou, James Adu Ansere, and Wenbo Zhang. A source location privacy protection scheme based on ring-loop routing for the IoT. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 142–150, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618311976.

Ghermezcheshmeh:2019:APE

[2704] Mojtaba Ghermezcheshmeh, Vahid Shah-Mansouri, and Mohammad Ghanbari. Analysis and performance evaluation of scalable video coding over heterogeneous cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 151–163, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618311654.

Salo:2019:DRI

[2705] Fadi Salo, Ali Bou Nassif, and Aleksander Essex. Dimensionality reduction with IG-PCA and ensemble classifier for network intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 164–175, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303037.

Marino:2019:ACN

[2706] Francesco Marino, Corrado Moiso, and Matteo Petracca. Automatic contract negotiation, service discovery and mutual authentication solutions: a survey on the enabling technologies of the forthcoming IoT ecosystems. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 176–195, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312167.

Kompara:2019:REM

[2707] Marko Kompara, SK Hafizul Islam, and Marko Hölbl. A robust and efficient mutual authentication and key agreement scheme with untraceability for WBANs. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):196–213, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303955.

Wahid:2019:HAC

[2708] Abdul Wahid, Humera Yasmeen, Munam Ali Shah, Masoom Alam, and Sayed Chhattan Shah. Holistic approach for coupling privacy with safety in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 214–230, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308405.

Ji:2019:FSS

[2709] Cun Ji, Chao Zhao, Shijun Liu, Chenglei Yang, Li Pan, Lei Wu, and Xiangxu Meng. A fast shapelet selection algorithm for time series classification. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 231–240, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312970.

Asghari:2019:ITA

[2710] Parvaneh Asghari, Amir Masoud Rahmani, and Hamid Haj Seyyed Javadi. Internet of Things applications: a systematic review. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):241–261, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305127.

Petreska:2019:BBP

[2711] Neda Petreska, Hussein Al-Zubaidy, Rudi Knorr, and James Gross. Boundbased power optimization for multihop heterogeneous wireless industrial networks under statistical delay constraints. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 262–279, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308831.

Schunter:2019:ESI

[2712] Matthias Schunter and Andreas Wespi. Editorial: Special issue on IoT security and privacy. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):280–282, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618313240.

Noor:2019:CRI

[2713] Mardiana binti Mohamad Noor and Wan Haslina Hassan. Current research on Internet of Things (IoT) security: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):283–294, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307035.

Hou:2019:SIT

[2714] Jianwei Hou, Leilei Qu, and Wenchang Shi. A survey on Internet of Things security from data perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 295–306, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306844.

Zhou:2019:PRI

[2715] Zheng Zhou, Weiming Zhang, Shangbin Li, and Nenghai Yu. Potential risk of IoT device supporting IR remote control. Computer Networks (Amsterdam, Netherlands: 1999), 148 (??):307–317, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306200.

Yang:2019:TAF

[2716] Kai Yang, Qiang Li, and Limin Sun. Towards automatic fingerprinting of IoT devices in the cyberspace. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 318–327, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306856.

Eldefrawy:2019:FSA

[2717] Mohamed Eldefrawy, Ismail Butun, Nuno Pereira, and Mikael Gidlund. Formal security analysis of Lo-RaWAN. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 328–339, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306145.

Liu:2019:PPR

[2718] Yi-Ning Liu, Yan-Ping Wang, Xiao-Fen Wang, Zhe Xia, and Jing-Fang Xu. Privacy-preserving raw data collection without a trusted authority for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 340–348, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306789.

Liu:2019:RFR

[2719] Yuxian Liu, Shaohua Tang, Hao-Tian Wu, and Xinglin Zhang. RTPT: a framework for real-time privacypreserving truth discovery on crowdsensed data streams. Computer Networks (Amsterdam, Netherlands: 1999), 148(??):349–360, January 15, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618306613.

Sarfraz:2019:PAI

[2720] Umair Sarfraz, Masoom Alam, Sherali Zeadally, and Abid Khan. Privacy aware IOTA ledger: Decentralized mixing and unlinkable IOTA transactions. Computer Networks (Amsterdam, Netherlands: 1999), 148(??): 361–372, January 15, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306972.

Anonymous:2019:EBb

[2721] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 149(??):ii, February 11, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619300568.

Li:2019:BFF

[2722] Qing Li, Yang Liu, Zhijie Zhu, Hengtong Li, and Yong Jiang. BOND: Flexible failure recovery in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):1–12, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312416.

Bernardini:2019:PPP

[2723] Cesar Bernardini, Samuel Marchal, Muhammad Rizwan Asghar, and Bruno Crispo. PrivICN: Privacy-preserving content retrieval in information-centric networking. Computer Networks (Amsterdam, Netherlands: 1999), 149(??):13– 28, February 11, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312179.

Gupta:2019:LAU

[2724] Ankur Gupta, Meenakshi Tripathi, Tabish Jamil Shaikh, and Aakar Sharma. A lightweight anonymous user authentication and key establishment scheme for wearable devices. Computer Networks (Amsterdam, Netherlands: 1999), 149(??): 29–42, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304389.

Steiner:2019:TMP

[2725] Rodrigo Vieira Steiner and Emil Lupu. Towards more practical software-based attestation. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):43–55, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307631.

Liu:2019:TTP

[2726] Tong Liu, Yanmin Zhu, and Liqun Huang. TGBA: a two-phase group buying based auction mechanism for recruiting workers in mobile crowd sensing. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):56–75, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830238X.

Zalghout:2019:CAP

[2727] Mohamad Zalghout, Ayman Khalil, Matthieu Crussiére, Samih Abdul-Nabi, and Jean-Francois Hélard. Context-aware and priority-based user association and resource allocation in heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 149(??): 76–92, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311812.

Sou:2019:FDP

[2728] Sok-Ian Sou, Ming-Ru Li, Shan-Huei Wang, and Meng-Hsun Tsai. File distribution via proximity group communications in LTE-advanced public safety networks. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):93–101, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312568.

Killi:2019:TIR

[2729] Bala Prakasa Rao Killi and Seela Veerabhadreswara Rao. Towards improving resilience of controller placement with minimum backup capacity in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):102–114, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309575.

Trevisan:2019:PPW

[2730] Martino Trevisan, Idilio Drago, and Marco Mellia. PAIN: a Passive Web performance indicator for ISPs. Computer Networks (Amsterdam, Netherlands: 1999), 149(??): 115–126, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830358X.

Xie:2019:CEL

[2731] Rongjun Xie, Ibrahim Khalil, Shahriar Badsha, and Mohammed Atiquzzaman. Collaborative extreme learning machine with a confidence interval for P2P learning in healthcare. Computer Networks (Amsterdam, Netherlands: 1999), 149(??): 127–143, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307278.

Qin:2019:EPR

[2732] Xueyang Qin, Xiaoming Wang, Liang Wang, Yaguang Lin, and Xinyan Wang. An efficient probabilistic routing scheme based on game theory in opportunistic networks. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):144–153, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312556.

Arfeen:2019:GLS

[2733] Asad Arfeen, Krzysztof Pawlikowski, Don McNickle, and Andreas Willig. Global and local scaling analysis of link streams in access and backbone core networks. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):154–172, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861831288X.

Wang:2019:OPB

[2734] Tong Wang, MengBo Tang, Houbing Song, Y. Cao, and Zaheer Khan. Opportunistic protocol based on social probability and resources efficiency for the intelligent and connected transportation system. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):173–186, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312994.

Linguaglossa:2019:HSD

[2735] Leonardo Linguaglossa, Dario Rossi, Salvatore Pontarelli, Dave Barach, Damjan Marjon, and Pierre Pfister. High-speed data plane and network functions virtualization by vectorizing packet processing. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):187–199, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312957.

Atzori:2019:SCI

[2736] L. Atzori, J. L. Bellido, R. Bolla, G. Genovese, A. Iera, A. Jara, C. Lombardo, and G. Morabito. SDN&NFV contribution to IoT objects virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 149(??): 200–212, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312933.

Qian:2019:FAM

[2737] Shiyou Qian, Jian Cao, Weichao Mao, Yanmin Zhu, Jiadi Yu, Minglu Li, and Jie Wang. A fast and antimatchability matching algorithm for content-based publish/subscribe systems. Computer Networks (Amsterdam, Netherlands: 1999), 149(??): 213–225, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313197.

Ray:2019:ITB

[2738] Partha Pratim Ray, Dinesh Dash, and Debashis De. Internet of Things-based real-time model study on e-healthcare: Device, message service and dew computing. Computer Networks (Amsterdam, Netherlands: 1999), 149(??): 226–239, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304973.

Gao:2019:EER

[2739] Xiaozheng Gao, Kai Yang, Nan Yang, and Jinsong Wu. Energy-efficient resource block assignment and power control for underlay device-to-device communications in multi-cell networks. Computer Networks (Amsterdam, Netherlands: 1999), 149(??): 240–251, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305231.

Leyva-Mayorga:2019:AAC

[2740] Israel Leyva-Mayorga, Miguel A.

Rodriguez-Hernandez, Vicent Pla, Jorge Martinez-Bauset, and Luis Tello-Oquendo. Adaptive access class barring for efficient mMTC. Computer Networks (Amsterdam, Netherlands: 1999), 149(??):252– 264, February 11, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308211.

Ramakrishnan:2019:AEE

[2741] S. Ramakrishnan, Subrat Kar, and Dharmaraja Selvamuthu. Analysis of energy efficiency in cloud based heterogeneous RAN with large-scale antenna systems. Computer Networks (Amsterdam, Netherlands: 1999), 149 (??):265–276, February 11, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618313264.

Anonymous:2019:EBc

[2742] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 150(??):ii, February 26, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619301215.

Cheng:2019:MIH

[2743] Siyao Cheng, Yingshu Li, Zhi Tian, Wei Cheng, and Xiuzhen Cheng. A model for integrating heterogeneous sensory data in IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 150 (??):1–14, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312982.

Wang:2019:EMI

[2744] You-Chiun Wang and Chien-Chun Huang. Efficient management of interference and power by jointly configuring ABS and DRX in LTE-A HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 150 (??):15–27, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307515.

Kumar:2019:IAQ

[2745] Puneet Kumar and Behnam Dezfouli. Implementation and analysis of QUIC for MQTT. Computer Networks (Amsterdam, Netherlands: 1999), 150 (??):28–45, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310776.

Hong:2019:SIT

[2746] Jin B. Hong, Armstrong Nhlabatsi, Dong Seong Kim, Alaa Hussein, Noora Fetais, and Khaled M. Khan. Systematic identification of threats in the cloud: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 150 (??):46–69, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308259.

Batista:2019:TIM

[2747] Pedro Batista, Ivanes Araújo, Neiva Linder, Kim Laraqui, and Aldebaro Klautau. Testbed for ICN media distribution over LTE radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 70–80, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861831380X.

Din:2019:EHA

[2748] Sadia Din, Anand Paul, and Abdul Rehman. 5G-enabled Hierarchical architecture for softwaredefined intelligent transportation system. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 81–89, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312969.

Li:2019:ISI

[2749] Kaiyang Li, Ling Tian, Wei Li, Guangchun Luo, and Zhipeng Cai. Incorporating social interaction into three-party game towards privacy protection in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 150 (??):90–101, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312945.

Nehra:2019:SSL

[2750] Ajay Nehra, Meenakshi Tripathi, Manoj Singh Gaur, Ramesh Babu Battula, and Chhagan Lal. SLDP: a secure and lightweight link discovery protocol for software defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 102–116, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307916.

Mharsi:2019:MPA

[2751] Niezi Mharsi and Makhlouf Hadji. A mathematical programming approach for full coverage hole optimization in Cloud Radio Access Networks. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 117–126, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307928.

Viard:2019:ACB

[2752] Tiphaine Viard and Raphaël Fournier-S’niehotta. Augmenting content-based rating prediction with link stream features. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 127–133, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618313215.

Akyildiz:2019:IST

[2753] Ian F. Akyildiz and Ahan Kak.

The Internet of Space Things/ CubeSats: a ubiquitous cyberphysical system for the connected world. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 134–149, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314191.

Hong:2019:ABB

[2754] Gongbing Hong, James Martin, and James Westall. Adaptive bandwidth binning for bandwidth management. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 150–169, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314452.

Hatanaka:2019:MLE

[2755] Takayuki Hatanaka and Takuji Tachibana.

LE-MRC: Low-energy based multiple routing configurations for fast failure recovery. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 170–178, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618302263.

Ayadi:2019:KTO

[2756] Aya Ayadi, Oussama Ghorbel, M. S. BenSalah, and Mohamed Abid. Kernelized technique for outliers detection to monitoring water pipeline based on WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 179–189, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307096.

Bai:2019:DWA

[2757] Sen Bai, Yunhao Liu, Zhenhua Li, and Xin Bai. Detecting wormhole attacks in 3D wireless ad hoc networks via 3D forbidden substructures. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 190–200, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618302317.

Liu:2019:NTS

[2758] Chuan-Gang Liu, I-Hsien Liu, Chang-De Lin, and Jung-Shian Li. A novel tag searching protocol with time efficiency and searching accuracy in RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 201–216, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619300325.

Abdel-Halim:2019:MPB

[2759] Islam Tharwat Abdel-Halim, Hossam Mahmoud Ahmed Fahmy, and Ayman M. Bahaa-El Din. Mobility prediction-based efficient clustering scheme for connected and automated vehicles in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 150(??):217– 233, February 26, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307175.

Cambiaso:2019:ISA

[2760] Enrico Cambiaso, Giovanni Chiola, and Maurizio Aiello. Introducing the SlowDrop Attack. Computer Networks (Amsterdam, Netherlands: 1999), 150 (??):234–249, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300210.

Li:2019:TTE

[2761] Binbin Li, Yuan He, Wenyuan Liu, and Lin Wang. Towards time-efficient localized polling for large-scale RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 150(??): 250–262, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303426.

Latapy:2019:LSM

[2762] Matthieu Latapy, Marco Fiore, and Artur Ziviani. Link streams: Methods and applications. Computer Networks (Amsterdam, Netherlands: 1999), 150 (??):263–265, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313598.

Li:2019:EGC

[2763] Feng Li, Shibo He, Jun Luo, Mohan Gurusamy, and Junshan Zhang. Editorial: Green computing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 150 (??):266–268, February 26, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312052.

Anonymous:2019:EBd

[2764] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 151(??):ii, March 14, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619302026.

eSilva:2019:TPC

[2765] Renan Fischer e Silva and Paul M. Carpenter. TCP proactive congestion control for east–west traffic: the marking threshold. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):1–11, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619300088.

Souissi:2019:MLS

[2766] Ilhem Souissi, Nadia Ben Azzouna, and Lamjed Ben Said. A multi-level study of information trust models in WSN-assisted IoT. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):12–30, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300258.

Salam:2019:SSR

[2767] Abdul Salam, Mehmet C. Vuran, and Suat Irmak. Di-Sense: In situ realtime permittivity estimation and soil moisture sensing using wireless underground communications. Computer Networks (Amsterdam, Netherlands: 1999), 151(??):31–41, March 14, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303141.

He:2019:OTA

[2768] Yu He, Lin Ma, Ruiting Zhou, Chuanhe Huang, and Zongpeng Li. Online task allocation in mobile cloud computing with budget constraints. Computer Networks (Amsterdam, Netherlands: 1999), 151(??):42–51, March 14, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304018.

Alnasser:2019:CSC

[2769] Aljawharah Alnasser, Hongjian Sun, and Jing Jiang. Cyber security challenges and solutions for V2X communications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):52–67, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306157.

Asvija:2019:SHA

[2770] B. Asvija, R. Eswari, and M. B. Bijoy. Security in hardware assisted virtualization for cloud computing — state of the art issues and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 151(??):68–92, March 14, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618302998.

eSilva:2019:SPS

[2771] Edmundo de Souza e Silva, Rosa M. M. Leão, Daniel S. Menasché, and Don Towsley. On the scalability of P2P swarming systems. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):93–113, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307126.

Li:2019:AMA

[2772] Xiaocui Li, Zhangbing Zhou, Junqi Guo, Shangguang Wang, and Junsheng Zhang. Aggregated multiattribute query processing in edge computing for industrial IoT applications. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):114–123, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619301008.

Kavitha:2019:QSL

[2773] V. R. Kavitha and M. Moorthi. A quality of service load balanced connected dominating set-stochastic diffusion search (CDS-SDS) network backbone for MANET. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):124–131, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307217.

Miri:2019:ERA

[2774] Mohammadhasan Miri, Kamal Mohamedpour, Yousef Darmani, Mahasweta Sarkar, and R. Lal Tummala. An efficient resource allocation algorithm based on vertex coloring to mitigate interference among coexisting WBANs. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):132–146, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300908.

daCosta:2019:ITS

[2775] Kelton A. P. da Costa, João P. Papa, Celso O. Lisboa, Roberto Munoz, and Victor Hugo C. de Albuquerque. Internet of Things: a survey on machine learning-based intrusion detection approaches. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):147–157, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308739.

Kist:2019:FFG

[2776] Maicon Kist, Juliano Araújo Wickboldt, Lisandro Zambenedetti Granville, Juergen Rochol, Luiz A. DaSilva, and Cristiano Bonato Both. Flexible finegrained baseband processing with network functions virtualization: Benefits and impacts. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):158–165, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619301033.

Guo:2019:XBP

[2777] Junqi Guo, Lan Yang, Rongfang Bie, Jiguo Yu, Yuan Gao, Yan Shen, and Anton Kos. An XGBoost-based physical fitness evaluation model using advanced feature selection and Bayesian hyper-parameter optimization for wearable running monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 151(??): 166–180, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300994.

Prabhukavin:2019:SSP

[2778] Balasubramanian Prabhu kavin and Sannasi Ganapathy. A secured storage and privacy-preserving model using CRT for providing security on cloud and IoT-based applications. Computer Networks (Amsterdam, Netherlands: 1999), 151(??): 181–190, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308880.

R:2019:RTV

[2779] Dinesh Jackson Samuel R, E. Fenil, Gunasekaran Manogaran, G. N. Vivekananda, M. Thanjaivadivel, S. Jeeva, and A. Ahilan. Real time violence detection framework for football stadium comprising of big data analysis and deep learning through bidirectional LSTM. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):191–200, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308521.

Krishnan:2019:EEN

[2780] Sivakumar Krishnan, S. Lokesh, and M. Ramya Devi. An efficient Elman neural network classifier with cloud supported Internet of Things structure for health monitoring system. Computer Networks (Amsterdam, Netherlands: 1999), 151(??): 201–210, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308752.

Thangaramya:2019:EAC

[2781] K. Thangaramya, K. Kulothungan, R. Logambigai, M. Selvi, Sannasi Ganapathy, and A. Kannan. Energy aware cluster and neuro-fuzzy based routing algorithm for wireless sensor networks in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):211–223, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307771.

AlFarraj:2019:NPA

[2782] Osama AlFarraj, Amr Tolba, Salem Alkhalaf, and Ahmad AlZubi. Neighbor predictive adaptive handoff algorithm for improving mobility management in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 151 (??):224–231, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618307242.

Yu:2019:LDL

[2783] Kan Yu, Yinglong Wang, Jiguo Yu, Dongxiao Yu, Xiuzhen Cheng, and Zhiguang Shan. Localized and distributed link scheduling algorithms in IoT under Rayleigh fading. Computer Networks (Amsterdam, Netherlands: 1999), 151(??): 232–244, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619300659.

Mosbah:2019:EDD

[2784] Aziza Ben Mosbah, Seid Eddine Hammami, Hassine Moungla, Hossam Afifi, and Ahmed E. Kamal. Enhancing device-to-device direct discovery based on predicted user density patterns. Computer Networks (Amsterdam, Netherlands: 1999), 151(??): 245–259, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300933.

Anonymous:2019:EBe

[2785] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 151(??):ii, March 14, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619302026.

Anonymous:2019:EBf

[2786] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):ii, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302750.

Younes:2019:SEC

[2787] Maram Bani Younes and Azzedine Boukerche. Safety and efficiency control protocol for highways using intelligent vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):1–11, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305267.

Li:2019:VNF

[2788] Defang Li, Peilin Hong, Kaiping Xue, and Jianing Pei. Virtual network function placement and resource optimization in NFV and edge computing enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):12–24, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305000.

Dourado:2019:DLI

[2789] Carlos M. J. M. Dourado, Jr., Suane Pires P. da Silva, Raul Victor M. da Nóbrega, Antonio Carlos da S. Barros, Pedo P. Rebouças, and Victor Hugo C. de Albuquerque. Deep learning IoT system for online stroke detection in skull computed tomography images. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):25–39, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830762X.

Barrachina-Munoz:2019:OOE

[2790] Sergio Barrachina-Muñoz, Francesc Wilhelmi, and Boris Bellalta. To overlap or not to overlap: Enabling channel bonding in high-density WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):40–53, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309745.

Lima:2019:WPI

[2791] Francisco H. M. B. Lima, Luiz F. M. Vieira, Marcos A. M. Vieira, Alex B. Vieira, and José Augusto M. Nacif. Water ping: ICMP for the Internet of Underwater Things. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):54–63, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300246.

Wu:2019:SAD

[2792] Zhijun Wu, Qingbo Pan, Meng Yue, and Liang Liu. Sequence alignment detection of TCP-targeted synchronous low-rate DoS attacks. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):64–77, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304043.

Tolba:2019:CAP

[2793] Amr Tolba. Content accessibility preference approach for improving service optimality in Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):78–86, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309216.

Madhja:2019:AWP

[2794] Adelina Madhja, Sotiris Nikoletseas, and Alexandros A. Voudouris. Adaptive wireless power transfer in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):87–97, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307370.

Al-Maitah:2019:OSU

[2795] Mohammed Al-Maitah, Ahmad Ali AlZubi, and Abdulaziz Alarifi. An optimal storage utilization technique for IoT devices using sequential machine learning. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):98–105, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308065.

AlZubi:2019:BFR

[2796] Ahmad Ali AlZubi, Mohammed Al-Maitah, and Abdulaziz Alarifi. A best-fit routing algorithm for nonredundant communication in largescale IoT based network. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):106–113, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307369.

Hu:2019:MRR

[2797] Yupeng Hu, Pengjie Ren, Wei Luo, Peng Zhan, and Xueqing Li. Multiresolution representation with recurrent neural networks application for streaming time series in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):114–132, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300660.

Leira:2019:PAG

[2798] Rafael Leira, Guillermo Julián-Moreno, Iván González, Francisco J. Gómez-Arribas, and Jorge E. López de Vergara. Performance assessment of 40 Gbit/s off-the-shelf network cards for virtual network probes in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):133–143, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930101X.

Baskar:2019:EPR

[2799] S. Baskar, S. Periyanayagi, P. Mohamed Shakeel, and V. R. Sarma Dhulipala. An energy persistent Range-dependent Regulated Transmission Communication model for vehicular network applications. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):144–153, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306777.

Dong:2019:UCP

[2800] Jiaqing Dong, Hao Yin, Chen Tian, Ahmed M. Abdelmoniem, Huapoing Zhou, Bo Bai, and Gong Zhang. Uranus: Congestion-proportionality among slices based on Weighted Virtual Congestion Control. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):154–166, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302585.

Liyanage:2019:CAT

[2801] Kushan Sudheera Kalupahana Liyanage, Maode Ma, and Peter Han Joo Chong. Connectivity aware tribrid routing framework for a generalized software defined vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):167–177, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830608X.

Shan:2019:SSC

[2802] Chun Shan, Shanshan Mei, Changzhen Hu, Liyuan Liu, and Limin Mao. Software structure characteristic measurement method based on weighted network. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):178–185, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619301124.

Rodriguez:2019:SHS

[2803] Pedro M Rodriguez, Aitor Lizeaga, Mikel Mendicute, and Iñaki Val. Spectrum handoff strategy for cognitive radio-based MAC for real-time industrial wireless sensor and actuator networks. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):186–198, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619301483.

Zhai:2019:HEC

[2804] Chao Zhai, Zhiyuan Yu, and Xinhua Wang. High-efficient cooperative relaying with wireless powered source and relay. Computer Networks (Amsterdam, Netherlands: 1999), 152 (??):199–209, April 7, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305747.

Xia:2019:ESL

[2805] Hui Xia, Chun qiang Hu, Fu Xiao, Xiang guo Cheng, and Zhen kuan Pan. An efficient social-like semantic-aware service discovery mechanism for largescale Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 152(??):210–220, April 7, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930177X.

Anonymous:2019:EBg

[2806] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 153(??):ii, April 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303172.

Hao:2019:FGD

[2807] Jialu Hao, Cheng Huang, Jianbing Ni, Hong Rong, Ming Xian, and Xuemin (Sherman) Shen. Finegrained data access control with attribute-hiding policy for cloud-based IoT. Computer Networks (Amsterdam, Netherlands: 1999), 153(??): 1–10, April 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619301793.

Oliveira:2019:ECE

[2808] Roberto M. Oliveira, Alex B. Vieira, and Moises V. Ribeiro. EPLC– CMAC: an enhanced cooperative MAC protocol for broadband PLC systems. Computer Networks (Amsterdam, Netherlands: 1999), 153 (??):11–22, April 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302099.

Kai:2019:ECA

[2809] Caihong Kai, Yuting Liang, Xinyue Hu, Zhengqiong Liu, and Lusheng Wang. An effective channel allocation algorithm to maximize system utility in heterogeneous DCB WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 153 (??):23–35, April 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300684.

Jindal:2019:SBB

[2810] Anish Jindal, Gagangeet Singh Aujla, and Neeraj Kumar. SURVIVOR: a blockchain based edge-as-a-service framework for secure energy trading in SDN-enabled vehicle-to-grid environment. Computer Networks (Amsterdam, Netherlands: 1999), 153 (??):36–48, April 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861831106X.

Li:2019:MAD

[2811] Yahui Li, Zhiliang Wang, Jiangyuan Yao, Xia Yin, Xingang Shi, Jianping Wu, and Han Zhang. MSAID: Automated detection of interference in multiple SDN applications. Computer Networks (Amsterdam, Netherlands: 1999), 153(??):49–62, April 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830536X.

Liang:2019:MDH

[2812] Yu Liang, Jidong Ge, Sheng Zhang, and Bin Luo. Modeling and deploying hybrid tenant requests with shared networklets. Computer Networks (Amsterdam, Netherlands: 1999), 153 (??):63–72, April 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302361.

Tsilopoulos:2019:ASS

[2813] Christos Tsilopoulos and George Xylomenos. Adaptive semi-stateless forwarding for Content-Centric Networks. Computer Networks (Amsterdam, Netherlands: 1999), 153 (??):73–85, April 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302536.

Godoi:2019:REP

[2814] Fabrício N. Godoi, Gustavo W.

Denardin, and Carlos H. Barriquello. Reliability enhancement of packet delivery in multi-hop wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 153 (??):86–91, April 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303529.

Song:2019:EEC

[2815] L. Song, K. K. Chai, Y. Chen, J. Loo, S. Jimaa, and Y. Iraqi. Energy efficient cooperative coalition selection in cluster-based capillary networks for CMIMO IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 153(??):92–102, April 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313811.

Lin:2019:QAD

[2816] Chuan-Sheng Lin and Sok-Ian Sou. QoS-aware dynamic bandwidth reallocation with deadline assurance for multipath data offloading. Computer Networks (Amsterdam, Netherlands: 1999), 153(??):103–112, April 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618302573.

Dhanvijay:2019:ITS

[2817] Mrinai M. Dhanvijay and Shailaja C. Patil. Internet of Things: a survey of enabling technologies in healthcare and its applications. Computer Networks (Amsterdam, Netherlands: 1999), 153 (??):113–131, April 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619302695.

Ghose:2019:EES

[2818] Debasish Ghose, Anders Frøytlog, and Frank Y. Li. Enabling early sleeping and early data transmission in wake-up radio-enabled IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 153(??):132–144, April 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310107.

Anonymous:2019:EBh

[2819] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 154(??):ii, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303433.

Muro:2019:RCT

[2820] Francisco-Javier Moreno Muro, Nina Skorin-Kapov, and Pablo Pavon-Marino. Revisiting core traffic growth in the presence of expanding CDNs. Computer Networks (Amsterdam, Netherlands: 1999), 154 (??):1–11, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309423.

Sayit:2019:DQS

[2821] Muge Sayit, Cihat Cetinkaya, Huseyin Ugur Yildiz, and Bulent Tavli. DASH-QoS: a scalable network layer service differentiation architecture for DASH over SDN. Computer Networks (Amsterdam, Netherlands: 1999), 154 (??):12–25, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309241.

Bruschi:2019:AAF

[2822] Roberto Bruschi, Florin Ciucu, and Thomas Zinner. Application areas and fundamental challenges in Network Functions Virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 154(??):26–27, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619302348.

Halder:2019:EPI

[2823] Subir Halder, Amrita Ghosal, and Mauro Conti. Efficient physical intrusion detection in Internet of Things: a Node deployment approach. Computer Networks (Amsterdam, Netherlands: 1999), 154(??):28–46, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309952.

Lv:2019:CSB

[2824] Cuicui Lv, Qiang Wang, Wenjie Yan, and Jia Li. Compressive sensingbased sequential data gathering in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 154 (??):47–59, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619302671.

Guo:2019:STE

[2825] Yingya Guo, Zhiliang Wang, Zhifeng Liu, Xia Yin, Xingang Shi, Jianping Wu, Yang Xu, and H. Jonathan Chao. SOTE: Traffic engineering in hybrid software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 154(??):60–72, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302603.

Yu:2019:NHL

[2826] Xiuwu Yu, Lixing Zhou, and Xiangyang Li. A novel hybrid localization scheme for deep mine based on wheel graph and chicken swarm optimization. Computer Networks (Amsterdam, Netherlands: 1999), 154 (??):73–78, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302142.

Kumar:2019:EEI

[2827] Shishupal Kumar, Nidhi Lal, and Vijay Kumar Chaurasiya. An energy efficient IPv6 packet delivery scheme for industrial IoT over G.9959 protocol based Wireless Sensor Network (WSN). Computer Networks (Amsterdam, Netherlands: 1999), 154 (??):79–87, May 8, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302634.

Ali:2019:TES

[2828] G. G. Md. Nawaz Ali, Kai Liu, Victor C. S. Lee, Peter H. J. Chong, Yong Liang Guan, and Jun Chen. Towards efficient and scalable implementation for coding-based on-demand data broadcast. Computer Networks (Amsterdam, Netherlands: 1999), 154 (??):88–104, May 8, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861930221X.

Anonymous:2019:EBi

[2829] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 155(??):ii, May 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619304256.

Kang:2019:ATE

[2830] Heedo Kang, Changhoon Yoon, and Seungwon Shin. Astraea: Towards an effective and usable application permission system for SDN. Computer Networks (Amsterdam, Netherlands: 1999), 155(??):1–14, May 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618306030.

Casares-Giner:2019:PEF

[2831] Vicente Casares-Giner, Jorge Martinez-Bauset, and Canek Portillo. Performance evaluation of framed slotted ALOHA with reservation packets and successive interference cancelation for M2M networks. Computer Networks (Amsterdam, Netherlands: 1999), 155(??):15–30, May 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308867.

Liu:2019:EDD

[2832] Zhizhu Liu, Yinqiao Xiong, Xin Liu, Wei Xie, and Peidong Zhu. 6Tree:

Efficient dynamic discovery of active addresses in the IPv6 address space. Computer Networks (Amsterdam, Netherlands: 1999), 155 (??):31–46, May 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312003.

Casazza:2019:ADN

[2833] Marco Casazza, Mathieu Bouet, and Stefano Secci. Availability-driven NFV orchestration. Computer Networks (Amsterdam, Netherlands: 1999), 155 (??):47–61, May 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930235X.

Zhang:2019:DAS

[2834] Zufan Zhang, Zhangyi Wang, Chenquan Gan, and Porui Zhang. A double auction scheme of resource allocation with social ties and sentiment classification for Device-to-Device communications. Computer Networks (Amsterdam, Netherlands: 1999), 155 (??):62–71, May 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304559.

Sukapuram:2019:PPP

[2835] Radhika Sukapuram and Gautam Barua. PPCU: Proportional per-packet consistent updates for SDNs using data plane time stamps. Computer Networks (Amsterdam, Netherlands: 1999), 155(??):72–86, May 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618310338.

Al-Nabhan:2019:HIB

[2836] Najla Al-Nabhan, Nadia Al-Aboody, and A. B. M. Alim Al Islam. A hybrid IoT-based approach for emergency evacuation. Computer Networks (Amsterdam, Netherlands: 1999), 155 (??):87–97, May 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830728X.

Zakariya:2019:OGT

[2837] Ahmed Y. Zakariya, Sherif I. Rabia, and Yasmine Abouelseoud. An optimized general target channel sequence for prioritized cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 155 (??):98–109, May 22, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618303116.

Yamini:2019:EEE

[2838] K. Anish Pon Yamini, K. Suthendran, and T. Arivoli. Enhancement of energy efficiency using a transition state MAC protocol for MANET. Computer Networks (Amsterdam, Netherlands: 1999), 155(??):110–118, May 22, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308090.

Anonymous:2019:EBj

[2839] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 156(??):ii, June 19, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305031.

Qi:2019:SBI

[2840] Yuying Qi, Xuanxia Yao, Tao Zhu, and Huansheng Ning. A semanticbased inference control algorithm for OWL repository privacy protection. Computer Networks (Amsterdam, Netherlands: 1999), 156 (??):1–8, June 19, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303731.

Smyth:2019:ADS

[2841] Dylan Smyth, Donna O’Shea, Victor Cionca, and Sean McSweeney. Attacking distributed software-defined networks by leveraging network state consistency. Computer Networks (Amsterdam, Netherlands: 1999), 156 (??):9–19, June 19, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830940X.

Shaverdian:2019:SEN

[2842] Ararat Shaverdian, Jagadish Ghimire, and Catherine Rosenberg. Simple and efficient network-aware user association rules for heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 156 (??):20–32, June 19, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304183.

Vishnuvarthan:2019:EED

[2843] R. Vishnuvarthan, R. Sakthivel, V. Bhanumathi, and K. Muralitharan. Energy-efficient data collection in strip-based wireless sensor networks with optimal speed mobile data collectors. Computer Networks (Amsterdam, Netherlands: 1999), 156 (??):33–40, June 19, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312544.

Peng:2019:MTO

[2844] Zhenlong Peng, Xiaolin Gui, Jian An, Tianjie Wu, and Ruowei Gui. Multi-task oriented data diffusion and transmission paradigm in crowdsensing based on city public traffic. Computer Networks (Amsterdam, Netherlands: 1999), 156(??):41–51, June 19, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830505X.

Sharafeddine:2019:DDM

[2845] Sanaa Sharafeddine and Rania Islambouli. On-demand deployment of multiple aerial base stations for traffic offloading and network recovery. Computer Networks (Amsterdam, Netherlands: 1999), 156 (??):52–61, June 19, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304328.

Araujo:2019:SMO

[2846] Francisco Renato C. Araújo, Antonio M. de Sousa, and Leobino N. Sampaio. SCaN-Mob: an opportunistic caching strategy to support producer mobility in named data wireless networking. Computer Networks (Amsterdam, Netherlands: 1999), 156 (??):62–74, June 19, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311599.

Anonymous:2019:EBk

[2847] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 157(??):ii, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930595X.

AL-Tam:2019:FSM

[2848] F. AL-Tam and N. Correia. Fractional switch migration in multi-controller software-defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 157(??):1–10, July 5, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930475X.

Bastos:2019:DBS

[2849] Ian Vilar Bastos and Igor Monteiro Moraes. A diversity-based search-and-routing approach for named-data networking. Computer Networks (Amsterdam, Netherlands: 1999), 157 (??):11–23, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311952.

Yan:2019:MCN

[2850] Zhiwei Yan and Jong-Hyouk Lee. Mobility capability negotiation for IPv6 based ubiquitous mobile Internet. Computer Networks (Amsterdam, Netherlands: 1999), 157 (??):24–28, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311496.

Huynh:2019:OID

[2851] Van-Van Huynh, Hoang-Sy Nguyen, Ly Tran Thai Hoc, Thanh-Sang Nguyen, and Miroslav Voznak. Optimization issues for data rate in energy harvesting relay-enabled cognitive sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 157(??):29–40, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307898.

Cheng:2019:AIM

[2852] Yujun Cheng, Dong Yang, Huachun Zhou, and Hongchao Wang. Adopting IEEE 802.11 MAC for industrial delaysensitive wireless control and monitoring applications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 157(??):41–67, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308016.

Xia:2019:HSS

[2853] Changqing Xia, Xi Jin, Linghe Kong, Chi Xu, and Peng Zeng. Heterogeneous slot scheduling for real-time industrial wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 157(??):68–77, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861831243X.

Hashim:2019:LML

[2854] Hashim A. Hashim and Mohammad A.

Abido. Location management in LTE networks using multi-objective particle swarm optimization. Computer Networks (Amsterdam, Netherlands: 1999), 157(??):78–88, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300891.

Ji:2019:JTS

[2855] Cun Ji, Chao Zhao, Li Pan, Shijun Liu, Chenglei Yang, and Xiangxu Meng. A just-in-time shapelet selection service for online time series classification. Computer Networks (Amsterdam, Netherlands: 1999), 157 (??):89–98, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304840.

Xiaofeng:2019:AAS

[2856] Lu Xiaofeng, Jiang Fangshuo, Zhou Xiao, Yi Shengwei, Sha Jing, and Pietro Lio. ASSCA: API sequence and statistics features combined architecture for malware detection. Computer Networks (Amsterdam, Netherlands: 1999), 157(??):99–111, July 5, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930461X.

Anithaashri:2019:SES

[2857] T. P. Anithaashri, G. Ravichandran, and R. Baskaran. Security enhancement for software defined network using game theoretical approach. Computer Networks (Amsterdam, Netherlands: 1999), 157 (??):112–121, July 5, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830820X.

Abdel-Basset:2019:NTB

[2858] Mohamed Abdel-Basset, Gunasekaran Manogaran, and Mai Mohamed. A neutrosophic theory based security approach for fog and mobile-edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 157 (??):122–132, July 5, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310727.

Shi:2019:ALB

[2859] Qingyu Shi, Fang Wang, Dan Feng, and Weibin Xie. Adaptive load balancing based on accurate congestion feedback for asymmetric topologies. Computer Networks (Amsterdam, Netherlands: 1999), 157(??):133–145, July 5, 2019.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311794.

Dong:2019:LAM

[2860] Enhuan Dong, Mingwei Xu, Xiaoming Fu, and Yu Cao. A loss aware MPTCP scheduler for highly lossy networks. Computer Networks (Amsterdam, Netherlands: 1999), 157 (??):146–158, July 5, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619301367.

Anonymous:2019:EBl

[2861] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 158(??):ii, July 20, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306619.

Ren:2019:HCC

[2862] Dewang Ren, Xiaolin Gui, Kaiyuan Zhang, and Jie Wu. Hybrid collaborative caching in mobile edge networks: an analytical approach. Computer Networks (Amsterdam, Netherlands: 1999), 158(??):1–16, July 20, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308375.

Hejja:2019:OOP

[2863] Khaled Hejja and Xavier Hesselbach.

Offline and online power aware resource allocation algorithms with migration and delay constraints. Computer Networks (Amsterdam, Netherlands: 1999), 158(??):17–34, July 20, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314063.

Hosseini:2019:HTD

[2864] Soodeh Hosseini and Mehrdad Azizi. The hybrid technique for DDoS detection with supervised learning algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):35–45, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306881.

Gajewski:2019:TTA

[2865] Mariusz Gajewski, Jordi Mongay Batalla, Albert Levi, Cengiz Togay, Constandinos X. Mavromoustakis, and George Mastorakis. Twotier anomaly detection based on traffic profiling of the home automation system. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):46–60, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311587.

Wang:2019:IMV

[2866] Huan Wang, Zhaolin Yuan, Yibin Chen, Bingyang Shen, and Aixiang Wu. An industrial missing values processing method based on generating model. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):61–68, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619301781.

Venkataraman:2019:DAA

[2867] N. L. Venkataraman and R. Kumar. Design and analysis of application specific network on chip for reliable custom topology. Computer Networks (Amsterdam, Netherlands: 1999), 158(??):69–76, July 20, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307254.

Xia:2019:MOH

[2868] Ming Xia, Kaikai Chi, Xiaoyan Wang, and Zhen Cheng. Mode-oriented hybrid programming of sensor network nodes for supporting rapid and flexible utility assembly. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):77–97, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307424.

Soualah:2019:OBA

[2869] Oussama Soualah, Marouen Mechtri, Chaima Ghribi, and Djamal Zeghlache. Online and batch algorithms for VNFs placement and chaining. Computer Networks (Amsterdam, Netherlands: 1999), 158(??):98–113, July 20, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302548.

Han:2019:NCH

[2870] Tao Han, Lijuan Zhang, Sandeep Pirbhulal, Wanqing Wu, and Victor Hugo C. de Albuquerque. A novel cluster head selection technique for edge-computing based IoMT systems. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):114–122, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618310764.

Agarwal:2019:SSM

[2871] Amit Agarwal and Durga Toshniwal. SmPFT: Social media based profile fusion technique for data enrichment. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):123–131, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305929.

Wu:2019:MUC

[2872] Jiehong Wu, Liangkai Zou, Liang Zhao, Ahmed Al-Dubai, Lewis Mackenzie, and Geyong Min. A multi-UAV clustering strategy for reducing insecure communication range. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):132–142, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312891.

Dias:2019:IAR

[2873] Klenilmar Lopes Dias, Mateus Almeida Pongelupe, Walmir Matos Caminhas, and Luciano de Errico. An innovative approach for real-time network traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):143–157, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305541.

Wu:2019:RRF

[2874] Bo Wu, Ke Xu, Qi Li, Bingyang Liu, Shoushou Ren, Fan Yang, Meng Shen, and Kui Ren. RFL: Robust fault localization on unreliable communication channels. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):158–174, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618304109.

Ehsanpour:2019:ESA

[2875] Mahsa Ehsanpour, Siavash Bayat, and Ali Mohammad Afshin Hemmatyar. An efficient and social-aware distributed in-network caching scheme in named data networks using matching theory. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):175–183, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305863.

Wei:2019:CPI

[2876] Mingkui Wei, Zhuo Lu, Yufei Tang, and Xiang Lu. Cyber and physical interactions to combat failure propagation in smart grid: Characterization, analysis and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 158 (??):184–192, July 20, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619301306.

Tian:2019:SDC

[2877] Bingchuan Tian, Chen Tian, Bingquan Wang, Bo Li, Zehao He, Haipeng Dai, Kexin Liu, Wanchun Dou, and Guihai Chen. Scheduling dependent coflows to minimize the total weighted job completion time in datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 158(??):193–205, July 20, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619302580.

Anonymous:2019:EBm

[2878] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 159(??):ii, August 4, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619307273.

Li:2019:CEB

[2879] Kun Li, Shengling Wang, and Xiuzhen Cheng. Crowdsourcee evaluation based on persuasion game. Computer Networks (Amsterdam, Netherlands: 1999), 159(??):1–9, August 4, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619304839.

Lassila:2019:OEA

[2880] Pasi Lassila, Misikir Eyob Gebrehiwot, and Samuli Aalto. Optimal energy-aware load balancing and base station switch-off control in 5G HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):10–22, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618307795.

Arfaoui:2019:CAA

[2881] Amel Arfaoui, Ali Kribeche, and Sidi-Mohammed Senouci. Context-aware anonymous authentication protocols in the Internet of Things dedicated to e-health applications. Computer Networks (Amsterdam, Netherlands: 1999), 159(??):23–36, August 4, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300465.

Uddin:2019:EEM

[2882] Md. Forkan Uddin. Energy efficiency maximization by joint transmission scheduling and resource allocation in downlink NOMA cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):37–50, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314099.

Yi:2019:MCD

[2883] Bo Yi, Xingwei Wang, Min Huang, and Anwei Dong. A multi-criteria decision approach for minimizing the influence of VNF migration. Computer Networks (Amsterdam, Netherlands: 1999), 159(??):51–62, August 4, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304748.

Yao:2019:ICA

[2884] Xuanxia Yao, Jiafei Wang, Mengyu Shen, Huafeng Kong, and Huansheng Ning. An improved clustering algorithm and its application in IoT data analysis. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):63–72, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304852.

Junior:2019:NSU

[2885] Eduardo P. M. Câmara Júnior, Luiz F. M. Vieira, and Marcos A. M. Vieira. 3DVS: Node scheduling in underwater sensor networks using 3D Voronoi diagrams. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):73–83, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930492X.

Anadiotis:2019:SWS

[2886] Angelos-Christos Anadiotis, Laura Galluccio, Sebastiano Milardo, Giacomo Morabito, and Sergio Palazzo. SD-WISE: a Software-Defined WIreless SEnsor network. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):84–95, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312192.

Caminero:2019:AER

[2887] Guillermo Caminero, Manuel Lopez-Martin, and Belen Carro. Adversarial environment reinforcement learning algorithm for intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):96–109, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311216.

Liu:2019:RRA

[2888] Zhixin Liu, Peng Zhang, Kit Yan Chan, Li Li, and Xinping Guan. Robust resource allocation for rates maximization using fuzzy estimation of dynamic channel states in OFDMA femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):110–124, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312593.

Nayyer:2019:LSC

[2889] Amit Nayyer, Aman Kumar Sharma, and Lalit Kumar Awasthi. Laman: a supervisor controller based scalable framework for software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):125–134, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309307.

Fasoulakis:2019:SIM

[2890] Michail Fasoulakis, Eirini Eleni Tsiropoulou, and Symeon Papavassiliou. Satisfy instead of maximize: Improving operation efficiency in wireless communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):135–146, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618313896.

Elhoseny:2019:EFC

[2891] Mohamed Elhoseny, Gui-Bin Bian, S. K. Lakshmanaprabu, K. Shankar, Amit Kumar Singh, and Wanqing Wu. Effective features to classify ovarian cancer data in Internet of Medical Things. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):147–156, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310144.

Bao:2019:CMC

[2892] Haizhou Bao, Chuanhe Huang, Zhongzheng Tang, Qihui Li, Qiufen Ni, Xiaodai Dong, and Bin Fu. Coded multicasting in cache-enabled vehicular ad hoc network. Computer Networks (Amsterdam, Netherlands: 1999), 159 (??):157–170, August 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314336.

Anonymous:2019:EBn

[2893] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 160(??):ii, September 4, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619308965.

Jenkins:2019:CNP

[2894] Sarah Jenkins. Computer Networks:

Publisher’s note. Computer Networks (Amsterdam, Netherlands: 1999), 160 (??):iii, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309028.

Anonymous:2019:PB

[2895] Anonymous. Passing the BA-TON. Computer Networks (Amsterdam, Netherlands: 1999), 160 (??):iv–v, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930903X.

Anonymous:2019:RB

[2896] Anonymous. Receiving the BA-TON. Computer Networks (Amsterdam, Netherlands: 1999), 160 (??):vi, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309041.

Ghannadrezaii:2019:MNC

[2897] Hossein Ghannadrezaii and Jean-François Bousquet. Maximizing network coverage in a multichannel shortrange underwater acoustic sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 160 (??):1–10, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306097.

Rahman:2019:EEO

[2898] Akhlaqur Rahman, Jiong Jin, Ashfaqur Rahman, Antonio Cricenti, Mahbuba Afrin, and Yu ning Dong. Energy-efficient optimal task offloading in cloud networked multi-robot systems. Computer Networks (Amsterdam, Netherlands: 1999), 160(??): 11–32, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306371.

Krishnan:2019:ECA

[2899] Muralitharan Krishnan, Sangwoon Yun, and Yoon Mo Jung. Enhanced clustering and ACO-based multiple mobile sinks for efficiency improvement of wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 160(??): 33–40, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300234.

Saxena:2019:AAD

[2900] Apoorv Saxena, Dieter Claeys, Herwig Bruneel, and Joris Walraevens. Analysis of the age of data in data backup systems. Computer Networks (Amsterdam, Netherlands: 1999), 160 (??):41–50, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308727.

Kale:2019:SDA

[2901] Preeti A. Kale and Manisha J.

Nene. Scheduling of data aggregation trees using local heuristics to enhance network lifetime in sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 160(??): 51–64, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306522.

Huang:2019:HTW

[2902] Gaofei Huang and Wanqing Tu. A high-throughput wireless-powered relay network with joint time and power allocations. Computer Networks (Amsterdam, Netherlands: 1999), 160(??): 65–76, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314087.

Gunleifsen:2019:DSI

[2903] Håkon Gunleifsen, Thomas Kemmerich, and Vasileios Gkioulos. Dynamic setup of IPsec VPNs in service function chaining. Computer Networks (Amsterdam, Netherlands: 1999), 160 (??):77–91, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300969.

Liu:2019:RPC

[2904] Zhixin Liu, Guochen Hou, Yang Liu, Xinbin Li, and Xinping Guan. Robust power control strategy based on hierarchical game with QoS provisioning in full-duplex femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 160(??): 92–104, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830481X.

Cicioglu:2019:HNI

[2905] Murtaza Cicioglu and Ali Çalhan. HUBsFLOW: a novel interface protocol for SDN-enabled WBANs. Computer Networks (Amsterdam, Netherlands: 1999), 160(??):105– 117, September 4, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619303573.

Bahtiyar:2019:MDM

[2906] Serif Bahtiyar, Mehmet Baris Yaman, and Can Yilmaz Altinigne. A multi-dimensional machine learning approach to predict advanced malware. Computer Networks (Amsterdam, Netherlands: 1999), 160(??): 118–129, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861831082X.

Ha:2019:ABU

[2907] Duc Thang Ha, Lila Boukhatem, Megumi Kaneko, Nhan Nguyen-Thanh, and Steven Martin. Adaptive beamforming and user association in heterogeneous cloud radio access networks: a mobility-aware performancecost trade-off. Computer Networks (Amsterdam, Netherlands: 1999), 160 (??):130–143, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830714X.

Ghasemi:2019:ENC

[2908] Maryam Ghasemi, Ibrahim Matta, and Flavio Esposito. The effect of (non-)competing brokers on the quality and price of differentiated Internet services. Computer Networks (Amsterdam, Netherlands: 1999), 160(??): 144–164, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307613.

Hajiheidari:2019:IDS

[2909] Somayye Hajiheidari, Karzan Wakil, Maryam Badri, and Nima Jafari Navimipour. Intrusion detection systems in the Internet of things: a comprehensive investigation. Computer Networks (Amsterdam, Netherlands: 1999), 160 (??):165–191, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306267.

Santhanavijayan:2019:MSO

[2910] A. Santhanavijayan and S. R. Balasundaram. Multi swarm optimization based automatic ontology for eassessment. Computer Networks (Amsterdam, Netherlands: 1999), 160(??): 192–199, September 4, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830896X.

Anonymous:2019:EBo

[2911] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 161(??):ii, October 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619310485.

Shehaj:2019:LDH

[2912] Marinela Shehaj, Dritan Nace, Ilya Kalesnikau, and Michał Pióro. Link dimensioning of hybrid FSO/fiber networks resilient to adverse weather conditions. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):1–13, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313926.

Pfitscher:2019:GPA

[2913] Ricardo José Pfitscher, Arthur Selle Jacobs, Luciano Zembruzki, Ricardo Luis dos Santos, Eder John Scheid, Muriel Figueredo Franco, Alberto Schaeffer-Filho, and Lisandro Zambenedetti Granville. Guiltiness: a practical approach for quantifying virtual network functions performance. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):14–31, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310508.

Chahal:2019:NSD

[2914] Manisha Chahal and Sandeep Harit.

Network selection and data dissemination in heterogeneous software-defined vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):32–44, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300866.

Alaluna:2019:SMC

[2915] Max Alaluna, Eric Vial, Nuno Neves, and Fernando M. V. Ramos. Secure multi-cloud network virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 161(??): 45–60, October 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312155.

Kocan:2019:MDM

[2916] Enis Kocan, Aleksandra Lopusina, and Milica Pejanovic-Djurisic. Macro diversity for mmWave cellular communications in indoor environment. Computer Networks (Amsterdam, Netherlands: 1999), 161(??):61–67, October 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618313100.

Leng:2019:LPE

[2917] Xue Leng, Kaiyu Hou, Yan Chen, Kai Bu, Libin Song, and You Li. A lightweight policy enforcement system for resource protection and management in the SDN-based cloud. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):68–81, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314129.

Liu:2019:EER

[2918] Zhixin Liu, Changjian Liang, Yazhou Yuan, and Xinping Guan. Energy efficient resource allocation based on relay selection and subcarrier pairing with channel uncertainty in cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):82–92, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309599.

Titouna:2019:DDO

[2919] Chafiq Titouna, Farid Naït-Abdesselam, and Ashfaq Khokhar. DODS: a distributed outlier detection scheme for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 161(??):93–101, October 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861930012X.

Liu:2019:RLB

[2920] Wenbin Liu, Leye Wang, En Wang, Yongjian Yang, Djamal Zeghlache, and Daqing Zhang. Reinforcement learning-based cell selection in sparse mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 161(??):102–114, October 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619303706.

Wang:2019:SRS

[2921] Jessie Hui Wang, Jilong Wang, Changqing An, and Qianli Zhang. A survey on resource scheduling for data transfers in inter-datacenter WANs. Computer Networks (Amsterdam, Netherlands: 1999), 161(??): 115–137, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303931.

Wei:2019:QLA

[2922] Zhenchun Wei, Fei Liu, Yan Zhang, Juan Xu, Jianjun Ji, and Zengwei Lyu. A Q-learning algorithm for task scheduling based on improved SVM in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 161(??): 138–149, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309289.

Yang:2019:PDR

[2923] Bin Yang, Zhenqiang Wu, Yulong Shen, and Xiaohong Jiang. Packet delivery ratio and energy consumption in multicast delay tolerant MANETs with power control. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):150–161, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312908.

Ak:2019:BPC

[2924] Elif Ak and Berk Canberk. BCDN: a proof of concept model for blockchainaided CDN orchestration and routing. Computer Networks (Amsterdam, Netherlands: 1999), 161(??): 162–171, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313549.

Huang:2019:RDF

[2925] Hailong Huang and Andrey V. Savkin. Reactive 3D deployment of a flying robotic network for surveillance of mobile targets. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):172–182, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619301604.

Antonioli:2019:ABS

[2926] Roberto P. Antonioli, Emanuel B. Rodrigues, Diego A. Sousa, Igor M. Guerreiro, Carlos F. M. e Silva, and Francisco R. P. Cavalcanti. Adaptive bearer split control for 5G multi-RAT scenarios with dual connectivity. Computer Networks (Amsterdam, Netherlands: 1999), 161(??): 183–196, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308527.

Wilmet:2019:ODI

[2927] Audrey Wilmet, Tiphaine Viard, Matthieu Latapy, and Robin Lamarche-Perrin. Outlier detection in IP traffic modelled as a link stream using the stability of degree distributions over time. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):197–209, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619301045.

W:2019:MSD

[2928] Wei W, Xu Xia, Marcin Wozniak, Xunli Fan, Robertas Damaševičius, and Ye Li. Multi-sink distributed power control algorithm for cyberphysical-systems in coal mine tunnels. Computer Networks (Amsterdam, Netherlands: 1999), 161(??): 210–219, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310673.

Liu:2019:RAS

[2929] Xin Liu, Ruisheng Zhang, and Mingqi Zhao. A robust authentication scheme with dynamic password for wireless body area networks. Computer Networks (Amsterdam, Netherlands: 1999), 161(??): 220–234, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309885.

Zhan:2019:FDN

[2930] Yufeng Zhan, Yuanqing Xia, and Athanasios V. Vasilakos. Future directions of networked control systems: a combination of cloud control and fog control approach. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):235–248, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308515.

Manogaran:2019:MLA

[2931] Gunasekaran Manogaran, Naveen Chilamkurti, and Ching-Hsien Hsu. Machine learning algorithms towards merging of mobile edge computing and Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):249–250, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308734.

Wang:2019:VTV

[2932] Sen Wang, Jun Bi, Jianping Wu, Athanasios V. Vasilakos, and Qilin Fan. VNE-TD: a virtual network embedding algorithm based on temporal-difference learning. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):251–263, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830584X.

Hu:2019:AAB

[2933] Tao Hu, Peng Yi, Julong Lan, Yuxiang Hu, and Penghao Sun. ACST:

Audit-based compromised switch tolerance for enhancing data plane robustness in software-defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 161(??): 264–280, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300180.

Sun:2019:IKS

[2934] Yunchuan Sun, Xiuzhen Cheng, Yu Bai, and Jiguo Yu. Information, knowledge, and semantics for interacting with Internet-of-Things. Computer Networks (Amsterdam, Netherlands: 1999), 161 (??):281–282, October 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308291.

Al-Salti:2019:ERG

[2935] Faiza Al-Salti, Nasser Alzeidi, Khaled Day, and Abderezak Touzene. An efficient and reliable grid-based routing protocol for UWSNs by exploiting minimum hop count. Computer Networks (Amsterdam, Netherlands: 1999), 162 (??):Article 106869, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302950.

Alfadhli:2019:LPA

[2936] Yahya Alfadhli, You-Wei Chen, Siming Liu, Shuyi Shen, Shuang Yao, Daniel Guidotti, Sufian Mitani, and Gee-Kung Chang. Latency performance analysis of low layers function split for URLLC applications in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106865, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619301343.

Anonymous:2019:EBp

[2937] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 162(??):Article 106911, ???? 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619312022.

Anonymous:2019:PNa

[2938] Anonymous. Publisher note. Computer Networks (Amsterdam, Netherlands: 1999), 162(??):Article 106912, ???? 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619312034.

Bagci:2019:SED

[2939] K. Tolga Bagci and A. Murat Tekalp. SDN-enabled distributed open exchange: Dynamic QoS-path optimization in multi-operator services. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106845, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618308387.

Boussada:2019:PPA

[2940] Rihab Boussada, Balkis Hamdane, Mohamed Elhoucine Elhdhili, and Leila Azouz Saidane. Privacypreserving aware data transmission for IoT-based e-health. Computer Networks (Amsterdam, Netherlands: 1999), 162(??):Article 106866, ???? 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619305730.

Brincat:2019:UBT

[2941] Alberto Attilio Brincat, Alfio Lombardo, Giacomo Morabito, and Salvatore Quattropani. On the use of Blockchain technologies in WiFi networks. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106855, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306073.

Chithaluru:2019:AAR

[2942] Premkumar Chithaluru, Rajeev Tiwari, and Kamal Kumar. AREOR

— adaptive ranking based energy efficient opportunistic routing scheme in wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 162(??):Article 106863, ???? 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861930430X.

Golchi:2019:HFI

[2943] Mahya Mohammadi Golchi, Shideh Saraeian, and Mehrnoosh Heydari. A hybrid of firefly and improved particle swarm optimization algorithms for load balancing in cloud environments: Performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106860, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313665.

Kantor:2019:SML

[2944] Mirosław Kantor, Edyta Biernacka, Piotr Boryło, Jerzy Domżał, Piotr Jurkiewicz, Miłosz Stypiński, and Robert Wójcik. A survey on multi-layer IP and optical software-defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106844, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618314312.

Khalily-Dermany:2019:CTC

[2945] M. Khalily-Dermany, M. J. Nadjafi-Arani, and S. Doostali. Combining topology control and network coding to optimize lifetime in wireless-sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106859, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308825.

Kheirkhah:2019:MTP

[2946] Morteza Kheirkhah, Ian Wakeman, and George Parisis. Multipath transport and packet spraying for efficient data delivery in data centres. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106852, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300374.

Morcos:2019:ERR

[2947] Mira Morcos, Jocelyne Elias, Fabio Martignon, Tijani Chahed, and Lin Chen. On efficient radio resource calendaring in cloud radio access network. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106862, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306832.

Oh:2019:COT

[2948] Seungmin Oh, Hyunchong Cho, Sang-Ha Kim, Wanseop Lee, and Euisin Lee. Continuous object tracking protocol with selective wakeup based on practical boundary prediction in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 162 (??):Article 106854, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310843. See

corrigendum [2997].

Pimpinella:2019:WWI

[2949] Andrea Pimpinella, Alessandro E. C.

Redondi, and Matteo Cesana. Walk this way! An IoT-based urban routing system for smart cities. Computer Networks (Amsterdam, Netherlands: 1999), 162(??):Article 106857, ???? 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313094.

Sicari:2019:BST

[2950] S. Sicari, A. Rizzardi, G. Piro, A. Coen-Porisini, and L. A. Grieco. Beyond the smart things: Towards the definition and the performance assessment of a secure architecture for the Internet of Nano-Things. Computer Networks (Amsterdam, Netherlands: 1999), 162 (??):Article 106856, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305875.

Tajiki:2019:JFR

[2951] Mohammad M. Tajiki, Mohammad Shojafar, Behzad Akbari, Stefano Salsano, Mauro Conti, and Mukesh Singhal. Joint failure recovery, fault prevention, and energy-efficient resource management for real-time SFC in fogsupported SDN. Computer Networks (Amsterdam, Netherlands: 1999), 162 (??):Article 106850, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308552.

Xie:2019:DDS

[2952] An Xie, Huawei Huang, Xiaoliang Wang, Song Guo, Zhuzhong Qian, and Sanglu Lu. Dual: Deploy stateful virtual network function chains by jointly allocating data-control traffic. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106868, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303962.

Yang:2019:IDR

[2953] Yan Yang, Xia Yin, Xingang Shi, Zhiliang Wang, Jiong He, Tom Z. J. Fu, and Marianne Winslett. Inter-domain routing bottlenecks and their aggravation. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106839, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310466.

Yu:2019:FDF

[2954] Yinbo Yu, Xing Li, Kai Bu, Yan Chen, and Jianfeng Yang. Falcon: Differential fault localization for SDN control plane. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106851, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300040.

Zhang:2019:CTM

[2955] Rui Zhang, Peng Xie, Chen Wang, Gaoyang Liu, and Shaohua Wan. Classifying transportation mode and speed from trajectory data via deep multiscale learning. Computer Networks (Amsterdam, Netherlands: 1999), 162 (??):Article 106861, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314397.

Zhang:2019:TSN

[2956] Shunliang Zhang, Yongming Wang, and Weihua Zhou. Towards secure 5G networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 162(??):Article 106871, ???? 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S138912861830817X.

Zhao:2019:CNN

[2957] Bobai Zhao, Dali Zhu, Tong Xi, Chenggang Jia, Shang Jiang, and Siye Wang. Convolutional neural network and dual-factor enhanced variational Bayes adaptive Kalman filter based indoor localization with Wi-Fi. Computer Networks (Amsterdam, Netherlands: 1999), 162(??): Article 106864, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618311691.

Akdogan:2019:SKA

[2958] Dilara Akdogan, Duygu Karaoglan Altop, and Albert Levi. Secure key agreement based on ordered biometric features. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106885, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313975.

Anonymous:2019:EBq

[2959] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 163(??):Article 106931, ???? 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619312782.

Arana:2019:AET

[2960] O. Arana, F. Garcia, and J. Gomez.

Analysis of the effectiveness of transmission power control as a location privacy technique. Computer Networks (Amsterdam, Netherlands: 1999), 163 (??):Article 106880, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311034.

Arfaoui:2019:GBA

[2961] Amel Arfaoui, Ali Kribeche, Sidi Mohammed Senouci, and Mohamed Hamdi. Game-based adaptive anomaly detection in wireless body area networks. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106870, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300015.

Asghar:2019:FPP

[2962] Hassan Jameel Asghar, Emiliano De Cristofaro, Guillaume Jourjon, Mohammed Ali Kaafar, Laurent Mathy, Luca Melis, Craig Russell, and Mang Yu. Fast privacypreserving network function outsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106893, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309873.

Chriki:2019:FCM

[2963] Amira Chriki, Haifa Touati, Hichem Snoussi, and Farouk Kamoun. FANET: Communication, mobility models and security issues. Computer Networks (Amsterdam, Netherlands: 1999), 163 (??):Article 106877, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309034.

Fouad:2019:OTW

[2964] Mohamed Mostafa Fouad, Ahmed Ibrahem Hafez, and Aboul Ella Hassanien. Optimizing topologies in wireless sensor networks: a comparative analysis between the Grey Wolves and the Chicken Swarm Optimization algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106882, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619300751.

Killi:2019:CPS

[2965] Bala Prakasa Rao Killi and Seela Veerabhadreswara Rao. Controller placement in software defined networks: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 163(??):Article 106883, ???? 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618313872.

Kim:2019:SSD

[2966] Yeonkeun Kim, Jaehyun Nam, Taejune Park, Sandra Scott-Hayward, and Seungwon Shin. SODA: a softwaredefined security framework for IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106889, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307522.

Li:2019:MMM

[2967] Shuangjuan Li and Hong Shen. Minimizing maximum movement of sensors for line barrier coverage in the plane. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106841, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308114.

Lin:2019:EEP

[2968] Jinting Lin, Lin Cui, Yuxiang Zhang, Fung Po Tso, and Quanlong Guan. Extensive evaluation on the performance and behaviour of TCP congestion control protocols under varied network scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 163 (??):Article 106872, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311265.

Longo:2019:AOE

[2969] Edoardo Longo, Alessandro E. C.

Redondi, and Matteo Cesana. Accurate occupancy estimation with WiFi and bluetooth/BLE packet capture. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106876, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618313045.

Mahmoudi:2019:UFH

[2970] Mohsen Mahmoudi, Karim Faez, and Abdorasoul Ghasemi. Uncoordinated frequency hopping scheme for defense against primary user emulation attack in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106884, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618313999.

Noura:2019:DRE

[2971] Hassan N. Noura, Reem Melki, Mohammad Malli, and Ali Chehab. Design and realization of efficient and secure multi-homed systems based on random linear network coding. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106886, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619301501.

Oliveira:2019:ECM

[2972] Roberto M. Oliveira, Lucas G. de Oliveira, Alex B. Vieira, and Moises V. Ribeiro. An enhanced cooperative MAC protocol for hybrid PLC/ wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 163 (??):Article 106878, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930338X.

Zafar:2019:IBB

[2973] Saniya Zafar, Rasheed Hussain, Fatima Hussain, and Sobia Jangsher. Interplay between Big Spectrum Data and Mobile Internet of Things: Current solutions and future challenges. Computer Networks (Amsterdam, Netherlands: 1999), 163(??): Article 106879, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307376.

Anonymous:2019:Da

[2974] Anonymous. 9 December 2019. Computer Networks (Amsterdam, Netherlands: 1999), 164(??):??, December 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2019:EBr

[2975] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 164(??):Article 106964, December 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128619314173.

Xu:2019:RCR

[2976] Hongli Xu, Jianchun Liu, Chen Qian, He Huang, and Chunming Qiao. Reducing controller response time with hybrid routing in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106891, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310478.

Li:2019:EEC

[2977] Chunlin Li, Jianhang Tang, Yang Zhang, Xin Yan, and Youlong Luo. Energy efficient computation offloading for nonorthogonal multiple access assisted mobile edge computing with energy harvesting devices. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106890, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308077.

Sarkar:2019:CKN

[2978] Pinaki Sarkar and Sukumar Nandi. A class of key-node indexed hash chains based key predistribution (KPS): Signed weighted graphs. Computer Networks (Amsterdam, Netherlands: 1999), 164(??):Article 106881, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303669.

Amodu:2019:TCA

[2979] Oluwatosin Ahmed Amodu, Mohamed Othman, Nor Kamariah Noordin, and Idawaty Ahmad. Transmission capacity analysis of relay-assisted D2D cellular networks with M2M coexistence. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106887, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618313690.

Li:2019:PSP

[2980] Wenjun Li, Dagang Li, Xinwei Liu, Ting Huang, Xianfeng Li, Wenxia Le, and Hui Li. A power-saving pre-classifier for TCAM-based IP lookup. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106898, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311927.

Phung:2019:MRA

[2981] Chi-Dung Phung, Benevid Felix Silva, Michele Nogueira, and Stefano Secci. MPTCP robustness against large-scale man-in-the-middle attacks. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106896, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304566.

Goto:2019:QAS

[2982] Yuki Goto, Bryan Ng, Winston K. G.

Seah, and Yutaka Takahashi. Queueing analysis of software defined network with realistic OpenFlow-based switch model. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106892, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619301902.

Benyamina:2019:ANE

[2983] Zakarya Benyamina, Khelifa Benahmed, and Fateh Bounaama. ANEL: a novel efficient and lightweight authentication scheme for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 164 (??):Article 106899, December 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619306012.

Conti:2019:BUB

[2984] Mauro Conti, Muhammad Hassan, and Chhagan Lal. BlockAuth: BlockChain based distributed producer authentication in ICN. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106888, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308308.

Wang:2019:EMT

[2985] Xia Wang, Jia Liu, Yanyan Wang, Xingyu Chen, and Lijun Chen. Efficient missing tag identification in blocker-enabled RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 164(??):Article 106894, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861930670X.

Selimi:2019:PEI

[2986] Mennan Selimi, Adisorn Lertsinsrubtavee, Arjuna Sathiaseelan, Llorenç Cerdà-Alabern, and Leandro Navarro. PiCasso: Enabling information-centric multi-tenancy at the edge of community mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 164(??):Article 106897, December 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618312787.

Noormohammadpour:2019:EID

[2987] Mohammad Noormohammadpour, Srikanth Kandula, Cauligi S. Raghavendra, and Sriram Rao. Efficient inter-datacenter bulk transfers with mixed completion time objectives. Computer Networks (Amsterdam, Netherlands: 1999), 164 (??):Article 106903, December 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305729.

Qiu:2019:EEL

[2988] Yuqing Qiu, Chung-Horng Lung, Samuel Ajila, and Pradeep Srivastava. Experimental evaluation of LXC container migration for cloudlets using multipath TCP. Computer Networks (Amsterdam, Netherlands: 1999), 164 (??):Article 106900, December 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302403.

Mharsi:2019:ECO

[2989] Niezi Mharsi and Makhlouf Hadji.

Edge computing optimization for efficient RRH–BBU assignment in cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 164(??):Article 106901, December 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619302816.

Rajesh:2019:CAS

[2990] G. Rajesh and Ashvini Chaturvedi. Correlation analysis and statistical characterization of heterogeneous sensor data in environmental sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106902, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309995.

Le:2019:FDR

[2991] Tuan Le and Mario Gerla. Fragmented data routing based on exponentially distributed contacts and inter-contact times in DTNs. Computer Networks (Amsterdam, Netherlands: 1999), 164 (??):Article 106917, December 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861830731X.

Saverimoutou:2019:MAF

[2992] Antoine Saverimoutou, Bertrand Mathieu, and Sandrine Vaton. A 6month analysis of factors impacting web browsing quality for QoE prediction. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106905, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307546.

Eugui:2019:AHF

[2993] David Eugui and José Alberto Hernández. Analysis of a hybrid Fixed-Elastic DBA with guaranteed fronthaul delay in XG(s)- PONs. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106907, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302117.

Priyadarsini:2019:ALB

[2994] Madhukrishna Priyadarsini, Joy Chandra Mukherjee, Padmalochan Bera, Shailesh Kumar, A. H. M. Jakaria, and M. Ashiqur Rahman. An adaptive load balancing scheme for software-defined network controllers. Computer Networks (Amsterdam, Netherlands: 1999), 164(??):Article 106918, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930252X.

Liang:2019:PIT

[2995] Xiaoyu Liang and Taieb Znati. On the performance of intelligent techniques for intensive and stealthy DDos detection. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106906, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302968.

Ahmed:2019:CNS

[2996] Syed Hassan Ahmed, Ali Kashif Bashir, Awais Ahmad, and Wael Guibene. Computer networks special issue on intelligent and connected transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106895, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311223.

Oh:2019:CCO

[2997] Seungmin Oh, Hyunchong Cho, Sang-Ha Kim, Wanseop Lee, and Euisin Lee. Corrigendum to “Continuous object tracking protocol with selective wakeup based on practical boundary prediction in wireless sensor networks” Computer Networks Volume 162 (2019) 106854. Computer Networks (Amsterdam, Netherlands: 1999), 164(??): Article 106938, December 9, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313192. See

[2948].

Anonymous:2019:Db

[2998] Anonymous. 24 December 2019. Computer Networks (Amsterdam, Netherlands: 1999), 165(??):??, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2019:EBs

[2999] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 165(??):Article 107000, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315154.

Anonymous:2019:PNb

[3000] Anonymous. Publisher note. Computer Networks (Amsterdam, Netherlands: 1999), 165(??):Article 107001, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315166.

Eroglu:2019:DAC

[3001] Alperen Eroglu, Okan Yaman, and Ertan Onur. Density-aware cellular coverage control: Interference-based density estimation. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106922, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305201.

Wu:2019:UAV

[3002] Qiyue Wu, Peng Sun, and Azzedine Boukerche. Unmanned aerial vehicleassisted energy-efficient data collection scheme for sustainable wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106927, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302853.

Costa:2019:ILP

[3003] José Maurício Costa, Pedro P. Paniago, Joaquim de Andrade, Thiago F. Noronha, and Marcos A. M. Vieira. Integer linear programming formulations for the variable data rate and variable channel bandwidth scheduling problem in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106939, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304578.

Gokhale:2019:QCT

[3004] Vineet Gokhale, Jayakrishnan Nair, Subhasis Chaudhuri, and Jan Fesl. On QoS-compliant telehaptic communication over shared networks. Computer Networks (Amsterdam, Netherlands: 1999), 165(??):Article 106935, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306322.

Asghar:2019:CIC

[3005] Muhammad Rizwan Asghar, Qinwen Hu, and Sherali Zeadally. Cybersecurity in industrial control systems: Issues, technologies, and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106946, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306292.

Yang:2019:MWC

[3006] Ze Yang and Kwan L. Yeung. Minimum weight controller tree design in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106949, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618302822.

Aceto:2019:MME

[3007] Giuseppe Aceto, Domenico Ciuonzo, Antonio Montieri, and Antonio Pescapè. MIMETIC: Mobile encrypted traffic classification using multimodal deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106944, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304669.

Gupta:2019:FPM

[3008] Lav Gupta, Tara Salman, Maede Zolanvari, Aiman Erbad, and Raj Jain. Fault and performance management in multi-cloud virtual network services using AI: a tutorial and a case study. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106950, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300131.

Ari:2019:RAS

[3009] Ado Adamou Abba Ari, Abdelhak Gueroui, Chafiq Titouna, Ousmane Thiare, and Zibouda Aliouat. Resource allocation scheme for 5G C-RAN: a Swarm Intelligence based approach. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106957, December 24, 2019.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310072.

Lopez:2019:EVS

[3010] Miguel López, Alberto Peinado, and Andrés Ortiz. An extensive validation of a SIR epidemic model to study the propagation of jamming attacks against IoT wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106945, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308990.

Abed:2019:SEE

[3011] Sa’ed Abed, Mohammad Al-Shayeji, and Fahad Ebrahim. A secure and energy-efficient platform for the integration of Wireless Sensor Networks and Mobile Cloud Computing. Computer Networks (Amsterdam, Netherlands: 1999), 165(??):Article 106956, December 24, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618301981.

Khandaker:2019:FTC

[3012] Faria Khandaker, Sharief Oteafy, Hossam S. Hassanein, and Hesham Farahat. A functional taxonomy of caching schemes: Towards guided designs in information-centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106937, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619301616.

Cheon:2019:SAM

[3013] Hye-Rim Cheon and Jae-Hyun Kim.

Social-aware mobile data offloading algorithm through small cell backhaul network: Direct and indirect user influence perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 165(??):Article 106951, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308588.

Bermudez:2019:LVS

[3014] H. F. Bermudez, J. M. Martinez-Caro, R. Sanchez-Iborra, J. L. Arciniegas, and M. D. Cano. Live videostreaming evaluation using the ITU-T P.1203 QoE model in LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106967, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304311.

Jia:2019:TDI

[3015] Siyuan Jia, Matthew Luckie, Bradley Huffaker, Ahmed Elmokashfi, Emile Aben, Kimberly Claffy, and Amogh Dhamdhere. Tracking the deployment of IPv6: Topology, routing and performance. Computer Networks (Amsterdam, Netherlands: 1999), 165(??): Article 106947, December 24, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618304092.